A Complete Bibliography of Publications in Computational Statistics & Data Analysis (2010–2019)

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

02 October 2019 Version 1.22

Title word cross-reference

1 [2434]. 2 [583, 378, 1031, 1043]. 2^{k} [356, 1591]. 2 × 2 [428]. 3 [970, 744, 1340, 1043]. 4 [1372]. ^{(}R) [3]. _{2} [1891]. α [628, 1857, 912]. AR(p) [1196]. ARCH(p) [50]. β [440, 673]. C [690, 2218]. D [1381, 183, 1388, 52, 1142, 2144, 2253, 2590, 1392, 137, 2574]. δ [1554]. g [770, 1459, 1766]. GARCH(1, 1) [1998]. h [1459, 1766]. I [1324, 87, 1664, 1392, 346, 1859]. I_{L}^{r} [313]. J [166]. K [72, 694, 177, 1582, 735, 1450, 1794, 1327, 132, 203, 2408, 194, 1232, 785, 119, 1283, 1048]. L [1372]. L^{1} [1888]. L_{1} [1100, 602, 2003]. L_{2} [211]. L_{∞} [1527]. Λ [1002]. M [235, 41, 343, 1182, 853, 1564, 161, 2046]. n [2046, 1732, 1918]. ν [1550]. P [1806, 1304, 2369, 1096, 760, 856, 1220, 647, 2334, 526, 1646, 649, 944, 1732, 2241, 1592, 1918]. P (T_{1} < T_{2}) [706]. ϕ [608]. ψ [745]. Q [223]. Q_{n} [1564]. R [488, 260]. S [453]. \sqrt{n} [2053]. T [2211, 242, 358, 923, 1597, 1531, 1313, 1944, 441, 1039, 1316, 1526, 249, 768, 1547, 909, 1776, 883, 2372, 1665, 2461, 1311, 422]. τ [1819]. U [2538, 198, 9, 145]. z [1298, 1575].

-and- [1766]. -Birnbaum [440, 673]. -consistent [2053]. -convex [260]. -copula [2372]. -copulas [242]. -D [970, 1340]. -dimensions [2574]. -distribution [909, 1311]. -distributions [358, 768, 1459, 422]. -divergence [608]. -errors [1531]. -estimated [161]. -estimates [1819]. -estimation [235].

-estimators [41, 343, 1564]. -factor [1665]. -GDP [1891]. -group [1048]. -mean [119]. -means [203, 1283, 690, 177, 1582]. -medians [735]. -moments [1372]. -monotone [1450, 1794, 341]. -monotonicity [2408]. -neighborhood [1002]. -norm [1100]. -optimal [2218, 1381, 2211, 52, 2253, 1392, 137, 313]. -optimality [183, 1388]. -out-of- [2046]. -parameter [1372]. -paths [453]. -penalised [1888]. -prime [72]. -prior [770]. -quantile [853]. -ratio [1597]. -records [1554]. -sample [1327, 194, 785, 694]. -series [628]. -small- [1732]. -Spline [1220, 1096, 856, 647]. -splines [1806, 1304, 2369]. -square [72]. -stable [1857, 912]. -statistics [1526, 2538, 9]. -test [1298, 883]. -tests [198]. -tube [1550]. -type [1182, 145]. -value [649, 944]. -values [2334, 526, 1575, 2241, 760]. -vine [1142, 2144]. -weak [745].

100K [3, 80].

2000s [1480]. 2nd [1302, 1476, 2225]. 2SLS [1973].

3rd [305, 1930, 858].

4th [2497].

51 [527]. 52 [611, 80]. 53 [59, 133]. 54 [1870, 790, 651]. 55 [1447, 2063]. 56 [1268]. 57 [1075].

74 [80].

ABC [2432, 2077]. ability [774, 1943]. abnormality [1920]. absence [134]. Absolute [688, 1881, 1767]. abundance [2066, 1345]. accelerated [2106, 1124, 601, 872, 957, 389, 517, 1348, 142, 173, 564, 833, 1067, 442]. Accelerating [2292, 1379, 636]. Acceleration [941, 328, 74, 1941]. acceptance [1656, 674]. accident [1432]. accuracy [1655, 1239, 2222, 1679, 1094]. Accurate [1649, 139, 2004, 2111, 760, 876, 1828]. ACOMCD [638]. acoustic [1596]. across [2328, 2256]. active [1540, 2457, 140]. activity [232, 1348]. acyclic [1717]. AD [1943]. adapted [2175]. Adaptive [1970, 1117, 488, 824, 1844, 1657, 771, 563, 2119, 2037, 1549, 2169, 2219, 684, 1984, 802, 2368, 662, 675, 405, 1744, 198, 1293, 2285, 1624, 1599, 657, 593, 1732, 2374, 376, 2170, 502, 12, 2207, 1641, 2370, 1773, 1041, 664, 1740, 2033, 1975, 1456]. added [1061]. addition [281]. Additive [117, 2315, 2327, 107, 1009, 1703, 2031, 1546, 687, 2462, 1328, 148, 1610, 777, 775, 2610, 1169, 1122, 2187, 2465, 749, 2567, 1135, 1611, 2591, 781, 2453, 521, 2181, 710, 2386, 2071, 2185, 1220, 2378, 2402, 1733, 346, 323, 2182, 1613]. additive-multiplicative [2185]. additivity [1200, 812]. address [1916]. Addressing [2411, 2250, 1000]. adequacy [1014, 159, 1653]. adjacent [947]. Adjusted [1709, 2584, 1610, 2425, 1987, 2249, 1080, 2153, 1335, 1465, 1195, 2578]. Adjusting [2133]. Adjustment [1287, 2505, 2273, 1832, 1273, 2526, 1694]. Adjustments [2538, 649, 2265]. adult [1262]. adults [1419]. Advances [1815, 1786, 2196, 1082, 1302, 1804, 2497]. affected [370]. affine [486]. Affymetrix [80, 3]. after [2385, 825, 1357, 665, 515]. aftershock [455]. against [95, 2490, 921, 1285]. age [2511, 1431]. aggregate [1482, 151, 1222, 366]. Aggregation [233, 1895, 808]. agreement [1088, 704, 17, 519]. agricultural [604]. Agriculture [1971]. aided [390]. AIDS [1795]. air [1865]. Akaike [668]. ALD [2257]. algebra [1379, 2356]. Algorithm [1448, 1761, 634, 690, 466, 658, 941, 1687, 1038, 879, 615, 735, 2539, 1644, 64, 783, 1297, 1682, 1983, 2310, 2029, 1103, 1260, 2481, 1211, 1801, 2542, 2563, 1330, 281, 329, 1272, 2457, 308, 463, 344, 976, 1675, 531, 328, 1114, 1843, 1287, 2441, 2305, 421, 2003, 1573, 1010, 501, 2221, 1657, 731, 1888, 1019, 1922, 150, 2577, 294, 920, 1907, 200, 1868, 1416, 256, 1751, 786, 736, 1612, 140, 74, 1392, 860, 2374, 532, 1406, 636, 2614, 1559, 302, 638, 1660, 2257, 1833, 859, 2070, 142, 137, 1032, 2356, 1461, 1569]. algorithmic [2262]. Algorithms [1385, 326, 1528, 2436, 1436, 759, 1802, 1374, 1175, 1367, 1770, 62, 1623, 1380, 1056, 2158, 844, 1822, 2084, 2205, 259, 818, 12, 1731, 640, 1210, 485, 1522, 1448, 1715]. alignment [2145, 119]. Allison [1074]. allocation [900, 336, 1465, 2370]. allocations [1770]. Allowing [425, 91, 1570, 1502, 1893]. alpha [1465]. Alternating [317, 2551, 775, 328, 1587, 935]. Alternative [2307, 2093, 541, 1073, 56]. Alternatives [1261, 95, 2332, 1008]. among [1194, 1056, 531, 1046]. amplitude [805, 1098]. amplitude-based [805]. Amplitudes [331]. Anal [1447, 59, 2063, 1268, 1870, 1075, 790, 527, 611, 133, 80]. analyses [923, 1655, 2133, 1730, 122]. Analysis [974, 1352, 1631, 1371, 1439, 2187, 2184, 2115, 385, 1849, 2341, 2329, 651, 1228, 1757, 1070, 1990, 266, 2383, 1446, 1180, 1119, 989, 1407, 2428, 882, 871, 1486, 1541, 1100, 326, 1528, 2559, 2278, 183, 696, 635, 1237, 187, 205, 107, 669, 1442, 1703, 2031, 2392, 2405, 928, 1867, 939, 1359, 1238, 1558, 1590, 622, 1164, 1277, 1804, 670, 2108, 2289, 604, 484, 967, 1055, 557, 2048, 423, 1374, 499, 524, 2470, 702, 101, 2132, 1313, 1495, 203, 1071, 1750, 1271, 1161, 1865, 779, 2136, 402, 2566, 82, 2454, 1581, 565, 1847, 1396, 87, 1903, 1332, 766, 1574, 2517, 951, 741]. analysis [587, 839, 2073, 328, 1639, 1584, 26, 154, 583, 780, 1472, 2149, 2139, 391, 444, 2591, 276, 1627, 1811, 1227, 836, 2021, 1276, 1128, 2388, 768, 597, 332, 1797, 1896, 2225, 2209, 518, 324, 251, 1337, 1188, 108, 981, 2363, 738, 1320, 134, 1235, 1033, 1294, 1341, 1031, 1814, 1465, 594, 1275, 257, 1917, 841, 2204, 733, 1863, 508, 2611, 400, 1343, 982, 1244, 618, 2614, 887, 1883, 302, 188, 1842, 719, 1206, 373, 303, 213, 1186, 1212, 1660, 2378, 407, 2258, 75, 1001, 1778, 2501, 24, 516, 564, 1600, 1582, 25, 147, 1893, 184, 174, 1720, 2078, 2515, 2581, 340, 955, 1409, 2509, 2468]. analytic [2105, 2204]. analytics [1719]. analyzers [1921, 553, 1547, 2373, 1665, 2613]. Analyzing [424, 715, 456, 722, 63, 1017, 2453, 2028]. Anatomical [1695]. Ancillarity [1503]. Ancillarity-sufficiency [1503]. Anderson [956]. Angle [2330, 1818]. Angle-based [2330]. angular [2083]. animal [1596, 2072, 370]. Anisotropic [484]. Annals [863, 1476, 1930]. annealing [1528, 934]. announcements [232]. anomalous [2447, 1028]. ANOVA [1289, 693, 2568, 2346, 179, 1096, 2488]. ant [638]. anthropometric [598]. anti [223]. anti- [223]. antimicrobial [1305]. anxiety [592]. any [425, 179, 20]. applicability [859]. applicable [2472]. Application [1398, 1726, 2408, 281, 2400, 2426, 1276, 2221, 1868, 598, 2035, 885, 2383, 2435, 1483, 830, 2491, 1995, 931, 879, 753, 850, 1289, 575, 2323, 212, 1442, 2509, 695, 2145, 237, 551, 1064, 2154, 1610, 874, 1404, 2177, 600, 524, 2132, 1385, 623, 1750, 1066, 649, 1869, 2350, 392, 2296, 1061, 1039, 1825, 1688, 1763, 2591, 854, 1555, 1319, 1795, 432, 918, 1508, 1079, 1509, 1693, 963, 2496, 413, 2344, 732, 1814, 458, 1228, 2434, 508, 109, 1978, 1891, 1753, 1279, 2012, 1773, 659, 353, 2163, 442, 1761, 2117]. Applications [903, 390, 262, 1251, 1834, 560, 986, 1133, 1057, 433, 838, 452, 1322, 27, 371, 499, 1113, 103, 193, 481, 2127, 1305, 285, 817, 520, 2149, 1664, 443, 1719, 710, 289, 651, 378, 907, 200, 1637, 379, 1031, 593, 1195, 42, 1559, 1595, 1186, 1459, 1840, 2488, 2603, 1990, 1975]. applied [1848, 1304, 2432, 1585, 768, 1191, 53, 2582, 117]. approach [2471, 803, 2093, 989, 2080, 338, 1931, 2559, 1308, 2390, 55, 2129, 158, 107, 273, 2031, 35, 43, 2602, 1558, 946, 622, 1055, 1708, 1435, 2516, 2137, 2319, 775, 2025, 2342, 702, 2199, 101, 1676, 896, 693, 2032, 1081, 2036, 1948, 2454, 1892, 165, 1132, 1596, 2350, 2400, 1680, 1017, 476, 533, 1125, 1609, 1291, 1953, 385, 337, 1073, 1741, 487, 1223, 1472, 2580, 1721, 1137, 1295, 1679, 2316, 1979, 2046, 1445, 410, 494, 394, 586, 1704, 252, 630, 1262, 1369, 409, 2536, 911, 1031, 2150, 1086, 1889, 1257, 733, 577, 1220, 947, 647, 259, 1677, 961, 1978, 808, 2382, 459]. approach [387, 1212, 1542, 1662, 1591, 2042, 46, 184, 1255, 2581, 1874, 1919, 2033]. approaches [939, 1902, 794, 1945, 566, 2517, 834, 293, 300, 2205, 1633, 1070, 331, 1549]. Approximate [1266, 1674, 2120, 415, 452, 730, 737, 2342, 588, 329, 470, 23, 309, 2077, 1411, 1040, 398, 326, 55, 1698, 1937, 1385, 900, 67, 1829, 1609, 2092, 2030, 2393, 1527, 2364, 2191]. approximated [1766]. Approximating [65, 956, 1973, 2598, 605]. Approximation [2350, 1968, 1597, 2452, 86, 2481, 1821, 708, 350, 1843, 2564, 1922, 292, 2051, 1589]. approximations [368, 1095, 2236, 2369, 766, 401, 1837, 320, 762, 1393]. arbitrage [295]. arbitrary [2342, 375]. ARCH [1975]. archaeological [2428]. archetypal [1708, 423]. archetype [2048]. archetypoid [2048]. Archetypoids [1708]. Archimedean [277, 493, 321, 1408]. area [473, 1850, 857, 1565, 850, 848, 852, 2391, 1546, 849, 1926, 851, 2324, 2273, 854, 111, 853, 1050, 985, 206, 2237, 48, 1742, 1145, 1715]. area-level [849]. areas [2102, 856]. ARFIMA [905, 1044]. arising [603]. arithmetic [542]. ARL [2446]. arm [183, 2492]. ARMA [1953, 2264, 2522, 642, 252, 982, 2325, 1961]. Arnold [1756]. Array [80, 1903, 3]. arrays [2298]. art [871]. article [609]. ascent [616].

ASIS [1503]. aspects [242, 920, 1423]. assemblage [822]. Assessing [268, 404, 1088, 2406, 1014, 774, 159, 1145, 486, 2206]. Assessment [704, 99, 964, 639, 1133, 701, 1328, 967, 2490, 900, 1355, 627, 1086, 556, 77, 1535, 2035]. assessments [1061]. asset [900]. assignment [2203]. assimilation [2434]. assistance [672]. assisted [850]. associated [923, 753, 1143, 1595, 1525]. Association [987, 497, 151, 1222, 703, 1268, 481, 2146, 1225, 1555, 1405, 1419, 14, 552]. assumptions [2511, 614]. Asymmetric [1299, 1550, 2498, 732, 897, 1462, 907, 626, 1978, 1210, 1372]. asymmetrically [909]. asymmetries [1619]. Asymmetry [1951, 1343, 1248]. Asymptotic [2545, 1424, 1428, 708, 2417, 14, 901, 910, 198, 883, 46]. Asymptotically [2039, 681, 2075, 1618]. Asymptotics [242, 1084]. Atlantic [1108]. atoms [2000]. attitudes [1726]. attributable [946]. attributed [171, 170]. attributes [1536]. AUC [1278]. augmentation [656, 1749]. augmented [1325]. Augmenting [1388]. Aumann [1361]. Australia [985]. Australian [2432]. author [2, 611]. autism [623]. auto [966]. auto-logistic [966]. autocorrelation [1972, 2446]. autocorrelation-consistent [1972]. autocorrelations [50]. autocovariances [1986]. autologistic [1674]. automata [1717]. Automated [236, 1409, 1516]. Automatic [2220, 1107, 2585, 1134, 501]. Automating [1271]. autoregression [812]. autoregressions [1991, 1512]. autoregressive [2094, 1325, 2318, 865, 1915, 2180, 1499, 1598, 337, 1691, 1585, 2474, 1972, 2117, 1170, 252, 545, 1030, 2604, 127, 1709]. autoregressive/unit [1972]. auxiliary [1703, 2185]. availability [2293]. average [1583, 674, 534, 821, 1507, 2248, 2183, 432, 1619, 2607, 1241]. averaged [581, 843, 1048]. averages [853]. Averaging [301, 1471, 1017, 1354, 295, 1357]. avoidance [1387]. Avoiding [216]. Azen [2469, 2468].

B [427, 2135, 2519, 1007]. B-spline [427, 2135]. B-splines [2519, 1007]. backfitting [1122]. background [2133]. Backward [2374, 329]. bagging [1649]. Balakrishnan [1075, 977]. Balanced [2395, 2423, 1441, 179]. Balancing [1678]. ball [2615]. band [2379]. bands [1617, 143]. Bandwidth [746, 1122, 2019, 1102, 1229, 2055, 588, 2089, 999, 1970, 675, 2076, 924, 1207, 2011, 1734, 1569, 1456]. Bank [627]. based [2093, 358, 2411, 338, 830, 1197, 2283, 269, 1038, 2267, 542, 45, 1925, 1398, 1307, 404, 850, 1696, 427, 2390, 2596, 212, 1762, 2129, 669, 2405, 1377, 1792, 946, 2601, 526, 238, 148, 1464, 554, 646, 2226, 2452, 1297, 471, 2310, 1923, 796, 465, 1567, 1790, 702, 132, 1051, 1941, 2444, 2481, 896, 66, 641, 799, 2202, 1554, 2089, 243, 2369, 666, 2361, 2032, 1149, 2250, 1201, 1735, 2454, 2066, 1315, 1309, 2355, 1756, 34, 284, 2360, 983, 16, 903, 2144, 1914, 514, 160, 817, 2398, 487, 583, 1293, 2441, 435, 1185, 1971, 1526, 2421, 2316, 1276, 410, 1168]. based [1668, 834, 194, 2474, 2230, 1097, 1334, 71, 1789, 856, 1509, 586, 1036, 378, 1463, 2433, 1339, 320, 2277, 1378, 1681, 1751, 1636, 2027, 384, 881, 1031, 167, 458, 2103, 619, 2279, 714, 1772, 84, 1780, 1819, 206, 579, 1626, 2068, 990, 1974, 508, 2513, 1516, 805, 1564, 1343, 341, 2357, 104, 1438, 532, 716, 1707, 537, 48, 1130, 661, 692, 53, 1303, 261, 1317, 638, 127, 573, 671, 1375, 1651, 2206, 2613, 2461, 962, 2402, 2330, 1562, 2541, 422, 1461, 1840, 1745, 157, 2033, 1743]. baseline [2299]. bases [1096]. Basic [1371]. basis [464, 401, 1863]. batch [1377]. batch-sequential [1377]. bathtub [1123, 1020, 1214, 93]. bathtub-shaped [1123, 1020]. Baumgartner [990]. Bayes [126, 1621, 1326, 465, 2273, 1847, 992, 2001, 726, 421, 2569, 1803, 1335, 1340, 300, 2349, 808, 1347, 207, 552]. Bayesian [1268, 266, 1381, 1155, 2020, 740, 91, 2056, 1813, 989, 83, 368, 338, 345, 964, 2561, 1471, 937, 1976, 823, 1325, 2396, 656, 2413, 1935, 1806, 548, 427, 364, 433, 838, 2552, 865, 55, 2186, 205, 158, 897, 1226, 1490, 1327, 1322, 35, 928, 43, 1698, 939, 1988, 695, 1531, 1937, 1328, 2135, 2412, 2376, 1085, 867, 604, 540, 996, 730, 737, 2177, 44, 1104, 703, 1902, 1706, 1623, 2375, 1746, 1329, 2369, 244, 1161, 1330, 1557, 1865, 1388, 1997, 2136, 2321, 1829, 914, 2454, 453, 943, 1596, 470, 69, 675, 2420, 2529, 2400, 908, 1680, 1017, 1332]. Bayesian [754, 741, 1154, 1609, 1857, 149, 1914, 562, 1157, 351, 621, 1603, 2058, 121, 1604, 1458, 1690, 1853, 1287, 2253, 2345, 1431, 2092, 2248, 2451, 2139, 2450, 2590, 1137, 444, 2571, 1432, 1460, 1979, 1430, 1719, 2152, 1198, 1657, 1079, 2212, 1334, 734, 2209, 518, 1336, 324, 1337, 293, 1188, 1339, 393, 1405, 252, 2164, 295, 296, 2496, 1449, 175, 2373, 339, 912, 1243, 72, 1235, 297, 2397, 1341, 192, 156, 1814, 1089, 594, 2030, 2393, 2530, 1513, 1425, 727, 1889, 2399, 1274, 1772, 1342, 1021, 2489, 1782, 1120, 1207, 2523, 2049, 2534, 1343, 947, 1244, 961, 1533, 110, 618, 2614]. Bayesian [20, 887, 1883, 188, 1084, 1978, 1965, 118, 720, 373, 387, 814, 945, 1856, 2448, 2555, 2364, 159, 1662, 2207, 2213, 1992, 2501, 516, 564, 590, 832, 2351, 793, 184, 323, 572, 1032, 117, 1592, 1697, 2078, 2416, 2515, 2504, 643, 1874, 2507, 1456, 592, 495]. BBNs [68]. beam [1693]. behave [1258]. Behavior [1738, 1242]. behaviour [1947, 1753]. Behrens [156]. Benchmark [818, 1374]. Benchmarking [761, 2105]. Bernoulli [2073, 2351]. Bernstein [2168, 1802, 1706, 649, 1678, 663, 1397, 837]. Bernstein-type [649]. Best [1721, 1643, 138]. beta [767, 1376, 1104, 1, 527, 747, 413, 94, 39, 495]. beta-generated [767]. better [1436]. between [1897, 1095, 2376, 86, 1374, 2132, 2580, 1678, 1536, 497, 20, 644, 221]. bi [2152]. bi-level [2152]. Bias [436, 555, 1420, 1781, 1646, 49, 1462, 1870, 910, 1080, 1273, 856, 1973, 2582, 216, 1709, 2507, 1943]. bias-adjusted [1080]. bias-corrected [856]. biased [2219, 64, 2593, 2495, 2134]. biased-coin [2219]. BIC [1724]. biclustering [1789, 1731]. big [1719]. big-data [1719]. Binary [146, 1323, 634, 338, 823, 1687, 18, 670, 447, 704, 1272, 1596, 2184, 685, 1147, 23, 121, 2301, 1679, 1602, 1627, 2243, 1336, 713, 384, 98, 2337, 1425, 2489, 2233, 1707, 774, 304, 2510, 1640, 1001, 723, 763, 2351]. Bingham [2120]. binned [310]. binning [91]. Binomial [1345, 2343, 1610, 430, 1063, 1688, 520, 1232, 963, 819, 123, 835, 610, 1737]. bioassay [1133, 158]. biological [1829, 444, 485]. biomarker [1326, 2296, 2347]. biomarkers [1643, 1247, 48, 2206, 721, 2208, 1715]. biomedical [2407]. biomolecular [102]. Biostatistics [2487]. biosurveillance [1032]. bipartite [382]. biplots [2444]. BIRCH [67]. Birnbaum [923, 979, 440, 646, 125, 412, 522, 1144, 1128, 40, 673, 750, 460, 1595, 1898, 958, 37, 24, 516]. Bispectral [1149]. Bispectral-based [1149]. Bivariate [2181, 931, 554, 447, 1655, 1631, 2013, 1902, 2445, 22, 224, 2542, 1161, 493, 1039, 992, 2189, 351, 780, 1078, 1526, 1010, 150, 192, 2489, 369, 2560, 1595, 1898, 1660, 408, 420]. Blackwellisation [70]. block [306, 2354, 176, 2572, 2544, 1387, 1330, 363, 1472, 1627, 2243, 1079, 2095, 2312]. block-wise [2243]. blocks [2512, 590]. BLUP [1702]. Board [2274, 2290, 2302, 2313, 2348, 2359, 2377, 6, 209, 225, 263, 29, 60, 78, 115, 136, 163, 180, 196, 314, 558, 569, 584, 396, 418, 445, 477, 489, 505, 529, 538, 612, 846, 861, 916, 632, 652, 677, 697, 724, 791, 806, 826, 971, 972, 1023, 1058, 1059, 1076, 1077, 1092, 1093, 1110, 1111, 1126, 1140, 1162, 1173, 1192, 1216, 1245, 1269, 1300, 1394, 1417, 1433, 1452, 1474, 1523, 1551, 1571, 1593, 1615, 1628, 1647, 1671, 1683, 1700, 1711, 1712, 1728, 1747, 1758, 1768, 1783, 1823, 1851, 1872]. Board [1885, 1899, 1909, 1928, 1981, 2005, 2014, 2044, 2064, 2085, 2098, 2112, 2125, 2142, 2155, 2171, 2194, 2231, 2245, 2246, 2322, 2336, 2389, 2401, 2422, 2449, 2459, 2467, 2486, 2503, 2524, 2537, 2546, 2556, 2575, 2592, 2606]. Boosting [211, 2198, 927, 1503, 1822, 2485]. Bootstrap [968, 801, 1361, 2460, 1957, 619, 1264, 1082, 1968, 2055, 272, 893, 220, 2557, 2452, 2547, 471, 655, 2199, 1136, 745, 363, 1094, 2076, 811, 877, 2046, 853, 1310, 1257, 1626, 318]. Bootstrapping [403, 2367, 1065, 2544, 1650, 457]. BOPA [943]. borrowing [1333]. both [1223, 2301, 1369]. bound [1047, 636, 147]. bound-inflated [147]. boundary [2100, 2295, 237, 1400, 1566]. bounded [1781]. Bounding [1608]. bounds [1821, 707]. brain [2132]. Branching [2205, 1801, 2167]. break [1494, 1570, 1502]. breaks [886, 1480, 1489, 238, 866, 246, 1196, 2528, 1965]. breast [836, 390, 442]. Bregman [1491, 451]. Bregman-proximal [1491]. bridge [2152]. bridging [2580]. British [1726]. broad [859]. broader [2547]. Brown [967, 1539]. bubble [1941]. build [2178]. building [1796]. Burr [413]. byproduct [2267].

C [1379]. C. [609]. C.-Y. [609]. calculating [106, 318]. calculation [942, 1451, 1048]. calculations [1412, 727]. Calibration [929, 795, 1702, 386, 454]. cancer [2491, 1995, 479, 587, 836, 390, 643, 2163, 442]. Cancho [1447]. Candecomp [1469]. Candecomp/Parafac [1469]. candidate [556]. cannot [900]. Canonical [2109, 1757, 1990, 2383, 1119, 635, 1277, 108]. capacity [2213]. capture [514, 370, 647, 1879]. capture-recapture [370, 1879]. car [2591, 1026, 734]. cardinality [66]. cardiovascular [2209, 598]. caries [1688]. Carlo [494, 1413, 890, 2239, 2437, 850, 33, 996, 1404, 1890, 1330, 1060, 165, 69, 70, 1857, 2110, 2583, 2522, 1137, 1803, 1797, 2577, 370, 736, 257, 2074, 2170, 168, 2597, 1662]. CART [2250]. CART-based [2250]. cascading [1483]. Case [795, 1866, 1904, 1702, 364, 995, 2289, 21, 41, 2201, 978, 2477, 2209, 1208, 1961, 718, 2347, 1210, 14, 2455, 925]. case-cohort [2289]. case-control [995, 1208, 14, 2455]. Case-deletion [795, 1702]. case-only [2455]. casewise [1911, 2161]. catastrophe [370]. CATCH [1406]. categorical [937, 1238, 524, 623, 191, 1138, 1017, 121, 2365, 2399, 982, 2614, 519, 945, 216, 1576]. categories [1239, 1115, 1562, 1733]. category [1176]. Causal [2203, 1811, 2084, 793]. Cause [2059, 2306]. Cause-specific [2059, 2306]. causes [1764]. CCA [286]. cDNA [659]. CECM [690]. cell [1248]. cellwise [1911, 2161, 1817]. censored [560, 933, 1621, 107, 1327, 2031, 2311, 2431, 2538, 429, 607, 2063, 2200, 2445, 566, 100, 1810, 155, 1805, 1305, 587, 976, 149, 1914, 2115, 1584, 844, 2173, 522, 824, 2514, 683, 1083, 1445, 2021, 1006, 673, 663, 1449, 682, 1771, 1414, 167, 1214, 1120, 1626, 1807, 2611, 816, 456, 1438, 2300, 1809, 1736, 765, 719, 517, 2378, 2613, 778, 1725, 184, 174, 2358, 1087, 605, 1919]. censoring [2056, 1420, 986, 979, 1656, 152, 2190, 277, 464, 1631, 706, 316, 336, 2059, 2149, 2567, 654, 150, 1577, 1129, 84, 1444, 2276, 1347, 564, 1204]. censorship [189]. census [850]. Center [1553, 2403, 2271, 2187]. Center-within-trial [1553]. centering [1907]. Central [627, 2403, 1526, 768, 1775]. centrality [2588]. centralized [342]. centred [345]. cervical [2163]. CFEnetwork [1476, 1930]. Chain [2583, 2437, 1404, 2597]. chains [1134, 1010, 736]. Change [1989, 923, 891, 2436, 43, 2544, 566, 2454, 1892, 1069, 336, 1227, 1508, 1889, 2447, 2325, 1016]. change-point [891, 2436, 43, 566, 1892, 1227]. changes [801, 162]. changing [2050, 1026]. Chao [514]. chaotic [2350]. Characterising [1329]. Characteristic [537, 1800, 903, 1777, 785, 1743]. characteristics [1848, 1202, 985]. characterization [1756]. chart [534, 954, 214, 157]. charts [578, 821, 87, 222, 1738, 488, 2083]. Checking [1653, 2576, 531, 905, 2580, 1831, 1591, 50]. chemical [2496]. Chi [2452, 2608]. Chi-square [2452, 2608]. Choice [1468, 746, 782, 1609, 2121, 395, 1130, 1303, 1249]. choices [138]. Cholesky [606, 2013, 1604, 2264, 2183]. chromatography [2073]. chronic [1716, 1015]. Cicchetti [1074]. circle [52]. circles [504]. circular [599, 1339, 924, 2035]. class [1118, 937, 1373, 1796, 756, 1237, 1689, 1386, 1646, 2038, 383, 1661, 948, 2514, 1763, 854, 438, 747, 2103, 1917, 888, 2513, 1053, 39, 692, 1739, 1535, 1087]. classes [2408, 2328]. Classical [1902, 834, 282, 2038, 2215]. Classification [1360, 1848, 2251, 1925, 512, 799, 1460, 117, 2414, 2471, 482, 803, 358, 974, 701, 479, 1644, 148, 1134, 2543, 1826, 676, 2157, 1557, 1000, 1291, 2249, 1295, 667, 1430, 1896, 2178, 1791, 1637, 2472, 192, 577, 980, 1043, 1406, 1403, 1156, 2349, 1358, 808, 1317, 1561, 382, 46]. classifications [121]. classifier [1545, 1640, 1840]. classifiers [1904, 1362, 1299, 121, 2513, 1145]. classifying [1080, 993]. Clayton [1439]. clickstream [1789]. clinic [1419]. Clinical [699, 2169, 2219, 1326, 2405, 1159, 1091, 710, 2178, 712, 2489, 2204, 556, 376, 718, 2207, 722, 1254]. clipped [191, 1774]. cloning [466, 1289, 2077, 952, 1742]. closed [1361, 1998, 1454, 647]. closed-form [1998]. closeness [830]. closure [1206]. cloud [2007]. clouds [1043]. clubs [256]. Cluster [1021, 1189, 183, 601, 499, 203, 2133, 1091, 1315, 510, 2203, 154, 496, 2421, 1294, 714, 424, 1686, 638, 75, 1600, 174, 1894]. cluster-specific [2203]. cluster-weighted [1315]. clustered [322, 589, 2334, 987, 2031, 601, 101, 1718, 2549, 1680, 1688, 202, 2189, 1584, 385, 2363, 1065, 456, 1842, 517, 407, 723, 763, 1701, 1630, 174]. Clustering [1304, 58, 1333, 1112, 1896, 1788, 2351, 2471, 1511, 2411, 2080, 338, 2491, 2572, 18, 1307, 1308, 735, 691, 759, 2145, 2029, 1790, 1103, 2562, 66, 2272, 1330, 1149, 2066, 1315, 1309, 2355, 510, 842, 16, 496, 2499, 2285, 768, 2230, 71, 1306, 1791, 614, 119, 1782, 1787, 805, 1707, 1303, 261, 1317, 2206, 2448, 2613, 2461, 1753, 485, 1992, 2410]. clusters [183, 1489, 974, 655, 1421, 1895, 2498, 1230, 1740, 1549]. clusterwise [279]. cluttered [2007]. co [2572, 2355, 1015, 1891]. co-clustering [2572, 2355]. co-morbidity [1015]. coarsened [374]. coauthor [2, 611]. coefficient [2478, 404, 1696, 89, 1088, 1689, 131, 582, 1926, 932, 965, 728, 1125, 1666, 2352, 348, 2438, 2386, 2276, 947, 1906, 1005, 1633, 814, 2521, 994, 2012, 1725, 1739, 580, 1859, 1576, 1745, 2419, 2603]. coefficients [1958, 104, 2326, 260, 2088, 2485, 1845, 515]. cognitive [1419]. Cohen [408]. cohort [2289]. coin [2219]. cointegrated [855, 545]. cointegration [238, 403, 1501]. colleagues [80]. colony [638]. column [2266]. COM [1197, 2327]. COM-Poisson [1197]. combination [1382, 1296, 2288, 543, 2400, 771, 1715]. combinations [1936, 1643, 1640]. combinatorial [66, 67]. combine [1836]. combined [772, 2222, 1185, 1310, 48, 213]. combiner [2118]. Combining [526, 299, 1206, 1298]. Comments [609]. common [1194, 2227, 271, 1029, 2000, 2256, 312, 1665, 90]. Community [2600, 2314]. Comorbidity [1716]. Comparative [390, 556, 1592, 890, 1497, 476, 161]. compare [503, 713]. Comparing [1864, 2038, 901, 1530, 81, 926, 450, 2217, 1147, 2114, 568, 106, 398, 519, 1048]. Comparison [635, 702, 623, 1496, 1056, 1860, 528, 487, 787, 1633, 1210, 473, 1447, 2565, 1095, 272, 2437, 1528, 2188, 363, 536, 287, 2114, 1868, 74, 104]. comparisons [86, 1545, 1039, 921, 707, 944, 2061, 1232, 1046, 2091, 449, 130, 1241, 331, 1694]. compartment [755]. compatibility [531]. compatible [467]. Competing [2189, 1864, 1764, 2186, 429, 1631, 199, 2115, 1987, 2059, 2363, 1130, 1601, 564]. competitive [2221]. complementary [426, 1447, 536]. complete [176, 2412, 590, 1409]. completion [2569, 1778]. complex [1625, 1164, 1410, 2324, 2426, 2225, 2600, 671]. complexity [1597]. compliers [2602]. Component [974, 2509, 1446, 1480, 1100, 1528, 2016, 1442, 1055, 44, 524, 2136, 1847, 766, 1744, 2371, 344, 555, 734, 2209, 981, 73, 1863, 102]. component-wise [73]. components [2545, 2227, 1471, 546, 271, 864, 153, 777, 766, 172, 328, 1722, 2153, 1390, 731, 1793, 2373, 1310, 379, 409, 1086, 1787, 1469, 952, 312, 1633, 2326]. componentwise [1396]. Composite [978, 1578, 2128, 2239, 2366, 1386, 1169, 2146, 1153, 2540, 2183, 630, 2150, 1086, 1465, 2257]. Compositional [2353, 282, 741, 2265, 834]. Compound [1543, 438, 732, 1053]. comprehensive [2073]. compression [2285]. Comput [1447, 59, 2063, 1268, 1870, 1075, 790, 527, 611, 133, 80]. Computation [830, 1656, 1331, 1429, 1986, 2218, 891, 1597, 2535, 1980, 325, 1698, 33, 904, 1901, 278, 1944, 1829, 2043, 1887, 970, 1039, 1609, 1114, 2092, 2047, 918, 448, 1871, 2030, 2393, 1828, 2614, 2364, 2356, 420]. Computational [227, 863, 1364, 1476, 1930, 356, 710, 651, 699, 920, 2205, 884, 1610, 242, 813, 2316, 332, 1588, 185, 147, 2468]. computations [305, 2081, 1854]. compute [300]. computed [1693]. computer [949, 1608, 27, 2222, 1338, 929, 390, 1841]. computer-aided [390]. Computing [1473, 86, 1389, 760, 642, 686, 466, 2343, 2539, 1322, 1937, 737, 307, 1063, 69, 308, 1675, 166, 1078, 1198, 1393, 137]. concave [1473, 1993, 2158, 2609]. concentration [1926, 280]. Concept [458, 2118]. Concordance [2326, 1088, 1633]. concrete [2223]. concurrent [921]. condition [876, 580]. Conditional [1101, 1762, 481, 668, 1266, 1511, 1864, 1932, 2168, 865, 2596, 467, 544, 607, 2063, 2543, 65, 1942, 1260, 566, 899, 1439, 1948, 165, 308, 1895, 1563, 2517, 901, 2265, 531, 23, 1691, 1585, 681, 2345, 1844, 1577, 2433, 1188, 8, 1423, 1714, 2268, 1517, 1963, 1403, 1978, 2446, 1537, 950, 2427, 50, 1975]. conditionally [2016, 1953, 812, 1249]. conditions [1109, 1947, 627, 113]. cone [1693, 1720]. cone-beam [1693]. Confidence [219, 15, 1862, 1427, 1681, 2027, 2074, 1961, 4, 133, 716, 48, 13, 267, 1473, 404, 752, 1643, 1361, 2424, 1689, 1113, 581, 1136, 1617, 2273, 143, 1239, 921, 707, 1147, 709, 984, 1273, 28, 384, 1257, 1046, 2091, 1807, 609, 106, 843, 1278, 2088, 962, 2456, 1461, 32]. confirmatory [1590, 276]. confounders [1811]. Confounding [1287, 2538]. conjoint [138]. Conjugate [1155, 1188]. connected [1787, 1542]. connectivity [1990]. consensus [2256]. considerations [395]. Consistency [687, 1465, 2255, 1534]. Consistency-adjusted [1465]. Consistent [1618, 2300, 944, 2053, 1972, 1042]. constancy [1399]. constant [904, 870, 1123, 1521]. constants [1332]. Constrained [690, 759, 1890, 2271, 38, 1754, 2166, 1682, 1175, 1103, 567, 1692, 1903, 2081, 276, 1227, 1668, 1335, 2577, 1663, 2118, 859, 1773, 1522, 2362]. Constraining [2587]. constraint [199, 1842, 778, 1634]. constraints [1398, 2289, 1770, 66, 1921, 463, 286, 393, 2312, 645, 610, 207]. constructing [1415, 1689]. Construction [2511, 1390, 1436, 1219, 1681, 716]. constructions [1673, 1509]. consumer [255]. contagion [234, 896]. containing [355]. Contaminated [1213, 2498, 1827]. contamination [1911, 2161, 1817]. content [171, 170]. Contents [7, 210, 226, 264, 30, 61, 79, 116, 135, 164, 181, 197, 315, 559, 570, 585, 397, 419, 446, 478, 490, 506, 530, 539, 613, 847, 862, 917, 633, 653, 678, 698, 725, 792, 807, 827, 1024, 1127, 1141, 1163, 1174, 1193, 1217, 1246, 1270, 1301, 1395, 1418, 1434, 1453, 1475, 1524, 1552, 1572, 1594, 1616, 1629, 1648, 1672, 1684, 1699, 1710, 1713, 1729, 1748, 1759, 1769, 1784, 1824, 1852, 1873, 1886, 1900, 1910, 1929, 1982, 2006, 2015, 2045, 2065, 2086, 2099, 2113, 2126, 2143, 2156, 2172, 2195, 2232, 2275, 2291, 2303]. context [695]. contingency [1923, 1225, 1240, 1913, 1591, 1562, 2410]. Continuous [2042, 2219, 35, 1150, 244, 2050, 68, 1139, 16, 1506, 1955, 2301, 1010, 2536, 2205, 341, 450, 1812, 335, 2067, 25, 2269, 2018]. continuous-GMM [1506]. continuous-time [1010]. Continuum [2380, 903, 1231]. contours [1078]. contraceptives [705]. contrastive [2110]. contrasts [143, 984]. contribution [2191]. control [2435, 879, 1057, 158, 995, 578, 821, 2229, 954, 87, 222, 921, 707, 944, 1738, 991, 2114, 1208, 449, 130, 14, 2455, 214, 1905, 157]. controlled [2201, 1801, 718]. controls [2405, 921]. convenience [813]. conventional [316]. convergence [1659, 256, 1534, 137]. Convergent [1761, 2305]. Conversion [121]. Convex [2224, 1215, 2342, 1624, 1455, 260]. convolution [2062]. convolutional [1040]. convolutions [2004]. Cook [1638]. Coordinate [616, 1114, 1888, 1392, 640, 1522]. coordinate-exchange [1392]. coordinate-wise [1522]. Copula [447, 645, 2402, 2018, 740, 2283, 1983, 1802, 277, 1939, 1673, 245, 2229, 2368, 1500, 434, 16, 2175, 2144, 528, 2092, 2054, 2181, 2587, 1509, 1408, 424, 1517, 1716, 2372, 422, 2515]. copulae [1064]. copulas [230, 913, 2596, 554, 1134, 242, 481, 2542, 493, 321, 1142, 2345, 2570, 907, 2096, 2608]. Core [36, 1029, 1903, 1341]. corporate [761]. correct [1943]. corrected [856]. Correcting [288]. correction [507, 237, 1781, 49, 1870, 403]. corrections [1264]. Correlated [753, 2553, 819, 1383, 1384, 2217, 2036, 191, 2038, 1429, 562, 984, 2453, 1575, 1370, 1019, 568, 630, 1449, 2218, 2279, 1749, 2233, 577, 1292, 2128]. Correlation [2132, 1757, 2122, 1990, 1511, 2383, 1119, 1932, 234, 864, 1541, 404, 635, 2539, 1088, 2311, 2544, 1277, 1816, 1689, 239, 582, 820, 1494, 899, 434, 1282, 2087, 1293, 1018, 108, 8, 2312, 424, 1906, 1633, 2326, 1978, 303, 814, 1600, 1893, 1739]. correlations [229, 1480, 769, 2425, 1942, 2483]. Correspondence [2428, 1320]. corresponding [17, 709, 57, 609]. Corrigenda [21]. Corrigendum [59, 1268, 1870, 790, 417]. cosine [599]. Cost [784, 395, 1838, 1398, 813, 1588, 778]. Cost-effective [784]. cost-effectiveness [1838]. Cost-efficiency [395]. count [2111, 2339, 989, 354, 19, 2174, 2129, 1586, 385, 2550, 1426, 1099, 2054, 1105, 2534, 616, 1893, 1045, 340, 975]. Counterfactual [878]. counterpart [1447, 536]. counting [639, 19]. counts [1915, 514]. course [1304, 395]. Court [2145]. Covariance [2494, 1771, 85, 857, 272, 864, 2315, 2407, 1605, 2013, 480, 342, 2361, 2307, 1948, 82, 1970, 1860, 983, 742, 1604, 2264, 749, 2062, 2441, 2047, 276, 2183, 1148, 401, 494, 1455, 614, 379, 1876, 2463, 1667, 1568, 2479, 1556, 2599, 1052, 2234, 2615, 2034]. Covariance-insured [2494]. Covariate [869, 1987, 303, 1259, 2219, 823, 2286, 2023, 2024, 2249, 2329, 2525, 1795, 253, 1876, 1470, 765, 335, 1451, 2358]. covariate-adaptive [2219]. Covariate-adjusted [1987, 2249]. covariates [1670, 1252, 1687, 2197, 2505, 380, 1703, 1996, 2431, 2538, 1723, 782, 622, 1746, 1178, 1069, 1832, 2341, 1223, 2540, 2230, 291, 1337, 1871, 1879, 189, 1881, 1224, 2415]. cover [980]. coverage [1273]. Cox [371, 800, 2021, 1780, 1208, 407, 1292]. CPCRA [1763]. Cramér [59, 5, 65, 785]. Credal [1362]. Credible [123]. Credit [696, 1478, 1483, 895, 2516, 117]. Crevecoeur [1020]. crime [2434]. crimes [1262]. crisis [1480, 887]. criteria [1484, 1258, 2038, 1952, 1290, 2365, 1600, 1082]. Criterion [703, 1268, 572, 2334, 1980, 2286, 2025, 662, 621, 2590, 2279, 2513, 1156, 159, 2252, 668]. critical [65]. criticism [251]. Cross [1179, 2095, 2208, 473, 1485, 2318, 272, 2490, 281, 308, 908, 1950, 766, 1285, 258, 1734]. cross-over [2490]. Cross-sectional [2208]. cross-sectionally [258]. Cross-validated [2095, 1734]. cross-validation [473, 1485, 2318, 272, 281, 766, 1285]. cross-validations [308]. crossed [317]. crossing [886]. crossover [660, 1602, 1627, 2049]. crossvalidation [1468]. CSDA [2468, 1785]. CUB [1560]. cumulative [1726, 2341, 1779, 84]. cure [1197, 2197, 753, 1806, 669, 987, 1902, 2369, 1805, 1584, 2514, 620, 2076, 2103, 2611, 173, 1067, 1919]. current [548, 1703, 2149, 781, 1660, 407, 37]. curve [473, 1695, 1936, 1546, 237, 1297, 2481, 443, 410, 501, 119, 406, 48, 1145, 1715]. curves [969, 1133, 1897, 36, 1800, 2509, 2069, 1545, 1066, 2265, 2574, 738, 456]. curvilinear [1636, 2439]. customer [1303]. CUSUM [2544, 578, 2083]. CUSUMs [2294]. cut [709, 2536, 1247, 609, 1525]. cut-off [1525]. cut-point [709, 1247, 609]. cut-points [2536]. cutaneous [753]. Cutpoint [335]. cycles [788].

D [583, 970, 744, 378, 1340, 1031, 2434, 1043]. DAG [1542]. daily [864]. Dantzig [935, 1751]. Darling [956]. Data [1349, 656, 1815, 1447, 1351, 59, 190, 1268, 1999, 1870, 1075, 2468, 2590, 1573, 651, 527, 2312, 2084, 611, 1749, 133, 129, 80, 2471, 141, 482, 1625, 2111, 219, 969, 1754, 803, 91, 2339, 1717, 1012, 989, 1267, 1321, 466, 354, 19, 933, 889, 2055, 2283, 937, 176, 968, 1838, 1995, 2428, 322, 2572, 2174, 1252, 546, 1038, 542, 1095, 1925, 753, 18, 1307, 589, 2007, 1621, 548, 427, 571, 1289, 2188, 2557, 1597, 433, 838, 1049, 865, 1726, 2323, 2423, 107, 606, 669, 1088, 987, 1009, 1327, 1442, 1703, 2031, 2407, 2418, 2311, 2538, 2424, 2379, 1304]. data [1359, 218, 691, 429, 1164, 2104, 2462, 2473, 2412, 802, 144, 866, 607, 2063, 349, 670, 2108, 1352, 1134, 551, 1159, 1983, 2154, 1402, 2105, 996, 2029, 796, 557, 1708, 1374, 601, 499, 2013, 2470, 41, 1051, 2482, 550, 101, 820, 1718, 1781, 2445, 641, 112, 623, 2549, 2151, 2309, 2157, 693, 15, 932, 366, 1750, 2568, 1945, 1066, 2050, 1586, 666, 2361, 1330, 566, 1557, 1353, 779, 2133, 68, 2271, 743, 1606, 67, 100, 402, 1947, 2250, 1439, 2187, 1331, 191, 82, 1000, 359, 1810, 282, 155, 953, 1139, 1805, 374, 2531, 1166, 2066, 87, 2400, 1903, 2057, 1309, 2355]. data [1680, 1017, 1305, 2184, 503, 383, 1574, 2517, 951, 842, 1154, 976, 2203, 2073, 1147, 744, 223, 708, 149, 1914, 2115, 845, 202, 2189, 2079, 1584, 351, 2192, 790, 23, 160, 26, 385, 844, 817, 1073, 1849, 1855, 2173, 2432, 2398, 2352, 2550, 2506, 749, 1585, 2426, 496, 824, 780, 1223, 1431, 1472, 1426, 2059, 1955, 1954, 2249, 2149, 2406, 2297, 2301, 2451, 2514, 2595, 2540, 2477, 1099, 178, 683, 1316, 1971, 2329, 991, 1763, 2308, 2525, 2591, 1094, 2083, 461, 1460, 781, 1602, 1627, 667, 2183, 1445, 2021, 2281, 330, 1719, 2159, 2453, 1560, 401, 2243, 1128, 2388, 2285, 768, 1837, 288, 834, 597, 2230, 1006, 1566, 494]. data [1789, 57, 1102, 1808, 2225, 2178, 251, 907, 963, 1530, 1339, 2367, 2365, 673, 663, 630, 1449, 310, 413, 355, 2121, 2344, 2472, 828, 2363, 1771, 1158, 1884, 437, 146, 2077, 738, 134, 1033, 409, 1636, 370, 2162, 1876, 1858, 785, 1423, 123, 384, 1411, 2150, 1814, 1776, 1214, 594, 2138, 657, 1275, 595, 2399, 1105, 841, 1228, 1780, 1021, 1065, 2434, 2022, 1960, 2204, 1863, 1916, 1626, 2068, 2439, 508, 1807, 2484, 2611, 577, 1401, 805, 816, 947, 625, 341, 1244, 1518, 980, 456, 1438, 835, 2320, 1642, 2300, 1809, 961, 1005, 1533, 1707, 1195, 993, 2600]. data [2614, 952, 1742, 519, 2326, 887, 1098, 1755, 1303, 2382, 118, 303, 213, 517, 1070, 1165, 1028, 1186, 1561, 1665, 1660, 1891, 2176, 2378, 2613, 2521, 2461, 407, 721, 2258, 2479, 485, 616, 2501, 37, 1589, 722, 1725, 1600, 2330, 2381, 1224, 2028, 25, 147, 408, 723, 763, 1701, 1630, 2119, 1893, 1045, 2351, 793, 1448, 184, 659, 2485, 174, 1461, 1840, 2078, 2122, 2358, 2581, 340, 1183, 1087, 975, 1409, 1877, 2128, 2414, 361, 1990, 1919, 1827, 1775, 1694, 1248, 2034]. data-analytic [2105]. data-cloning [2077]. data-dependent [1331]. Data-driven [190, 2590, 2312, 2084, 129, 503, 1863]. data-rich [1945]. database [195]. databases [1410]. dataset [701]. datasets [757, 735, 596, 730, 2564, 379, 1190]. dating [162]. DC [1203]. DCC [1942]. de-escalation [705]. Dealing [2354, 1237, 437]. death [1601]. decision [1360, 2471, 1875]. decisions [900, 743]. deck [1642]. decomposable [813]. decomposition [1407, 606, 2013, 1495, 22, 224, 223, 1316, 1979, 1037, 97]. decompositions [2568, 1034]. decreasing [649, 1123, 93]. decrement [2511]. Default [2520, 296, 882, 1335]. defective [2103]. define [1708]. definite [292, 1074]. deformable [1761]. Degeneracy [463]. degradation [2259]. degree [1678]. degrees [1692, 2372, 664]. Delaunay [357]. Deletion [629, 795, 1702, 281, 2123, 1866]. demand [252, 255]. dementia [1601]. demonstration [1469]. denoised [414]. denoising [662]. densities [1473, 2016, 452, 237, 1993, 351, 1295, 1054, 1776, 261]. Density [2100, 1165, 83, 2055, 1925, 31, 1308, 51, 973, 1450, 1835, 349, 588, 1781, 641, 1706, 794, 2089, 649, 2000, 2019, 49, 1462, 1870, 453, 165, 999, 1596, 1847, 675, 745, 308, 706, 1598, 344, 533, 2053, 421, 1400, 1678, 1128, 194, 1102, 1588, 1994, 924, 310, 645, 1423, 2609, 2030, 1207, 922, 715, 2011, 980, 1963, 1809, 2466, 1397, 1278, 201, 353, 509, 1569, 1456]. dental [1688]. Dependence [2570, 2283, 1351, 550, 2000, 745, 1951, 1142, 880, 907, 2312, 1517, 1518, 647, 1233]. dependency [1821, 808]. Dependent [491, 2304, 1333, 219, 1967, 2055, 1383, 893, 544, 2442, 1654, 1631, 1865, 1331, 1895, 2149, 2021, 253, 1229, 785, 771, 258, 424, 835, 519, 1451, 2555, 1045, 215]. dependently [796]. Depth [2068, 2349, 2379, 554, 1901, 1078, 991, 461, 2474, 1927, 934, 2600, 420]. Depth-based [2068, 2474]. Depth-weighted [2349]. Derivation [342]. derivative [1220]. derivatives [1428, 738]. Deriving [2105]. descent [1114, 1888, 640, 1522]. Description [1033, 2068]. descriptive [666, 841]. Design [212, 447, 2210, 1398, 1383, 850, 1415, 2405, 27, 2376, 2289, 1385, 2224, 1380, 1060, 1047, 944, 1602, 1627, 287, 194, 2212, 2344, 1536, 2223, 1274, 1772, 2204, 1716, 1520, 304, 2207, 2208, 1254]. Design-based [212, 850]. designed [35, 996, 1391]. Designing [2222]. designs [1381, 88, 1382, 306, 2219, 1324, 187, 782, 2262, 1540, 1384, 2217, 2211, 2310, 549, 1386, 705, 1387, 2490, 2563, 1271, 2324, 1388, 1389, 2288, 52, 1396, 784, 2220, 182, 2214, 154, 145, 2299, 2590, 1390, 1365, 2221, 395, 2397, 2218, 1470, 1392, 2049, 2534, 940, 2215, 20, 1393, 161, 2213, 1041, 137, 2415, 313, 2455, 822, 2216]. Destructive [1902, 753]. detect [1091, 391]. Detecting [271, 275, 374, 1166, 1779, 2528, 721, 1467, 1923, 1200, 990, 157, 1740]. Detection [2267, 866, 1586, 1043, 2435, 1252, 1687, 2436, 360, 212, 43, 1494, 1071, 243, 204, 246, 943, 953, 2238, 2073, 2421, 1555, 2384, 1281, 501, 292, 1230, 1320, 390, 598, 1889, 2447, 2600, 638, 1459, 1284, 1549]. Determinant [85]. determination [823, 35, 131, 582, 775, 1066, 2296, 296, 1536, 1438]. determine [1255]. determining [250, 2536]. deterministic [1841]. deterrence [1262]. DetMM [1622]. DetS [1622]. developing [1049]. Developments [2212]. deviance [1980, 1156, 159]. deviation [1767]. device [658]. diagnosis [1276]. Diagnostic [1831, 750, 50, 1702, 1655, 2123, 1239, 1115, 905, 1679, 2281, 290, 713, 390, 99, 723]. Diagnostics [948, 795, 364, 1226, 474, 430, 1006, 334, 1221, 629, 1866, 718, 1072]. DiceOptim [1378]. dichotomized [825]. dichotomous [2012]. Dickey [1570, 201]. dictionaries [2527]. DIF [1071]. difference [190, 2146, 1239, 2573, 1536, 2138, 81, 1438, 454, 2317]. differenced [2075]. differences [1610, 707, 223, 410, 2399]. different [1258, 331]. differential [1862, 2409, 2571, 254, 1663, 439, 1677]. differentially [391]. diffusion [58, 2043, 728, 339, 2447, 1003, 930, 1346]. digital [1031, 573]. digitised [2522]. Dimension [1446, 1265, 105, 2166, 1761, 1488, 1698, 1352, 2394, 2387, 1138, 2127, 2380, 2192, 1896, 1356, 2084, 2124, 579, 2160, 2109, 993, 988, 1765, 189, 2028, 2252, 1082, 1894, 2615, 2419, 1827, 2033]. Dimensional [1349, 2471, 1625, 1904, 2198, 1307, 1597, 433, 838, 2407, 2311, 1723, 1816, 802, 777, 2470, 480, 1051, 550, 1176, 1203, 1169, 1353, 2229, 2368, 2566, 2238, 2296, 2073, 562, 2041, 2079, 2192, 1096, 2001, 2173, 2240, 1722, 2047, 2308, 2438, 1669, 2243, 1037, 789, 1637, 1884, 448, 439, 379, 1858, 1871, 1612, 102, 934, 508, 577, 2589, 1707, 2518, 1437, 1667, 1358, 2579, 213, 1665, 721, 1190, 616, 2070, 1556, 2234, 2485, 1845, 2122, 2395, 2034]. Dimensionality [349, 1135, 1611, 1033]. dimensions [2478, 604, 112, 790, 2092, 2458, 2284, 2574, 2570, 2440, 2526]. Direct [762, 634, 598, 206, 1375]. Directed [2463]. direction [944, 1587, 935, 2551, 105]. direction-mixed [944]. Directional [2473, 1485, 2508]. directions [2380]. directly [2516, 106]. Dirichlet [1376, 491, 1342, 2287, 42, 2176, 169, 822]. disconnected [1387]. discontinuities [2603]. discover [1999]. Discovering [1542, 1640]. discovery [757, 377, 1555, 525, 793]. discrepancy [1338]. Discrete [1415, 1717, 2174, 1794, 2270, 783, 996, 1215, 1150, 2408, 1892, 784, 1705, 531, 1253, 1955, 1094, 1097, 1188, 2096, 2121, 828, 1470, 982, 1730, 2170, 1697, 2018]. discrete-continuous [2018]. discrete-state [1730]. discrete-time [784, 1470, 1730]. discrete-valued [982]. discretely [58, 2043, 2612]. discretised [2553]. Discretization [1375]. Discretization-based [1375]. discretizing [335]. Discriminant [1276, 326, 1528, 2278, 2392, 779, 2566, 1896, 518, 1917, 841, 1186, 1778, 955]. discriminating [2376, 2211, 20]. discrimination [1754, 749, 2212]. discriminative [46]. disease [734, 2209, 716, 1015, 1549]. diseases [1716, 2207]. disentangling [864]. disequilibrium [425]. disorder [623]. disparity [1146, 643]. dispersion [354, 205, 222, 356, 2285, 617, 1916, 952, 1634, 2017]. Dissimilarity [842, 1923, 347, 146]. dissimilarity-based [1923]. Distance [45, 483, 160, 803, 2539, 973, 2311, 799, 2050, 591, 817, 1668, 2433, 1638, 2268, 2513, 2607, 1044]. Distance-based [45, 160, 817, 2433]. distance-constrained [1668]. distances [1167, 221]. distortion [1068, 1653, 2182]. Distributed [69, 342, 2282, 1247, 260]. Distribution [821, 1251, 2056, 1968, 1372, 230, 2545, 1834, 560, 658, 770, 2120, 2168, 2396, 1447, 126, 2343, 1440, 372, 575, 2129, 152, 897, 2598, 440, 2601, 646, 1492, 86, 1215, 739, 1631, 193, 95, 1143, 2309, 2408, 1554, 956, 1161, 92, 199, 186, 1063, 1428, 2081, 902, 513, 149, 992, 1914, 1287, 522, 1123, 1318, 520, 654, 1664, 2525, 536, 2046, 1678, 1575, 388, 40, 289, 651, 1042, 880, 909, 150, 1339, 1788, 673, 413, 1908, 94, 1884, 732, 1858, 167, 1214, 257, 595, 84, 369, 1120, 1007, 93, 2589, 816, 625, 1312, 42, 20, 1595, 2372]. distribution [10, 958, 671, 1651, 1210, 1521, 2417, 24, 516, 1766, 1562, 1311, 2241, 605, 221, 1658, 1796, 1985]. Distribution-free [821, 2601, 2525, 2046]. distributions [266, 1323, 767, 358, 1321, 269, 1101, 2174, 624, 427, 364, 650, 2596, 1794, 5, 59, 2602, 429, 540, 1424, 2342, 1313, 2202, 675, 1166, 1860, 901, 1705, 1857, 1532, 531, 1157, 375, 2001, 1777, 2571, 1979, 936, 2243, 768, 2010, 878, 1097, 438, 2498, 2008, 72, 1340, 1714, 1040, 2103, 603, 1257, 1053, 4, 133, 1736, 109, 502, 460, 373, 387, 2244, 2461, 169, 1459, 353, 422]. disturbances [432]. divergence [608, 1598, 2110, 2573, 2365, 1466]. divergences [166]. diverging [1469, 1881]. diversified [728]. diversity [208]. divisive [842]. DNA [757]. do [1258]. document [676]. dollar [232]. Domain [1666, 1467, 715, 1190]. domains [27, 465, 855]. dominance [1366]. Donoho [661, 420]. dose [2169, 705, 1752, 2415]. dose-response [2415]. Double [1634, 2443, 70, 1862, 1221, 123, 384, 1697]. double-sampling [123]. Doubly [788, 840, 587, 976, 749, 683, 1102, 1808, 185, 2116, 1807, 719, 184, 605]. doubly-censored [184]. Dow [228]. down [1123, 93]. Dr [700]. Drift [1000, 2118, 1346]. drifts [954, 563]. driven [1934, 190, 503, 2590, 2312, 2084, 1863, 129]. drop [82]. drop-outs [82]. dropout [1255]. drug [1382, 2288]. DSGE [1325]. Dual [2287, 165, 1692, 2441, 2212]. Dual-semiparametric [2287]. dual-tree [165]. due [1420, 216]. duplication [599]. duration [1654, 1691, 1585, 1212]. duration-dependent [1654]. durations [2400, 1856, 2480]. Durbin [615, 1760]. during [1480]. dwell [375]. dwell-time [375]. Dynamic [898, 874, 1977, 1514, 1286, 1990, 495, 1511, 1479, 1267, 788, 2354, 231, 1932, 2413, 879, 1440, 1289, 2462, 866, 2226, 2452, 831, 868, 245, 899, 1947, 463, 2360, 533, 1989, 910, 1504, 2583, 2352, 1954, 431, 2605, 1338, 8, 2121, 1973, 458, 1089, 762, 1960, 1517, 2333, 1244, 1978]. dynamical [2409, 2483]. dynamics [352, 873, 2101, 637]. DySco [1479].

earlier [128]. Early [211, 2169, 44, 712]. earthquakes [1028]. easily [596]. ECB [627]. ECF [1588]. ECG [1276]. echelon [1424]. ECM [1612, 1559, 127]. ECM-based [127]. ecology [2066]. Econometric [871, 879]. Econometrics [227, 1966, 1976, 884, 262, 863, 1476, 1930]. economic [1482, 1329]. Edge [1281, 507, 1414]. edges [1542]. Edgeworth [962]. Editorial [2274, 2290, 2302, 2313, 2348, 2359, 2377, 2497, 2470, 679, 6, 209, 225, 263, 29, 60, 78, 115, 136, 163, 180, 196, 314, 558, 569, 584, 396, 418, 445, 477, 489, 505, 529, 538, 612, 846, 861, 916, 632, 652, 677, 697, 724, 791, 806, 826, 971, 972, 1023, 1058, 1059, 1076, 1077, 1092, 1093, 1110, 1111, 1126, 1140, 1162, 1173, 1192, 1216, 1245, 1269, 1300, 1394, 1417, 1433, 1452, 1474, 1523, 1551, 1571, 1593, 1615, 1628, 1647, 1671, 1683, 1700, 1711, 1712, 1728, 1747, 1758, 1768, 1783]. Editorial [1823, 1851, 1872, 1885, 1899, 1909, 1928, 1981, 2005, 2014, 2044, 2064, 2085, 2098, 2112, 2125, 2142, 2155, 2171, 2194, 2231, 2245, 2246, 2322, 2336, 2389, 2401, 2422, 2449, 2459, 2467, 2486, 2503, 2524, 2537, 2546, 2556, 2575, 2592, 2606]. educational [1061]. Effect [1868, 91, 949, 1085, 739, 511, 2187, 392, 706, 179, 2041, 1741, 1287, 1185, 1599, 1956, 1958, 2027, 1264, 2560, 2170, 776, 2526, 1576, 1943]. effective [237, 784, 1752, 1431, 2178, 1765, 1990]. effectiveness [1838]. effects [1381, 1012, 2545, 589, 183, 1726, 205, 1326, 2298, 317, 1969, 1304, 2286, 646, 2226, 1655, 551, 1402, 1539, 868, 1152, 2254, 1260, 1902, 2050, 374, 2558, 2184, 1429, 1154, 182, 1282, 1157, 202, 2189, 23, 742, 1604, 1458, 1954, 455, 1779, 2147, 2097, 2571, 2525, 854, 620, 1795, 667, 2554, 1006, 1370, 1234, 2026, 1907, 764, 355, 334, 439, 1959, 1612, 472, 1772, 947, 450, 104, 1393, 118, 408, 832, 2415, 77, 664, 1840, 2358]. efficiencies [1756]. efficiency [998, 782, 967, 1146, 1668, 332, 395, 46]. Efficient [740, 942, 634, 1933, 893, 1422, 2327, 606, 2585, 33, 245, 2043, 1139, 965, 908, 1061, 1505, 1096, 254, 1425, 259, 1561, 1614, 1766, 422, 1761, 890, 2436, 1377, 2286, 1424, 278, 1387, 2443, 1389, 2273, 1752, 562, 1288, 681, 2441, 781, 1635, 2388, 2131, 1727, 727, 1007, 1564, 2340, 1403]. Efficiently [321]. EGARCH [1498]. eigen [97]. eigen-decomposition [97]. eigenvalue [1316]. elderly [1716]. electricity [874, 252]. electron [344]. element [339]. elements [1903, 221]. Eliminating [1420]. ellipse [840]. ellipses [504]. elliptical [1151, 334, 1985, 1595]. EM-like [142]. EM-type [2158]. email [676]. embarrassing [2248]. embedded [2574]. embedding [1297]. embeddings [2050]. emphasis [2068]. Empirical [639, 234, 1643, 2586, 2569, 1579, 1807, 53, 454, 2495, 1183, 1932, 2418, 596, 2025, 2542, 666, 1673, 2273, 965, 1117, 901, 16, 1691, 726, 1777, 1112, 1034, 586, 525, 2277, 1158, 2162, 2483, 104, 2582, 1278, 2247, 2521, 1224, 2317, 1845, 1767, 2603]. employment [1292]. Emulating [2605]. emulation [458]. emulators [949]. encompassing [937, 201]. endogeneity [878]. endpoint [1737]. endpoint-inflated [1737]. endpoints [784, 1086, 1465, 1553, 2337, 1470, 556]. energy [1500]. Engel [237]. enhanced [102]. Enhancement [405]. Enron [170]. Ensemble [2471, 482, 148, 2028, 1649, 1035]. ensembles [467, 1362]. Entropy [131, 1201, 1691, 615, 582, 1735, 1639, 547, 978, 2212, 1714, 502, 53, 1591]. Entropy-based [1201]. environment [809, 75, 2455]. environmental [1133, 953, 370]. Epanechnikov [2270]. epidemic [455, 1334, 1016]. epidemic-type [455]. epidemics [1614]. epidemiology [2376, 710]. epigenetic [113]. EPMS [1428]. Epsilon [960, 941]. equal [183]. equality [19, 220, 2361, 1777, 1602, 1335, 1405, 711, 1884, 645, 694, 104, 428]. Equating [219, 1746]. equation [1226, 511, 2200, 2350, 901, 2409, 2329, 1704, 2164, 1973, 2138, 947]. equations [1351, 622, 1618, 2394, 2151, 2087, 1073, 2571, 254, 1663, 1677, 2340, 519, 1633, 335, 475, 718, 2582, 2521, 2078]. equicorrelation [1977]. equispaced [1321]. equity [1951]. equivalence [365, 1526]. Erratum [2063, 1075, 651, 527, 133]. Error [360, 857, 272, 650, 236, 800, 1580, 2032, 186, 1069, 2350, 283, 403, 284, 902, 2329, 853, 2209, 909, 1831, 1243, 1876, 1423, 1207, 1879, 1861, 1195, 1013, 1595, 2075, 1662, 1210, 1881, 346, 1569, 2395, 515, 2035]. error-based [2032]. error-prone [1069]. errors [2106, 1151, 2090, 1531, 1493, 1116, 2036, 908, 875, 948, 1972, 2095, 619, 2279, 127, 10, 1165, 1187, 1249, 580, 1068, 1653, 2182, 1745]. errors-in-variables [2090, 580, 1653]. escalation [705]. estimability [1289]. estimate [2390, 1421, 999, 1675, 1632, 370, 140, 2241]. estimated [619, 715, 2326, 161]. estimates [1479, 466, 2148, 1051, 1429, 288, 1264, 1819, 980, 1438, 169]. Estimating [1487, 367, 2407, 464, 1363, 2101, 92, 728, 2057, 2588, 1239, 2595, 1185, 2147, 2097, 574, 856, 1908, 828, 409, 2589, 217, 10, 1739, 515, 473, 968, 1351, 2129, 2298, 622, 1064, 2394, 1802, 2200, 1706, 2151, 2288, 1596, 921, 2087, 1073, 2329, 1679, 1390, 642, 386, 150, 1704, 1310, 1536, 2340, 2256, 519, 1633, 335, 475, 718, 2582, 2521, 2078, 2480]. Estimation [399, 1932, 2259, 54, 542, 2007, 364, 1124, 1590, 2255, 241, 2476, 550, 366, 402, 1060, 1581, 155, 1500, 1115, 855, 875, 2522, 411, 2297, 1971, 2573, 2593, 660, 858, 1106, 2130, 764, 525, 738, 1214, 469, 816, 1879, 1963, 1736, 460, 459, 37, 1224, 2193, 2541, 1068, 2182, 2429, 1877, 2603, 1775, 2533, 885, 1266, 1967, 1761, 2056, 1259, 1012, 83, 773, 1372, 1481, 628, 923, 2561, 684, 2055, 2283, 1670, 968, 1674, 2120, 1850, 233, 892, 2100, 2197, 1133, 2090, 2594, 746, 235, 864, 31, 1806, 850, 89, 650, 848, 51, 852, 2391, 55, 212, 1762, 2327, 973]. estimation [1450, 1794, 606, 701, 987, 1490, 1643, 2190, 2418, 1835, 1605, 317, 2602, 1491, 27, 381, 2379, 946, 47, 1816, 2104, 237, 64, 1984, 783, 357, 867, 2226, 1297, 1682, 1539, 540, 1404, 1215, 1920, 849, 465, 1926, 2023, 2177, 242, 1150, 1567, 2013, 851, 1116, 588, 2460, 870, 1152, 1260, 1781, 430, 2039, 1921, 641, 2549, 2309, 2408, 1136, 208, 139, 794, 1746, 1801, 800, 2089, 2332, 1580, 2324, 2443, 566, 1673, 245, 899, 1122, 1946, 2019, 247, 49, 1462, 1870, 453, 2260, 307, 165, 953, 1139, 675, 1993, 2465, 2040, 2558, 363, 908, 2519, 2296, 608, 248]. estimation [1061, 591, 1752, 1503, 951, 284, 476, 533, 1125, 983, 1825, 910, 1504, 1857, 744, 182, 1506, 992, 905, 350, 709, 1288, 351, 528, 2583, 2001, 681, 1318, 1911, 2161, 906, 2053, 1954, 2441, 2301, 2567, 1135, 966, 435, 1172, 1721, 1083, 854, 111, 2076, 276, 781, 1319, 1627, 1227, 421, 432, 1400, 1146, 1678, 401, 853, 1050, 2131, 2474, 1354, 1455, 1972, 637, 2212, 1370, 553, 1102, 1184, 1170, 1042, 1577, 2008, 1704, 1588, 1994, 924, 1, 527, 614, 682, 1793, 310, 545, 2121, 2277, 254, 1663, 1229, 1529, 1771, 1973, 439, 2077, 959, 1030, 1876, 645, 1423, 688, 1714, 2609]. estimation [1022, 2124, 311, 1275, 2030, 2268, 672, 2279, 257, 714, 1425, 206, 1960, 1408, 2116, 2328, 1207, 2185, 2484, 1866, 1220, 2604, 625, 2082, 922, 341, 2011, 1924, 980, 1043, 609, 2320, 1677, 2340, 1809, 1686, 1817, 537, 48, 1403, 376, 1742, 312, 1559, 1397, 302, 502, 1003, 2607, 118, 454, 958, 1165, 1607, 1568, 2088, 1833, 1662, 1210, 304, 859, 169, 2479, 1001, 1881, 776, 1459, 2599, 1052, 2234, 2456, 2548, 2362, 1044, 422, 1067, 1569, 1697, 2037, 2427, 975, 1745, 2395, 97, 2507, 13, 1456, 2612, 2034, 1346]. estimations [1179, 152]. Estimator [639, 1296, 1968, 1011, 998, 414, 648, 649, 1947, 1970, 1428, 2293, 1166, 1598, 1820, 514, 194, 1922, 206, 85, 1534, 2466, 661, 260, 1347, 1520, 1709, 420]. Estimators [857, 2106, 1511, 1420, 2227, 795, 1702, 2168, 1350, 756, 2535, 607, 2063, 554, 1618, 1424, 480, 278, 41, 2254, 1646, 343, 2307, 1735, 1622, 745, 1860, 2508, 1504, 1836, 1182, 1998, 1080, 1128, 853, 2587, 856, 1959, 2268, 39, 1564, 2165, 835, 1828, 1667, 2075, 353, 605]. euro [232]. euro/dollar [232]. European [627]. evaluate [35, 1841]. evaluating [757, 2318, 1829]. Evaluation [954, 1599, 890, 1485, 2598, 596, 1410, 2254, 1211, 1553, 888, 2170, 2035]. evenness [208]. event [1267, 1608, 2283, 1995, 1009, 702, 2201, 15, 2187, 2451, 1184, 1465, 657, 1780, 556, 1212, 1601, 1733]. events [1118, 639, 2198, 330, 568, 1733, 644]. Evidence [1027, 234, 714]. evidence-based [714]. evidential [690]. evolution [2574]. evolutionary [1330, 69, 1573, 294]. evolving [1190]. EWMA [222, 1738, 563].

Exact

[326, 2423, 240, 1901, 247, 52, 709, 984, 1526, 981, 2534, 609, 719, 1041, 420, 1967, 319, 86, 1821, 1389, 179, 1675, 760, 2010, 2337, 1889]. Exactly [1078]. examination [1797]. Examining [2250, 1262]. example [2346]. Excel [114]. excess [2244]. exchange [1509, 232, 1392, 162]. exchangeabilities [1344]. exchangeable [749]. exciting [2434]. exemplified [185]. exercise [98]. expansion [1749]. expansions [910, 962]. Expectation [2029, 2522, 1874, 1801, 920, 1761]. Expectation-Maximization [2029, 1874, 1801]. expectations [873]. expected [897, 1825, 681, 1086, 1046, 1419]. expectile [2323, 2137, 1825, 680, 2269]. expenditure [856]. expensive [1841]. expensive-to-evaluate [1841]. experiment [170]. Experimental [1382, 473, 2262, 2288, 1390, 1365, 395, 2397, 2223, 1274, 1772, 793]. Experiments [2210, 1608, 176, 27, 447, 996, 1374, 581, 1200, 1380, 1389, 1060, 1056, 336, 2222, 2253, 2149, 1391, 287, 1392, 940, 138, 1255]. experts [1799, 1338]. explaining [1049, 769]. Exploratory [1867, 1574, 251, 1419, 2559, 302]. exploring [1971, 887, 2382]. Exponential [1372, 658, 426, 2396, 1447, 372, 152, 274, 21, 2069, 1161, 199, 1892, 1950, 2110, 1123, 536, 936, 2114, 1336, 2268, 2333, 1736, 1347, 130, 1241, 353]. exponential-family [2069, 2110]. exponential-type [1123]. exponentiality [1756]. exponentially [534, 821, 954]. exponentiated [100]. ExPosition [1407]. exposure [1432]. exposures [898]. expressed [391]. expression [1304, 943, 732, 192, 1186, 485]. Extended [2333, 1519, 1440, 92, 1144, 388, 1019, 1053]. Extending [202, 2092]. extension [354, 514, 369, 1001]. extensions [1082]. external [1693]. Extracting [1029]. Extraction [1025, 2319, 96, 1027]. extrapolation [74, 1734]. extremal [1690]. extreme [436, 2482, 1646, 500, 880, 1577, 1908, 938, 1343, 460]. extreme-value [436]. extremes [737, 2432, 2384, 1871, 1342, 1856].

F [1447, 609]. face [1211]. facilitate [1907]. Factor [2425, 1275, 684, 1325, 1541, 2298, 1969, 1590, 2452, 1200, 1921, 1581, 2273, 951, 1604, 2264, 431, 276, 2183, 2243, 553, 1547, 2373, 300, 1514, 577, 1924, 302, 1842, 1665, 2613, 1001, 2191, 1409, 248]. Factor-adjusted [2425]. factor-augmented [1325]. Factor-GMM [248]. Factorial [177, 2216, 581, 2512, 2563, 1056, 2220, 1582]. factorials [356]. factorizable [2323]. factorization [309, 2448]. factors [1483, 1621, 273, 1540, 874, 52, 2558, 309, 1335, 598, 2191, 552]. failure [141, 2106, 1420, 107, 669, 987, 1124, 2031, 2431, 1867, 2424, 601, 92, 199, 1123, 780, 2514, 1020, 389, 2021, 369, 93, 816, 517, 2378, 778, 142, 173, 564, 833, 174, 1087, 442]. failure-time [141]. Faithfulness [1705]. false [757, 377, 1555, 525]. Familial [195, 670]. families [425, 830, 1944, 944]. Family [1466, 1197, 499, 2069, 1892, 2110, 1585, 2243, 2365, 2608, 2103, 2333, 168, 1347, 353]. Fan [1293]. FAST [331, 2315, 2535, 1980, 1277, 1328, 1770, 2236, 2369, 1441, 165, 1887, 284, 9, 1227, 448, 1871, 2609, 657, 1527, 1564, 1828, 1743, 1687, 2343, 735, 2539, 1464, 1103, 2089, 1821, 308, 34, 1114, 333, 1588, 1994, 1730, 2070, 2356, 318]. fastKDE [1994]. FastPCS [1250]. fat [1498, 288]. fatigue [440, 2223]. Fay [1850, 1050, 1290]. FDR [2229, 963, 1905]. Feature [362, 2438, 2518, 2012, 2471, 18, 2311, 1352, 1363, 2566, 2079, 2240, 1205, 1844, 290, 1186, 2338]. feature-screening [2566]. features [623, 171, 1353, 2038, 1198, 170]. feedback [1242, 1943]. Fernández [575]. fewer [1148]. FFT [2089, 2004]. FFT-based [2089]. fibre [2223]. fibre-reinforced [2223]. Fiducial [1097, 1364, 487]. Fiducial-based [1097]. FIEGARCH [1236]. field [368, 2464, 604, 2224, 1305, 1430, 1693, 1243, 129]. fields [1398, 2437, 1652, 2002, 744, 618, 2356]. Fifth [227]. filling [1461]. filter [1266, 1116, 1497, 1657, 355, 339, 1454, 619, 2338]. Filtering [1025, 1026, 1829, 1034, 786, 919]. filters [1035, 342, 2282, 2167]. finance [903, 262]. Financial [863, 1476, 1930, 627, 1478, 1480, 234, 1442, 1064, 896, 243, 1428, 1585, 1128, 907, 8, 1518, 887, 1964, 2480]. find [2221]. Finding [67, 1751, 1247, 1469, 1250, 2310, 1047, 935, 2397, 1636]. Finite [126, 1795, 877, 672, 2502, 1717, 931, 2174, 701, 1792, 242, 2254, 1313, 2307, 2508, 1853, 984, 1979, 410, 1668, 1427, 339, 1310, 300, 1787, 2500, 2417, 776, 353, 2335, 32]. Finite-sample [126, 877, 242, 2508, 1668, 32]. first [863, 1260, 1385, 910, 1504, 514, 816, 169, 778]. first-failure [816, 778]. first-order [910, 1504]. Fisher [126, 33, 1821, 779, 1599, 1958, 156, 2170, 168]. fit [840, 964, 830, 1167, 995, 5, 59, 1723, 277, 796, 95, 1882, 186, 2346, 347, 547, 978, 1253, 178, 410, 586, 1797, 2608, 84, 1812, 53, 671, 304, 2578, 1743]. Fitting [601, 2179, 829, 872, 2071, 804, 1687, 1175, 2342, 1272, 504, 541, 356, 920, 657, 2530, 762, 860, 1312, 1204]. five [984]. Fixed [589, 1842, 408, 1012, 144, 2226, 2187, 179, 1954, 1751, 1612, 456]. Fixed-effects [408, 1012]. fixed-point [1751]. fixed-size [144]. flaws [114]. Flexible [1252, 1806, 2138, 2340, 2502, 1197, 542, 2369, 2036, 145, 500, 2577, 298, 598, 1916, 1516, 450, 502, 1753, 1569]. Floating [391]. flows [1049]. fluctuation [1521]. fluorescence [285, 2163]. fMRI [2323]. focused [1391]. folding [2028]. follow [587, 1206, 2208]. follow-up [587, 1206, 2208]. following [1264]. food [856]. forecast [1485, 1936, 1295, 261]. Forecasting [1480, 889, 1952, 249, 1510, 252, 631, 1471, 1325, 1935, 2565, 894, 897, 1654, 1969, 1371, 1945, 2510]. forecasts [228, 543, 1887, 888]. foreign [232]. forest [2471, 1878, 574]. Forests [1150, 1081, 1606, 1189, 1750]. form [2598, 798, 239, 240, 1293, 1998, 880]. formal [1374]. formative [2319]. forms [86]. formulas [306]. formulation [2441]. formulations [1439]. forward [329, 2374]. Founder [2468]. four [984]. Fourier [2211, 1520, 331]. fractality [249]. fraction [669, 987, 1584, 2611]. fractional [788, 2512, 1203, 2260, 2511, 98, 450]. fractional-norm [1203]. fractionally [1508]. fragmentation [1730]. Frailty [726, 1995, 1124, 2445, 1660, 407, 142, 1292]. frame [1464, 2466, 833]. framework [1796, 2262, 1374, 44, 2610, 1000, 2178, 1516]. frameworks [1505]. free [1931, 1167, 2601, 540, 821, 1829, 2079, 2525, 2046, 2440, 912, 1521]. freedom [1692, 2372, 664]. Freeman [2428]. frequency [1848, 1467, 514, 1432, 2223, 2484, 715]. Frequentist [1882, 301]. Fridley [80]. friendly [1377]. front [1391]. frontier [322, 2082]. frontiers [913]. FS [818]. FSAVE [1765]. FSIR [1765]. FSR [492]. Full [1813, 566, 999, 57]. Fuller [1570]. fully [1772]. Function [2265, 228, 2343, 746, 1422, 1450, 1835, 1792, 1984, 607, 2063, 464, 1296, 114, 1063, 2043, 728, 903, 541, 2261, 1123, 2062, 1678, 401, 2212, 627, 1036, 785, 1714, 1425, 84, 424, 922, 537, 2579, 10, 303, 671, 2479, 2532, 644, 1743]. function-based [903]. Functional [557, 2048, 2249, 2030, 665, 1283, 1582, 509, 803, 1259, 2227, 546, 45, 89, 825, 2392, 2379, 2473, 2476, 2470, 1051, 2482, 2069, 22, 224, 693, 1750, 2361, 1970, 2040, 1309, 541, 496, 2451, 2477, 2097, 461, 2458, 2159, 1168, 597, 2474, 602, 1566, 2344, 2472, 738, 1033, 1423, 1411, 1419, 2068, 1207, 805, 2353, 993, 1098, 1846, 1206, 1765, 1891, 2475, 2479, 1992, 2548, 1284, 2488, 1990, 2163]. Functionally [1396]. functionals [2443, 1183]. functions [88, 1864, 1308, 635, 2186, 2270, 1464, 349, 447, 483, 1492, 2547, 2025, 1620, 1706, 1746, 2361, 2000, 1847, 308, 2519, 1182, 2341, 1777, 981, 2374, 715, 1206, 1880, 2480]. funds [898]. Further [86, 42]. Fused [2338, 1587, 1868, 1839, 2551, 451]. Fusing [2600]. Fusion [171, 170, 786, 2513]. future [729]. Fuzzy [691, 693, 696, 695, 689, 694, 1534].

G [1447]. gains [1478]. galaxy [974].

Gamma [1985, 2073, 1247, 1660, 2106, 1251, 1124, 407, 1280, 142, 2445, 2010]. Gamma-frailty [1660, 407, 2445]. gamma-normal [1251]. GAMs [134]. gap [1661, 809, 644, 75]. GARCH [1961, 1481, 1933, 235, 1487, 1488, 1490, 1491, 867, 1496, 246, 902, 875, 1953, 877, 257, 1514, 1517, 2325, 1963]. GARCH-type [1487]. garrote [1685]. gas [2073]. Gauss [2295, 970]. Gaussian [1967, 2411, 949, 1179, 2464, 1422, 1800, 1652, 1983, 2292, 813, 1676, 2039, 1921, 2236, 244, 329, 2036, 2229, 2368, 463, 829, 2002, 1953, 844, 2522, 2345, 333, 2092, 1281, 918, 2574, 1837, 2095, 71, 731, 2605, 1234, 734, 1599, 1788, 416, 409, 881, 657, 2030, 1247, 1787, 2328, 469, 1686, 2560, 618, 2170, 129, 2475, 90, 2356, 498]. Gaussianity [1463]. Gaussians [1753]. GCV [2279]. GDP [1952, 1891]. GEE [1435, 1688, 1530, 1255]. geepack [1530]. Gegenbauer [1938]. Gene [479, 1304, 732, 192, 508, 1344, 1186, 721, 485, 2455]. gene-environment [2455]. GeneChip [3, 80]. General [1796, 1917, 228, 436, 2442, 27, 65, 2324, 2576, 307, 872, 1836, 824, 410, 2072, 747, 2277, 736, 39, 2614, 1645, 2582, 2078]. general-purpose [736]. generalisation [1011]. Generalised [1152, 2220, 1855, 1888, 296, 626, 2213]. generalization [371, 41, 193, 201]. generalizations [21, 40]. Generalized [767, 2594, 589, 1915, 5, 59, 2424, 695, 2154, 2444, 2087, 1955, 1722, 431, 936, 2453, 1128, 880, 545, 1757, 1737, 519, 718, 958, 2521, 1634, 346, 572, 353, 664, 1990, 1118, 88, 2106, 2533, 810, 964, 1834, 466, 1171, 1351, 270, 753, 1468, 571, 2327, 152, 452, 1124, 2104, 148, 867, 484, 131, 582, 2023, 2460, 103, 2151, 1944, 2610, 1364, 1692, 1117, 2238, 377, 1429, 2371, 1639, 742, 1073, 1604, 2341, 2511, 333, 1107, 1611, 2590, 2525, 1635, 2003, 521, 289, 651, 1234, 1896, 293, 2051, 1416, 1908, 1221, 94, 1638, 1235, 672, 140, 369, 56, 2185, 860, 2523, 2049, 947]. generalized [625, 2333, 4, 133, 961, 952, 1013, 1633, 765, 335, 1443, 1072, 2461, 2402, 169, 2012, 76, 1368, 668, 105, 1087, 1918, 1658, 2010]. generally [984]. generated [767]. Generating [2295, 1376, 333, 2262, 596, 2563]. generation [2220, 1375]. Generative [1098, 1038, 1156, 46]. genes [391]. genetic [1869, 1555, 14]. genetics [2383]. GENLIN [1530]. GENMOD [1530]. genome [703, 1268, 552]. genome-wide [703, 1268, 552]. genomic [2173]. genomics [1867]. genotype [2146]. genotype-phenotype [2146]. genotypes [425, 3, 80]. Geoadditive [680]. geometric [1447, 865, 212, 357, 1554, 2304, 536, 373]. German [850]. GH [909]. Gibbs [2141, 1933, 2437, 467, 2366, 357, 1396, 685, 1341]. GICC [1739]. Gini [1926, 1974]. Gini-based [1974]. given [1895]. GLMM [470, 2420]. Global [713, 2580]. Globally [267, 2305]. globular [974]. GMM [1947, 1946, 248, 903, 1506, 1954, 1960, 1866]. gold [823, 1239]. Gompertz [2101, 816]. good [869, 2563]. Goodness [186, 2346, 1253, 178, 410, 586, 2608, 1812, 830, 1167, 5, 59, 277, 796, 1882, 84, 53, 671, 2578, 1743].

Goodness-of-fit [186, 2346, 178, 410, 586, 2608, 1812, 830, 1167, 5, 59, 277, 796, 1882, 53, 671, 2578, 1743]. GPS [1028]. GPU [1348]. Gradient [2198, 999, 105]. grading [250]. Graph [1884, 2491, 2552, 280, 171, 2321, 1219, 1291, 2110, 170, 2268]. Graph-theoretic [1884]. Grapham [12]. Graphical [2570, 12, 32, 1150, 813, 2229, 829, 1281, 1188, 1004, 2328, 1686, 75, 1739, 2362]. graphs [2132, 171, 170]. greater [1918]. Greedy [2457]. Green [1118]. grid [1057, 2412, 320]. gridded [112, 790]. Group [1457, 1083, 2347, 640, 964, 2385, 367, 1558, 1159, 2146, 1741, 2173, 1587, 2152, 1868, 2051, 2399, 577, 1358, 1048, 2509]. group-parallel [1159]. Grouped [1750, 1546, 402, 1887, 595, 860, 1461]. grouping [2552, 1160, 2281]. groupwise [1818]. growth [755, 1482, 1941, 1952, 443, 1319, 410, 1306, 1022, 406, 1559]. GSCANO [1990]. Guaranteed [2446]. Guide [981]. guided [2169, 1605, 1809].

h.c [700]. habitat [134]. HAC [2307]. Hadamard [2260]. Haenszel [835]. half [560, 960, 729, 461, 94]. half-logistic [560]. half-normal [960, 729, 94]. half-region [461]. Halfline [1008]. halfspace [1901]. haplotype [497]. hard [1341]. hard-core [1341]. hardcore [357]. Hastings [1563, 726, 1007, 259]. Hawkes [2357, 1833]. hazard [1813, 923, 2186, 702, 566, 1532, 2059, 1445, 1337, 1704, 1229, 2242, 84, 1120, 2185, 1809]. hazard-based [702]. hazards [548, 107, 1703, 2031, 2190, 2289, 371, 1682, 2445, 1661, 976, 1458, 348, 957, 2567, 1135, 781, 2386, 1812, 1451, 1660, 2378, 1292, 1067]. HCV [551]. HDLSS [1754]. healthcare [1410]. heaped [2174]. heart [1639, 715, 2488]. heat [2482]. heavier [95]. heavy [266, 2056, 650, 801, 908, 1073, 948, 2571, 877, 909, 1961, 373, 387]. heavy-tailed [266, 650, 801, 908, 1073, 948, 2571, 877, 909, 373, 387]. hedge [898]. Hellinger [591]. Helmert [970]. Herriot [1850, 1050, 1290]. heterogeneity [98, 647, 822, 1859]. heterogeneous [571, 779, 1869, 1499, 2517, 2595, 2591, 2370]. heteroscedastic [788, 650, 474, 1288, 812, 10, 1187, 1745, 1975]. Heteroscedasticity [2033, 752, 544, 798, 1385, 2114, 130, 1241]. heteroskedastic [2060, 1249]. heteroskedasticity [240, 1282, 1513]. heuristic [1952, 2397]. Heuristics [858, 250]. HHCART [1875]. Hidden [375, 63, 2145, 783, 1200, 1949, 1227, 1837, 1693, 2072, 1040, 1345, 2555, 2501, 1796]. Hierarchical [2561, 2141, 465, 359, 16, 1863, 2510, 989, 83, 1308, 1326, 352, 1706, 943, 543, 1887, 337, 1073, 2564, 1234, 57, 1907, 736, 1856, 159, 1854, 207, 2351, 1880, 1724]. hierarchical-likelihood [337]. hierarchically [313]. High [1349, 757, 1904, 2470, 2229, 2002, 1668, 2234, 2471, 1625, 2478, 1761, 2198, 1307, 433, 838, 2407, 2311, 1723, 1816, 802, 1352, 777, 1379, 480, 1051, 550, 1176, 1203, 1169, 1353, 2368, 2238, 2296, 562, 2041, 2192, 2001, 2173, 2398, 2092, 1135, 1722, 2047, 1669, 2458, 2284, 2243, 394, 1356, 2570, 2440, 1637, 1884, 448, 439, 379, 1858, 1612, 508, 2484, 577, 2589, 1707, 1437, 1667, 1358, 213, 1665, 721, 1190, 616, 2526, 2485, 1845, 1894, 2395, 2034]. high-dimension [1894]. High-dimensional [2470, 2471, 2198, 1307, 433, 2407, 1723, 1816, 480, 550, 1203, 1169, 1353, 2238, 562, 2041, 2001, 2173, 2047, 2243, 448, 1858, 508, 577, 2589, 1707, 1437, 1667, 213, 1665, 721, 1190, 616, 2485, 1845, 2395, 2034]. high-frequency [2484]. high-performance [1379]. High-throughput [757, 2398, 394]. Higher [910, 112, 2490, 790, 934, 1084]. Higher-order [910, 2490, 1084]. highest [533]. highly [1441, 2590, 760, 1065, 1564]. Hilbert [1925, 2252]. Hill [1011, 1166]. histogram [510, 351]. histogram-valued [510]. histograms [1583, 299]. historical [158, 2405, 921, 2204, 761]. history [1015]. HIV [352]. HMMs [1028]. Homogeneity [2238, 1101, 2244, 1280, 2417]. homogeneous [2101]. homoscedastic [291]. Hong [432, 1016]. Hopfield [2562]. Horizon [1956]. hot [1642]. hourly [874]. housing [432]. HPV [2177]. hsmm [63]. Huber [343]. Huberized [433, 661]. Human [3, 80, 769]. hurdle [1101]. Hybrid [986, 552, 2471, 1656, 433, 1631, 1472, 654, 370, 1214, 1736, 1765]. hydrology [458]. HYGARCH [905]. Hyper [504, 1179, 2039, 1105]. hyper-parameter [2039]. hyper-parameters [1179]. hyper-Poisson [1105]. hyperbolic [998, 867, 1944, 1896, 2263, 2461]. Hypercube [1350]. hypergeometric [822]. hypergeometric-Dirichlet [822]. hypergraphs [1705]. Hyperparameter [980]. hypertension [392]. hypnograms [1730]. hypotheses [2239, 550, 1892, 522, 1335, 1905]. Hypothesis [2500, 190, 799, 2368, 978, 1913, 2047, 1080, 1526, 768, 1575, 963, 713, 2242, 2182, 2429].

ice [1029]. ICS [2435]. identical [221]. identifiability [2104]. Identification [1912, 2272, 836, 2386, 2052, 1695, 2529, 2296, 1611, 290, 256, 102, 2347, 721]. identified [1716]. Identifying [2383, 496, 8, 1787, 1028, 798, 623, 291]. identity [166]. ideological [2145]. ignorable [2594, 2203, 2406, 1319, 2022]. II [560, 979, 429, 2517, 336, 149, 522, 712, 732, 167, 84, 2489, 1736, 719, 1347]. II/III [712]. III [712]. illnesses [1016]. illustration [933]. image [799, 662, 200]. imagery [809]. images [2179, 583, 618, 573]. imaging [2383, 2206, 2208]. imbalance [2157]. imbalanced [1904, 330]. immigration [1833]. Impact [77, 2281, 887]. imperialist [2221]. implementation [1377, 2261, 1378]. implementations [71, 168]. implemented [596]. Implementing [72, 1760, 712]. importance [937, 1750, 872, 1505, 760, 1355, 254, 1529, 318]. imprecise [1366]. imprecision [1359]. improve [203, 1822]. Improved [979, 480, 1946, 2282, 533, 1504, 1240, 125, 1146, 2010, 1079, 1, 527, 39, 1755, 1443, 962, 1511, 440, 648, 2307, 1047, 2329, 1686]. improves [1344]. Improving [1436, 782, 2157, 1734]. Imputation [282, 1252, 2423, 107, 2031, 1996, 317, 748, 1402, 2023, 2476, 524, 743, 1814, 809, 1357, 1642, 561, 1195, 1755, 216, 173, 77, 1919]. impute [2250]. imputed [218, 2151, 1776]. Imputing [2178, 1855, 1642]. IMSE [2224]. In-sample [627, 2565]. INAR [1935, 1209, 1543]. incidence [2341, 2076]. including [782]. Inclusion [1256, 1019]. Income [1979, 237]. incompatibility [1448]. incompatible [467]. incomplete [2424, 41, 1387, 743, 839, 1913, 2406, 1316, 1627, 494, 828, 1275, 2614, 2461, 216, 659, 1409]. incorporate [2405]. Incorporating [158, 1519]. Incorporation [1344]. increased [1028]. increasing [748, 465, 1123, 2242, 93, 1448]. incremental [1190]. increments [469]. independence [1420, 2601, 1493, 1895, 1240, 2284, 1188, 888, 2252]. Independent [974, 2509, 2039, 1180, 624, 2016, 1298, 1442, 1055, 1157, 966, 2010, 81, 1749, 1863, 719, 122]. index [2056, 2331, 151, 1222, 892, 2601, 1920, 775, 1646, 2288, 965, 1117, 1069, 2531, 728, 1153, 347, 709, 1288, 2345, 348, 2297, 2573, 2593, 2554, 637, 1577, 764, 1235, 1681, 1208, 609, 1607, 2247, 994, 1279, 1224, 1525, 2427, 1877, 2404]. indexed [1066, 1068]. indexes [510]. indicators [2400, 853, 2276]. Indices [82, 1797, 485]. Indirect [1487, 421, 881, 1506, 255]. individual [1049, 2558, 223, 1043]. individual-level [1049, 2558]. individualized [782]. Induced [957, 101, 1396, 682, 2397, 1880]. industrial [1848]. inequality [1979]. Inference [755, 1177, 575, 352, 2430, 371, 199, 1985, 706, 2409, 654, 40, 614, 1414, 595, 1401, 765, 1651, 2134, 1360, 1118, 2111, 740, 2331, 1813, 345, 560, 933, 1197, 979, 937, 1838, 2413, 126, 270, 2505, 753, 1487, 1289, 960, 51, 2585, 1698, 695, 2403, 2452, 1159, 240, 730, 588, 1143, 1554, 2369, 244, 1066, 171, 1364, 914, 2036, 470, 729, 2203, 2175, 535, 1688, 149, 1914, 1157, 23, 1073, 1987, 125, 1911, 2345, 2248, 1137, 2571, 1432, 876, 877, 1578, 1430, 1079, 1334, 1579, 393, 2440, 1538, 997, 1727, 2084, 912, 1243, 28, 170, 2162, 1340, 1411, 881, 2399]. inference [1021, 2280, 2483, 1120, 462, 1516, 2357, 1963, 952, 1013, 1742, 1443, 127, 814, 1187, 2521, 1753, 1766, 1525, 2495, 77, 323, 580, 1183, 1767, 1775, 592, 1658]. Inferences [316, 90, 965, 1117, 1776, 1613]. inferential [986]. inferiority [942, 2492, 2202, 2406, 568, 2370, 723]. inferring [1199]. infimum [1570]. Infinite [1499, 1196, 1721, 1036]. Infinite-order [1499]. infinite-variance [1196]. inflated [1101, 1412, 380, 2430, 1242, 430, 1688, 385, 1314, 1664, 747, 1916, 1737, 120, 147, 346]. inflation [1486, 241, 873, 1952, 1099, 1508]. Influence [1006, 1072, 122, 364, 2390, 430, 441, 2164, 334, 379, 1470, 400, 475, 2078]. influential [2267, 271, 275, 1166, 2123]. Information [551, 1258, 703, 1268, 1290, 572, 930, 1484, 1980, 158, 33, 1363, 342, 1952, 621, 1333, 1172, 1365, 642, 1366, 1599, 2365, 1310, 1022, 311, 2185, 2600, 1403, 1883, 2170, 138, 2347, 159, 2213, 1001, 1591, 516, 668]. informative [2190, 2108, 601, 174, 340, 975]. informatively [2378]. Informativeness [2281]. infrared [838]. inhomogeneous [2366]. Initial [2533, 1440, 1947]. Initializing [731]. INLA [1198, 2209, 762]. inlier [1146]. innovations [1487, 877, 249]. inpainting [200]. input [1383, 27]. input-dependent [1383]. insertion [1038]. instability [1969]. instance [1730]. instant [2293]. instrumental [2602, 1354]. instruments [248]. insurance [931]. insured [2494]. integer [1770, 1389, 1998, 950, 50]. integer-valued [1770, 1998, 950, 50]. Integral [1393]. integrals [1652]. integrated [1508, 762]. integrating [779]. Integrative [2051, 836]. integro [2138]. integro-difference [2138]. intensity [882, 894]. inter [367, 424]. inter-cluster [424]. inter-group [367]. interacting [1353, 1184, 736]. Interaction [1846, 2572, 367, 1049, 357, 1402, 1620, 1091, 1022, 721, 2455]. interactions [2298]. intercept [2520, 2057, 720]. intercepts [1318]. interdirections [1927]. Interest [1495, 586]. interlaboratory [472]. internal [250]. interpolation [576, 168]. Interpretable [993]. Interpretation [46]. interpurchase [868, 2400]. interpurchase-timing [2400]. Interquantile [1263]. intersect [1545]. Interval [1971, 376, 542, 1597, 107, 152, 1643, 2031, 2431, 2424, 607, 2063, 769, 1152, 2445, 666, 2271, 100, 2273, 49, 1870, 1805, 143, 1305, 1574, 533, 709, 1584, 2059, 2514, 2567, 2021, 38, 1, 527, 1449, 1129, 1681, 888, 1626, 2611, 609, 716, 48, 958, 2088, 2378, 2456, 174, 2358, 1087, 1919, 515, 13]. interval-censored [107, 2031, 2431, 607, 2063, 2445, 100, 1805, 1584, 2514, 2021, 1449, 2611, 2378, 174, 2358, 1087]. Interval-valued [1971, 2271, 1574, 38]. intervals [267, 219, 1473, 404, 1652, 1689, 21, 581, 1136, 15, 1862, 1239, 921, 1147, 984, 1097, 1273, 1427, 1957, 2027, 384, 1257, 1046, 2091, 2074, 1961, 4, 133, 106, 843, 1278]. intervention [1091, 1262, 1255]. Interview [2468]. interweaving [1503]. Intra [1600, 1689, 2038, 1739]. intra-class [1689, 2038, 1739]. Intra-cluster [1600]. intraclass [404, 1088]. intractable [2376, 1332]. Intradaily [231]. intrinsic [1621, 967, 1539]. Introduction [1211]. intuition [2248]. invariant [1573, 486]. invasive [1698]. Inverse [2228, 327, 1544, 1201, 2529, 149, 1073, 333, 2441, 1455, 1788, 81, 459, 1283, 90, 13, 2034]. inversion [1377, 333]. invertibility [1507]. Investigating [969, 1374, 32]. investigation [126, 1094, 1592]. Investigations [963]. ionosphere [344]. ionospheric [1066]. irregular [1415, 675, 299]. irregularly [759, 2472, 926, 1740]. irregularly-shaped [759, 1740]. IRT [345, 964]. ischemic [2488]. isotonic [2602]. Issue [227, 1815, 2487, 2497, 1786, 1476, 858, 699, 1349, 305, 1976, 863, 884, 1302, 1966, 848, 1322, 689, 1359, 1164, 1804, 2210, 2470, 1367, 679, 62, 1380, 1930, 2225, 1025, 98, 265]. Issues [1785, 1364, 356]. Italian [2188]. Item [1071, 1373, 2022, 592]. Iterated [2023, 1678]. iterations [1678]. Iterative [561, 64, 103, 421, 2003, 860, 1854]. Iteratively [1975]. IV [901].

Jackknife

[2418, 1845, 1767, 1777, 1414, 2317]. James [203]. Japan [252]. Jeffreys [2332]. JM [2139]. Joint [228, 802, 1763, 401, 2388, 1234, 416, 2328, 25, 1525, 2419, 230, 1687, 2431, 967, 1921, 2271, 392, 1680, 2306, 2451, 356, 1795, 1455, 617, 2344, 1512, 657, 2604, 1533]. joint-quantile [1680]. Jones [228]. judgment [648, 1538, 2193]. Jump [864, 1416, 2437, 1404, 444, 1907, 2603, 1346]. jumping [2420]. jumps [2101, 1953].

Kalman [1266, 1116, 1035, 1497, 1027, 1657, 786, 1454, 619]. Kaplan [1438]. Kappa [1701, 408, 1630]. Kendall [1420, 2543, 155, 2284]. Kernel [1644, 275, 34, 1231, 576, 1757, 2383, 838, 2270, 237, 144, 1218, 588, 2089, 779, 2019, 49, 1870, 165, 999, 675, 745, 308, 330, 1400, 1128, 194, 2587, 1102, 518, 1994, 1229, 2312, 2011, 980, 1828, 2109, 692, 1456].

Kernel-based [34, 692]. kernel-type [49, 1870]. kernels [2436, 687, 1462]. kinetic [2553]. kinetics [2496]. Kolmogorov [2516, 2070]. Kong [432, 1016]. Koopman [1760]. Koziol [1118]. kriging [1377, 1378]. kriging-based [1378]. KrigInv [1377]. Krylov [2071]. Kullback [166]. Kuramoto [2350]. kurtosis [544, 1940, 511].

L [609]. label [1363, 2321]. labor [1049]. labour [2154, 2410]. lack [1723, 347, 1004, 304]. lack-of-fit [1723]. LAD [773, 2040, 605]. LAD-LASSO [773]. lag [403, 260]. lagged [1419]. Lai [609]. lambda [4, 133]. land [980]. land-cover [980]. Laplace [651, 897, 95, 2369, 872, 289, 2498, 1799, 732, 1619, 595, 762, 472, 1312, 1210]. Laplacian [2552, 362]. Large [809, 1095, 735, 2407, 2598, 1867, 2140, 144, 730, 67, 1860, 248, 829, 2583, 2595, 2564, 1079, 1632, 1793, 711, 1871, 311, 1274, 2071, 1732, 2484, 532, 883, 2555, 1556, 451, 1052]. large- [1732]. large-dimensional [1556]. large-scale [311, 2071, 532, 2555]. largest [752]. Lasso [2568, 1999, 2260, 286, 2570, 1839, 451, 1548, 1546, 1531, 1692, 2173, 2305, 1868, 2051, 1285, 1004, 1686, 1358, 1279, 1975, 773, 2610, 1774, 824, 1587, 2374]. lasso-penalized [2305]. LASSO-type [2610]. late [1480, 44]. late-2000s [1480]. Latent [2464, 443, 1797, 1021, 634, 1373, 1670, 2572, 571, 1980, 1237, 1698, 939, 2145, 319, 2319, 1242, 2309, 1331, 191, 470, 383, 1763, 1591, 2501, 793, 1535]. latent-variable [793]. later [128]. law [2424]. layer [2007, 2257]. Learned [927]. Learned-loss [927]. Learning [1316, 73, 692, 1448, 2491, 1398, 2390, 1374, 1997, 2457, 34, 1705, 1070, 1542, 590, 2381, 1409]. learning-based [1398]. Least [1450, 47, 1215, 800, 1266, 1818, 1350, 2319, 307, 1299, 308, 504, 910, 1675, 328, 667, 432, 758, 292, 1881]. Least-squares [1215, 308, 910]. Left [933, 1914, 2115, 2059, 1771, 1444, 2611]. left-censored [1771]. left-censoring [1444]. left-truncated [2611]. legacy [1964]. legislative [1228]. Leibler [166]. length [1033, 2495]. length-biased [2495]. lesson [1016]. level [886, 1565, 1049, 187, 2298, 849, 1260, 2512, 2563, 2558, 749, 1099, 2152, 985, 1553, 1040, 556]. levels [2367, 28]. leverage [1085, 1498, 908, 1950, 909, 1515, 387]. Levinson [615]. liability [1810]. life [1109, 440, 816, 2541]. lifetime [426, 372, 674, 1144, 413, 603]. lifetimes [1197]. like [2137, 1029, 142]. Likelihood [639, 1179, 540, 2586, 2564, 2440, 912, 384, 2357, 952, 1013, 1742, 1658, 1967, 1479, 890, 1481, 2545, 466, 933, 1197, 2283, 891, 1674, 2120, 126, 1422, 1487, 1643, 2190, 2418, 352, 783, 554, 2025, 2342, 831, 1801, 566, 1882, 1829, 247, 1581, 2043, 1993, 965, 1117, 2146, 2517, 1061, 1429, 403, 951, 476, 1504, 1507, 528, 337, 1073, 2506, 125, 1223, 2567, 2540, 2147, 1679, 1777, 1365, 1578, 1445, 1668, 1830, 1112, 1370, 553, 2008, 1579, 630, 997, 2277, 254, 1529, 437, 2077, 1310, 2162, 645, 2150, 299, 2483, 2116, 462, 1807, 1516, 1261, 104, 2320, 1677, 1443, 53, 2582]. likelihood [127, 454, 1028, 1278, 1568, 2247, 2521, 407, 1280, 2417, 1766, 1224, 2317, 833, 1709, 2234, 2495, 1845, 353, 422, 2427, 1183, 1767, 1918, 157, 498, 1549]. Likelihood-based [384, 1197, 104]. Likelihood-free [540, 2440, 912, 1829]. likelihoods [2366, 2376, 1331]. limit [1469, 1521, 1461]. Limited [1001, 1591]. limits [1252, 1687, 953, 2035]. Lindley [1143]. line [1617]. lineage [1248]. Linear [1151, 749, 1354, 2026, 1347, 572, 88, 267, 2533, 810, 466, 1702, 54, 1171, 2090, 542, 270, 2505, 45, 1468, 589, 326, 2007, 1440, 2278, 183, 2391, 452, 1643, 1996, 2392, 2255, 2104, 866, 646, 1618, 551, 2154, 1379, 131, 582, 2023, 2476, 2460, 1385, 1313, 279, 1685, 2236, 204, 1169, 1457, 2576, 329, 1617, 1659, 2127, 1735, 307, 1805, 954, 1069, 2040, 2531, 392, 441, 1860, 1315, 143, 2238, 1061, 1429, 2371, 533, 1125, 901, 1288, 742, 1604, 1853, 1855, 2264, 1107, 1611, 1018, 1722, 2147, 1721, 2097, 2525, 2591, 2305, 2593, 1635, 2458, 2003, 2453, 1168, 1006, 290, 1888]. linear [1355, 291, 2605, 1234, 1036, 38, 1799, 1579, 293, 673, 296, 1221, 416, 1638, 1235, 2162, 1876, 645, 1423, 688, 1714, 1612, 672, 2279, 140, 1917, 1819, 369, 56, 733, 1120, 1208, 860, 462, 2276, 629, 2523, 2049, 2165, 2333, 1261, 104, 563, 961, 2518, 952, 1013, 765, 1443, 261, 1072, 127, 665, 1186, 1187, 1607, 1891, 2213, 1537, 1773, 1368, 668, 1715, 2415, 1739, 2356, 122, 580, 664, 1068, 2182, 2429, 955, 1613, 1877, 442]. linearity [1477]. linearization [1260]. Linearized [1587]. linearly [1692, 260]. link [88, 1996, 2430, 1336, 1425]. link-preserving [1996]. linkage [839]. linkages [234, 16]. Linking [1964]. links [1323]. lists [2256]. Liu [86, 1927]. LMS [818]. load [874]. loading [1744, 2223]. Local [2141, 441, 310, 1714, 475, 1854, 45, 1049, 51, 1237, 870, 1225, 1666, 1843, 2580, 2147, 2164, 2051, 525, 997, 2386, 1423, 1392, 400, 2523, 1121, 1542, 159, 2078, 1968]. Local-moment [310]. Localized [886, 1848]. Locally [662, 1116, 2465, 2607]. Location [2584, 1753, 830, 54, 674, 534, 821, 2610, 343, 1132, 1622, 2508, 198, 2161, 356, 2131, 2181, 959, 2116, 1883, 2402, 130, 1241]. Location-adjusted [2584]. location-scale [830, 1883]. location-shifted [2116]. location/dispersion [356]. loci [55]. log [2106, 1473, 646, 674, 100, 2019, 1993, 2158, 347, 1979, 673, 2609, 1046, 1963, 1087]. log-concave [1473, 1993, 2158, 2609]. log-density [2019]. log-exponentiated [100]. log-gamma [2106]. log-GARCH-X [1963]. log-linear [646, 673]. log-location [674]. log-multiplicative [347]. log-normal [1979, 1046]. logarithm [2234]. Logic [1640]. Logistic [1178, 1160, 1381, 560, 2520, 1726, 995, 1996, 2457, 392, 1114, 1318, 966, 330, 1566, 2160, 1861, 1191, 1195, 1393, 720, 832, 1840, 2504]. logit [782, 1779, 98, 1645]. LogitBoost [2180]. loglinear [946, 347]. loglog [98]. Lognormal [1197, 1565, 1257]. logrank [523, 2547]. LOL [1356]. Lomax [429]. Long [1764, 1486, 911, 1480, 1484, 233, 769, 1938, 1497, 1970, 1499, 533, 202, 1849, 249, 637, 1986, 1337, 1956, 1515]. long-memory [1499, 637]. long-range [769, 1986]. long-run [1480, 1956]. Long-term [1764, 202, 1337]. longer [2233]. Longitudinal [26, 1717, 989, 2219, 1687, 1351, 1696, 606, 2431, 1304, 2462, 2108, 2013, 820, 1718, 932, 1586, 2133, 1091, 82, 2184, 1073, 1849, 2352, 2550, 1107, 1223, 2297, 2301, 2451, 2540, 178, 2329, 1763, 443, 667, 1635, 2183, 401, 1837, 2230, 1530, 1339, 1632, 2367, 128, 2162, 1876, 657, 1470, 2022, 1401, 947, 341, 1244, 961, 1005, 1533, 2326, 718, 303, 1561, 2521, 2501, 1589, 722, 1600, 2381, 25, 1045, 2485, 1840, 2581]. Loss [124, 228, 2454, 927, 2040, 1172, 2374]. loss-based [2454]. Louzada [1447]. low [309, 2579, 2266]. low-dimensional [2579]. low-rank [309]. Lower [1461, 636]. lower-bound [636]. LTS [1922, 818].

M [1447, 1555]. M-regression [1555]. MA [880]. machine [433, 1203, 1299, 34, 362, 667, 1430, 1827]. machines [479, 687, 2140, 1108]. macroeconomic [1489, 1945]. Mahalanobis [1920]. Mahdi [1075]. main [2298]. maintenance [967, 1539, 1136]. majorization [302]. make [1776]. Making [1545, 1190]. MALA [2141]. MALA-within-Gibbs [2141]. mammograms [1031]. management [894]. Managing [1939]. Manhattan [825]. manifest [319]. manifold [1297]. manifolds [2100]. Manly [2335]. MANOVA [11]. Mantel [835]. many [109]. map [1398, 109]. Mapping [3, 344, 80, 2251, 734]. MAR [82]. Marginal [669, 1404, 2320, 890, 937, 891, 54, 2292, 1175, 601, 831, 2553, 281, 1882, 1439, 1331, 1609, 57, 1704, 1188, 1529, 1871, 763, 644]. Marginalized [2550, 772, 26, 385, 1849]. marginally [2026]. marginals [1532]. margins [1147]. mark [464, 1812]. mark-specific [1812]. marked [507]. marker [2281, 1229]. markers [836, 1640]. Market [234, 1478, 2154, 896, 1219, 432, 232, 887, 2410]. Marketing [1971, 2400, 1191]. markets [871, 1049]. Markov [2039, 1481, 1670, 1933, 1095, 1398, 2437, 63, 2295, 2145, 783, 1134, 1404, 1940, 900, 1949, 375, 2583, 1227, 1837, 1010, 2117, 1957, 828, 911, 736, 1040, 1345, 618, 2555, 2597, 1601, 2501, 2356, 2612]. Markov-modulation [2612]. Markov-switching [1481, 1940]. Markovian [1272, 808]. marks [1864]. Marshall [1631, 1161, 992, 780, 2103, 797]. martingale [190, 2573]. mass [2073, 109]. massive [625]. massively [1411]. Master [2469]. matched [995, 704, 708, 568, 723, 763, 1630, 14]. matched-pair [723, 763, 1630]. matches [2410]. Matching [2410, 2052, 1238, 483]. matrices [857, 1541, 1816, 1330, 2047, 2564, 276, 642, 1573, 2439, 1817, 1556, 1052, 1456, 2034]. Matrix [307, 1950, 1778, 305, 2390, 2407, 1605, 33, 2013, 480, 2568, 2307, 999, 1860, 1574, 983, 223, 742, 1604, 2264, 2161, 2441, 1148, 2569, 1455, 1599, 614, 1771, 379, 2484, 1667, 2170, 1568, 2448, 1074, 75, 2028, 2599, 2234]. matrix-logarithm [2234]. matrix-valued [2028]. MAVE [1131]. Max [198]. Max-type [198]. maximin [2221]. Maximisation [2522]. Maximization [2029, 1874, 615, 1801, 920, 1761]. maximize [2222, 1715]. maximizing [1546, 2516]. Maximum [1479, 1481, 639, 1179, 783, 1801, 1993, 2567, 1370, 553, 2008, 498, 466, 1674, 2120, 1487, 2190, 2342, 247, 1061, 1429, 1504, 528, 547, 978, 2432, 1445, 1668, 1830, 2277, 254, 2077, 2116, 980, 2320, 1677, 502, 127, 407, 1766, 1709, 422]. may [1448]. MCEM [634]. MCMC [2292, 2420, 1503, 344, 2248, 444, 1719, 1907, 341, 1348, 1210, 1614, 184, 1346]. MCMC-based [341]. Mean [1213, 2586, 1430, 1243, 368, 1109, 2385, 43, 1361, 2601, 802, 464, 648, 1646, 2443, 2566, 749, 401, 853, 617, 1876, 619, 119, 1732, 629, 2589, 1437, 459, 1568, 449, 1634, 2338, 207, 90, 2541, 214, 1894, 1183, 1767, 1918, 157, 1775]. mean-covariance [401]. mean-field [368]. mean-of-order [1646]. mean-shift [629]. mean-variance [2566, 2338]. means [690, 752, 203, 1039, 487, 984, 393, 711, 1257, 177, 1283, 449, 1582]. measure [365, 208, 1225, 379, 2027, 919, 573]. measured [3, 80]. measurement [650, 706, 283, 1855, 2329, 2209, 1243, 1876, 786, 1879, 1861, 1195, 1013, 1964, 1595, 10, 1068, 1048, 2182, 2395]. measurements [2199, 1676, 704, 715]. measures [415, 1088, 1655, 582, 481, 1500, 842, 182, 1690, 597, 789, 1908, 128, 598, 841, 556, 1519]. Measuring [272, 2426]. Mechanism [458, 2104, 2203, 2406, 797]. Mechanism-based [458]. mechanisms [411, 603]. Median [2226, 15, 1078, 2046, 454]. Median-based [2226]. medians [735, 1181]. mediation [1811]. mediator [1811]. mediator-outcome [1811]. Medical [2196, 433, 35, 1679]. medicine [723]. medium [1325]. meets [1497]. Meier [1438]. melanoma [753]. members [425]. membership [1061]. memberships [692]. memory [1484, 233, 1486, 1938, 1497, 1499, 249, 637, 911, 1515]. merits [1469]. Meta [702, 2204, 1180, 1655, 1152, 719, 2456, 422]. meta-analyses [1655]. Meta-analysis [702, 1180, 719]. Meta-analytic-predictive [2204]. meta-regression [702, 2456]. method [773, 1372, 2343, 2557, 2418, 1491, 1544, 1937, 748, 655, 2177, 524, 1200, 1421, 1829, 2127, 1832, 179, 1587, 1293, 2540, 391, 2308, 935, 1635, 386, 290, 1355, 1042, 1037, 963, 1588, 1994, 108, 545, 1637, 339, 1959, 1636, 2393, 1924, 104, 2256, 1191, 2614, 2170, 1765, 2597, 1992, 776, 2551, 173, 451, 793, 1067, 2504]. methodologies [757]. methodology [1386]. methods [890, 2060, 327, 2197, 2565, 1095, 1925, 272, 1322, 2598, 1835, 1590, 492, 507, 526, 551, 1689, 1610, 596, 86, 1410, 679, 1781, 623, 1149, 1462, 282, 363, 1952, 2426, 2092, 2580, 287, 710, 1822, 296, 981, 1378, 1320, 881, 1465, 594, 74, 2071, 1626, 341, 106, 561, 717, 1755, 1199, 2088, 2475, 1725, 2028, 147, 2414, 605, 32]. methylation [757]. Metropolis [1563, 1657, 1007, 259, 12]. MGARCH [1511]. microarray [377, 768, 395, 732, 659]. microsimulation [985]. Microsoft [114]. microstructure [1520]. migration [895]. minimal [1621, 2011]. minimax [2310, 2221, 1827]. minimising [963]. minimization [1639, 1527, 2532]. Minimum [973, 608, 1598, 1820, 1146, 2268, 85, 2607, 1044, 799, 591, 1172, 1740]. Mining [1815, 68, 1351, 1821, 1000, 34]. mis [82, 811]. mis-specification [82, 811]. misclassification [1862]. misclassified [823, 2154]. Mises [126, 59, 5, 65, 785]. mismeasured [2197, 1795]. Missing [2525, 425, 1864, 54, 1171, 2594, 380, 1996, 2418, 622, 2104, 1983, 2394, 1402, 2023, 2476, 112, 2490, 1265, 2443, 2576, 1606, 2250, 82, 282, 1178, 2400, 1832, 383, 708, 2079, 790, 1073, 1855, 2341, 1223, 1472, 2406, 2540, 2477, 2329, 1795, 2178, 1579, 1876, 1776, 301, 2276, 1401, 2215, 2320, 1642, 2518, 1755, 1665, 1224, 25, 1733, 2078, 1183, 2216]. missingness [1319, 341, 722]. misspecification [1179, 374, 2426, 1779, 555, 1600, 930]. misspecified [88, 608, 1014]. Mixed [1104, 1868, 1323, 1381, 2545, 466, 1702, 2102, 1468, 571, 183, 205, 1304, 782, 2286, 783, 646, 551, 739, 1150, 2460, 1260, 2563, 204, 754, 1061, 1429, 944, 1157, 23, 742, 1096, 1604, 1458, 1855, 1318, 1955, 2097, 2571, 2525, 854, 1795, 667, 2554, 2388, 1837, 1006, 1599, 2026, 296, 334, 416, 439, 1638, 409, 2150, 1612, 1513, 1772, 56, 733, 629, 1261, 1716, 1191, 961, 1736, 952, 1013, 2170, 118, 127, 2206, 2475, 776, 25, 120, 832, 668, 1709, 323, 1739, 122, 664, 1840, 2018]. mixed-effect [776]. mixed-effects [205, 1157, 23, 1458, 2097, 2571, 1795, 1006, 334, 439, 664]. mixed-type [2388, 2150]. mixing [1490, 2342, 1213, 1318, 217, 215]. Mixture [1796, 1790, 1035, 1786, 1142, 1306, 2150, 1891, 1724, 1323, 2502, 91, 1813, 804, 269, 931, 2174, 1302, 2464, 624, 571, 2016, 2129, 1436, 195, 1792, 1531, 2497, 44, 2342, 831, 2332, 1985, 1993, 1305, 1505, 202, 844, 2001, 1853, 1333, 1314, 1316, 2305, 555, 2076, 1460, 1319, 1795, 1979, 500, 1830, 2587, 71, 731, 1799, 920, 525, 73, 300, 1342, 1782, 1787, 185, 815, 2589, 217, 1312, 1642, 2287, 1317, 2176, 2257, 2500, 1280, 2417, 2501, 776, 2042, 173, 2381, 1311, 2351, 793, 353, 1067, 643, 2335]. Mixture-based [1790]. mixture-of-regressions [2502]. Mixtures [1321, 817, 918, 2498, 1547, 1791, 1338, 134, 1665, 2461, 1798, 215, 266, 358, 83, 269, 427, 364, 1800, 973, 1313, 1921, 491, 2304, 2158, 463, 983, 768, 553, 734, 1797, 1896, 1793, 2373, 1320, 1310, 1030, 1776, 2116, 387, 2613, 1753, 2035]. MLE [463, 2110, 185]. MM [1682, 1114, 288, 2614, 925]. Modal [2562, 2521]. modalities [1237]. Mode [2420, 571, 1139, 1820, 1034, 2614]. mode-sharing [2614]. Model [1349, 2339, 358, 1447, 2437, 1307, 380, 946, 1218, 581, 2481, 2576, 2066, 124, 1315, 1309, 2355, 685, 2214, 2079, 2580, 536, 834, 2230, 1789, 291, 2096, 257, 1021, 301, 1357, 579, 1626, 940, 1707, 843, 1303, 2206, 2613, 2213, 1562, 1284, 1381, 1477, 1118, 885, 1266, 755, 2020, 1761, 91, 2411, 83, 338, 1481, 1482, 345, 964, 639, 1179, 2227, 1864, 1471, 436, 426, 968, 1674, 1565, 2572, 656, 2197, 54, 1038, 542, 1796, 1486, 1440, 548, 650, 696, 960, 838, 865, 895, 1726, 2391, 2186, 2327, 2129, 107, 825, 1124, 995, 1009, 1326, 1703, 2031, 2190, 474]. model [1546, 939, 1969, 2009, 782, 2145, 1328, 153, 1085, 2289, 646, 371, 1655, 2452, 551, 1682, 769, 739, 44, 2476, 1150, 775, 1790, 1116, 1152, 820, 1213, 1385, 66, 2445, 1706, 2101, 280, 800, 1329, 2369, 2024, 204, 1865, 245, 281, 329, 199, 2271, 100, 1272, 914, 2187, 2229, 2368, 2493, 191, 1810, 943, 1132, 2465, 52, 374, 965, 2040, 2531, 392, 2529, 772, 908, 1017, 1305, 2238, 706, 383, 2517, 741, 587, 729, 533, 976, 1609, 1666, 1282, 905, 202, 1288, 2189, 385, 1458, 1741, 1853, 2432, 412, 1144, 2426, 2522, 957, 1431, 1998, 2092, 2139, 966, 1314, 455, 1018, 1137, 1020]. model [1083, 1202, 1763, 811, 1227, 667, 2183, 2458, 500, 401, 1290, 2574, 287, 1106, 1830, 2474, 1354, 1079, 2212, 71, 1693, 1797, 909, 1037, 251, 1831, 2072, 1014, 1338, 1339, 1510, 1405, 253, 673, 1, 527, 295, 296, 1449, 437, 1369, 985, 1341, 298, 2223, 598, 672, 1425, 1342, 1105, 1780, 206, 2280, 1514, 1782, 1916, 185, 1207, 1208, 2185, 2276, 2484, 2523, 577, 1343, 1517, 217, 1312, 1924, 1962, 1812, 1906, 1191, 610, 1737, 519, 1633, 774, 1156, 1898, 1978, 808, 1965, 1317, 720, 373, 517, 1278, 1451, 1660, 1856, 2378, 1601, 2461, 189, 407, 159, 1662, 1210, 2510, 1591]. model [2017, 37, 142, 173, 950, 832, 1204, 1733, 2351, 833, 1709, 2241, 664, 930, 822, 1067, 1569, 1697, 1653, 2269, 2427, 2416, 2358, 643, 1576, 1724, 2266, 2419, 2404, 1919, 2033, 2018, 442, 592, 1943]. Model-averaged [581, 843]. Model-based [358, 1307, 946, 2481, 2066, 1315, 1309, 2355, 834, 2230, 1789, 579, 2206, 2613, 2411, 338, 2129, 2452, 66, 71, 206, 1317, 2461]. model-fit [964]. Model-robust [940]. modeled [618]. Modeling [229, 1482, 544, 1654, 868, 513, 742, 2062, 1560, 1619, 1419, 424, 1730, 113, 303, 1016, 1045, 2480, 644, 141, 2094, 269, 1252, 2259, 624, 427, 1308, 440, 1938, 1242, 2069, 1062, 1985, 587, 434, 1142, 845, 2192, 1144, 726, 178, 443, 2183, 2243, 1566, 494, 617, 2138, 2103, 1559, 1665, 1348, 1537, 408, 2335, 592]. modelled [447]. Modelling [1717, 631, 850, 894, 913, 1489, 1800, 2549, 1432, 858, 388, 1345, 2174, 36, 1422, 874, 1948, 2306, 855, 2209, 2570, 450, 262, 117]. Models [986, 1786, 1634, 572, 266, 1323, 599, 88, 2533, 1151, 1967, 740, 2331, 634, 1479, 949, 1012, 1813, 368, 788, 338, 810, 1483, 2545, 466, 804, 1484, 1197, 1702, 1764, 937, 1373, 1670, 2354, 891, 1932, 1995, 2141, 1325, 322, 1850, 931, 1687, 2413, 1171, 2090, 1933, 2594, 1935, 1934, 879, 270, 2505, 1302, 1796, 2464, 753, 2102, 235, 1468, 1806, 589, 63, 571, 1487, 364, 89, 1488, 2188, 1049, 1980, 1412, 2016, 205, 452, 467, 669, 701, 1226, 1490, 1915, 1996, 2431, 2585, 1698, 317, 687, 1792, 236, 381, 1304, 1399, 352, 1723, 2286]. models [492, 507, 2255, 2104, 2462, 2366, 2430, 64, 148, 144, 319, 866, 2376, 783, 867, 1384, 2217, 2211, 2226, 2452, 551, 1610, 777, 240, 241, 1424, 730, 2497, 1258, 277, 796, 131, 582, 849, 1175, 601, 2023, 1912, 2319, 2342, 674, 851, 702, 831, 813, 2460, 511, 1942, 1260, 1902, 430, 2224, 1496, 2553, 2236, 244, 2610, 1580, 1169, 899, 1122, 281, 1498, 1829, 402, 1947, 1946, 246, 1439, 2036, 1948, 2346, 2127, 247, 359, 307, 1949, 2457, 1805, 2158, 1117, 1069, 2350, 1499, 441, 1315, 2057, 143, 1153, 1832, 608, 1332, 1895, 283, 1598, 1661, 754, 1061, 1429, 403]. models [1503, 951, 2371, 829, 1125, 1989, 910, 1504, 902, 872, 2110, 347, 2041, 1157, 375, 1603, 1953, 528, 2583, 23, 160, 26, 337, 844, 742, 817, 1096, 1604, 1690, 1691, 1849, 2001, 1855, 2264, 2341, 2432, 2409, 2398, 2352, 2550, 1585, 1318, 2345, 411, 348, 824, 948, 780, 906, 1107, 1587, 2053, 1954, 2297, 2451, 2514, 2567, 1135, 1611, 1333, 1316, 1779, 1722, 1971, 2590, 1721, 2097, 2571, 2573, 2525, 2591, 2305, 876, 854, 431, 620, 2076, 356, 1460, 1319, 1795, 2438, 2593, 2384, 877, 1281, 1635, 1445, 2554, 642, 410, 2003, 2453, 1560, 1050, 1837, 521, 2054, 2181, 710, 1508, 1006, 602, 2587, 332, 1334, 71, 731]. models [1888, 291, 1509, 1413, 1234, 57, 2117, 586, 734, 1170, 880, 324, 1036, 38, 1306, 1599, 2026, 293, 1338, 1188, 920, 1632, 1907, 663, 747, 2164, 296, 2496, 175, 764, 1129, 1793, 355, 545, 2121, 828, 334, 1221, 929, 912, 1973, 416, 439, 1638, 819, 1235, 1030, 409, 2162, 1876, 2386, 688, 881, 98, 2150, 73, 458, 1089, 1040, 657, 2103, 2218, 1612, 2268, 1513, 2279, 140, 300, 762, 472, 1274, 1772, 1819, 2022, 1960, 56, 2071, 733, 1787, 815, 2328, 860, 462, 629, 2523, 39, 2049, 1516, 2604, 2205, 2165, 947, 982, 2082, 2325, 2333, 1244, 1345]. models [1261, 647, 1879, 1861, 104, 1963, 2320, 2300, 1841, 406, 1686, 961, 1005, 1533, 2518, 537, 2560, 20, 1393, 952, 1013, 883, 765, 626, 1883, 1098, 1645, 2170, 1130, 1846, 99, 750, 1443, 1519, 1084, 692, 12, 1595, 1072, 1317, 1731, 118, 720, 127, 387, 1187, 1607, 1891, 2176, 2257, 2555, 2521, 159, 2500, 161, 304, 2402, 994, 1280, 2417, 2213, 1591, 2012, 1881, 2191, 2501, 776, 1589, 1725, 2578, 564, 1292, 1854, 2330, 2381, 1224, 207, 120, 1368, 1015, 793, 346, 668, 2415, 2495, 2456, 313, 323, 1739, 2362, 122, 580, 664, 930, 1068, 1535, 2182, 2429, 2504, 1613, 1745, 1877, 2603, 2395, 605, 50, 2033, 2134]. models [495, 1868]. Moderately [1774]. modern [1500]. modification [1146]. modifications [113]. Modified [1484, 703, 1268, 2506, 572, 1834, 2566, 1563, 1843, 990, 2417]. modifiers [1576]. modulation [2612]. molecular [2398, 1460]. Moment [1972, 1195, 2081, 876, 310, 2185]. moments [1372, 560, 1937, 953, 545, 169]. Monitoring [1148, 1848, 1057, 578, 534, 821, 222, 2489, 563, 1568, 162, 214]. Monotone [1548, 1133, 548, 1450, 1794, 1399, 437, 341]. monotonic [532]. monotonicity [2408, 137]. Monotony [1659]. Monte [1890, 2583, 890, 2239, 2437, 850, 33, 996, 1404, 1330, 1060, 165, 69, 70, 1857, 2110, 2522, 1137, 494, 1803, 1413, 1797, 2577, 370, 736, 257, 2074, 2170, 168, 2597, 1662]. monthly [855]. Moran [1859]. morbidity [1015]. mortality [2188, 1654, 44, 2306, 2209]. Most [1090, 291]. mostly [268]. motivated [2562]. motor [1404]. movement [2072]. Moving [2544, 534, 821, 954, 1507, 2183, 1619, 2607]. MT [1477]. MT-STAR [1477]. multgee [1530]. Multi [1659, 1995, 913, 2538, 775, 17, 2123, 2438, 2499, 1871, 1917, 1863, 102, 945, 822]. multi-class [1917]. multi-dimensional [1871, 102]. multi-index [775]. multi-output [913]. multi-partitions [2499]. multi-rater [17]. multi-regime [945]. multi-resolution [1863]. multi-response [2438]. multi-row [2123]. multi-sample [2538]. multi-species [822]. multi-type [1995]. Multi-way [1659]. multiband [980]. multicategorical [2024]. multiclass [1644, 1826, 1566]. multicollinearity [1351, 1130, 2526]. multicycle [1612]. Multidimensional [2022, 1181, 1373, 2450, 1994, 1158, 738, 1518, 382]. multidimensionally [2444]. multifactor [1071]. MultiLCIRT [1373]. Multilevel [1726, 2364, 468, 772, 2301, 2540, 1306, 774]. multilinear [359]. multilogit [1644]. Multimix [2038]. multimodal [842]. multimodal-valued [842]. multimodality [969]. multinomial [608, 2091, 2233, 1645, 1358, 408]. multinomials [398]. Multiobjective [1841]. Multiple [43, 1494, 2061, 768, 1232, 1130, 173, 1694, 2383, 1848, 964, 2174, 1252, 1687, 2436, 838, 1237, 107, 325, 801, 2031, 1988, 748, 2425, 1176, 1029, 2201, 1750, 743, 2040, 503, 1061, 707, 944, 2511, 435, 1679, 1391, 2114, 1986, 963, 2367, 686, 1416, 2337, 2280, 1046, 1357, 2328, 258, 2256, 428, 486, 497, 719, 1965, 638, 113, 1856, 2244, 2555, 216, 449, 130, 1241, 1641, 1015, 1715, 2548, 77, 1905, 1919]. multiple-output [686]. multiple-try [1416]. multiplicative [236, 1781, 49, 1870, 347, 2003, 1831, 2185, 919, 2429]. multipliers [1587, 2551]. Multiply [1855, 2418, 218, 2151, 1776, 1642]. multiply-imputed [2151]. multiresolution [565, 662]. multiresponse [1429]. multisample [1884]. multiscale [518, 1031, 618]. multistratum [2299, 2590]. Multivariate [1119, 1995, 1850, 624, 2323, 1491, 670, 1242, 1895, 2360, 1532, 2161, 1320, 2560, 2163, 141, 969, 228, 358, 2435, 2219, 2283, 1407, 546, 2505, 235, 18, 1488, 360, 2016, 669, 1703, 43, 5, 59, 1590, 1544, 274, 2473, 2108, 540, 465, 2342, 600, 588, 1940, 480, 1781, 1313, 1944, 2089, 1439, 1138, 1139, 675, 1622, 2066, 463, 908, 1309, 1154, 983, 1820, 1147, 172, 96, 844, 749, 2301, 2451, 1137, 991, 1664, 1008, 1669, 1276, 2131, 1624, 2230, 1413, 1234, 2117, 2498, 1788, 175, 929, 1209, 448, 416, 2608, 156, 2609, 257, 2399, 1342, 1514, 369, 2357, 1963, 1533, 486]. multivariate [1195, 1250, 1003, 2372, 118, 127, 1753, 2501, 1733, 1739, 214, 822, 2037, 2404, 1775, 2018, 1456]. multiway [359]. mutations [2205]. mutual [1363, 1403].

Nadarajah [1075]. Natural [1888, 101, 2397, 1015]. NBUE [600]. NDARMA [1233]. near [838, 1721, 1400, 2010]. near-exact [2010]. Nearest [2148, 1755]. nearly [467, 1384]. needed [2492]. Negative [1345, 1781, 430, 1688, 520, 2003, 409, 2448, 2004]. negatively [1292]. neglected [98]. neighbor [2148, 1755]. neighborhood [1002, 129]. Neighbourhood [1960]. Nelson [914]. Nested [1720, 367, 2013, 321, 1429, 1096, 762, 1408, 1208, 42]. Network [2278, 2407, 2562, 2132, 2282, 121, 1814, 1021, 1199, 1773]. network-constrained [1773]. Networks [1997, 2471, 2561, 2491, 2390, 2180, 2588, 1079, 1632, 1230, 2, 611, 2600, 590]. neuroimaging [1990]. news [232]. Newton [970]. Neyman [1293]. NGC [974]. NMR [102]. NN [132]. no [823, 82]. Noise [1822, 1037, 102, 1383, 1363, 2133, 2001, 919, 1520]. Noisy [1378, 631, 2437, 2179, 1636, 2484]. nominal [26]. nomination [1427]. Non [2492, 2202, 1735, 2273, 853, 261, 219, 942, 176, 1932, 2594, 2391, 2412, 607, 2063, 240, 2025, 2254, 1421, 1781, 641, 2101, 2236, 244, 68, 247, 2293, 1860, 1429, 2203, 744, 2406, 1526, 1319, 2003, 768, 2095, 2605, 568, 1129, 409, 1536, 881, 926, 2022, 1863, 2523, 341, 1978, 2448, 1210, 2004, 2402, 2370, 723, 77]. Non-area-specific [2273]. non-central [1526, 768]. non-Gaussian [2236, 244, 2095, 409, 881]. non-homogeneous [2101]. non-ignorable [2594, 2203, 2406, 1319, 2022]. Non-inferiority [2492, 2202, 942, 2406, 568, 2370, 723]. Non-linear [261, 2391, 2236, 2605]. non-local [2523]. non-negative [1781, 2003, 2448, 2004]. non-nested [1429]. non-normal [176, 2202, 1860, 1210, 77]. non-normality [240]. non-orthogonal [1863]. non-overlap [219]. Non-parametric [1735, 853, 2412, 607, 2063, 2254, 1421, 68, 2293, 261, 1978]. non-random [1129, 341, 2402]. non-randomized [1536]. non-scalar [1932]. non-smooth [2025]. non-stationary [641, 247, 744, 926]. noncausal [1510]. nonconvex [860, 2374, 640, 2551]. nonignorable [2104, 2576, 1017, 1832, 1579, 1401, 1642, 722, 25, 2078]. nonignorably [383, 1183]. Nonlinear [327, 1825, 1560, 339, 885, 1266, 2545, 879, 364, 2585, 47, 414, 524, 1260, 2549, 1580, 1060, 1857, 1157, 812, 1721, 1795, 1006, 1599, 334, 1274, 99, 750, 1868]. nonlinearities [627, 911]. nonlocal [919]. nonmonotone [1401]. Nonnegative [1462, 1279, 1685, 2244, 1720]. Nonnegative-lasso [1279]. nonnormal [739, 2250, 1797]. Nonparametric [1180, 628, 31, 2016, 1794, 1835, 1546, 1781, 1136, 1066, 1673, 2321, 49, 1870, 744, 1318, 1426, 1205, 2076, 1400, 194, 1808, 1577, 663, 2121, 959, 938, 1408, 922, 1406, 1667, 2579, 407, 2479, 1360, 1259, 83, 19, 1133, 1935, 746, 2557, 51, 973, 2129, 1327, 1988, 1558, 622, 1328, 64, 2527, 596, 777, 2342, 2199, 2309, 794, 1746, 534, 1557, 1865, 1882, 186, 2346, 1225, 17, 308, 754, 476, 1820, 182, 351, 1955, 435, 2308, 1578, 2026, 310, 1243, 2536, 206, 2068, 1207, 1809, 110, 719, 573, 837, 2075, 2475, 832, 833, 1569, 1346]. nonparametric-test-based [573]. nonrandomized [2024]. nonresponse [1017]. nonspherical [432]. nonstationarity [852]. Nonstationary [494, 1467, 2002, 2062, 883, 2037]. Norbert [700]. norm [684, 1100, 1605, 1203, 2361, 2260, 1522]. Normal [91, 2073, 266, 1323, 268, 1251, 345, 1321, 1471, 269, 176, 404, 624, 1440, 364, 960, 2598, 1618, 851, 480, 193, 2202, 470, 1860, 1075, 741, 729, 1039, 983, 1157, 2001, 787, 1137, 984, 555, 1979, 289, 651, 1339, 1788, 711, 94, 448, 192, 1454, 1513, 1046, 185, 815, 2589, 977, 2326, 1210, 1562, 77, 1918, 2035]. normal-based [404]. Normal-Gamma [2073]. normal-Laplace [651, 289]. normal/independent [122]. normality [240, 486, 2263]. normalized [513]. normalizing [1332]. normals [525]. Northern [825]. note [368, 2318, 1447, 64, 2069, 1760, 334, 1004, 610]. Notes [2469, 1602]. notice [611]. novel [1923, 705, 1999, 2531, 2256, 2614, 1591, 2581, 2504]. novelty [2384]. NP [1611]. NP-dimensionality [1611]. nuclear [425, 1605, 2035]. nuFFT [1588]. nugget [949, 756]. nuisance [1084, 1521]. nuisance-free [1521]. null [550, 1570, 525, 2417, 2241]. number [1483, 1764, 273, 2492, 1404, 655, 1421, 2408, 1949, 1047, 1903, 766, 172, 496, 2595, 1678, 731, 1793, 2373, 711, 1310, 1776, 1274, 2372, 1881, 2191]. Numbers [2469, 1376]. Numerical [2598, 1492, 2261, 1350, 121, 1094].

Obituary [700]. Objective [1161, 741, 2139, 1084, 2372, 564, 974, 541, 1588, 1994]. objectives [1391]. objects [510]. oblique [1581, 1875]. observation [2108, 1426, 340, 975]. observational [468, 2203, 1780, 454, 793]. observations [271, 275, 1384, 2217, 2375, 2123, 1223, 1148, 2218, 301, 2215, 2244, 1992, 2216]. observed [58, 1580, 2043, 2477, 28, 2612]. observer [704]. Obtaining [138]. occupancy [2517]. odds [1222, 2492, 835, 692]. off [1525]. offs [1956]. older [1419]. Olkin [780, 1631, 1161, 992, 2103, 797]. OLS [1176, 901]. One [1012, 591, 1340, 449, 130, 1241, 658, 183, 1597, 112, 139, 693, 953, 1862, 790, 1008, 2243, 747, 128, 1807, 1564, 11, 2317, 2488]. one- [1807]. one-dimensional [1597]. one-factor [2243]. One-sample [1340, 2317]. one-shot [658]. One-sided [449, 953, 1008]. One-stage [130]. One-step [1012, 591, 1564]. one-way [693, 1862, 11, 2488]. Online [641, 2558, 2159, 2390, 870, 96]. only [1070, 2455]. open [1879]. operating [1800]. operator [1751, 2479]. operator-regularized [2479]. operators [1678]. Optimal [2219, 1324, 879, 1057, 1383, 2376, 2217, 674, 705, 2375, 2288, 543, 970, 336, 944, 1365, 1391, 2337, 2223, 2279, 1186, 1520, 2448, 2258, 2415, 1381, 840, 830, 1762, 2262, 1384, 2211, 2105, 2310, 240, 1770, 2224, 1136, 578, 1060, 52, 1047, 709, 2253, 2221, 2397, 2218, 1247, 1470, 1392, 2215, 609, 988, 2075, 2176, 1525, 137, 313]. optimality [183, 1388, 1659]. optimally [1298]. Optimisation [1184, 1952, 85].

Optimization [250, 858, 1422, 1415, 775, 2342, 66, 1843, 2496, 1378, 1841, 638, 1522]. Optimized [144, 145]. Optimizing [2298, 1254, 636]. optimum [2169, 1656, 20, 1393]. option [479, 906, 1513]. optional [1204]. options [1027]. Order [2505, 218, 2280, 560, 492, 526, 1260, 1385, 139, 2490, 1646, 1056, 2273, 52, 1499, 403, 910, 1504, 1971, 393, 1405, 1727, 1858, 1084]. order-based [526]. Ordered [28, 1238, 1790, 191, 1240, 1256, 140, 532, 398, 945, 313]. ordering [1913, 2176]. orders [1109]. Ordinal [1640, 2020, 1696, 2430, 2387, 2309, 68, 2355, 1878, 1239, 1115, 951, 1154, 1849, 2506, 178, 443, 2061, 1530, 2027, 1749, 469, 692, 2258, 1562, 1630]. ordinary [2571]. ordination [2066]. orienting [1542]. Ornstein [881, 244, 537]. orthogonal [187, 2298, 1271, 1722, 1863, 459]. Orthogonality [1745]. Orthogonality-projection-based [1745]. orthologs [394]. orthostatic [392]. oscillating [1464]. other [1095, 702, 343]. our [549]. out-of-sample [900, 627]. outcome [825, 2602, 1178, 2506, 1811, 2388]. outcomes [702, 1073, 2173, 1185, 712, 2489, 2510, 1640]. Outlier [292, 2435, 360, 1167, 212, 943, 1555, 290, 1459]. outlier-free [1167]. outliers [2383, 798, 243, 204, 1586, 2230, 2498, 188, 1250]. outlyingness [2473, 661]. output [913, 1495, 686, 929]. outs [82]. over-dispersion [354]. over-reported [123, 384]. overdispersed [2111, 1764, 772, 134, 1105]. overdispersion [88, 989, 1586, 513, 1099, 1543, 950, 120, 1045]. Overfitting [2373, 2411]. overlap [219, 2088].

P [941]. P-EM [941]. package [1479, 1373, 63, 1376, 712, 1378]. PAGE [583]. pair [1673, 453, 1509, 723, 763, 1630]. pair-copula [1673, 1509]. paired [482, 2406, 194, 456, 1438, 659]. paired-data [2406]. pairs [708]. Pairwise [1267, 951, 2433, 1134, 2000, 1223, 2004]. Panel [1501, 1012, 338, 1483, 19, 889, 1726, 866, 2226, 1493, 1947, 1946, 1650, 1991, 2583, 1426, 1954, 355, 1960, 258, 883, 887, 616, 1893, 340, 975, 2414]. panels [1502, 1370]. parabolas [2179]. Parafac [1469]. Parallel [736, 1159, 2248, 1202, 1719, 71, 1411, 2416, 2039]. parallelisation [259]. Parameter [1372, 1483, 152, 1260, 1580, 966, 1022, 2325, 1559, 1003, 169, 2612, 1381, 773, 1931, 770, 2520, 1440, 1289, 1490, 2424, 2289, 1939, 2039, 1062, 2296, 587, 285, 2212, 1042, 1663, 1214, 73, 311, 1275, 1425, 1749, 2333, 1677, 636, 302, 1347, 1662, 1614, 2266]. parameter-free [1931]. parameterization [345, 1541, 405]. parameterized [1316]. parameters [2561, 1179, 968, 2413, 2584, 794, 2324, 1659, 405, 316, 1857, 1738, 488, 2595, 1972, 150, 1405, 2130, 959, 619, 140, 1782, 462, 816, 1924, 1084, 994, 130, 1241, 1254, 1459]. Parametric [1838, 811, 2088, 639, 2259, 2104, 2412, 607, 2063, 371, 1539, 777, 796, 65, 242, 2254, 1421, 68, 1060, 1735, 2293, 363, 1305, 1094, 853, 2054, 57, 2577, 614, 1129, 1262, 1369, 1257, 2300, 261, 1978, 1212, 2241, 2495, 930, 1653]. Parametrically [1809]. parametrization [2260, 1342]. Pareto [1113, 1166, 1391, 1908, 167, 84, 625, 1658]. Pareto-type [1166]. Parsimonious [1541, 2146, 1317, 71]. part [1912]. Partial [2052, 2127, 434, 2255, 2531, 1125, 2253, 1365, 2593, 2453, 2162, 716, 961, 48, 1883, 692, 516, 1145, 1715, 2182, 2429, 442]. partial-linear [2531]. Partially [1805, 348, 1611, 2477, 1208, 1537, 54, 2090, 270, 2505, 1996, 2255, 2154, 1646, 1169, 2127, 1069, 441, 1125, 1288, 2591, 2458, 1256, 1235, 1681, 1876, 2027, 688, 2279, 462, 2276, 532, 2518, 502, 1187, 1607, 580, 1613, 1877]. particle [2235, 2390, 1415, 2282, 744, 2167, 1184, 859]. particle-learning-based [2390]. partition [324]. partitioned [159]. partitioning [2093, 1402, 1706, 743]. partitions [2499]. partly [2567]. parts [2477, 1924]. past [1036]. path [166, 2003]. paths [453]. pathway [2398]. pathway-based [2398]. pathways [1015]. patients [392, 587, 1701]. Pattern [132, 1194, 1320, 1839, 1341, 469, 1642]. pattern-detection [1320]. pattern-mixture [1642]. patterned [157]. Patterns [1716, 2283, 444, 2384, 109]. PCA [275, 676, 1114]. PDF [2295]. peak [2073]. Pearson [2017]. peeling [1820, 1822]. penalised [1888, 2530]. penalization [1912, 2610, 213]. Penalized [1350, 89, 2366, 499, 1620, 2160, 1702, 1351, 2151, 567, 1581, 476, 350, 1843, 2567, 2305, 1445, 1868, 645, 299, 860, 2513, 450, 1677, 1568, 2118, 640, 616, 1537, 2119, 2480]. penalizing [2234]. penalties [1605]. penalty [272, 635, 2519, 1455, 688, 1220, 1880]. percentage [856]. percentile [1617]. percentiles [460]. perfect [1090, 216]. perform [1448]. Performance [1655, 1129, 1932, 404, 607, 2063, 1379, 2254, 1545, 1211, 2157, 2250, 761, 502, 2446, 46]. performing [1206]. period [1602]. periodic [2094, 615, 874, 247]. Permutation [2442, 659, 179, 16]. persistence [801, 1508]. persistencies [1986]. persistent [1947]. person [195]. perspective [1360, 1104, 714]. perturbation [599]. pharmaceutical [2496]. pharmacoepidemiologic [1410]. pharmacogenetic [717]. Pharmacokinetically [2169]. pharmacokinetics [1868]. Phase [805, 2169, 1324, 2379, 87, 712, 1392, 2489, 1098]. phenotype [623, 2146]. phi [2365]. phi-divergence [2365]. phylogenetic [1460, 2574]. physical [2222, 2496]. physician [1701]. physics [2529, 1663]. piece [1883]. piecewise [904, 643]. piling [1754]. Pitman [830]. pixels [799]. PL [1649]. PL-bagging [1649]. plain [2223]. planning [1120]. plans [1656, 674, 778]. plant [1559, 2035]. Plasmode [1410]. platform [395]. plot [306, 17, 1392, 940]. plots [810, 251]. plotting [830]. PLS [1618, 1659]. PLS2 [1659]. plug [924, 980]. plug-in [924, 980]. Point [2384, 923, 891, 2007, 2436, 367, 43, 507, 2366, 2543, 1920, 566, 2454, 1892, 709, 1227, 789, 1, 527, 1751, 1341, 1089, 1247, 1889, 456, 1043, 609, 1016, 1525]. point-of-view [2543]. points [830, 2267, 92, 1166, 1047, 336, 2536]. Poisson [1197, 1101, 931, 2343, 2102, 372, 1412, 2327, 1915, 739, 1902, 1063, 1862, 1598, 1314, 1664, 710, 1793, 297, 1105, 1543, 1906, 2326, 373, 776, 25, 120]. Poisson-exponential [372]. Poisson-hurdle [1101]. Polarization [1295]. policies [1398]. policy [2254]. political [817]. pollution [2529]. polygonal [2008]. Polynomial [114, 1706, 1675, 1666, 1678, 1284]. polynomials [2168, 663, 450, 1397, 837]. polythetic [510]. polytomous [1701]. pooling [1194]. Population [70, 1864, 2188, 2408, 1441, 1857, 514, 2595, 1137, 1538, 711, 370, 1858, 2268, 672, 647, 1879, 2067, 2193, 2612]. populations [2166, 787, 2595, 2114, 1427, 985, 90]. portfolio [231, 578, 621, 1978, 1522]. portfolios [1500, 1951]. Position [1021]. positive [1479, 31, 1074]. positivity [286]. possibilistic [691]. possible [88]. possibly [2025, 248, 1014]. Post [386, 648, 1538, 2393, 2193, 1255]. post-intervention [1255]. post-processing [2393]. post-stratification [648, 2193]. Post-stratified [386, 1538]. posterior [317, 2562, 1227, 251, 1040, 1591]. posteriori [980]. posteriors [1529]. potential [2602, 128]. potentially [601]. Potts [1693]. poverty [2251, 2102, 849, 465, 853]. Power [1412, 1143, 2047, 568, 1834, 426, 2396, 753, 325, 2424, 2547, 582, 869, 1132, 915, 1598, 198, 412, 936, 438, 1336, 1086, 1053, 1041, 1466, 2035, 1372]. power-law [2424]. Powered [2248]. powerful [1090, 2146, 391, 1905]. powers [14]. Practical [357, 521, 439, 838, 1922, 379, 1423, 1469]. practice [1762, 1054]. PRAM [1573]. pre [2298, 981, 310, 1255, 515, 2163]. pre- [1255]. pre-binned [310]. pre-cancer [2163]. pre-selection [981]. pre-specified [2298]. pre-test [515]. precision [1817]. predict [392]. predicted [394]. Predicting [1936, 938]. Prediction [2502, 2596, 1652, 1878, 172, 1108, 167, 857, 1565, 2318, 2198, 272, 2509, 479, 2476, 2460, 2321, 284, 149, 2222, 2398, 1355, 1036, 2496, 1957, 619, 1444, 2237, 1191, 216, 2356]. predictions [2058]. Predictive [729, 1360, 1468, 452, 1558, 730, 582, 587, 1830, 713, 72, 2489, 2204, 774, 1156, 721, 1591]. predictor [1669, 1861, 450, 215]. predictor-dependent [215]. predictors [45, 2387, 1242, 1138, 562, 2438, 140, 860, 2475, 1368, 2548]. preference [1391]. prefetching [259]. presence [886, 1656, 1402, 277, 1631, 1363, 1196, 2175, 1911, 2149, 1795, 637, 1501, 1958, 134, 1776, 84, 1444, 919, 10, 1070, 1520, 2446]. presence-only [1070]. presence/absence [134]. preserving [1996]. pretest [523]. prevalence [1536, 716, 1212]. prevalences [1536]. prevention [1995]. preventive [967, 1539]. preweighted [2272]. prices [1027]. pricing [1428, 906, 295, 1513]. prime [72, 590]. Principal [546, 2153, 1827, 969, 1446, 2227, 271, 1100, 1528, 524, 2136, 1847, 766, 1744, 2371, 344, 172, 328, 1182, 108, 981, 379, 2160, 312, 2419]. Principle [124, 343, 1206]. Prior [1991, 1335, 770, 2552, 33, 2135, 2173, 1287, 1693, 2589, 2600, 2582, 201, 1880]. prior-induced [1880]. priori [311]. prioritized [391]. priors [1155, 937, 2520, 2332, 2304, 1396, 1160, 2107, 2523, 2560, 1883, 2372]. privacy [1573]. Probabilistic [1471, 2331, 1717, 696, 2472, 2223]. probabilities [1608, 1298, 1039, 1227, 1273, 1256, 448, 469, 1879, 763]. probability [1308, 2493, 706, 1073, 1588]. probable [2035]. probit [368, 338, 656, 670, 1154, 1413, 98, 1244, 1739]. probit-linear [1739]. problem [2052, 1488, 939, 523, 1060, 198, 156, 1191]. Problems [858, 327, 512, 67, 2041, 1843, 785, 1732, 2551]. procedure [1477, 2009, 1464, 1923, 1387, 2368, 2566, 1047, 1836, 2450, 2114, 1642, 398, 486, 1641, 1854, 1905, 318]. procedures [1118, 2267, 325, 2175, 1219, 708, 944, 40, 72, 593, 563, 265, 99, 750]. process [949, 1848, 639, 1057, 1422, 865, 2424, 939, 507, 2366, 730, 2039, 2101, 222, 510, 16, 1691, 2384, 2574, 339, 1619, 2030, 2357, 2287, 373, 1568, 2176, 2475, 1833, 214, 822, 2612, 2345]. processes [1482, 628, 1179, 19, 1383, 893, 2295, 367, 58, 1938, 2292, 1940, 22, 224, 1801, 244, 491, 2036, 2043, 1507, 1426, 1236, 2605, 586, 1463, 1957, 1209, 416, 881, 1022, 1089, 2434, 1543, 469, 2205, 563, 2607, 1233, 1292, 1044, 498, 1975, 1346]. processing [2393]. processors [1411]. Procrustes [484]. product [2260, 2010]. production [769]. products [2496]. professional [2306]. Professor [700, 2468]. Profile [580, 1237, 1117, 1797]. profile-empirical-likelihood [1117]. profiles [1304, 1923]. prognosis [836]. prognostic [2206, 2208]. programming [2310, 1203, 1389, 1721, 292]. progress [1133]. progression [643]. progressive [152, 1631, 654, 816]. Progressively [429, 778, 560, 167, 1120, 1736]. projected [496, 1339]. projection [132, 2375, 2249, 2240, 1078, 2316, 1463, 293, 934, 1745, 420]. projection-based [132]. projections [1051, 2444, 2300]. projective [2127]. promotion [868, 2369, 620]. prone [1069]. propensity [468, 1832, 2426, 1694]. Properties [1251, 960, 193, 2332, 520, 1753, 1447, 319, 242, 2307, 412, 536, 1209, 1543, 42, 32]. property [2255, 600]. proportion [550, 660, 217, 2241, 2193]. proportional [548, 2190, 2289, 371, 1682, 2445, 976, 1532, 1458, 348, 1135, 1445, 1812, 692, 1451, 1660, 189, 1292, 1854, 1877]. proportionality [1912, 1556]. proportions [849, 568, 1232, 424, 215]. proposal [2141, 1007]. Proposals [1057]. Proschan [967, 1539]. prostate [643]. proteomic [2491]. proximal [1491]. proximity [223, 1751]. proxy [1642]. pruning [2093, 1649]. pseudo [2292, 2553, 1609, 1679, 657]. pseudo-adaptive [657]. pseudo-likelihood [1679]. pseudo-marginal [2292, 2553, 1609]. pseudolikelihood [1676]. Psycho [1374]. Psycho- [1374]. Public [232]. pulse [1029]. pulse-like [1029]. pure [1100]. purely [182]. purpose [2212, 736]. Pursuit [2462, 2316]. puzzle [1483].

QR [2279]. QR-based [2279]. Quadratic [350, 973, 2598, 86, 1389, 1843, 2051, 1423, 1861, 636]. quadrature [1837, 1599, 657, 2170]. qualitative [52]. quality [2435, 639, 1865]. quantal [158]. quantify [1782]. Quantifying [2482, 1662]. Quantile [820, 683, 2458, 2276, 109, 1725, 1155, 2020, 810, 684, 2396, 1762, 1835, 1723, 2462, 2135, 2412, 540, 2200, 1676, 679, 2549, 1169, 1869, 2493, 1132, 1680, 2519, 685, 1153, 1263, 2144, 2265, 1987, 2297, 991, 2183, 853, 878, 758, 1014, 686, 682, 2530, 938, 625, 922, 1005, 110, 2257, 1798, 2548, 323, 1461, 1576, 1874, 2128, 2419, 2404, 361]. quantile-filling [1461]. quantiles [212, 787, 853, 386, 1577, 1427, 1538, 1414, 4, 133, 1537, 2317]. quantitative [2188, 55, 1869, 1185, 497]. quantization [1762]. Quasi [520, 2067, 2427, 1420, 1424, 2043, 2147, 76, 833, 1709]. quasi-efficient [1424]. quasi-independence [1420]. Quasi-likelihood [2427, 2147, 833]. quasi-maximum [1709]. Quasi-negative [520]. quasi-regression [76]. Quasi-systematic [2067]. QuEST [2261]. quoting [232].

R [80, 1372, 1373, 1407, 63, 1376, 1379, 1367, 1371, 34, 1530, 712]. R-geepack [1530]. R-multgee [1530]. R-repolr [1530]. racial [643]. radial [2466]. radius [367]. RADWT [2527]. Randle [1927]. Random [1583, 1150, 1081, 1606, 2000, 1878, 620, 1589, 399, 886, 2056, 2464, 1398, 2437, 589, 2295, 1726, 2423, 1376, 317, 1361, 2286, 2135, 1652, 1655, 739, 2476, 1051, 1152, 1494, 1902, 2224, 1750, 171, 1265, 2443, 1091, 565, 374, 392, 2057, 706, 2588, 754, 1429, 1154, 2002, 1282, 2110, 2041, 202, 2189, 2079, 742, 1604, 1253, 1741, 1318, 333, 455, 1779, 2525, 2010, 574, 1355, 1370, 1234, 150, 324, 1577, 2026, 2072, 1463, 1907, 764, 1129, 355, 170, 694, 2268, 472, 2276, 2233, 947, 341, 2518, 1534, 618, 1437, 1393, 12, 118, 720, 1375, 2402, 169, 832, 77, 2356]. random [498, 221]. random-effect [2041, 1741]. random-effects [374, 1154, 1779]. random-field [2224]. random-intercept [2057]. random-projection [1463]. randomized [176, 183, 1091, 1427, 1536, 714, 424, 1254]. randomly [361]. range [865, 365, 769, 2271, 1986, 1341]. Rank [101, 124, 283, 2081, 517, 1038, 2594, 1435, 2025, 1090, 534, 198, 845, 1018, 2083, 2294, 309, 1256, 406, 1368, 2122, 1087, 2266]. ranked [471, 2256, 2541]. Ranking [708, 1488, 160, 817, 382, 2330]. rankings [1090, 1344]. Rao [70]. rare [1608, 2198, 330, 1184, 2207]. rate [757, 923, 1197, 753, 1009, 1902, 1495, 92, 377, 855, 1639, 1123, 1020, 620, 586, 525, 2103, 369, 1120, 93, 715, 137, 162, 644]. ratemaking [931]. rater [17]. rates [360, 1654, 465, 2187, 1862, 1432, 1555, 1508, 1509, 106, 1534, 1733, 346]. rating [250, 1560]. ratings [1478]. Ratio [1972, 2545, 1222, 1597, 1643, 2492, 706, 1862, 403, 1507, 2506, 660, 1112, 1704, 1310, 384, 2242, 1261, 1278, 201, 1280, 2417, 2191, 763, 930, 1918, 157, 498, 1549]. ratio/information [1310]. ratios [1882, 1337, 123, 1257, 835, 428, 53]. rBeta2009 [1376]. RBF [2471]. RC [1405]. RcppArmadillo [1379]. re [1342]. re-parametrization [1342]. reaction [627]. Real [1969, 246, 241]. Real-time [1969]. realised [1519]. reality [695]. Realized [1515, 631, 229, 1939, 911, 2528]. really [98]. reasons [722]. Recalibration [2393]. recall [1780]. recall-based [1780]. recapture [370, 647, 1879]. receiver [1800]. reciprocity [2354]. recognition [132, 1211]. reconciled [1887]. reconsidered [177]. reconstruction [1398, 676]. record [1651]. records [1180, 1554]. recovery [1839]. rectangular [223]. Recurrence [2447]. recurrent [1118, 639, 1995, 1009, 15, 2187, 1661, 1733, 644]. recursion [1830]. Recursive [1259, 904, 1402, 743, 2093, 1486, 735, 2560, 1032]. Reduced [2025, 1368, 1646, 406, 177]. reduced-bias [1646]. Reduced-rank [1368]. Reducing [1588, 1973]. Reduction [1349, 2166, 1446, 1057, 349, 2394, 2387, 1265, 1138, 2127, 1462, 2380, 1952, 2192, 1896, 2084, 1033, 2124, 579, 102, 2160, 2109, 993, 2582, 988, 1765, 189, 1614, 2252, 1082, 2615, 105, 2419, 1827, 2033]. reduction-based [2033]. Reference [24, 1883]. referenda [2510]. refinements [963, 938]. reflecting [2295]. reflective [2319]. reformulation [1222]. Regarding [80]. Regime [1518, 544, 249, 2447, 1026, 945]. regime-dependent [544]. regime-switching [249]. regimens [1984]. regimes [2105, 162]. Region [583, 752, 1491, 27, 533, 461]. Region-based [583]. regions [1415, 1113, 145, 686, 1807, 962, 32]. register [850]. register-assisted [850]. registration [2481, 2579, 1992]. Regression [1703, 2289, 402, 839, 2149, 1780, 1906, 1660, 1640, 2532, 2526, 174, 267, 2106, 885, 1151, 1818, 1155, 2020, 1259, 1119, 773, 2060, 804, 684, 436, 823, 931, 2396, 656, 2594, 1897, 45, 1696, 364, 183, 89, 838, 1726, 1412, 2323, 2327, 512, 380, 474, 2602, 1792, 381, 1723, 1544, 2286, 622, 492, 414, 2255, 1328, 237, 153, 2135, 2412, 2527, 2108, 2217, 2211, 371, 2154, 2148, 777, 240, 582, 1218, 2137, 2025, 1116, 1104, 702, 278, 1152, 2200, 101, 820, 1676, 1718, 1620, 679, 1385, 1313, 430, 279, 2549, 1685, 1296, 800, 1169, 1457, 2576, 1617, 1659, 2271, 100]. regression [567, 1439, 2036, 1869, 2368, 1060, 186, 2346, 1201, 1735, 1062, 1178, 2457, 2158, 1550, 143, 1680, 2519, 685, 1153, 1263, 1832, 1196, 283, 1744, 2371, 476, 1825, 910, 1504, 535, 1666, 1688, 1160, 2144, 541, 1231, 2341, 412, 1144, 1318, 1911, 348, 948, 1431, 1472, 2053, 2059, 1314, 435, 683, 1018, 1722, 2097, 2305, 389, 1555, 781, 2183, 2021, 2458, 330, 2388, 1168, 1106, 1830, 878, 1624, 602, 1566, 574, 290, 1888, 1355, 1808, 1336, 38, 758, 2577, 1579, 1014, 294, 673, 663, 1, 527, 747, 686, 682, 1793, 2363, 1243, 819, 1414, 256, 297, 98, 2030, 390, 598, 2279, 1425, 1105]. regression [1819, 938, 2071, 1419, 1207, 1208, 462, 1527, 2276, 2611, 2160, 39, 2165, 680, 1312, 1861, 2287, 532, 2353, 561, 1195, 110, 1393, 818, 765, 1130, 1846, 774, 99, 750, 718, 1121, 1595, 1898, 720, 262, 517, 945, 665, 837, 1283, 2075, 2118, 2257, 2475, 2521, 640, 161, 616, 37, 1589, 1725, 2551, 2381, 147, 76, 1773, 1311, 1893, 833, 2456, 2548, 313, 323, 1845, 1068, 1284, 1569, 1653, 1840, 2504, 340, 1874, 2128, 2266, 2419, 605, 361, 97, 2033, 2163, 2228]. regressions [2502, 2520, 1531, 874, 1138, 2123, 1853, 125, 2580, 1791, 814, 1798, 215]. regressive [1709]. regular [1064, 1997, 2220, 2096, 299]. Regularization [311, 1576, 684, 802, 602, 508]. Regularized [1869, 2604, 406, 1881, 1277, 2003, 1121, 2479]. reinforced [2223]. Reinforcement [1398]. reject [479]. rejection [2334, 2364]. related [1109, 2604]. relational [1228]. Relations [560]. relationship [1374, 1632, 497]. relationships [2, 611]. relative [1984, 2032, 1591, 1881]. relaxation [2224, 2409]. RelaxMCD [85]. relevant [2272, 291]. Reliability [967, 399, 1656, 2259, 1431, 2139, 778, 1461]. REML [404]. REML-based [404]. remote [809]. Removing [1026]. renewal [2357, 1833]. repairable [1136]. Repeatability [80]. Repeated [597, 415, 1088, 2199, 1676, 704, 2400, 182, 841, 1048]. repeated-measurement [1048]. repeated-measures [415]. replacement [834, 629]. replicates [2512]. replication [2253]. repolr [1530]. reported [123, 384]. Representation [223]. representations [670, 2312, 2439]. representative [1708]. Reproducibility [3, 80, 1344]. reproducing [2436]. reproductive [205]. repulsion [1341]. require [1274]. requirements [1091, 568]. resampling [2267, 2127, 284, 605]. resampling-based [2267]. Research [699, 2250]. reservoir [2597]. Residual [2167, 733, 1109, 1691, 574, 1714, 2017, 2541, 50]. resistance [1305]. Resistant [1051]. resolution [2002, 1863]. respiratory [1016]. Response [1071, 2370, 2219, 1373, 1687, 838, 1399, 1544, 2023, 2476, 1386, 2069, 2024, 1265, 2443, 2576, 52, 729, 533, 541, 1223, 1083, 2438, 1579, 2022, 2237, 2353, 376, 304, 1041, 2415, 1697, 1183, 1877, 1775, 2163, 592, 2216]. Response-adaptive [2370]. responses [1323, 634, 2219, 54, 1171, 2594, 1696, 447, 2394, 2476, 82, 1878, 587, 172, 2079, 2301, 2540, 2329, 443, 2061, 1669, 2054, 297, 1876, 1401, 1716, 2518, 497, 118, 2017, 2012, 1048, 2404]. restricted [2505, 2506, 1405, 2280, 837, 580]. restriction [515]. results [986, 549]. Retracted [224, 22, 2]. Retraction [611]. retrospective [1632]. return [907, 1956]. returns [1064, 1942]. reversed [1532]. Reversible [1416, 1346, 2437, 1404, 444, 1907]. Review [74, 1307, 1001]. Revisiting [1945]. reweighted [1975]. rho [1521]. RHSBoost [2157]. rich [1945]. richness [1794]. Ridge [1636, 2034, 770, 1218]. Ridge-based [1636]. right [933, 2311, 2538, 277, 464, 1914, 2115, 824, 1445, 663, 456, 1438, 2300]. right-censored [663, 456, 1438, 2300]. rigidities [241]. Risk [897, 1133, 898, 894, 2492, 946, 1984, 1610, 2177, 1939, 2482, 1500, 707, 2189, 1956, 1908, 123, 384, 598, 81, 938, 428, 1964, 2532, 117, 1825, 1961]. risk-return [1956]. risks [1864, 2186, 429, 1631, 674, 199, 2115, 1987, 2059, 2363]. river [1345]. RMCMC [1603]. road [1432]. Robust [266, 88, 2106, 1818, 2060, 804, 2227, 269, 882, 1815, 1171, 2090, 270, 1897, 235, 1488, 1167, 2311, 474, 2379, 274, 554, 279, 112, 1685, 2490, 280, 953, 2040, 1752, 983, 1639, 96, 790, 285, 1073, 1911, 1998, 2299, 1172, 876, 389, 276, 1319, 330, 1168, 2131, 2474, 1170, 617, 2472, 2124, 298, 841, 1819, 462, 258, 2165, 1312, 2082, 1817, 11, 1199, 1607, 1459, 1204, 1131, 1311, 2456, 2455, 2416, 2615, 975, 267, 2111, 2533, 886, 1012, 773, 892, 746, 864, 2535, 2418, 381, 1164, 1297, 423, 278, 1103, 1921, 2443, 2136, 282, 2158, 284, 1820, 535, 2214]. robust [1953, 1635, 834, 2225, 294, 1727, 1425, 1527, 1564, 940, 647, 561, 818, 302, 265, 373, 161, 776, 14, 2581, 552, 2216]. robustly [715]. Robustness [2197, 2215, 304, 998, 2314, 748, 702, 2508, 1668]. ROC [473, 1546, 1558, 1545, 1115, 390, 598, 1145, 1715, 2507]. role [1921]. Roman [1447]. Root [815, 1477, 1931, 2009, 1492, 869, 1570, 1650, 1972, 1512, 1974, 258, 1965, 1249]. roots [1502]. Rosenblatt [2360]. rotations [1340]. rough [696]. rounded [834]. rounding [1165]. row [2123, 2266]. rule [2585, 1821, 924, 657]. rules [132, 1697]. run [1480, 1056, 1970, 533, 1956]. RV [2478]. RWM [2141]. RWM-within-Gibbs [2141].

Saddlepoint [2111, 1578]. SAEM [783, 1868]. safe [2140]. safety [2035]. sales [2591, 1026]. SAMC [1332]. Sample [1689, 1091, 2296, 1536, 714, 1438, 671, 1451, 1048, 399, 268, 2331, 823, 2385, 126, 2565, 1621, 1412, 2423, 325, 1327, 2538, 35, 2585, 523, 1352, 242, 2254, 900, 2307, 2250, 2123, 587, 729, 2508, 708, 198, 411, 1202, 876, 877, 1668, 1106, 2054, 194, 627, 568, 379, 1340, 785, 123, 694, 1120, 1437, 53, 1375, 2402, 169, 2070, 2317, 1709, 1255, 1697, 1894, 32]. sampled [2472]. sampler [1396, 685, 1563, 726]. samples [1194, 1180, 483, 471, 535, 2583, 522, 1080, 1538, 167, 81, 2268, 2242, 2091, 1807, 1736, 719]. Sampling [822, 1761, 634, 795, 1702, 937, 1656, 1933, 1398, 64, 674, 1770, 2324, 1441, 321, 1862, 872, 1505, 166, 2248, 444, 760, 1719, 1427, 1632, 1256, 254, 1529, 2312, 370, 123, 384, 736, 81, 1274, 593, 2233, 2170, 188, 2597, 2364, 2067, 778, 331, 2495, 2541, 1569, 318, 13, 2039]. Sandwich [1623]. SAR [2320]. Saralees [1075]. SARS [1016]. SAS [1530]. SAS-GENMOD [1530]. satellite [980]. saturated [187]. Saunders [923, 979, 440, 646, 125, 412, 522, 1144, 1128, 40, 673, 750, 460, 1595, 1898, 958, 37, 24, 516]. Saunders- [923]. Savage [201]. SCAD [350]. scalable [2286, 1730]. scalar [1932, 2584, 1792, 2476, 1620, 1296]. scalar-on-function [1792, 1296]. scalar-on-functions [1620]. Scale [565, 266, 1323, 830, 269, 1325, 746, 364, 1867, 2424, 2140, 674, 1495, 2610, 1132, 1989, 2001, 2181, 1830, 997, 959, 2536, 1871, 311, 2071, 1564, 532, 1883, 2382, 387, 2555, 2402, 1753, 2042, 451]. scale-mixture [1830]. scale-space [997]. scaled [2444]. scales [1855, 250]. scaling [103, 2450, 1320, 2268, 2258, 382, 1854]. scan [2506, 2421, 638, 1641, 1466]. scans [2344]. scatter [1622, 2161, 2131]. scenario [1944]. scheduling [2344]. scheme [1621, 1857, 123, 2176]. schemes [2553, 828]. Schindler [990]. Schisterman [609]. Schmidt [2252]. SCOMDY [1509]. Score [1099, 120, 1479, 468, 1934, 1832, 283, 707, 2426, 1405, 1842, 1694]. scores [2388, 1924]. scoring [2334, 696, 2516, 168, 988]. Screening [1540, 2311, 2494, 2566, 2079, 2253, 2240, 1205, 1844, 2153, 2308, 2438, 394, 2518, 2012, 2338, 2122, 2488]. sea [1108]. seamless [712]. search [195, 1329, 562, 1391, 1392]. Searching [1194]. Seasonal [1236, 2094, 1492]. seasonality [1286, 1026]. Second [62, 699, 492, 2273, 52, 1727, 169]. second-order [492, 2273, 52]. sectional [2208]. sectionally [258]. seed [2110]. seeded [1082]. seemingly [262, 814]. segmentation [200, 416, 532]. segmentation-based [532]. segmented [512]. selected [172, 2130]. Selecting [1064, 766, 273, 1353, 2038, 1156]. Selection [1349, 655, 1612, 1600, 1625, 2169, 1155, 2020, 2339, 1119, 83, 773, 1484, 684, 770, 2055, 231, 1350, 2267, 326, 1528, 2552, 2186, 380, 1009, 2585, 479, 939, 236, 218, 2286, 1937, 1352, 1134, 551, 1218, 1363, 2151, 1623, 1685, 1296, 794, 280, 1169, 1457, 1122, 281, 2032, 1081, 1606, 2019, 2368, 1062, 1970, 2040, 2420, 124, 1878, 685, 1263, 383, 1666, 1991, 845, 562, 2041, 1288, 621, 2192, 285, 362, 1458, 1690, 1741, 1853, 2173, 2432, 154, 411, 1107, 1954, 2107, 2301, 2540, 1135, 2450, 1083, 2076, 1319, 1635, 2152, 1290, 287, 521, 1106, 2054, 602, 1566, 290, 291, 1102, 1797, 1356]. selection [293, 108, 924, 295, 981, 175, 2096, 712, 1229, 297, 2397, 2386, 2124, 298, 603, 1264, 727, 1357, 815, 2513, 2523, 2011, 1005, 1406, 1403, 519, 1633, 335, 1645, 265, 129, 1978, 1965, 720, 303, 213, 1186, 1734, 2521, 2402, 2213, 832, 1131, 1880, 664, 1697, 2416, 2504, 1576, 1613, 1724, 1877, 2419]. selector [2089]. selectors [999, 935, 1751]. Self [2201, 786, 2434]. Self-controlled [2201]. self-exciting [2434]. Self-tuning [786]. SEM [875]. Semi [1539, 751, 2054, 57, 2259, 63, 2186, 371, 242, 2254, 363, 1305, 1721, 854, 518, 292, 1262, 2363, 1156, 1070, 1212, 2495, 1745]. semi- [2254]. semi-competing [2186, 2363]. semi-definite [292]. semi-infinite [1721]. semi-Markov [63]. semi-mixed [854]. Semi-parametric [1539, 2054, 57, 2259, 371, 242, 363, 1305, 1262, 1212, 2495]. Semi-supervised [751, 518, 1156, 1070]. semi-varying [1745]. semidefinite [2310]. Semiparametric [322, 1934, 825, 2431, 2108, 777, 2443, 567, 493, 1584, 1954, 2554, 255, 2363, 2611, 1533, 1278, 2378, 2042, 207, 1893, 2358, 340, 2507, 427, 756, 205, 606, 669, 2430, 1085, 1055, 679, 245, 1122, 2036, 1650, 1680, 1154, 1125, 1288, 351, 528, 2341, 957, 1431, 2053, 2514, 1830, 2587, 1337, 2164, 1449, 2116, 217, 2287, 961, 2560, 616, 142, 173, 1733, 1067, 1919, 97, 442]. Semiparametrically [381]. semivariogram [756]. semivarying [2521]. sense [600, 1190]. sensed [809]. Sensible [2392]. sensitive [2024, 1202, 1536]. Sensitivity [587, 331, 402, 1525]. sensor [2282]. Separable [955, 749]. separately [1178]. separation [1055, 1098]. separators [590]. sequence [801, 1196, 1855, 455, 1460, 128, 498, 221]. Sequential [2413, 996, 2083, 1413, 1030, 2434, 717, 1625, 1324, 2385, 1377, 939, 1056, 872, 1526, 2294, 1803, 2577, 1889, 2074, 1542, 1032, 1840]. sequentially [996, 1890]. Serial [71, 1233, 1493, 239, 1293]. serially [1370]. seriation [1190]. Series [176, 1966, 399, 2339, 628, 1834, 1484, 2318, 753, 31, 1467, 1915, 274, 2135, 1029, 2201, 243, 1149, 247, 1970, 2465, 2531, 543, 1887, 503, 983, 855, 96, 1849, 1971, 1080, 1295, 1276, 936, 1986, 2117, 438, 2178, 8, 1262, 2121, 1681, 170, 2027, 926, 258, 1053, 2534, 1286, 982, 716, 261, 459, 950, 2362, 2037, 2515, 2414, 1975]. Service [1971]. Set [80, 542, 36, 696, 1361, 471, 3, 984, 789, 1256, 140, 508, 1344, 1534, 497, 1448, 2541]. sets [1095, 1361, 689, 144, 1374, 1944, 1066, 67, 248, 1855, 2595, 288, 123, 694, 1776]. setting [2375, 1454]. settings [1667]. several [1194, 2361, 487, 787, 2114, 1257, 2091, 90]. sex [1262]. Shape [923, 36, 837, 557, 2610, 567, 1892, 2181, 1214, 2402, 859]. shape-constrained [567, 859]. shaped [759, 1123, 1020, 1740]. shapes [1043]. Shared [1999]. sharing [2614]. Sharpening [535]. Shewhart [87]. Shewhart-type [87]. shift [204, 629]. shifted [1583, 2498, 2116]. shifts [157]. short [1480, 233, 1502, 1337, 715, 2208]. short- [1337]. short-run [1480]. short-term [715, 2208]. shortfall [897, 1825, 681]. shot [658]. Should [2250]. Shrinkage [1135, 508, 885, 1119, 415, 1464, 480, 203, 2040, 1860, 2558, 1263, 751, 2001, 2107, 1019, 1959, 688, 1002, 636, 1667]. Shrinkage-based [508]. Shtarkov [2058]. sided [21, 953, 1008, 449]. sides [1369]. Siegel [914]. Sieve [2190]. Signal [1025, 96, 1721]. signaling [2491]. signals [1848, 1029, 565, 1276, 2256, 1028]. Signed [2294, 2594, 534]. signed-rank [2594, 534]. significance [219, 2478, 1669, 994, 346]. significant [549]. SIMD [1719]. similar [8]. Similarity [324, 970, 111, 919, 573]. Simple [237, 900, 605, 267, 2080, 2343, 542, 1011, 2009, 1941, 199, 1735, 143, 336, 531, 812, 1913, 660, 1959, 1715]. simplex [610]. Simplicial [1847, 1927]. Simplified [141].

Simulated [934, 1528, 1937, 1677]. Simulating [985, 2233, 1075, 977]. Simulation [435, 332, 1334, 1772, 989, 357, 1410, 1122, 1829, 1760, 2002, 1291, 1837, 1129, 98, 458, 733, 106, 77, 221]. Simulation-based [435, 1334, 1772]. simulations [712]. simulator [1841]. simulators [2605]. Simultaneous [684, 2527, 2452, 1071, 1617, 143, 2519, 707, 1147, 2192, 2301, 1273, 2277, 1876, 2530, 1257, 1046, 2091, 2513, 109, 1568, 2412, 2066, 921, 901, 984, 290, 1973, 1924, 302, 1992]. simultaneously [2298, 798, 1178, 713, 214, 2266]. single [965, 1117, 1069, 2531, 1153, 1288, 2345, 348, 2297, 2573, 2593, 2554, 764, 1235, 1208, 1607, 994, 1224, 1068, 2427, 1877, 2404]. single-index [965, 1117, 1069, 2531, 1153, 1288, 2297, 2573, 764, 1235, 1607, 994, 1224, 1877, 2404]. single-indexed [1068]. singly [1240]. singly-ordered [1240]. Singular [1371, 1407, 22, 224]. Sinh [1595]. sinusoidal [1037]. SIR [579]. Sivashinsky [2350]. six [855]. sixth [884]. Size [412, 399, 823, 1412, 325, 35, 2585, 144, 1689, 601, 869, 2512, 1091, 2296, 514, 2595, 568, 1536, 2027, 2268, 714, 1438, 1451, 1448, 174, 1048, 1255]. sizes [942, 183]. skew [345, 269, 2396, 624, 1440, 364, 851, 193, 1944, 470, 1075, 1137, 1547, 909, 175, 1454, 595, 977, 1317, 2461, 1562, 122]. skew- [1944, 1547, 2461]. skew-Laplace [595]. skew-normal [345, 269, 624, 1440, 364, 1075, 1137, 977, 122]. skew-normal/independent [122]. skewed [1838, 867, 2061]. skewing [797]. Skewness [1940, 2316, 436, 544, 1498, 1961]. Skewness-based [2316]. slash [1151]. slash-elliptical [1151]. sleep [1730]. Sliced [2228, 1544, 1201]. slopes [1091, 1318]. Small [2331, 1565, 852, 2391, 849, 1926, 851, 2324, 111, 1050, 206, 1325, 1850, 857, 2102, 850, 848, 1352, 483, 465, 1066, 2250, 2123, 535, 708, 1738, 1080, 876, 854, 853, 856, 985, 1776, 1732, 2237, 1742, 1709]. small-sample [2123]. Smallest [1113, 752]. SMCVB [1803]. SMILE [1923]. Smirnov [2516, 2070]. Smooth [2168, 351, 85, 190, 2025, 2536, 10, 945]. Smoothed [2055, 1018, 2247, 2317, 185, 2116]. smoother [1760, 1027, 346]. Smoothing [2346, 1054, 2315, 928, 274, 1984, 101, 112, 794, 2236, 343, 405, 1738, 790, 285, 1096, 1182, 957, 1955, 2580, 2147, 332, 518, 682, 73, 2119]. smoothly [1490]. snowball [2268]. Social [2, 611, 2390, 1726, 1632]. socio [2306]. socio-professional [2306]. soft [485]. software [1367, 62, 2139]. solution [2058, 156, 2257]. solutions [67]. Solving [1522]. Some [507, 319, 2175, 1209, 593, 990, 607, 2063, 412, 1097]. sort [1038]. source [1055, 2529]. South [2432]. Sovereign [1478]. space [1323, 1266, 2413, 2585, 1938, 1948, 565, 332, 880, 1037, 355, 997, 1089, 1228, 2382, 1549]. spaced [2178, 926]. spaces [1847, 2579]. Spanish [2410]. spanning [759, 2011, 1740]. Sparse [1754, 2094, 1816, 1203, 1903, 1744, 2371, 2398, 286, 2117, 1121, 1358, 262, 988, 2548, 2080, 2491, 635, 1923, 2272, 2069, 2136, 1062, 2260, 829, 1114, 1587, 2441, 1281, 1917, 2160, 2165, 1199, 1186, 1591, 2351, 2485, 572, 2356, 2266, 97, 2509]. Sparse-Group [2509]. sparse-specific [1923]. sparsity [1869, 2386]. Spatial [1049, 2285, 1444, 466, 889, 656, 1095, 1057, 2557, 367, 852, 759, 47, 1984, 604, 730, 737, 470, 363, 1282, 2583, 2432, 2506, 966, 2564, 2421, 111, 431, 2384, 2554, 1791, 252, 985, 1411, 594, 1782, 2068, 638, 1856, 1709, 1466, 2358, 1740]. spatially [2391, 1066, 191, 405, 1370, 630, 1449, 1749]. spatio [507, 2482, 1050, 1022, 2138]. spatio-temporal [507, 2482, 1050, 1022, 2138]. spatiotemporal [1989]. spatiotemporally [2350]. Spearman [1521]. Special [1349, 305, 1976, 227, 1815, 1966, 848, 1322, 2487, 689, 1359, 1164, 1804, 2210, 2497, 1367, 1380, 1786, 1785, 858, 265, 884, 1302, 2470, 679, 62, 2225, 1025]. species [1794, 1698, 1345, 822]. specific [798, 1923, 2273, 2306, 2203, 2059, 1812]. Specification [902, 468, 2009, 1496, 800, 1329, 82, 2360, 811, 56]. specifications [65]. specificity [1525]. specified [2298]. speckled [437]. spectra [378]. Spectral [1931, 2491, 906, 1858, 715, 2037, 2515]. spectrometry [2073, 109]. spectroscopy [838, 1557, 285, 2163]. Spectrum [1371, 623]. speed [1471, 894]. Speeding [1527]. sphere [2062]. spherical [2382]. spiked [2484]. spillovers [234]. Spline [1220, 833, 427, 89, 928, 2135, 2346, 1125, 1096, 1182, 1106, 856, 1791, 647, 665]. splines [1548, 1806, 548, 1304, 2369, 567, 405, 2519, 1168, 1624, 2530, 1007, 450, 110, 1537, 2119]. Split [845, 451, 306, 1392, 940, 1962]. split-plot [306, 1392, 940]. Split-SV [1962]. spread [1483]. spreads [1495]. SPSS [1530]. SPSS-GENLIN [1530]. spurious [1196]. square [2452, 72, 2608, 1562, 2410]. squared [853, 619]. squares [1266, 1350, 1450, 47, 1215, 2319, 800, 307, 1299, 308, 504, 910, 1675, 328, 667, 432, 292]. stabilisation [70]. stability [1350, 234, 1352, 1512, 498]. stabilization [802]. stabilized [2087]. Stabilizing [1285]. Stable [1605, 1610, 1150, 1487, 1504, 1857, 513, 906, 912, 1943]. stage [2239, 939, 242, 2517, 528, 435, 2114, 130, 1241]. stage-sequential [939]. stagewise [2532, 2374]. Stahel [661, 420]. Stan [2469]. standard [823, 1239, 1972, 561]. Standardization [666]. standardized [2221, 106, 1130]. Stanley [2468]. STAR [1477]. start [1254]. Stat [80]. State [1323, 1938, 1266, 2413, 1440, 2585, 464, 1948, 375, 332, 880, 2072, 1730, 1032]. state-space [1266]. states [1489, 1949, 619]. static [2413, 901, 2191]. Stationary [745, 1650, 2055, 1116, 641, 247, 2465, 744, 1463, 926, 469, 2607, 2515]. Statist [1447, 59, 1268, 1870, 1075, 790, 527, 611, 133]. statistic [2428, 2516, 139, 11, 638, 1701, 1630]. Statistical [1159, 1554, 171, 1219, 780, 789, 1538, 997, 1025, 170, 982, 1187, 221, 1248, 219, 2052, 327, 1289, 2405, 1359, 1238, 2452, 1367, 62, 662, 1332, 583, 394, 1230, 2521, 1041]. Statistics [1815, 2196, 2468, 651, 699, 2545, 560, 305, 2584, 2538, 689, 1937, 666, 363, 2506, 9, 1971, 1526, 2421, 990, 2447, 1591, 1641, 1466]. Statstic [2063]. status [548, 1703, 2149, 781, 1632, 1660, 407, 37]. steady [1489]. Steel [575]. Stein [480, 203, 1637, 1667]. Stein-type [480, 1667]. step [1012, 1670, 199, 2465, 1153, 591, 336, 944, 2280, 1564, 2485]. step-stress [199, 336, 2280]. step-up [944]. Stepwise [400, 312, 1839, 561]. stereotype [1790]. stickiness [932]. Stochastic [914, 1366, 909, 1543, 387, 1015, 266, 755, 1761, 740, 1482, 2354, 322, 1486, 2559, 2188, 913, 670, 2481, 22, 224, 2553, 1329, 244, 1580, 908, 1950, 1503, 872, 1977, 562, 337, 906, 1079, 254, 1663, 439, 881, 1515, 1516, 2082, 1677, 537, 1559, 626, 1519, 1187, 930]. stock [728, 1027, 887]. stocks [228]. stopping [211]. Storey [963]. strategies [141, 656, 1377, 307, 291, 303]. strategy [1503, 154, 185]. stratification [648, 2193]. Stratified [710, 1770, 1441, 2167, 386, 1538, 835, 1812]. stratum [2262]. streaming [641, 2542, 2558]. streams [2057]. stress [2444, 199, 336, 2280]. stress-based [2444]. Strict [137]. stroke [825, 392]. Strong [1534]. Structural [895, 590, 886, 1480, 1226, 238, 866, 1618, 241, 511, 281, 246, 1069, 1196, 1502, 855, 2164, 1512, 947, 2528, 428, 1965, 573, 162].

Structure [1997, 2491, 2286, 2462, 1912, 1494, 1581, 2087, 749, 1135, 1611, 1779, 1956, 2312, 1408, 1518, 2382, 303, 1600]. Structured [1880, 1990, 2136, 2260, 309, 1030, 2071]. structures [367, 1636, 1199, 930]. student [867, 2228, 441, 249, 909]. Student- [441, 249]. Studies [1773, 1382, 468, 1324, 850, 995, 1298, 35, 2105, 703, 1268, 377, 817, 1555, 1795, 568, 472, 1470, 1780, 102, 1208, 717, 454, 14, 1048, 552]. study [890, 989, 770, 2428, 1497, 1122, 476, 198, 1763, 1837, 2209, 1129, 1158, 390, 106, 497, 718, 161, 925, 2163, 1248, 825]. studying [2102]. sub [2166]. sub-populations [2166]. subacute [392]. subblocks [966]. Subgroup [2591]. subgroups [623]. Subject [2162, 1252, 1687, 954, 1903, 1879, 1861, 563, 1893, 1068]. Subject-wise [2162]. submodel [1350]. subprime [887]. subsamples [238]. subsampling [1734]. subsampling-extrapolation [1734]. subset [1326, 1457, 391]. subsets [379]. subspace [2535, 496, 2249, 1355, 2071, 1775]. subspaces [1437, 1894]. substitution [281]. subtypes [2272]. successful [1959]. sufficiency [1503]. Sufficient [2166, 1138, 457, 2252, 1109, 2394, 2124, 2160, 988, 189, 2028, 2615, 1827]. suitability [134]. summary [1185]. sup [684]. sup-norm [684]. superiority [1897, 2547]. superpositions [244]. Supersaturated [549, 1540, 1388, 1396, 154, 287]. Supervised [803, 2387, 382, 2380, 751, 518, 1156, 1070]. support [433, 479, 687, 2140, 1203, 1331, 1299, 1550, 362, 1108, 667, 1430, 1827]. Supreme [2145]. supremum [2361]. supremum-norm [2361]. SURE [1052]. SURE-tuned [1052]. surface [1386, 52, 1115, 1108, 2507, 2216]. surfaces [2559]. surprise [1690]. surrogacy [556]. surrogate [141, 1925, 743, 128, 1553, 556]. surrogates [9]. survey [795, 1702, 968, 1726, 2423, 2426, 188]. surveys [1202, 671]. survival [1864, 1764, 1252, 2197, 1806, 427, 1804, 607, 2063, 1159, 601, 15, 402, 1439, 784, 2214, 202, 2189, 1584, 948, 1185, 1763, 2308, 1811, 1445, 2209, 630, 1449, 1129, 2363, 1054, 1814, 1470, 2185, 1879, 456, 1533, 1607, 2258, 184, 2122, 1694]. surviving [669]. survivor [202]. susceptibility [444]. SV [1962]. SVM [2137]. SVM-like [2137]. swarm [1415, 859]. switches [1518]. switching [1481, 1933, 153, 1159, 1940, 2236, 900, 329, 2147, 249, 2117, 911]. Switzerland [2209]. symbolic [666, 2271, 1574, 842, 38]. Symmetric [599, 1984, 126, 1487, 2315, 2026, 1340, 1072]. symmetrical [99]. symmetrically [729]. symmetry [1321, 2403, 471, 2046, 2433, 1004]. synchronization [509]. synthesis [2496]. synthetic [596, 2440]. system [399, 1136, 2293, 1603, 390, 313]. Systematic [1663, 2067]. systematically [2477]. systems [1289, 2409, 1184, 1003].

tables [1923, 1225, 1240, 2511, 1913, 428, 1591, 1562, 2410]. Tail [637, 2056, 892, 956, 953, 1951, 1577, 907, 1908, 1343, 1753]. tailed [266, 2056, 650, 801, 95, 908, 1073, 948, 2571, 877, 909, 373, 387]. tails [1482, 1498, 1961]. Tales [953]. tandem [109]. Tang [86]. tapering [1052]. target [2007, 27, 1542]. target-layer [2007]. tau [1420, 2543, 155, 2284]. technique [1310, 1007]. techniques [473, 2250, 1500, 985, 390, 1734]. Teimouri [1075]. temperature [2432, 1108, 2117]. tempered [513, 906]. tempering [2235]. template [1761, 2379, 1297]. temporal [2145, 507, 2029, 2482, 1050, 2178, 1022, 2138, 1190, 1614, 1015]. tensor [2568]. term [1764, 1620, 202, 1337, 1956, 715, 2208]. Terms [80, 1145]. terrorist [1348]. tessellations [357]. Test [95, 1627, 104, 1280, 331, 1477, 1625, 2478, 886, 942, 425, 2545, 354, 19, 220, 1289, 1467, 1298, 2544, 1723, 2601, 471, 277, 796, 1941, 896, 2202, 693, 1746, 2361, 1821, 2346, 17, 503, 403, 1507, 547, 978, 1253, 1691, 812, 1080, 991, 1526, 1777, 194, 586, 1463, 963, 713, 1405, 1512, 84, 1974, 1732, 1261, 1812, 2300, 398, 486, 1437, 883, 573, 671, 1249, 1521, 2417, 2191, 2578, 659, 1845, 930, 1894, 2182, 2429, 2488, 1918, 2603, 552, 157, 498, 1740, 515, 592]. Testing [995, 1399, 2403, 153, 1493, 511, 2408, 2361, 1949, 1069, 1502, 182, 1282, 1953, 522, 1913, 1669, 2284, 1508, 289, 651, 711, 1858, 2242, 428, 2244, 994, 1556, 723, 763, 162, 1321, 1931, 658, 1897, 1167, 325, 365, 801, 2492, 1988, 2009, 2547, 2425, 239, 799, 2368, 2146, 179, 1293, 2297, 2047, 1602, 2046, 410, 768, 1575, 1335, 963, 1501, 1543, 818, 1965, 2347, 2555, 2500, 304, 590, 2033]. Tests [238, 1756, 2111, 2239, 1101, 1171, 1327, 35, 2442, 5, 59, 218, 319, 554, 1492, 2547, 65, 190, 1435, 869, 1090, 1496, 1882, 1570, 186, 1650, 283, 2360, 902, 198, 1240, 412, 1426, 1099, 178, 1679, 1008, 1578, 2433, 713, 1158, 1884, 785, 2608, 694, 2337, 1819, 888, 771, 990, 508, 258, 2325, 53, 1206, 2206, 2070, 1600, 120, 346, 14, 2455, 1087, 1743]. text [34]. textual [2572]. textured [2559]. their [1321, 452, 481, 1756, 2384, 57, 738, 1086, 593]. theoretic [1291, 1884]. Theoretical [1423]. Theory [1834, 1071, 918, 1762, 695, 2482, 708, 295, 938]. Third [1786, 1025]. Three [1670, 1558, 582, 242, 2186, 187, 2492, 604, 1903, 143, 1239, 1115, 587, 1039, 514, 749, 984, 1602, 1042, 1469, 1190, 1032, 1592]. three-arm [2492]. three-dimensional [1190]. Three-group [1558]. three-level [187, 749]. three-parameter [587, 1042]. Three-stage [242]. three-state [1032]. Three-step [1670]. three-way [1903, 1469]. Threshold [1634, 544, 1810, 1690, 1343, 2269]. thresholding [239, 2095]. thresholds [465, 253]. throughput [757, 2398, 394]. Tian [609]. tick [229]. tick-by-tick [229]. ties [2175]. tight [1821]. Tikhonov [311]. Time [230, 1966, 2352, 1403, 2404, 141, 2106, 1967, 740, 2339, 1267, 1848, 1484, 2283, 2318, 1351, 31, 107, 669, 987, 1124, 1490, 1467, 1915, 2031, 2431, 1867, 1969, 1304, 274, 769, 874, 849, 601, 702, 1495, 1029, 1136, 2553, 15, 243, 2369, 1865, 199, 1149, 246, 1091, 247, 1970, 2465, 2531, 543, 1887, 2184, 503, 784, 983, 1675, 855, 375, 96, 780, 2297, 2451, 2514, 1971, 2147, 1080, 1295, 389, 620, 2021, 1276, 1010, 2230, 1986, 2117, 880, 2178, 1501, 8, 1958, 355, 1262, 2121, 395, 170, 1465, 1089, 657, 1470, 1228, 926, 1780, 1419, 556, 258, 2534, 2205, 1286, 947]. time [982, 647, 456, 1730, 261, 517, 1451, 2378, 142, 173, 950, 1045, 833, 2362, 174, 1048, 2037, 2358, 2515, 1087, 2414, 1975, 442, 1549, 1442]. time-averaged [1048]. time-course [1304]. time-dependence [880]. time-dependent [1967, 1451, 1045]. time-discretised [2553]. Time-dynamic [2352]. Time-efficient [1403]. time-series [1262]. time-to-event [702, 2451, 1465, 1780, 556]. Time-varying [230, 2404, 740, 1490, 874, 2465, 2184, 2297, 2147, 2230, 1958, 2358]. times [2108, 868, 1661, 1274, 340, 644, 975]. timing [2400]. Titterington [137]. Tobit [1590, 685, 1083, 253, 1210, 1589]. Tolerance [2035, 21, 1097]. tomography [1693]. tool [1999]. tools [964, 2405]. topic [2572]. topology [2407]. total [2519, 2212, 645]. toxicology [2288]. Trace [2266]. Tracking [2118, 2282, 744, 339, 1279]. tractable [2499]. trade [1956]. trade-offs [1956]. tradeoff [46]. training [1621, 1448]. trait [55, 319, 1591]. traits [1869]. trajectories [2007, 2306, 1306, 1419]. Transdimensional [1698, 1803, 1609]. transects [1596]. transfer [1036]. Transform [1369, 2360]. Transformation [1567, 669, 1226, 2431, 1805, 824, 2514, 1018, 797, 665, 2495, 2134, 2335]. Transformation-based [1567]. transformations [2202, 2263]. Transformed [2191, 1132, 1661, 2234, 2414]. Transforming [2237]. Transition [1796, 945, 1601]. transitions [2154, 1294]. transmission [425]. treat [2492]. treated [693]. treatment [183, 1159, 2254, 2490, 706, 2147, 393, 712, 1959, 1264, 104, 113, 454, 721, 449, 2370, 2526]. treatment-by-gene [721]. treatments [365, 2061]. Tree [1696, 117, 1436, 759, 165, 845, 160, 256, 1814, 1408, 2011, 1875, 2414, 1740]. Tree-based [1696]. trees [1360, 1446, 2251, 1718, 1272, 541, 146]. Trend [1034, 1066, 1570, 1501, 14]. Trendline [114]. trends [47, 1329, 855, 1432, 990]. tri [1659]. tri-linear [1659]. trial [1995, 2385, 2492, 604, 2177, 705, 660, 1553, 556, 722]. trial-level [1553, 556]. trials [2169, 2219, 183, 1326, 2405, 1159, 1091, 784, 2214, 712, 714, 2489, 2204, 556, 424, 376, 718, 2207, 1254, 2370]. triangular [2008, 1536]. trimmed [66, 307, 729, 1675, 758, 292]. trimming [279, 1921, 617, 298]. triplot [1826]. tropical [1108]. true [550, 2241]. truncated [933, 1471, 1541, 796, 2200, 566, 1914, 2115, 844, 1102, 1808, 595, 2611, 361]. truncation [1420, 2442, 2059]. trust [1491]. trust-region [1491]. try [1416]. tube [1550]. Tucker2 [1903]. Tukey [2428, 1964, 1766]. tumor [755, 2101]. tumorigenicity [2149]. tuned [1052]. Tuning [1062, 786]. turning [92]. TVICA [1442]. twin [1810]. Two [1175, 21, 2512, 2465, 1153, 2517, 577, 2257, 1190, 1715, 2485, 1381, 755, 2239, 1897, 2298, 195, 2262, 153, 523, 1912, 44, 1200, 1260, 1781, 2490, 2250, 1225, 503, 1862, 2073, 198, 1282, 1240, 528, 1096, 411, 1099, 435, 1777, 2114, 2209, 568, 1037, 713, 789, 1536, 785, 123, 1214, 1040, 81, 2242, 1392, 1807, 2333, 398, 1437, 952, 1883, 814, 1347, 2070, 1556, 1368, 1894]. two- [1807]. two-compartment [755]. two-component [44, 2209]. two-dimensional [2073, 1096, 1037, 2070]. two-factor [2298, 1200]. Two-group [577]. Two-layer [2257]. Two-level [2512, 2298, 1260, 1099, 1040]. two-parameter [1214, 2333, 1347]. two-part [1912]. two-person [195]. two-phase [1392]. two-piece [1883]. two-sample [523, 198, 123, 1437, 1894]. Two-sided [21]. Two-stage [2517, 2239, 528, 435, 2114]. Two-step [2465, 2485]. two-stratum [2262]. two-treatment [2490]. Two-way [1190, 1225, 1240]. Tyler [41]. Type [1664, 84, 795, 979, 1995, 1487, 152, 1435, 480, 800, 2610, 1580, 649, 49, 1870, 2158, 1166, 87, 316, 535, 1688, 198, 1507, 1738, 547, 978, 1182, 522, 1123, 145, 455, 2388, 732, 2150, 1736, 1667, 1347, 2206, 346, 1975, 560, 429, 336, 149, 167, 719, 564]. type-I [152, 316, 564]. type-II [522, 1736, 560, 429, 336, 149, 167, 719]. types [2201].

U.S. [2145]. Uhlenbeck [244, 881, 537]. Uhlenbeck-based [537]. UK [873]. ultra [2311, 1669, 2458, 2570]. ultra-high [2311, 2458]. Ultrahigh [2240, 2566, 2079, 2308, 2438, 2518, 2122]. ultrahigh-dimensional [2566, 2308, 2518, 2122]. ultrasound [2344]. unbalanced [179, 1065, 2328]. Unbiased [1718, 76, 2093, 2329, 1721]. uncensored [1327]. uncertain [2590, 1432]. Uncertainty [701, 1238, 2140, 241, 2460, 1942, 257]. Uncertainty-safe [2140]. unconditional [398]. undercounts [850]. underdispersed [1105]. underfitting [2411]. underreported [297]. Understanding [331, 2559, 1338]. undirected [1230]. unemployment [855]. unequal [183, 1426]. unevenly [2178]. Unified [103, 147, 393, 1378, 603]. Uniform [27, 1415]. uniformity [789, 1158]. unifying [1892]. Unimodal [1397, 453]. unique [1924]. uniqueness [1181]. unit [1477, 1931, 1565, 2009, 1492, 1404, 869, 1570, 49, 1870, 1650, 1502, 1972, 1512, 1974, 258, 1965, 1249]. unit-level [1565]. unitary [1185]. univariate [1477, 1321, 2505, 1892, 1963]. universally [1384]. unknown [1483, 1698, 2286, 2403, 239, 240, 821, 403, 1293, 2046, 731, 2373, 2312, 1207, 1963, 722, 822]. unobserved [1483, 2400, 1811, 619]. unrelated [262, 814]. unreplicated [1200, 356]. unrestricted [1947]. unscented [1266, 1027]. Unsupervised [1134, 2381, 808]. Update [306, 1130]. updating [1603]. uplift [1999]. upper [1821, 1047]. upside [1123, 93]. upside-down [1123, 93]. urban [2434]. Usage [453]. Use [1332, 80, 1321, 921, 2041, 525, 1529, 2204, 2068]. used [552]. useful [1945]. usefulness [1485]. user [1377]. user-friendly [1377]. Using [1996, 66, 2110, 2058, 1355, 1632, 1907, 128, 370, 1776, 266, 2383, 1717, 83, 1372, 1848, 770, 1471, 1222, 2168, 2396, 1351, 2437, 1468, 1541, 1806, 2007, 624, 1936, 1289, 696, 913, 1800, 973, 1794, 365, 1436, 479, 687, 2379, 1304, 523, 33, 2135, 526, 783, 670, 1134, 2292, 239, 730, 1055, 1912, 1631, 674, 851, 702, 813, 2482, 1260, 1781, 1313, 1495, 2039, 342, 623, 2369, 244, 2050, 1997, 743, 1081, 1606, 1389, 67, 2036, 1948, 1462, 2260, 282, 2558, 1500, 2519, 685, 766, 1862, 1952, 729, 872, 1609, 1157, 26, 1073, 1690, 1849, 1855, 2173]. using [2409, 2506, 2345, 1431, 1293, 2249, 2107, 1080, 991, 2571, 760, 1460, 1391, 1979, 286, 421, 2284, 1573, 388, 1657, 1803, 856, 2605, 1797, 2209, 880, 909, 292, 1599, 1588, 2365, 663, 2496, 525, 997, 712, 416, 2077, 1320, 28, 1536, 123, 384, 1214, 1022, 311, 657, 1612, 81, 2268, 1513, 2279, 762, 206, 2071, 733, 1007, 2523, 2589, 816, 982, 715, 2011, 259, 4, 133, 2300, 2287, 561, 110, 1437, 952, 1742, 1098, 1397, 2170, 373, 387, 1028, 988, 1734, 1660, 2118, 2448, 2597, 159, 1280, 1537, 2252, 1311, 2241, 1739, 1461, 1894, 2480, 105, 2414, 50]. usual [1261]. utility [754, 2212, 255].

V [1447]. vaccine [2177]. validated

[2095, 1681, 2027, 1734]. validating [1118]. Validation [1179, 2314, 141, 473, 1485, 2318, 272, 281, 766, 985, 1285, 716]. validations [308]. validity [2318, 510, 485]. Value [897, 436, 1407, 2482, 22, 224, 1646, 649, 1061, 944, 500, 880, 938, 1343, 2206, 1825, 1961]. value-added [1061]. Value-at-Risk [897, 1825, 1961]. valued [1479, 1770, 2271, 1574, 510, 842, 1998, 1971, 38, 982, 950, 2028, 50]. values [2334, 526, 65, 2023, 112, 2490, 1606, 282, 790, 1855, 760, 2281, 1575, 2178, 713, 1046, 2237, 1651, 1665, 25, 2241, 1592]. VAR [1946, 2001, 1510]. variability [2379, 1639, 1285, 715, 214]. Variable [1009, 2151, 1169, 1606, 2368, 1288, 2173, 2107, 2540, 602, 1566, 1005, 1645, 213, 1535, 1625, 1155, 634, 773, 770, 2267, 326, 1528, 571, 2552, 1980, 2186, 2602, 236, 464, 1623, 1685, 1750, 2032, 1081, 191, 2420, 1878, 1263, 1952, 845, 562, 2041, 2192, 488, 1458, 1741, 1853, 1107, 1954, 2301, 1135, 2450, 1083, 2153, 1635, 2152, 521, 1355, 1019, 293, 108, 295, 981, 297, 2386, 2124, 727, 2513, 1406, 1403, 265, 1542, 2521, 2501, 2578, 832, 1131, 1880, 793, 2504, 1613, 1877, 2419]. variable-adjusted [2578]. variables [2090, 1489, 2598, 1698, 1238, 414, 2319, 1150, 524, 1494, 1176, 2024, 2250, 1331, 470, 1895, 16, 541, 121, 1253, 1148, 2010, 1354, 1019, 38, 1046, 2233, 1861, 450, 945, 216, 169, 580, 1653]. Variance [1511, 1213, 682, 2545, 968, 234, 1486, 864, 928, 2601, 802, 1271, 204, 2566, 954, 70, 363, 1196, 1504, 1390, 574, 1538, 409, 2204, 2589, 1633, 2326, 1206, 303, 2075, 2338, 207, 2427, 157, 515]. variances [2141, 220, 555, 694, 2370, 2356]. variant [1202]. variants [1454, 593, 1559]. variates [2295, 1075, 333, 977]. variation [2559, 87, 2519, 2367, 645]. Variational [594, 1021, 808, 1731, 1613, 368, 2029, 1430, 1657, 1803, 320, 1243, 2504]. variogram [2340]. variograms [1177]. various [828]. VARMA [1967, 1424]. VARs [900]. varying [740, 230, 1696, 1490, 1442, 874, 2465, 965, 2184, 1125, 1666, 2352, 348, 2297, 2147, 2438, 2230, 1958, 2386, 1341, 1419, 2276, 1005, 994, 2012, 1725, 2485, 580, 2358, 1576, 1745, 2419, 2404, 2603]. varying-coefficient [965, 1125, 2276, 994, 1725, 580, 2419]. vector [2094, 1325, 433, 479, 687, 236, 2140, 1203, 1299, 1550, 1991, 362, 749, 1108, 667, 1430, 2117, 1170, 1831, 545, 1732, 2604, 1568, 1368, 1827]. vectors [1376, 43, 2233, 2004]. velocimetry [744]. velocity [744]. verification [2507]. version [1624, 1416]. versions [703, 1268, 572]. versus [941, 1487, 183, 1553]. very [67, 829, 818]. via [2471, 1266, 358, 2491, 684, 1252, 606, 897, 2311, 1491, 352, 655, 132, 2481, 2272, 1203, 1586, 1882, 1829, 1060, 1581, 374, 2529, 1315, 405, 1666, 121, 2409, 726, 2092, 2240, 2540, 1316, 2147, 1721, 444, 1777, 2573, 836, 2388, 2285, 878, 602, 1566, 71, 1034, 1896, 920, 682, 1529, 73, 74, 424, 2233, 636, 1559, 1633, 213, 1070, 1542, 1568, 2613, 2004, 2213, 1614, 1522, 1880, 822, 2615, 1724]. Victor [700]. view [2543, 603]. Villasenor [1756]. Vine [2283, 907, 1517, 2596, 1064, 1142, 2144, 2570, 2096]. Vine-copula [1517]. vines [1541, 1997]. virtual [695]. visits [1419]. visualisation [1826]. visualising [68]. visualization [1574, 997, 109, 1190, 382, 75]. Visualizing [2050, 2439]. volatile [871]. volatilities [904, 1942, 870]. Volatility [914, 266, 1478, 631, 1934, 1085, 244, 908, 1950, 1503, 872, 1977, 337, 906, 1027, 249, 909, 911, 881, 2484, 1515, 1516, 2528, 537, 626, 1519, 387]. volume [1115, 1820]. Voronoi [357]. vote [1294]. voting [1228]. vs [970, 925].

wage [241]. wages [878]. Wald

[2584, 1435, 535, 2421, 843]. Wald-based [2421]. Wald-type [1435, 535]. walk [886, 2588, 2072, 12]. warping [1267]. Wavelet [1792, 243, 1835, 239, 896, 641, 751, 1293, 2095, 378, 1031, 1002, 2463, 2466, 2548, 2414]. Wavelet-based [1792, 243, 896, 641, 1293, 378, 1031]. wavelet-transformed [2414]. Wavelets [738, 1495]. waves [2482]. way [693, 1659, 1225, 1903, 1862, 1282, 1240, 1469, 11, 1190, 2488]. weak [745, 248, 642]. weakly [893]. wear [967, 1539]. wear-out [967, 1539]. weather [894, 915]. Web [2405]. Web-based [2405]. weed [1398]. Weibull [1834, 933, 44, 1631, 92, 100, 316, 587, 476, 149, 992, 1914, 780, 654, 438, 1042, 150, 1053, 1651, 1204]. weight [2250, 1160]. Weighted [773, 2235, 423, 779, 432, 2483, 1630, 2599, 803, 1298, 381, 622, 1902, 534, 821, 954, 1315, 817, 1073, 2248, 330, 2051, 786, 2287, 1633, 2349, 408, 353, 361]. weighting [18, 1959, 1074, 1589]. weights [578, 2304, 318]. Weiß [990]. wells [918]. where [568]. Whittle [1968]. Who [2410]. whom [2410]. wide [703, 1268, 552]. Widely [1036]. Wilcoxon [523]. Wild [2547, 2199]. wind [1471, 894]. window [2177, 1641]. wireless [2282]. wise [2243, 2162, 73, 1522]. withdrawal [1254]. within [2141, 1378, 1553, 502, 1893, 833]. within-subject [1893]. without [798, 1239, 2611]. women [705]. working [820, 2087, 966]. world [237, 728, 1508]. worse [1448]. worst [138].

X [1963]. XII [413].

Y. [609]. yield [1936, 2509]. Youden [609, 709, 1681, 2247, 1525]. Yule [388].

Zero [1314, 1101, 1412, 380, 2430, 1242, 430, 1903, 1688, 385, 1099, 1664, 747, 1916, 428, 2244, 120, 147]. Zero-inflated [1314, 1101, 1412, 380, 1242, 430, 1688, 385, 1664, 1916, 120, 147]. zero-inflation [1099]. zero-or-one [747]. zeros [834]. Zhang [86].

References

Ospina:2006:IPI

[1] Raydonal Ospina, Francisco Cribari-Neto, and Klaus L. P. Vasconcellos. Improved point and interval estimation for a beta regression model. Computational Statistics & Data Analysis, 51(2):960–981, November 15, 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947305002549. See

erratum [527].

Said:2008:RSN

[2] Yasmin H. Said, Edward J. Wegman, Walid K. Sharabati, and John T. Rigsby. RETRACTED: Social networks of author–coauthor relationships. Computational Statistics & Data Analysis, 52(4):2177–2184, January 10, 2008. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947307002861. See

retraction [611].

Fridley:2008:RGM

[3] Brooke L. Fridley, Stephen T. Turner, Arlene B. Chapman, Andrei S. Rodin, Eric Boerwinkle, and Kent R. Bailey. Reproducibility of genotypes as measured by the Affymetrix GeneChip^{(}R) 100K Human Mapping Array set. Computational Statistics & Data Analysis, 52(12):5367– 5374, August 15, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308002818. See

comments [80].

Su:2009:CIQ

[4] Steve Su. Confidence intervals for quantiles using generalized lambda distributions. Computational Statistics & Data Analysis, 53(9):3324– 3333, July 1, 2009. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309000437. See

erratum [133].

Chiu:2009:GCM

[5] Sung Nok Chiu and Kwong Ip Liu. Generalized Cramér–von Mises goodness-of-fit tests for multivariate distributions. Computational Statistics & Data Analysis, 53(11):3817– 3834, September 1, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794730900139X. See

corrigendum [59].

Anonymous:2010:EBa

[6] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 54 (1):v–vii, January 1, 2010. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309003363.

Anonymous:2010:Ca

[7] Anonymous. Contents. Computational Statistics & Data Analysis, 54 (1):viii–ix, January 1, 2010. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309003375.

Otranto:2010:IFT

[8] Edoardo Otranto. Identifying financial time series with similar dynamic conditional correlation. Computational Statistics & Data Analysis, 54 (1):1–15, January 1, 2010. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309002710.

Lin:2010:FSS

[9] N. Lin and R. Xi. Fast surrogates of U-statistics. Computational Statistics & Data Analysis, 54(1):16– 24, January 1, 2010. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309002801.

Wang:2010:ESD

[10] Xiao-Feng Wang, Zhaozhi Fan, and Bin Wang. Estimating smooth distribution function in the presence of heteroscedastic measurement errors. Computational Statistics & Data Analysis, 54(1):25–36, January 1, 2010. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309003028.

Todorov:2010:RSO

[11] Valentin Todorov and Peter Filzmoser. Robust statistic for the one-way MANOVA. Computational Statistics & Data Analysis, 54(1):37– 48, January 1, 2010. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309003090.

Vihola:2010:GGM

[12] Matti Vihola. Grapham: Graphical models with adaptive random walk Metropolis algorithms. Computational Statistics & Data Analysis, 54(1):49– 54, January 1, 2010. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309003193.

Zou:2010:CIE

[13] G. Y. Zou. Confidence interval estimation under inverse sampling. Computational Statistics & Data Analysis, 54 (1):55–64, January 1, 2010. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309002564.

Zang:2010:APM

[14] Yong Zang, Wing Kam Fung, and Gang Zheng. Asymptotic powers for matched trend tests and robust matched trend tests in case-control genetic association studies. Computational Statistics & Data Analysis, 54 (1):65–77, January 1, 2010. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309002576.

Gonzalez:2010:CIM

[15] Juan R. Gonzalez, Edsel A. Peña, and Pedro Delicado. Confidence intervals for median survival time with recurrent event data. Computational Statistics & Data Analysis, 54(1):78– 89, January 1, 2010. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309002588.

Kojadinovic:2010:HCC

[16] Ivan Kojadinovic. Hierarchical clustering of continuous variables based on the empirical copula process and permutation linkages. Computational Statistics & Data Analysis, 54(1):90– 108, January 1, 2010. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794730900259X.

Hutson:2010:MRN

[17] Alan D. Hutson. A multi-rater nonparametric test of agreement and corresponding agreement plot. Computational Statistics & Data Analysis, 54 (1):109–119, January 1, 2010. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309002606.

Bouguila:2010:MBD

[18] Nizar Bouguila. On multivariate binary data clustering and feature weighting. Computational Statistics & Data Analysis, 54(1):120–134, January 1, 2010. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309002618.

Balakrishnan:2010:NTE

[19] N. Balakrishnan and Xingqiu Zhao. A nonparametric test for the equality of counting processes with panel count data. Computational Statistics & Data Analysis, 54(1):135–142, January 1, 2010. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794730900262X.

Tommasi:2010:BOD

[20] C. Tommasi and J. López-Fidalgo. Bayesian optimum designs for discriminating between models with any distribution. Computational Statistics & Data Analysis, 54(1):143–150, January 1, 2010. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309002655.

Fernandez:2010:TST

[21] Arturo J. Fernández. Two-sided tolerance intervals in the exponential case: Corrigenda and generalizations. Computational Statistics & Data Analysis, 54(1):151–162, January 1, 2010. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309002679.

Gervini:2010:RFSa

[22] Daniel Gervini. RETRACTED: the functional singular value decomposition for bivariate stochastic processes. Computational Statistics & Data Analysis, 54(1):163–172, January 1, 2010. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309002692.

Lee:2010:ACI

[23] Woojoo Lee, Jian Qing Shi, and Youngjo Lee. Approximate conditional inference in mixed-effects models with binary data. Computational Statistics & Data Analysis, 54(1):173– 184, January 1, 2010. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309002709.

Xu:2010:RAB

[24] Ancha Xu and Yincai Tang. Reference analysis for Birnbaum–Saunders distribution. Computational Statistics & Data Analysis, 54(1):185–192, January 1, 2010. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309002734.

Yang:2010:JAM

[25] Ying Yang and Jian Kang. Joint analysis of mixed Poisson and continuous longitudinal data with nonignorable missing values. Computational Statistics & Data Analysis, 54 (1):193–207, January 1, 2010. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947309002771.

Lee:2010:LND

[26] Keunbaik Lee and Donald Mercante. Longitudinal nominal data analysis using marginalized models. Computational Statistics & Data Analysis, 54 (1):208–218, January 1, 2010. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947309002795.

Chuang:2010:UDG

[27] S. C. Chuang and Y. C. Hung. Uniform design over general input domains with applications to target region estimation in computer experiments. Computational Statistics & Data Analysis, 54 (1):219–232, January 1, 2010. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016794730900293X.

Polansky:2010:OIU

[28] Alan M. Polansky. Ordered inference using observed confidence levels. Computational Statistics & Data Analysis, 54(1):233–244, January 1, 2010. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947309003041.

Anonymous:2010:EBe

[29] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 54 (2):iii–v, February 1, 2010. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309003880.

Anonymous:2010:Ce

[30] Anonymous. Contents. Computational Statistics & Data Analysis, 54 (2):vi–vii, February 1, 2010. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309003892.

Bouezmarni:2010:NDE

[31] Taoufik Bouezmarni and Jeroen V. K. Rombouts. Nonparametric density estimation for positive time series. Computational Statistics & Data Analysis, 54(2):245–261, February 1, 2010. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309003077.

dePeretti:2010:GMI

[32] Christian de Peretti and Carole Siani. Graphical methods for investigating the finite-sample properties of confidence regions. Computational Statistics & Data Analysis, 54(2):262–271, February 1, 2010. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309003284.

Das:2010:EMC

[33] Sonjoy Das, James C. Spall, and Roger Ghanem. Efficient Monte Carlo computation of Fisher information matrix using prior information. Computational Statistics & Data Analysis, 54(2):272– 289, February 1, 2010. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309003442.

Karatzoglou:2010:KBM

[34] Alexandros Karatzoglou and Ingo Feinerer. Kernel-based machine learning for fast text mining in R. Computational Statistics & Data Analysis, 54(2):290–297, February 1, 2010. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0167947309003533.

Cheng:2010:BAS

[35] Dunlei Cheng, Adam J. Branscum, and James D. Stamey. A Bayesian approach to sample size determination for studies designed to evaluate continuous medical tests. Computational Statistics & Data Analysis, 54(2):298– 307, February 1, 2010. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309003545.

Boudaoud:2010:CSM

[36] S. Boudaoud, H. Rix, and O. Meste. Core shape modelling of a set of curves. Computational Statistics & Data Analysis, 54(2):308–325, February 1, 2010. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309002746.

Xiao:2010:EBS

[37] Qingchu Xiao, Zaiming Liu, N. Balakrishnan, and Xuewen Lu. Estimation of the Birnbaum–Saunders regression model with current status data. Computational Statistics & Data Analysis, 54(2):326–332, February 1, 2010. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309003053.

Neto:2010:CLR

[38] Eufrásio de A. Lima Neto and Francisco de A. T. de Carvalho. Constrained linear regression models for symbolic interval-valued variables. Computational Statistics & Data Analysis, 54(2):333–347, February 1, 2010. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309003065.

Simas:2010:IEG

[39] Alexandre B. Simas, Wagner Barreto-Souza, and Andréa V. Rocha. Improved estimators for a general class of beta regression models. Computational Statistics & Data Analysis, 54(2):348– 366, February 1, 2010. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309003107.

Meintanis:2010:IPB

[40] Simos G. Meintanis. Inference procedures for the Birnbaum–Saunders distribution and its generalizations. Computational Statistics & Data Analysis, 54(2):367–373, February 1, 2010. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0167947309003132.

Frahm:2010:GTE

[41] Gabriel Frahm and Uwe Jaekel. A generalization of Tyler’s Mestimators to the case of incomplete data. Computational Statistics & Data Analysis, 54(2):374–393, February 1, 2010. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309003144.

Tian:2010:FPN

[42] Guo-Liang Tian, Man-Lai Tang, Kam Chuen Yuen, and Kai Wang Ng. Further properties and new applications of the nested Dirichlet distribution. Computational Statistics & Data Analysis, 54(2):394–405, February 1, 2010. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309003156.

Cheon:2010:MCP

[43] Sooyoung Cheon and Jaehee Kim. Multiple change-point detection of multivariate mean vectors with the Bayesian approach. Computational Statistics & Data Analysis, 54(2):406– 415, February 1, 2010. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309003181.

Farcomeni:2010:TCW

[44] Alessio Farcomeni and Alessandra Nardi. A two-component Weibull mixture to model early and late mortality in a Bayesian framework. Computational Statistics & Data Analysis, 54(2):416–428, February 1, 2010. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309003235.

Boj:2010:DBL

[45] Eva Boj, Pedro Delicado, and Josep Fortiana. Distance-based local linear regression for functional predictors. Computational Statistics & Data Analysis, 54(2):429–437, February 1, 2010. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309003260.

Xue:2010:GDT

[46] Jing-Hao Xue and D. Michael Titterington. On the generative– discriminative tradeoff approach: Interpretation, asymptotic efficiency and classification performance. Computational Statistics & Data Analysis, 54(2):438–451, February 1, 2010. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0167947309003272.

Crujeiras:2010:LSE

[47] Rosa M. Crujeiras and Ingrid Van Keilegom. Least squares estimation of nonlinear spatial trends. Computational Statistics & Data Analysis, 54(2):452– 465, February 1, 2010. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309003302.

Tian:2010:CIE

[48] Lili Tian. Confidence interval estimation of partial area under curve based on combined biomarkers. Computational Statistics & Data Analysis, 54(2):466–472, February 1, 2010. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0167947309003429.

Hirukawa:2010:NMB

[49] Masayuki Hirukawa. Nonparametric multiplicative bias correction for kernel-type density estimation on the unit interval. Computational Statistics & Data Analysis, 54(2):473–495, February 1, 2010. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309003430. See

corrigendum [1870].

Zhu:2010:DCI

[50] Fukang Zhu and Dehui Wang. Diagnostic checking integer-valued ARCH(p) models using conditional residual autocorrelations. Computational Statistics & Data Analysis, 54(2):496–508, February 1, 2010. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309003454.

Chan:2010:NLI

[51] Ngai-Hang Chan, Thomas C. M. Lee, and Liang Peng. On nonparametric local inference for density estimation. Computational Statistics & Data Analysis, 54(2):509–515, February 1, 2010. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794730900351X.

Huang:2010:EOD

[52] Mong-Na Lo Huang, Chuan-Pin Lee, Ray-Bing Chen, and Thomas Klein. Exact D-optimal designs for a secondorder response surface model on a circle with qualitative factors. Computational Statistics & Data Analysis, 54(2):516– 530, February 1, 2010. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309003521.

Vexler:2010:ELR

[53] Albert Vexler and Gregory Gurevich. Empirical likelihood ratios applied to goodness-of-fit tests based on sample entropy. Computational Statistics & Data Analysis, 54(2):531–545, February 1, 2010. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309003557.

Bianco:2010:EML

[54] Ana Bianco, Graciela Boente, Wenceslao González-Manteiga, and Ana Pérez-González. Estimation of the marginal location under a partially linear model with missing responses. Computational Statistics & Data Analysis, 54(2):546–564, February 1, 2010. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309003582.

Chang:2010:ABA

[55] Yu-Ling Chang, Fei Zou, and Fred A. Wright. An approximate Bayesian approach for quantitative trait loci estimation. Computational Statistics & Data Analysis, 54(2):565–574, February 1, 2010. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309003594.

Sartori:2010:ASG

[56] N. Sartori, T. A. Severini, and E. Marras. An alternative specification of generalized linear mixed models. Computational Statistics & Data Analysis, 54(2):575–584, February 1, 2010. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309003600.

Molenberghs:2010:SPM

[57] Geert Molenberghs and Michael G. Kenward. Semi-parametric marginal models for hierarchical data and their corresponding full models. Computational Statistics & Data Analysis, 54(2):585–597, February 1, 2010. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0167947309003703.

DeGregorio:2010:CDO

[58] Alessandro De Gregorio and Stefano Maria Iacus. Clustering of discretely observed diffusion processes. Computational Statistics & Data Analysis, 54(2):598–606, February 1, 2010. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309003752.

Chiu:2010:CGC

[59] Sung Nok Chiu and Kwong Ip Liu. Corrigendum to “Generalized Cramér–von Mises goodness-of-fit tests for multivariate distributions” [comput. statist. data anal. 53 (2009) 3817–3834]. Computational Statistics & Data Analysis, 54(2):607, February 1, 2010. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309003296. See

[5].

Anonymous:2010:EBf

[60] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 54(3):iii–v, March 1, 2010. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309004198.

Anonymous:2010:Cf

[61] Anonymous. Contents. Computational Statistics & Data Analysis, 54 (3):vi–vii, March 1, 2010. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309004204.

Gatu:2010:SSI

[62] Cristian Gatu and B. D. McCullough. Second special issue on statistical algorithms and software. Computational Statistics & Data Analysis, 54 (3):609–610, March 1, 2010. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947309004125.

Bulla:2010:HRP

[63] Jan Bulla, Ingo Bulla, and Oleg Nenadić. hsmm — an R package for analyzing hidden semi-Markov models. Computational Statistics & Data Analysis, 54(3):611–619, March 1, 2010. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016794730800426X.

Davidov:2010:NIA

[64] Ori Davidov and George Iliopoulos. A note on an iterative algorithm for nonparametric estimation in biased sampling models. Computational Statistics & Data Analysis, 54 (3):620–624, March 1, 2010. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947308004052.

Escanciano:2010:ACV

[65] Juan Carlos Escanciano and David T. Jacho-Chávez. Approximating the critical values of Cramér–von Mises tests in general parametric conditional specifications. Computational Statistics & Data Analysis, 54(3):625– 636, March 1, 2010. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308002867.

Gallegos:2010:UCO

[66] María Teresa Gallegos and Gunter Ritter. Using combinatorial optimization in model-based trimmed clustering with cardinality constraints. Computational Statistics & Data Analysis, 54 (3):637–654, March 1, 2010. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309003119.

Harrington:2010:FAS

[67] Justin Harrington and Matias Salibián-Barrera. Finding approximate solutions to combinatorial problems with very large data sets using BIRCH. Computational Statistics & Data Analysis, 54(3):655– 667, March 1, 2010. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308003873.

Hanea:2010:MVO

[68] A. M. Hanea, D. Kurowicka, R. M. Cooke, and D. A. Ababei. Mining and visualising ordinal data with nonparametric continuous BBNs. Computational Statistics & Data Analysis, 54 (3):668–687, March 1, 2010. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308004568.

Hu:2010:DEM

[69] Bo Hu and Kam-Wah Tsui. Distributed evolutionary Monte Carlo for Bayesian computing. Computational Statistics & Data Analysis, 54 (3):688–697, March 1, 2010. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794730800491X.

Iacobucci:2010:VSP

[70] Alessandra Iacobucci, Jean-Michel Marin, and Christian Robert. On variance stabilisation in population Monte Carlo by double Rao– Blackwellisation. Computational Statistics & Data Analysis, 54(3):698– 710, March 1, 2010. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308004489.

McNicholas:2010:SPI

[71] P. D. McNicholas, T. B. Murphy, A. F. McDaid, and D. Frost. Serial and parallel implementations of model-based clustering via parsimonious Gaussian mixture models. Computational Statistics & Data Analysis, 54(3):711–723, March 1, 2010. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309000632.

Poitevineau:2010:IBP

[72] Jacques Poitevineau and Bruno Lecoutre. Implementing Bayesian predictive procedures: the K-prime and K-square distributions. Computational Statistics & Data Analysis, 54 (3):724–731, March 1, 2010. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308005410.

Reddy:2010:LMM

[73] Chandan K. Reddy and Bala Rajaratnam. Learning mixture models via component-wise parameter smoothing. Computational Statistics & Data Analysis, 54(3):732– 749, March 1, 2010. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309001625.

Saadaoui:2010:AEA

[74] Foued Saâdaoui. Acceleration of the EM algorithm via extrapolation methods: Review, comparison and new methods. Computational Statistics & Data Analysis, 54(3):750– 766, March 1, 2010. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308005422.

Wu:2010:GGE

[75] Han-Ming Wu, Yin-Jing Tien, and Chun houh Chen. GAP: a graphical environment for matrix visualization and cluster analysis. Computational Statistics & Data Analysis, 54 (3):767–778, March 1, 2010. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308004349.

Yang:2010:UGQ

[76] Guijun Yang, Zhigang Wang, and Wei Deng. Unbiased generalized quasiregression. Computational Statistics & Data Analysis, 54(3):779– 789, March 1, 2010. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016794730800251X.

Yucel:2010:INN

[77] Recai M. Yucel and Hakan Demirtas. Impact of non-normal random effects on inference by multiple imputation: a simulation assessment. Computational Statistics & Data Analysis, 54 (3):790–801, March 1, 2010. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947309000346.

Anonymous:2010:EBg

[78] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 54(4):iii–v, April 1, 2010. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947309004496.

Anonymous:2010:Cg

[79] Anonymous. Contents. Computational Statistics & Data Analysis, 54 (4):vi–viii, April 1, 2010. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947309004502.

Vonthein:2010:UTR

[80] Reinhard Vonthein and Andreas Ziegler. On the use of the terms repeatability and reproducibility regarding “Reproducibility of genotypes as measured by the Affymetrix GeneChip(R) 100K Human Mapping Array Set” by Fridley and colleagues (2008) Comput. Stat. Data Anal. 52:5367–74. Computational Statistics & Data Analysis, 54(4): 803, April 1, 2010. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309004447. See

[3].

Rohmel:2010:CTI

[81] Joachim Röhmel. Comparing two independent samples using the risk difference under inverse sampling. Computational Statistics & Data Analysis, 54 (4):804–805, April 1, 2010. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309003806.

Hines:2010:ICM

[82] R. J. O’Hara Hines and W. G. S. Hines. Indices for covariance misspecification in longitudinal data analysis with no missing responses and with MAR drop-outs. Computational Statistics & Data Analysis, 54(4):806– 815, April 1, 2010. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309004046.

Argiento:2010:BDE

[83] Raffaele Argiento, Alessandra Guglielmi, and Antonio Pievatolo. Bayesian density estimation and model selection using nonparametric hierarchical mixtures. Computational Statistics & Data Analysis, 54(4):816– 832, April 1, 2010. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947309004058.

Saldana-Zepeda:2010:GFT

[84] Dayna P. Saldaña-Zepeda, Humberto Vaquera-Huerta, and Barry C. Arnold. A goodness of fit test for the Pareto distribution in the presence of type II censoring, based on the cumulative hazard function. Computational Statistics & Data Analysis, 54(4):833– 842, April 1, 2010. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947309004071.

Schyns:2010:RSO

[85] M. Schyns, G. Haesbroeck, and F. Critchley. RelaxMCD: Smooth optimisation for the minimum covariance determinant estimator. Computational Statistics & Data Analysis, 54 (4):843–857, April 1, 2010. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947309004083.

Duchesne:2010:CDQ

[86] Pierre Duchesne and Pierre Lafaye De Micheaux. Computing the distribution of quadratic forms: Further comparisons between the Liu–Tang–Zhang approximation and exact methods. Computational Statistics & Data Analysis, 54(4):858–862, April 1, 2010. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309004381.

Human:2010:STC

[87] S. W. Human, S. Chakraborti, and C. F. Smit. Shewhart-type control charts for variation in phase I data analysis. Computational Statistics & Data Analysis, 54(4):863– 874, April 1, 2010. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309003569.

Adewale:2010:RDG

[88] Adeniyi J. Adewale and Xiaojian Xu. Robust designs for generalized linear models with possible overdispersion and misspecified link functions. Computational Statistics & Data Analysis, 54(4):875–890, April 1, 2010. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309003624.

Cao:2010:PSE

[89] Yanrong Cao, Haiqun Lin, Tracy Z. Wu, and Yan Yu. Penalized spline estimation for functional coefficient regression models. Computational Statistics & Data Analysis, 54(4):891– 905, April 1, 2010. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947309003661.

Ye:2010:ICM

[90] Ren-Dao Ye, Tie-Feng Ma, and Song-Gui Wang. Inferences on the common mean of several inverse Gaussian populations. Computational Statistics & Data Analysis, 54(4):906– 915, April 1, 2010. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947309003697.

Alston:2010:AED

[91] C. L. Alston and K. L. Mengersen. Allowing for the effect of data binning in a Bayesian normal mixture model. Computational Statistics & Data Analysis, 54(4):916–923, April 1, 2010. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947309003739.

Gupta:2010:ETP

[92] Ramesh C. Gupta, Sergey Lvin, and Cheng Peng. Estimating turning points of the failure rate of the extended Weibull distribution. Computational Statistics & Data Analysis, 54 (4):924–934, April 1, 2010. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947309003740.

Silva:2010:NDD

[93] Rodrigo B. Silva, Wagner Barreto-Souza, and Gauss M. Cordeiro. A new distribution with decreasing, increasing and upside-down bathtub failure rate. Computational Statistics & Data Analysis, 54(4):935– 944, April 1, 2010. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309003764.

Pescim:2010:BGH

[94] Rodrigo R. Pescim, Clarice G. B. Demétrio, Gauss M. Cordeiro, Edwin M. M. Ortega, and Mariana R. Urbano. The beta generalized halfnormal distribution. Computational Statistics & Data Analysis, 54(4):945– 957, April 1, 2010. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309003776.

Gel:2010:TFL

[95] Yulia R. Gel. Test of fit for a Laplace distribution against heavier tailed alternatives. Computational Statistics & Data Analysis, 54(4):958– 965, April 1, 2010. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309003788.

Lanius:2010:ROS

[96] Vivian Lanius and Ursula Gather. Robust online signal extraction from multivariate time series. Computational Statistics & Data Analysis, 54 (4):966–975, April 1, 2010. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794730900379X.

Zhu:2010:SED

[97] Li-Ping Zhu, Zhou Yu, and Li-Xing Zhu. A sparse eigen-decomposition estimation in semiparametric regression. Computational Statistics & Data Analysis, 54(4):976– 986, April 1, 2010. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947309003818.

Ramalho:2010:NHR

[98] Esmeralda A. Ramalho and Joaquim J. S. Ramalho. Is neglected heterogeneity really an issue in binary and fractional regression models? a simulation exercise for logit, probit and loglog models. Computational Statistics & Data Analysis, 54(4):987– 1001, April 1, 2010. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016794730900382X.

Vanegas:2010:ADP

[99] Luis Hernando Vanegas and Francisco José A. Cysneiros. Assessment of diagnostic procedures in symmetrical nonlinear regression models. Computational Statistics & Data Analysis, 54 (4):1002–1016, April 1, 2010. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947309003831.

Hashimoto:2010:LEW

[100] Elizabeth M. Hashimoto, Edwin M. M. Ortega, Vicente G. Cancho, and Gauss M. Cordeiro. The log-exponentiated Weibull regression model for interval-censored data. Computational Statistics & Data Analysis, 54(4):1017–1035, April 1, 2010. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309003958.

Fu:2010:RRA

[101] Liya Fu, You-Gan Wang, and Zhidong Bai. Rank regression for analysis of clustered data: a natural induced smoothing approach. Computational Statistics & Data Analysis, 54 (4):1036–1050, April 1, 2010. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794730900396X.

Serban:2010:NRE

[102] Nicoleta Serban. Noise reduction for enhanced component identification in multi-dimensional biomolecular NMR studies. Computational Statistics & Data Analysis, 54(4):1051– 1065, April 1, 2010. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309003971.

Gao:2010:UGI

[103] Wei Gao, Ning-Zhong Shi, Man-Lai Tang, Lianyan Fu, and Guoliang Tian. Unified generalized iterative scaling and its applications. Computational Statistics & Data Analysis, 54(4):1066– 1078, April 1, 2010. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947309003983.

Su:2010:ELB

[104] Haiyan Su and Hua Liang. An empirical likelihood-based method for comparison of treatment effects — test of equality of coefficients in linear models. Computational Statistics & Data Analysis, 54(4):1079– 1088, April 1, 2010. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947309003995.

Zhao:2010:DRU

[105] Junlong Zhao and Xiuli Zhao. Dimension reduction using the generalized gradient direction. Computational Statistics & Data Analysis, 54 (4):1089–1102, April 1, 2010. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947309004009.

Swift:2010:SSC

[106] Michael Bruce Swift. A simulation study comparing methods for calculating confidence intervals for directly standardized rates. Computational Statistics & Data Analysis, 54 (4):1103–1108, April 1, 2010. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947309004022.

Chen:2010:MIA

[107] Ling Chen and Jianguo Sun. A multiple imputation approach to the analysis of interval-censored failure time data with the additive hazards model. Computational Statistics & Data Analysis, 54 (4):1109–1116, April 1, 2010. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309004034.

Ogura:2010:VSM

[108] Toru Ogura. A variable selection method in principal canonical correlation analysis. Computational Statistics & Data Analysis, 54(4):1117– 1123, April 1, 2010. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309003247.

Tseng:2010:QMS

[109] George C. Tseng. Quantile map: Simultaneous visualization of patterns in many distributions with application to tandem mass spectrometry. Computational Statistics & Data Analysis, 54 (4):1124–1137, April 1, 2010. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794730900317X.

Thompson:2010:BNQ

[110] Paul Thompson, Yuzhi Cai, Rana Moyeed, Dominic Reeve, and Julian Stander. Bayesian nonparametric quantile regression using splines. Computational Statistics & Data Analysis, 54(4):1138–1150, April 1, 2010. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794730900320X.

Longford:2010:SAE

[111] Nicholas T. Longford. Small area estimation with spatial similarity. Computational Statistics & Data Analysis, 54 (4):1151–1166, April 1, 2010. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947309003211.

Garcia:2010:RSG

[112] Damien Garcia. Robust smoothing of gridded data in one and higher dimensions with missing values. Computational Statistics & Data Analysis, 54 (4):1167–1178, April 1, 2010. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947309003491. See

corrigendum [790].

Wang:2010:MEM

[113] Dong Wang. Modeling epigenetic modifications under multiple treatment conditions. Computational Statistics & Data Analysis, 54(4):1179– 1189, April 1, 2010. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016794730900365X.

Hargreaves:2010:PTF

[114] Bruce R. Hargreaves and Thomas P. McWilliams. Polynomial trendline function flaws in Microsoft Excel. Computational Statistics & Data Analysis, 54(4):1190–1196, April 1, 2010. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309004010.

Anonymous:2010:EBh

[115] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 54(5):iii–v, May 1, 2010. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310000137.

Anonymous:2010:Ch

[116] Anonymous. Contents. Computational Statistics & Data Analysis, 54 (5):vi–vii, May 1, 2010. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310000149.

Zhang:2010:BAC

[117] Junni L. Zhang and Wolfgang K. Härdle. The Bayesian additive classification tree applied to credit risk modelling. Computational Statistics & Data Analysis, 54(5):1197– 1205, May 1, 2010. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309004356.

Wagner:2010:BER

[118] Helga Wagner and Regina Tüchler. Bayesian estimation of random effects models for multivariate responses of mixed data. Computational Statistics & Data Analysis, 54(5):1206– 1218, May 1, 2010. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309004575.

Sangalli:2010:MAC

[119] Laura M. Sangalli, Piercesare Secchi, Simone Vantini, and Valeria Vitelli. k-mean alignment for curve clustering. Computational Statistics & Data Analysis, 54(5):1219– 1233, May 1, 2010. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309004605.

Yang:2010:STO

[120] Zhao Yang, James W. Hardin, and Cheryl L. Addy. Score tests for overdispersion in zero-inflated Poisson mixed models. Computational Statistics & Data Analysis, 54(5): 1234–1246, May 1, 2010. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309004137.

Lee:2010:CCV

[121] Namgil Lee and Jong-Min Kim. Conversion of categorical variables into numerical variables via Bayesian network classifiers for binary classifications. Computational Statistics & Data Analysis, 54(5):1247– 1265, May 1, 2010. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794730900406X.

Zeller:2010:IAS

[122] Camila B. Zeller, Filidor V. Labra, Victor H. Lachos, and N. Balakrishnan. Influence analyses of skew-normal/independent linear mixed models. Computational Statistics & Data Analysis, 54(5):1266– 1280, May 1, 2010. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309004113.

Rahardja:2010:CSR

[123] Dewi Rahardja and Dean M. Young. Credible sets for risk ratios in overreported two-sample binomial data using the double-sampling scheme. Computational Statistics & Data Analysis, 54(5):1281–1287, May 1, 2010. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0167947309004149.

Hutter:2010:MSL

[124] Marcus Hutter and Minh-Ngoc Tran. Model selection with the loss rank principle. Computational Statistics & Data Analysis, 54(5):1288– 1306, May 1, 2010. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309004265.

Lemonte:2010:ILI

[125] Artur J. Lemonte, Silvia L. P. Ferrari, and Francisco Cribari-Neto. Improved likelihood inference in Birnbaum– Saunders regressions. Computational Statistics & Data Analysis, 54(5): 1307–1316, May 1, 2010. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947309004289.

Bingham:2010:FSI

[126] Melissa A. Bingham, Daniel J. Nordman, and Stephen B. Vardeman. Finite-sample investigation of likelihood and Bayes inference for the symmetric von Mises–Fisher distribution. Computational Statistics & Data Analysis, 54(5):1317– 1327, May 1, 2010. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947309004332.

Wang:2010:EBM

[127] Wan-Lun Wang and Tsai-Hung Fan. ECM-based maximum likelihood inference for multivariate linear mixed models with autoregressive errors. Computational Statistics & Data Analysis, 54 (5):1328–1341, May 1, 2010. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947309004344.

Pryseley:2010:UEM

[128] Assam Pryseley, Abel Tilahun, Ariel Alonso, and Geert Molenberghs. Using earlier measures in a longitudinal sequence as a potential surrogate for a later one. Computational Statistics & Data Analysis, 54(5): 1342–1354, May 1, 2010. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016794730900437X.

Verzelen:2010:DDN

[129] Nicolas Verzelen. Data-driven neighborhood selection of a Gaussian field. Computational Statistics & Data Analysis, 54(5):1355–1371, May 1, 2010. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309004393.

Wu:2010:OSM

[130] Shu-Fei Wu, Ying-Po Lin, and Yuh-Ru Yu. One-stage multiple comparisons with the control for exponential location parameters under heteroscedasticity. Computational Statistics & Data Analysis, 54(5):1372– 1380, May 1, 2010. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794730900440X.

Eshima:2010:ECD

[131] Nobuoki Eshima and Minoru Tabata. Entropy coefficient of determination for generalized linear models. Computational Statistics & Data Analysis, 54 (5):1381–1389, May 1, 2010. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309004411.

Fraiman:2010:PRP

[132] Ricardo Fraiman, Ana Justel, and Marcela Svarc. Pattern recognition via projection-based kNN rules. Computational Statistics & Data Analysis, 54 (5):1390–1403, May 1, 2010. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947309004617.

Su:2010:ECI

[133] Steve Su. Erratum to: “Confidence intervals for quantiles using generalized lambda distributions” [comput. statist. data anal. 53 (2009) 3324–3333]. Computational Statistics & Data Analysis, 54(5):1404, May 1, 2010. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947309004241. See

[4].

Pleydell:2010:MGH

[134] David R. J. Pleydell and Stéphane Chrétien. Mixtures of GAMs for habitat suitability analysis with overdispersed presence/absence data. Computational Statistics & Data Analysis, 54 (5):1405–1418, May 1, 2010. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947309004277.

Anonymous:2010:Ci

[135] Anonymous. Contents. Computational Statistics & Data Analysis, 54 (6):iii–iv, June 1, 2010. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947310000654.

Anonymous:2010:EBi

[136] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 54(6):v–vii, June 1, 2010. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310000642.

Yu:2010:SMC

[137] Yaming Yu. Strict monotonicity and convergence rate of Titterington’s algorithm for computing Doptimal designs. Computational Statistics & Data Analysis, 54(6):1419– 1425, June 1, 2010. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310000411.

Vermeulen:2010:OMI

[138] Bart Vermeulen, Peter Goos, and Martina Vandebroek. Obtaining more information from conjoint experiments by best–worst choices. Computational Statistics & Data Analysis, 54 (6):1426–1433, June 1, 2010. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310000034.

Glen:2010:AEO

[139] Andrew G. Glen. Accurate estimation with one order statistic. Computational Statistics & Data Analysis, 54 (6):1434–1441, June 1, 2010. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310000277.

Rufibach:2010:ASA

[140] Kaspar Rufibach. An active set algorithm to estimate parameters in generalized linear models with ordered predictors. Computational Statistics & Data Analysis, 54(6):1442– 1456, June 1, 2010. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310000290.

Abrahantes:2010:SMS

[141] José Cortiñas Abrahantes and Tomasz Burzykowski. Simplified modeling strategies for surrogate validation with multivariate failure-time data. Computational Statistics & Data Analysis, 54 (6):1457–1466, June 1, 2010. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310000319.

Xu:2010:ELA

[142] Linzhi Xu and Jiajia Zhang. An EM-like algorithm for the semiparametric accelerated failure time gamma frailty model. Computational Statistics & Data Analysis, 54(6):1467– 1474, June 1, 2010. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310000368.

Jamshidian:2010:SCB

[143] Mortaza Jamshidian, Wei Liu, and Frank Bretz. Simultaneous confidence bands for all contrasts of three or more simple linear regression models over an interval. Computational Statistics & Data Analysis, 54(6): 1475–1483, June 1, 2010. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794731000037X.

DeBrabanter:2010:OFS

[144] K. De Brabanter, J. De Brabanter, J. A. K. Suykens, and B. De Moor. Optimized fixed-size kernel models for large data sets. Computational Statistics & Data Analysis, 54 (6):1484–1504, June 1, 2010. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310000393.

Lin:2010:OTD

[145] D. K. J. Lin, C. Sharpe, and P. Winker. Optimized U-type designs on flexible regions. Computational Statistics & Data Analysis, 54(6):1505– 1515, June 1, 2010. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310000472.

Piccarreta:2010:BTD

[146] Raffaella Piccarreta. Binary trees for dissimilarity data. Computational Statistics & Data Analysis, 54(6): 1516–1524, June 1, 2010. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309004630.

Yang:2010:UCM

[147] Yan Yang and Douglas Simpson. Unified computational methods for regression analysis of zero-inflated and bound-inflated data. Computational Statistics & Data Analysis, 54(6): 1525–1534, June 1, 2010. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947309004642.

DeBock:2010:ECB

[148] Koen W. De Bock, Kristof Coussement, and Dirk Van den Poel. Ensemble classification based on generalized additive models. Computational Statistics & Data Analysis, 54 (6):1535–1546, June 1, 2010. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947309004654.

Kundu:2010:BIP

[149] Debasis Kundu and Hatem Howlader. Bayesian inference and prediction of the inverse Weibull distribution for Type-II censored data. Computational Statistics & Data Analysis, 54 (6):1547–1558, June 1, 2010. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947310000046.

Nandi:2010:EAE

[150] Swagata Nandi and Isha Dewan. An EM algorithm for estimating the parameters of bivariate Weibull distribution under random censoring. Computational Statistics & Data Analysis, 54 (6):1559–1569, June 1, 2010. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947310000058.

Beh:2010:AAI

[151] Eric J. Beh. The aggregate association index. Computational Statistics & Data Analysis, 54(6):1570– 1580, June 1, 2010. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310000071.

Chen:2010:PEG

[152] D. G. Chen and Y. L. Lio. Parameter estimations for generalized exponential distribution under progressive type-I interval censoring. Computational Statistics & Data Analysis, 54 (6):1581–1591, June 1, 2010. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310000083.

Dannemann:2010:TTC

[153] Jörn Dannemann and Hajo Holzmann. Testing for two components in a switching regression model. Computational Statistics & Data Analysis, 54(6): 1592–1604, June 1, 2010. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310000241.

Li:2010:CAS

[154] Peng Li, Shengli Zhao, and Runchu Zhang. A cluster analysis selection strategy for supersaturated designs. Computational Statistics & Data Analysis, 54(6):1605– 1612, June 1, 2010. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947310000253.

Hsieh:2010:EKT

[155] Jin-Jian Hsieh. Estimation of Kendall’s tau from censored data. Computational Statistics & Data Analysis, 54 (6):1613–1621, June 1, 2010. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947310000307.

Ramos:2010:BSM

[156] Patrícia de Siqueira Ramos and Daniel Furtado Ferreira. A Bayesian solution to the multivariate Behrens– Fisher problem. Computational Statistics & Data Analysis, 54(6):1622– 1633, June 1, 2010. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947310000344.

Zhou:2010:CCB

[157] Qin Zhou, Yunzhao Luo, and Zhaojun Wang. A control chart based on likelihood ratio test for detecting patterned mean and variance shifts. Computational Statistics & Data Analysis, 54 (6):1634–1645, June 1, 2010. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947310000356.

Chen:2010:IHC

[158] D. G. Chen. Incorporating historical control information into quantal bioassay with Bayesian approach. Computational Statistics & Data Analysis, 54 (6):1646–1656, June 1, 2010. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310000381.

Wheeler:2010:ALM

[159] David C. Wheeler, DeMarc A. Hickson, and Lance A. Waller. Assessing local model adequacy in Bayesian hierarchical models using the partitioned deviance information criterion. Computational Statistics & Data Analysis, 54 (6):1657–1671, June 1, 2010. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794731000040X.

Lee:2010:DBT

[160] Paul H. Lee and Philip L. H. Yu. Distance-based tree models for ranking data. Computational Statistics & Data Analysis, 54(6):1672– 1682, June 1, 2010. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310000423.

Wiens:2010:CSR

[161] Douglas P. Wiens and Eden K. H. Wu. A comparative study of robust designs for M-estimated regression models. Computational Statistics & Data Analysis, 54(6):1683– 1695, June 1, 2010. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310000484.

Zeileis:2010:TMD

[162] Achim Zeileis, Ajay Shah, and Ila Patnaik. Testing, monitoring, and dating structural changes in exchange rate regimes. Computational Statistics & Data Analysis, 54(6):1696– 1706, June 1, 2010. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947309004435.

Anonymous:2010:EBj

[163] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 54(7):iii–v, July 1, 2010. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016794731000099X.

Anonymous:2010:Cj

[164] Anonymous. Contents. Computational Statistics & Data Analysis, 54 (7):vi–vii, July 1, 2010. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947310001003.

Holmes:2010:FKC

[165] Michael P. Holmes, Alexander G. Gray, and Charles Lee Isbell, Jr. Fast kernel conditional density estimation: a dualtree Monte Carlo approach. Computational Statistics & Data Analysis, 54 (7):1707–1718, July 1, 2010. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947310000265.

Lefebvre:2010:PSI

[166] Geneviève Lefebvre, Russell Steele, and Alain C. Vandal. A path sampling identity for computing the Kullback– Leibler and J divergences. Computational Statistics & Data Analysis, 54 (7):1719–1731, July 1, 2010. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310000332.

Raqab:2010:PPD

[167] Mohammad Z. Raqab, A. Asgharzadeh, and R. Valiollahi. Prediction for Pareto distribution based on progressively Type-II censored samples. Computational Statistics & Data Analysis, 54(7):1732– 1743, July 1, 2010. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310000563.

Wang:2010:FSI

[168] Yong Wang. Fisher scoring: an interpolation family and its Monte Carlo implementations. Computational Statistics & Data Analysis, 54 (7):1744–1755, July 1, 2010. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310000575.

Wong:2010:PEG

[169] Tzu-Tsung Wong. Parameter estimation for generalized Dirichlet distributions from the sample estimates of the first and the second moments of random variables. Computational Statistics & Data Analysis, 54(7): 1756–1765, July 1, 2010. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310000599.

Priebe:2010:SIA

[170] Carey E. Priebe, Youngser Park, David J. Marchette, John M. Conroy, John Grothendieck, and Allen L. Gorin. Statistical inference on attributed random graphs: Fusion of graph features and content: an experiment on time series of Enron graphs. Computational Statistics & Data Analysis, 54(7):1766– 1776, July 1, 2010. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794731000023X.

Grothendieck:2010:SIA

[171] John Grothendieck, Carey E. Priebe, and Allen L. Gorin. Statistical inference on attributed random graphs: Fusion of graph features and content. Computational Statistics & Data Analysis, 54(7):1777– 1790, July 1, 2010. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310000320.

Koch:2010:PMR

[172] Inge Koch and Kanta Naito. Prediction of multivariate responses with a selected number of principal components. Computational Statistics & Data Analysis, 54(7):1791– 1807, July 1, 2010. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310000459.

Xu:2010:MIM

[173] Linzhi Xu and Jiajia Zhang. Multiple imputation method for the semiparametric accelerated failure time mixture cure model. Computational Statistics & Data Analysis, 54(7): 1808–1816, July 1, 2010. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310000502.

Zhang:2010:RAC

[174] Xinyan Zhang and Jianguo Sun. Regression analysis of clustered intervalcensored failure time data with informative cluster size. Computational Statistics & Data Analysis, 54(7): 1817–1823, July 1, 2010. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310000526.

Panagiotelis:2010:BSS

[175] Anastasios Panagiotelis and Michael Smith. Bayesian skew selection for multivariate models. Computational Statistics & Data Analysis, 54(7): 1824–1839, July 1, 2010. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310000551.

Bathke:2010:SRC

[176] Arne C. Bathke, Solomon W. Harrar, Haiyan Wang, Ke Zhang, and Hans-Peter Piepho. Series of randomized complete block experiments with non-normal data. Computational Statistics & Data Analysis, 54 (7):1840–1857, July 1, 2010. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310000587.

Timmerman:2010:FRM

[177] Marieke E. Timmerman, Eva Ceulemans, Henk A. L. Kiers, and Maurizio Vichi. Factorial and reduced Kmeans reconsidered. Computational Statistics & Data Analysis, 54(7): 1858–1871, July 1, 2010. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310000745.

Lin:2010:GFT

[178] Kuo-Chin Lin. Goodness-of-fit tests for modeling longitudinal ordinal data. Computational Statistics & Data Analysis, 54(7):1872– 1880, July 1, 2010. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310000782.

Kherad-Pajouh:2010:EPM

[179] Sara Kherad-Pajouh and Olivier Renaud. An exact permutation method for testing any effect in balanced and unbalanced fixed effect ANOVA. Computational Statistics & Data Analysis, 54(7):1881–1893, July 1, 2010. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310000800.

Anonymous:2010:EBk

[180] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 54(8):iii–v, August 1, 2010. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310001362.

Anonymous:2010:Ck

[181] Anonymous. Contents. Computational Statistics & Data Analysis, 54 (8):vi–vii, August 1, 2010. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310001374.

Konietschke:2010:TEP

[182] F. Konietschke, A. C. Bathke, L. A. Hothorn, and E. Brunner. Testing and estimation of purely nonparametric effects in repeated measures designs. Computational Statistics & Data Analysis, 54(8):1895– 1905, August 1, 2010. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310000848.

Candel:2010:OUV

[183] Math J. J. M. Candel and Gerard J. P. Van Breukelen. D-optimality of unequal versus equal cluster sizes for mixed effects linear regression analysis of randomized trials with clusters in one treatment arm. Computational Statistics & Data Analysis, 54(8):1906– 1920, August 1, 2010. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016794731000085X.

Yu:2010:BMA

[184] Binbing Yu. A Bayesian MCMC approach to survival analysis with doubly-censored data. Computational Statistics & Data Analysis, 54(8):1921– 1929, August 1, 2010. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947310000903.

Seo:2010:CSD

[185] Byungtae Seo and Bruce G. Lindsay. A computational strategy for doubly smoothed MLE exemplified in the normal mixture model. Computational Statistics & Data Analysis, 54(8):1930– 1941, August 1, 2010. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947310000927.

Heuchenne:2010:GFT

[186] Cédric Heuchenne and Ingrid Van Keilegom. Goodness-of-fit tests for the error distribution in nonparametric regression. Computational Statistics & Data Analysis, 54(8):1942– 1951, August 1, 2010. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947310000757.

Chen:2010:ATL

[187] Ying Chen, Chi Kin Chan, and Bartholomew P. K. Leung. An analysis of three-level orthogonal saturated designs. Computational Statistics & Data Analysis, 54(8):1952– 1961, August 1, 2010. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310000836.

Unnikrishnan:2010:BAO

[188] N. K. Unnikrishnan. Bayesian analysis for outliers in survey sampling. Computational Statistics & Data Analysis, 54(8):1962–1974, August 1, 2010. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310000861.

Wen:2010:SDR

[189] Xuerong Meggie Wen. On sufficient dimension reduction for proportional censorship model with covariates. Computational Statistics & Data Analysis, 54(8):1975–1982, August 1, 2010. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310000873.

Escanciano:2010:DDS

[190] Juan Carlos Escanciano and Silvia Mayoral. Data-driven smooth tests for the martingale difference hypothesis. Computational Statistics & Data Analysis, 54(8):1983– 1998, August 1, 2010. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310000885.

Higgs:2010:CLV

[191] Megan Dailey Higgs and Jennifer A. Hoeting. A clipped latent variable model for spatially correlated ordered categorical data. Computational Statistics & Data Analysis, 54(8):1999– 2011, August 1, 2010. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947310000897.

Ramos:2010:BCB

[192] Sandra Ramos, Antónia Amaral Turkman, and Marília Antunes. Bayesian classification for bivariate normal gene expression. Computational Statistics & Data Analysis, 54(8):2012– 2020, August 1, 2010. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947310000915.

Garcia:2010:NSG

[193] V. J. García, E. Gómez-Déniz, and F. J. Vázquez-Polo. A new skew generalization of the normal distribution: Properties and applications. Computational Statistics & Data Analysis, 54(8):2021–2034, August 1, 2010. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947310000940.

Martinez-Camblor:2010:NST

[194] Pablo Martínez-Camblor. Nonparametric k-sample test based on kernel density estimator for paired design. Computational Statistics & Data Analysis, 54(8):2035–2045, August 1, 2010. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310001118.

Chung:2010:FDS

[195] Yuk-Ka Chung, Wing K. Fung, and Yue-Qing Hu. Familial database search on two-person mixture. Computational Statistics & Data Analysis, 54(8):2046– 2051, August 1, 2010. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310000939.

Anonymous:2010:EBl

[196] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 54 (9):iii–v, September 1, 2010. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310001726.

Anonymous:2010:Cl

[197] Anonymous. Contents. Computational Statistics & Data Analysis, 54 (9):vi, September 1, 2010. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310001738.

Kossler:2010:MTR

[198] Wolfgang Kössler. Max-type rank tests, U-tests, and adaptive tests for the twosample location problem — an asymptotic power study. Computational Statistics & Data Analysis, 54(9):2053– 2065, September 1, 2010. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310001167.

Han:2010:ISS

[199] Donghoon Han and N. Balakrishnan. Inference for a simple step-stress model with competing risks for failure from the exponential distribution under time constraint. Computational Statistics & Data Analysis, 54(9):2066– 2081, September 1, 2010. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310001179.

Ojeda:2010:NIS

[200] Silvia Ojeda, Ronny Vallejos, and Oscar Bustos. A new image segmentation algorithm with applications to image inpainting. Computational Statistics & Data Analysis, 54(9):2082– 2093, September 1, 2010. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310001234.

Wetzels:2010:EPG

[201] Ruud Wetzels, Raoul P. P. P. Grasman, and Eric-Jan Wagenmakers. An encompassing prior generalization of the Savage–Dickey density ratio. Computational Statistics & Data Analysis, 54(9):2094–2102, September 1, 2010. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310001180.

Lai:2010:ELT

[202] Xin Lai and Kelvin K. W. Yau. Extending the long-term survivor mixture model with random effects for clustered survival data. Computational Statistics & Data Analysis, 54(9):2103– 2112, September 1, 2010. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310001192.

Gao:2010:JSS

[203] Jinxin Gao and David B. Hitchcock. James–Stein shrinkage to improve kmeans cluster analysis. Computational Statistics & Data Analysis, 54(9):2113– 2127, September 1, 2010. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310001209.

Gumedze:2010:VSM

[204] Freedom N. Gumedze, Sue J. Welham, Beverley J. Gogel, and Robin Thompson. A variance shift model for detection of outliers in the linear mixed model. Computational Statistics & Data Analysis, 54(9):2128– 2144, September 1, 2010. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310001210.

Chen:2010:BAS

[205] Xue-Dong Chen and Nian-Sheng Tang. Bayesian analysis of semiparametric reproductive dispersion mixed-effects models. Computational Statistics & Data Analysis, 54(9):2145–2158, September 1, 2010. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310001246.

Salvati:2010:SAE

[206] Nicola Salvati, Hukum Chandra, M. Giovanna Ranalli, and Ray Chambers. Small area estimation using a nonparametric model-based direct estimator. Computational Statistics & Data Analysis, 54(9):2159–2171, September 1, 2010. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310001258.

Yang:2010:SBH

[207] Mingan Yang, David B. Dunson, and Donna Baird. Semiparametric Bayes hierarchical models with mean and variance constraints. Computational Statistics & Data Analysis, 54(9):2172– 2186, September 1, 2010. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310001271.

Ginebra:2010:MEE

[208] Josep Ginebra and Xavier Puig. On the measure and the estimation of evenness and diversity. Computational Statistics & Data Analysis, 54(9):2187– 2201, September 1, 2010. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310001313.

Anonymous:2010:EBb

[209] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 54 (10):iii–v, October 1, 2010. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794731000201X.

Anonymous:2010:Cb

[210] Anonymous. Contents. Computational Statistics & Data Analysis, 54 (10):vi–vii, October 1, 2010. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310002021.

Chang:2010:ESB

[211] Yuan-Chin Ivan Chang, Yufen Huang, and Yu-Pai Huang. Early stopping in L_{2} boosting. Computational Statistics & Data Analysis, 54(10):2203– 2213, October 1, 2010. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794731000126X.

Chaouch:2010:DBE

[212] Mohamed Chaouch and Camelia Goga. Design-based estimation for geometric quantiles with application to outlier detection. Computational Statistics & Data Analysis, 54(10):2214– 2229, October 1, 2010. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310001088.

Wang:2010:VSC

[213] Xiaoming Wang, Taesung Park, and K. C. Carriere. Variable selection via combined penalization for high-dimensional data analysis. Computational Statistics & Data Analysis, 54(10):2230–2243, October 1, 2010. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310001283.

Zhang:2010:MCC

[214] Jiujun Zhang, Zhonghua Li, and Zhaojun Wang. A multivariate control chart for simultaneously monitoring process mean and variability. Computational Statistics & Data Analysis, 54(10): 2244–2252, October 1, 2010. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310001295.

Young:2010:MRP

[215] D. S. Young and D. R. Hunter. Mixtures of regressions with predictordependent mixing proportions. Computational Statistics & Data Analysis, 54(10):2253–2266, October 1, 2010. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0167947310001465.

White:2010:ABD

[216] Ian R. White, Rhian Daniel, and Patrick Royston. Avoiding bias due to perfect prediction in multiple imputation of incomplete categorical variables. Computational Statistics & Data Analysis, 54(10):2267– 2275, October 1, 2010. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310001490.

Song:2010:EMP

[217] Seongjoo Song, Dan L. Nicolae, and Jongwoo Song. Estimating the mixing proportion in a semiparametric mixture model. Computational Statistics & Data Analysis, 54(10):2276– 2283, October 1, 2010. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310001520.

Consentino:2010:OST

[218] Fabrizio Consentino and Gerda Claeskens. Order selection tests with multiply imputed data. Computational Statistics & Data Analysis, 54(10):2284– 2295, October 1, 2010. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310001544.

Afshartous:2010:CID

[219] David Afshartous and Richard A. Preston. Confidence intervals for dependent data: Equating non-overlap with statistical significance. Computational Statistics & Data Analysis, 54(10): 2296–2305, October 1, 2010. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310001568.

Cahoy:2010:BTE

[220] Dexter O. Cahoy. A bootstrap test for equality of variances. Computational Statistics & Data Analysis, 54(10): 2306–2316, October 1, 2010. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016794731000157X.

Zornig:2010:SSD

[221] Peter Zörnig. Statistical simulation and the distribution of distances between identical elements in a random sequence. Computational Statistics & Data Analysis, 54(10):2317– 2327, October 1, 2010. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016794731000006X.

Huwang:2010:NEC

[222] Longcheen Huwang, Chun-Jung Huang, and Yi-Hua Tina Wang. New EWMA control charts for monitoring process dispersion. Computational Statistics & Data Analysis, 54(10):2328– 2342, October 1, 2010. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947310001131.

Kohn:2010:RID

[223] Hans-Friedrich Köhn. Representation of individual differences in rectangular proximity data through anti-Q matrix decomposition. Computational Statistics & Data Analysis, 54(10): 2343–2357, October 1, 2010. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310001143.

Gervini:2010:RFSb

[224] Daniel Gervini. Retracted: the functional singular value decomposition for bivariate stochastic processes. Computational Statistics & Data Analysis, 54 (10):2358, October 1, 2010. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310002094.

Anonymous:2010:EBc

[225] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 54 (11):iii–v, November 1, 2010. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310002288.

Anonymous:2010:Cc

[226] Anonymous. Contents. Computational Statistics & Data Analysis, 54 (11):vi–vii, November 1, 2010. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794731000229X.

Belsley:2010:FSI

[227] David A. Belsley, Pierre Duchesne, George Kapetanios, Erricos John Kontoghiorghes, Marc Paolella, and Herman K. van Dijk. The fifth special issue on computational econometrics. Computational Statistics & Data Analysis, 54(11):2359, November 1, 2010. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947310002343.

Alp:2010:JFD

[228] Tansel Alp and Matei Demetrescu. Joint forecasts of Dow Jones stocks under general multivariate loss function. Computational Statistics & Data Analysis, 54(11):2360–2371, November 1, 2010. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947309003570.

Audrino:2010:MTT

[229] Francesco Audrino and Fulvio Corsi. Modeling tick-by-tick realized correlations. Computational Statistics & Data Analysis, 54(11):2372–2382, November 1, 2010. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947309003636.

Ausin:2010:TVJ

[230] M. Concepcion Ausin and Hedibert F. Lopes. Time-varying joint distribution through copulas. Computational Statistics & Data Analysis, 54(11): 2383–2399, November 1, 2010. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947309001042.

Bauwens:2010:IDP

[231] Luc Bauwens, Walid Ben Omrane, and Erick Rengifo. Intradaily dynamic portfolio selection. Computational Statistics & Data Analysis, 54(11):

2400–2418, November 1, 2010. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309002266.

Omrane:2010:PNA

[232] Walid Ben Omrane and Andréas Heinen. Public news announcements and quoting activity in the euro/dollar foreign exchange market. Computational Statistics & Data Analysis, 54(11):2419–2431, November 1, 2010. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309004290.

Beran:2010:SLM

[233] Jan Beran, Martin Schützner, and Sucharita Ghosh. From short to long memory: Aggregation and estimation. Computational Statistics & Data Analysis, 54(11):2432–2442, November 1, 2010. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309002059.

Billio:2010:MLV

[234] Monica Billio and Massimiliano Caporin. Market linkages, variance spillovers, and correlation stability: Empirical evidence of financial contagion. Computational Statistics & Data Analysis, 54(11):2443–2458, November 1, 2010. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309001297.

Boudt:2010:REM

[235] Kris Boudt and Christophe Croux. Robust M-estimation of multivariate GARCH models. Computational Statistics & Data Analysis, 54(11): 2459–2469, November 1, 2010. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947309004101.

Cipollini:2010:AVS

[236] Fabrizio Cipollini and Giampiero M. Gallo. Automated variable selection in vector multiplicative error models. Computational Statistics & Data Analysis, 54(11):2470–2486, November 1, 2010. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016794730900276X.

Dai:2010:SEB

[237] J. Dai and S. Sperlich. Simple and effective boundary correction for kernel densities and regression with an application to the world income and Engel curve estimation. Computational Statistics & Data Analysis, 54(11): 2487–2497, November 1, 2010. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947310001519.

Davidson:2010:TCS

[238] James Davidson and Andrea Monticini. Tests for cointegration with structural breaks based on subsamples. Computational Statistics & Data Analysis, 54(11):2498–2511, November 1, 2010. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310000435.

Duchesne:2010:TSC

[239] Pierre Duchesne, Linyuan Li, and Jill Vandermeerschen. On testing for serial correlation of unknown form using wavelet thresholding. Computational Statistics & Data Analysis, 54(11): 2512–2531, November 1, 2010. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794730900423X.

Dufour:2010:EOI

[240] Jean-Marie Dufour and Abderrahim Taamouti. Exact optimal inference in regression models under heteroskedasticity and non-normality of unknown form. Computational Statistics & Data Analysis, 54(11):2532– 2553, November 1, 2010. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309003715.

Dufour:2010:EUS

[241] Jean-Marie Dufour, Lynda Khalaf, and Maral Kichian. Estimation uncertainty in structural inflation models with real wage rigidities. Computational Statistics & Data Analysis, 54(11):2554– 2561, November 1, 2010. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794730900142X.

Fantazzini:2010:TSS

[242] Dean Fantazzini. Three-stage semiparametric estimation of T -copulas: Asymptotics, finite-sample properties and computational aspects. Computational Statistics & Data Analysis, 54(11):2562–2579, November 1, 2010. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0167947309000371.

Grane:2010:WBD

[243] Aurea Grané and Helena Veiga. Wavelet-based detection of outliers in financial time series. Computational Statistics & Data Analysis, 54(11): 2580–2593, November 1, 2010. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947309004629.

Griffin:2010:BIS

[244] J. E. Griffin and M. F. J. Steel. Bayesian inference with stochastic volatility models using continuous superpositions of non-Gaussian Ornstein–Uhlenbeck processes. Computational Statistics & Data Analysis, 54(11):2594–2608, November 1, 2010. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0167947309002382.

Hafner:2010:EES

[245] Christian M. Hafner and Olga Reznikova. Efficient estimation of a semiparametric dynamic copula model. Computational Statistics & Data Analysis, 54(11):2609–2627, November 1, 2010. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310000289.

He:2010:RTD

[246] Zhongfang He and John M. Maheu. Real time detection of structural breaks in GARCH models. Computational Statistics & Data Analysis, 54(11): 2628–2640, November 1, 2010. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309003685.

Hindrayanto:2010:EML

[247] Irma Hindrayanto, Siem Jan Koopman, and Marius Ooms. Exact maximum likelihood estimation for non-stationary periodic time series models. Computational Statistics & Data Analysis, 54(11):2641–2654, November 1, 2010. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310001556.

Kapetanios:2010:FGE

[248] George Kapetanios and Massimiliano Marcellino. Factor-GMM estimation with large sets of possibly weak instruments. Computational Statistics & Data Analysis, 54(11):2655–2675, November 1, 2010. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310001532.

Lux:2010:FVU

[249] Thomas Lux and Leonardo Morales-Arias. Forecasting volatility under fractality, regime-switching, long memory and Student-t innovations. Computational Statistics & Data Analysis, 54(11):2676–2692, November 1, 2010. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0167947310001076.

Lyra:2010:OHD

[250] M. Lyra, J. Paha, S. Paterlini, and P. Winker. Optimization heuristics for determining internal rating grading scales. Computational Statistics & Data Analysis, 54(11):2693– 2706, November 1, 2010. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947309001005.

Naylor:2010:EDA

[251] J. C. Naylor, A. R. Tremayne, and J. M. Marriott. Exploratory data analysis and model criticism with posterior plots. Computational Statistics & Data Analysis, 54(11):2707– 2720, November 1, 2010. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947309000735.

Ohtsuka:2010:FED

[252] Yoshihiro Ohtsuka, Takashi Oga, and Kazuhiko Kakamu. Forecasting electricity demand in Japan: a Bayesian spatial autoregressive ARMA approach. Computational Statistics & Data Analysis, 54(11):2721–2735, November 1, 2010. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794730900231X.

Omori:2010:TMC

[253] Yasuhiro Omori and Koji Miyawaki. Tobit model with covariate dependent thresholds. Computational Statistics & Data Analysis, 54(11):2736– 2752, November 1, 2010. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309000383.

Pastorello:2010:EIS

[254] S. Pastorello and E. Rossi. Efficient importance sampling maximum likelihood estimation of stochastic differential equations. Computational Statistics & Data Analysis, 54(11):2753– 2762, November 1, 2010. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310000496.

Pendakur:2010:SIU

[255] Krishna Pendakur, Michael Scholz, and Stefan Sperlich. Semiparametric indirect utility and consumer demand. Computational Statistics & Data Analysis, 54(11):2763–2775, November 1, 2010. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310001489.

Postiglione:2010:RTA

[256] Paolo Postiglione, Roberto Benedetti, and Giovanni Lafratta. A regression tree algorithm for the identification of convergence clubs. Computational Statistics & Data Analysis, 54(11): 2776–2785, November 1, 2010. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947309001583.

Rossi:2010:MDU

[257] E. Rossi and F. Spazzini. Model and distribution uncertainty in multivariate GARCH estimation: a Monte Carlo analysis. Computational Statistics & Data Analysis, 54(11):2786– 2800, November 1, 2010. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947309002345.

Shin:2010:RPU

[258] Dong Wan Shin and Sangun Park. Robust panel unit root tests for cross-sectionally dependent multiple time series. Computational Statistics & Data Analysis, 54(11):2801– 2813, November 1, 2010. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947310000769.

Strid:2010:EPM

[259] Ingvar Strid. Efficient parallelisation of Metropolis–Hastings algorithms using a prefetching approach. Computational Statistics & Data Analysis, 54(11): 2814–2835, November 1, 2010. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309004320.

Vassiliou:2010:LDL

[260] E. E. Vassiliou and I. C. Demetriou. A linearly distributed lag estimator with r-convex coefficients. Computational Statistics & Data Analysis, 54(11): 2836–2849, November 1, 2010. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310000447.

Vilar:2010:NLT

[261] J. A. Vilar, A. M. Alonso, and J. M. Vilar. Non-linear time series clustering based on non-parametric forecast densities. Computational Statistics & Data Analysis, 54(11):2850– 2865, November 1, 2010. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794730900067X.

Wang:2010:SSU

[262] Hao Wang. Sparse seemingly unrelated regression modelling: Applications in finance and econometrics. Computational Statistics & Data Analysis, 54(11):2866–2877, November 1, 2010. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310001301.

Anonymous:2010:EBd

[263] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 54 (12):iii–v, December 1, 2010. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947310002938.

Anonymous:2010:Cd

[264] Anonymous. Contents. Computational Statistics & Data Analysis, 54(12): vi–viii, December 1, 2010. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016794731000294X.

VanAelst:2010:SIV

[265] Stefan Van Aelst, Roy Welsch, and Ruben H. Zamar. Special issue on variable selection and robust procedures. Computational Statistics & Data Analysis, 54(12):2879–2882, December 1, 2010. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947310002999.

Abanto-Valle:2010:RBA

[266] C. A. Abanto-Valle, D. Bandyopadhyay, V. H. Lachos, and I. Enriquez. Robust Bayesian analysis of heavy-tailed stochastic volatility models using scale mixtures of normal distributions. Computational Statistics & Data Analysis, 54(12):2883– 2898, December 1, 2010. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947309002448.

Adrover:2010:GRC

[267] Jorge Adrover and Matias Salibian-Barrera. Globally robust confidence intervals for simple linear regression. Computational Statistics & Data Analysis, 54(12):2899–2913, December 1, 2010. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309001923.

Alvarez-Esteban:2010:AWS

[268] Pedro C. Alvarez-Esteban, Eustasio del Barrio, Juan A. Cuesta-Albertos, and Carlos Matrán. Assessing when a sample is mostly normal. Computational Statistics & Data Analysis, 54(12): 2914–2925, December 1, 2010. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309004423.

Basso:2010:RMM

[269] Rodrigo M. Basso, Víctor H. Lachos, Celso Rômulo Barbosa Cabral, and Pulak Ghosh. Robust mixture modeling based on scale mixtures of skewnormal distributions. Computational Statistics & Data Analysis, 54(12): 2926–2941, December 1, 2010. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309003612.

Boente:2010:RIG

[270] Graciela Boente and Daniela Rodriguez. Robust inference in generalized partially linear models. Computational Statistics & Data Analysis, 54(12):2942–2966, December 1, 2010. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0167947310002367.

Boente:2010:DIO

[271] Graciela Boente, Ana M. Pires, and Isabel M. Rodrigues. Detecting influential observations in principal components and common principal components. Computational Statistics & Data Analysis, 54(12):2967–2975, December 1, 2010. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310000022.

Borra:2010:MPE

[272] Simone Borra and Agostino Di Ciaccio. Measuring the prediction error. a comparison of cross-validation, bootstrap and covariance penalty methods. Computational Statistics & Data Analysis, 54(12):2976–2989, December 1, 2010. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310001064.

Chen:2010:NAS

[273] Yin-Ping Chen, Hsin-Cheng Huang, and I-Ping Tu. A new approach for selecting the number of factors. Computational Statistics & Data Analysis, 54(12):2990–2998, December 1, 2010. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309003727.

Croux:2010:RES

[274] Christophe Croux, Sarah Gelper, and Koen Mahieu. Robust exponential smoothing of multivariate time series. Computational Statistics & Data Analysis, 54(12):2999– 3006, December 1, 2010. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309001728.

Debruyne:2010:DIO

[275] Michiel Debruyne, Mia Hubert, and Johan Van Horebeek. Detecting influential observations in kernel PCA. Computational Statistics & Data Analysis, 54(12):3007–3019, December 1, 2010. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309003120.

Lozeron:2010:REC

[276] E. Dupuis Lozeron and M. P. Victoria-Feser. Robust estimation of constrained covariance matrices for confirmatory factor analysis. Computational Statistics & Data Analysis, 54(12): 3020–3032, December 1, 2010. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309003089.

Emura:2010:GFT

[277] Takeshi Emura, Chien-Wei Lin, and Weijing Wang. A goodness-of-fit test for Archimedean copula models in the presence of right censoring. Computational Statistics & Data Analysis, 54(12):3033–3043, December 1, 2010. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947310001155.

Flores:2010:ECR

[278] Salvador Flores. On the efficient computation of robust regression estimators. Computational Statistics & Data Analysis, 54(12):3044–3056, December 1, 2010. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947310001222.

Garcia-Escudero:2010:RCL

[279] L. A. García-Escudero, A. Gordaliza, A. Mayo-Iscar, and R. San Martín. Robust clusterwise linear regression through trimming. Computational Statistics & Data Analysis, 54(12): 3057–3069, December 1, 2010. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947309002497.

Gottard:2010:RCG

[280] Anna Gottard and Simona Pacillo. Robust concentration graph model selection. Computational Statistics & Data Analysis, 54(12):3070–3079, December 1, 2010. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947308005550.

Haight:2010:CVD

[281] Thaddeus J. Haight, Yue Wang, Mark J. van der Laan, and Ira B. Tager. A cross-validation deletion– substitution–addition model selection algorithm: Application to marginal structural models. Computational Statistics & Data Analysis, 54(12): 3080–3094, December 1, 2010. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310000514.

Hron:2010:IMV

[282] K. Hron, M. Templ, and P. Filzmoser. Imputation of missing values for compositional data using classical and robust methods. Computational Statistics & Data Analysis, 54(12):3095– 3107, December 1, 2010. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309004368.

Jureckova:2010:RTR

[283] Jana Jurečková, Jan Picek, and A. K. Md. Ehsanes Saleh. Rank tests and regression rank score tests in measurement error models. Computational Statistics & Data Analysis, 54(12): 3108–3120, December 1, 2010. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794730900303X.

Khan:2010:FRE

[284] Jafar A. Khan, Stefan Van Aelst, and Ruben H. Zamar. Fast robust estimation of prediction error based on resampling. Computational Statistics & Data Analysis, 54(12):3121– 3130, December 1, 2010. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310000460.

Lee:2010:RSS

[285] Jong Soo Lee and Dennis D. Cox. Robust smoothing: Smoothing parameter selection and applications to fluorescence spectroscopy. Computational Statistics & Data Analysis, 54(12): 3131–3143, December 1, 2010. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309002758.

Lykou:2010:SCU

[286] Anastasia Lykou and Joe Whittaker. Sparse CCA using a lasso with positivity constraints. Computational Statistics & Data Analysis, 54(12):3144– 3157, December 1, 2010. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309002783.

Marley:2010:CDM

[287] Christopher J. Marley and David C. Woods. A comparison of design and model selection methods for supersaturated experiments. Computational Statistics & Data Analysis, 54(12): 3158–3167, December 1, 2010. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310000824.

Maronna:2010:CME

[288] Ricardo A. Maronna and Victor J. Yohai. Correcting MM estimates for “fat” data sets. Computational Statistics & Data Analysis, 54(12):3168– 3173, December 1, 2010. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309003314.

Meintanis:2010:TGN

[289] Simos G. Meintanis and Efthimios Tsionas. Testing for the generalized normal-Laplace distribution with applications. Computational Statistics & Data Analysis, 54(12):3174– 3180, December 1, 2010. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309002187. See

erratum [651].

Menjoge:2010:DMS

[290] Rajiv S. Menjoge and Roy E. Welsch. A diagnostic method for simultaneous feature selection and outlier identification in linear regression. Computational Statistics & Data Analysis, 54(12): 3181–3193, December 1, 2010. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310000794.

Min:2010:MSS

[291] Aleksey Min, Hajo Holzmann, and Claudia Czado. Model selection strategies for identifying most relevant covariates in homoscedastic linear models. Computational Statistics & Data Analysis, 54(12):3194– 3211, December 1, 2010. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309003223.

Nguyen:2010:ODL

[292] T. D. Nguyen and R. Welsch. Outlier detection and least trimmed squares approximation using semi-definite programming. Computational Statistics & Data Analysis, 54(12):3212– 3226, December 1, 2010. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309003673.

Nott:2010:BPA

[293] David J. Nott and Chenlei Leng. Bayesian projection approaches to variable selection in generalized linear models. Computational Statistics & Data Analysis, 54(12):3227– 3241, December 1, 2010. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794731000054X.

Nunkesser:2010:EAR

[294] Robin Nunkesser and Oliver Morell. An evolutionary algorithm for robust regression. Computational Statistics & Data Analysis, 54(12):3242– 3248, December 1, 2010. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310001623.

Ouysse:2010:BVS

[295] Rachida Ouysse and Robert Kohn. Bayesian variable selection and model averaging in the arbitrage pricing theory model. Computational Statistics & Data Analysis, 54(12):3249– 3268, December 1, 2010. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309003648.

Overstall:2010:DBM

[296] Antony M. Overstall and Jonathan J. Forster. Default Bayesian model determination methods for generalised linear mixed models. Computational Statistics & Data Analysis, 54(12): 3269–3288, December 1, 2010. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310001106.

Powers:2010:BVS

[297] Stephanie Powers, Richard Gerlach, and James Stamey. Bayesian variable selection for Poisson regression with underreported responses. Computational Statistics & Data Analysis, 54(12):3289–3299, December 1, 2010. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0167947310001477.

Riani:2010:RMS

[298] Marco Riani and Anthony C. Atkinson. Robust model selection with flexible trimming. Computational Statistics & Data Analysis, 54(12):3300– 3312, December 1, 2010. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016794731000109X.

Rozenholc:2010:CRI

[299] Yves Rozenholc, Thoralf Mildenberger, and Ursula Gather. Combining regular and irregular histograms by penalized likelihood. Computational Statistics & Data Analysis, 54(12):3313– 3323, December 1, 2010. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947310001660.

Rufo:2010:NAC

[300] M. J. Rufo, J. Martín, and C. J. Pérez. New approaches to compute Bayes factor in finite mixture models. Computational Statistics & Data Analysis, 54(12):3324–3335, December 1, 2010. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947310001891.

Schomaker:2010:FMA

[301] Michael Schomaker, Alan T. K. Wan, and Christian Heumann. Frequentist model averaging with missing observations. Computational Statistics & Data Analysis, 54(12):3336– 3347, December 1, 2010. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947309002667.

Unkel:2010:MAS

[302] S. Unkel and N. T. Trendafilov. A majorization algorithm for simultaneous parameter estimation in robust exploratory factor analysis. Computational Statistics & Data Analysis, 54(12):3348–3358, December 1, 2010. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310000538.

Wang:2010:MSL

[303] You-Gan Wang and Lin-Yee Hin. Modeling strategies in longitudinal data analysis: Covariate, variance function and correlation structure selection. Computational Statistics & Data Analysis, 54(12):3359–3370, December 1, 2010. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309004095.

Wiens:2010:RDT

[304] Douglas P. Wiens. Robustness of design for the testing of lack of fit and for estimation in binary response models. Computational Statistics & Data Analysis, 54(12):3371– 3378, December 1, 2010. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309000838.

Barlow:2010:SIM

[305] Jesse Barlow, Lars Eldén, and Paolo Foschi. 3rd special issue on matrix computations and statistics. Computational Statistics & Data Analysis, 54(12):3379–3380, December 1, 2010. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310003002.

Arnouts:2010:UFS

[306] Heidi Arnouts and Peter Goos. Update formulas for split-plot and block designs. Computational Statistics & Data Analysis, 54(12):3381–3391, December 1, 2010. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309004253.

Hofmann:2010:MSC

[307] Marc Hofmann and Erricos John Kontoghiorghes. Matrix strategies for computing the least trimmed squares estimation of the general linear and SUR models. Computational Statistics & Data Analysis, 54(12):3392– 3403, December 1, 2010. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310001817.

Ichimura:2010:FAC

[308] Tsuyoshi Ichimura and Daisuke Fukuda. A fast algorithm for computing leastsquares cross-validations for nonparametric conditional kernel density functions. Computational Statistics & Data Analysis, 54(12):3404–3410, December 1, 2010. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309003168.

Markovsky:2010:ALR

[309] Ivan Markovsky and Mahesan Niranjan. Approximate low-rank factorization with structured factors. Computational Statistics & Data Analysis, 54(12):3411–3420, December 1, 2010. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309002321.

Papkov:2010:LMN

[310] Galen I. Papkov and David W. Scott. Local-moment nonparametric density estimation of pre-binned data. Computational Statistics & Data Analysis, 54(12):3421–3429, December 1, 2010. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310000770.

Renaut:2010:RPE

[311] Rosemary A. Renaut, Iveta Hnětynková, and Jodi Mead. Regularization parameter estimation for large-scale Tikhonov regularization using a priori information. Computational Statistics & Data Analysis, 54(12):3430– 3445, December 1, 2010. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309002278.

Trendafilov:2010:SEC

[312] Nickolay T. Trendafilov. Stepwise estimation of common principal components. Computational Statistics & Data Analysis, 54(12):3446–3457, December 1, 2010. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016794731000112X.

Yue:2010:ODH

[313] Rong-Xian Yue and Xin Liu. I_{L}^{r}- optimal designs for a hierarchically ordered system of regression models. Computational Statistics & Data Analysis, 54(12):3458–3465, December 1, 2010. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947309003259.

Anonymous:2011:EBa

[314] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 55 (1):iv–vi, January 1, 2011. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947310003336.

Anonymous:2011:Ca

[315] Anonymous. Contents. Computational Statistics & Data Analysis, 55 (1):vii–x, January 1, 2011. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947310003348.

Joarder:2011:IWP

[316] Avijit Joarder, Hare Krishna, and Debasis Kundu. Inferences on Weibull parameters with conventional type-I censoring. Computational Statistics & Data Analysis, 55(1):1–11, January 1, 2011. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310001507.

Cho:2011:AIP

[317] S.-J. Cho and S. Rabe-Hesketh. Alternating imputation posterior estimation of models with crossed random effects. Computational Statistics & Data Analysis, 55(1):12–25, January 1, 2011. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794731000160X.

Zhou:2011:FPC

[318] Hua Zhou and Kenneth Lange. A fast procedure for calculating importance weights in bootstrap sampling. Computational Statistics & Data Analysis, 55(1):26–33, January 1, 2011. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310001647.

DeGooijer:2011:SET

[319] Jan G. De Gooijer and Ao Yuan. Some exact tests for manifest properties of latent trait models. Computational Statistics & Data Analysis, 55 (1):34–44, January 1, 2011. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310001672.

Ormerod:2011:GBV

[320] John T. Ormerod. Grid based variational approximations. Computational Statistics & Data Analysis, 55 (1):45–56, January 1, 2011. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947310001829.

Hofert:2011:ESN

[321] Marius Hofert. Efficiently sampling nested Archimedean copulas. Computational Statistics & Data Analysis, 55 (1):57–70, January 1, 2011. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947310001830.

Bellio:2011:SSF

[322] Ruggero Bellio and Luca Grassetti. Semiparametric stochastic frontier models for clustered data. Computational Statistics & Data Analysis, 55 (1):71–83, January 1, 2011. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947310001866.

Yue:2011:BIA

[323] Yu Ryan Yue and Håvard Rue. Bayesian inference for additive mixed quantile regression models. Computational Statistics & Data Analysis, 55 (1):84–96, January 1, 2011. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947310001933.

Navarrete:2011:SAB

[324] Carlos A. Navarrete and Fernando A. Quintana. Similarity analysis in Bayesian random partition models. Computational Statistics & Data Analysis, 55(1):97–109, January 1, 2011. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794731000215X.

Chen:2011:PSS

[325] Jie Chen, Jianfeng Luo, Kenneth Liu, and Devan V. Mehrotra. On power and sample size computation for multiple testing procedures. Computational Statistics & Data Analysis, 55 (1):110–122, January 1, 2011. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310002355.

Brusco:2011:EAA

[326] Michael J. Brusco and Douglas Steinley. Exact and approximate algorithms for variable selection in linear discriminant analysis. Computational Statistics & Data Analysis, 55 (1):123–131, January 1, 2011. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310002380.

Barbillon:2011:NMI

[327] Pierre Barbillon, Gilles Celeux, Agnès Grimaud, Yannick Lefebvre, and Étienne De Rocquigny. Nonlinear methods for inverse statistical problems. Computational Statistics & Data Analysis, 55(1):132–142, January 1, 2011. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947310002422.

Kuroda:2011:AAL

[328] Masahiro Kuroda, Yuichi Mori, Masaya Iizuka, and Michio Sakakihara. Acceleration of the alternating least squares algorithm for principal components analysis. Computational Statistics & Data Analysis, 55 (1):143–153, January 1, 2011. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947310002458.

Hammer:2011:AFB

[329] Hugo Hammer and Håkon Tjelmeland. Approximate forward–backward algorithm for a switching linear Gaussian model. Computational Statistics & Data Analysis, 55(1):154–167, January 1, 2011. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947310002525.

Maalouf:2011:RWK

[330] Maher Maalouf and Theodore B. Trafalis. Robust weighted kernel logistic regression in imbalanced and rare events data. Computational Statistics & Data Analysis, 55(1):168– 183, January 1, 2011. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947310002598.

Xu:2011:UCD

[331] Chonggang Xu and George Gertner. Understanding and comparisons of different sampling approaches for the Fourier amplitudes sensitivity test (FAST). Computational Statistics & Data Analysis, 55(1):184– 198, January 1, 2011. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310002756.

McCausland:2011:SSS

[332] William J. McCausland, Shirley Miller, and Denis Pelletier. Simulation smoothing for state–space models: a computational efficiency analysis. Computational Statistics & Data Analysis, 55(1):199–212, January 1, 2011. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310002823.

Leydold:2011:GGI

[333] Josef Leydold and Wolfgang Hörmann. Generating generalized inverse Gaussian random variates by fast inversion. Computational Statistics & Data Analysis, 55(1):213–217, January 1, 2011. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310002847.

Patriota:2011:NID

[334] Alexandre G. Patriota. A note on influence diagnostics in nonlinear mixedeffects elliptical models. Computational Statistics & Data Analysis, 55 (1):218–225, January 1, 2011. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947310002513.

Tunes-da-Silva:2011:CSD

[335] Gisela Tunes da Silva and John P. Klein. Cutpoint selection for discretizing a continuous covariate for generalized estimating equations. Computational Statistics & Data Analysis, 55 (1):226–235, January 1, 2011. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947310000812.

Kateri:2011:OAC

[336] Maria Kateri, Udo Kamps, and Narayanaswamy Balakrishnan. Optimal allocation of change points in simple step-stress experiments under Type-II censoring. Computational Statistics & Data Analysis, 55(1):236– 247, January 1, 2011. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947310001581.

Lee:2011:HLA

[337] Woojoo Lee, Johan Lim, Youngjo Lee, and Joan del Castillo. The hierarchicallikelihood approach to autoregressive stochastic volatility models. Computational Statistics & Data Analysis, 55 (1):248–260, January 1, 2011. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947310001593.

Assmann:2011:BAM

[338] Christian Aßmann and Jens Boysen-Hogrefe. A Bayesian approach to model-based clustering for binary panel probit models. Computational Statistics & Data Analysis, 55(1):261–279, January 1, 2011. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310001611.

Pedersen:2011:NTD

[339] M. W. Pedersen, U. H. Thygesen, and H. Madsen. Nonlinear tracking in a diffusion process with a Bayesian filter and the finite element method. Computational Statistics & Data Analysis, 55 (1):280–290, January 1, 2011. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310001635.

Zhao:2011:SRA

[340] Xingqiu Zhao and Xingwei Tong. Semiparametric regression analysis of panel count data with informative observation times. Computational Statistics & Data Analysis, 55(1):291–300, January 1, 2011. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310001659.

Sotto:2011:MBE

[341] Cristina Sotto, Caroline Beunckens, Geert Molenberghs, and Michael G. Kenward. MCMC-based estimation methods for continuous longitudinal data with non-random (non)-monotone missingness. Computational Statistics & Data Analysis, 55(1):301–311, January 1, 2011. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947310001842.

Garcia-Ligero:2011:DCD

[342] M. J. García-Ligero, A. Hermoso-Carazo, and J. Linares-Pérez. Derivation of centralized and distributed filters using covariance information. Computational Statistics & Data Analysis, 55(1):312–323, January 1, 2011. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947310001854.

Hampel:2011:SPH

[343] Frank Hampel, Christian Hennig, and Elvezio Ronchetti. A smoothing principle for the Huber and other location M-estimators. Computational Statistics & Data Analysis, 55(1):324– 337, January 1, 2011. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947310001878.

Khorsheed:2011:MED

[344] Eman Khorsheed, Merrilee Hurn, and Christopher Jennison. Mapping electron density in the ionosphere: a principal component MCMC algorithm. Computational Statistics & Data Analysis, 55(1):338–352, January 1, 2011. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016794731000188X.

Azevedo:2011:BIS

[345] Caio L. N. Azevedo, Heleno Bolfarine, and Dalton F. Andrade. Bayesian inference for a skew-normal IRT model under the centred parameterization. Computational Statistics & Data Analysis, 55(1):353–365, January 1, 2011. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310001908.

Young:2011:GAM

[346] Robin L. Young, Janice Weinberg, Verónica Vieira, Al Ozonoff, and Thomas F. Webster. Generalized additive models and inflated type I error rates of smoother significance tests. Computational Statistics & Data Analysis, 55(1):366–374, January 1, 2011. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794731000191X.

Kuha:2011:IDL

[347] Jouni Kuha and David Firth. On the index of dissimilarity for lack of fit in loglinear and log-multiplicative models. Computational Statistics & Data Analysis, 55(1):375–388, January 1, 2011. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310001921.

Li:2011:PVC

[348] Jianbo Li and Riquan Zhang. Partially varying coefficient single index proportional hazards regression models. Computational Statistics & Data Analysis, 55(1):389–400, January 1, 2011. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310001945.

Delicado:2011:DRW

[349] P. Delicado. Dimensionality reduction when data are density functions. Computational Statistics & Data Analysis, 55(1):401–420, January 1, 2011. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310001957.

Kwon:2011:QAS

[350] Sunghoon Kwon, Hosik Choi, and Yongdai Kim. Quadratic approximation on SCAD penalized estimation. Computational Statistics & Data Analysis, 55(1):421–428, January 1, 2011. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310001969.

Lambert:2011:SSN

[351] Philippe Lambert. Smooth semiparametric and nonparametric Bayesian estimation of bivariate densities from bivariate histogram data. Computational Statistics & Data Analysis, 55 (1):429–445, January 1, 2011. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310002124.

Commenges:2011:IHD

[352] D. Commenges, D. Jolly, J. Drylewicz, H. Putter, and R. Thiébaut. Inference in HIV dynamics models via hierarchical likelihood. Computational Statistics & Data Analysis, 55(1):446– 456, January 1, 2011. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310002136.

Zhan:2011:GWL

[353] Tingting Zhan, Inna Chevoneva, and Boris Iglewicz. Generalized weighted likelihood density estimators with application to finite mixture of exponential family distributions. Computational Statistics & Data Analysis, 55 (1):457–465, January 1, 2011. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310002148.

Baksh:2011:EDT

[354] M. Fazil Baksh, Dankmar Böhning, and Rattana Lerdsuwansri. An extension of an over-dispersion test for count data. Computational Statistics & Data Analysis, 55(1):466–474, January 1, 2011. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310002161.

Parent:2011:STF

[355] Olivier Parent and James P. LeSage. A space–time filter for panel data models containing random effects. Computational Statistics & Data Analysis, 55 (1):475–490, January 1, 2011. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310002173.

Loughin:2011:CIF

[356] Thomas M. Loughin and Jorge E. Rodríguez. Computational issues with fitting joint location/dispersion models in unreplicated 2^{k} factorials. Computational Statistics & Data Analysis, 55 (1):491–497, January 1, 2011. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310002185.

Dereudre:2011:PSE

[357] D. Dereudre and F. Lavancier. Practical simulation and estimation for Gibbs Delaunay–Voronoi tessellations with geometric hardcore interaction. Computational Statistics & Data Analysis, 55(1):498–519, January 1, 2011. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310002197.

Andrews:2011:MBC

[358] Jeffrey L. Andrews, Paul D. McNicholas, and Sanjeena Subedi. Modelbased classification via mixtures of multivariate t-distributions. Computational Statistics & Data Analysis, 55 (1):520–529, January 1, 2011. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310002203.

Hoff:2011:HMM

[359] Peter D. Hoff. Hierarchical multilinear models for multiway data. Computational Statistics & Data Analysis, 55 (1):530–543, January 1, 2011. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310002215.

Cerioli:2011:ERM

[360] Andrea Cerioli and Alessio Farcomeni. Error rates for multivariate outlier detection. Computational Statistics & Data Analysis, 55(1):544–553, January 1, 2011. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310002227.

Zhou:2011:WQR

[361] Weihua Zhou. A weighted quantile regression for randomly truncated data. Computational Statistics & Data Analysis, 55(1):554–566, January 1, 2011. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310002239.

Lee:2011:FSL

[362] Sangjun Lee, Changyi Park, and Ja-Yong Koo. Feature selection in the Laplacian support vector machine. Computational Statistics & Data Analysis, 55(1):567–577, January 1, 2011. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310002379.

Iranpanah:2011:CBS

[363] N. Iranpanah, M. Mohammadzadeh, and C. C. Taylor. A comparison of block and semi-parametric bootstrap methods for variance estimation in spatial statistics. Computational Statistics & Data Analysis, 55 (1):578–587, January 1, 2011. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947310002434.

Cancho:2011:BNR

[364] Vicente G. Cancho, Dipak K. Dey, Victor H. Lachos, and Marinho G. Andrade. Bayesian nonlinear regression models with scale mixtures of skewnormal distributions: Estimation and case influence diagnostics. Computational Statistics & Data Analysis, 55 (1):588–602, January 1, 2011. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947310002446.

Chen:2011:TET

[365] Hubert J. Chen, Miin-Jye Wen, and Chia-Jui Chuang. On testing the equivalence of treatments using the measure of range. Computational Statistics & Data Analysis, 55(1):603– 614, January 1, 2011. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016794731000246X.

Gouno:2011:EAD

[366] E. Gouno, L. Courtrai, and M. Fredette. Estimation from aggregate data. Computational Statistics & Data Analysis, 55(1):615–626, January 1, 2011. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310002471.

Chadoeuf:2011:EIG

[367] J. Chadœuf, G. Certain, E. Bellier, A. Bar-Hen, P. Couteron, P. Monestiez, and V. Bretagnolle. Estimating inter-group interaction radius for point processes with nested spatial structures. Computational Statistics & Data Analysis, 55(1):627–640, January 1, 2011. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310002483.

Armagan:2011:NMF

[368] Artin Armagan and Russell L. Zaretzki. A note on mean-field variational approximations in Bayesian probit models. Computational Statistics & Data Analysis, 55(1):641–643, January 1, 2011. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310002495.

Sarhan:2011:BGL

[369] Ammar M. Sarhan, David C. Hamilton, Bruce Smith, and Debasis Kundu. The bivariate generalized linear failure rate distribution and its multivariate extension. Computational Statistics & Data Analysis, 55(1):644– 654, January 1, 2011. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947310002501.

Qian:2011:UCR

[370] Guoqi Qian, Ning Li, and Richard Huggins. Using capture-recapture data and hybrid Monte Carlo sampling to estimate an animal population affected by an environmental catastrophe. Computational Statistics & Data Analysis, 55 (1):655–666, January 1, 2011. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947310002537.

Devarajan:2011:SPG

[371] Karthik Devarajan and Nader Ebrahimi. A semi-parametric generalization of the Cox proportional hazards regression model: Inference and applications. Computational Statistics & Data Analysis, 55(1):667–676, January 1, 2011. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947310002549.

Cancho:2011:PEL

[372] Vicente G. Cancho, Franscisco Louzada-Neto, and Gladys D. C. Barriga. The Poisson-exponential lifetime distribution. Computational Statistics & Data Analysis, 55(1):677–686, January 1, 2011. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947310002550.

Wan:2011:BAR

[373] Wai-Yin Wan and Jennifer So-Kuen Chan. Bayesian analysis of robust Poisson geometric process model using heavy-tailed distributions. Computational Statistics & Data Analysis, 55 (1):687–702, January 1, 2011. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310002562.

Huang:2011:DRE

[374] Xianzheng Huang. Detecting randomeffects model misspecification via coarsened data. Computational Statistics & Data Analysis, 55(1):703–714, January 1, 2011. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310002574.

Langrock:2011:HMM

[375] R. Langrock and W. Zucchini. Hidden Markov models with arbitrary state dwell-time distributions. Computational Statistics & Data Analysis, 55 (1):715–724, January 1, 2011. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310002604.

Tolusso:2011:IER

[376] David Tolusso and Xikui Wang. Interval estimation for response adaptive clinical trials. Computational Statistics & Data Analysis, 55(1):725– 730, January 1, 2011. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947310002616.

Kang:2011:GFD

[377] Moonsu Kang and Heuiju Chun. A generalized false discovery rate in microarray studies. Computational Statistics & Data Analysis, 55(1):731– 737, January 1, 2011. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947310002628.

Nicolis:2011:WBS

[378] Orietta Nicolis, Pepa Ramírez-Cobo, and Brani Vidakovic. 2D wavelet-based spectra with applications. Computational Statistics & Data Analysis, 55 (1):738–751, January 1, 2011. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947310002653.

Prendergast:2011:NPI

[379] Luke A. Prendergast and Connie Li Wai Suen. A new and practical influence measure for subsets of covariance matrix sample principal components with applications to high dimensional datasets. Computational Statistics & Data Analysis, 55(1):752– 764, January 1, 2011. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947310002677.

Chen:2011:MSZ

[380] Xue-Dong Chen and Ying-Zi Fu. Model selection for zero-inflated regression with missing covariates. Computational Statistics & Data Analysis, 55 (1):765–773, January 1, 2011. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310002689.

Cizek:2011:SWR

[381] Pavel Čižek. Semiparametrically weighted robust estimation of regression models. Computational Statistics & Data Analysis, 55(1):774–788, January 1, 2011. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310002690.

Witten:2011:SMS

[382] Daniela M. Witten and Robert Tibshirani. Supervised multidimensional scaling for visualization, classification, and bipartite ranking. Computational Statistics & Data Analysis, 55(1):789– 801, January 1, 2011. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310002732.

Jung:2011:LCS

[383] Hyekyung Jung, Joseph L. Schafer, and Byungtae Seo. A latent class selection model for nonignorably missing data. Computational Statistics & Data Analysis, 55(1):802–812, January 1, 2011. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310002744.

Rahardja:2011:LBC

[384] Dewi Rahardja and Dean M. Young. Likelihood-based confidence intervals for the risk ratio using double sampling with over-reported binary data. Computational Statistics & Data Analysis, 55(1):813–823, January 1, 2011. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016794731000277X.

Lee:2011:AZI

[385] Keunbaik Lee, Yongsung Joo, Joon Jin Song, and Dee Wood Harper. Analysis of zero-inflated clustered count data: a marginalized model approach. Computational Statistics & Data Analysis, 55 (1):824–837, January 1, 2011. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947310002781.

Martinez:2011:PSC

[386] S. Martínez, M. Rueda, A. Arcos, H. Martínez, and I. Sánchez-Borrego. Post-stratified calibration method for estimating quantiles. Computational Statistics & Data Analysis, 55(1):838– 851, January 1, 2011. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947310002793.

Wang:2011:SVM

[387] Joanna J. J. Wang, Jennifer S. K. Chan, and S. T. Boris Choy. Stochastic volatility models with leverage and heavy-tailed distributions: a Bayesian approach using scale mixtures. Computational Statistics & Data Analysis, 55 (1):852–862, January 1, 2011. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310002811.

Martinez-Rodriguez:2011:MUE

[388] A. M. Martínez-Rodríguez, A. J. Sáez-Castillo, and A. Conde-Sánchez. Modelling using an extended Yule distribution. Computational Statistics & Data Analysis, 55(1):863–873, January 1, 2011. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310002872.

Locatelli:2011:RAF

[389] Isabella Locatelli, Alfio Marazzi, and Victor J. Yohai. Robust accelerated failure time regression. Computational Statistics & Data Analysis, 55 (1):874–887, January 1, 2011. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310002963.

Rodriguez-Alvarez:2011:CSR

[390] María Xosé Rodríguez-Álvarez, Pablo G. Tahoces, Carmen Cadarso-Suárez, and María José Lado. Comparative study of ROC regression techniques — applications for the computeraided diagnostic system in breast cancer detection. Computational Statistics & Data Analysis, 55(1):888–902, January 1, 2011. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947310002975.

Lin:2011:FPS

[391] Wan-Yu Lin and Wen-Chung Lee. Floating prioritized subset analysis: a powerful method to detect differentially expressed genes. Computational Statistics & Data Analysis, 55 (1):903–913, January 1, 2011. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947310003105.

Hwang:2011:JML

[392] Yi-Ting Hwang, Hao-Yun Tsai, Yeu-Jhy Chang, Hsun-Chih Kuo, and Chun-Chao Wang. The joint model of the logistic model and linear random effect model — an application to predict orthostatic hypertension for subacute stroke patients. Computational Statistics & Data Analysis, 55 (1):914–923, January 1, 2011. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947310003117.

Oh:2011:UBI

[393] Man-Suk Oh and Dong Wan Shin. A unified Bayesian inference on treatment means with order constraints. Computational Statistics & Data Analysis, 55 (1):924–934, January 1, 2011. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947310003130.

Min:2011:SAH

[394] Jeong Eun Min, Matthew D. Whiteside, Fiona S. L. Brinkman, Brad Mc-Neney, and Jinko Graham. A statistical approach to high-throughput screening of predicted orthologs. Computational Statistics & Data Analysis, 55 (1):935–943, January 1, 2011. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310003178.

Passos:2011:CEC

[395] Valéria Lima Passos, Frans E. S. Tan, and Martijn P. F. Berger. Costefficiency considerations in the choice of a microarray platform for time course experimental designs. Computational Statistics & Data Analysis, 55 (1):944–954, January 1, 2011. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310002410.

Anonymous:2011:EBe

[396] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 55 (2):iii–v, February 1, 2011. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310003865.

Anonymous:2011:Ce

[397] Anonymous. Contents. Computational Statistics & Data Analysis, 55 (2):vi–vii, February 1, 2011. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310003877.

Tang:2011:AUT

[398] Man-Lai Tang, Wai-Yin Poon, Leevan Ling, Yijie Liao, and Hang-Wai Chui. Approximate unconditional test procedure for comparing two ordered multinomials. Computational Statistics & Data Analysis, 55(2):955– 963, February 1, 2011. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947310003245.

Al-Mutairi:2011:ERS

[399] D. K. Al-Mutairi, M. E. Ghitany, and Ramesh C. Gupta. Estimation of reliability in a series system with random sample size. Computational Statistics & Data Analysis, 55(2):964– 972, February 1, 2011. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947310003142.

Shi:2011:SLI

[400] Lei Shi and Mei Huang. Stepwise local influence analysis. Computational Statistics & Data Analysis, 55(2):973– 982, February 1, 2011. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947310003166.

Mao:2011:JEM

[401] Jie Mao, Zhongyi Zhu, and Wing K. Fung. Joint estimation of meancovariance model for longitudinal data with basis function approximations. Computational Statistics & Data Analysis, 55(2):983–992, February 1, 2011. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794731000318X.

Hashimoto:2011:RMG

[402] Elizabeth M. Hashimoto, Edwin M. M. Ortega, Gilberto A. Paula, and Mauricio L. Barreto. Regression models for grouped survival data: Estimation and sensitivity analysis. Computational Statistics & Data Analysis, 55(2):993– 1007, February 1, 2011. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310003191.

Kascha:2011:BLR

[403] Christian Kascha and Carsten Trenkler. Bootstrapping the likelihood ratio cointegration test in error correction models with unknown lag order. Computational Statistics & Data Analysis, 55(2):1008– 1017, February 1, 2011. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310003208.

Burch:2011:APN

[404] Brent D. Burch. Assessing the performance of normal-based and REML-based confidence intervals for the intraclass correlation coefficient. Computational Statistics & Data Analysis, 55(2):1018–1028, February 1, 2011. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947310003221.

Jang:2011:ESA

[405] Dongik Jang and Hee-Seok Oh. Enhancement of spatially adaptive smoothing splines via parameterization of smoothing parameters. Computational Statistics & Data Analysis, 55(2):1029–1040, February 1, 2011. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947310003233.

Takane:2011:RRR

[406] Yoshio Takane, Kwanghee Jung, and Heungsun Hwang. Regularized reduced rank growth curve models. Computational Statistics & Data Analysis, 55(2):1041–1052, February 1, 2011. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947310003270.

Wen:2011:NML

[407] Chi-Chung Wen and Yi-Hau Chen. Nonparametric maximum likelihood analysis of clustered current status data with the gamma-frailty Cox model. Computational Statistics & Data Analysis, 55(2):1053– 1060, February 1, 2011. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947310003282.

Yang:2011:FEM

[408] Jingyun Yang and Vernon M. Chinchilli. Fixed-effects modeling of Cohen’s weighted kappa for bivariate multinomial data. Computational Statistics & Data Analysis, 55(2):1061– 1070, February 1, 2011. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310003427.

Pryseley:2011:ENV

[409] Assam Pryseley, Clotaire Tchonlafi, Geert Verbeke, and Geert Molenberghs. Estimating negative variance components from Gaussian and non-Gaussian data: a mixed models approach. Computational Statistics & Data Analysis, 55(2):1071– 1085, February 1, 2011. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310003452.

Mandal:2011:GFT

[410] A. Mandal, W. T. Huang, S. K. Bhandari, and A. Basu. Goodnessof-fit testing in growth curve models: a general approach based on finite differences. Computational Statistics & Data Analysis, 55(2):1086– 1098, February 1, 2011. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310003464.

Li:2011:ESS

[411] Phillip Li. Estimation of sample selection models with two selection mechanisms. Computational Statistics & Data Analysis, 55(2):1099– 1108, February 1, 2011. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794731000349X.

Lemonte:2011:SPP

[412] Artur J. Lemonte and Silvia L. P. Ferrari. Size and power properties of some tests in the Birnbaum–Saunders regression model. Computational Statistics & Data Analysis, 55(2):1109– 1117, February 1, 2011. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310003518.

Paranaiba:2011:BBX

[413] Patrícia F. Paranaíba, Edwin M. M. Ortega, Gauss M. Cordeiro, and Rodrigo R. Pescim. The beta Burr XII distribution with application to lifetime data. Computational Statistics & Data Analysis, 55(2):1118– 1136, February 1, 2011. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794731000352X.

Cui:2011:NRE

[414] Hengjian Cui and Tao Hu. On nonlinear regression estimator with denoised variables. Computational Statistics & Data Analysis, 55(2):1137– 1149, February 1, 2011. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310003543.

Brentnall:2011:ARM

[415] Adam R. Brentnall, Martin J. Crowder, and David J. Hand. Approximate repeated-measures shrinkage. Computational Statistics & Data Analysis, 55(2):1150–1159, February 1, 2011. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0167947310003579.

Picard:2011:JSM

[416] F. Picard, E. Lebarbier, E. Budinskà, and S. Robin. Joint segmentation of multivariate Gaussian processes using mixed linear models. Computational Statistics & Data Analysis, 55(2):1160– 1170, February 1, 2011. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310003580.

Ren:2011:C

[417] Dianxu Ren. Corrigendum. Computational Statistics & Data Analysis, 55 (2):1171, February 1, 2011. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310003932.

Anonymous:2011:EBf

[418] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 55(3):iii–v, March 1, 2011. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310004226.

Anonymous:2011:Cf

[419] Anonymous. Contents. Computational Statistics & Data Analysis, 55 (3):vi–vii, March 1, 2011. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310004238.

Zuo:2011:ECB

[420] Yijun Zuo and Shaoyong Lai. Exact computation of bivariate projection depth and the Stahel–Donoho estimator. Computational Statistics & Data Analysis, 55(3):1173– 1179, March 1, 2011. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310003531.

Ma:2011:IDE

[421] Jun Ma. Indirect density estimation using the iterative Bayes algorithm. Computational Statistics & Data Analysis, 55(3):1180– 1195, March 1, 2011. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310003610.

Zhang:2011:EML

[422] Ran Zhang, Claudia Czado, and Aleksey Min. Efficient maximum likelihood estimation of copula based meta t-distributions. Computational Statistics & Data Analysis, 55(3):1196– 1214, March 1, 2011. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310003701.

Eugster:2011:WRA

[423] Manuel J. A. Eugster and Friedrich Leisch. Weighted and robust archetypal analysis. Computational Statistics & Data Analysis, 55(3):1215– 1225, March 1, 2011. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310004056.

Shoukri:2011:ADP

[424] Mohamed M. Shoukri, Pranesh Kumar, and Dilek Colak. Analyzing dependent proportions in cluster randomized trials: Modeling inter-cluster correlation via copula function. Computational Statistics & Data Analysis, 55 (3):1226–1235, March 1, 2011. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310003257.

Alpargu:2011:AMG

[425] Gülhan Alpargu. Allowing for missing genotypes in any members of the nuclear families in transmission disequilibrium test. Computational Statistics & Data Analysis, 55(3):1236– 1249, March 1, 2011. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310003476.

Barriga:2011:CEP

[426] Gladys D. C. Barriga, Franscisco Louzada-Neto, and Vicente G. Cancho. The complementary exponential power lifetime model. Computational Statistics & Data Analysis, 55 (3):1250–1259, March 1, 2011. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310003488.

Cai:2011:BSM

[427] Bo Cai and Renate Meyer. Bayesian semiparametric modeling of survival data based on mixtures of B-spline distributions. Computational Statistics & Data Analysis, 55(3):1260– 1272, March 1, 2011. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310003506.

Tang:2011:TER

[428] Nian-Sheng Tang and Shao-Ping Jiang. Testing equality of risk ratios in multiple 2 × 2 tables with structural zero. Computational Statistics & Data Analysis, 55(3):1273– 1284, March 1, 2011. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310003592.

Cramer:2011:PTI

[429] Erhard Cramer and Anja Bettina Schmiedt. Progressively Type-II censored competing risks data from Lomax distributions. Computational Statistics & Data Analysis, 55(3):1285– 1303, March 1, 2011. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310003609.

Garay:2011:EID

[430] Aldo M. Garay, Elizabeth M. Hashimoto, Edwin M. M. Ortega, and Víctor H. Lachos. On estimation and influence diagnostics for zero-inflated negative binomial regression models. Computational Statistics & Data Analysis, 55 (3):1304–1318, March 1, 2011. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310003622.

Lopes:2011:GSD

[431] Hedibert Freitas Lopes, Dani Gamerman, and Esther Salazar. Generalized spatial dynamic factor models. Computational Statistics & Data Analysis, 55 (3):1319–1330, March 1, 2011. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310003634.

Magnus:2011:WAL

[432] Jan R. Magnus, Alan T. K. Wan, and Xinyu Zhang. Weighted average least squares estimation with nonspherical disturbances and an application to the Hong Kong housing market. Computational Statistics & Data Analysis, 55 (3):1331–1341, March 1, 2011. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794731000366X.

Chakraborty:2011:BHH

[433] Sounak Chakraborty and Ruixin Guo. A Bayesian hybrid Huberized support vector machine and its applications in high-dimensional medical data. Computational Statistics & Data Analysis, 55(3):1342– 1356, March 1, 2011. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310003671.

Kim:2011:PCC

[434] Jong-Min Kim, Yoon-Sung Jung, Taeryon Choi, and Engin A. Sungur. Partial correlation with copula modeling. Computational Statistics & Data Analysis, 55(3):1357– 1366, March 1, 2011. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310003683.

Lin:2011:SBT

[435] Lu Lin, Qi Zhang, Feng Li, and Xia Cui. Simulation-based two-stage estimation for multiple nonparametric regression. Computational Statistics & Data Analysis, 55(3):1367– 1378, March 1, 2011. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310003695.

Barreto-Souza:2011:BSG

[436] Wagner Barreto-Souza and Klaus L. P. Vasconcellos. Bias and skewness in a general extreme-value regression model. Computational Statistics & Data Analysis, 55(3):1379– 1393, March 1, 2011. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310003713.

Pianto:2011:DML

[437] Donald M. Pianto and Francisco Cribari-Neto. Dealing with monotone likelihood in a model for speckled data. Computational Statistics & Data Analysis, 55(3):1394– 1409, March 1, 2011. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310003725.

Morais:2011:CCW

[438] Alice Lemos Morais and Wagner Barreto-Souza. A compound class of Weibull and power series distributions. Computational Statistics & Data Analysis, 55(3):1410– 1425, March 1, 2011. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310003737.

Picchini:2011:PEH

[439] Umberto Picchini and Susanne Ditlevsen. Practical estimation of high dimensional stochastic differential mixedeffects models. Computational Statistics & Data Analysis, 55(3):1426– 1444, March 1, 2011. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310003774.

Cordeiro:2011:BSD

[440] Gauss M. Cordeiro and Artur J. Lemonte. The β-Birnbaum–Saunders distribution: an improved distribution for fatigue life modeling. Computational Statistics & Data Analysis, 55 (3):1445–1461, March 1, 2011. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947310003816.

Ibacache-Pulgar:2011:LIS

[441] Germán Ibacache-Pulgar and Gilberto A. Paula. Local influence for Studentt partially linear models. Computational Statistics & Data Analysis, 55 (3):1462–1478, March 1, 2011. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016794731000397X.

Zou:2011:SAF

[442] Yubo Zou, Jiajia Zhang, and Guoyou Qin. A semiparametric accelerated failure time partial linear model and its application to breast cancer. Computational Statistics & Data Analysis, 55 (3):1479–1487, March 1, 2011. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947310004007.

Lu:2011:LGC

[443] Tong-Yu Lu, Wai-Yin Poon, and Yim-Fan Tsang. Latent growth curve modeling for longitudinal ordinal responses with applications. Computational Statistics & Data Analysis, 55 (3):1488–1497, March 1, 2011. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947310004020.

Liu:2011:BAP

[444] Rui-Yin Liu, Jian Tao, Ning-Zhong Shi, and Xuming He. Bayesian analysis of the patterns of biological susceptibility via reversible jump MCMC sampling. Computational Statistics & Data Analysis, 55(3):1498– 1508, March 1, 2011. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310004044.

Anonymous:2011:EBg

[445] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 55(4):iii–v, April 1, 2011. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310004664.

Anonymous:2011:Cg

[446] Anonymous. Contents. Computational Statistics & Data Analysis, 55 (4):vi–vii, April 1, 2011. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310004676.

Denman:2011:DEB

[447] N. G. Denman, J. M. McGree, J. A. Eccleston, and S. B. Duffull. Design of experiments for bivariate binary responses modelled by copula functions. Computational Statistics & Data Analysis, 55(4):1509– 1520, April 1, 2011. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310003129.

Phinikettos:2011:FCH

[448] Ioannis Phinikettos and Axel Gandy. Fast computation of high-dimensional multivariate normal probabilities. Computational Statistics & Data Analysis, 55(4):1521–1529, April 1, 2011. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0167947310003798.

Wright:2011:OSM

[449] F. T. Wright and Kane Nashimoto. One-sided multiple comparisons for treatment means with a control mean. Computational Statistics & Data Analysis, 55(4):1530–1539, April 1, 2011. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947310004019.

Strasak:2011:CPS

[450] Alexander M. Strasak, Nikolaus Umlauf, Ruth M. Pfeiffer, and Stefan Lang. Comparing penalized splines and fractional polynomials for flexible modelling of the effects of continuous predictor variables. Computational Statistics & Data Analysis, 55 (4):1540–1551, April 1, 2011. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016794731000407X.

Ye:2011:SBM

[451] Gui-Bo Ye and Xiaohui Xie. Split Bregman method for large scale fused lasso. Computational Statistics & Data Analysis, 55(4):1552– 1569, April 1, 2011. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310004093.

Chen:2011:APD

[452] Min Chen and Xinlei Wang. Approximate predictive densities and their applications in generalized linear models. Computational Statistics & Data Analysis, 55(4):1570– 1580, April 1, 2011. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310004330.

Ho:2011:UPP

[453] Man-Wai Ho. Usage of a pair of Spaths in Bayesian estimation of a unimodal density. Computational Statistics & Data Analysis, 55(4):1581– 1595, April 1, 2011. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310004354.

Wang:2011:ELC

[454] Qihua Wang and Peng Lai. Empirical likelihood calibration estimation for the median treatment difference in observational studies. Computational Statistics & Data Analysis, 55 (4):1596–1609, April 1, 2011. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310003555.

Lin:2011:REE

[455] Feng-Chang Lin. A random effects epidemic-type aftershock sequence model. Computational Statistics & Data Analysis, 55(4):1610– 1616, April 1, 2011. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310003749.

Su:2011:ASC

[456] Pei-Fang Su, Yunchan Chi, Chung-I Li, Yu Shyr, and Yi-De Liao. Analyzing survival curves at a fixed point in time for paired and clustered right-censored data. Computational Statistics & Data Analysis, 55 (4):1617–1628, April 1, 2011. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310003750.

Singh:2011:SB

[457] Sarjinder Singh and Stephen A. Sedory. Sufficient bootstrapping. Computational Statistics & Data Analysis, 55 (4):1629–1637, April 1, 2011. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310003981.

Reichert:2011:MBE

[458] P. Reichert, G. White, M. J. Bayarri, and E. B. Pitman. Mechanismbased emulation of dynamic simulation models: Concept and application in hydrology. Computational Statistics & Data Analysis, 55(4):1638– 1655, April 1, 2011. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310003993.

Wang:2011:EIM

[459] Qin Wang and Xiangrong Yin. Estimation of inverse mean: an orthogonal series approach. Computational Statistics & Data Analysis, 55 (4):1656–1664, April 1, 2011. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794731000410X.

Vilca:2011:EEP

[460] Filidor Vilca, Lucia Santana, Víctor Leiva, and N. Balakrishnan. Estimation of extreme percentiles in Birnbaum–Saunders distributions. Computational Statistics & Data Analysis, 55(4):1665–1678, April 1, 2011. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0167947310004111.

Lopez-Pintado:2011:HRD

[461] Sara López-Pintado and Juan Romo. A half-region depth for functional data. Computational Statistics & Data Analysis, 55(4):1679–1695, April 1, 2011. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310004123.

Shen:2011:RLI

[462] Chung-Wei Shen, Tsung-Shan Tsou, and N. Balakrishnan. Robust likelihood inference for regression parameters in partially linear models. Computational Statistics & Data Analysis, 55 (4):1696–1714, April 1, 2011. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310004135.

Ingrassia:2011:DEA

[463] Salvatore Ingrassia and Roberto Rocci. Degeneracy of the EM algorithm for the MLE of multivariate Gaussian mixtures and dynamic constraints. Computational Statistics & Data Analysis, 55 (4):1715–1725, April 1, 2011. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310004147.

Fang:2011:EMM

[464] Hong-Bin Fang, Jiantian Wang, Dianliang Deng, and Man-Lai Tang. Estimating the mean of a mark variable under right censoring on the basis of a state function. Computational Statistics & Data Analysis, 55 (4):1726–1735, April 1, 2011. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310004160.

Fabrizi:2011:HBM

[465] Enrico Fabrizi, Maria Rosaria Ferrante, Silvia Pacei, and Carlo Trivisano. Hierarchical Bayes multivariate estimation of poverty rates based on increasing thresholds for small domains. Computational Statistics & Data Analysis, 55 (4):1736–1747, April 1, 2011. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310004299.

Baghishani:2011:DCA

[466] Hossein Baghishani and Mohsen Mohammadzadeh. A data cloning algorithm for computing maximum likelihood estimates in spatial generalized linear mixed models. Computational Statistics & Data Analysis, 55 (4):1748–1759, April 1, 2011. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310004329.

Chen:2011:GEN

[467] Shyh-Huei Chen, Edward H. Ip, and Yuchung J. Wang. Gibbs ensembles for nearly compatible and incompatible conditional models. Computational Statistics & Data Analysis, 55 (4):1760–1769, April 1, 2011. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310004342.

Arpino:2011:SPS

[468] Bruno Arpino and Fabrizia Mealli. The specification of the propensity score in multilevel observational studies. Computational Statistics & Data Analysis, 55(4):1770– 1780, April 1, 2011. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310004366.

Sinn:2011:EOP

[469] Mathieu Sinn and Karsten Keller. Estimation of ordinal pattern probabilities in Gaussian processes with stationary increments. Computational Statistics & Data Analysis, 55(4):1781– 1790, April 1, 2011. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310004378.

Hosseini:2011:ABI

[470] Fatemeh Hosseini, Jo Eidsvik, and Mohsen Mohammadzadeh. Approximate Bayesian inference in spatial GLMM with skew normal latent variables. Computational Statistics & Data Analysis, 55(4):1791– 1806, April 1, 2011. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310004391.

Drikvandi:2011:BTS

[471] Reza Drikvandi, Reza Modarres, and Abdullah H. Jalilian. A bootstrap test for symmetry based on ranked set samples. Computational Statistics & Data Analysis, 55(4):1807– 1814, April 1, 2011. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310004408.

Rukhin:2011:LRE

[472] Andrew L. Rukhin and Antonio Possolo. Laplace random effects models for interlaboratory studies. Computational Statistics & Data Analysis, 55 (4):1815–1827, April 1, 2011. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310004445.

Airola:2011:ECC

[473] Antti Airola, Tapio Pahikkala, Willem Waegeman, Bernard De Baets, and Tapio Salakoski. An experimental comparison of cross-validation techniques for estimating the area under the ROC curve. Computational Statistics & Data Analysis, 55(4):1828– 1844, April 1, 2011. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310004469.

Cheng:2011:RDH

[474] Tsung-Chi Cheng. Robust diagnostics for the heteroscedastic regression model. Computational Statistics & Data Analysis, 55(4):1845– 1866, April 1, 2011. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310004524.

Venezuela:2011:LIE

[475] Maria Kelly Venezuela, Mônica Carneiro Sandoval, and Denise Aparecida Botter. Local influence in estimating equations. Computational Statistics & Data Analysis, 55(4):1867– 1883, April 1, 2011. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310004081.

Kim:2011:CSN

[476] Young-Ju Kim. A comparative study of nonparametric estimation in Weibull regression: a penalized likelihood approach. Computational Statistics & Data Analysis, 55(4):1884– 1896, April 1, 2011. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947310004172.

Anonymous:2011:EBh

[477] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 55(5):iii–v, May 1, 2011. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947311000156.

Anonymous:2011:Ch

[478] Anonymous. Contents. Computational Statistics & Data Analysis, 55 (5):vi, May 1, 2011. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947311000168.

Choi:2011:GSP

[479] Hosik Choi, Donghwa Yeo, Sunghoon Kwon, and Yongdai Kim. Gene selection and prediction for cancer classification using support vector machines with a reject option. Computational Statistics & Data Analysis, 55 (5):1897–1908, May 1, 2011. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947310004573.

Fisher:2011:IST

[480] Thomas J. Fisher and Xiaoqian Sun. Improved Stein-type shrinkage estimators for the high-dimensional multivariate normal covariance matrix. Computational Statistics & Data Analysis, 55 (5):1909–1918, May 1, 2011. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310004743.

Gijbels:2011:CCA

[481] Irène Gijbels, Noël Veraverbeke, and Marel Omelka. Conditional copulas, association measures and their applications. Computational Statistics & Data Analysis, 55(5):1919– 1932, May 1, 2011. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794731000438X.

Adler:2011:ECP

[482] Werner Adler, Alexander Brenning, Sergej Potapov, Matthias Schmid, and Berthold Lausen. Ensemble classification of paired data. Computational Statistics & Data Analysis, 55 (5):1933–1941, May 1, 2011. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310004457.

Dettmann:2011:DFM

[483] E. Dettmann, C. Becker, and C. Schmeißer. Distance functions for matching in small samples. Computational Statistics & Data Analysis, 55(5):1942– 1960, May 1, 2011. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310004500.

Dosse:2011:AGP

[484] Mohammed Bennani Dosse, Henk A. L. Kiers, and Jos M. F. Ten Berge. Anisotropic generalized Procrustes analysis. Computational Statistics & Data Analysis, 55(5):1961– 1968, May 1, 2011. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016794731000455X.

Wu:2011:BVI

[485] Han-Ming Wu. On biological validity indices for soft clustering algorithms for gene expression data. Computational Statistics & Data Analysis, 55 (5):1969–1979, May 1, 2011. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947310004597.

Tenreiro:2011:AIM

[486] Carlos Tenreiro. An affine invariant multiple test procedure for assessing multivariate normality. Computational Statistics & Data Analysis, 55 (5):1980–1992, May 1, 2011. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947310004603.

Li:2011:CSM

[487] Xinmin Li, Juan Wang, and Hua Liang. Comparison of several means: a fiducial based approach. Computational Statistics & Data Analysis, 55 (5):1993–2002, May 1, 2011. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310004779.

Lee:2011:ACV

[488] Pei-Hsi Lee. Adaptive R charts with variable parameters. Computational Statistics & Data Analysis, 55(5): 2003–2010, May 1, 2011. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310004548.

Anonymous:2011:EBi

[489] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 55(6):iii–v, June 1, 2011. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947311000612.

Anonymous:2011:Ci

[490] Anonymous. Contents. Computational Statistics & Data Analysis, 55 (6):vi–vii, June 1, 2011. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947311000624.

Hatjispyros:2011:DMD

[491] Spyridon J. Hatjispyros, Theodoros Nicoleris, and Stephen G. Walker. Dependent mixtures of Dirichlet processes. Computational Statistics & Data Analysis, 55(6):2011– 2025, June 1, 2011. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310004615.

Crews:2011:FMS

[492] Hugh B. Crews, Dennis D. Boos, and Leonard A. Stefanski. FSR methods for second-order regression models. Computational Statistics & Data Analysis, 55(6):2026– 2037, June 1, 2011. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947311000259.

Hernandez-Lobato:2011:SBA

[493] José Miguel Hernández-Lobato and Alberto Suárez. Semiparametric bivariate Archimedean copulas. Computational Statistics & Data Analysis, 55 (6):2038–2058, June 1, 2011. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947311000375.

Matsuo:2011:NCM

[494] Tomoko Matsuo, Douglas W. Nychka, and Debashis Paul. Nonstationary covariance modeling for incomplete data: Monte Carlo EM approach. Computational Statistics & Data Analysis, 55 (6):2059–2073, June 1, 2011. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310004585.

da-Silva:2011:DBB

[495] C. Q. da Silva, H. S. Migon, and L. T. Correia. Dynamic Bayesian beta models. Computational Statistics & Data Analysis, 55(6):2074– 2089, June 1, 2011. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310004792.

Li:2011:ICN

[496] Pai-Ling Li and Jeng-Min Chiou. Identifying cluster number for subspace projected functional data clustering. Computational Statistics & Data Analysis, 55(6):2090–2103, June 1, 2011. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794731100003X.

Tomita:2011:ASR

[497] Makoto Tomita, Noboru Hashimoto, and Yutaka Tanaka. Association study for the relationship between a haplotype or haplotype set and multiple quantitative responses. Computational Statistics & Data Analysis, 55 (6):2104–2113, June 1, 2011. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947311000041.

Zhou:2011:MLR

[498] Jie Zhou. Maximum likelihood ratio test for the stability of sequence of Gaussian random processes. Computational Statistics & Data Analysis, 55 (6):2114–2127, June 1, 2011. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947311000053.

Fang:2011:PCA

[499] Yixin Fang and Junhui Wang. Penalized cluster analysis with applications to family data. Computational Statistics & Data Analysis, 55 (6):2128–2136, June 1, 2011. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947311000065.

MacDonald:2011:FEV

[500] A. MacDonald, C. J. Scarrott, D. Lee, B. Darlow, M. Reale, and G. Russell. A flexible extreme value mixture model. Computational Statistics & Data Analysis, 55(6):2137– 2157, June 1, 2011. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947311000077.

Martinez:2011:AAC

[501] Z. Martínez and C. Ludeña. An algorithm for automatic curve detection. Computational Statistics & Data Analysis, 55(6):2158– 2171, June 1, 2011. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947311000107.

Usta:2011:PFM

[502] Ilhan Usta and Yeliz Mert Kantar. On the performance of the flexible maximum entropy distributions within partially adaptive estimation. Computational Statistics & Data Analysis, 55 (6):2172–2182, June 1, 2011. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947311000260.

Jin:2011:DDT

[503] Lei Jin. A data-driven test to compare two or multiple time series. Computational Statistics & Data Analysis, 55 (6):2183–2196, June 1, 2011. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947311000296.

Kanatani:2011:HLS

[504] Kenichi Kanatani and Prasanna Rangarajan. Hyper least squares fitting of circles and ellipses. Computational Statistics & Data Analysis, 55 (6):2197–2208, June 1, 2011. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310004809.

Anonymous:2011:EBj

[505] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 55(7):iii–v, July 1, 2011. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947311000892.

Anonymous:2011:Cj

[506] Anonymous. Contents. Computational Statistics & Data Analysis, 55 (7):vi–vii, July 1, 2011. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947311000909.

Cronie:2011:SEC

[507] Ottmar Cronie and Aila Särkkä. Some edge correction methods for marked spatio-temporal point process models. Computational Statistics & Data Analysis, 55(7):2209– 2220, July 1, 2011. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947311000417.

Shen:2011:SBR

[508] Yanfeng Shen, Zhengyan Lin, and Jun Zhu. Shrinkage-based regularization tests for high-dimensional data with application to gene set analysis. Computational Statistics & Data Analysis, 55(7):2221–2233, July 1, 2011. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947311000028.

Zhang:2011:FDS

[509] Zhen Zhang and Hans-Georg Müller. Functional density synchronization. Computational Statistics & Data Analysis, 55(7):2234–2249, July 1, 2011. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0167947311000090.

Kim:2011:PCP

[510] Jaejik Kim and L. Billard. A polythetic clustering process and cluster validity indexes for histogramvalued objects. Computational Statistics & Data Analysis, 55(7):2250– 2262, July 1, 2011. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947311000272.

Foss:2011:TSE

[511] Tron Foss, Karl G. Jöreskog, and Ulf H. Olsson. Testing structural equation models: the effect of kurtosis. Computational Statistics & Data Analysis, 55 (7):2263–2275, July 1, 2011. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947311000284.

Chen:2011:CSR

[512] Cathy W. S. Chen, Jennifer S. K. Chan, Mike K. P. So, and Kevin K. M. Lee. Classification in segmented regression problems. Computational Statistics & Data Analysis, 55 (7):2276–2287, July 1, 2011. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947311000314.

Kolossiatis:2011:MON

[513] M. Kolossiatis, J. E. Griffin, and M. F. J. Steel. Modeling overdispersion with the normalized tempered stable distribution. Computational Statistics & Data Analysis, 55(7): 2288–2301, July 1, 2011. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947311000351.

Lanumteang:2011:ECE

[514] K. Lanumteang and D. Böhning. An extension of Chao’s estimator of population size based on the first three capture frequency counts. Computational Statistics & Data Analysis, 55 (7):2302–2311, July 1, 2011. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947311000363.

Zhu:2011:EEV

[515] Rong Zhu and Sherry Z. F. Zhou. Estimating the error variance after a pre-test for an interval restriction on the coefficients. Computational Statistics & Data Analysis, 55(7): 2312–2323, July 1, 2011. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947311000387.

Xu:2011:BAB

[516] Ancha Xu and Yincai Tang. Bayesian analysis of Birnbaum–Saunders distribution with partial information. Computational Statistics & Data Analysis, 55(7):2324–2333, July 1, 2011. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947311000405.

Wang:2011:RRA

[517] You-Gan Wang and Liya Fu. Rank regression for accelerated failure time model with clustered and censored data. Computational Statistics & Data Analysis, 55(7):2334– 2343, July 1, 2011. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947311000429.

Mukhopadhyay:2011:BMS

[518] Subhadeep Mukhopadhyay and Anil K. Ghosh. Bayesian multiscale smoothing in supervised and semi-supervised kernel discriminant analysis. Computational Statistics & Data Analysis, 55 (7):2344–2353, July 1, 2011. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947311000466.

Tsai:2011:GEE

[519] Miao-Yu Tsai, Jung-Feng Wang, and Jia-Ling Wu. Generalized estimating equations with model selection for comparing dependent categorical agreement data. Computational Statistics & Data Analysis, 55(7):2354– 2362, July 1, 2011. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947311000478.

Li:2011:QNB

[520] Shubiao Li, Fang Yang, Felix Famoye, Carl Lee, and Dennis Black. Quasinegative binomial distribution: Properties and applications. Computational Statistics & Data Analysis, 55 (7):2363–2371, July 1, 2011. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794731100048X.

Marra:2011:PVS

[521] Giampiero Marra and Simon N. Wood. Practical variable selection for generalized additive models. Computational Statistics & Data Analysis, 55 (7):2372–2387, July 1, 2011. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947311000491.

Lemonte:2011:THB

[522] Artur J. Lemonte and Silvia L. P. Ferrari. Testing hypotheses in the Birnbaum–Saunders distribution under type-II censored samples. Computational Statistics & Data Analysis, 55 (7):2388–2399, July 1, 2011. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947311000508.

Darilay:2011:PUL

[523] Annie Tordilla Darilay and Joshua D. Naranjo. A pretest for using logrank or Wilcoxon in the two-sample problem. Computational Statistics & Data Analysis, 55(7):2400– 2409, July 1, 2011. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016794731100051X.

Ferrari:2011:IMC

[524] Pier Alda Ferrari, Paola Annoni, Alessandro Barbiero, and Giancarlo Manzi. An imputation method for categorical variables with application to nonlinear principal component analysis. Computational Statistics & Data Analysis, 55(7):2410– 2420, July 1, 2011. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947311000521.

Park:2011:EEN

[525] DoHwan Park, Junyong Park, Xiaosong Zhong, and Michel Sadelain. Estimation of empirical null using a mixture of normals and its use in local false discovery rate. Computational Statistics & Data Analysis, 55 (7):2421–2432, July 1, 2011. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947311000545.

Davidov:2011:CVU

[526] Ori Davidov. Combining p-values using order-based methods. Computational Statistics & Data Analysis, 55 (7):2433–2444, July 1, 2011. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947311000430.

Ospina:2011:EIP

[527] Raydonal Ospina, Francisco Cribari-Neto, and Klaus L. P. Vasconcellos. Erratum to “Improved point and interval estimation for a beta regression model” [comput. statist. data anal. 51 (2006) 960–981]. Computational Statistics & Data Analysis, 55(7):2445, July 1, 2011. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947311000740. See

[1].

Lawless:2011:CSM

[528] Jerald F. Lawless and Yildiz E. Yilmaz. Comparison of semiparametric maximum likelihood estimation and two-stage semiparametric estimation in copula models. Computational Statistics & Data Analysis, 55 (7):2446–2455, July 1, 2011. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947311000533.

Anonymous:2011:EBk

[529] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 55(8):iii–v, August 1, 2011. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947311001204.

Anonymous:2011:Ck

[530] Anonymous. Contents. Computational Statistics & Data Analysis, 55 (8):vi, August 1, 2011. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947311001216.

Kuo:2011:SAC

[531] Kun-Lin Kuo and Yuchung J. Wang. A simple algorithm for checking compatibility among discrete conditional distributions. Computational Statistics & Data Analysis, 55(8):2457– 2462, August 1, 2011. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947311000764.

Sysoev:2011:SBA

[532] O. Sysoev, O. Burdakov, and A. Grimvall. A segmentation-based algorithm for large-scale partially ordered monotonic regression. Computational Statistics & Data Analysis, 55(8):2463– 2476, August 1, 2011. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947311000806.

Kim:2011:IIE

[533] Jae H. Kim, Iain Fraser, and Rob J. Hyndman. Improved interval estimation of long run response from a dynamic linear model: a highest density region approach. Computational Statistics & Data Analysis, 55(8):2477– 2489, August 1, 2011. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794731100082X.

Graham:2011:NEW

[534] M. A. Graham, S. Chakraborti, and S. W. Human. A nonparametric exponentially weighted moving average signed-rank chart for monitoring location. Computational Statistics & Data Analysis, 55(8):2490– 2503, August 1, 2011. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947311000727.

Koller:2011:SWT

[535] Manuel Koller and Werner A. Stahel. Sharpening Wald-type inference in robust regression for small samples. Computational Statistics & Data Analysis, 55(8):2504–2515, August 1, 2011. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947311000739.

Louzada:2011:CEG

[536] Francisco Louzada, Mari Roman, and Vicente G. Cancho. The complementary exponential geometric distribution: Model, properties, and a comparison with its counterpart. Computational Statistics & Data Analysis, 55(8):2516–2524, August 1, 2011. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947311000776. See

note [1447].

Taufer:2011:CFE

[537] Emanuele Taufer, Nikolai Leonenko, and Marco Bee. Characteristic function estimation of Ornstein– Uhlenbeck-based stochastic volatility models. Computational Statistics & Data Analysis, 55(8):2525– 2539, August 1, 2011. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947311000752.

Anonymous:2011:EBl

[538] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 55 (9):iii–v, September 1, 2011. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947311001575.

Anonymous:2011:Cl

[539] Anonymous. Contents. Computational Statistics & Data Analysis, 55 (9):vi–vii, September 1, 2011. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947311001587.

Drovandi:2011:LFB

[540] Christopher C. Drovandi and Anthony N. Pettitt. Likelihood-free Bayesian estimation of multivariate quantile distributions. Computational Statistics & Data Analysis, 55(9):2541– 2556, September 1, 2011. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947311001125.

Lane:2011:AOF

[541] Stephen E. Lane and Andrew P. Robinson. An alternative objective function for fitting regression trees to functional response variables. Computational Statistics & Data Analysis, 55(9):2557– 2567, September 1, 2011. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947311001137.

Blanco-Fernandez:2011:EFS

[542] Angela Blanco-Fernández, Norberto Corral, and Gil González-Rodríguez. Estimation of a flexible simple linear model for interval data based on set arithmetic. Computational Statistics & Data Analysis, 55(9):2568– 2578, September 1, 2011. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947311000843.

Hyndman:2011:OCF

[543] Rob J. Hyndman, Roman A. Ahmed, George Athanasopoulos, and Han Lin Shang. Optimal combination forecasts for hierarchical time series. Computational Statistics & Data Analysis, 55(9):2579–2589, September 1, 2011. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0167947311000971.

Cheng:2011:MTC

[544] Xixin Cheng, W. K. Li, Philip L. H. Yu, Xuan Zhou, Chao Wang, and P. H. Lo. Modeling threshold conditional heteroscedasticity with regime-dependent skewness and kurtosis. Computational Statistics & Data Analysis, 55(9):2590– 2604, September 1, 2011. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947311000995.

Park:2011:GMM

[545] Suk K. Park, Sung K. Ahn, and Sinsup Cho. Generalized method of moments estimation for cointegrated vector autoregressive models. Computational Statistics & Data Analysis, 55(9):2605– 2618, September 1, 2011. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947311001010.

Berrendero:2011:PCM

[546] J. R. Berrendero, A. Justel, and M. Svarc. Principal components for multivariate functional data. Computational Statistics & Data Analysis, 55(9):2619–2634, September 1, 2011. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0167947311001022.

Lee:2011:MET

[547] Sangyeol Lee, Ilia Vonta, and Alex Karagrigoriou. A maximum entropy type test of fit. Computational Statistics & Data Analysis, 55(9):2635– 2643, September 1, 2011. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947311001034.

Cai:2011:BPH

[548] Bo Cai, Xiaoyan Lin, and Lianming Wang. Bayesian proportional hazards model for current status data with monotone splines. Computational Statistics & Data Analysis, 55(9):2644– 2651, September 1, 2011. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947311001046.

Edwards:2011:SDO

[549] David J. Edwards and Robert W. Mee. Supersaturated designs: Are our results significant? Computational Statistics & Data Analysis, 55(9):2652– 2664, September 1, 2011. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016794731100106X.

Friguet:2011:EPT

[550] Chloé Friguet and David Causeur. Estimation of the proportion of true null hypotheses in high-dimensional data under dependence. Computational Statistics & Data Analysis, 55(9):2665– 2676, September 1, 2011. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947311001071.

Dimova:2011:IMM

[551] Rositsa B. Dimova, Marianthi Markatou, and Andrew H. Talal. Information methods for model selection in linear mixed effects models with application to HCV data. Computational Statistics & Data Analysis, 55(9):2677– 2697, September 1, 2011. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947311001101.

Zheng:2011:HBF

[552] Gang Zheng, Ao Yuan, and Neal Jeffries. Hybrid Bayes factors for genomewide association studies when a robust test is used. Computational Statistics & Data Analysis, 55(9):2698– 2711, September 1, 2011. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947311001149.

Montanari:2011:MLE

[553] Angela Montanari and Cinzia Viroli. Maximum likelihood estimation of mixtures of factor analyzers. Computational Statistics & Data Analysis, 55(9):2712–2723, September 1, 2011. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947311001289.

Denecke:2011:RET

[554] Liesa Denecke and Christine H. Müller. Robust estimators and tests for bivariate copulas based on likelihood depth. Computational Statistics & Data Analysis, 55(9):2724–2738, September 1, 2011. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947311001320.

Lo:2011:BMC

[555] Yungtai Lo. Bias from misspecification of the component variances in a normal mixture. Computational Statistics & Data Analysis, 55(9):2739– 2747, September 1, 2011. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947311001368.

Shi:2011:CAT

[556] Qian Shi, Lindsay A. Renfro, Brian M. Bot, Tomasz Burzykowski, Marc Buyse, and Daniel J. Sargent. Comparative assessment of trial-level surrogacy measures for candidate timeto-event surrogate endpoints in clinical trials. Computational Statistics & Data Analysis, 55(9):2748–2757, September 1, 2011. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947311001058.

Epifanio:2011:FDA

[557] Irene Epifanio and Noelia Ventura-Campos. Functional data analysis in shape analysis. Computational Statistics & Data Analysis, 55(9):2758– 2773, September 1, 2011. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947311001307.

Anonymous:2011:EBb

[558] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 55 (10):iii–v, October 1, 2011. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947311001885.

Anonymous:2011:Cb

[559] Anonymous. Contents. Computational Statistics & Data Analysis, 55 (10):vi, October 1, 2011. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947311001897.

Balakrishnan:2011:RMP

[560] N. Balakrishnan and H. M. Saleh. Relations for moments of progressively Type-II censored order statistics from half-logistic distribution with applications to inference. Computational Statistics & Data Analysis, 55(10): 2775–2792, October 1, 2011. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947311001319.

Templ:2011:ISR

[561] Matthias Templ, Alexander Kowarik, and Peter Filzmoser. Iterative stepwise regression imputation using standard and robust methods. Computational Statistics & Data Analysis, 55(10): 2793–2806, October 1, 2011. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947311001411.

Kwon:2011:ESS

[562] Deukwoo Kwon, Maria Teresa Landi, Marina Vannucci, Haleem J. Issaq, DaRue Prieto, and Ruth M. Pfeiffer. An efficient stochastic search for Bayesian variable selection with high-dimensional correlated predictors. Computational Statistics & Data Analysis, 55(10):2807– 2818, October 1, 2011. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947311001484.

Su:2011:AEP

[563] Yan Su, Lianjie Shu, and Kwok-Leung Tsui. Adaptive EWMA procedures for monitoring processes subject to linear drifts. Computational Statistics & Data Analysis, 55(10):2819– 2829, October 1, 2011. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794731100137X.

Xu:2011:OBA

[564] Ancha Xu and Yincai Tang. Objective Bayesian analysis of accelerated competing failure models under Type-I censoring. Computational Statistics & Data Analysis, 55(10): 2830–2839, October 1, 2011. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947311001381.

Holmstrom:2011:SSM

[565] Lasse Holmström, Leena Pasanen, Reinhard Furrer, and Stephan R. Sain. Scale space multiresolution analysis of random signals. Computational Statistics & Data Analysis, 55(10): 2840–2855, October 1, 2011. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016794731100140X.

Gurler:2011:FCL

[566] Ülkü Gürler and C. Deniz Yenigün. Full and conditional likelihood approaches for hazard change-point estimation with truncated and censored data. Computational Statistics & Data Analysis, 55(10):2856– 2870, October 1, 2011. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947311001435.

Hazelton:2011:SRS

[567] Martin L. Hazelton and Berwin A. Turlach. Semiparametric regression with shape-constrained penalized splines. Computational Statistics & Data Analysis, 55(10):2871– 2879, October 1, 2011. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947311001472.

Nam:2011:PSS

[568] Jun mo Nam. Power and sample size requirements for non-inferiority in studies comparing two matched proportions where the events are correlated. Computational Statistics & Data Analysis, 55(10):2880– 2887, October 1, 2011. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947311001496.

Anonymous:2011:EBc

[569] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 55 (11):iii–v, November 1, 2011. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947311002271.

Anonymous:2011:Cc

[570] Anonymous. Contents. Computational Statistics & Data Analysis, 55 (11):vi–vii, November 1, 2011. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947311002283.

Cai:2011:MGL

[571] Jing-Heng Cai, Xin-Yuan Song, Kwok-Hap Lam, and Edward Hak-Sing Ip. A mixture of generalized latent variable models for mixed mode and heterogeneous data. Computational Statistics & Data Analysis, 55(11):2889– 2907, November 1, 2011. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947311001770.

Zak-Szatkowska:2011:MVB

[572] Małgorzata Żak-Szatkowska and Małgorzata Bogdan. Modified versions of the Bayesian information criterion for sparse generalized linear models. Computational Statistics & Data Analysis, 55(11):2908–2924, November 1, 2011. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947311001459.

Wang:2011:NTB

[573] Haiyan Wang, Diego Maldonado, and Sharad Silwal. A nonparametric-test-based structural similarity measure for digital images. Computational Statistics & Data Analysis, 55(11):2925– 2936, November 1, 2011. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947311001502.

Mendez:2011:ERV

[574] Guillermo Mendez and Sharon Lohr. Estimating residual variance in random forest regression. Computational Statistics & Data Analysis, 55(11): 2937–2950, November 1, 2011. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947311001514.

Castillo:2011:FSD

[575] Nabor O. Castillo, Héctor W. Gómez, Víctor Leiva, and Antonio Sanhueza. On the Fernández–Steel distribution: Inference and application. Computational Statistics & Data Analysis, 55(11):2951–2961, November 1, 2011. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947311001526.

Muhlenstadt:2011:KI

[576] Thomas Mühlenstädt and Sonja Kuhnt. Kernel interpolation. Computational Statistics & Data Analysis, 55(11):2962–2974, November 1, 2011. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0167947311001678.

Silva:2011:TGC

[577] A. Pedro Duarte Silva. Two-group classification with high-dimensional correlated data: a factor model approach. Computational Statistics & Data Analysis, 55(11):2975–2990, November 1, 2011. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794731100168X.

Golosnoy:2011:CCC

[578] Vasyl Golosnoy, Sergiy Ragulin, and Wolfgang Schmid. CUSUM control charts for monitoring optimal portfolio weights. Computational Statistics & Data Analysis, 55(11):2991– 3009, November 1, 2011. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947311001708.

Scrucca:2011:MBS

[579] Luca Scrucca. Model-based SIR for dimension reduction. Computational Statistics & Data Analysis, 55(11): 3010–3026, November 1, 2011. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947311001721.

Zhang:2011:PIP

[580] Weiwei Zhang, Gaorong Li, and Liugen Xue. Profile inference on partially linear varying-coefficient errorsin-variables models under restricted condition. Computational Statistics & Data Analysis, 55(11):3027–3040, November 1, 2011. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947311001782.

Fletcher:2011:MAC

[581] David Fletcher and Peter W. Dillingham. Model-averaged confidence intervals for factorial experiments. Computational Statistics & Data Analysis, 55(11):3041–3048, November 1, 2011. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0167947311001800.

Eshima:2011:TPP

[582] Nobuoki Eshima and Minoru Tabata. Three predictive power measures for generalized linear models: the entropy coefficient of determination, the entropy correlation coefficient and the regression correlation coefficient. Computational Statistics & Data Analysis, 55(11):3049–3058, November 1, 2011. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947311001976.

Li:2011:RBS

[583] Feng Li, Françoise Seillier-Moiseiwitsch, and Valeriy R. Korostyshevskiy. Region-based statistical analysis of 2D PAGE images. Computational Statistics & Data Analysis, 55(11):3059– 3072, November 1, 2011. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947311001794.

Anonymous:2011:EBd

[584] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 55 (12):iii–v, December 1, 2011. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947311002726.

Anonymous:2011:Cd

[585] Anonymous. Contents. Computational Statistics & Data Analysis, 55 (12):vi–vii, December 1, 2011. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947311002738.

Monsalve-Cobis:2011:GFT

[586] Abelardo Monsalve-Cobis, Wenceslao González-Manteiga, and Manuel Febrero-Bande. Goodness-of-fit test for interest rate models: an approach based on empirical processes. Computational Statistics & Data Analysis, 55(12):3073–3092, December 1, 2011. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947311002052.

Khan:2011:SAP

[587] Hafiz M. R. Khan, Ahmed Albatineh, Saeed Alshahrani, Nadine Jenkins, and Nasar U. Ahmed. Sensitivity analysis of predictive modeling for responses from the threeparameter Weibull model with a followup doubly censored sample of cancer patients. Computational Statistics & Data Analysis, 55(12):3093– 3103, December 1, 2011. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947311001836.

Filippone:2011:AIB

[588] Maurizio Filippone and Guido Sanguinetti. Approximate inference of the bandwidth in multivariate kernel density estimation. Computational Statistics & Data Analysis, 55(12):3104– 3122, December 1, 2011. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947311002088.

Brostrom:2011:GLM

[589] Göran Broström and Henrik Holmberg. Generalized linear models with clustered data: Fixed and random effects models. Computational Statistics & Data Analysis, 55(12):3123– 3134, December 1, 2011. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947311002143.

Xu:2011:SLB

[590] Ping-Feng Xu, Jianhua Guo, and Man-Lai Tang. Structural learning for Bayesian networks by testing complete separators in prime blocks. Computational Statistics & Data Analysis, 55(12):3135–3147, December 1, 2011. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947311002209.

Karunamuni:2011:OSM

[591] Rohana J. Karunamuni and Jingjing Wu. One-step minimum Hellinger distance estimation. Computational Statistics & Data Analysis, 55(12): 3148–3164, December 1, 2011. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947311002465.

da-Silva:2011:BII

[592] C. Q. da Silva and A. E. Gomes. Bayesian inference for an item response model for modeling test anxiety. Computational Statistics & Data Analysis, 55(12):3165–3182, December 1, 2011. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947311001824.

Sengupta:2011:SVA

[593] Raghu Nandan Sengupta and Angana Sengupta. Some variants of adaptive sampling procedures and their applications. Computational Statistics & Data Analysis, 55(12):3183– 3196, December 1, 2011. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794731100199X.

Ren:2011:VBM

[594] Qian Ren, Sudipto Banerjee, Andrew O. Finley, and James S. Hodges. Variational Bayesian methods for spatial data analysis. Computational Statistics & Data Analysis, 55(12): 3197–3217, December 1, 2011. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947311002003.

Rubio:2011:IGD

[595] F. J. Rubio and M. F. J. Steel. Inference for grouped data with a truncated skew-Laplace distribution. Computational Statistics & Data Analysis, 55(12):3218–3231, December 1, 2011. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947311002039.

Drechsler:2011:EEE

[596] Jörg Drechsler and Jerome P. Reiter. An empirical evaluation of easily implemented, nonparametric methods for generating synthetic datasets. Computational Statistics & Data Analysis, 55(12):3232–3243, December 1, 2011. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947311002076.

Martinez-Camblor:2011:RMA

[597] Pablo Martínez-Camblor and Norberto Corral. Repeated measures analysis for functional data. Computational Statistics & Data Analysis, 55(12):3244– 3256, December 1, 2011. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947311002106.

Rodriguez-Alvarez:2011:NFD

[598] María Xosé Rodríguez-Álvarez, Javier Roca-Pardiñas, and Carmen Cadarso-Suárez. A new flexible direct ROC regression model: Application to the detection of cardiovascular risk factors by anthropometric measures. Computational Statistics & Data Analysis, 55(12):3257– 3270, December 1, 2011. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947311002118.

Abe:2011:SCM

[599] Toshihiro Abe and Arthur Pewsey. Symmetric circular models through duplication and cosine perturbation. Computational Statistics & Data Analysis, 55(12):3271–3282, December 1, 2011. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794731100212X.

Fernandez-Ponce:2011:NNP

[600] J. M. Fernández-Ponce, F. Pellerey, and M. R. Rodríguez-Griñolo. On a new NBUE property in multivariate sense: an application. Computational Statistics & Data Analysis, 55(12): 3283–3294, December 1, 2011. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947311002131.

Fan:2011:FMA

[601] Jie Fan and Somnath Datta. Fitting marginal accelerated failure time models to clustered survival data with potentially informative cluster size. Computational Statistics & Data Analysis, 55(12):3295–3303, December 1, 2011. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947311002180.

Matsui:2011:VSF

[602] Hidetoshi Matsui and Sadanori Konishi. Variable selection for functional regression models via the L_{1} regularization. Computational Statistics & Data Analysis, 55(12):3304– 3310, December 1, 2011. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947311002192.

Rodrigues:2011:UVL

[603] Josemar Rodrigues, N. Balakrishnan, Gauss M. Cordeiro, and Mário de Castro. A unified view on lifetime distributions arising from selection mechanisms. Computational Statistics & Data Analysis, 55(12):3311– 3319, December 1, 2011. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947311002210.

Donald:2011:BAA

[604] Margaret Donald, Clair L. Alston, Rick R. Young, and Kerrie L. Mengersen. A Bayesian analysis of an agricultural field trial with three spatial dimensions. Computational Statistics & Data Analysis, 55(12):3320– 3332, December 1, 2011. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947311002398.

Zhou:2011:SRM

[605] Xiuqing Zhou and Jin Zhao. Simple resampling methods of approximating the distribution of LAD estimators for doubly censored regression models. Computational Statistics & Data Analysis, 55(12):3333– 3343, December 1, 2011. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947311002404.

Chen:2011:ESE

[606] Ziqi Chen, Ning-Zhong Shi, Wei Gao, and Man-Lai Tang. Efficient semiparametric estimation via Cholesky decomposition for longitudinal data. Computational Statistics & Data Analysis, 55(12):3344– 3354, December 1, 2011. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947311002428.

Dehghan:2011:PSN

[607] Mohammad Hossein Dehghan and Thierry Duchesne. On the performance of some non-parametric estimators of the conditional survival function with interval-censored data. Computational Statistics & Data Analysis, 55(12): 3355–3364, December 1, 2011. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947311002441. See

erratum [2063].

Jimenez-Gamero:2011:MDE

[608] M. D. Jiménez-Gamero, R. Pino-Mejías, V. Alba-Fernández, and J. L. Moreno-Rebollo. Minimum ϕdivergence estimation in misspecified multinomial models. Computational Statistics & Data Analysis, 55(12): 3365–3378, December 1, 2011. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947311002520.

Subtil:2011:CAC

[609] Fabien Subtil. Comments on the article of C.-Y. Lai, L. Tian, and E.

F. Schisterman on the “Exact confidence interval estimation for the Youden index and its corresponding optimal cut-point”. Computational Statistics & Data Analysis, 55(12): 3379–3380, December 1, 2011. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947311002027. See

[709].

Tian:2011:NBM

[610] Guo-Liang Tian, Kai Wang Ng, and Philip L. H. Yu. A note on the binomial model with simplex constraints. Computational Statistics & Data Analysis, 55(12):3381–3385, December 1, 2011. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947311002064.

Said:2011:RNS

[611] Y. H. Said, E. J. Wegman, W. K. Sharabati, and J. T. Rigsby. Retraction notice to “Social networks of author–coauthor relationships” [comput. statist. data anal. 52 (4) 2177–2184]. Computational Statistics & Data Analysis, 55(12):3386, December 1, 2011. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947311002799. See

[2].

Anonymous:2012:EBa

[612] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 56(1):iii–v, January 1, 2012. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947311002957.

Anonymous:2012:Ca

[613] Anonymous. Contents. Computational Statistics & Data Analysis, 56 (1):vi–vii, January 1, 2012. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947311002969.

Packalen:2012:IAC

[614] Mikko Packalen and Tony S. Wirjanto. Inference about clustering and parametric assumptions in covariance matrix estimation. Computational Statistics & Data Analysis, 56 (1):1–14, January 1, 2012. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947311002519.

Boshnakov:2012:PLD

[615] Georgi N. Boshnakov and Sophie Lambert-Lacroix. A periodic Levinson– Durbin algorithm for entropy maximization. Computational Statistics & Data Analysis, 56(1):15–24, January 1, 2012. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947311002556.

Wu:2012:CAP

[616] Tong Tong Wu and Xin He. Coordinate ascent for penalized semiparametric regression on high-dimensional panel count data. Computational Statistics & Data Analysis, 56(1):25– 33, January 1, 2012. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794731100257X.

Neykov:2012:RJM

[617] N. M. Neykov, P. Filzmoser, and P. N. Neytchev. Robust joint modeling of mean and dispersion through trimming. Computational Statistics & Data Analysis, 56(1):34–48, January 1, 2012. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947311002611.

Thon:2012:BMA

[618] Kevin Thon, Håvard Rue, Stein Olav Skrøvseth, and Fred Godtliebsen. Bayesian multiscale analysis of images modeled as Gaussian Markov random fields. Computational Statistics & Data Analysis, 56(1):49–61, January 1, 2012. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947311002635.

Rodriguez:2012:BPM

[619] Alejandro Rodríguez and Esther Ruiz. Bootstrap prediction mean squared errors of unobserved states based on the Kalman filter with estimated parameters. Computational Statistics & Data Analysis, 56(1):62–74, January 1, 2012. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947311002647.

Lopes:2012:REP

[620] Celia Mendes Carvalho Lopes and Heleno Bolfarine. Random effects in promotion time cure rate models. Computational Statistics & Data Analysis, 56 (1):75–87, January 1, 2012. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947311001745.

Lan:2012:BIC

[621] Wei Lan, Hansheng Wang, and Chih-Ling Tsai. A Bayesian information criterion for portfolio selection. Computational Statistics & Data Analysis, 56 (1):88–99, January 1, 2012. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947311002155.

Creemers:2012:NAW

[622] An Creemers, Marc Aerts, Niel Hens, and Geert Molenberghs. A nonparametric approach to weighted estimating equations for regression analysis with missing covariates. Computational Statistics & Data Analysis, 56 (1):100–113, January 1, 2012. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947311002167.

Gebregziabher:2012:CMI

[623] Mulugeta Gebregziabher, Matthew S. Shotwell, Jane M. Charles, and Joyce S. Nicholas. Comparison of methods for identifying phenotype subgroups using categorical features data with application to autism spectrum disorder. Computational Statistics & Data Analysis, 56(1):114–125, January 1, 2012. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947311002179.

Cabral:2012:MMM

[624] Celso Rômulo Barbosa Cabral, Víctor Hugo Lachos, and Marcos O. Prates. Multivariate mixture modeling using skew-normal independent distributions. Computational Statistics & Data Analysis, 56(1):126–142, January 1, 2012. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794731100243X.

Song:2012:QEM

[625] Jongwoo Song and Seongjoo Song. A quantile estimation for massive data with generalized Pareto distribution. Computational Statistics & Data Analysis, 56(1):143–150, January 1, 2012. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947311002477.

Tsiotas:2012:GAS

[626] Georgios Tsiotas. On generalised asymmetric stochastic volatility models. Computational Statistics & Data Analysis, 56(1):151–172, January 1, 2012. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947311002489.

Milas:2012:FCN

[627] Costas Milas and Ruthira Naraidoo. Financial conditions and nonlinearities in the European Central Bank (ECB) reaction function: In-sample and out-of-sample assessment. Computational Statistics & Data Analysis, 56 (1):173–189, January 1, 2012. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947311002490.

Aydogdu:2012:NES

[628] Halil Aydoğdu and Mahmut Kara. Nonparametric estimation in α-series processes. Computational Statistics & Data Analysis, 56(1):190–201, January 1, 2012. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947311002544.

Shi:2012:DRM

[629] Lei Shi and Gemai Chen. Deletion, replacement and mean-shift for diagnostics in linear mixed models. Computational Statistics & Data Analysis, 56 (1):202–208, January 1, 2012. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947311002568.

Paik:2012:CLA

[630] Jane Paik and Zhiliang Ying. A composite likelihood approach for spatially correlated survival data. Computational Statistics & Data Analysis, 56 (1):209–216, January 1, 2012. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947311002581.

Asai:2012:MFN

[631] Manabu Asai, Michael McAleer, and Marcelo C. Medeiros. Modelling and forecasting noisy realized volatility. Computational Statistics & Data Analysis, 56(1):217–230, January 1, 2012. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947311002416.

Anonymous:2012:EBe

[632] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 56 (2):iii–v, February 1, 2012. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947311003392.

Anonymous:2012:Ce

[633] Anonymous. Contents. Computational Statistics & Data Analysis, 56 (2):vi–vii, February 1, 2012. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947311003409.

An:2012:EDS

[634] Xinming An and Peter M. Bentler. Efficient direct sampling MCEM algorithm for latent variable models with binary responses. Computational Statistics & Data Analysis, 56(2):231– 244, February 1, 2012. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947311002453.

Chalise:2012:CPF

[635] Prabhakar Chalise and Brooke L. Fridley. Comparison of penalty functions for sparse canonical correlation analysis. Computational Statistics & Data Analysis, 56(2):245–254, February 1, 2012. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947311002660.

Tian:2012:AQL

[636] Guo-Liang Tian, Man-Lai Tang, and Chunling Liu. Accelerating the quadratic lower-bound algorithm via optimizing the shrinkage parameter. Computational Statistics & Data Analysis, 56(2):255–265, February 1, 2012. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947311002672.

McElroy:2012:TIE

[637] Tucker McElroy and Agnieszka Jach. Tail index estimation in the presence of long-memory dynamics. Computational Statistics & Data Analysis, 56(2):266–282, February 1, 2012. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0167947311002854.

Wan:2012:AMC

[638] You Wan, Tao Pei, Chenghu Zhou, Yong Jiang, Chenxu Qu, and Youlin Qiao. ACOMCD: a multiple cluster detection algorithm based on the spatial scan statistic and ant colony optimization. Computational Statistics & Data Analysis, 56(2):283–296, February 1, 2012. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947311002866.

Babykina:2012:EAM

[639] Génia Babykina and Vincent Couallier. Empirical assessment of the maximum likelihood estimator quality in a parametric counting process model for recurrent events. Computational Statistics & Data Analysis, 56(2):297– 315, February 1, 2012. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794731100288X.

Wei:2012:GCD

[640] Fengrong Wei and Hongxiao Zhu. Group coordinate descent algorithms for nonconvex penalized regression. Computational Statistics & Data Analysis, 56(2):316–326, February 1, 2012. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0167947311003069.

Garcia-Trevino:2012:OWB

[641] E. S. García-Treviño and J. A. Barria. Online wavelet-based density estimation for non-stationary streaming data. Computational Statistics & Data Analysis, 56(2):327–344, February 1, 2012. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947311003082.

Mainassara:2012:CEI

[642] Y. Boubacar Mainassara, M. Carbon, and C. Francq. Computing and estimating information matrices of weak ARMA models. Computational Statistics & Data Analysis, 56(2):345– 361, February 1, 2012. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794731100260X.

Zhao:2012:BPM

[643] L. Zhao and M. Banerjee. Bayesian piecewise mixture model for racial disparity in prostate cancer progression. Computational Statistics & Data Analysis, 56(2):362–369, February 1, 2012. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947311002659.

Zhao:2012:MGT

[644] Xiaobing Zhao and Xian Zhou. Modeling gap times between recurrent events by marginal rate function. Computational Statistics & Data Analysis, 56(2):370–383, February 1, 2012. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947311002829.

Qu:2012:CDE

[645] Leming Qu and Wotao Yin. Copula density estimation by total variation penalized likelihood with linear equality constraints. Computational Statistics & Data Analysis, 56(2):384– 398, February 1, 2012. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947311002830.

Desmond:2012:MEL

[646] A. F. Desmond, Carlos L. Cíntora González, R. S. Singh, and Xuewen Lu. A mixed effects log-linear model based on the Birnbaum–Saunders distribution. Computational Statistics & Data Analysis, 56(2):399–407, February 1, 2012. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947311002842.

Stoklosa:2012:RSA

[647] Jakub Stoklosa and Richard M. Huggins. A robust P -spline approach to closed population capture–recapture models with time dependence and heterogeneity. Computational Statistics & Data Analysis, 56(2):408–417, February 1, 2012. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947311002891.

Frey:2012:IME

[648] Jesse Frey and Timothy G. Feeman. An improved mean estimator for judgment post-stratification. Computational Statistics & Data Analysis, 56(2):418– 426, February 1, 2012. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947311003057.

Han:2012:BTE

[649] Bing Han and Siddhartha R. Dalal. A Bernstein-type estimator for decreasing density with application to p-value adjustments. Computational Statistics & Data Analysis, 56(2):427– 437, February 1, 2012. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947311003094.

Cao:2012:EHM

[650] Chun-Zheng Cao, Jin-Guan Lin, and Xiao-Xin Zhu. On estimation of a heteroscedastic measurement error model under heavy-tailed distributions. Computational Statistics & Data Analysis, 56(2):438–448, February 1, 2012. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947311003100.

Meintanis:2012:ETG

[651] Simos G. Meintanis and Efthimios Tsionas. Erratum to: “Testing for the generalized normal-Laplace distribution with applications” [computational statistics and data analysis 54 (2010) 3174–3180]. Computational Statistics & Data Analysis, 56 (2):449, February 1, 2012. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794731100346X. See

[289].

Anonymous:2012:EBf

[652] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 56(3):iii–v, March 1, 2012. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794731100380X.

Anonymous:2012:Cf

[653] Anonymous. Contents. Computational Statistics & Data Analysis, 56 (3):vi–vii, March 1, 2012. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947311003811.

Lin:2012:IWD

[654] Chien-Tai Lin, Cheng-Chieh Chou, and Yen-Lung Huang. Inference for the Weibull distribution with progressive hybrid censoring. Computational Statistics & Data Analysis, 56 (3):451–467, March 1, 2012. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947311003203.

Fang:2012:SNC

[655] Yixin Fang and Junhui Wang. Selection of the number of clusters via the bootstrap method. Computational Statistics & Data Analysis, 56 (3):468–477, March 1, 2012. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947311003215.

Berrett:2012:DAS

[656] Candace Berrett and Catherine A. Calder. Data augmentation strategies for the Bayesian spatial probit regression model. Computational Statistics & Data Analysis, 56(3):478– 490, March 1, 2012. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947311003227.

Rizopoulos:2012:FFJ

[657] Dimitris Rizopoulos. Fast fitting of joint models for longitudinal and event time data using a pseudo-adaptive Gaussian quadrature rule. Computational Statistics & Data Analysis, 56 (3):491–501, March 1, 2012. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947311003264.

Balakrishnan:2012:EAO

[658] N. Balakrishnan and M. H. Ling. EM algorithm for one-shot device testing under the exponential distribution. Computational Statistics & Data Analysis, 56(3):502– 509, March 1, 2012. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794731100329X.

Yu:2012:PTI

[659] Donghyeon Yu, Johan Lim, Feng Liang, Kyunga Kim, Byung Soo Kim, and Woncheol Jang. Permutation test for incomplete paired data with application to cDNA microarray data. Computational Statistics & Data Analysis, 56(3):510–521, March 1, 2012. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947311003112.

Lui:2012:EPR

[660] Kung-Jong Lui and Kuang-Chao Chang. Estimation of the proportion ratio under a simple crossover trial. Computational Statistics & Data Analysis, 56(3):522–530, March 1, 2012. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947311003124.

VanAelst:2012:SDE

[661] S. Van Aelst, E. Vandervieren, and G. Willems. A Stahel–Donoho estimator based on Huberized outlyingness. Computational Statistics & Data Analysis, 56(3):531– 542, March 1, 2012. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947311003136.

Hotz:2012:LAI

[662] Thomas Hotz, Philipp Marnitz, Rahel Stichtenoth, Laurie Davies, Zakhar Kabluchko, and Axel Munk. Locally adaptive image denoising by a statistical multiresolution criterion. Computational Statistics & Data Analysis, 56 (3):543–558, March 1, 2012. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947311003173.

Osman:2012:NRM

[663] Muhtarjan Osman and Sujit K. Ghosh. Nonparametric regression models for right-censored data using Bernstein polynomials. Computational Statistics & Data Analysis, 56(3):559– 573, March 1, 2012. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947311003185.

Zhang:2012:GDF

[664] Bo Zhang, Xiaotong Shen, and Sunni L. Mumford. Generalized degrees of freedom and adaptive model selection in linear mixedeffects models. Computational Statistics & Data Analysis, 56(3):574– 586, March 1, 2012. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947311003197.

Wang:2012:FLR

[665] Guochang Wang, Nan Lin, and Baoxue Zhang. Functional linear regression after spline transformation. Computational Statistics & Data Analysis, 56 (3):587–601, March 1, 2012. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947311003240.

Guo:2012:SIS

[666] Junpeng Guo, Wenhua Li, Chenhua Li, and Sa Gao. Standardization of interval symbolic data based on the empirical descriptive statistics. Computational Statistics & Data Analysis, 56 (3):602–610, March 1, 2012. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947311003252.

Luts:2012:MEL

[667] Jan Luts, Geert Molenberghs, Geert Verbeke, Sabine Van Huffel, and Johan A. K. Suykens. A mixed effects least squares support vector machine model for classification of longitudinal data. Computational Statistics & Data Analysis, 56(3):611– 628, March 1, 2012. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947311003276.

Yu:2012:CAI

[668] Dalei Yu and Kelvin K. W. Yau. Conditional Akaike information criterion for generalized linear mixed models. Computational Statistics & Data Analysis, 56(3):629–644, March 1, 2012. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947311003318.

Chen:2012:MAM

[669] Chyong-Mei Chen and Tai-Fang C. Lu. Marginal analysis of multivariate failure time data with a surviving fraction based on semiparametric transformation cure models. Computational Statistics & Data Analysis, 56 (3):645–655, March 1, 2012. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016794731100332X.

Deng:2012:MPA

[670] Yihao Deng, Roy T. Sabo, and N. Rao Chaganty. Multivariate probit analysis of binary familial data using stochastic representations. Computational Statistics & Data Analysis, 56 (3):656–663, March 1, 2012. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947311003331.

Wang:2012:SDF

[671] Jianqiang C. Wang. Sample distribution function based goodness-of-fit test for complex surveys. Computational Statistics & Data Analysis, 56 (3):664–679, March 1, 2012. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947311003343.

Rondon:2012:FPE

[672] Luz Marina Rondon, Luis Hernando Vanegas, and Cristiano Ferraz. Finite population estimation under generalized linear model assistance. Computational Statistics & Data Analysis, 56 (3):680–697, March 1, 2012. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947311003483.

Ortega:2012:LLR

[673] Edwin M. M. Ortega, Gauss M. Cordeiro, and Artur J. Lemonte. A log-linear regression model for the β-Birnbaum–Saunders distribution with censored data. Computational Statistics & Data Analysis, 56(3):698– 718, March 1, 2012. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947311003495.

Fernandez:2012:OAS

[674] Arturo J. Fernández and Carlos J. Pérez-González. Optimal acceptance sampling plans for log-location– scale lifetime models using average risks. Computational Statistics & Data Analysis, 56(3):719– 731, March 1, 2012. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947311003513.

Hu:2012:BAB

[675] Shuowen Hu, D. S. Poskitt, and Xibin Zhang. Bayesian adaptive bandwidth kernel density estimation of irregular multivariate distributions. Computational Statistics & Data Analysis, 56 (3):732–740, March 1, 2012. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947311003537.

Gomez:2012:PDR

[676] Juan Carlos Gomez and Marie-Francine Moens. PCA document reconstruction for email classification. Computational Statistics & Data Analysis, 56(3):741–751, March 1, 2012. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947311003549.

Anonymous:2012:EBg

[677] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 56(4):iii–v, April 1, 2012. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947311004373.

Anonymous:2012:Cg

[678] Anonymous. Contents. Computational Statistics & Data Analysis, 56 (4):vi–vii, April 1, 2012. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947311004385.

Fung:2012:ESI

[679] Wing-Kam Fung, Xuming He, Mia Hubert, Stephen Portnoy, and Huixia Judy Wang. Editorial for the special issue on quantile regression and semiparametric methods. Computational Statistics & Data Analysis, 56(4):753– 754, April 1, 2012. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947311004427.

Sobotka:2012:GER

[680] Fabian Sobotka and Thomas Kneib. Geoadditive expectile regression. Computational Statistics & Data Analysis, 56(4):755–767, April 1, 2012. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947310004433.

Leorato:2012:AEE

[681] Samantha Leorato, Franco Peracchi, and Andrei V. Tanase. Asymptotically efficient estimation of the conditional expected shortfall. Computational Statistics & Data Analysis, 56 (4):768–784, April 1, 2012. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794731100079X.

Pang:2012:VEC

[682] Lei Pang, Wenbin Lu, and Huixia Judy Wang. Variance estimation in censored quantile regression via induced smoothing. Computational Statistics & Data Analysis, 56(4):785– 796, April 1, 2012. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310004068.

Lin:2012:QRD

[683] Guixian Lin, Xuming He, and Stephen Portnoy. Quantile regression with doubly censored data. Computational Statistics & Data Analysis, 56 (4):797–812, April 1, 2012. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947311001009.

Bang:2012:SEF

[684] Sungwan Bang and Myoungshic Jhun. Simultaneous estimation and factor selection in quantile regression via adaptive sup-norm regularization. Computational Statistics & Data Analysis, 56 (4):813–826, April 1, 2012. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947311000454.

Ji:2012:MSB

[685] Yonggang Ji, Nan Lin, and Baoxue Zhang. Model selection in binary and Tobit quantile regression using the Gibbs sampler. Computational Statistics & Data Analysis, 56(4):827– 839, April 1, 2012. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947311003604.

Paindaveine:2012:CMO

[686] Davy Paindaveine and Miroslav Šiman. Computing multiple-output regression quantile regions. Computational Statistics & Data Analysis, 56(4):840– 853, April 1, 2012. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947310004421.

Christmann:2012:CSV

[687] Andreas Christmann and Robert Hable. Consistency of support vector machines using additive kernels for additive models. Computational Statistics & Data Analysis, 56(4):854– 873, April 1, 2012. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947311001356.

Raheem:2012:APS

[688] S. M. Enayetur Raheem, S. Ejaz Ahmed, and Kjell A. Doksum. Absolute penalty and shrinkage estimation in partially linear models. Computational Statistics & Data Analysis, 56 (4):874–891, April 1, 2012. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947311003525.

Colubi:2012:SIF

[689] Ana Colubi and Didier Dubois. Special issue on fuzzy sets in statistics. Computational Statistics & Data Analysis, 56 (4):892–893, April 1, 2012. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947311004464.

Antoine:2012:CCE

[690] V. Antoine, B. Quost, M.-H. Masson, and T. Denœux. CECM: Constrained evidential C-means algorithm. Computational Statistics & Data Analysis, 56 (4):894–914, April 1, 2012. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310003646.

Coppi:2012:FPC

[691] Renato Coppi, Pierpaolo D’Urso, and Paolo Giordani. Fuzzy and possibilistic clustering for fuzzy data. Computational Statistics & Data Analysis, 56 (4):915–927, April 1, 2012. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310003567.

Verwaeren:2012:LPO

[692] Jan Verwaeren, Willem Waegeman, and Bernard De Baets. Learning partial ordinal class memberships with kernel-based proportional odds models. Computational Statistics & Data Analysis, 56(4):928– 942, April 1, 2012. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947310004755.

Gonzalez-Rodriguez:2012:FDT

[693] Gil González-Rodríguez, Ana Colubi, and María Ángeles Gil. Fuzzy data treated as functional data: a oneway ANOVA test approach. Computational Statistics & Data Analysis, 56 (4):943–955, April 1, 2012. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947310002586.

Ramos-Guajardo:2012:STE

[694] Ana Belén Ramos-Guajardo and María Asunción Lubiano. K-sample tests for equality of variances of random fuzzy sets. Computational Statistics & Data Analysis, 56(4):956– 966, April 1, 2012. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947310004536.

Coletti:2012:GBI

[695] Giulianella Coletti, Osvaldo Gervasi, Sergio Tasso, and Barbara Vantaggi. Generalized Bayesian inference in a fuzzy context: From theory to a virtual reality application. Computational Statistics & Data Analysis, 56 (4):967–980, April 1, 2012. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947311002374.

Capotorti:2012:CSA

[696] Andrea Capotorti and Eva Barbanera. Credit scoring analysis using a fuzzy probabilistic rough set model. Computational Statistics & Data Analysis, 56 (4):981–994, April 1, 2012. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947311002532.

Anonymous:2012:EBh

[697] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 56(5):iii–v, May 1, 2012. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947312000199.

Anonymous:2012:Ch

[698] Anonymous. Contents. Computational Statistics & Data Analysis, 56 (5):vi–vii, May 1, 2012. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947312000205.

Mittlbock:2012:SIC

[699] Martina Mittlböck, Lutz Edler, Michael LeBlanc, Joyce Niland, and Koos Zwinderman. Second issue for computational statistics for clinical research. Computational Statistics & Data Analysis, 56(5):995– 997, May 1, 2012. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947312000217.

Edler:2012:OPD

[700] Lutz Edler. Obituary for Professor Dr. Dr. h.c. Norbert Victor (1940–2011). Computational Statistics & Data Analysis, 56(5):998– 1015, May 1, 2012. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947312000229.

Chen:2012:UEF

[701] Weijie Chen, Waleed A. Yousef, Brandon D. Gallas, Elizabeth R. Hsu, Samir Lababidi, Rong Tang, Gene A. Pennello, W. Fraser Symmans, and Lajos Pusztai. Uncertainty estimation with a finite dataset in the assessment of classification models. Computational Statistics & Data Analysis, 56 (5):1016–1027, May 1, 2012. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794731100209X.

Fiocco:2012:MAT

[702] Marta Fiocco, Theo Stijnen, and Hein Putter. Meta-analysis of timeto-event outcomes using a hazardbased approach: Comparison with other models, robustness and metaregression. Computational Statistics & Data Analysis, 56(5):1028– 1037, May 1, 2012. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947311001757.

Frommlet:2012:MVB

[703] Florian Frommlet, Felix Ruhaltinger, Piotr Twaróg, and Małgorzata Bogdan. Modified versions of Bayesian information criterion for genome-wide association studies. Computational Statistics & Data Analysis, 56(5): 1038–1051, May 1, 2012. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794731100171X. See

corrigendum [1268].

Gao:2012:AOA

[704] Jingjing Gao, Yi Pan, and Michael Haber. Assessment of observer agreement for matched repeated binary measurements. Computational Statistics & Data Analysis, 56(5):1052– 1060, May 1, 2012. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947311004026.

Gerlinger:2012:ODE

[705] Christoph Gerlinger, Harald Siedentop, Oke Gerke, Ilka Schellschmidt, and Jan Endrikat. Optimal dose de-escalation trial designs for novel contraceptives in women. Computational Statistics & Data Analysis, 56 (5):1061–1068, May 1, 2012. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947311002908.

Jiang:2012:IPM

[706] Shan Jiang and Dongsheng Tu. Inference on the probability P (T_{1} < T_{2}) as a measurement of treatment effect under a density ratio model and random censoring. Computational Statistics & Data Analysis, 56 (5):1069–1078, May 1, 2012. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947311000569.

Klingenberg:2012:SSC

[707] Bernhard Klingenberg. Simultaneous score confidence bounds for risk differences in multiple comparisons to a control. Computational Statistics & Data Analysis, 56(5):1079– 1089, May 1, 2012. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947311000442.

Konietschke:2012:RPM

[708] F. Konietschke, S. W. Harrar, K. Lange, and E. Brunner. Ranking procedures for matched pairs with missing data — asymptotic theory and a small sample approximation. Computational Statistics & Data Analysis, 56 (5):1090–1102, May 1, 2012. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947311001150.

Lai:2012:ECI

[709] Chin-Ying Lai, Lili Tian, and Enrique F. Schisterman. Exact confidence interval estimation for the Youden index and its corresponding optimal cut-point. Computational Statistics & Data Analysis, 56(5): 1103–1114, May 1, 2012. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310004512. See

[609].

Marschner:2012:SAP

[710] Ian C. Marschner, Alexandra C. Gillett, and Rachel L. O’Connell. Stratified additive Poisson models: Computational methods and applications in clinical epidemiology. Computational Statistics & Data Analysis, 56 (5):1115–1130, May 1, 2012. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947311002878.

Park:2012:TEL

[711] Junyong Park and DoHwan Park. Testing the equality of a large number of normal population means. Computational Statistics & Data Analysis, 56 (5):1131–1149, May 1, 2012. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947311003161.

Parsons:2012:RPI

[712] Nick Parsons, Tim Friede, Susan Todd, Elsa Valdes Marquez, Jeremy Chataway, Richard Nicholas, and Nigel Stallard. An R package for implementing simulations for seamless phase II/ III clinical trials using early outcomes for treatment selection. Computational Statistics & Data Analysis, 56 (5):1150–1160, May 1, 2012. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947310004159.

Nofuentes:2012:GHT

[713] José Antonio Roldán Nofuentes, Juan de Dios Luna del Castillo, and Miguel Ángel Montero Alonso. Global hypothesis test to simultaneously compare the predictive values of two binary diagnostic tests. Computational Statistics & Data Analysis, 56(5): 1161–1173, May 1, 2012. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947311002040.

Rotondi:2012:SSE

[714] Michael Rotondi and Allan Donner. Sample size estimation in cluster randomized trials: an evidence-based perspective. Computational Statistics & Data Analysis, 56(5):1174– 1187, May 1, 2012. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947310004780.

Spangl:2012:AST

[715] B. Spangl and R. Dutter. Analyzing short-term measurements of heart rate variability in the frequency domain using robustly estimated spectral density functions. Computational Statistics & Data Analysis, 56 (5):1188–1199, May 1, 2012. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947311001769.

Tang:2012:CIC

[716] Man-Lai Tang, Shi-Fang Qiu, and Wai-Yin Poon. Confidence interval construction for disease prevalence based on partial validation series. Computational Statistics & Data Analysis, 56 (5):1200–1220, May 1, 2012. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947311000557.

Todd:2012:SMP

[717] Susan Todd, M. Fazil Baksh, and John Whitehead. Sequential methods for pharmacogenetic studies. Computational Statistics & Data Analysis, 56 (5):1221–1231, May 1, 2012. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947311000788.

Vens:2012:GEE

[718] Maren Vens and Andreas Ziegler. Generalized estimating equations and regression diagnostics for longitudinal controlled clinical trials: a case study. Computational Statistics & Data Analysis, 56(5):1232– 1242, May 1, 2012. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947311001393.

Volterman:2012:ENM

[719] William Volterman, N. Balakrishnan, and Erhard Cramer. Exact nonparametric meta-analysis for multiple independent doubly Type-II censored samples. Computational Statistics & Data Analysis, 56(5):1243– 1255, May 1, 2012. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947311003690.

Wagner:2012:BMS

[720] Helga Wagner and Christine Duller. Bayesian model selection for logistic regression models with random intercept. Computational Statistics & Data Analysis, 56(5):1256– 1274, May 1, 2012. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947311002507.

Werft:2012:IPB

[721] W. Werft, A. Benner, and A. Kopp-Schneider. On the identification of predictive biomarkers: Detecting treatment-by-gene interaction in highdimensional data. Computational Statistics & Data Analysis, 56(5): 1275–1286, May 1, 2012. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310004470.

Xie:2012:ALC

[722] Hui Xie. Analyzing longitudinal clinical trial data with nonignorable missingness and unknown missingness reasons. Computational Statistics & Data Analysis, 56(5):1287– 1300, May 1, 2012. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310004494.

Yang:2012:TNI

[723] Zhao Yang, Xuezheng Sun, and James W. Hardin. Testing noninferiority for clustered matched-pair binary data in diagnostic medicine. Computational Statistics & Data Analysis, 56(5):1301–1320, May 1, 2012. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0167947311002222.

Anonymous:2012:EBi

[724] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 56(6):iii–v, June 2012. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794731200045X.

Anonymous:2012:Ci

[725] Anonymous. Contents. Computational Statistics & Data Analysis, 56 (6):vi–x, June 2012. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947312000461.

Levine:2012:FME

[726] Richard A. Levine, Juanjuan Fan, Pamela Ohman Strickland, and Shaban Demirel. Frailty modeling via the empirical Bayes–Hastings sampler. Computational Statistics & Data Analysis, 56(6):1303–1318, June 2012. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947311003239.

Ruggieri:2012:ECB

[727] Eric Ruggieri and Charles E. Lawrence. On efficient calculations for Bayesian variable selection. Computational Statistics & Data Analysis, 56(6): 1319–1332, June 2012. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947311003574.

Ignatieva:2012:EDC

[728] Katja Ignatieva and Eckhard Platen. Estimating the diffusion coefficient function for a diversified world stock index. Computational Statistics & Data Analysis, 56(6):1333– 1349, June 2012. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947311003616.

Khan:2012:PIF

[729] Hafiz M. R. Khan. Predictive inference for a future response using symmetrically trimmed sample from the half-normal model. Computational Statistics & Data Analysis, 56 (6):1350–1361, June 2012. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947311003665.

Eidsvik:2012:ABI

[730] Jo Eidsvik, Andrew O. Finley, Sudipto Banerjee, and Håvard Rue. Approximate Bayesian inference for large spatial datasets using predictive process models. Computational Statistics & Data Analysis, 56(6):1362– 1380, June 2012. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947311003938.

Melnykov:2012:IEA

[731] Volodymyr Melnykov and Igor Melnykov. Initializing the EM algorithm in Gaussian mixture models with an unknown number of components. Computational Statistics & Data Analysis, 56(6):1381–1395, June 2012. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947311003963.

Punathumparambath:2012:ATI

[732] Bindu Punathumparambath, Sangita Kulathinal, and Sebastian George. Asymmetric type II compound Laplace distribution and its application to microarray gene expression. Computational Statistics & Data Analysis, 56 (6):1396–1404, June 2012. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947311004014.

Schutzenmeister:2012:RAL

[733] André Schützenmeister and Hans-Peter Piepho. Residual analysis of linear mixed models using a simulation approach. Computational Statistics & Data Analysis, 56(6):1405– 1416, June 2012. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947311004038.

Moraga:2012:GCM

[734] Paula Moraga and Andrew B. Lawson. Gaussian component mixtures and CAR models in Bayesian disease mapping. Computational Statistics & Data Analysis, 56(6):1417– 1433, June 2012. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947311004087.

Cardot:2012:FRA

[735] Hervé Cardot, Peggy Cénac, and Jean-Marie Monnez. A fast and recursive algorithm for clustering large datasets with k-medians. Computational Statistics & Data Analysis, 56 (6):1434–1449, June 2012. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947311004178.

Rigat:2012:PHS

[736] F. Rigat and A. Mira. Parallel hierarchical sampling: a generalpurpose interacting Markov chains Monte Carlo algorithm. Computational Statistics & Data Analysis, 56 (6):1450–1467, June 2012. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016794731100418X.

Erhardt:2012:ABC

[737] Robert J. Erhardt and Richard L. Smith. Approximate Bayesian computing for spatial extremes. Computational Statistics & Data Analysis, 56 (6):1468–1481, June 2012. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947311004245.

Pigoli:2012:WFD

[738] Davide Pigoli and Laura M. Sangalli. Wavelets in functional data analysis: Estimation of multidimensional curves and their derivatives. Computational Statistics & Data Analysis, 56 (6):1482–1498, June 2012. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947311004518.

Fabio:2012:PMM

[739] Lizandra C. Fabio, Gilberto A. Paula, and Mário de Castro. A Poisson mixed model with nonnormal random effect distribution. Computational Statistics & Data Analysis, 56 (6):1499–1510, June 2012. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947311004233.

Almeida:2012:EBI

[740] Carlos Almeida and Claudia Czado. Efficient Bayesian inference for stochastic time-varying copula models. Computational Statistics & Data Analysis, 56(6):1511–1527, June 2012. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947311003148.

Kazianka:2012:OBA

[741] Hannes Kazianka. Objective Bayesian analysis for the normal compositional model. Computational Statistics & Data Analysis, 56(6):1528– 1544, June 2012. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016794731100315X.

Lee:2012:MRE

[742] Keunbaik Lee, JungBok Lee, Joseph Hagan, and Jae Keun Yoo. Modeling the random effects covariance matrix for generalized linear mixed models. Computational Statistics & Data Analysis, 56(6):1545– 1551, June 2012. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947311003306.

Hapfelmeier:2012:RPI

[743] A. Hapfelmeier, T. Hothorn, and K. Ulm. Recursive partitioning on incomplete data using surrogate decisions and multiple imputation. Computational Statistics & Data Analysis, 56 (6):1552–1565, June 2012. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947311003550.

Kohler:2012:NEN

[744] Michael Kohler and Adam Krzyżak. Nonparametric estimation of nonstationary velocity fields from 3D particle tracking velocimetry data. Computational Statistics & Data Analysis, 56(6):1566–1580, June 2012. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947311003562.

Hwang:2012:SBK

[745] Eunju Hwang and Dong Wan Shin. Stationary bootstrap for kernel density estimators under ψ-weak dependence. Computational Statistics & Data Analysis, 56(6):1581– 1593, June 2012. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947311003586.

Boente:2012:BCR

[746] Graciela Boente, Marcelo Ruiz, and Ruben H. Zamar. Bandwidth choice for robust nonparametric scale function estimation. Computational Statistics & Data Analysis, 56(6):1594– 1608, June 2012. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947311003598.

Ospina:2012:GCZ

[747] Raydonal Ospina and Silvia L. P. Ferrari. A general class of zero-or-one inflated beta regression models. Computational Statistics & Data Analysis, 56(6):1609–1623, June 2012. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947311003628.

Daniel:2012:MIR

[748] Rhian M. Daniel and Michael G. Kenward. A method for increasing the robustness of multiple imputation. Computational Statistics & Data Analysis, 56(6):1624–1643, June 2012. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016794731100363X.

Leiva:2012:LDT

[749] Ricardo Leiva and Anuradha Roy. Linear discrimination for three-level multivariate data with a separable additive mean vector and a doubly exchangeable covariance structure. Computational Statistics & Data Analysis, 56 (6):1644–1661, June 2012. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947311003641.

Vanegas:2012:DPB

[750] Luis Hernando Vanegas, Luz Marina Rondón, and Francisco José A. Cysneiros. Diagnostic procedures in Birnbaum–Saunders nonlinear regression models. Computational Statistics & Data Analysis, 56(6):1662– 1680, June 2012. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947311003653.

Lee:2012:SSW

[751] Kichun Lee and Brani Vidakovic. Semisupervised wavelet shrinkage. Computational Statistics & Data Analysis, 56 (6):1681–1691, June 2012. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947311003677.

Chen:2012:CRL

[752] Hubert J. Chen and Shu-Fei Wu. A confidence region for the largest and the smallest means under heteroscedasticity. Computational Statistics & Data Analysis, 56(6):1692– 1702, June 2012. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947311003689.

Borges:2012:CDG

[753] Patrick Borges, Josemar Rodrigues, and Narayanaswamy Balakrishnan. Correlated destructive generalized power series cure rate models and associated inference with an application to a cutaneous melanoma data. Computational Statistics & Data Analysis, 56(6):1703–1713, June 2012. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947311003707.

Karabatsos:2012:BNM

[754] George Karabatsos and Stephen G. Walker. Bayesian nonparametric mixed random utility models. Computational Statistics & Data Analysis, 56 (6):1714–1722, June 2012. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947311003719.

Albano:2012:IST

[755] Giuseppina Albano, Virginia Giorno, Patricia Román-Román, and Francisco Torres-Ruiz. Inference on a stochastic two-compartment model in tumor growth. Computational Statistics & Data Analysis, 56(6):1723– 1736, June 2012. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947311003732.

Carmack:2012:NCS

[756] Patrick S. Carmack, Jeffrey S. Spence, William R. Schucany, Richard F. Gunst, Qihua Lin, and Robert W. Haley. A new class of semiparametric semivariogram and nugget estimators. Computational Statistics & Data Analysis, 56(6):1737– 1747, June 2012. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947311003744.

Asomaning:2012:HTD

[757] N. Asomaning and K. J. Archer. High-throughput DNA methylation datasets for evaluating false discovery rate methodologies. Computational Statistics & Data Analysis, 56 (6):1748–1756, June 2012. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947311003914.

Neykov:2012:LTQ

[758] N. M. Neykov, P. Čižek, P. Filzmoser, and P. N. Neytchev. The least trimmed quantile regression. Computational Statistics & Data Analysis, 56 (6):1757–1770, June 2012. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794731100394X.

Costa:2012:CST

[759] Marcelo Azevedo Costa, Renato Martins Assunção, and Martin Kulldorff. Constrained spanning tree algorithms for irregularly-shaped spatial clustering. Computational Statistics & Data Analysis, 56(6):1771– 1783, June 2012. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947311003951.

Lloyd:2012:CHA

[760] Chris J. Lloyd. Computing highly accurate or exact P -values using importance sampling. Computational Statistics & Data Analysis, 56(6): 1784–1794, June 2012. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947311003975.

Scott:2012:BHC

[761] James G. Scott. Benchmarking historical corporate performance. Computational Statistics & Data Analysis, 56(6):1795–1807, June 2012. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947311003987.

Ruiz-Cardenas:2012:DFD

[762] Ramiro Ruiz-Cárdenas, Elias T. Krainski, and Håvard Rue. Direct fitting of dynamic models using integrated nested Laplace approximations — INLA. Computational Statistics & Data Analysis, 56(6):1808– 1828, June 2012. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947311003999.

Yang:2012:TRM

[763] Zhao Yang, Xuezheng Sun, and James W. Hardin. Testing ratio of marginal probabilities in clustered matched-pair binary data. Computational Statistics & Data Analysis, 56 (6):1829–1836, June 2012. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947311004002.

Pang:2012:ESI

[764] Zhen Pang and Liugen Xue. Estimation for the single-index models with random effects. Computational Statistics & Data Analysis, 56 (6):1837–1853, June 2012. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016794731100404X.

Tsimikas:2012:IGL

[765] John V. Tsimikas, Leonidas E. Bantis, and Stelios D. Georgiou. Inference in generalized linear regression models with a censored covariate. Computational Statistics & Data Analysis, 56(6):1854–1868, June 2012. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947311004075.

Josse:2012:SNC

[766] Julie Josse and François Husson. Selecting the number of components in principal component analysis using cross-validation approximations. Computational Statistics & Data Analysis, 56(6):1869–1879, June 2012. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947311004099.

Alexander:2012:GBG

[767] Carol Alexander, Gauss M. Cordeiro, Edwin M. M. Ortega, and José María Sarabia. Generalized beta-generated distributions. Computational Statistics & Data Analysis, 56(6):1880– 1897, June 2012. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947311004129.

Marin:2012:MHT

[768] J. M. Marín and M. T. Rodríguez-Bernal. Multiple hypothesis testing and clustering with mixtures of noncentral t-distributions applied in microarray data analysis. Computational Statistics & Data Analysis, 56 (6):1898–1907, June 2012. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947311004130.

Diniz:2012:NME

[769] Ana Diniz, João Barreiros, and Nuno Crato. A new model for explaining long-range correlations in human time interval production. Computational Statistics & Data Analysis, 56 (6):1908–1919, June 2012. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947311004142.

Baragatti:2012:SVS

[770] M. Baragatti and D. Pommeret. A study of variable selection using g-prior distribution with ridge parameter. Computational Statistics & Data Analysis, 56(6):1920– 1934, June 2012. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947311004154.

Sexton:2012:ACD

[771] Joseph Sexton, Rune Blomhoff, Anette Karlsen, and Petter Laake. Adaptive combination of dependent tests. Computational Statistics & Data Analysis, 56(6):1935–1943, June 2012. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947311004166.

Iddi:2012:COM

[772] Samuel Iddi and Geert Molenberghs. A combined overdispersed and marginalized multilevel model. Computational Statistics & Data Analysis, 56 (6):1944–1951, June 2012. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947311004191.

Arslan:2012:WLL

[773] Olcay Arslan. Weighted LAD-LASSO method for robust parameter estimation and variable selection in regression. Computational Statistics & Data Analysis, 56(6):1952– 1965, June 2012. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947311004208.

VanOirbeek:2012:APA

[774] R. Van Oirbeek and E. Lesaffre. Assessing the predictive ability of a multilevel binary regression model. Computational Statistics & Data Analysis, 56(6):1966–1980, June 2012. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947311004221.

Feng:2012:ADO

[775] Zhenghui Feng and Lixing Zhu. An alternating determination–optimization approach for an additive multiindex model. Computational Statistics & Data Analysis, 56(6):1981– 1993, June 2012. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947311004257.

Xiang:2012:REM

[776] Liming Xiang, Kelvin K. W. Yau, and Andy H. Lee. The robust estimation method for a finite mixture of Poisson mixed-effect models. Computational Statistics & Data Analysis, 56 (6):1994–2005, June 2012. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947311004270.

Du:2012:SRM

[777] Pang Du, Guang Cheng, and Hua Liang. Semiparametric regression models with additive nonparametric components and high dimensional parametric components. Computational Statistics & Data Analysis, 56(6): 2006–2017, June 2012. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947311004282.

Wu:2012:PFF

[778] Shuo-Jye Wu and Syuan-Rong Huang. Progressively first-failure censored reliability sampling plans with cost constraint. Computational Statistics & Data Analysis, 56(6):2018– 2030, June 2012. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947311004294.

Hamid:2012:WKF

[779] Jemila S. Hamid, Celia M. T. Greenwood, and Joseph Beyene. Weighted kernel Fisher discriminant analysis for integrating heterogeneous data. Computational Statistics & Data Analysis, 56(6):2031–2040, June 2012. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947311004300.

Li:2012:SAB

[780] Yang Li, Jianguo Sun, and Shuguang Song. Statistical analysis of bivariate failure time data with Marshall– Olkin Weibull models. Computational Statistics & Data Analysis, 56 (6):2041–2050, June 2012. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947311004312.

Lu:2012:EEA

[781] Xuewen Lu and Peter X.-K. Song. On efficient estimation in additive hazards regression with current status data. Computational Statistics & Data Analysis, 56(6):2051– 2058, June 2012. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947311004324.

Crabbe:2012:IEI

[782] M. Crabbe and M. Vandebroek. Improving the efficiency of individualized designs for the mixed logit choice model by including covariates. Computational Statistics & Data Analysis, 56(6):2059–2072, June 2012. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947311004506.

Delattre:2012:MLE

[783] M. Delattre and M. Lavielle. Maximum likelihood estimation in discrete mixed hidden Markov models using the SAEM algorithm. Computational Statistics & Data Analysis, 56 (6):2073–2085, June 2012. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794731100452X.

Jozwiak:2012:CED

[784] Katarzyna Jóżwiak and Mirjam Moerbeek. Cost-effective designs for trials with discrete-time survival endpoints. Computational Statistics & Data Analysis, 56(6):2086– 2096, June 2012. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947311004531.

Quessy:2012:CMC

[785] Jean-François Quessy and François Éthier. Cramér–von Mises and characteristic function tests for the two and k-sample problems with dependent data. Computational Statistics & Data Analysis, 56(6):2097– 2111, June 2012. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947311004567.

Ran:2012:STW

[786] Chenjian Ran and Zili Deng. Selftuning weighted measurement fusion Kalman filtering algorithm. Computational Statistics & Data Analysis, 56 (6):2112–2128, June 2012. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947312000035.

Li:2012:CQS

[787] Xinmin Li, Lili Tian, Juan Wang, and Josephia R. Muindi. Comparison of quantiles for several normal populations. Computational Statistics & Data Analysis, 56(6):2129– 2138, June 2012. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947312000047.

Artiach:2012:DFM

[788] Miguel Artiach and Josu Arteche. Doubly fractional models for dynamic heteroscedastic cycles. Computational Statistics & Data Analysis, 56(6): 2139–2158, June 2012. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947311003756.

Ong:2012:SMT

[789] Meng Sang Ong, Ye Chow Kuang, and Melanie Po-Leen Ooi. Statistical measures of two dimensional point set uniformity. Computational Statistics & Data Analysis, 56(6): 2159–2181, June 2012. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947311004269.

LeTarnec:2012:CRS

[790] Louis Le Tarnec and Damien Garcia. Corrigendum to “Robust smoothing of gridded data in one and higher dimensions with missing values” [comput. statist. data anal. 54 (2010) 1167–1178]. Computational Statistics & Data Analysis, 56(6): 2182, June 2012. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016794731100421X. See

[112].

Anonymous:2012:EBj

[791] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 56(7):iii–v, July 2012. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947312000849.

Anonymous:2012:Cj

[792] Anonymous. Contents. Computational Statistics & Data Analysis, 56 (7):vi–vii, July 2012. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947312000850.

Yoo:2012:BMC

[793] Changwon Yoo. The Bayesian method for causal discovery of latent-variable models from a mixture of experimental and observational data. Computational Statistics & Data Analysis, 56 (7):2183–2205, July 2012. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947312000291.

Golyandina:2012:NAN

[794] Nina Golyandina, Andrey Pepelyshev, and Ansgar Steland. New approaches to nonparametric density estimation and selection of smoothing parameters. Computational Statistics & Data Analysis, 56(7):2206– 2218, July 2012. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947311004543.

Barranco-Chamorro:2012:CDT

[795] I. Barranco-Chamorro, M. D. Jiménez-Gamero, J. L. Moreno-Rebollo, and J. M. Muñoz-Pichardo. Case-deletion type diagnostics for calibration estimators in survey sampling. Computational Statistics & Data Analysis, 56 (7):2219–2236, July 2012. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947311004555.

Emura:2012:GFT

[796] Takeshi Emura and Yoshihiko Konno. A goodness-of-fit test for parametric models based on dependently truncated data. Computational Statistics & Data Analysis, 56(7):2237– 2250, July 2012. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947312000023.

Rubio:2012:MOT

[797] F. J. Rubio and M. F. J. Steel. On the Marshall–Olkin transformation as a skewing mechanism. Computational Statistics & Data Analysis, 56 (7):2251–2257, July 2012. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947312000059.

Cheng:2012:SIO

[798] Tsung-Chi Cheng. On simultaneously identifying outliers and heteroscedasticity without specific form. Computational Statistics & Data Analysis, 56 (7):2258–2272, July 2012. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947312000060.

Ghimire:2012:CIP

[799] Santosh Ghimire and Haiyan Wang. Classification of image pixels based on minimum distance and hypothesis testing. Computational Statistics & Data Analysis, 56(7):2273– 2287, July 2012. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947312000072.

Gradowska:2012:LST

[800] P. L. Gradowska and R. M. Cooke. Least squares type estimation for Cox regression model and specification error. Computational Statistics & Data Analysis, 56(7):2288– 2302, July 2012. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947312000084.

Chen:2012:BTM

[801] Zhanshou Chen, Zi Jin, Zheng Tian, and Peiyan Qi. Bootstrap testing multiple changes in persistence for a heavy-tailed sequence. Computational Statistics & Data Analysis, 56 (7):2303–2316, July 2012. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947312000308.

Dazard:2012:JAM

[802] Jean-Eudes Dazard and J. Sunil Rao. Joint adaptive mean–variance regularization and variance stabilization of high dimensional data. Computational Statistics & Data Analysis, 56 (7):2317–2333, July 2012. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947312000321.

Alonso:2012:SCF

[803] Andrés M. Alonso, David Casado, and Juan Romo. Supervised classification for functional data: a weighted distance approach. Computational Statistics & Data Analysis, 56(7): 2334–2346, July 2012. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947312000333.

Bai:2012:RFM

[804] Xiuqin Bai, Weixin Yao, and John E. Boyer. Robust fitting of mixture regression models. Computational Statistics & Data Analysis, 56(7): 2347–2359, July 2012. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947312000369.

Slaets:2012:PAB

[805] Leen Slaets, Gerda Claeskens, and Mia Hubert. Phase and amplitude-based clustering for functional data. Computational Statistics & Data Analysis, 56(7):2360–2374, July 2012. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947312000370.

Anonymous:2012:EBk

[806] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 56(8):iii–v, August 2012. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947312001211.

Anonymous:2012:Ck

[807] Anonymous. Contents. Computational Statistics & Data Analysis, 56 (8):vi–vii, August 2012. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947312001223.

Volant:2012:VBA

[808] Stevenn Volant, Marie-Laure Martin Magniette, and Stéphane Robin. Variational Bayes approach for model aggregation in unsupervised classification with Markovian dependency. Computational Statistics & Data Analysis, 56 (8):2375–2387, August 2012. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947312000710.

Rulloni:2012:LGI

[809] Valeria Rulloni, Oscar Bustos, and Ana Georgina Flesia. Large gap imputation in remote sensed imagery of the environment. Computational Statistics & Data Analysis, 56(8): 2388–2403, August 2012. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947312001041.

Augustin:2012:QQP

[810] Nicole H. Augustin, Erik-André Sauleau, and Simon N. Wood. On quantile quantile plots for generalized linear models. Computational Statistics & Data Analysis, 56(8): 2404–2409, August 2012. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947312000692.

Lu:2012:PBU

[811] H. Y. Kevin Lu and G. Alastair Young. Parametric bootstrap under model mis-specification. Computational Statistics & Data Analysis, 56 (8):2410–2420, August 2012. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947312000382.

Levine:2012:SAT

[812] Michael Levine and Jinguang (Tony) Li. A simple additivity test for conditionally heteroscedastic nonlinear autoregression. Computational Statistics & Data Analysis, 56(8):2421– 2429, August 2012. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947312000394.

Fitch:2012:CUD

[813] A. Marie Fitch and Beatrix Jones. The cost of using decomposable Gaussian graphical models for computational convenience. Computational Statistics & Data Analysis, 56(8): 2430–2441, August 2012. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947312000400.

Wang:2012:BIC

[814] Min Wang and Xiaoqian Sun. Bayesian inference for the correlation coefficient in two seemingly unrelated regressions. Computational Statistics & Data Analysis, 56(8):2442– 2453, August 2012. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947312000618.

Seo:2012:RSN

[815] Byungtae Seo and Daeyoung Kim. Root selection in normal mixture models. Computational Statistics & Data Analysis, 56(8):2454– 2470, August 2012. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016794731200062X.

Soliman:2012:EPL

[816] Ahmed A. Soliman, Ahmed H. Abd-Ellah, Naser A. Abou-Elheggag, and Gamal A. Abd-Elmougod. Estimation of the parameters of life for Gompertz distribution using progressive first-failure censored data. Computational Statistics & Data Analysis, 56 (8):2471–2485, August 2012. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947312000655.

Lee:2012:MWD

[817] Paul H. Lee and Philip L. H. Yu. Mixtures of weighted distance-based models for ranking data with applications in political studies. Computational Statistics & Data Analysis, 56 (8):2486–2500, August 2012. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947312000679.

Torti:2012:BTA

[818] Francesca Torti, Domenico Perrotta, Anthony C. Atkinson, and Marco Riani. Benchmark testing of algorithms for very robust regression: FS, LMS and LTS. Computational Statistics & Data Analysis, 56(8):2501– 2512, August 2012. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947312000680.

Pires:2012:CBR

[819] Rubiane M. Pires and Carlos A. R. Diniz. Correlated binomial regression models. Computational Statistics & Data Analysis, 56(8):2513– 2525, August 2012. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947312000709.

Fu:2012:QRL

[820] Liya Fu and You-Gan Wang. Quantile regression for longitudinal data with a working correlation model. Computational Statistics & Data Analysis, 56 (8):2526–2538, August 2012. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947312000722.

Graham:2012:DFE

[821] M. A. Graham, A. Mukherjee, and S. Chakraborti. Distribution-free exponentially weighted moving average control charts for monitoring unknown location. Computational Statistics & Data Analysis, 56(8):2539– 2561, August 2012. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947312000801.

Zhang:2012:SDM

[822] Hongmei Zhang, Kaushik Ghosh, and Pulak Ghosh. Sampling designs via a multivariate hypergeometric-Dirichlet process model for a multispecies assemblage with unknown heterogeneity. Computational Statistics & Data Analysis, 56(8):2562– 2573, August 2012. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947312000953.

Beavers:2012:BSS

[823] Daniel P. Beavers and James D. Stamey. Bayesian sample size determination for binary regression with a misclassified covariate and no gold standard. Computational Statistics & Data Analysis, 56(8):2574– 2582, August 2012. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947312000965.

Li:2012:ALG

[824] Jianbo Li and Minggao Gu. Adaptive LASSO for general transformation models with right censored data. Computational Statistics & Data Analysis, 56(8):2583– 2597, August 2012. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947312001053.

Chen:2012:SMD

[825] Huaihou Chen, Myunghee Cho Paik, Mandip S. Dhamoon, Yeseon Park Moon, Joshua Willey, Ralph L. Sacco, and Mitchell S. V. Elkind. Semiparametric model for the dichotomized functional outcome after stroke: the Northern Manhattan Study. Computational Statistics & Data Analysis, 56 (8):2598–2608, August 2012. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947312000667.

Anonymous:2012:EBl

[826] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 56(9):iii–v, September 2012. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947312001569.

Anonymous:2012:Cl

[827] Anonymous. Contents. Computational Statistics & Data Analysis, 56 (9):vi–vii, September 2012. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947312001570.

Pasanisi:2012:EDM

[828] Alberto Pasanisi, Shuai Fu, and Nicolas Bousquet. Estimating discrete Markov models from various incomplete data schemes. Computational Statistics & Data Analysis, 56(9):2609– 2625, September 2012. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947312001090.

Kiiveri:2012:FVL

[829] Harri Kiiveri and Frank de Hoog. Fitting very large sparse Gaussian graphical models. Computational Statistics & Data Analysis, 56(9):2626– 2636, September 2012. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947312000746.

Balakrishnan:2012:COP

[830] N. Balakrishnan, K. F. Davies, J. P. Keating, and R. L. Mason. Computation of optimal plotting points based on Pitman closeness with an application to goodness-of-fit for locationscale families. Computational Statistics & Data Analysis, 56(9):2637– 2649, September 2012. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794731200103X.

Fiorentini:2012:MLD

[831] G. Fiorentini, C. Planas, and A. Rossi. The marginal likelihood of dynamic mixture models. Computational Statistics & Data Analysis, 56(9):2650– 2662, September 2012. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947312001338.

Yang:2012:BVS

[832] Mingan Yang. Bayesian variable selection for logistic mixed model with nonparametric random effects. Computational Statistics & Data Analysis, 56(9):2663–2674, September 2012. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794731100449X.

Yu:2012:SNQ

[833] Lili Yu and Karl E. Peace. Spline nonparametric quasi-likelihood regression within the frame of the accelerated failure time model. Computational Statistics & Data Analysis, 56(9):2675– 2687, September 2012. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947312000795.

Martin-Fernandez:2012:MBR

[834] J. A. Martín-Fernández, K. Hron, M. Templ, P. Filzmoser, and J. Palarea-Albaladejo. Model-based replacement of rounded zeros in compositional data: Classical and robust approaches. Computational Statistics & Data Analysis, 56(9):2688– 2704, September 2012. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947312000941.

Suesse:2012:MHE

[835] Thomas Suesse and Ivy Liu. Mantel– Haenszel estimators of odds ratios for stratified dependent binomial data. Computational Statistics & Data Analysis, 56(9):2705– 2717, September 2012. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947312000977.

Ma:2012:IBC

[836] Shuangge Ma, Ying Dai, Jian Huang, and Yang Xie. Identification of breast cancer prognosis markers via integrative analysis. Computational Statistics & Data Analysis, 56(9):2718– 2728, September 2012. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947312000990.

Wang:2012:SRN

[837] J. Wang and S. K. Ghosh. Shape restricted nonparametric regression with Bernstein polynomials. Computational Statistics & Data Analysis, 56(9):2729– 2741, September 2012. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947312001004.

Chakraborty:2012:BMR

[838] Sounak Chakraborty. Bayesian multiple response kernel regression model for high dimensional data and its practical applications in near infrared spectroscopy. Computational Statistics & Data Analysis, 56(9):2742– 2755, September 2012. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947312001016.

Kim:2012:RAU

[839] Gunky Kim and Raymond Chambers. Regression analysis under incomplete linkage. Computational Statistics & Data Analysis, 56(9):2756– 2770, September 2012. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947312001089.

Al-Sharadqah:2012:DOE

[840] A. Al-Sharadqah and N. Chernov. A doubly optimal ellipse fit. Computational Statistics & Data Analysis, 56(9):2771–2781, September 2012. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0167947312001107.

Sajobi:2012:RDD

[841] Tolulope T. Sajobi, Lisa M. Lix, Bolanle M. Dansu, William Laverty, and Longhai Li. Robust descriptive discriminant analysis for repeated measures data. Computational Statistics & Data Analysis, 56(9):2782– 2794, September 2012. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947312001119.

Kim:2012:DMD

[842] Jaejik Kim and L. Billard. Dissimilarity measures and divisive clustering for symbolic multimodalvalued data. Computational Statistics & Data Analysis, 56(9):2795– 2808, September 2012. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947312001132.

Turek:2012:MAW

[843] Daniel Turek and David Fletcher. Model-averaged Wald confidence intervals. Computational Statistics & Data Analysis, 56(9):2809–2815, September 2012. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947312001144.

Lee:2012:EAM

[844] Gyemin Lee and Clayton Scott. EM algorithms for multivariate Gaussian mixture models with truncated and censored data. Computational Statistics & Data Analysis, 56(9):2816– 2829, September 2012. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947312001156.

Kung:2012:SVS

[845] Yi-Hung Kung, Chang-Ting Lin, and Yu-Shan Shih. Split variable selection for tree modeling on rank data. Computational Statistics & Data Analysis, 56(9):2830– 2836, September 2012. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947312001168.

Anonymous:2012:EBb

[846] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 56(10):iii–v, October 2012. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947312001958.

Anonymous:2012:Cb

[847] Anonymous. Contents. Computational Statistics & Data Analysis, 56 (10):vi–vii, October 2012. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794731200196X.

Chambers:2012:SIS

[848] Ray Chambers, J. N. K. Rao, Domingo Morales, and María Dolores Ugarte. Special issue on small area estimation. Computational Statistics & Data Analysis, 56(10):2837– 2839, October 2012. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947312001892.

Esteban:2012:SAE

[849] M. D. Esteban, D. Morales, A. Pérez, and L. Santamaría. Small area estimation of poverty proportions under area-level time models. Computational Statistics & Data Analysis, 56(10): 2840–2855, October 2012. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947311003720.

Burgard:2012:MUD

[850] Jan Pablo Burgard and Ralf T. Münnich. Modelling over and undercounts for design-based Monte Carlo studies in small area estimation: an application to the German registerassisted census. Computational Statistics & Data Analysis, 56(10):2856– 2863, October 2012. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310004305.

Ferraz:2012:SAE

[851] V. R. S. Ferraz and F. A. S. Moura. Small area estimation using skew normal models. Computational Statistics & Data Analysis, 56(10):2864– 2874, October 2012. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947311002593.

Chandra:2012:SAE

[852] Hukum Chandra, Nicola Salvati, Ray Chambers, and Nikos Tzavidis. Small area estimation under spatial nonstationarity. Computational Statistics & Data Analysis, 56(10):2875– 2888, October 2012. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947312000734.

Marchetti:2012:NPB

[853] Stefano Marchetti, Nikos Tzavidis, and Monica Pratesi. Non-parametric bootstrap mean squared error estimation for M-quantile estimators of small area averages, quantiles and poverty indicators. Computational Statistics & Data Analysis, 56(10): 2889–2902, October 2012. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947312000631.

Lombardia:2012:NCS

[854] María José Lombardía and Stefan Sperlich. A new class of semi-mixed effects models and its application in small area estimation. Computational Statistics & Data Analysis, 56(10): 2903–2917, October 2012. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947312000357.

Krieg:2012:EMU

[855] Sabine Krieg and Jan A. van den Brakel. Estimation of the monthly unemployment rate for six domains through structural time series modelling with cointegrated trends. Computational Statistics & Data Analysis, 56(10):2918–2933, October 2012. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0167947312000783.

Militino:2012:EPF

[856] A. F. Militino, T. Goicoa, and M. D. Ugarte. Estimating the percentage of food expenditure in small areas using bias-corrected P -spline based estimators. Computational Statistics & Data Analysis, 56(10):2934– 2948, October 2012. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794731200028X.

Berg:2012:EEC

[857] Emily J. Berg and Wayne A. Fuller. Estimators of error covariance matrices for small area prediction. Computational Statistics & Data Analysis, 56(10):2949– 2962, October 2012. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947312001120.

Maringer:2012:SIO

[858] Dietmar Maringer, Sandra Paterlini, and Peter Winker. The 3rd special issue on optimization heuristics in estimation and modelling problems. Computational Statistics & Data Analysis, 56(10):2963– 2964, October 2012. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947312001909.

Wolters:2012:PSA

[859] Mark A. Wolters. A particle swarm algorithm with broad applicability in shape-constrained estimation. Computational Statistics & Data Analysis, 56(10):2965– 2975, October 2012. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947311004063.

She:2012:IAF

[860] Yiyuan She. An iterative algorithm for fitting nonconvex penalized generalized linear models with grouped predictors. Computational Statistics & Data Analysis, 56(10):2976– 2990, October 2012. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947311004105.

Anonymous:2012:EBc

[861] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 56(11):iii–v, November 2012. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947312002290.

Anonymous:2012:Cc

[862] Anonymous. Contents. Computational Statistics & Data Analysis, 56 (11):vi–ix, November 2012. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947312002307.

Belsley:2012:ACF

[863] D. A. Belsley, E. J. Kontoghiorghes, H. K. Van Dijk, Luc Bauwens, David A. Belsley, Erricos John Kontoghiorghes, Siem Jan Koopman, Michael McAleer, Herman K. van Dijk, Alessandra Amendola, Monica Billio, Christophe Croux, Cathy W. S. Chen, Russell Davidson, Pierre Duchesne, Paolo Foschi, Christian Francq, Ana-Maria Fuertes, Gary Koop, Lynda Khalaf, Marc Paolella, et al. The Annals of Computational and Financial Econometrics, first issue. Computational Statistics & Data Analysis, 56(11): 2991–2992, November 2012. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947312001648.

Boudt:2012:JRD

[864] Kris Boudt, Jonathan Cornelissen, and Christophe Croux. Jump robust daily covariance estimation by disentangling variance and correlation components. Computational Statistics & Data Analysis, 56(11):2993– 3005, November 2012. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947311001691.

Chan:2012:BCA

[865] J. S. K. Chan, C. P. Y. Lam, P. L. H. Yu, S. T. B. Choy, and C. W. S. Chen. A Bayesian conditional autoregressive geometric process model for range data. Computational Statistics & Data Analysis, 56(11):3006– 3019, November 2012. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947311000089.

DeWachter:2012:DSB

[866] Stefan De Wachter and Elias Tzavalis. Detection of structural breaks in linear dynamic panel data models. Computational Statistics & Data Analysis, 56(11):3020–3034, November 2012. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947312001077.

Deschamps:2012:BEG

[867] Philippe J. Deschamps. Bayesian estimation of generalized hyperbolic skewed student GARCH models. Computational Statistics & Data Analysis, 56(11):3035–3054, November 2012. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947311003926.

Fok:2012:MDE

[868] Dennis Fok, Richard Paap, and Philip Hans Franses. Modeling dynamic effects of promotion on interpurchase times. Computational Statistics & Data Analysis, 56(11):3055– 3069, November 2012. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794731200148X.

Fossati:2012:CUR

[869] Sebastian Fossati. Covariate unit root tests with good size and power. Computational Statistics & Data Analysis, 56(11):3070–3079, November 2012. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947311001988.

Fried:2012:OEL

[870] Roland Fried. On the online estimation of local constant volatilities. Computational Statistics & Data Analysis, 56(11):3080–3090, November 2012. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947311000715.

Bocart:2012:EAV

[871] Fabian Y. R. P. Bocart and Christian M. Hafner. Econometric analysis of volatile art markets. Computational Statistics & Data Analysis, 56(11): 3091–3104, November 2012. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947311003902.

Kleppe:2012:FGS

[872] Tore Selland Kleppe and Hans Julius Skaug. Fitting general stochastic volatility models using Laplace accelerated sequential importance sampling. Computational Statistics & Data Analysis, 56(11):3105– 3119, November 2012. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947311001733.

Gefang:2012:DUU

[873] Deborah Gefang, Gary Koop, and Simon M. Potter. The dynamics of UK and US inflation expectations. Computational Statistics & Data Analysis, 56(11):3120– 3133, November 2012. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947311002623.

Dordonnat:2012:DFP

[874] Virginie Dordonnat, Siem Jan Koopman, and Marius Ooms. Dynamic factors in periodic time-varying regressions with an application to hourly electricity load modelling. Computational Statistics & Data Analysis, 56(11): 3134–3152, November 2012. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947311001290.

Krishnakumar:2012:ESG

[875] Jaya Krishnakumar, Andi Kabili, and Ilir Roko. Estimation of SEM with GARCH errors. Computational Statistics & Data Analysis, 56(11):3153– 3181, November 2012. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947312001326.

Lo:2012:RSS

[876] Serigne N. Lô and Elvezio Ronchetti. Robust small sample accurate inference in moment condition models. Computational Statistics & Data Analysis, 56(11):3182–3197, November 2012. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947311000399.

Luger:2012:FSB

[877] Richard Luger. Finite-sample bootstrap inference in GARCH models with heavy-tailed innovations. Computational Statistics & Data Analysis, 56(11):3198–3211, November 2012. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0167947311003070.

Martinez-Sanchis:2012:CDW

[878] Elena Martinez-Sanchis, Juan Mora, and Ilker Kandemir. Counterfactual distributions of wages via quantile regression with endogeneity. Computational Statistics & Data Analysis, 56(11):3212–3229, November 2012. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947312001065.

Blueschke-Nikolaeva:2012:OCN

[879] V. Blueschke-Nikolaeva, D. Blueschke, and R. Neck. Optimal control of nonlinear dynamic econometric models: an algorithm and an application. Computational Statistics & Data Analysis, 56(11):3230– 3240, November 2012. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947311001095.

Nakajima:2012:GEV

[880] Jouchi Nakajima, Tsuyoshi Kunihama, Yasuhiro Omori, and Sylvia Frühwirth-Schnatter. Generalized extreme value distribution with time-dependence using the AR and MA models in state space form. Computational Statistics & Data Analysis, 56(11):3241– 3259, November 2012. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947311001460.

Raknerud:2012:IIM

[881] Arvid Raknerud and Øivind Skare. Indirect inference methods for stochastic volatility models based on non-Gaussian Ornstein–Uhlenbeck processes. Computational Statistics & Data Analysis, 56(11):3260–3275, November 2012. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947311000302.

Bellini:2012:RAD

[882] Tiziano Bellini and Marco Riani. Robust analysis of default intensity. Computational Statistics & Data Analysis, 56(11):3276–3285, November 2012. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947311000983.

Trapani:2012:ATL

[883] Lorenzo Trapani. On the asymptotic t-test for large nonstationary panel models. Computational Statistics & Data Analysis, 56(11):3286– 3306, November 2012. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947311000831.

Belsley:2012:SSI

[884] David A. Belsley, Cathy W. S. Chen, Christian Francq, Giampiero Gallo, Lynda Khalaf, Erricos John Kontoghiorghes, and Herman K. van Dijk. The sixth special issue on computational econometrics. Computational Statistics & Data Analysis, 56(11): 3307–3308, November 2012. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016794731200165X.

Ahmed:2012:ASE

[885] S. Ejaz Ahmed and Christopher J. Nicol. An application of shrinkage estimation to the nonlinear regression model. Computational Statistics & Data Analysis, 56(11):3309– 3321, November 2012. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947310003099.

Alexeev:2012:LLC

[886] Vitali Alexeev and Alex Maynard. Localized level crossing random walk test robust to the presence of structural breaks. Computational Statistics & Data Analysis, 56(11):3322– 3344, November 2012. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310002719.

Tsay:2012:BPD

[887] Ruey S. Tsay and Tomohiro Ando. Bayesian panel data analysis for exploring the impact of subprime financial crisis on the US stock market. Computational Statistics & Data Analysis, 56(11):3345–3365, November 2012. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310004561.

Santos:2012:NCI

[888] P. Araújo Santos and M. I. Fraga Alves. A new class of independence tests for interval forecasts evaluation. Computational Statistics & Data Analysis, 56(11):3366– 3380, November 2012. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310003762.

Baltagi:2012:FSP

[889] Badi H. Baltagi, Georges Bresson, and Alain Pirotte. Forecasting with spatial panel data. Computational Statistics & Data Analysis, 56(11): 3381–3397, November 2012. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794731000321X.

Ardia:2012:CSM

[890] David Ardia, Nalan Baştürk, Lennart Hoogerheide, and Herman K. van Dijk. A comparative study of Monte Carlo methods for efficient evaluation of marginal likelihood. Computational Statistics & Data Analysis, 56(11): 3398–3414, November 2012. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947310003440.

Bauwens:2012:MLC

[891] Luc Bauwens and Jeroen V. K. Rombouts. On marginal likelihood computation in change-point models. Computational Statistics & Data Analysis, 56(11):3415–3429, November 2012. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947310002707.

Beran:2012:RTI

[892] Jan Beran and Dieter Schell. On robust tail index estimation. Computational Statistics & Data Analysis, 56(11): 3430–3443, November 2012. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947310002392.

Bravo:2012:EBW

[893] Francesco Bravo and Federico Crudu. Efficient bootstrap with weakly dependent processes. Computational Statistics & Data Analysis, 56(11): 3444–3458, November 2012. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310003087.

Caporin:2012:MFW

[894] Massimiliano Caporin and Juliusz Preś. Modelling and forecasting wind speed intensity for weather risk management. Computational Statistics & Data Analysis, 56(11):3459– 3476, November 2012. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310002641.

Chan:2012:SMC

[895] Ngai Hang Chan, Hoi Ying Wong, and Jing Zhao. Structural model of credit migration. Computational Statistics & Data Analysis, 56(11): 3477–3490, November 2012. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310004032.

Gallegati:2012:WBA

[896] Marco Gallegati. A wavelet-based approach to test for financial market contagion. Computational Statistics & Data Analysis, 56(11):3491– 3497, November 2012. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310004317.

Chen:2012:BVR

[897] Qian Chen, Richard Gerlach, and Zudi Lu. Bayesian value-at-risk and expected shortfall forecasting via the asymmetric Laplace distribution. Computational Statistics & Data Analysis, 56(11):3498– 3516, November 2012. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794731000263X.

Billio:2012:DRE

[898] Monica Billio, Mila Getmansky, and Loriana Pelizzon. Dynamic risk exposures in hedge funds. Computational Statistics & Data Analysis, 56(11): 3517–3532, November 2012. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310003439.

Hafner:2012:EDC

[899] Christian M. Hafner and Olga Reznikova. On the estimation of dynamic conditional correlation models. Computational Statistics & Data Analysis, 56(11):3533–3545, November 2012. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310003658.

Guidolin:2012:SVC

[900] Massimo Guidolin and Stuart Hyde. Simple VARs cannot approximate Markov switching asset allocation decisions: an out-of-sample assessment. Computational Statistics & Data Analysis, 56(11):3546– 3566, November 2012. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310003804.

Kiviet:2012:CAE

[901] Jan F. Kiviet and Jerzy Niemczyk. Comparing the asymptotic and empirical (un)conditional distributions of OLS and IV in a linear static simultaneous equation. Computational Statistics & Data Analysis, 56(11): 3567–3586, November 2012. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310003154.

Klar:2012:STE

[902] B. Klar, F. Lindner, and S. G. Meintanis. Specification tests for the error distribution in GARCH models. Computational Statistics & Data Analysis, 56(11):3587–3598, November 2012. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310002409.

Kotchoni:2012:ACF

[903] Rachidi Kotchoni. Applications of the characteristic function-based continuum GMM in finance. Computational Statistics & Data Analysis, 56(11): 3599–3622, November 2012. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310003269.

Davies:2012:RCP

[904] Laurie Davies, Christian Höhenrieder, and Walter Krämer. Recursive computation of piecewise constant volatilities. Computational Statistics & Data Analysis, 56(11):3623– 3631, November 2012. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310002720.

Kwan:2012:EDC

[905] Wilson Kwan, Wai Keung Li, and Guodong Li. On the estimation and diagnostic checking of the ARFIMA– HYGARCH model. Computational Statistics & Data Analysis, 56(11): 3632–3644, November 2012. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310002835.

Li:2012:SET

[906] Junye Li, Carlo Favero, and Fulvio Ortu. A spectral estimation of tempered stable stochastic volatility models and option pricing. Computational Statistics & Data Analysis, 56(11): 3645–3658, November 2012. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794731000441X.

Nikoloulopoulos:2012:VCA

[907] Aristidis K. Nikoloulopoulos, Harry Joe, and Haijun Li. Vine copulas with asymmetric tail dependence and applications to financial return data. Computational Statistics & Data Analysis, 56(11):3659– 3673, November 2012. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310002951.

Ishihara:2012:EBE

[908] Tsunehiro Ishihara and Yasuhiro Omori. Efficient Bayesian estimation of a multivariate stochastic volatility model with cross leverage and heavy-tailed errors. Computational Statistics & Data Analysis, 56(11): 3674–3689, November 2012. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310002884.

Nakajima:2012:SVM

[909] Jouchi Nakajima and Yasuhiro Omori. Stochastic volatility model with leverage and asymmetrically heavy-tailed error using GH skew Student’s tdistribution. Computational Statistics & Data Analysis, 56(11):3690– 3704, November 2012. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310002859.

Kiviet:2012:HOA

[910] Jan F. Kiviet and Garry D. A. Phillips. Higher-order asymptotic expansions of the least-squares estimation bias in first-order dynamic regression models. Computational Statistics & Data Analysis, 56(11):3705– 3729, November 2012. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310002860.

Raggi:2012:LMN

[911] Davide Raggi and Silvano Bordignon. Long memory and nonlinearities in realized volatility: a Markov switching approach. Computational Statistics & Data Analysis, 56(11):3730– 3742, November 2012. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947310004767.

Peters:2012:LFB

[912] G. W. Peters, S. A. Sisson, and Y. Fan. Likelihood-free Bayesian inference for α-stable models. Computational Statistics & Data Analysis, 56(11):3743– 3756, November 2012. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947310003786.

Carta:2012:MMO

[913] Alessandro Carta and Mark F. J. Steel. Modelling multi-output stochastic frontiers using copulas. Computational Statistics & Data Analysis, 56(11): 3757–3773, November 2012. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016794731000280X.

Hautsch:2012:BIS

[914] Nikolaus Hautsch and Fuyu Yang. Bayesian inference in a stochastic volatility Nelson–Siegel model. Computational Statistics & Data Analysis, 56(11):3774–3792, November 2012. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0167947310002768.

Huurman:2012:PW

[915] Christian Huurman, Francesco Ravazzolo, and Chen Zhou. The power of weather. Computational Statistics & Data Analysis, 56(11):3793– 3807, November 2012. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947310002665.

Anonymous:2012:EBd

[916] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 56(12):iii–v, December 2012. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794731200271X.

Anonymous:2012:Cd

[917] Anonymous. Contents. Computational Statistics & Data Analysis, 56 (12):vi–ix, December 2012. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947312002721.

Manolopoulou:2012:MGW

[918] Ioanna Manolopoulou, Thomas B. Kepler, and Daniel M. Merl. Mixtures of Gaussian wells: Theory, computation, and application. Computational Statistics & Data Analysis, 56(12): 3809–3820, December 2012. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947312001740.

Teuber:2012:NSM

[919] T. Teuber and A. Lang. A new similarity measure for nonlocal filtering in the presence of multiplicative noise. Computational Statistics & Data Analysis, 56(12):3821– 3842, December 2012. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947312002083.

OHagan:2012:CAF

[920] Adrian O’Hagan, Thomas Brendan Murphy, and Isobel Claire Gormley. Computational aspects of fitting mixture models via the expectation– maximization algorithm. Computational Statistics & Data Analysis, 56(12):3843–3864, December 2012. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0167947312002101.

Kitsche:2012:UHC

[921] A. Kitsche, L. A. Hothorn, and F. Schaarschmidt. The use of historical controls in estimating simultaneous confidence intervals for comparisons against a concurrent control. Computational Statistics & Data Analysis, 56(12):3865–3875, December 2012. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947312002095.

Soni:2012:NEQ

[922] Pooja Soni, Isha Dewan, and Kanchan Jain. Nonparametric estimation of quantile density function. Computational Statistics & Data Analysis, 56(12):3876–3886, December 2012. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947312001788.

Azevedo:2012:SCP

[923] Cecilia Azevedo, Víctor Leiva, Emilia Athayde, and N. Balakrishnan. Shape and change point analyses of the Birnbaum–Saunders-t hazard rate and associated estimation. Computational Statistics & Data Analysis, 56(12): 3887–3897, December 2012. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947312002046.

Oliveira:2012:PRB

[924] M. Oliveira, R. M. Crujeiras, and A. Rodríguez-Casal. A plug-in rule for bandwidth selection in circular density estimation. Computational Statistics & Data Analysis, 56(12): 3898–3908, December 2012. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947312002204.

Zhou:2012:EVM

[925] Hua Zhou and Yiwen Zhang. EM vs MM: a case study. Computational Statistics & Data Analysis, 56(12): 3909–3920, December 2012. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947312002174.

Salcedo:2012:CNS

[926] Gladys E. Salcedo, Rogério F. Porto, and Pedro A. Morettin. Comparing non-stationary and irregularly spaced time series. Computational Statistics & Data Analysis, 56(12):3921– 3934, December 2012. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947312002216.

Hooker:2012:LLB

[927] Giles Hooker and James O. Ramsay. Learned-loss boosting. Computational Statistics & Data Analysis, 56(12): 3935–3944, December 2012. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947312002186.

Cheng:2012:BSS

[928] Chin-I. Cheng and Paul L. Speckman. Bayesian smoothing spline analysis of variance. Computational Statistics & Data Analysis, 56(12):3945– 3958, December 2012. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947312002198.

Paulo:2012:CCM

[929] Rui Paulo, Gonzalo García-Donato, and Jesús Palomo. Calibration of computer models with multivariate output. Computational Statistics & Data Analysis, 56(12):3959– 3974, December 2012. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947312002228.

Zhang:2012:IRT

[930] Shulin Zhang, Peter X.-K. Song, Daimin Shi, and Qian M. Zhou. Information ratio test for model misspecification on parametric structures in stochastic diffusion models. Computational Statistics & Data Analysis, 56(12):3975–3987, December 2012. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947312002125.

Bermudez:2012:FMB

[931] Lluís Bermúdez and Dimitris Karlis. A finite mixture of bivariate Poisson regression models with an application to insurance ratemaking. Computational Statistics & Data Analysis, 56(12): 3988–3999, December 2012. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947312002150.

Gottlieb:2012:SCL

[932] Andrea Gottlieb and Hans-Georg Müller. A stickiness coefficient for longitudinal data. Computational Statistics & Data Analysis, 56(12):4000– 4010, December 2012. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947312001351.

Balakrishnan:2012:LTR

[933] N. Balakrishnan and Debanjan Mitra. Left truncated and right censored Weibull data and likelihood inference with an illustration. Computational Statistics & Data Analysis, 56(12): 4011–4025, December 2012. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947312001880.

Shao:2012:SAH

[934] Wei Shao and Yijun Zuo. Simulated annealing for higher dimensional projection depth. Computational Statistics & Data Analysis, 56(12):4026– 4036, December 2012. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947312001843.

Lu:2012:ADM

[935] Zhaosong Lu, Ting Kei Pong, and Yong Zhang. An alternating direction method for finding Dantzig selectors. Computational Statistics & Data Analysis, 56(12):4037– 4046, December 2012. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947312001867.

Mahmoudi:2012:GEP

[936] Eisa Mahmoudi and Ali Akbar Jafari. Generalized exponential–power series distributions. Computational Statistics & Data Analysis, 56(12): 4047–4066, December 2012. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947312001727.

Bartolucci:2012:BIT

[937] Francesco Bartolucci, Luisa Scaccia, and Alessio Farcomeni. Bayesian inference through encompassing priors and importance sampling for a class of marginal models for categorical data. Computational Statistics & Data Analysis, 56(12):4067– 4080, December 2012. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947312001661.

Schaumburg:2012:PEV

[938] Julia Schaumburg. Predicting extreme value at risk: Nonparametric quantile regression with refinements from extreme value theory. Computational Statistics & Data Analysis, 56(12): 4081–4096, December 2012. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947312001429.

Chung:2012:BAM

[939] Hwan Chung and Hsiu-Ching Chang. Bayesian approaches to the model selection problem in the analysis of latent stage-sequential process. Computational Statistics & Data Analysis, 56(12):4097–4110, December 2012. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947312001417.

Smucker:2012:MRD

[940] Byran J. Smucker, Enrique del Castillo, and James L. Rosenberger. Modelrobust designs for split-plot experiments. Computational Statistics & Data Analysis, 56(12):4111– 4121, December 2012. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947312001363.

Berlinet:2012:AEA

[941] A. F. Berlinet and Ch. Roland. Acceleration of the EM algorithm: P-EM versus epsilon algorithm. Computational Statistics & Data Analysis, 56(12): 4122–4137, December 2012. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947312001314.

Almendra-Arao:2012:ECT

[942] Félix Almendra-Arao. Efficient calculation of test sizes for noninferiority. Computational Statistics & Data Analysis, 56(12):4138– 4145, December 2012. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947311004051.

Hong:2012:BBH

[943] Zhaoping Hong and Heng Lian. BOPA: a Bayesian hierarchical model for outlier expression detection. Computational Statistics & Data Analysis, 56(12):4146–4156, December 2012. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0167947312001855.

Kwong:2012:ODV

[944] Koon Shing Kwong and Siu Hung Cheung. Optimal design for p-value consistent step-up procedures for multiple comparisons with a control in directionmixed families. Computational Statistics & Data Analysis, 56(12):4157– 4164, December 2012. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947312001831.

Wang:2012:BMR

[945] Jianqiang C. Wang and Scott H. Holan. Bayesian multi-regime smooth transition regression with ordered categorical variables. Computational Statistics & Data Analysis, 56(12):4165– 4179, December 2012. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794731200182X.

Cox:2012:MBE

[946] Christopher Cox and Xiuhong Li. Model-based estimation of the attributable risk: a loglinear approach. Computational Statistics & Data Analysis, 56(12):4180–4189, December 2012. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947312001818.

Song:2012:BAG

[947] Xin-Yuan Song, Nian-Sheng Tang, and Sy-Miin Chow. A Bayesian approach for generalized random coefficient structural equation models for longitudinal data with adjacent time effects. Computational Statistics & Data Analysis, 56(12):4190– 4203, December 2012. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947312001806.

Li:2012:DCS

[948] Ai-Ping Li and Feng-Chang Xie. Diagnostics for a class of survival regression models with heavytailed errors. Computational Statistics & Data Analysis, 56(12):4204– 4214, December 2012. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794731200179X.

Andrianakis:2012:ENG

[949] Ioannis Andrianakis and Peter G. Challenor. The effect of the nugget on Gaussian process emulators of computer models. Computational Statistics & Data Analysis, 56(12):4215– 4228, December 2012. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947312001879.

Xu:2012:MIV

[950] Hai-Yan Xu, Min Xie, Thong Ngee Goh, and Xiuju Fu. A model for integer-valued time series with conditional overdispersion. Computational Statistics & Data Analysis, 56(12): 4229–4242, December 2012. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947312001752.

Katsikatsou:2012:PLE

[951] Myrsini Katsikatsou, Irini Moustaki, Fan Yang-Wallentin, and Karl G. Jöreskog. Pairwise likelihood estimation for factor analysis models with ordinal data. Computational Statistics & Data Analysis, 56(12):4243– 4258, December 2012. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947312001739.

Torabi:2012:LIG

[952] Mahmoud Torabi. Likelihood inference in generalized linear mixed models with two components of dispersion using data cloning. Computational Statistics & Data Analysis, 56(12): 4259–4265, December 2012. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947312001715.

Hsieh:2012:TTR

[953] Ping-Hung Hsieh. Tales from the tail: Robust estimation of moments of environmental data with one-sided detection limits. Computational Statistics & Data Analysis, 56(12):4266– 4277, December 2012. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947312001703.

Huang:2012:EEW

[954] Wenpo Huang, Lianjie Shu, and Wei Jiang. Evaluation of exponentially weighted moving variance control chart subject to linear drifts. Computational Statistics & Data Analysis, 56(12): 4278–4289, December 2012. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947312001776.

Zhao:2012:SLD

[955] Jianhua Zhao, Philip L. H. Yu, Lei Shi, and Shulan Li. Separable linear discriminant analysis. Computational Statistics & Data Analysis, 56(12): 4290–4300, December 2012. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947312001636.

Grace:2012:ATA

[956] Adam W. Grace and Ian A. Wood. Approximating the tail of the Anderson– Darling distribution. Computational Statistics & Data Analysis, 56(12): 4301–4311, December 2012. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794731200151X.

Li:2012:ISS

[957] Haifen Li, Jiajia Zhang, and Yincai Tang. Induced smoothing for the semiparametric accelerated hazards model. Computational Statistics & Data Analysis, 56(12):4312– 4319, December 2012. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947312001491.

Wang:2012:GIE

[958] Bing Xing Wang. Generalized interval estimation for the Birnbaum– Saunders distribution. Computational Statistics & Data Analysis, 56(12): 4320–4326, December 2012. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947312001508.

Potgieter:2012:NEL

[959] C. J. Potgieter and F. Lombard. Nonparametric estimation of location and scale parameters. Computational Statistics & Data Analysis, 56(12): 4327–4337, December 2012. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947312001478.

Castro:2012:EHN

[960] Luis M. Castro, Héctor W. Gómez, and Maria Valenzuela. Epsilon half-normal model: Properties and inference. Computational Statistics & Data Analysis, 56(12):4338– 4347, December 2012. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947312001466.

Tang:2012:SBA

[961] Nian-Sheng Tang and Xing-De Duan. A semiparametric Bayesian approach to generalized partial linear mixed models for longitudinal data. Computational Statistics & Data Analysis, 56(12):4348–4365, December 2012. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0167947312001442.

Withers:2012:ICR

[962] Christopher S. Withers and Saralees Nadarajah. Improved confidence regions based on Edgeworth expansions. Computational Statistics & Data Analysis, 56(12):4366– 4380, December 2012. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947312001454.

Nixon:2012:IRS

[963] John H. Nixon. Investigations into refinements of Storey’s method of multiple hypothesis testing minimising the FDR, and its application to test binomial data. Computational Statistics & Data Analysis, 56(12):4381– 4398, December 2012. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947312001697.

Azevedo:2012:BGM

[964] Caio L. N. Azevedo, Dalton F. Andrade, and Jean-Paul Fox. A Bayesian generalized multiple group IRT model with model-fit assessment tools. Computational Statistics & Data Analysis, 56(12):4399– 4412, December 2012. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947312001430.

Huang:2012:EIV

[965] Zhensheng Huang. Efficient inferences on the varying-coefficient single-index model with empirical likelihood. Computational Statistics & Data Analysis, 56(12):4413– 4420, December 2012. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947312001673.

Lim:2012:PES

[966] Johan Lim, Kiseop Lee, Donghyeon Yu, Haiyan Liu, and Michael Sherman. Parameter estimation in the spatial auto-logistic model with working independent subblocks. Computational Statistics & Data Analysis, 56(12): 4421–4432, December 2012. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947312001399.

Doyen:2012:RAJ

[967] L. Doyen. Reliability analysis and joint assessment of Brown–Proschan preventive maintenance efficiency and intrinsic wear-out. Computational Statistics & Data Analysis, 56(12):4433– 4449, December 2012. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947312001387.

Beaumont:2012:BVE

[968] Jean-François Beaumont and Anne-Sophie Charest. Bootstrap variance estimation with survey data when estimating model parameters. Computational Statistics & Data Analysis, 56(12):4450–4461, December 2012. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947312001375.

Ahmed:2012:IMM

[969] Murat O. Ahmed and Guenther Walther. Investigating the multimodality of multivariate data with principal curves. Computational Statistics & Data Analysis, 56(12):4462– 4469, December 2012. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947312001028.

Kanatani:2012:OCD

[970] Kenichi Kanatani and Hirotaka Niitsuma. Optimal computation of 3-D similarity: Gauss–Newton vs. Gauss–Helmert. Computational Statistics & Data Analysis, 56(12):4470– 4483, December 2012. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947312001405.

Anonymous:2013:EBa

[971] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 57(1):iii–v, January 2013. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947312003295.

Anonymous:2013:EBb

[972] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 57(1):vi–ix, January 2013. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947312003301.

Chee:2013:MQD

[973] Chew-Seng Chee and Yong Wang. Minimum quadratic distance density estimation using nonparametric mixtures. Computational Statistics & Data Analysis, 57(1):1– 16, January 2013. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947312002447.

Chattopadhyay:2013:ICA

[974] Asis Kumar Chattopadhyay, Saptarshi Mondal, and Tanuka Chattopadhyay. Independent component analysis for the objective classification of globular clusters of the galaxy NGC 5128. Computational Statistics & Data Analysis, 57(1):17–32, January 2013. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947312002484.

Zhao:2013:REP

[975] Xingqiu Zhao, Xingwei Tong, and Jianguo Sun. Robust estimation for panel count data with informative observation times. Computational Statistics & Data Analysis, 57 (1):33–40, January 2013. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947312002149.

Kim:2013:EAP

[976] Yongdai Kim, Joungyoun Kim, and Woncheol Jang. An EM algorithm for the proportional hazards model with doubly censored data. Computational Statistics & Data Analysis, 57(1):41–51, January 2013. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947312002411.

Teimouri:2013:SBS

[977] Mahdi Teimouri and Saralees Nadarajah. On simulating Balakrishnan skew-normal variates. Computational Statistics & Data Analysis, 57(1): 52–58, January 2013. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947312002496. See

erratum [1075].

Lee:2013:MET

[978] Sangyeol Lee. A maximum entropy type test of fit: Composite hypothesis case. Computational Statistics & Data Analysis, 57(1): 59–67, January 2013. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947312002460.

Barreto:2013:IBS

[979] Larissa Santana Barreto, Audrey H. M. A. Cysneiros, and Francisco Cribari-Neto. Improved Birnbaum– Saunders inference under type II censoring. Computational Statistics & Data Analysis, 57(1):68– 81, January 2013. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947312002459.

Stover:2013:HEP

[980] Jason H. Stover and Matthew C. Ulm. Hyperparameter estimation and plug-in kernel density estimates for maximum a posteriori land-cover classification with multiband satellite data. Computational Statistics & Data Analysis, 57(1):82– 94, January 2013. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947312002502.

Pacheco:2013:EMV

[981] Joaquín Pacheco, Silvia Casado, and Santiago Porras. Exact methods for variable selection in principal component analysis: Guide functions and pre-selection. Computational Statistics & Data Analysis, 57(1):95– 111, January 2013. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947312002551.

Song:2013:SAD

[982] Peter X.-K. Song, R. Keith Freeland, Atanu Biswas, and Shulin Zhang. Statistical analysis of discretevalued time series using categorical ARMA models. Computational Statistics & Data Analysis, 57(1):112– 124, January 2013. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947312002435.

Kim:2013:REC

[983] Byungsoo Kim and Sangyeol Lee. Robust estimation for the covariance matrix of multivariate time series based on normal mixtures. Computational Statistics & Data Analysis, 57(1):125– 140, January 2013. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947312002538.

Liu:2013:ESC

[984] W. Liu, P. Ah-Kine, F. Bretz, and A. J. Hayter. Exact simultaneous confidence intervals for a finite set of contrasts of three, four or five generally correlated normal means. Computational Statistics & Data Analysis, 57 (1):141–148, January 2013. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947312002472.

Rahman:2013:SCP

[985] Azizur Rahman, Ann Harding, Robert Tanton, and Shuangzhe Liu. Simulating the characteristics of populations at the small area level: New validation techniques for a spatial microsimulation model in Australia. Computational Statistics & Data Analysis, 57 (1):149–165, January 2013. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947312002599.

Balakrishnan:2013:HCM

[986] N. Balakrishnan and Debasis Kundu. Hybrid censoring: Models, inferential results and applications. Computational Statistics & Data Analysis, 57 (1):166–209, January 2013. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947312001685.

Chen:2013:AEC

[987] Chyong-Mei Chen, Tai-Fang C. Lu, and Chao-Min Hsu. Association estimation for clustered failure time data with a cure fraction. Computational Statistics & Data Analysis, 57 (1):210–222, January 2013. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947312002575.

Wang:2013:SSD

[988] Tao Wang and Lixing Zhu. Sparse sufficient dimension reduction using optimal scoring. Computational Statistics & Data Analysis, 57(1):223– 232, January 2013. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947312002563.

Aregay:2013:HBA

[989] Mehreteab Aregay, Ziv Shkedy, and Geert Molenberghs. A hierarchical Bayesian approach for the analysis of longitudinal count data with overdispersion: a simulation study. Computational Statistics & Data Analysis, 57 (1):233–245, January 2013. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947312002617.

Shan:2013:STD

[990] Guogen Shan, Changxing Ma, Alan D. Hutson, and Gregory E. Wilding. Some tests for detecting trends based on the modified Baumgartner–Weiß– Schindler statistics. Computational Statistics & Data Analysis, 57(1):246– 261, January 2013. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947312002514.

Liu:2013:MCQ

[991] Zhenyu Liu, Reza Modarres, and Mengta Yang. A multivariate control quantile test using data depth. Computational Statistics & Data Analysis, 57 (1):262–270, January 2013. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016794731200254X.

Kundu:2013:BEM

[992] Debasis Kundu and Arjun K. Gupta. Bayes estimation for the Marshall– Olkin bivariate Weibull distribution. Computational Statistics & Data Analysis, 57(1):271–281, January 2013. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0167947312002423.

Tian:2013:IDR

[993] Tian Siva Tian and Gareth M. James. Interpretable dimension reduction for classifying functional data. Computational Statistics & Data Analysis, 57 (1):282–296, January 2013. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947312002587.

Wong:2013:TSI

[994] Heung Wong, Riquan Zhang, Bartholomew Leung, and Zhensheng Huang. Testing the significance of index parameters in varying-coefficient singleindex models. Computational Statistics & Data Analysis, 57(1):297– 308, January 2013. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947312002666.

Chen:2013:TFL

[995] Li-Ching Chen and Jiun-Yi Wang. Testing the fit of the logistic model for matched case-control studies. Computational Statistics & Data Analysis, 57 (1):309–319, January 2013. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947312002654.

Drovandi:2013:SMC

[996] Christopher C. Drovandi, James M. McGree, and Anthony N. Pettitt. Sequential Monte Carlo for Bayesian sequentially designed experiments for discrete data. Computational Statistics & Data Analysis, 57(1):320– 335, January 2013. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947312002137.

Park:2013:SIV

[997] Cheolwoo Park and Jib Huh. Statistical inference and visualization in scalespace using local likelihood. Computational Statistics & Data Analysis, 57 (1):336–348, January 2013. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947312002642.

Bruffaerts:2013:RHE

[998] C. Bruffaerts, B. De Rock, and C. Dehon. The robustness of the hyperbolic efficiency estimator. Computational Statistics & Data Analysis, 57 (1):349–363, January 2013. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947312002769.

Horova:2013:FBM

[999] Ivana Horová, Jan Koláček, and Kamila Vopatová. Full bandwidth matrix selectors for gradient kernel density estimate. Computational Statistics & Data Analysis, 57(1):364– 376, January 2013. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947312002782.

Hofer:2013:DMD

[1000] Vera Hofer and Georg Krempl. Drift mining in data: a framework for addressing drift in classification. Computational Statistics & Data Analysis, 57 (1):377–391, January 2013. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947312002812.

Wu:2013:LIE

[1001] Jianmin Wu and Peter M. Bentler.

Limited information estimation in binary factor analysis: a review and extension. Computational Statistics & Data Analysis, 57(1):392– 403, January 2013. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947312002630.

Remenyi:2013:NWS

[1002] Norbert Reményi and Brani Vidakovic. Λ-neighborhood wavelet shrinkage. Computational Statistics & Data Analysis, 57(1):404– 416, January 2013. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947312002824.

Varughese:2013:PEM

[1003] Melvin M. Varughese. Parameter estimation for multivariate diffusion systems. Computational Statistics & Data Analysis, 57(1):417– 428, January 2013. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947312002848.

Rolfs:2013:NLS

[1004] Benjamin T. Rolfs and Bala Rajaratnam. A note on the lack of symmetry in the graphical lasso. Computational Statistics & Data Analysis, 57 (1):429–434, January 2013. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947312002873.

Tang:2013:VSQ

[1005] Yanlin Tang, Huixia Judy Wang, and Zhongyi Zhu. Variable selection in quantile varying coefficient models with longitudinal data. Computational Statistics & Data Analysis, 57 (1):435–449, January 2013. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947312002897.

Matos:2013:IDL

[1006] Larissa A. Matos, Victor H. Lachos, N. Balakrishnan, and Filidor V. Labra. Influence diagnostics in linear and nonlinear mixed-effects models with censored data. Computational Statistics & Data Analysis, 57 (1):450–464, January 2013. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947312002629.

Shao:2013:EPD

[1007] Wei Shao, Guangbao Guo, Fanyu Meng, and Shuqin Jia. An efficient proposal distribution for Metropolis– Hastings using a B-splines technique. Computational Statistics & Data Analysis, 57(1):465–478, January 2013. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0167947312002885.

Lu:2013:HTM

[1008] Zeng-Hua Lu. Halfline tests for multivariate one-sided alternatives. Computational Statistics & Data Analysis, 57 (1):479–490, January 2013. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947312002903.

Chen:2013:VSA

[1009] Xiaolin Chen and Qihua Wang. Variable selection in the additive rate model for recurrent event data. Computational Statistics & Data Analysis, 57 (1):491–503, January 2013. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947312002605.

Mark:2013:EAC

[1010] Brian L. Mark and Yariv Ephraim.

An EM algorithm for continuous-time bivariate Markov chains. Computational Statistics & Data Analysis, 57 (1):504–517, January 2013. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947312002915.

Brilhante:2013:SGH

[1011] M. Fátima Brilhante, M. Ivette Gomes, and Dinis Pestana. A simple generalisation of the Hill estimator. Computational Statistics & Data Analysis, 57 (1):518–535, January 2013. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947312002939.

Aquaro:2013:OSR

[1012] M. Aquaro and P. Čižek. Onestep robust estimation of fixed-effects panel data models. Computational Statistics & Data Analysis, 57(1):536– 548, January 2013. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947312002757.

Torabi:2013:LIG

[1013] Mahmoud Torabi. Likelihood inference in generalized linear mixed measurement error models. Computational Statistics & Data Analysis, 57 (1):549–557, January 2013. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947312002927.

Noh:2013:AMA

[1014] Hohsuk Noh, Anouar El Ghouch, and Ingrid Van Keilegom. Assessing model adequacy in possibly misspecified quantile regression. Computational Statistics & Data Analysis, 57 (1):558–569, January 2013. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947312002940.

Yen:2013:SMM

[1015] Amy Ming-Fang Yen and Hsiu-Hsi Chen. Stochastic models for multiple pathways of temporal natural history on co-morbidity of chronic disease. Computational Statistics & Data Analysis, 57(1):570– 588, January 2013. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947312002836.

Wong:2013:MRI

[1016] Heung Wong, Quanxi Shao, and Wai cheung Ip. Modeling respiratory illnesses with change point: a lesson from the SARS epidemic in Hong Kong. Computational Statistics & Data Analysis, 57(1):589– 599, January 2013. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947312003039.

Janicki:2013:BMA

[1017] Ryan Janicki and Donald Malec. A Bayesian model averaging approach to analyzing categorical data with nonignorable nonresponse. Computational Statistics & Data Analysis, 57 (1):600–614, January 2013. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947312003027.

Lin:2013:SRC

[1018] Huazhen Lin and Heng Peng. Smoothed rank correlation of the linear transformation regression model. Computational Statistics & Data Analysis, 57 (1):615–630, January 2013. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947312002861.

Mkhadri:2013:EVI

[1019] Abdallah Mkhadri and Mohamed Ouhourane. An extended variable inclusion and shrinkage algorithm for correlated variables. Computational Statistics & Data Analysis, 57(1):631– 644, January 2013. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947312002976.

Liu:2013:CBS

[1020] Junfeng Liu and Yi Wang. On Crevecoeur’s bathtub-shaped failure rate model. Computational Statistics & Data Analysis, 57(1):645– 660, January 2013. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947312003052.

Salter-Townshend:2013:VBI

[1021] Michael Salter-Townshend and Thomas Brendan Murphy. Variational Bayesian inference for the latent position cluster model for network data. Computational Statistics & Data Analysis, 57 (1):661–671, January 2013. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947312003076.

Redenbach:2013:PEG

[1022] Claudia Redenbach and Aila Särkkä. Parameter estimation for growth interaction processes using spatio-temporal information. Computational Statistics & Data Analysis, 57(1):672– 683, January 2013. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794731200309X.

Anonymous:2013:EBc

[1023] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 58(??):iii–v, February 2013. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947312003672.

Anonymous:2013:Ca

[1024] Anonymous. Contents. Computational Statistics & Data Analysis, 58 (??):vi–viii, February 2013. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947312003684.

Pollock:2013:TSI

[1025] D. Stephen G. Pollock, Tommaso Proietti, Esther Ruiz, and Howard Weinert. The third special issue on statistical signal extraction and filtering. Computational Statistics & Data Analysis, 58(??):1–3, February 2013. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947312003040.

Thornton:2013:RSU

[1026] Michael A. Thornton. Removing seasonality under a changing regime: Filtering new car sales. Computational Statistics & Data Analysis, 58 (??):4–14, February 2013. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947311002386.

Li:2013:UKS

[1027] Junye Li. An unscented Kalman smoother for volatility extraction: Evidence from stock prices and options. Computational Statistics & Data Analysis, 58(??):15– 26, February 2013. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947311002015.

Wang:2013:IAS

[1028] Ting Wang and Mark Bebbington. Identifying anomalous signals in GPS data using HMMs: an increased likelihood of earthquakes? Computational Statistics & Data Analysis, 58 (??):27–44, February 2013. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947311003501.

Gazeaux:2013:ECP

[1029] Julien Gazeaux, Deborah Batista, Caspar M. Ammann, Philippe Naveau, Cyrille Jégat, and Chaochao Gao. Extracting common pulse-like signals from multiple ice core time series. Computational Statistics & Data Analysis, 58(??):45–57, February 2013. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947312000643.

Prado:2013:SEM

[1030] Raquel Prado. Sequential estimation of mixtures of structured autoregressive models. Computational Statistics & Data Analysis, 58(??):

58–70, February 2013. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947311001083.

Ramirez-Cobo:2013:WBM

[1031] Pepa Ramírez-Cobo and Brani Vidakovic. A 2D wavelet-based multiscale approach with applications to the analysis of digital mammograms. Computational Statistics & Data Analysis, 58(??):71–81, February 2013. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947311003288.

Zamba:2013:TSR

[1032] K. D. Zamba, Panagiotis Tsiamyrtzis, and Douglas M. Hawkins. A threestate recursive sequential Bayesian algorithm for biosurveillance. Computational Statistics & Data Analysis, 58 (??):82–97, February 2013. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947311001447.

Poskitt:2013:DLD

[1033] D. S. Poskitt and Arivalzahan Sengarapillai. Description length and dimensionality reduction in functional data analysis. Computational Statistics & Data Analysis, 58(??):98– 113, February 2013. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947311001113.

Moghtaderi:2013:TFE

[1034] Azadeh Moghtaderi, Patrick Flandrin, and Pierre Borgnat. Trend filtering via empirical mode decompositions. Computational Statistics & Data Analysis, 58(??):114– 126, February 2013. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947311001812.

Frei:2013:MEK

[1035] Marco Frei and Hans R. Künsch. Mixture ensemble Kalman filters. Computational Statistics & Data Analysis, 58 (??):127–138, February 2013. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947311001423.

Navarro-Moreno:2013:WLP

[1036] Jesús Navarro-Moreno, Javier Moreno-Kaiser, Rosa María Fernández-Alcalá, and Juan Carlos Ruiz-Molina. Widely linear prediction for transfer function models based on the infinite past. Computational Statistics & Data Analysis, 58(??):139– 146, February 2013. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947310004482.

Nandi:2013:NSD

[1037] Swagata Nandi, Debasis Kundu, and Rajesh Kumar Srivastava. Noise space decomposition method for twodimensional sinusoidal model. Computational Statistics & Data Analysis, 58 (??):147–161, February 2013. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947311000818.

Biernacki:2013:GMR

[1038] Christophe Biernacki and Julien Jacques. A generative model for rank data based on insertion sort algorithm. Computational Statistics & Data Analysis, 58(??):162– 176, February 2013. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947312003118.

Kim:2013:CBN

[1039] Jongphil Kim. The computation of bivariate normal and t probabilities, with application to comparisons of three normal means. Computational Statistics & Data Analysis, 58(??): 177–186, February 2013. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947312003180.

Rimstad:2013:APD

[1040] Kjartan Rimstad and Henning Omre.

Approximate posterior distributions for convolutional two-level hidden Markov models. Computational Statistics & Data Analysis, 58(??):187– 200, February 2013. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947312003210.

Yi:2013:ESP

[1041] Yanqing Yi. Exact statistical power for response adaptive designs. Computational Statistics & Data Analysis, 58 (??):201–209, February 2013. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947312003234.

Nagatsuka:2013:CME

[1042] Hideki Nagatsuka, Toshinari Kamakura, and N. Balakrishnan. A consistent method of estimation for the three-parameter Weibull distribution. Computational Statistics & Data Analysis, 58(??):210– 226, February 2013. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947312003349.

Su:2013:DCE

[1043] J. Su, A. Srivastava, and F. W. Huffer. Detection, classification and estimation of individual shapes in 2D and 3D point clouds. Computational Statistics & Data Analysis, 58(??): 227–241, February 2013. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947312003374.

Zevallos:2013:MDE

[1044] Mauricio Zevallos and Wilfredo Palma. Minimum distance estimation of ARFIMA processes. Computational Statistics & Data Analysis, 58(??): 242–256, February 2013. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947312003088.

Ye:2013:MTD

[1045] Fei Ye, Chen Yue, and Ying Yang. Modeling time-dependent overdispersion in longitudinal count data. Computational Statistics & Data Analysis, 58(??):257–264, February 2013. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794731200312X.

Schaarschmidt:2013:SCI

[1046] Frank Schaarschmidt. Simultaneous confidence intervals for multiple comparisons among expected values of log-normal variables. Computational Statistics & Data Analysis, 58(??): 265–275, February 2013. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947312003143.

Hyun:2013:PFI

[1047] Seung Won Hyun, Min Yang, and Nancy Flournoy. A procedure for finding an improved upper bound on the number of optimal design points. Computational Statistics & Data Analysis, 58(??):276– 282, February 2013. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947312003155.

Zhang:2013:SSC

[1048] Song Zhang and Chul Ahn. Sample size calculation for comparing time-averaged responses in k-group repeated-measurement studies. Computational Statistics & Data Analysis, 58(??):283–291, February 2013. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0167947312003167.

Chakraborty:2013:SIM

[1049] A. Chakraborty, M. A. Beamonte, A. E. Gelfand, M. P. Alonso, P. Gargallo, and M. Salvador. Spatial interaction models with individual-level data for explaining labor flows and developing local labor markets. Computational Statistics & Data Analysis, 58 (??):292–307, February 2013. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947312003209.

Marhuenda:2013:SAE

[1050] Yolanda Marhuenda, Isabel Molina, and Domingo Morales. Small area estimation with spatio-temporal Fay– Herriot models. Computational Statistics & Data Analysis, 58(??):308– 325, February 2013. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947312003222.

Fraiman:2013:REH

[1051] Ricardo Fraiman and Marcela Svarc. Resistant estimates for high dimensional and functional data based on random projections. Computational Statistics & Data Analysis, 58(??): 326–338, February 2013. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947312003350.

Yi:2013:STT

[1052] Feng Yi and Hui Zou. SURE-tuned tapering estimation of large covariance matrices. Computational Statistics & Data Analysis, 58(??):339– 351, February 2013. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947312003362.

Silva:2013:CCE

[1053] Rodrigo B. Silva, Marcelo Bourguignon, Cícero R. B. Dias, and Gauss M. Cordeiro. The compound class of extended Weibull power series distributions. Computational Statistics & Data Analysis, 58(??):352– 367, February 2013. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947312003386.

Perez:2013:SSD

[1054] M. Luz Gámiz Pérez, María Dolores Martínez Miranda, and Jens Perch Nielsen. Smoothing survival densities in practice. Computational Statistics & Data Analysis, 58(??):368– 382, February 2013. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947312003404.

Eloyan:2013:SAS

[1055] Ani Eloyan and Sujit K. Ghosh. A semiparametric approach to source separation using independent component analysis. Computational Statistics & Data Analysis, 58(??):383– 396, February 2013. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947312003416.

Hilow:2013:CAR

[1056] Hisham Hilow. Comparison among run order algorithms for sequential factorial experiments. Computational Statistics & Data Analysis, 58(??): 397–406, February 2013. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947312003428.

Borgoni:2013:ORS

[1057] Riccardo Borgoni, Luigi Radaelli, Valeria Tritto, and Diego Zappa. Optimal reduction of a spatial monitoring grid: Proposals and applications in process control. Computational Statistics & Data Analysis, 58 (??):407–419, February 2013. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947312003106.

Anonymous:2013:EBd

[1058] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 59(??):iii–v, March 2013. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947312004070.

Anonymous:2013:EBe

[1059] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 59(??):vi–vii, March 2013. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947312004082.

Hertel:2013:EOD

[1060] Ida Hertel and Michael Kohler. Estimation of the optimal design of a nonlinear parametric regression problem via Monte Carlo experiments. Computational Statistics & Data Analysis, 59(??):1–12, March 2013. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794731200343X.

Karl:2013:EML

[1061] Andrew T. Karl, Yan Yang, and Sharon L. Lohr. Efficient maximum likelihood estimation of multiple membership linear mixed models, with an application to educational value-added assessments. Computational Statistics & Data Analysis, 59 (??):13–27, March 2013. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947312003544.

Hirose:2013:TPS

[1062] Kei Hirose, Shohei Tateishi, and Sadanori Konishi. Tuning parameter selection in sparse regression modeling. Computational Statistics & Data Analysis, 59(??):28– 40, March 2013. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947312003556.

Hong:2013:CDF

[1063] Yili Hong. On computing the distribution function for the Poisson binomial distribution. Computational Statistics & Data Analysis, 59(??): 41–51, March 2013. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947312003568.

Dissmann:2013:SER

[1064] J. Dißmann, E. C. Brechmann, C. Czado, and D. Kurowicka. Selecting and estimating regular vine copulae and application to financial returns. Computational Statistics & Data Analysis, 59(??):52– 69, March 2013. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947312003131.

Samanta:2013:BHU

[1065] Mayukh Samanta and A. H. Welsh. Bootstrapping for highly unbalanced clustered data. Computational Statistics & Data Analysis, 59(??):70– 81, March 2013. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947312003246.

Gromenko:2013:NIS

[1066] Oleksandr Gromenko and Piotr Kokoszka.

Nonparametric inference in small data sets of spatially indexed curves with application to ionospheric trend determination. Computational Statistics & Data Analysis, 59(??):82– 94, March 2013. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947312003453.

Zhang:2013:NSE

[1067] Jiajia Zhang, Yingwei Peng, and Haifen Li. A new semiparametric estimation method for accelerated hazards mixture cure model. Computational Statistics & Data Analysis, 59 (??):95–102, March 2013. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947312003465.

Zhang:2013:ELR

[1068] Jun Zhang, Yujie Gai, and Ping Wu.

Estimation in linear regression models with measurement errors subject to single-indexed distortion. Computational Statistics & Data Analysis, 59 (??):103–120, March 2013. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947312003519.

Huang:2013:TSC

[1069] Zhensheng Huang, Zhen Pang, and Tao Hu. Testing structural change in partially linear single-index models with error-prone linear covariates. Computational Statistics & Data Analysis, 59 (??):121–133, March 2013. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947312003520.

Wang:2013:APO

[1070] Junhui Wang and Yixin Fang. Analysis of presence-only data via semisupervised learning approaches. Computational Statistics & Data Analysis, 59(??):134–143, March 2013. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016794731200357X.

Goncalves:2013:SMD

[1071] F. B. Gonçalves, D. Gamerman, and T. M. Soares. Simultaneous multifactor DIF analysis and detection in item response theory. Computational Statistics & Data Analysis, 59 (??):144–160, March 2013. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947312003611.

Villegas:2013:IDG

[1072] Cristian Villegas, Gilberto A. Paula, Francisco José A. Cysneiros, and Manuel Galea. Influence diagnostics in generalized symmetric linear models. Computational Statistics & Data Analysis, 59(??):161– 170, March 2013. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947312003623.

Lee:2013:RIU

[1073] Donghwan Lee, Youngjo Lee, Myunghee Cho Paik, and Michael G. Kenward. Robust inference using hierarchical likelihood approach for heavy-tailed longitudinal outcomes with missing data: an alternative to inverse probability weighted generalized estimating equations. Computational Statistics & Data Analysis, 59(??):171–179, March 2013. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947312003817.

Warrens:2013:CAW

[1074] Matthijs J. Warrens. The Cicchetti– Allison weighting matrix is positive definite. Computational Statistics & Data Analysis, 59(??):180– 182, March 2013. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947312003489.

Jones:2013:ESB

[1075] M. C. Jones. Erratum to: “On simulating Balakrishnan skew-normal variates” by Mahdi Teimouri and Saralees Nadarajah [Comput. Statist. Data Anal. 57 (2013) 52–58]. Computational Statistics & Data Analysis, 59 (??):183–184, March 2013. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947312003507. See

[977].

Anonymous:2013:EBf

[1076] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 60(??):iii–v, April 2013. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016794731200429X.

Anonymous:2013:EBg

[1077] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 60(??):vi–vii, April 2013. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947312004306.

Liu:2013:ECB

[1078] Xiaohui Liu, Yijun Zuo, and Zhizhong Wang. Exactly computing bivariate projection depth contours and median. Computational Statistics & Data Analysis, 60(??):1–11, April 2013. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947312003842.

McDaid:2013:IBI

[1079] Aaron F. McDaid, Thomas Brendan Murphy, Nial Friel, and Neil J. Hurley. Improved Bayesian inference for the stochastic block model with application to large networks. Computational Statistics & Data Analysis, 60(??):12–31, April 2013. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947312003891.

Liu:2013:HTU

[1080] Shen Liu and Elizabeth Ann Maharaj. A hypothesis test using biasadjusted AR estimators for classifying time series in small samples. Computational Statistics & Data Analysis, 60(??):32–49, April 2013. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947312004161.

Hapfelmeier:2013:NVS

[1081] A. Hapfelmeier and K. Ulm. A new variable selection approach using random forests. Computational Statistics & Data Analysis, 60(??): 50–69, April 2013. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947312003490.

Yoo:2013:ASD

[1082] Jae Keun Yoo. Advances in seeded dimension reduction: Bootstrap criteria and extensions. Computational Statistics & Data Analysis, 60(??): 70–79, April 2013. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947312003532.

Liu:2013:GVS

[1083] Xianhui Liu, Zhanfeng Wang, and Yaohua Wu. Group variable selection and estimation in the Tobit censored response model. Computational Statistics & Data Analysis, 60(??):80–89, April 2013. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947312003878.

Ventura:2013:OBH

[1084] Laura Ventura, Nicola Sartori, and Walter Racugno. Objective Bayesian higher-order asymptotics in models with nuisance parameters. Computational Statistics & Data Analysis, 60(??):90–96, April 2013. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947312003908.

Delatola:2013:BSM

[1085] E.-I. Delatola and J. E. Griffin. A Bayesian semiparametric model for volatility with a leverage effect. Computational Statistics & Data Analysis, 60(??):97–110, April 2013. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016794731200391X.

Rauch:2013:EPA

[1086] G. Rauch and M. Kieser. An expected power approach for the assessment of composite endpoints and their components. Computational Statistics & Data Analysis, 60(??): 111–122, April 2013. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947312003921.

Zhao:2013:NCG

[1087] Xingqiu Zhao, Ran Duan, Qiang Zhao, and Jianguo Sun. A new class of generalized log rank tests for intervalcensored failure time data. Computational Statistics & Data Analysis, 60 (??):123–131, April 2013. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947312003933.

Chen:2013:AAI

[1088] Chia-Cheng Chen and Huiman X.

Barnhart. Assessing agreement with intraclass correlation coefficient and concordance correlation coefficient for data with repeated measures. Computational Statistics & Data Analysis, 60(??):132–145, April 2013. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947312003957.

Reis:2013:BDM

[1089] Edna A. Reis, Dani Gamerman, Marina S. Paez, and Thiago G. Martins. Bayesian dynamic models for space–time point processes. Computational Statistics & Data Analysis, 60(??):146–156, April 2013. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947312003994.

Frey:2013:MPR

[1090] Jesse Frey and Le Wang. Most powerful rank tests for perfect rankings. Computational Statistics & Data Analysis, 60(??):157–168, April 2013. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947312004148.

Heo:2013:SSR

[1091] Moonseong Heo, Xiaonan Xue, and Mimi Y. Kim. Sample size requirements to detect an intervention by time interaction in longitudinal cluster randomized clinical trials with random slopes. Computational Statistics & Data Analysis, 60 (??):169–178, April 2013. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947312004185.

Anonymous:2013:EBh

[1092] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 61(??):iii–v, May 2013. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947313000133.

Anonymous:2013:EBi

[1093] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 61(??):vi–vii, May 2013. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947313000145.

Lloyd:2013:NIA

[1094] Chris J. Lloyd. A numerical investigation of the accuracy of parametric bootstrap for discrete data. Computational Statistics & Data Analysis, 61(??):1–6, May 2013. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947312003441.

Bolin:2013:CBM

[1095] David Bolin and Finn Lindgren. A comparison between Markov approximations and other methods for large spatial data sets. Computational Statistics & Data Analysis, 61(??): 7–21, May 2013. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947312004021.

Lee:2013:ETD

[1096] Dae-Jin Lee, María Durbán, and Paul Eilers. Efficient two-dimensional smoothing with P -spline ANOVA mixed models and nested bases. Computational Statistics & Data Analysis, 61(??):22–37, May 2013. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794731200415X.

Mathew:2013:FBT

[1097] Thomas Mathew and Derek S. Young.

Fiducial-based tolerance intervals for some discrete distributions. Computational Statistics & Data Analysis, 61(??):38–49, May 2013. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947312004173.

Tucker:2013:GMF

[1098] J. Derek Tucker, Wei Wu, and Anuj Srivastava. Generative models for functional data using phase and amplitude separation. Computational Statistics & Data Analysis, 61(??): 50–66, May 2013. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947312004227.

Lim:2013:STZ

[1099] Hwa Kyung Lim, Juwon Song, and Byoung Cheol Jung. Score tests for zero-inflation and overdispersion in two-level count data. Computational Statistics & Data Analysis, 61 (??):67–82, May 2013. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947312003970.

Brooks:2013:PNP

[1100] J. P. Brooks, J. H. Dulá, and E. L. Boone. A pure L_{1}-norm principal component analysis. Computational Statistics & Data Analysis, 61 (??):83–98, May 2013. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947312003982.

Bedrick:2013:CTH

[1101] Edward J. Bedrick and Anwar Hossain. Conditional tests for homogeneity of zero-inflated Poisson and Poisson-hurdle distributions. Computational Statistics & Data Analysis, 61(??):99–106, May 2013. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947312004008.

Moreira:2013:BSK

[1102] C. Moreira and I. Van Keilegom. Bandwidth selection for kernel density estimation with doubly truncated data. Computational Statistics & Data Analysis, 61(??):107– 123, May 2013. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947312004197.

Fritz:2013:FAR

[1103] Heinrich Fritz, Luis A. García-Escudero, and Agustín Mayo-Iscar. A fast algorithm for robust constrained clustering. Computational Statistics & Data Analysis, 61(??): 124–136, May 2013. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947312004203.

Figueroa-Zuniga:2013:MBR

[1104] Jorge I. Figueroa-Zúñiga, Reinaldo B.

Arellano-Valle, and Silvia L. P. Ferrari. Mixed beta regression: a Bayesian perspective. Computational Statistics & Data Analysis, 61(??): 137–147, May 2013. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947312004239.

Saez-Castillo:2013:HPR

[1105] A. J. Sáez-Castillo and A. Conde-Sánchez. A hyper-Poisson regression model for overdispersed and underdispersed count data. Computational Statistics & Data Analysis, 61 (??):148–157, May 2013. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947312004434.

Marra:2013:ERS

[1106] Giampiero Marra and Rosalba Radice.

Estimation of a regression spline sample selection model. Computational Statistics & Data Analysis, 61(??): 158–173, May 2013. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947312004446.

Li:2013:AVS

[1107] Gaorong Li, Heng Lian, Sanying Feng, and Lixing Zhu. Automatic variable selection for longitudinal generalized linear models. Computational Statistics & Data Analysis, 61(??): 174–186, May 2013. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947312004495.

Lins:2013:PSS

[1108] Isis Didier Lins, Moacyr Araujo, Márcio das Chagas Moura, Marcus André Silva, and Enrique López Droguett. Prediction of sea surface temperature in the tropical Atlantic by support vector machines. Computational Statistics & Data Analysis, 61(??):187–198, May 2013. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947312004240.

Belzunce:2013:SCM

[1109] Félix Belzunce, Carolina Martínez-Riquelme, and José M. Ruiz. On sufficient conditions for mean residual life and related orders. Computational Statistics & Data Analysis, 61(??):199–210, May 2013. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947312004392.

Anonymous:2013:EBj

[1110] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 62(??):iii–v, June 2013. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313000637.

Anonymous:2013:EBk

[1111] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 62(??):vi–vii, June 2013. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313000649.

Melnykov:2013:CTE

[1112] Volodymyr Melnykov and Gang Shen.

Clustering through empirical likelihood ratio. Computational Statistics & Data Analysis, 62(??):1– 10, June 2013. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947312004458.

Fernandez:2013:SPC

[1113] Arturo J. Fernández. Smallest Pareto confidence regions and applications. Computational Statistics & Data Analysis, 62(??):11–25, June 2013. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947312004501.

Lee:2013:CDM

[1114] Seokho Lee and Jianhua Z. Huang. A coordinate descent MM algorithm for fast computation of sparse logistic PCA. Computational Statistics & Data Analysis, 62(??):26– 38, June 2013. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313000029.

Kang:2013:EVU

[1115] Le Kang and Lili Tian. Estimation of the volume under the ROC surface with three ordinal diagnostic categories. Computational Statistics & Data Analysis, 62(??):39–51, June 2013. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313000054.

Ferreira:2013:KFE

[1116] Guillermo Ferreira, Alejandro Rodríguez, and Bernardo Lagos. Kalman filter estimation for a regression model with locally stationary errors. Computational Statistics & Data Analysis, 62(??):52–69, June 2013. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313000066.

Huang:2013:APE

[1117] Zhensheng Huang, Zhen Pang, and Riquan Zhang. Adaptive profileempirical-likelihood inferences for generalized single-index models. Computational Statistics & Data Analysis, 62(??):70–82, June 2013. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947312004409.

Adekpedjou:2013:CIP

[1118] Akim Adekpedjou, Russell Stocker, and Withanage A. De Mel. A class of inference procedures for validating the generalized Koziol–Green model with recurrent events. Computational Statistics & Data Analysis, 62 (??):83–92, June 2013. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947312004483.

An:2013:MRS

[1119] Baiguo An, Jianhua Guo, and Hansheng Wang. Multivariate regression shrinkage and selection by canonical correlation analysis. Computational Statistics & Data Analysis, 62 (??):93–107, June 2013. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947312004513.

Sen:2013:BPI

[1120] Ananda Sen, Nandini Kannan, and Debasis Kundu. Bayesian planning and inference of a progressively censored sample from linear hazard rate distribution. Computational Statistics & Data Analysis, 62 (??):108–121, June 2013. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313000042.

Vidaurre:2013:SRL

[1121] Diego Vidaurre, Concha Bielza, and Pedro Larrañaga. Sparse regularized local regression. Computational Statistics & Data Analysis, 62(??): 122–135, June 2013. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947313000236.

Haggstrom:2013:BSB

[1122] Jenny Häggström. Bandwidth selection for backfitting estimation of semiparametric additive models: a simulation study. Computational Statistics & Data Analysis, 62(??): 136–148, June 2013. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016794731300025X.

Lemonte:2013:NET

[1123] Artur J. Lemonte. A new exponentialtype distribution with constant, decreasing, increasing, upside-down bathtub and bathtub-shaped failure rate function. Computational Statistics & Data Analysis, 62(??):149– 170, June 2013. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947313000261.

Chen:2013:EAF

[1124] Pengcheng Chen, Jiajia Zhang, and Riquan Zhang. Estimation of the accelerated failure time frailty model under generalized gamma frailty. Computational Statistics & Data Analysis, 62(??):171–180, June 2013. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313000315.

Kim:2013:PSA

[1125] Young-Ju Kim. A partial spline approach for semiparametric estimation of varying-coefficient partially linear models. Computational Statistics & Data Analysis, 62(??):181– 187, June 2013. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313000078.

Anonymous:2013:EBl

[1126] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 63(??):iii–v, July 2013. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313000960.

Anonymous:2013:Cb

[1127] Anonymous. Contents. Computational Statistics & Data Analysis, 63 (??):vi, July 2013. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313000972.

Marchant:2013:GBS

[1128] Carolina Marchant, Karine Bertin, Víctor Leiva, and Helton Saulo. Generalized Birnbaum–Saunders kernel density estimators and an analysis of financial data. Computational Statistics & Data Analysis, 63 (??):1–15, July 2013. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947313000285.

Pantazis:2013:PPS

[1129] Nikos Pantazis, Michael G. Kenward, and Giota Touloumi. Performance of parametric survival models under non-random interval censoring: a simulation study. Computational Statistics & Data Analysis, 63(??):16–30, July 2013. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947313000297.

Ueki:2013:MCC

[1130] Masao Ueki and Yoshinori Kawasaki. Multiple choice from competing regression models under multicollinearity based on standardized update. Computational Statistics & Data Analysis, 63(??):31–41, July 2013. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947313000340.

Yao:2013:RVS

[1131] Weixin Yao and Qin Wang. Robust variable selection through MAVE. Computational Statistics & Data Analysis, 63(??):42–49, July 2013. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947313000364.

Hong:2013:QAP

[1132] Hyokyoung Grace Hong. A quantile approach to the power transformed location–scale model. Computational Statistics & Data Analysis, 63(??):50–62, July 2013. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313000376.

Bhattacharya:2013:RPN

[1133] Rabi Bhattacharya and Lizhen Lin.

Recent progress in the nonparametric estimation of monotone curves — with applications to bioassay and environmental risk assessment. Computational Statistics & Data Analysis, 63(??):63–80, July 2013. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313000388.

Derrode:2013:UDC

[1134] Stéphane Derrode and Wojciech Pieczynski. Unsupervised data classification using pairwise Markov chains with automatic copulas selection. Computational Statistics & Data Analysis, 63(??):81–98, July 2013. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794731300042X.

Lian:2013:SVS

[1135] Heng Lian, Jianbo Li, and Yuao Hu. Shrinkage variable selection and estimation in proportional hazards models with additive structure and high dimensionality. Computational Statistics & Data Analysis, 63(??):99–112, July 2013. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313000479.

Gilardoni:2013:NEB

[1136] Gustavo L. Gilardoni, Maristela D. de Oliveira, and Enrico A. Colosimo. Nonparametric estimation and bootstrap confidence intervals for the optimal maintenance time of a repairable system. Computational Statistics & Data Analysis, 63(??): 113–124, July 2013. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313000509.

Liseo:2013:BIM

[1137] Brunero Liseo and Antonio Parisi. Bayesian inference for the multivariate skew-normal model: a population Monte Carlo approach. Computational Statistics & Data Analysis, 63(??):125–138, July 2013. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313000510.

Hilafu:2013:SDR

[1138] Haileab Hilafu and Xiangrong Yin. Sufficient dimension reduction in multivariate regressions with categorical predictors. Computational Statistics & Data Analysis, 63(??):139– 147, July 2013. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313000583.

Hsu:2013:EEM

[1139] Chih-Yuan Hsu and Tiee-Jian Wu. Efficient estimation of the mode of continuous multivariate data. Computational Statistics & Data Analysis, 63(??):148–159, July 2013. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313000339.

Anonymous:2013:EBm

[1140] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 64(??):iii–v, August 2013. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313001461.

Anonymous:2013:Cc

[1141] Anonymous. Contents. Computational Statistics & Data Analysis, 64 (??):vi–vii, August 2013. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313001473.

Kim:2013:MVC

[1142] Daeyoung Kim, Jong-Min Kim, Shu-Min Liao, and Yoon-Sung Jung. Mixture of D-vine copulas for modeling dependence. Computational Statistics & Data Analysis, 64(??): 1–19, August 2013. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313000741.

Ghitany:2013:PLD

[1143] M. E. Ghitany, D. K. Al-Mutairi, N. Balakrishnan, and L. J. Al-Enezi. Power Lindley distribution and associated inference. Computational Statistics & Data Analysis, 64(??): 20–33, August 2013. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313000820.

Lemonte:2013:NEB

[1144] Artur J. Lemonte. A new extended Birnbaum–Saunders regression model for lifetime modeling. Computational Statistics & Data Analysis, 64(??): 34–50, August 2013. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313000819.

Yousef:2013:ACT

[1145] Waleed A. Yousef. Assessing classifiers in terms of the partial area under the ROC curve. Computational Statistics & Data Analysis, 64 (??):51–70, August 2013. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313000881.

Mandal:2013:MDE

[1146] Abhijit Mandal and Ayanendranath Basu. Minimum disparity estimation: Improved efficiency through inlier modification. Computational Statistics & Data Analysis, 64(??): 71–86, August 2013. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313000868.

Klingenberg:2013:SCI

[1147] Bernhard Klingenberg and Ville Satopää. Simultaneous confidence intervals for comparing margins of multivariate binary data. Computational Statistics & Data Analysis, 64(??): 87–98, August 2013. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313000728.

Maboudou-Tchao:2013:MCM

[1148] Edgard M. Maboudou-Tchao and Vincent Agboto. Monitoring the covariance matrix with fewer observations than variables. Computational Statistics & Data Analysis, 64(??): 99–112, August 2013. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313000844.

Harvill:2013:BBM

[1149] Jane L. Harvill, Nalini Ravishanker, and Bonnie K. Ray. Bispectralbased methods for clustering time series. Computational Statistics & Data Analysis, 64(??):113–131, August 2013. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0167947313000893.

Fellinghauer:2013:SGM

[1150] Bernd Fellinghauer, Peter Bühlmann, Martin Ryffel, Michael von Rhein, and Jan D. Reinhardt. Stable graphical model estimation with random forests for discrete, continuous, and mixed variables. Computational Statistics & Data Analysis, 64(??):132– 152, August 2013. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313000789.

Alcantara:2013:LRM

[1151] Izabel Cristina Alcantara and Francisco José A. Cysneiros. Linear regression models with slashelliptical errors. Computational Statistics & Data Analysis, 64(??):153– 164, August 2013. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313000856.

Friedrich:2013:GIE

[1152] Thomas Friedrich and Guido Knapp. Generalised interval estimation in the random effects meta regression model. Computational Statistics & Data Analysis, 64(??):165– 179, August 2013. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313001138.

Jiang:2013:TSC

[1153] Rong Jiang, Zhan-Gong Zhou, Wei-Min Qian, and Yong Chen. Two step composite quantile regression for single-index models. Computational Statistics & Data Analysis, 64(??): 180–191, August 2013. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313001163.

Kim:2013:BMP

[1154] Jung Seek Kim and Brian T. Ratchford. A Bayesian multivariate probit for ordinal data with semiparametric random-effects. Computational Statistics & Data Analysis, 64(??): 192–208, August 2013. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313001023.

Alhamzawi:2013:CPV

[1155] Rahim Alhamzawi and Keming Yu.

Conjugate priors and variable selection for Bayesian quantile regression. Computational Statistics & Data Analysis, 64(??):209–219, August 2013. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947312000345.

Vandewalle:2013:PDC

[1156] Vincent Vandewalle, Christophe Biernacki, Gilles Celeux, and Gérard Govaert. A predictive deviance criterion for selecting a generative model in semi-supervised classification. Computational Statistics & Data Analysis, 64 (??):220–236, August 2013. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313000546.

Lachos:2013:BIN

[1157] Victor H. Lachos, Luis M. Castro, and Dipak K. Dey. Bayesian inference in nonlinear mixed-effects models using normal independent distributions. Computational Statistics & Data Analysis, 64(??):237– 252, August 2013. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313000558.

Petrie:2013:EST

[1158] Adam Petrie and Thomas R. Willemain. An empirical study of tests for uniformity in multidimensional data. Computational Statistics & Data Analysis, 64(??):253– 268, August 2013. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313000571.

Ding:2013:SIS

[1159] Chang Ding, Siu keung Tse, and Ronghai Yang. Statistical inference on survival data in group-parallel clinical trials with treatment switching. Computational Statistics & Data Analysis, 64 (??):269–280, August 2013. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313000716.

Korzen:2013:LRW

[1160] M. Korzeń, S. Jaroszewicz, and P. Klesk. Logistic regression with weight grouping priors. Computational Statistics & Data Analysis, 64 (??):281–298, August 2013. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313001151.

Guan:2013:OBA

[1161] Qiang Guan, Yincai Tang, and Ancha Xu. Objective Bayesian analysis for bivariate Marshall–Olkin exponential distribution. Computational Statistics & Data Analysis, 64(??): 299–313, August 2013. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313001230.

Anonymous:2013:EBn

[1162] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 65 (??):iii–v, September 2013. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313001734.

Anonymous:2013:Cd

[1163] Anonymous. Contents. Computational Statistics & Data Analysis, 65 (??):vi, September 2013. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313001746.

Croux:2013:SIR

[1164] Christophe Croux, Elvezio Ronchetti, Matías Salibián-Barrera, and Stefan Van Aelst. Special issue on robust analysis of complex data. Computational Statistics & Data Analysis, 65 (??):1–3, September 2013. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313001370.

Wang:2013:DED

[1165] B. Wang and W. Wertelecki. Density estimation for data with rounding errors. Computational Statistics & Data Analysis, 65(??):4–12, September 2013. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947312000989.

Hubert:2013:DID

[1166] Mia Hubert, Goedele Dierckx, and Dina Vanpaemel. Detecting influential data points for the Hill estimator in Pareto-type distributions. Computational Statistics & Data Analysis, 65 (??):13–28, September 2013. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794731200285X.

Cerioli:2013:RDO

[1167] Andrea Cerioli, Alessio Farcomeni, and Marco Riani. Robust distances for outlier-free goodness-of-fit testing. Computational Statistics & Data Analysis, 65(??):29– 45, September 2013. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794731200134X.

Maronna:2013:RFL

[1168] Ricardo A. Maronna and Victor J. Yohai. Robust functional linear regression based on splines. Computational Statistics & Data Analysis, 65 (??):46–55, September 2013. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947311004117.

Guo:2013:VSH

[1169] Jie Guo, Manlai Tang, Maozai Tian, and Kai Zhu. Variable selection in high-dimensional partially linear additive models for composite quantile regression. Computational Statistics & Data Analysis, 65(??):56– 67, September 2013. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313001199.

Muler:2013:REV

[1170] Nora Muler and Víctor J. Yohai. Robust estimation for vector autoregressive models. Computational Statistics & Data Analysis, 65(??):68– 79, September 2013. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794731200093X.

Bianco:2013:RTG

[1171] Ana M. Bianco, Graciela Boente, and Isabel M. Rodrigues. Robust tests in generalized linear models with missing responses. Computational Statistics & Data Analysis, 65(??): 80–97, September 2013. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947312002071.

Lind:2013:RMI

[1172] John C. Lind, Douglas P. Wiens, and Victor J. Yohai. Robust minimum information loss estimation. Computational Statistics & Data Analysis, 65 (??):98–112, September 2013. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947312002526.

Anonymous:2013:EBo

[1173] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 66(??):iii–v, October 2013. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947313001928.

Anonymous:2013:Ce

[1174] Anonymous. Contents. Computational Statistics & Data Analysis, 66 (??):vi–vii, October 2013. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016794731300193X.

Evans:2013:TAF

[1175] R. J. Evans and A. Forcina. Two algorithms for fitting constrained marginal models. Computational Statistics & Data Analysis, 66(??): 1–7, October 2013. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947313000455.

Gaure:2013:OMH

[1176] Simen Gaure. OLS with multiple high dimensional category variables. Computational Statistics & Data Analysis, 66(??):8–18, October 2013. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313001266.

Bowman:2013:IV

[1177] Adrian W. Bowman and Rosa M. Crujeiras. Inference for variograms. Computational Statistics & Data Analysis, 66(??):19–31, October 2013. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313000832.

Hsieh:2013:LRO

[1178] Shu-Hui Hsieh, Chin-Shang Li, and Shen-Ming Lee. Logistic regression with outcome and covariates missing separately or simultaneously. Computational Statistics & Data Analysis, 66 (??):32–54, October 2013. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313001059.

Bachoc:2013:CVM

[1179] François Bachoc. Cross validation and maximum likelihood estimations of hyper-parameters of Gaussian processes with model misspecification. Computational Statistics & Data Analysis, 66(??):55– 69, October 2013. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313001187.

Amini:2013:NMA

[1180] Morteza Amini and N. Balakrishnan. Nonparametric meta-analysis of independent samples of records. Computational Statistics & Data Analysis, 66 (??):70–81, October 2013. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947313001217.

Zuo:2013:MMU

[1181] Yijun Zuo. Multidimensional medians and uniqueness. Computational Statistics & Data Analysis, 66(??): 82–88, October 2013. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947313001229.

Lee:2013:TSS

[1182] Seokho Lee, Hyejin Shin, and Nedret Billor. M-type smoothing spline estimators for principal functions. Computational Statistics & Data Analysis, 66 (??):89–100, October 2013. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947313001242.

Zhao:2013:ELI

[1183] Hui Zhao, Pu-Ying Zhao, and Nian-Sheng Tang. Empirical likelihood inference for mean functionals with nonignorably missing response data. Computational Statistics & Data Analysis, 66 (??):101–116, October 2013. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947313001254.

Morio:2013:OIP

[1184] Jérôme Morio, Damien Jacquemart, Mathieu Balesdent, and Julien Marzat. Optimisation of interacting particle systems for rare event estimation. Computational Statistics & Data Analysis, 66(??):117– 128, October 2013. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313001278.

Lin:2013:EUE

[1185] Huazhen Lin, Yi Li, and Ming T.

Tan. Estimating a unitary effect summary based on combined survival and quantitative outcomes. Computational Statistics & Data Analysis, 66 (??):129–139, October 2013. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313001308.

Wang:2013:OFS

[1186] Cheng Wang, Longbing Cao, and Baiqi Miao. Optimal feature selection for sparse linear discriminant analysis and its applications in gene expression data. Computational Statistics & Data Analysis, 66(??):140– 149, October 2013. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313001333.

Wang:2013:SIP

[1187] Xiaoguang Wang, Dawei Lu, and Lixin Song. Statistical inference for partially linear stochastic models with heteroscedastic errors. Computational Statistics & Data Analysis, 66(??): 150–160, October 2013. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313001345.

Ntzoufras:2013:CCC

[1188] Ioannis Ntzoufras and Claudia Tarantola. Conjugate and conditional conjugate Bayesian analysis of discrete graphical models of marginal independence. Computational Statistics & Data Analysis, 66(??):161– 177, October 2013. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313001357.

Yan:2013:CF

[1189] Donghui Yan, Aiyou Chen, and Michael I. Jordan. Cluster forests. Computational Statistics & Data Analysis, 66(??):178–192, October 2013. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0167947313001400.

Wittek:2013:TWI

[1190] Peter Wittek. Two-way incremental seriation in the temporal domain with three-dimensional visualization: Making sense of evolving high-dimensional datasets. Computational Statistics & Data Analysis, 66 (??):193–201, October 2013. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794731300128X.

Tamura:2013:NPM

[1191] Karin Ayumi Tamura and Viviana Giampaoli. New prediction method for the mixed logistic model applied in a marketing problem. Computational Statistics & Data Analysis, 66 (??):202–216, October 2013. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313001369.

Anonymous:2013:EBp

[1192] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 67(??):iii–v, November 2013. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313002387.

Anonymous:2013:Cf

[1193] Anonymous. Contents. Computational Statistics & Data Analysis, 67 (??):vi–vii, November 2013. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313002399.

Alvarez-Esteban:2013:SCP

[1194] P. C. Álvarez-Esteban, E. del Barrio, J. A. Cuesta-Albertos, and C. Matrán. Searching for a common pooling pattern among several samples. Computational Statistics & Data Analysis, 67 (??):1–14, November 2013. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313001564.

Thomas:2013:MAI

[1195] Laine Thomas, Leonard A. Stefanski, and Marie Davidian. Moment adjusted imputation for multivariate measurement error data with applications to logistic regression. Computational Statistics & Data Analysis, 67 (??):15–24, November 2013. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313001588.

Jin:2013:SRI

[1196] Hao Jin, Jinsuo Zhang, Si Zhang, and Cong Yu. The spurious regression of AR(p) infinite-variance sequence in the presence of structural breaks. Computational Statistics & Data Analysis, 67 (??):25–40, November 2013. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313001527.

Balakrishnan:2013:LLL

[1197] N. Balakrishnan and Suvra Pal. Lognormal lifetimes and likelihood-based inference for flexible cure rate models based on COM-Poisson family. Computational Statistics & Data Analysis, 67 (??):41–67, November 2013. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313001606.

Martins:2013:BCI

[1198] Thiago G. Martins, Daniel Simpson, Finn Lindgren, and Håvard Rue. Bayesian computing with INLA: New features. Computational Statistics & Data Analysis, 67(??):68– 83, November 2013. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313001552.

Vinciotti:2013:RMI

[1199] Veronica Vinciotti and Hussein Hashem.

Robust methods for inferring sparse network structures. Computational Statistics & Data Analysis, 67(??):84– 94, November 2013. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313001655.

Franck:2013:MDH

[1200] Christopher T. Franck, Dahlia M.

Nielsen, and Jason A. Osborne. A method for detecting hidden additivity in two-factor unreplicated experiments. Computational Statistics & Data Analysis, 67(??):95– 104, November 2013. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313001618.

Hino:2013:EBS

[1201] Hideitsu Hino, Keigo Wakayama, and Noboru Murata. Entropy-based sliced inverse regression. Computational Statistics & Data Analysis, 67(??): 105–114, November 2013. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313002028.

Liu:2013:VPM

[1202] Yin Liu and Guo-Liang Tian. A variant of the parallel model for sample surveys with sensitive characteristics. Computational Statistics & Data Analysis, 67 (??):115–135, November 2013. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947313001643.

Guan:2013:SHD

[1203] Wei Guan and Alexander Gray. Sparse high-dimensional fractional-norm support vector machine via DC programming. Computational Statistics & Data Analysis, 67(??):136– 148, November 2013. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947313000352.

Yang:2013:RFW

[1204] Jingjing Yang and David W. Scott. Robust fitting of a Weibull model with optional censoring. Computational Statistics & Data Analysis, 67 (??):149–161, November 2013. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947313001825.

Lin:2013:NFS

[1205] Lu Lin, Jing Sun, and Lixing Zhu. Nonparametric feature screening. Computational Statistics & Data Analysis, 67 (??):162–174, November 2013. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947313002016.

Vsevolozhskaya:2013:CFC

[1206] O. A. Vsevolozhskaya, M. C. Greenwood, G. J. Bellante, S. L. Powell, R. L. Lawrence, and K. S. Repasky. Combining functions and the closure principle for performing follow-up tests in functional analysis of variance. Computational Statistics & Data Analysis, 67 (??):175–184, November 2013. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313001667.

Shang:2013:BBE

[1207] Han Lin Shang. Bayesian bandwidth estimation for a nonparametric functional regression model with unknown error density. Computational Statistics & Data Analysis, 67(??): 185–198, November 2013. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313001679.

Shang:2013:PLS

[1208] Shulian Shang, Mengling Liu, Anne Zeleniuch-Jacquotte, Tess V. Clendenen, Vittorio Krogh, Goran Hallmans, and Wenbin Lu. Partially linear single index Cox regression model in nested case-control studies. Computational Statistics & Data Analysis, 67(??): 199–212, November 2013. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313001849.

Pedeli:2013:SPM

[1209] Xanthi Pedeli and Dimitris Karlis.

Some properties of multivariate INAR(1) processes. Computational Statistics & Data Analysis, 67(??):213– 225, November 2013. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313002065.

Wichitaksorn:2013:CMA

[1210] Nuttanan Wichitaksorn and Hiroki Tsurumi. Comparison of MCMC algorithms for the estimation of Tobit model with non-normal error: the case of asymmetric Laplace distribution. Computational Statistics & Data Analysis, 67(??):226– 235, November 2013. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313002156.

Givens:2013:IFR

[1211] G. H. Givens, J. R. Beveridge, P. J. Phillips, B. Draper, Y. M. Lui, and D. Bolme. Introduction to face recognition and evaluation of algorithm performance. Computational Statistics & Data Analysis, 67(??):236– 247, November 2013. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313002120.

Wang:2013:SPA

[1212] Jixian Wang and George Quartey. A semi-parametric approach to analysis of event duration and prevalence. Computational Statistics & Data Analysis, 67(??):248–257, November 2013. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313002107.

Fung:2013:CVM

[1213] Thomas Fung, Joanna J. J. Wang, and Eugene Seneta. Contaminated variance–mean mixing model. Computational Statistics & Data Analysis, 67 (??):258–267, November 2013. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313002119.

Rastogi:2013:EUH

[1214] Manoj Kumar Rastogi and Yogesh Mani Tripathi. Estimation using hybrid censored data from a two-parameter distribution with bathtub shape. Computational Statistics & Data Analysis, 67(??):268– 281, November 2013. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313002090.

Durot:2013:LSE

[1215] Cécile Durot, Sylvie Huet, François Koladjo, and Stéphane Robin. Leastsquares estimation of a convex discrete distribution. Computational Statistics & Data Analysis, 67(??): 282–298, November 2013. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794731300162X.

Anonymous:2013:EBq

[1216] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 68(??):iii–v, December 2013. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947313003009.

Anonymous:2013:Cg

[1217] Anonymous. Contents. Computational Statistics & Data Analysis, 68 (??):vi–vii, December 2013. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947313003010.

Exterkate:2013:MSK

[1218] Peter Exterkate. Model selection in kernel ridge regression. Computational Statistics & Data Analysis, 68 (??):1–16, December 2013. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947313002181.

Koldanov:2013:SPM

[1219] Alexander P. Koldanov, Petr A. Koldanov, Valeriy A. Kalyagin, and Panos M. Pardalos. Statistical procedures for the market graph construction. Computational Statistics & Data Analysis, 68(??):17– 29, December 2013. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016794731300217X.

Simpkin:2013:APS

[1220] Andrew Simpkin and John Newell. An additive penalty P -spline approach to derivative estimation. Computational Statistics & Data Analysis, 68 (??):30–43, December 2013. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313002193.

Paula:2013:DDG

[1221] Gilberto A. Paula. On diagnostics in double generalized linear models. Computational Statistics & Data Analysis, 68(??):44–51, December 2013. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794731300220X.

Beh:2013:RAA

[1222] Eric J. Beh, Duy Tran, and Irene L.

Hudson. A reformulation of the aggregate association index using the odds ratio. Computational Statistics & Data Analysis, 68(??):52– 65, December 2013. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313002211.

Li:2013:PLA

[1223] Haocheng Li and Grace Y. Yi. A pairwise likelihood approach for longitudinal data with missing observations in both response and covariates. Computational Statistics & Data Analysis, 68 (??):66–81, December 2013. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313002132.

Xue:2013:EEL

[1224] Liugen Xue. Estimation and empirical likelihood for single-index models with missing data in the covariates. Computational Statistics & Data Analysis, 68 (??):82–97, December 2013. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313002296.

Hui:2013:NML

[1225] Francis K. C. Hui and Gery Geenens. A nonparametric measure of local association for two-way contingency tables. Computational Statistics & Data Analysis, 68(??):98– 110, December 2013. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313002259.

Chen:2013:BDT

[1226] Ji Chen, Pengfei Liu, and Xinyuan Song. Bayesian diagnostics of transformation structural equation models. Computational Statistics & Data Analysis, 68(??):111–128, December 2013. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313002247.

Luong:2013:FEP

[1227] The Minh Luong, Yves Rozenholc, and Gregory Nuel. Fast estimation of posterior probabilities in changepoint analysis through a constrained hidden Markov model. Computational Statistics & Data Analysis, 68(??): 129–140, December 2013. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313002326.

Salazar:2013:AST

[1228] Esther Salazar, David B. Dunson, and Lawrence Carin. Analysis of space– time relational data with application to legislative voting. Computational Statistics & Data Analysis, 68(??): 141–154, December 2013. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313002302.

Perez:2013:BSM

[1229] María Luz Gámiz Pérez, Lena Janys, María Dolores Martínez Miranda, and Jens Perch Nielsen. Bandwidth selection in marker dependent kernel hazard estimation. Computational Statistics & Data Analysis, 68(??): 155–169, December 2013. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313002223.

Perry:2013:SDC

[1230] Marcus B. Perry, Gregory V. Michaelson, and M. Alan Ballard. On the statistical detection of clusters in undirected networks. Computational Statistics & Data Analysis, 68(??): 170–189, December 2013. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313002314.

Lee:2013:KCR

[1231] Myung Hee Lee and Yufeng Liu. Kernel continuum regression. Computational Statistics & Data Analysis, 68 (??):190–201, December 2013. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313002284.

Nashimoto:2013:MCB

[1232] Kane Nashimoto, Kristin M. Haldeman, and Christopher M. Tait. Multiple comparisons of k binomial proportions. Computational Statistics & Data Analysis, 68(??):202– 212, December 2013. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313002478.

Weiss:2013:SDN

[1233] Christian H. Weiß. Serial dependence of NDARMA processes. Computational Statistics & Data Analysis, 68 (??):213–238, December 2013. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313002545.

Molas:2013:JHG

[1234] Marek Molas, Maengseok Noh, Youngjo Lee, and Emmanuel Lesaffre. Joint hierarchical generalized linear models with multivariate Gaussian random effects. Computational Statistics & Data Analysis, 68(??):239– 250, December 2013. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313002569.

Poon:2013:BAG

[1235] Wai-Yin Poon and Hai-Bin Wang.

Bayesian analysis of generalized partially linear single-index models. Computational Statistics & Data Analysis, 68(??):251–261, December 2013. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0167947313002636.

Lopes:2013:SFP

[1236] Sílvia R. C. Lopes and Taiane S.

Prass. Seasonal FIEGARCH processes. Computational Statistics & Data Analysis, 68(??):262–295, December 2013. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313002466.

Chang:2013:DML

[1237] Hsiu-Ching Chang and Hwan Chung.

Dealing with multiple local modalities in latent class profile analysis. Computational Statistics & Data Analysis, 68 (??):296–310, December 2013. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313002612.

Conti:2013:UAS

[1238] Pier Luigi Conti, Daniela Marella, and Mauro Scanu. Uncertainty analysis for statistical matching of ordered categorical variables. Computational Statistics & Data Analysis, 68(??): 311–325, December 2013. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947313002491.

Kang:2013:ECI

[1239] Le Kang, Chengjie Xiong, and Lili Tian. Estimating confidence intervals for the difference in diagnostic accuracy with three ordinal diagnostic categories without a gold standard. Computational Statistics & Data Analysis, 68 (??):326–338, December 2013. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947313002521.

Lang:2013:ITI

[1240] Joseph B. Lang and Maria Iannario. Improved tests of independence in singly-ordered two-way contingency tables. Computational Statistics & Data Analysis, 68(??):339– 351, December 2013. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947313002260.

Wu:2013:OSM

[1241] Shu-Fei Wu. One stage multiple comparisons with the average for exponential location parameters under heteroscedasticity. Computational Statistics & Data Analysis, 68(??):352– 360, December 2013. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947313002508.

Fox:2013:MZI

[1242] Jean-Paul Fox. Multivariate zeroinflated modeling with latent predictors: Modeling feedback behavior. Computational Statistics & Data Analysis, 68(??):361–374, December 2013. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794731300248X.

Pham:2013:MFV

[1243] Tung H. Pham, John T. Ormerod, and M. P. Wand. Mean field variational Bayesian inference for nonparametric regression with measurement error. Computational Statistics & Data Analysis, 68(??):375– 387, December 2013. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313002594.

Soyer:2013:BDP

[1244] Refik Soyer and Minje Sung. Bayesian dynamic probit models for the analysis of longitudinal data. Computational Statistics & Data Analysis, 68 (??):388–398, December 2013. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313001576.

Anonymous:2014:EBa

[1245] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 69(??):iii–v, January 2014. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313003307.

Anonymous:2014:Ca

[1246] Anonymous. Contents. Computational Statistics & Data Analysis, 69 (??):vi–vii, January 2014. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947313003319.

Rota:2014:FOC

[1247] Matteo Rota and Laura Antolini. Finding the optimal cut-point for Gaussian and gamma distributed biomarkers. Computational Statistics & Data Analysis, 69(??):1– 14, January 2014. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947313002600.

deSaporta:2014:SSA

[1248] Benoîte de Saporta, Anne Gégout-Petit, and Laurence Marsalle. Statistical study of asymmetry in cell lineage data. Computational Statistics & Data Analysis, 69(??):15– 39, January 2014. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947313002703.

Westerlund:2014:CTU

[1249] Joakim Westerlund. On the choice of test for a unit root when the errors are conditionally heteroskedastic. Computational Statistics & Data Analysis, 69 (??):40–53, January 2014. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947313002673.

Vakili:2014:FMO

[1250] Kaveh Vakili and Eric Schmitt.

Finding multivariate outliers with FastPCS. Computational Statistics & Data Analysis, 69(??):54– 66, January 2014. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313002661.

Alzaatreh:2014:GND

[1251] Ayman Alzaatreh, Felix Famoye, and Carl Lee. The gamma-normal distribution: Properties and applications. Computational Statistics & Data Analysis, 69(??):67– 80, January 2014. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313002806.

Bernhardt:2014:FMS

[1252] Paul W. Bernhardt, Huixia Judy Wang, and Daowen Zhang. Flexible modeling of survival data with covariates subject to detection limits via multiple imputation. Computational Statistics & Data Analysis, 69 (??):81–91, January 2014. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313002727.

Lee:2014:GFT

[1253] Sangyeol Lee. Goodness of fit test for discrete random variables. Computational Statistics & Data Analysis, 69 (??):92–100, January 2014. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947313002715.

Xiong:2014:OPC

[1254] Chengjie Xiong, Jingqin Luo, Feng Gao, and John C. Morris. Optimizing parameters in clinical trials with a randomized start or withdrawal design. Computational Statistics & Data Analysis, 69(??):101– 113, January 2014. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947313002582.

Zhang:2014:GAD

[1255] Song Zhang, Jing Cao, and Chul Ahn. A GEE approach to determine sample size for pre- and post-intervention experiments with dropout. Computational Statistics & Data Analysis, 69 (??):114–121, January 2014. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016794731300282X.

Ozturk:2014:IPP

[1256] Omer Ozturk and Mohammad Jafari Jozani. Inclusion probabilities in partially rank ordered set sampling. Computational Statistics & Data Analysis, 69(??):122– 132, January 2014. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016794731300279X.

Sadooghi-Alvandi:2014:SCI

[1257] S. M. Sadooghi-Alvandi and A. Malekzadeh.

Simultaneous confidence intervals for ratios of means of several lognormal distributions: a parametric bootstrap approach. Computational Statistics & Data Analysis, 69(??):133– 140, January 2014. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313002855.

Emiliano:2014:ICH

[1258] Paulo C. Emiliano, Mário J. F. Vivanco, and Fortunato S. de Menezes. Information criteria: How do they behave in different models? Computational Statistics & Data Analysis, 69 (??):141–153, January 2014. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313002776.

Amiri:2014:REN

[1259] Aboubacar Amiri, Christophe Crambes, and Baba Thiam. Recursive estimation of nonparametric regression with functional covariate. Computational Statistics & Data Analysis, 69 (??):154–172, January 2014. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313002752.

Fu:2014:PET

[1260] Liyong Fu, Mingliang Wang, Yuancai Lei, and Shouzheng Tang. Parameter estimation of two-level nonlinear mixed effects models using first order conditional linearization and the EM algorithm. Computational Statistics & Data Analysis, 69(??): 173–183, January 2014. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313002831.

Stein:2014:AUL

[1261] Markus Chagas Stein, Michel Ferreira da Silva, and Luiz Henrique Duczmal. Alternatives to the usual likelihood ratio test in mixed linear models. Computational Statistics & Data Analysis, 69 (??):184–197, January 2014. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313002880.

Park:2014:EDA

[1262] Jin-Hong Park, Dipankar Bandyopadhyay, and Elizabeth Letourneau. Examining deterrence of adult sex crimes: a semi-parametric intervention timeseries approach. Computational Statistics & Data Analysis, 69(??):198– 207, January 2014. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313002909.

Jiang:2014:ISV

[1263] Liewen Jiang, Howard D. Bondell, and Huixia Judy Wang. Interquantile shrinkage and variable selection in quantile regression. Computational Statistics & Data Analysis, 69(??): 208–219, January 2014. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313002922.

Rosenkranz:2014:BCT

[1264] Gerd K. Rosenkranz. Bootstrap corrections of treatment effect estimates following selection. Computational Statistics & Data Analysis, 69(??): 220–227, January 2014. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313003071.

Guo:2014:DRM

[1265] Xu Guo, Tao Wang, Wangli Xu, and Lixing Zhu. Dimension reduction with missing response at random. Computational Statistics & Data Analysis, 69 (??):228–242, January 2014. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313002879.

Ahn:2014:ACL

[1266] Kwang Woo Ahn and Kung-Sik Chan. Approximate conditional least squares estimation of a nonlinear state-space model via an unscented Kalman filter. Computational Statistics & Data Analysis, 69(??):243– 254, January 2014. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313002843.

Arribas-Gil:2014:PDT

[1267] Ana Arribas-Gil and Hans-Georg Müller. Pairwise dynamic time warping for event data. Computational Statistics & Data Analysis, 69(??):

255–268, January 2014. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947313003083.

Frommlet:2014:CMV

[1268] Florian Frommlet, Felix Ruhaltinger, Piotr Twaróg, and Małgorzata Bogdan. Corrigendum to “Modified versions of Bayesian Information Criterion for genome-wide association studies” [comput. statist. data anal. 56 (5) (2012) 1038–1051]. Computational Statistics & Data Analysis, 69 (??):269, January 2014. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947313002910. See

[703].

Anonymous:2014:EBb

[1269] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 70(??):iii–v, February 2014. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947313003836.

Anonymous:2014:Cb

[1270] Anonymous. Contents. Computational Statistics & Data Analysis, 70 (??):vi–vii, February 2014. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947313003848.

Grossmann:2014:AAV

[1271] Heiko Großmann. Automating the analysis of variance of orthogonal designs. Computational Statistics & Data Analysis, 70(??):1– 18, February 2014. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313003113.

Hautphenne:2014:EAM

[1272] Sophie Hautphenne and Mark Fackrell.

An EM algorithm for the model fitting of Markovian binary trees. Computational Statistics & Data Analysis, 70 (??):19–34, February 2014. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313003125.

Menendez:2014:SAB

[1273] P. Menéndez, Y. Fan, P. H. Garthwaite, and S. A. Sisson. Simultaneous adjustment of bias and coverage probabilities for confidence intervals. Computational Statistics & Data Analysis, 70 (??):35–44, February 2014. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313003137.

Ryan:2014:TBE

[1274] Elizabeth G. Ryan, Christopher C.

Drovandi, M. Helen Thompson, and Anthony N. Pettitt. Towards Bayesian experimental design for nonlinear models that require a large number of sampling times. Computational Statistics & Data Analysis, 70(??):45– 60, February 2014. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947313003149.

Roberts:2014:FAP

[1275] W. J. J. Roberts. Factor analysis parameter estimation from incomplete data. Computational Statistics & Data Analysis, 70(??):61– 66, February 2014. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947313003150.

Maharaj:2014:DAM

[1276] Elizabeth Ann Maharaj and Andrés M. Alonso. Discriminant analysis of multivariate time series: Application to diagnosis based on ECG signals. Computational Statistics & Data Analysis, 70(??):67–87, February 2014. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947313003216.

Cruz-Cano:2014:FRC

[1277] Raul Cruz-Cano and Mei-Ling Ting Lee. Fast regularized canonical correlation analysis. Computational Statistics & Data Analysis, 70(??):88– 100, February 2014. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947313003447.

Wang:2014:SEL

[1278] Suohong Wang and Biao Zhang. Semiparametric empirical likelihood confidence intervals for AUC under a density ratio model. Computational Statistics & Data Analysis, 70(??):

101–115, February 2014. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313002934.

Wu:2014:NLA

[1279] Lan Wu, Yuehan Yang, and Hanzhong Liu. Nonnegative-lasso and application in index tracking. Computational Statistics & Data Analysis, 70 (??):116–126, February 2014. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313003095.

Wong:2014:THG

[1280] Tony Siu Tung Wong and Wai Keung Li. Test for homogeneity in gamma mixture models using likelihood ratio. Computational Statistics & Data Analysis, 70(??):127– 137, February 2014. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313003162.

Luo:2014:EDS

[1281] Shan Luo and Zehua Chen. Edge detection in sparse Gaussian graphical models. Computational Statistics & Data Analysis, 70(??):138– 152, February 2014. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313003174.

Kouassi:2014:THS

[1282] Eugene Kouassi, Mbodja Mougoué, Joel Sango, J. M. Bosson Brou, Claude M. O. Amba, and Afeez Adebare Salisu. Testing for heteroskedasticity and spatial correlation in a two way random effects model. Computational Statistics & Data Analysis, 70(??): 153–171, February 2014. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313003186.

Wang:2014:FMI

[1283] Guochang Wang, Nan Lin, and Baoxue Zhang. Functional k-means inverse regression. Computational Statistics & Data Analysis, 70(??):172– 182, February 2014. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313003198.

Zhang:2014:MDF

[1284] Tao Zhang, Qingzhao Zhang, and Qihua Wang. Model detection for functional polynomial regression. Computational Statistics & Data Analysis, 70 (??):183–197, February 2014. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313003228.

Roberts:2014:SLA

[1285] S. Roberts and G. Nowak. Stabilizing the lasso against cross-validation variability. Computational Statistics & Data Analysis, 70(??):198– 211, February 2014. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794731300323X.

So:2014:DST

[1286] Mike K. P. So and Ray S. W.

Chung. Dynamic seasonality in time series. Computational Statistics & Data Analysis, 70(??):212– 226, February 2014. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313003253.

Lefebvre:2014:EPD

[1287] Geneviève Lefebvre, Juli Atherton, and Denis Talbot. The effect of the prior distribution in the Bayesian adjustment for confounding algorithm. Computational Statistics & Data Analysis, 70(??):227– 240, February 2014. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313003344.

Lai:2014:VSS

[1288] Peng Lai, Qihua Wang, and Xiao-Hua Zhou. Variable selection and semiparametric efficient estimation for the heteroscedastic partially linear single-index model. Computational Statistics & Data Analysis, 70(??): 241–256, February 2014. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313003368.

Campbell:2014:ATP

[1289] David Campbell and Subhash Lele.

An ANOVA test for parameter estimability using data cloning with application to statistical inference for dynamic systems. Computational Statistics & Data Analysis, 70(??):257– 267, February 2014. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794731300337X.

Marhuenda:2014:ICF

[1290] Yolanda Marhuenda, Domingo Morales, and María del Carmen Pardo. Information criteria for Fay–Herriot model selection. Computational Statistics & Data Analysis, 70(??):268– 280, February 2014. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794731300340X.

Kouritzin:2014:GTA

[1291] Michael A. Kouritzin, Fraser Newton, and Biao Wu. A graph theoretic approach to simulation and classification. Computational Statistics & Data Analysis, 70(??):281– 294, February 2014. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313003502.

Xu:2014:CPH

[1292] Wenjing Xu, Qing Pan, and Joseph L. Gastwirth. Cox proportional hazards models with frailty for negatively correlated employment processes. Computational Statistics & Data Analysis, 70 (??):295–307, February 2014. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313003514.

Li:2014:WBT

[1293] Linyuan Li, Shan Yao, and Pierre Duchesne. On wavelet-based testing for serial correlation of unknown form using Fan’s adaptive Neyman method. Computational Statistics & Data Analysis, 70(??):308– 327, February 2014. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313003563.

Puig:2014:CAV

[1294] Xavier Puig and Josep Ginebra. A cluster analysis of vote transitions. Computational Statistics & Data Analysis, 70 (??):328–344, February 2014. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313003599.

Liu:2014:PFD

[1295] Shen Liu, Elizabeth Ann Maharaj, and Brett Inder. Polarization of forecast densities: a new approach to time series classification. Computational Statistics & Data Analysis, 70 (??):345–361, February 2014. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313003617.

Goldsmith:2014:ESC

[1296] Jeff Goldsmith and Fabian Scheipl. Estimator selection and combination in scalar-on-function regression. Computational Statistics & Data Analysis, 70 (??):362–372, February 2014. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947313003629.

Dimeglio:2014:RAT

[1297] Chloé Dimeglio, Santiago Gallón, Jean-Michel Loubes, and Elie Maza. A robust algorithm for template curve estimation based on manifold embedding. Computational Statistics & Data Analysis, 70(??):373– 386, February 2014. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016794731300368X.

Chen:2014:OWT

[1298] Zhongxue Chen and Saralees Nadarajah. On the optimally weighted ztest for combining probabilities from independent studies. Computational Statistics & Data Analysis, 70(??): 387–394, February 2014. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947313003204.

Huang:2014:ALS

[1299] Xiaolin Huang, Lei Shi, and Johan A. K. Suykens. Asymmetric least squares support vector machine classifiers. Computational Statistics & Data Analysis, 70(??):395– 405, February 2014. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947313003393.

Anonymous:2014:EBc

[1300] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 71(??):iii–v, March 2014. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794731300412X.

Anonymous:2014:Cc

[1301] Anonymous. Contents. Computational Statistics & Data Analysis, 71 (??):vi–xii, March 2014. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313004131.

Bohning:2014:SIA

[1302] Dankmar Böhning, Christian Hennig, Geoffrey J. McLachlan, and Paul D. McNicholas. The 2nd special issue on advances in mixture models. Computational Statistics & Data Analysis, 71(??):1–2, March 2014. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313003630.

Vicari:2014:MBC

[1303] Donatella Vicari and Marco Alfó.

Model based clustering of customer choice data. Computational Statistics & Data Analysis, 71(??):3– 13, March 2014. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313003381.

Coffey:2014:CLP

[1304] N. Coffey, J. Hinde, and E. Holian. Clustering longitudinal profiles using P -splines and mixed effects models applied to time-course gene expression data. Computational Statistics & Data Analysis, 71(??): 14–29, March 2014. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794731300131X.

Jaspers:2014:NSP

[1305] Stijn Jaspers, Marc Aerts, Geert Verbeke, and Pierre-Alexandre Beloeil. A new semi-parametric mixture model for interval censored data, with applications in the field of antimicrobial resistance. Computational Statistics & Data Analysis, 71(??): 30–42, March 2014. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794731300039X.

Ng:2014:MMC

[1306] S. K. Ng and G. J. McLachlan. Mixture models for clustering multilevel growth trajectories. Computational Statistics & Data Analysis, 71(??): 43–51, March 2014. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947312004410.

Bouveyron:2014:MBC

[1307] Charles Bouveyron and Camille Brunet-Saumard. Model-based clustering of high-dimensional data: a review. Computational Statistics & Data Analysis, 71(??):52– 78, March 2014. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947312004422.

Calo:2014:HMA

[1308] Daniela G. Calò, Angela Montanari, and Cinzia Viroli. A hierarchical modeling approach for clustering probability density functions. Computational Statistics & Data Analysis, 71 (??):79–91, March 2014. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313001540.

Jacques:2014:MBC

[1309] Julien Jacques and Cristian Preda.

Model-based clustering for multivariate functional data. Computational Statistics & Data Analysis, 71(??): 92–106, March 2014. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947312004380.

Polymenis:2014:CLR

[1310] Athanase Polymenis. A combined likelihood ratio/information ratio bootstrap technique for estimating the number of components in finite mixtures. Computational Statistics & Data Analysis, 71(??):107– 115, March 2014. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313000431.

Yao:2014:RMR

[1311] Weixin Yao, Yan Wei, and Chun Yu. Robust mixture regression using the t-distribution. Computational Statistics & Data Analysis, 71(??): 116–127, March 2014. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947313002648.

Song:2014:RMR

[1312] Weixing Song, Weixin Yao, and Yanru Xing. Robust mixture regression model fitting by Laplace distribution. Computational Statistics & Data Analysis, 71(??):128–137, March 2014. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947313002442.

Galimberti:2014:MLR

[1313] Giuliano Galimberti and Gabriele Soffritti. A multivariate linear regression analysis using finite mixtures of t distributions. Computational Statistics & Data Analysis, 71(??): 138–150, March 2014. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947313000327.

Lim:2014:ZIP

[1314] Hwa Kyung Lim, Wai Keung Li, and Philip L. H. Yu. Zero-inflated Poisson regression mixture model. Computational Statistics & Data Analysis, 71 (??):151–158, March 2014. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313002338.

Ingrassia:2014:MBC

[1315] Salvatore Ingrassia, Simona C. Minotti, and Antonio Punzo. Model-based clustering via linear cluster-weighted models. Computational Statistics & Data Analysis, 71(??):159– 182, March 2014. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794731300056X.

Lin:2014:LID

[1316] Tsung-I Lin. Learning from incomplete data via parameterized t mixture models through eigenvalue decomposition. Computational Statistics & Data Analysis, 71(??):183– 195, March 2014. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313000765.

Vrbik:2014:PSM

[1317] Irene Vrbik and Paul D. McNicholas. Parsimonious skew mixture models for model-based clustering and classification. Computational Statistics & Data Analysis, 71 (??):196–210, March 2014. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313002533.

Lesperance:2014:NEM

[1318] Mary Lesperance, Rabih Saab, and John Neuhaus. Nonparametric estimation of the mixing distribution in logistic regression mixed models with random intercepts and slopes. Computational Statistics & Data Analysis, 71(??):211–219, March 2014. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313001874.

Lu:2014:RGM

[1319] Zhenqiu (Laura) Lu and Zhiyong Zhang. Robust growth mixture models with non-ignorable missingness: Models, estimation, selection, and application. Computational Statistics & Data Analysis, 71(??):220– 240, March 2014. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313002818.

Pledger:2014:MMU

[1320] Shirley Pledger and Richard Arnold. Multivariate methods using mixtures: Correspondence analysis, scaling and pattern-detection. Computational Statistics & Data Analysis, 71(??): 241–261, March 2014. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313001862.

Bacci:2014:MEN

[1321] Silvia Bacci and Francesco Bartolucci. Mixtures of equispaced normal distributions and their use for testing symmetry with univariate data. Computational Statistics & Data Analysis, 71 (??):262–272, March 2014. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313000303.

Chen:2014:SIB

[1322] Cathy W. S. Chen, David Dunson, Sylvia Frühwirth-Schnatter, and Stephen G. Walker. Special issue on Bayesian computing, methods and applications. Computational Statistics & Data Analysis, 71(??): 273, March 2014. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313003642.

Abanto-Valle:2014:SSM

[1323] Carlos A. Abanto-Valle and Dipak K.

Dey. State space mixed models for binary responses with scale mixture of normal distributions links. Computational Statistics & Data Analysis, 71 (??):274–287, March 2014. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313000248.

Azriel:2014:OSD

[1324] David Azriel. Optimal sequential designs in phase I studies. Computational Statistics & Data Analysis, 71 (??):288–297, March 2014. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313001837.

Bekiros:2014:BFS

[1325] Stelios D. Bekiros and Alessia Paccagnini.

Bayesian forecasting with small and medium scale factor-augmented vector autoregressive DSGE models. Computational Statistics & Data Analysis, 71 (??):298–323, March 2014. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313003423.

Chen:2014:HBM

[1326] Bingshu E. Chen, Wenyu Jiang, and Dongsheng Tu. A hierarchical Bayes model for biomarker subset effects in clinical trials. Computational Statistics & Data Analysis, 71(??):324– 334, March 2014. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313002004.

Chen:2014:BNS

[1327] Yuhui Chen and Timothy E. Hanson. Bayesian nonparametric ksample tests for censored and uncensored data. Computational Statistics & Data Analysis, 71(??):335– 346, March 2014. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947312003945.

Curtis:2014:FBM

[1328] S. McKay Curtis, Sayantan Banerjee, and Subhashis Ghosal. Fast Bayesian model assessment for nonparametric additive regression. Computational Statistics & Data Analysis, 71(??):347–358, March 2014. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313001850.

Grassi:2014:CET

[1329] S. Grassi and T. Proietti. Characterising economic trends by Bayesian stochastic model specification search. Computational Statistics & Data Analysis, 71(??):359–374, March 2014. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0167947313000807.

Gupta:2014:EMC

[1330] Mayetri Gupta. An evolutionary Monte Carlo algorithm for Bayesian block clustering of data matrices. Computational Statistics & Data Analysis, 71 (??):375–391, March 2014. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794731300251X.

Heaps:2014:CML

[1331] Sarah E. Heaps, Richard J. Boys, and Malcolm Farrow. Computation of marginal likelihoods with data-dependent support for latent variables. Computational Statistics & Data Analysis, 71(??):392– 401, March 2014. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313002788.

Jin:2014:USB

[1332] Ick Hoon Jin and Faming Liang. Use of SAMC for Bayesian analysis of statistical models with intractable normalizing constants. Computational Statistics & Data Analysis, 71(??): 402–416, March 2014. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947312002770.

Lijoi:2014:DMM

[1333] Antonio Lijoi, Bernardo Nipoti, and Igor Prünster. Dependent mixture models: Clustering and borrowing information. Computational Statistics & Data Analysis, 71(??):417– 433, March 2014. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947313002272.

McKinley:2014:SBB

[1334] Trevelyan J. McKinley, Joshua V. Ross, Rob Deardon, and Alex R. Cook. Simulation-based Bayesian inference for epidemic models. Computational Statistics & Data Analysis, 71 (??):434–447, March 2014. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016794731200446X.

Mulder:2014:PAD

[1335] Joris Mulder. Prior adjusted default Bayes factors for testing (in)equality constrained hypotheses. Computational Statistics & Data Analysis, 71 (??):448–463, March 2014. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947313002624.

Naranjo:2014:BBR

[1336] L. Naranjo, J. Martín, and C. J. Pérez. Bayesian binary regression with exponential power link. Computational Statistics & Data Analysis, 71 (??):464–476, March 2014. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947312002964.

Nieto-Barajas:2014:BSA

[1337] Luis E. Nieto-Barajas. Bayesian semiparametric analysis of shortand long-term hazard ratios with covariates. Computational Statistics & Data Analysis, 71(??):477– 490, March 2014. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794731300114X.

Nott:2014:MEU

[1338] David J. Nott, Lucy Marshall, Mark Fielding, and Shie-Yui Liong. Mixtures of experts for understanding model discrepancy in dynamic computer models. Computational Statistics & Data Analysis, 71(??):491– 505, March 2014. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313001631.

Nunez-Antonio:2014:BML

[1339] Gabriel Nuñez-Antonio and Eduardo Gutiérrez-Peña. A Bayesian model for longitudinal circular data based on the projected normal distribution. Computational Statistics & Data Analysis, 71(??):506–519, March 2014. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016794731200299X.

Qiu:2014:OSB

[1340] Yu Qiu, Daniel J. Nordman, and Stephen B. Vardeman. One-sample Bayes inference for symmetric distributions of 3-D rotations. Computational Statistics & Data Analysis, 71 (??):520–529, March 2014. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947313000480.

Rajala:2014:BAG

[1341] T. Rajala and A. Penttinen. Bayesian analysis of a Gibbs hard-core point pattern model with varying repulsion range. Computational Statistics & Data Analysis, 71(??):530– 541, March 2014. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947312003179.

Sabourin:2014:BDM

[1342] Anne Sabourin and Philippe Naveau. Bayesian Dirichlet mixture model for multivariate extremes: a reparametrization. Computational Statistics & Data Analysis, 71(??): 542–567, March 2014. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947313001680.

So:2014:BAT

[1343] Mike K. P. So and Raymond K. S. Chan. Bayesian analysis of tail asymmetry based on a threshold extreme value model. Computational Statistics & Data Analysis, 71(??): 568–587, March 2014. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313000522.

Soneson:2014:IGE

[1344] Charlotte Soneson and Magnus Fontes.

Incorporation of gene exchangeabilities improves the reproducibility of gene set rankings. Computational Statistics & Data Analysis, 71(??): 588–598, March 2014. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947312003003.

Spezia:2014:MSA

[1345] L. Spezia, S. L. Cooksley, M. J.

Brewer, D. Donnelly, and A. Tree. Modelling species abundance in a river by negative binomial hidden Markov models. Computational Statistics & Data Analysis, 71(??):599– 614, March 2014. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313003411.

vanderMeulen:2014:RJM

[1346] Frank van der Meulen, Moritz Schauer, and Harry van Zanten. Reversible jump MCMC for nonparametric drift estimation for diffusion processes. Computational Statistics & Data Analysis, 71 (??):615–632, March 2014. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016794731300090X.

Wang:2014:LBE

[1347] Lichun Wang and Radhey S. Singh. Linear Bayes estimator for the twoparameter exponential family under type II censoring. Computational Statistics & Data Analysis, 71(??): 633–642, March 2014. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016794731300265X.

White:2014:GAM

[1348] Gentry White and Michael D. Porter. GPU accelerated MCMC for modeling terrorist activity. Computational Statistics & Data Analysis, 71 (??):643–651, March 2014. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947313001291.

Ahmed:2014:SIM

[1349] S. Ejaz Ahmed, Gerda Claeskens, Hidetoshi Shimodaira, and Stefan Van Aelst. Special issue on model selection and high dimensional data reduction. Computational Statistics & Data Analysis, 71(??):652– 653, March 2014. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947313004258.

Beran:2014:HEP

[1350] Rudolf Beran. Hypercube estimators: Penalized least squares, submodel selection, and numerical stability. Computational Statistics & Data Analysis, 71(??):654–666, March 2014. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313002077.

Blommaert:2014:DML

[1351] A. Blommaert, N. Hens, and Ph. Beutels. Data mining for longitudinal data under multicollinearity and time dependence using penalized generalized estimating equations. Computational Statistics & Data Analysis, 71 (??):667–680, March 2014. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313000790.

Dernoncourt:2014:AFS

[1352] David Dernoncourt, Blaise Hanczar, and Jean-Daniel Zucker. Analysis of feature selection stability on high dimension and small sample data. Computational Statistics & Data Analysis, 71(??):681–693, March 2014. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313002570.

Hall:2014:SIF

[1353] Peter Hall and Jing-Hao Xue. On selecting interacting features from highdimensional data. Computational Statistics & Data Analysis, 71(??): 694–708, March 2014. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794731200360X.

Martins:2014:LIV

[1354] Luis F. Martins and Vasco J. Gabriel. Linear instrumental variables model averaging estimation. Computational Statistics & Data Analysis, 71(??): 709–724, March 2014. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947313001813.

Mielniczuk:2014:URS

[1355] Jan Mielniczuk and Paweł Teisseyre. Using random subspace method for prediction and variable importance assessment in linear regression. Computational Statistics & Data Analysis, 71 (??):725–742, March 2014. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947312003477.

Mougeot:2014:LSH

[1356] M. Mougeot, D. Picard, and K. Tribouley. LOL selection in high dimension. Computational Statistics & Data Analysis, 71(??):743– 757, March 2014. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947312001764.

Schomaker:2014:MSM

[1357] Michael Schomaker and Christian Heumann. Model selection and model averaging after multiple imputation. Computational Statistics & Data Analysis, 71(??):758– 770, March 2014. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794731300073X.

Vincent:2014:SGL

[1358] Martin Vincent and Niels Richard Hansen. Sparse group lasso and high dimensional multinomial classification. Computational Statistics & Data Analysis, 71(??):771– 786, March 2014. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313002168.

Colubi:2014:SII

[1359] Ana Colubi and Thierry Denoeux.

Special issue on imprecision in statistical data analysis. Computational Statistics & Data Analysis, 71(??): 787–788, March 2014. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313003666.

Abellan:2014:CDT

[1360] Joaquín Abellán, Rebecca M. Baker, Frank P. A. Coolen, Richard J. Crossman, and Andrés R. Masegosa. Classification with decision trees from a nonparametric predictive inference perspective. Computational Statistics & Data Analysis, 71(??):789– 802, March 2014. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313000534.

Choirat:2014:BCS

[1361] Christine Choirat and Raffaello Seri. Bootstrap confidence sets for the Aumann mean of a random closed set. Computational Statistics & Data Analysis, 71(??):803–817, March 2014. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0167947312003830.

Corani:2014:CEC

[1362] G. Corani and A. Antonucci. Credal ensembles of classifiers. Computational Statistics & Data Analysis, 71 (??):818–831, March 2014. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016794731200401X.

Frenay:2014:EMI

[1363] Benoît Frénay, Gauthier Doquire, and Michel Verleysen. Estimating mutual information for feature selection in the presence of label noise. Computational Statistics & Data Analysis, 71 (??):832–848, March 2014. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016794731300159X.

Hannig:2014:CIG

[1364] Jan Hannig, Randy C. S. Lai, and Thomas C. M. Lee. Computational issues of generalized fiducial inference. Computational Statistics & Data Analysis, 71(??):849– 858, March 2014. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313000911.

Lopez-Fidalgo:2014:OED

[1365] J. López-Fidalgo and M. J. Rivas-López. Optimal experimental designs for partial likelihood information. Computational Statistics & Data Analysis, 71(??):859– 867, March 2014. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947312003593.

Montes:2014:SDI

[1366] Ignacio Montes, Enrique Miranda, and Susana Montes. Stochastic dominance with imprecise information. Computational Statistics & Data Analysis, 71 (??):868–886, March 2014. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947312003192.

Filzmoser:2014:SIS

[1367] Peter Filzmoser, Cristian Gatu, and Achim Zeileis. Special issue on statistical algorithms and software in R. Computational Statistics & Data Analysis, 71(??):887–888, March 2014. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313003654.

Yee:2014:RRV

[1368] Thomas W. Yee. Reduced-rank vector generalized linear models with two linear predictors. Computational Statistics & Data Analysis, 71(??):

889–902, March 2014. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313000273.

Polpo:2014:TBS

[1369] A. Polpo, C. P. de Campos, D. Sinha, S. Lipsitz, and J. Lin. Transform both sides model: a parametric approach. Computational Statistics & Data Analysis, 71(??):903– 913, March 2014. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313002685.

Millo:2014:MLE

[1370] Giovanni Millo. Maximum likelihood estimation of spatially and serially correlated panels with random effects. Computational Statistics & Data Analysis, 71(??):914– 933, March 2014. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313002697.

Golyandina:2014:BSS

[1371] Nina Golyandina and Anton Korobeynikov. Basic singular spectrum analysis and forecasting with R. Computational Statistics & Data Analysis, 71(??):934–954, March 2014. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313001394.

Asquith:2014:PEP

[1372] William H. Asquith. Parameter estimation for the 4-parameter Asymmetric Exponential Power distribution by the method of L-moments using R. Computational Statistics & Data Analysis, 71(??):955–970, March 2014. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947312004471.

Bartolucci:2014:MRP

[1373] Francesco Bartolucci, Silvia Bacci, and Michela Gnaldi. MultiLCIRT: an R package for multidimensional latent class item response models. Computational Statistics & Data Analysis, 71 (??):971–985, March 2014. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313002053.

Eugster:2014:PAB

[1374] Manuel J. A. Eugster, Friedrich Leisch, and Carolin Strobl. (psycho-)analysis of benchmark experiments: a formal framework for investigating the relationship between data sets and learning algorithms. Computational Statistics & Data Analysis, 71(??): 986–1000, March 2014. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313002946.

Wang:2014:DBD

[1375] Liqun Wang and Chel Hee Lee.

Discretization-based direct random sample generation. Computational Statistics & Data Analysis, 71(??): 1001–1010, March 2014. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313002235.

Cheng:2014:GBR

[1376] Ching-Wei Cheng, Ying-Chao Hung, and Narayanaswamy Balakrishnan. Generating beta random numbers and Dirichlet random vectors in R: the package rBeta2009. Computational Statistics & Data Analysis, 71(??): 1011–1020, March 2014. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313000753.

Chevalier:2014:KEU

[1377] Clément Chevalier, Victor Picheny, and David Ginsbourger. KrigInv: an efficient and user-friendly implementation of batch-sequential inversion strategies based on kriging. Computational Statistics & Data Analysis, 71 (??):1021–1034, March 2014. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313001060.

Picheny:2014:NKB

[1378] Victor Picheny and David Ginsbourger. Noisy kriging-based optimization methods: a unified implementation within the DiceOptim package. Computational Statistics & Data Analysis, 71(??):1035– 1053, March 2014. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313001205.

Eddelbuettel:2014:RAR

[1379] Dirk Eddelbuettel and Conrad Sanderson. RcppArmadillo: Accelerating R with high-performance C++ linear algebra. Computational Statistics & Data Analysis, 71(??):1054– 1063, March 2014. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313000492.

Gilmour:2014:SIA

[1380] Steven Gilmour and Roger W. Payne.

Special issue on algorithms for design of experiments. Computational Statistics & Data Analysis, 71(??): 1064–1065, March 2014. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794731300354X.

Abebe:2014:BOD

[1381] Haftom T. Abebe, Frans E. S. Tan, Gerard J. P. Van Breukelen, and Martijn P. F. Berger. Bayesian D-optimal designs for the two parameter logistic mixed effects model. Computational Statistics & Data Analysis, 71 (??):1066–1076, March 2014. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313002867.

Almohaimeed:2014:EDD

[1382] B. Almohaimeed and A. N. Donev.

Experimental designs for drug combination studies. Computational Statistics & Data Analysis, 71(??):1077– 1087, March 2014. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794731300008X.

Boukouvalas:2014:ODC

[1383] A. Boukouvalas, D. Cornford, and M. Stehlík. Optimal design for correlated processes with inputdependent noise. Computational Statistics & Data Analysis, 71(??):1088– 1102, March 2014. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313003484.

Dette:2014:NUO

[1384] Holger Dette, Andrey Pepelyshev, and Anatoly Zhigljavsky. ‘Nearly’ universally optimal designs for models with correlated observations. Computational Statistics & Data Analysis, 71 (??):1103–1112, March 2014. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313000467.

Gaffke:2014:AAL

[1385] N. Gaffke, U. Graßhoff, and R. Schwabe. Algorithms for approximate linear regression design with application to a first order model with heteroscedasticity. Computational Statistics & Data Analysis, 71(??):1113– 1123, March 2014. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313002740.

Georgiou:2014:CCD

[1386] Stelios D. Georgiou, Stella Stylianou, and Manohar Aggarwal. A class of composite designs for response surface methodology. Computational Statistics & Data Analysis, 71(??):1124– 1133, March 2014. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313001084.

Godolphin:2014:EPA

[1387] J. D. Godolphin and H. R. Warren. An efficient procedure for the avoidance of disconnected incomplete block designs. Computational Statistics & Data Analysis, 71(??):1134– 1146, March 2014. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313003496.

Gutman:2014:ASD

[1388] Alex J. Gutman, Edward D. White, Dennis K. J. Lin, and Raymond R. Hill. Augmenting supersaturated designs with Bayesian D-optimality. Computational Statistics & Data Analysis, 71 (??):1147–1158, March 2014. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313003241.

Harman:2014:CEE

[1389] Radoslav Harman and Lenka Filová.

Computing efficient exact designs of experiments using integer quadratic programming. Computational Statistics & Data Analysis, 71(??):1159– 1167, March 2014. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313000777.

Loeza-Serrano:2014:CED

[1390] S. Loeza-Serrano and A. N. Donev. Construction of experimental designs for estimating variance components. Computational Statistics & Data Analysis, 71(??):1168– 1177, March 2014. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947312003581.

Lu:2014:ODE

[1391] Lu Lu, Christine M. Anderson-Cook, and Dennis K. J. Lin. Optimal designed experiments using a Pareto front search for focused preference of multiple objectives. Computational Statistics & Data Analysis, 71(??): 1178–1192, March 2014. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313001382.

Sambo:2014:CET

[1392] Francesco Sambo, Matteo Borrotti, and Kalliopi Mylona. A coordinateexchange two-phase local search algorithm for the D- and I-optimal designs of split-plot experiments. Computational Statistics & Data Analysis, 71 (??):1193–1207, March 2014. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313001175.

Tommasi:2014:IAC

[1393] C. Tommasi, J. M. Rodríguez-Díaz, and M. T. Santos-Martín. Integral approximations for computing optimum designs in random effects logistic regression models. Computational Statistics & Data Analysis, 71 (??):1208–1220, March 2014. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794731200223X.

Anonymous:2014:EBd

[1394] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 72(??):iii–v, April 2014. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313004647.

Anonymous:2014:Cd

[1395] Anonymous. Contents. Computational Statistics & Data Analysis, 72 (??):vi–vii, April 2014. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313004659.

Huang:2014:FIP

[1396] Hengzhen Huang, Jinyu Yang, and Min-Qian Liu. Functionally induced priors for componentwise Gibbs sampler in the analysis of supersaturated designs. Computational Statistics & Data Analysis, 72(??): 1–12, April 2014. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313003769.

Turnbull:2014:UDE

[1397] Bradley C. Turnbull and Sujit K. Ghosh. Unimodal density estimation using Bernstein polynomials. Computational Statistics & Data Analysis, 72(??):13–29, April 2014. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947313003757.

Bonneau:2014:RLB

[1398] Mathieu Bonneau, Sabrina Gaba, Nathalie Peyrard, and Régis Sabbadin. Reinforcement learning-based design of sampling policies under cost constraints in Markov random fields: Application to weed map reconstruction. Computational Statistics & Data Analysis, 72(??):30–44, April 2014. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947313003551.

Colubi:2014:TCM

[1399] Ana Colubi, J. Santos Domínguez-Menchero, and Gil González-Rodríguez. Testing constancy in monotone response models. Computational Statistics & Data Analysis, 72(??):45– 56, April 2014. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947313003976.

Malec:2014:NKD

[1400] Peter Malec and Melanie Schienle. Nonparametric kernel density estimation near the boundary. Computational Statistics & Data Analysis, 72(??):57–76, April 2014. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313003770.

Sinha:2014:ILD

[1401] Sanjoy K. Sinha, Amit Kaushal, and Wenzhong Xiao. Inference for longitudinal data with nonignorable nonmonotone missing responses. Computational Statistics & Data Analysis, 72(??):77–91, April 2014. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313003952.

Doove:2014:RPM

[1402] L. L. Doove, S. Van Buuren, and E. Dusseldorp. Recursive partitioning for missing data imputation in the presence of interaction effects. Computational Statistics & Data Analysis, 72(??):92–104, April 2014. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313003939.

Todorov:2014:TEE

[1403] Diman Todorov and Rossi Setchi.

Time-efficient estimation of conditional mutual information for variable selection in classification. Computational Statistics & Data Analysis, 72 (??):105–127, April 2014. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313003940.

Drovandi:2014:MRJ

[1404] Christopher C. Drovandi, Anthony N. Pettitt, Robert D. Henderson, and Pamela A. McCombe. Marginal reversible jump Markov chain Monte Carlo with application to motor unit number estimation. Computational Statistics & Data Analysis, 72(??): 128–146, April 2014. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947313004015.

Oh:2014:BTE

[1405] Man-Suk Oh. Bayesian test on equality of score parameters in the order restricted RC association model. Computational Statistics & Data Analysis, 72(??):147–157, April 2014. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947313004477.

Tang:2014:NVS

[1406] Shijie Tang, Lisha Chen, Kam-Wah Tsui, and Kjell Doksum. Nonparametric variable selection and classification: the CATCH algorithm. Computational Statistics & Data Analysis, 72(??):158–175, April 2014. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947313003782.

Beaton:2014:EMA

[1407] Derek Beaton, Cherise R. Chin Fatt, and Hervé Abdi. An ExPosition of multivariate analysis with the singular value decomposition in R. Computational Statistics & Data Analysis, 72(??):176–189, April 2014. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313004441.

Segers:2014:NET

[1408] Johan Segers and Nathan Uyttendaele. Nonparametric estimation of the tree structure of a nested Archimedean copula. Computational Statistics & Data Analysis, 72(??): 190–204, April 2014. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313003964.

Zhao:2014:ALF

[1409] Jianhua Zhao and Lei Shi. Automated learning of factor analysis with complete and incomplete data. Computational Statistics & Data Analysis, 72 (??):205–218, April 2014. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313004465.

Franklin:2014:PSE

[1410] Jessica M. Franklin, Sebastian Schneeweiß, Jennifer M. Polinski, and Jeremy A. Rassen. Plasmode simulation for the evaluation of pharmacoepidemiologic methods in complex healthcare databases. Computational Statistics & Data Analysis, 72(??):219– 226, April 2014. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313003721.

Raket:2014:AIS

[1411] Lars Lau Rakêt and Bo Markussen. Approximate inference for spatial functional data on massively parallel processors. Computational Statistics & Data Analysis, 72(??):227– 240, April 2014. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947313003708.

Channouf:2014:PSS

[1412] Nabil Channouf, Marc Fredette, and Brenda MacGibbon. Power and sample size calculations for Poisson and zero-inflated Poisson regression models. Computational Statistics & Data Analysis, 72(??):241– 251, April 2014. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947313003538.

Moffa:2014:SMC

[1413] Giusi Moffa and Jack Kuipers. Sequential Monte Carlo EM for multivariate probit models. Computational Statistics & Data Analysis, 72 (??):252–272, April 2014. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947313003733.

Portnoy:2014:JEI

[1414] Stephen Portnoy. The jackknife’s edge: Inference for censored regression quantiles. Computational Statistics & Data Analysis, 72(??): 273–281, April 2014. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794731300371X.

Chen:2014:DPS

[1415] Ray-Bing Chen, Yen-Wen Hsu, Ying Hung, and Weichung Wang. Discrete particle swarm optimization for constructing uniform design on irregular regions. Computational Statistics & Data Analysis, 72(??):282– 297, April 2014. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313003691.

Pandolfi:2014:GMT

[1416] Silvia Pandolfi, Francesco Bartolucci, and Nial Friel. A generalized multipletry version of the reversible jump algorithm. Computational Statistics & Data Analysis, 72(??):298– 314, April 2014. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313003605.

Anonymous:2014:EBe

[1417] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 73(??):iii–v, May 2014. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947314000073.

Anonymous:2014:Ce

[1418] Anonymous. Contents. Computational Statistics & Data Analysis, 73 (??):vi–vii, May 2014. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947314000085.

Senturk:2014:ETV

[1419] Damla Şentürk, Samiran Ghosh, and Danh V. Nguyen. Exploratory time varying lagged regression: Modeling association of cognitive and functional trajectories with expected clinic visits in older adults. Computational Statistics & Data Analysis, 73 (??):1–15, May 2014. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794731300399X.

Austin:2014:EBD

[1420] Matthew D. Austin and Rebecca A. Betensky. Eliminating bias due to censoring in Kendall’s tau estimators for quasi-independence of truncation and failure. Computational Statistics & Data Analysis, 73(??): 16–26, May 2014. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313004568.

Fujita:2014:NPM

[1421] André Fujita, Daniel Y. Takahashi, and Alexandre G. Patriota. A nonparametric method to estimate the number of clusters. Computational Statistics & Data Analysis, 73(??): 27–39, May 2014. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313004507.

Butler:2014:EOL

[1422] A. Butler, R. D. Haynes, T. D.

Humphries, and P. Ranjan. Efficient optimization of the likelihood function in Gaussian process modelling. Computational Statistics & Data Analysis, 73(??):40–52, May 2014. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313004556.

Rachdi:2014:TPA

[1423] Mustapha Rachdi, Ali Laksaci, Jacques Demongeot, Abdel Abdali, and Fethi Madani. Theoretical and practical aspects of the quadratic error in the local linear estimation of the conditional density for functional data. Computational Statistics & Data Analysis, 73(??):53–68, May 2014. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313004490.

Dufour:2014:ADQ

[1424] Jean-Marie Dufour and Tarek Jouini.

Asymptotic distributions for quasiefficient estimators in echelon VARMA models. Computational Statistics & Data Analysis, 73(??):69– 86, May 2014. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313004003.

Roy:2014:EEL

[1425] Vivekananda Roy. Efficient estimation of the link function parameter in a robust Bayesian binary regression model. Computational Statistics & Data Analysis, 73(??): 87–102, May 2014. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313004519.

Li:2014:NTP

[1426] Yang Li, Hui Zhao, Jianguo Sun, and KyungMann Kim. Nonparametric tests for panel count data with unequal observation processes. Computational Statistics & Data Analysis, 73(??):103–111, May 2014. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313004520.

Nourmohammadi:2014:CIQ

[1427] Mohammad Nourmohammadi, Mohammad Jafari Jozani, and Brad C. Johnson. Confidence intervals for quantiles in finite populations with randomized nomination sampling. Computational Statistics & Data Analysis, 73(??):112–128, May 2014. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313004581.

Huang:2014:ADE

[1428] Shih-Feng Huang and Ya-Ting Tu. Asymptotic distribution of the EPMS estimator for financial derivatives pricing. Computational Statistics & Data Analysis, 73(??):129– 145, May 2014. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313004593.

Karl:2014:CML

[1429] Andrew T. Karl, Yan Yang, and Sharon L. Lohr. Computation of maximum likelihood estimates for multiresponse generalized linear mixed models with non-nested, correlated random effects. Computational Statistics & Data Analysis, 73(??):146– 162, May 2014. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794731300457X.

Luts:2014:MFV

[1430] Jan Luts and John T. Ormerod.

Mean field variational Bayesian inference for support vector machine classification. Computational Statistics & Data Analysis, 73(??):163– 176, May 2014. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313003988.

Li:2014:BSR

[1431] Li Li and Timothy E. Hanson. A Bayesian semiparametric regression model for reliability data using effective age. Computational Statistics & Data Analysis, 73(??):177– 188, May 2014. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313004532.

Lloyd:2014:MTR

[1432] Louise K. Lloyd and Jonathan J. Forster. Modelling trends in road accident frequency — Bayesian inference for rates with uncertain exposure. Computational Statistics & Data Analysis, 73(??):189– 204, May 2014. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947313003745.

Anonymous:2014:EBf

[1433] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 74(??):iii–v, June 2014. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947314000413.

Anonymous:2014:Cf

[1434] Anonymous. Contents. Computational Statistics & Data Analysis, 74 (??):vi–vii, June 2014. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947314000425.

Fan:2014:WTR

[1435] Chunpeng Fan and Donghui Zhang. Wald-type rank tests: a GEE approach. Computational Statistics & Data Analysis, 74(??):1– 16, June 2014. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947313004738.

Chen:2014:IMT

[1436] Shu-Chuan (Grace) Chen and Bruce Lindsay. Improving mixture tree construction using better EM algorithms. Computational Statistics & Data Analysis, 74(??):17–25, June 2014. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313004489.

Thulin:2014:HDT

[1437] Måns Thulin. A high-dimensional two-sample test for the mean using random subspaces. Computational Statistics & Data Analysis, 74 (??):26–38, June 2014. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313004726.

Su:2014:SSD

[1438] Pei-Fang Su, Chung-I Li, and Yu Shyr.

Sample size determination for paired right-censored data based on the difference of Kaplan–Meier estimates. Computational Statistics & Data Analysis, 74(??):39–51, June 2014. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313004751.

He:2014:AMS

[1439] W. He. Analysis of multivariate survival data with Clayton regression models under conditional and marginal formulations. Computational Statistics & Data Analysis, 74(??): 52–63, June 2014. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947314000115.

Cabral:2014:DLM

[1440] Celso Rômulo Barbosa Cabral, Cibele Queiroz da Silva, and Helio S. Migon. A dynamic linear model with extended skew-normal for the initial distribution of the state parameter. Computational Statistics & Data Analysis, 74(??):64–80, June 2014. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947314000024.

Hasler:2014:FBS

[1441] Caren Hasler and Yves Tillé. Fast balanced sampling for highly stratified population. Computational Statistics & Data Analysis, 74(??): 81–94, June 2014. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016794731300474X.

Chen:2014:TTV

[1442] Ray-Bing Chen, Ying Chen, and Wolfgang K. Härdle. TVICA — Time varying independent component analysis and its application to financial data. Computational Statistics & Data Analysis, 74(??): 95–109, June 2014. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947314000127.

Vargas:2014:ILI

[1443] Tiago M. Vargas, Silvia L. P. Ferrari, and Artur J. Lemonte. Improved likelihood inference in generalized linear models. Computational Statistics & Data Analysis, 74(??): 110–124, June 2014. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313004714.

Schelin:2014:SPP

[1444] Lina Schelin and Sara Sjöstedt de Luna. Spatial prediction in the presence of left-censoring. Computational Statistics & Data Analysis, 74 (??):125–141, June 2014. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947314000152.

Ma:2014:MPL

[1445] Jun Ma, Stephane Heritier, and Serigne N. Lô. On the maximum penalized likelihood approach for proportional hazard models with right censored survival data. Computational Statistics & Data Analysis, 74 (??):142–156, June 2014. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947314000164.

Alfaro:2014:DRP

[1446] Carlos A. Alfaro, Burcu Aydın, Carlos E. Valencia, Elizabeth Bullitt, and Alim Ladha. Dimension reduction in principal component analysis for trees. Computational Statistics & Data Analysis, 74(??):157– 179, June 2014. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313004763.

Bidram:2014:NLF

[1447] Hamid Bidram. A note on “Louzada, F., Roman, M., Cancho, V. G., 2011. “The complementary exponential geometric distribution: Model, properties, and a comparison with its counterpart”. Comput. Statist. Data Anal., 55, 2516–2524”. Computational Statistics & Data Analysis, 74 (??):180, June 2014. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947314000176. See

[536].

Yousef:2014:LAM

[1448] Waleed A. Yousef and Subrata Kundu. Learning algorithms may perform worse with increasing training set size: Algorithm–data incompatibility. Computational Statistics & Data Analysis, 74(??):181–197, June 2014. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313002089.

Pan:2014:BSM

[1449] Chun Pan, Bo Cai, Lianming Wang, and Xiaoyan Lin. Bayesian semiparametric model for spatially correlated interval-censored survival data. Computational Statistics & Data Analysis, 74(??):198–208, June 2014. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313004544.

Chee:2014:LSE

[1450] Chew-Seng Chee and Yong Wang.

Least squares estimation of a kmonotone density function. Computational Statistics & Data Analysis, 74(??):209–216, June 2014. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947314000188.

Wang:2014:SSC

[1451] Songfeng Wang, Jiajia Zhang, and Wenbin Lu. Sample size calculation for the proportional hazards model with a time-dependent covariate. Computational Statistics & Data Analysis, 74(??):217–227, June 2014. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947314000292.

Anonymous:2014:EBg

[1452] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 75(??):iii–v, July 2014. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947314000711.

Anonymous:2014:Cg

[1453] Anonymous. Contents. Computational Statistics & Data Analysis, 75 (??):vi–vii, July 2014. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947314000723.

Rezaie:2014:KFV

[1454] Javad Rezaie and Jo Eidsvik. Kalman filter variants in the closed skew normal setting. Computational Statistics & Data Analysis, 75(??): 1–14, July 2014. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947314000255.

Maurya:2014:JCP

[1455] Ashwini Maurya. A joint convex penalty for inverse covariance matrix estimation. Computational Statistics & Data Analysis, 75(??): 15–27, July 2014. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947314000267.

Zougab:2014:BEA

[1456] Nabil Zougab, Smail Adjabi, and Célestin C. Kokonendji. Bayesian estimation of adaptive bandwidth matrices in multivariate kernel density estimation. Computational Statistics & Data Analysis, 75(??): 28–38, July 2014. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947314000322.

Guo:2014:GSS

[1457] Yi Guo, Mark Berman, and Junbin Gao. Group subset selection for linear regression. Computational Statistics & Data Analysis, 75(??):

39–52, July 2014. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794731400036X.

Lee:2014:BVS

[1458] Kyeong Eun Lee, Yongku Kim, and Ronghui Xu. Bayesian variable selection under the proportional hazards mixed-effects model. Computational Statistics & Data Analysis, 75(??):53–65, July 2014. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947314000498.

Xu:2014:REP

[1459] Yihuan Xu, Boris Iglewicz, and Inna Chervoneva. Robust estimation of the parameters of g- and h-distributions, with applications to outlier detection. Computational Statistics & Data Analysis, 75(??):66–80, July 2014. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947314000140.

Loza-Reyes:2014:CMS

[1460] E. Loza-Reyes, M. A. Hurn, and A. Robinson. Classification of molecular sequence data using Bayesian phylogenetic mixture models. Computational Statistics & Data Analysis, 75(??):81–95, July 2014. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794731400019X.

Zhang:2014:LCL

[1461] Mimi Zhang, Qingpei Hu, Min Xie, and Dan Yu. Lower confidence limit for reliability based on grouped data using a quantile-filling algorithm. Computational Statistics & Data Analysis, 75(??):96–111, July 2014. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947314000218.

Hirukawa:2014:NBR

[1462] Masayuki Hirukawa and Mari Sakudo. Nonnegative bias reduction methods for density estimation using asymmetric kernels. Computational Statistics & Data Analysis, 75(??):112– 123, July 2014. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947314000231.

Nieto-Reyes:2014:RPB

[1463] Alicia Nieto-Reyes, Juan Antonio Cuesta-Albertos, and Fabrice Gamboa. A random-projection based test of Gaussianity for stationary processes. Computational Statistics & Data Analysis, 75(??):124– 141, July 2014. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947314000243.

DeCanditiis:2014:FBS

[1464] Daniela De Canditiis. A frame based shrinkage procedure for fast oscillating functions. Computational Statistics & Data Analysis, 75(??):

142–150, July 2014. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947314000279.

Rauch:2014:CAA

[1465] G. Rauch, M. Wirths, and M. Kieser.

Consistency-adjusted alpha allocation methods for a time-to-event analysis of composite endpoints. Computational Statistics & Data Analysis, 75(??):151–161, July 2014. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947314000280.

Zhang:2014:FPD

[1466] Tonglin Zhang and Ge Lin. Family of power divergence spatial scan statistics. Computational Statistics & Data Analysis, 75(??):162– 178, July 2014. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947314000346.

Chen:2014:FDT

[1467] Yen-Hung Chen and Nan-Jung Hsu.

A frequency domain test for detecting nonstationary time series. Computational Statistics & Data Analysis, 75(??):179–189, July 2014. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947314000462.

Braun:2014:CGL

[1468] Julia Braun, Daniel Sabanés Bové, and Leonhard Held. Choice of generalized linear mixed models using predictive crossvalidation. Computational Statistics & Data Analysis, 75 (??):190–202, July 2014. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947314000486.

Stegeman:2014:FLD

[1469] Alwin Stegeman. Finding the limit of diverging components in three-way Candecomp/Parafac — a demonstration of its practical merits. Computational Statistics & Data Analysis, 75(??):203–216, July 2014. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947314000504.

Safarkhani:2014:ICO

[1470] Maryam Safarkhani and Mirjam Moerbeek. The influence of a covariate on optimal designs in longitudinal studies with discrete-time survival endpoints. Computational Statistics & Data Analysis, 75(??): 217–226, July 2014. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947314000528.

Baran:2014:PWS

[1471] Sándor Baran. Probabilistic wind speed forecasting using Bayesian model averaging with truncated normal components. Computational Statistics & Data Analysis, 75(??): 227–238, July 2014. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794731400053X.

Li:2014:HAR

[1472] Zhengbang Li, Qizhai Li, Chien-Pai Han, and Bo Li. A hybrid approach for regression analysis with block missing data. Computational Statistics & Data Analysis, 75(??): 239–247, July 2014. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947314000541.

Azadbakhsh:2014:CCI

[1473] Mahdis Azadbakhsh, Hanna Jankowski, and Xin Gao. Computing confidence intervals for log-concave densities. Computational Statistics & Data Analysis, 75(??):248– 264, July 2014. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947314000334.

Anonymous:2014:EBh

[1474] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 76(??):iii–v, August 2014. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947314001066.

Anonymous:2014:Ch

[1475] Anonymous. Contents. Computational Statistics & Data Analysis, 76 (??):vi–ix, August 2014. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947314001078.

Kontoghiorghes:2014:CAC

[1476] Erricos J. Kontoghiorghes, Herman K. Van Dijk, David A. Belsley, Tim Bollerslev, Francis X. Diebold, Jean-Marie Dufour, Robert Engle, Andrew Harvey, Siem Jan Koopman, Hashem Pesaran, Peter C. B. Phillips, Richard J. Smith, Mike West, Qiwei Yao, Alessandra Amendola, Monica Billio, Cathy W. S. Chen, Carl Chiarella, Ana Colubi, Manfred Deistler, Christian Francq, et al. CFEnetwork: the Annals of Computational and Financial Econometrics: 2nd issue. Computational Statistics & Data Analysis, 76(??):1–3, August 2014. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947314001133.

Addo:2014:UMS

[1477] Peter Martey Addo, Monica Billio, and Dominique Guégan. The univariate MT-STAR model and a new linearity and unit root test procedure. Computational Statistics & Data Analysis, 76(??):4–19, August 2014. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947314000139.

Afonso:2014:SCR

[1478] António Afonso, Pedro Gomes, and Abderrahim Taamouti. Sovereign credit ratings, market volatility, and financial gains. Computational Statistics & Data Analysis, 76(??):20– 33, August 2014. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313003526.

Andres:2014:MLE

[1479] Philipp Andres. Maximum likelihood estimates for positive valued dynamic score models; the DySco package. Computational Statistics & Data Analysis, 76(??):34–42, August 2014. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313004027.

Audrino:2014:FCD

[1480] Francesco Audrino. Forecasting correlations during the late-2000s financial crisis: the short-run component, the long-run component, and structural breaks. Computational Statistics & Data Analysis, 76(??): 43–60, August 2014. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313002144.

Augustyniak:2014:MLE

[1481] Maciej Augustyniak. Maximum likelihood estimation of the Markovswitching GARCH model. Computational Statistics & Data Analysis, 76 (??):61–75, August 2014. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313000418.

Auray:2014:MTA

[1482] Stéphane Auray, Aurélien Eyquem, and Frédéric Jouneau-Sion. Modeling tails of aggregate economic processes in a stochastic growth model. Computational Statistics & Data Analysis, 76(??):76–94, August 2014. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947314000516.

Bada:2014:PCP

[1483] Oualid Bada and Alois Kneip. Parameter cascading for panel models with unknown number of unobserved factors: an application to the credit spread puzzle. Computational Statistics & Data Analysis, 76(??): 95–115, August 2014. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313004453.

Baillie:2014:MIC

[1484] Richard T. Baillie, George Kapetanios, and Fotis Papailias. Modified information criteria and selection of long memory time series models. Computational Statistics & Data Analysis, 76 (??):116–131, August 2014. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313001539.

Bergmeir:2014:UCV

[1485] Christoph Bergmeir, Mauro Costantini, and José M. Benítez. On the usefulness of cross-validation for directional forecast evaluation. Computational Statistics & Data Analysis, 76 (??):132–143, August 2014. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947314000310.

Bos:2014:LMS

[1486] Charles S. Bos, Siem Jan Koopman, and Marius Ooms. Long memory with stochastic variance model: a recursive analysis for US inflation. Computational Statistics & Data Analysis, 76 (??):144–157, August 2014. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947312004215.

Calzolari:2014:EGT

[1487] Giorgio Calzolari, Roxana Halbleib, and Alessandro Parrini. Estimating GARCH-type models with symmetric stable innovations: Indirect inference versus maximum likelihood. Computational Statistics & Data Analysis, 76 (??):158–171, August 2014. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313002739.

Caporin:2014:RRM

[1488] Massimiliano Caporin and Michael McAleer. Robust ranking of multivariate GARCH models by problem dimension. Computational Statistics & Data Analysis, 76(??):172– 185, August 2014. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947312002113.

Chan:2014:MBC

[1489] Joshua C. C. Chan and Gary Koop.

Modelling breaks and clusters in the steady states of macroeconomic variables. Computational Statistics & Data Analysis, 76(??):186– 193, August 2014. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313001801.

Chen:2014:BES

[1490] Cathy W. S. Chen, Richard Gerlach, and Edward M. H. Lin. Bayesian estimation of smoothly mixing time-varying parameter GARCH models. Computational Statistics & Data Analysis, 76(??):194– 209, August 2014. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313003435.

Chretien:2014:MGE

[1491] Stéphane Chrétien and Juan-Pablo Ortega. Multivariate GARCH estimation via a Bregman-proximal trustregion method. Computational Statistics & Data Analysis, 76(??):210– 236, August 2014. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794731200388X.

Diaz-Emparanza:2014:NDF

[1492] Ignacio Diaz-Emparanza. Numerical distribution functions for seasonal unit root tests. Computational Statistics & Data Analysis, 76(??): 237–247, August 2014. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313001047.

Du:2014:TSI

[1493] Zaichao Du. Testing for serial independence of panel errors. Computational Statistics & Data Analysis, 76 (??):248–261, August 2014. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313002764.

Galeano:2014:MBD

[1494] Pedro Galeano and Dominik Wied.

Multiple break detection in the correlation structure of random variables. Computational Statistics & Data Analysis, 76(??):262–282, August 2014. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S016794731300087X.

Gallegati:2014:IRS

[1495] Marco Gallegati, James B. Ramsey, and Willi Semmler. Interest rate spreads and output: a time scale decomposition analysis using wavelets. Computational Statistics & Data Analysis, 76(??):283– 290, August 2014. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947314000644.

Ghoudi:2014:CST

[1496] Kilani Ghoudi and Bruno Rémillard.

Comparison of specification tests for GARCH models. Computational Statistics & Data Analysis, 76(??):

291–300, August 2014. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313001072.

Grassi:2014:WLM

[1497] Stefano Grassi and Paolo Santucci de Magistris. When long memory meets the Kalman filter: a comparative study. Computational Statistics & Data Analysis, 76(??):301– 319, August 2014. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947312003866.

Harvey:2014:EMF

[1498] Andrew Harvey and Genaro Sucarrat. EGARCH models with fat tails, skewness and leverage. Computational Statistics & Data Analysis, 76 (??):320–338, August 2014. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313003460.

Hwang:2014:IOL

[1499] Eunju Hwang and Dong Wan Shin. Infinite-order, long-memory heterogeneous autoregressive models. Computational Statistics & Data Analysis, 76 (??):339–358, August 2014. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794731300306X.

Jaschke:2014:ERM

[1500] Stefan Jäschke. Estimation of risk measures in energy portfolios using modern copula techniques. Computational Statistics & Data Analysis, 76 (??):359–376, August 2014. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947314000309.

Orsal:2014:PCT

[1501] Deniz Dilan Karaman Örsal and Bernd Droge. Panel cointegration testing in the presence of a time trend. Computational Statistics & Data Analysis, 76 (??):377–390, August 2014. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947312002162.

Karavias:2014:TUR

[1502] Yiannis Karavias and Elias Tzavalis.

Testing for unit roots in short panels allowing for a structural break. Computational Statistics & Data Analysis, 76(??):391–407, August 2014. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947312003829.

Kastner:2014:ASI

[1503] Gregor Kastner and Sylvia Frühwirth-Schnatter. Ancillarity-sufficiency interweaving strategy (ASIS) for boosting MCMC estimation of stochastic volatility models. Computational Statistics & Data Analysis, 76(??): 408–423, August 2014. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313000030.

Kiviet:2014:IVE

[1504] Jan F. Kiviet and Garry D. A. Phillips. Improved variance estimation of maximum likelihood estimators in stable first-order dynamic regression models. Computational Statistics & Data Analysis, 76(??):424– 448, August 2014. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947313003459.

Kleppe:2014:EIS

[1505] Tore Selland Kleppe and Roman Liesenfeld. Efficient importance sampling in mixture frameworks. Computational Statistics & Data Analysis, 76 (??):449–463, August 2014. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947313000406.

Kotchoni:2014:ICG

[1506] Rachidi Kotchoni. The indirect continuous-GMM estimation. Computational Statistics & Data Analysis, 76 (??):464–488, August 2014. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947313003472.

Larsson:2014:LRT

[1507] Rolf Larsson. A likelihood ratio type test for invertibility in moving average processes. Computational Statistics & Data Analysis, 76(??):489– 501, August 2014. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947314000656.

Martins:2014:TPC

[1508] Luis F. Martins and Paulo M. M.

Rodrigues. Testing for persistence change in fractionally integrated models: an application to world inflation rates. Computational Statistics & Data Analysis, 76(??):502– 522, August 2014. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947312002952.

Min:2014:SMB

[1509] Aleksey Min and Claudia Czado.

SCOMDY models based on pair-copula constructions with application to exchange rates. Computational Statistics & Data Analysis, 76(??):523– 535, August 2014. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947312003064.

Nyberg:2014:FNV

[1510] Henri Nyberg and Pentti Saikkonen. Forecasting with a noncausal VAR model. Computational Statistics & Data Analysis, 76(??):536– 555, August 2014. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313003678.

Aielli:2014:VCI

[1511] Gian Piero Aielli and Massimiliano Caporin. Variance clustering improved dynamic conditional correlation MGARCH estimators. Computational Statistics & Data Analysis, 76 (??):556–576, August 2014. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313000443.

Pitarakis:2014:JTS

[1512] Jean-Yves Pitarakis. A joint test for structural stability and a unit root in autoregressions. Computational Statistics & Data Analysis, 76 (??):577–587, August 2014. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947312003015.

Rombouts:2014:BOP

[1513] Jeroen V. K. Rombouts and Lars Stentoft. Bayesian option pricing using mixed normal heteroskedasticity models. Computational Statistics & Data Analysis, 76(??):588– 605, August 2014. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313002454.

Santos:2014:DFM

[1514] André A. P. Santos and Guilherme V. Moura. Dynamic factor multivariate GARCH model. Computational Statistics & Data Analysis, 76(??): 606–617, August 2014. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947312003398.

Shirota:2014:RSV

[1515] Shinichiro Shirota, Takayuki Hizu, and Yasuhiro Omori. Realized stochastic volatility with leverage and long memory. Computational Statistics & Data Analysis, 76(??):618– 641, August 2014. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313003101.

Skaug:2014:FAL

[1516] Hans J. Skaug and Jun Yu. A flexible and automated likelihood based framework for inference in stochastic volatility models. Computational Statistics & Data Analysis, 76(??): 642–654, August 2014. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313003587.

So:2014:VCG

[1517] Mike K. P. So and Cherry Y. T. Yeung. Vine-copula GARCH model with dynamic conditional dependence. Computational Statistics & Data Analysis, 76(??):655–671, August 2014. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313002958.

Stober:2014:RSD

[1518] Jakob Stöber and Claudia Czado.

Regime switches in the dependence structure of multidimensional financial data. Computational Statistics & Data Analysis, 76(??):672– 686, August 2014. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947313001321.

Venter:2014:ESV

[1519] J. H. Venter and P. J. de Jongh. Extended stochastic volatility models incorporating realised measures. Computational Statistics & Data Analysis, 76 (??):687–707, August 2014. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947312003969.

Wang:2014:ODF

[1520] Fangfang Wang. Optimal design of Fourier estimator in the presence of microstructure noise. Computational Statistics & Data Analysis, 76(??): 708–722, August 2014. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947313002892.

Wied:2014:FTC

[1521] Dominik Wied, Herold Dehling, Maarten van Kampen, and Daniel Vogel. A fluctuation test for constant Spearman’s rho with nuisancefree limit distribution. Computational Statistics & Data Analysis, 76 (??):723–736, August 2014. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947313001035.

Yen:2014:SNC

[1522] Yu-Min Yen and Tso-Jung Yen. Solving norm constrained portfolio optimization via coordinate-wise descent algorithms. Computational Statistics & Data Analysis, 76(??):737– 759, August 2014. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947313002557.

Anonymous:2014:EBi

[1523] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 77 (??):iii–v, September 2014. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947314001273.

Anonymous:2014:Ci

[1524] Anonymous. Contents. Computational Statistics & Data Analysis, 77 (??):vi–vii, September 2014. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947314001285.

Yin:2014:JIA

[1525] Jingjing Yin and Lili Tian. Joint inference about sensitivity and specificity at the optimal cut-off point associated with Youden index. Computational Statistics & Data Analysis, 77 (??):1–13, September 2014. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947314000450.

Liu:2014:EST

[1526] Fang Liu and Qing Li. Exact sequential test of equivalence hypothesis based on bivariate non-central t-statistics. Computational Statistics & Data Analysis, 77(??):14– 24, September 2014. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947314000474.

Shen:2014:FAM

[1527] Fumin Shen, Chunhua Shen, Rhys Hill, Anton van den Hengel, and Zhenmin Tang. Fast approximate L_{∞} minimization: Speeding up robust regression. Computational Statistics & Data Analysis, 77(??):25– 37, September 2014. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947314000589.

Brusco:2014:CSA

[1528] Michael J. Brusco. A comparison of simulated annealing algorithms for variable selection in principal component analysis and discriminant analysis. Computational Statistics & Data Analysis, 77(??): 38–53, September 2014. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947314000668.

Perrakis:2014:UMP

[1529] Konstantinos Perrakis, Ioannis Ntzoufras, and Efthymios G. Tsionas. On the use of marginal posteriors in marginal likelihood estimation via importance sampling. Computational Statistics & Data Analysis, 77(??): 54–69, September 2014. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947314000814.

Nooraee:2014:GLO

[1530] Nazanin Nooraee, Geert Molenberghs, and Edwin R. van den Heuvel. GEE for longitudinal ordinal data: Comparing R-geepack, R-multgee, R-repolr, SAS-GENMOD, SPSS-GENLIN. Computational Statistics & Data Analysis, 77 (??):70–83, September 2014. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947314000863.

Cozzini:2014:BML

[1531] Alberto Cozzini, Ajay Jasra, Giovanni Montana, and Adam Persing. A Bayesian mixture of lasso regressions with t-errors. Computational Statistics & Data Analysis, 77(??): 84–97, September 2014. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947314000954.

Kundu:2014:MDP

[1532] Debasis Kundu, Manuel Franco, and Juana-Maria Vivo. Multivariate distributions with proportional reversed hazard marginals. Computational Statistics & Data Analysis, 77(??): 98–112, September 2014. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947314000358.

Tang:2014:SBJ

[1533] Nian-Sheng Tang, An-Min Tang, and Dong-Dong Pan. Semiparametric Bayesian joint models of multivariate longitudinal and survival data. Computational Statistics & Data Analysis, 77(??):113– 129, September 2014. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947314000553.

Teran:2014:SCR

[1534] Pedro Terán and Miguel López-Díaz. Strong consistency and rates of convergence for a random estimator of a fuzzy set. Computational Statistics & Data Analysis, 77(??):130– 145, September 2014. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947314000565.

Zhang:2014:VAL

[1535] Q. Zhang and E. H. Ip. Variable assessment in latent class models. Computational Statistics & Data Analysis, 77(??):146–156, September 2014. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947314000577.

Qiu:2014:SSD

[1536] Shi-Fang Qiu, G. Y. Zou, and Man-Lai Tang. Sample size determination for estimating prevalence and a difference between two prevalences of sensitive attributes using the non-randomized triangular design. Computational Statistics & Data Analysis, 77(??):157– 169, September 2014. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947314000590.

Wu:2014:PLM

[1537] Chaojiang Wu and Yan Yu. Partially linear modeling of conditional quantiles using penalized splines. Computational Statistics & Data Analysis, 77(??): 170–187, September 2014. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947314000607.

Ozturk:2014:SIP

[1538] Omer Ozturk. Statistical inference for population quantiles and variance in judgment post-stratified samples. Computational Statistics & Data Analysis, 77(??):188– 205, September 2014. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947314000619.

Doyen:2014:SPE

[1539] L. Doyen. Semi-parametric estimation of Brown–Proschan preventive maintenance effects and intrinsic wear-out. Computational Statistics & Data Analysis, 77(??):206– 222, September 2014. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947314000620.

Das:2014:SAF

[1540] Ujjwal Das, Sudhir Gupta, and Shuva Gupta. Screening active factors in supersaturated designs. Computational Statistics & Data Analysis, 77(??): 223–232, September 2014. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947314000632.

Brechmann:2014:PPC

[1541] Eike C. Brechmann and Harry Joe. Parsimonious parameterization of correlation matrices using truncated vines and factor analysis. Computational Statistics & Data Analysis, 77(??): 233–251, September 2014. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947314000796.

Wang:2014:DOE

[1542] Changzhang Wang, You Zhou, Qiang Zhao, and Zhi Geng. Discovering and orienting the edges connected to a target variable in a DAG via a sequential local learning approach. Computational Statistics & Data Analysis, 77(??): 252–266, September 2014. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947314000802.

Schweer:2014:CPI

[1543] Sebastian Schweer and Christian H. Weiß. Compound Poisson INAR(1) processes: Stochastic properties and testing for overdispersion. Computational Statistics & Data Analysis, 77(??):267–284, September 2014.

CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0167947314000826.

Coudret:2014:NSI

[1544] R. Coudret, S. Girard, and J. Saracco.

A new sliced inverse regression method for multivariate response. Computational Statistics & Data Analysis, 77(??):285–299, September 2014. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0167947314000838.

Gigliarano:2014:MCP

[1545] Chiara Gigliarano, Silvia Figini, and Pietro Muliere. Making classifier performance comparisons when ROC curves intersect. Computational Statistics & Data Analysis, 77(??):300– 312, September 2014. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947314000851.

Choi:2014:NAM

[1546] Sungwoo Choi and Junyong Park.

Nonparametric additive model with grouped lasso and maximizing area under the ROC curve. Computational Statistics & Data Analysis, 77(??): 313–325, September 2014. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947314000875.

Murray:2014:MSF

[1547] Paula M. Murray, Ryan P. Browne, and Paul D. McNicholas. Mixtures of skew-t factor analyzers. Computational Statistics & Data Analysis, 77(??): 326–335, September 2014. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947314000899.

Bergersen:2014:MSL

[1548] Linn Cecilie Bergersen, Kukatharmini Tharmaratnam, and Ingrid K. Glad. Monotone splines lasso. Computational Statistics & Data Analysis, 77(??): 336–351, September 2014. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947314000905.

deLima:2014:ALR

[1549] Max Sousa de Lima and Luiz Henrique Duczmal. Adaptive likelihood ratio approaches for the detection of space– time disease clusters. Computational Statistics & Data Analysis, 77(??): 352–370, September 2014. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947314000929.

Huang:2014:ATS

[1550] Xiaolin Huang, Lei Shi, Kristiaan Pelckmans, and Johan A. K. Suykens. Asymmetric ν-tube support vector regression. Computational Statistics & Data Analysis, 77(??):371– 382, September 2014. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947314000930.

Anonymous:2014:EBj

[1551] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 78(??):iii–v, October 2014. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947314001595.

Anonymous:2014:Cj

[1552] Anonymous. Contents. Computational Statistics & Data Analysis, 78 (??):vi–vii, October 2014. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947314001601.

Renfro:2014:CWT

[1553] Lindsay A. Renfro, Qian Shi, Yuan Xue, Junlong Li, Hongwei Shang, and Daniel J. Sargent. Center-within-trial versus trial-level evaluation of surrogate endpoints. Computational Statistics & Data Analysis, 78(??): 1–20, October 2014. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947314000887.

Gouet:2014:SIG

[1554] Raúl Gouet, F. Javier López, Lina P.

Maldonado, and Gerardo Sanz. Statistical inference for the geometric distribution based on δ-records. Computational Statistics & Data Analysis, 78(??):21–32, October 2014. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794731400098X.

Lourenco:2014:MRF

[1555] V. M. Lourenço and A. M. Pires. Mregression, false discovery rates and outlier detection with application to genetic association studies. Computational Statistics & Data Analysis, 78 (??):33–42, October 2014. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947314000966.

Xu:2014:TPT

[1556] Lin Xu, Baisen Liu, Shurong Zheng, and Shaokun Bao. Testing proportionality of two large-dimensional covariance matrices. Computational Statistics & Data Analysis, 78(??): 43–55, October 2014. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947314000917.

Gutierrez:2014:BNC

[1557] Luis Gutiérrez, Eduardo Gutiérrez-Peña, and Ramsés H. Mena. Bayesian nonparametric classification for spectroscopy data. Computational Statistics & Data Analysis, 78(??):56– 68, October 2014. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947314001170.

Coolen-Maturi:2014:TGR

[1558] Tahani Coolen-Maturi, Faiza F. Elkhafifi, and Frank P. A. Coolen. Three-group ROC analysis: a nonparametric predictive approach. Computational Statistics & Data Analysis, 78 (??):69–81, October 2014. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947314001017.

Trevezas:2014:PES

[1559] S. Trevezas, S. Malefaki, and P.- H. Cournêde. Parameter estimation via stochastic variants of the ECM algorithm with applications to plant growth modeling. Computational Statistics & Data Analysis, 78 (??):82–99, October 2014. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947314001005.

Manisera:2014:MRD

[1560] Marica Manisera and Paola Zuccolotto.

Modeling rating data with nonlinear CUB models. Computational Statistics & Data Analysis, 78(??):100– 118, October 2014. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947314000978.

Wang:2014:ECL

[1561] Xianlong Wang and Annie Qu. Efficient classification for longitudinal data. Computational Statistics & Data Analysis, 78(??):119– 134, October 2014. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947314001157.

Yamamoto:2014:MBS

[1562] Kouji Yamamoto and Hidetoshi Murakami. Model based on skew normal distribution for square contingency tables with ordinal categories. Computational Statistics & Data Analysis, 78 (??):135–140, October 2014. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947314001145.

Johnson:2014:MCM

[1563] Alicia A. Johnson and James M. Flegal. A modified conditional Metropolis– Hastings sampler. Computational Statistics & Data Analysis, 78(??): 141–152, October 2014. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947314001169.

Smirnov:2014:FHE

[1564] Pavel O. Smirnov and Georgy L. Shevlyakov. Fast highly efficient and robust one-step M-estimators of scale based on Q_{n}. Computational Statistics & Data Analysis, 78(??):153– 158, October 2014. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947314001200.

Berg:2014:SAP

[1565] Emily Berg and Hukum Chandra. Small area prediction for a unitlevel lognormal model. Computational Statistics & Data Analysis, 78 (??):159–175, October 2014. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794731400084X.

Matsui:2014:VBS

[1566] Hidetoshi Matsui. Variable and boundary selection for functional data via multiclass logistic regression modeling. Computational Statistics & Data Analysis, 78(??):176– 185, October 2014. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947314001224.

Feng:2014:TBE

[1567] Zhenghui Feng, Tao Wang, and Lixing Zhu. Transformation-based estimation. Computational Statistics & Data Analysis, 78(??):186– 205, October 2014. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794731400139X.

Wang:2014:SMP

[1568] Kaibo Wang, Arthur B. Yeh, and Bo Li. Simultaneous monitoring of process mean vector and covariance matrix via penalized likelihood estimation. Computational Statistics & Data Analysis, 78(??):206– 217, October 2014. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947314001388.

Zhang:2014:SAB

[1569] Xibin Zhang, Maxwell L. King, and Han Lin Shang. A sampling algorithm for bandwidth estimation in a nonparametric regression model with a flexible error density. Computational Statistics & Data Analysis, 78 (??):218–234, October 2014. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947314001376.

Harvey:2014:IDF

[1570] David I. Harvey, Stephen J. Leybourne, and A. M. Robert Taylor. On infimum Dickey–Fuller unit root tests allowing for a trend break under the null. Computational Statistics & Data Analysis, 78(??):235– 242, October 2014. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947312003854.

Anonymous:2014:EBk

[1571] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 79(??):iii–v, November 2014. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947314001959.

Anonymous:2014:Ck

[1572] Anonymous. Contents. Computational Statistics & Data Analysis, 79 (??):vi–vii, November 2014. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947314001960.

Mares:2014:DPU

[1573] Jordi Marés and Natalie Shlomo.

Data privacy using an evolutionary algorithm for invariant PRAM matrices. Computational Statistics & Data Analysis, 79(??):1– 13, November 2014. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947314001406.

Kao:2014:EDA

[1574] Chiun-How Kao, Junji Nakano, Sheau-Hue Shieh, Yin-Jing Tien, Han-Ming Wu, Chuan kai Yang, and Chun houh Chen. Exploratory data analysis of interval-valued symbolic data with matrix visualization. Computational Statistics & Data Analysis, 79 (??):14–29, November 2014. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947314001194.

Martinez-Camblor:2014:CVD

[1575] Pablo Martínez-Camblor. On correlated z-values distribution in hypothesis testing. Computational Statistics & Data Analysis, 79(??):30– 43, November 2014. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947314001467.

Zhao:2014:RMS

[1576] Weihua Zhao, Riquan Zhang, and Jicai Liu. Regularization and model selection for quantile varying coefficient model with categorical effect modifiers. Computational Statistics & Data Analysis, 79(??):44– 62, November 2014. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947314001431.

Ndao:2014:NEC

[1577] Pathé Ndao, Aliou Diop, and Jean-François Dupuy. Nonparametric estimation of the conditional tail index and extreme quantiles under random censoring. Computational Statistics & Data Analysis, 79(??):63– 79, November 2014. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947314001479.

Lunardon:2014:CLI

[1578] Nicola Lunardon and Elvezio Ronchetti. Composite likelihood inference by nonparametric saddlepoint tests. Computational Statistics & Data Analysis, 79 (??):80–90, November 2014. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947314001546.

Niu:2014:ELI

[1579] Cuizhen Niu, Xu Guo, Wangli Xu, and Lixing Zhu. Empirical likelihood inference in linear regression with nonignorable missing response. Computational Statistics & Data Analysis, 79 (??):91–112, November 2014. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947314001455.

Gu:2014:PET

[1580] Wei Gu, Hulin Wu, Hongyu Miao, and Hongqi Xue. Parameter estimation for a type of nonlinear stochastic models observed with error. Computational Statistics & Data Analysis, 79 (??):113–119, November 2014. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947314001534.

Hirose:2014:EOS

[1581] Kei Hirose and Michio Yamamoto.

Estimation of an oblique structure via penalized likelihood factor analysis. Computational Statistics & Data Analysis, 79(??):120– 132, November 2014. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947314001510.

Yamamoto:2014:FFM

[1582] Michio Yamamoto and Yoshikazu Terada. Functional factorial Kmeans analysis. Computational Statistics & Data Analysis, 79(??):133– 148, November 2014. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947314001509.

Bourel:2014:RAS

[1583] M. Bourel, R. Fraiman, and B. Ghattas. Random average shifted histograms. Computational Statistics & Data Analysis, 79(??):149– 164, November 2014. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947314001443.

Lam:2014:SAC

[1584] K. F. Lam and Kin-Yau Wong. Semiparametric analysis of clustered interval-censored survival data with a cure fraction. Computational Statistics & Data Analysis, 79(??):165– 174, November 2014. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947314001698.

Leiva:2014:FAC

[1585] Víctor Leiva, Helton Saulo, Jeremias Leão, and Carolina Marchant. A family of autoregressive conditional duration models applied to financial data. Computational Statistics & Data Analysis, 79(??):175– 191, November 2014. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947314001662.

Gumedze:2014:DOL

[1586] Freedom N. Gumedze and Tinashe D. Chatora. Detection of outliers in longitudinal count data via overdispersion. Computational Statistics & Data Analysis, 79(??):192–202, November 2014. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947314001686.

Li:2014:LAD

[1587] Xinxin Li, Lili Mo, Xiaoming Yuan, and Jianzhong Zhang. Linearized alternating direction method of multipliers for sparse group and fused LASSO models. Computational Statistics & Data Analysis, 79(??):203– 221, November 2014. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947314001674.

OBrien:2014:RCC

[1588] Travis A. O’Brien, William D. Collins, Sara A. Rauscher, and Todd D. Ringler. Reducing the computational cost of the ECF using a nuFFT: a fast and objective probability density estimation method. Computational Statistics & Data Analysis, 79 (??):222–234, November 2014. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794731400173X.

Xiao:2014:RWA

[1589] L. Q. Xiao, B. Hou, Z. F. Wang, and Y. H. Wu. Random weighting approximation for Tobit regression models with longitudinal data. Computational Statistics & Data Analysis, 79 (??):235–247, November 2014. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947314001704.

Costa:2014:EMM

[1590] D. R. Costa, V. H. Lachos, J. L. Bazan, and C. L. N. Azevedo. Estimation methods for multivariate Tobit confirmatory factor analysis. Computational Statistics & Data Analysis, 79 (??):248–260, November 2014. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947314001716.

Wu:2014:NRE

[1591] Huiping Wu, Ka-Veng Yuen, and Shing-On Leung. A novel relative entropy–posterior predictive model checking approach with limited information statistics for latent trait models in sparse 2^{k} contingency tables. Computational Statistics & Data Analysis, 79(??):261– 276, November 2014. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947314001753.

Zhang:2014:CIT

[1592] Junni L. Zhang. Comparative investigation of three Bayesian p values. Computational Statistics & Data Analysis, 79(??):277–291, November 2014. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947314001522.

Anonymous:2014:EBl

[1593] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 80(??):iii–v, December 2014. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947314002230.

Anonymous:2014:Cl

[1594] Anonymous. Contents. Computational Statistics & Data Analysis, 80 (??):vi–vii, December 2014. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947314002242.

Vilca:2014:BSE

[1595] Filidor Vilca, N. Balakrishnan, and Camila Borelli Zeller. The bivariate Sinh–Elliptical distribution with applications to Birnbaum–Saunders distribution and associated regression and measurement error models. Computational Statistics & Data Analysis, 80 (??):1–16, December 2014. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947314001728.

Horrocks:2014:BAE

[1596] Julie Horrocks and Matthew Rueffer. A Bayesian approach to estimating animal density from binary acoustic transects. Computational Statistics & Data Analysis, 80(??):17– 25, December 2014. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947314001765.

Cerny:2014:CCA

[1597] Michal Çerný and Milan Hladík. The complexity of computation and approximation of the t-ratio over onedimensional interval data. Computational Statistics & Data Analysis, 80 (??):26–43, December 2014. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947314001789.

Kang:2014:MDP

[1598] Jiwon Kang and Sangyeol Lee. Minimum density power divergence estimator for Poisson autoregressive models. Computational Statistics & Data Analysis, 80(??):44– 56, December 2014. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947314001807.

Nguyen:2014:EFI

[1599] Thu Thuy Nguyen and France Mentré. Evaluation of the Fisher information matrix in nonlinear mixed effect models using adaptive Gaussian quadrature. Computational Statistics & Data Analysis, 80(??):57– 69, December 2014. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947314001820.

Xu:2014:ICC

[1600] Jianwen Xu and You-Gan Wang. Intracluster correlation structure in longitudinal data analysis: Selection criteria and misspecification tests. Computational Statistics & Data Analysis, 80 (??):70–77, December 2014. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947314001844.

Wei:2014:MTM

[1601] Shaoceng Wei, Liou Xu, and Richard J.

Kryscio. Markov transition model to dementia with death as a competing event. Computational Statistics & Data Analysis, 80(??):78– 88, December 2014. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947314001856.

Lui:2014:NTE

[1602] Kung-Jong Lui, William G. Cumberland, and Kuang-Chao Chang. Notes on testing equality in binary data under a three period crossover design. Computational Statistics & Data Analysis, 80(??):89–98, December 2014. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947314001868.

Lau:2014:RSU

[1603] F. Din-Houn Lau and Axel Gandy.

RMCMC: a system for updating Bayesian models. Computational Statistics & Data Analysis, 80(??): 99–110, December 2014. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947314001819.

Lee:2014:BCF

[1604] Keunbaik Lee and Jae Keun Yoo.

Bayesian Cholesky factor models in random effects covariance matrix for generalized linear mixed models. Computational Statistics & Data Analysis, 80(??):111–116, December 2014. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016794731400187X.

Chi:2014:SEC

[1605] Eric C. Chi and Kenneth Lange. Stable estimation of a covariance matrix guided by nuclear norm penalties. Computational Statistics & Data Analysis, 80(??):117– 128, December 2014. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947314001893.

Hapfelmeier:2014:VSR

[1606] A. Hapfelmeier and K. Ulm. Variable selection by random forests using data with missing values. Computational Statistics & Data Analysis, 80 (??):129–139, December 2014. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947314001881.

Wang:2014:RES

[1607] Xiaoguang Wang and Xinyong Shi. Robust estimation for survival partially linear single-index models. Computational Statistics & Data Analysis, 80 (??):140–152, December 2014. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947314001996.

Auffray:2014:BRE

[1608] Yves Auffray, Pierre Barbillon, and Jean-Michel Marin. Bounding rare event probabilities in computer experiments. Computational Statistics & Data Analysis, 80(??):153– 166, December 2014. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947314002023.

Kobayashi:2014:TAB

[1609] Genya Kobayashi. A transdimensional approximate Bayesian computation using the pseudo-marginal approach for model choice. Computational Statistics & Data Analysis, 80 (??):167–183, December 2014. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947314002047.

Donoghoe:2014:SCM

[1610] Mark W. Donoghoe and Ian C.

Marschner. Stable computational methods for additive binomial models with application to adjusted risk differences. Computational Statistics & Data Analysis, 80(??):184– 196, December 2014. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794731400190X.

Lian:2014:PLS

[1611] Heng Lian, Pang Du, YuanZhang Li, and Hua Liang. Partially linear structure identification in generalized additive models with NPdimensionality. Computational Statistics & Data Analysis, 80(??):197– 208, December 2014. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016794731400200X.

Rohart:2014:SFE

[1612] Florian Rohart, Magali San Cristobal, and Béatrice Laurent. Selection of fixed effects in high dimensional linear mixed models using a multicycle ECM algorithm. Computational Statistics & Data Analysis, 80(??): 209–222, December 2014. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947314002011.

Zhao:2014:VIP

[1613] Kaifeng Zhao and Heng Lian. Variational inferences for partially linear additive models with variable selection. Computational Statistics & Data Analysis, 80(??):223– 239, December 2014. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947314002072.

Xiang:2014:EMT

[1614] Fei Xiang and Peter Neal. Efficient MCMC for temporal epidemics via parameter reduction. Computational Statistics & Data Analysis, 80(??): 240–250, December 2014. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947314002060.

Anonymous:2015:EBa

[1615] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 81(??):iii–vi, January 2015. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947314002539.

Anonymous:2015:Ca

[1616] Anonymous. Contents. Computational Statistics & Data Analysis, 81 (??):vii, January 2015. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947314002540.

Han:2015:SCB

[1617] Yang Han, Wei Liu, Frank Bretz, and Fang Wan. Simultaneous confidence bands for a percentile line in linear regression. Computational Statistics & Data Analysis, 81(??): 1–9, January 2015. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947314002138.

Dijkstra:2015:CAN

[1618] Theo K. Dijkstra and Jörg Henseler.

Consistent and asymptotically normal PLS estimators for linear structural equations. Computational Statistics & Data Analysis, 81(??):10– 23, January 2015. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947314002126.

Raillard:2015:MPA

[1619] Nicolas Raillard, Marc Prevosto, and Pierre Ailliot. Modeling process asymmetries with Laplace moving average. Computational Statistics & Data Analysis, 81(??):24– 37, January 2015. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794731400214X.

Fuchs:2015:PSF

[1620] Karen Fuchs, Fabian Scheipl, and Sonja Greven. Penalized scalar-on-functions regression with interaction term. Computational Statistics & Data Analysis, 81(??):38–51, January 2015. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947314002059.

Cabras:2015:NMT

[1621] Stefano Cabras, Maria Eugenia Castellanos, and Silvia Perra. A new minimal training sample scheme for intrinsic Bayes factors in censored data. Computational Statistics & Data Analysis, 81(??):52–63, January 2015. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947314002163.

Hubert:2015:DDE

[1622] Mia Hubert, Peter Rousseeuw, Dina Vanpaemel, and Tim Verdonck. The DetS and DetMM estimators for multivariate location and scatter. Computational Statistics & Data Analysis, 81(??):64–75, January 2015. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947314002175.

Ghosh:2015:SAB

[1623] Joyee Ghosh and Aixin Tan. Sandwich algorithms for Bayesian variable selection. Computational Statistics & Data Analysis, 81(??):76– 88, January 2015. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947314002187.

Martinez:2015:CVM

[1624] Diana L. Martinez, Dachuan T. Shih, Victoria C. P. Chen, and Seoung Bum Kim. A convex version of multivariate adaptive regression splines. Computational Statistics & Data Analysis, 81 (??):89–106, January 2015. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947314002291.

Adragni:2015:STV

[1625] Kofi P. Adragni and Moumita Karmakar. A sequential test for variable selection in high dimensional complex data. Computational Statistics & Data Analysis, 81(??):107– 120, January 2015. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947314002308.

Sen:2015:MBB

[1626] Bodhisattva Sen and Gongjun Xu.

Model based bootstrap methods for interval censored data. Computational Statistics & Data Analysis, 81 (??):121–129, January 2015. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947314002114.

Lui:2015:TEB

[1627] Kung-Jong Lui and Kuang-Chao Chang. Test and estimation in binary data analysis under an incomplete block crossover design. Computational Statistics & Data Analysis, 81 (??):130–138, January 2015. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016794731400231X.

Anonymous:2015:EBb

[1628] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 82(??):iii–vi, February 2015. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947314002874.

Anonymous:2015:Cb

[1629] Anonymous. Contents. Computational Statistics & Data Analysis, 82 (??):vii–viii, February 2015. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947314002886.

Yang:2015:WKS

[1630] Zhao Yang and Ming Zhou. Weighted kappa statistic for clustered matchedpair ordinal data. Computational Statistics & Data Analysis, 82(??): 1–18, February 2015. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947314002369.

Feizjavadian:2015:ADC

[1631] S. H. Feizjavadian and R. Hashemi.

Analysis of dependent competing risks in the presence of progressive hybrid censoring using Marshall–Olkin bivariate Weibull distribution. Computational Statistics & Data Analysis, 82 (??):19–34, February 2015. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947314002345.

OMalley:2015:URS

[1632] A. James O’Malley and Sudeshna Paul. Using retrospective sampling to estimate models of relationship status in large longitudinal social networks. Computational Statistics & Data Analysis, 82(??): 35–46, February 2015. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947314002333.

Tsai:2015:CCC

[1633] Miao-Yu Tsai. Comparison of concordance correlation coefficient via variance components, generalized estimating equations and weighted approaches with model selection. Computational Statistics & Data Analysis, 82(??):47–58, February 2015. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947314002370.

Wu:2015:DGT

[1634] K. Y. K. Wu and W. K. Li. Double generalized threshold models with constraint on the dispersion by the mean. Computational Statistics & Data Analysis, 82(??): 59–73, February 2015. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947314002357.

Lv:2015:ERV

[1635] Jing Lv, Hu Yang, and Chaohui Guo. An efficient and robust variable selection method for longitudinal generalized linear models. Computational Statistics & Data Analysis, 82 (??):74–88, February 2015. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947314002382.

Pulkkinen:2015:RBM

[1636] Seppo Pulkkinen. Ridge-based method for finding curvilinear structures from noisy data. Computational Statistics & Data Analysis, 82(??):89– 109, February 2015. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947314002394.

Park:2015:SMH

[1637] Junyong Park and DoHwan Park. Stein’s method in high dimensional classification and applications. Computational Statistics & Data Analysis, 82 (??):110–125, February 2015. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947314002412.

Pinho:2015:CDG

[1638] Luis Gustavo B. Pinho, Juvêncio S.

Nobre, and Julio M. Singer. Cook’s distance for generalized linear mixed models. Computational Statistics & Data Analysis, 82(??):126– 136, February 2015. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947314002400.

LaVecchia:2015:RHR

[1639] Davide La Vecchia, Lorenzo Camponovo, and Davide Ferrari. Robust heart rate variability analysis by generalized entropy minimization. Computational Statistics & Data Analysis, 82 (??):137–151, February 2015. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794731400259X.

Wolf:2015:OLR

[1640] Bethany J. Wolf, Elizabeth H. Slate, and Elizabeth G. Hill. Ordinal logic regression: a classifier for discovering combinations of binary markers for ordinal outcomes. Computational Statistics & Data Analysis, 82(??): 152–163, February 2015. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794731400245X.

Wu:2015:NAP

[1641] Tung-Lung Wu and Joseph Glaz. A new adaptive procedure for multiple window scan statistics. Computational Statistics & Data Analysis, 82 (??):164–172, February 2015. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016794731400262X.

Sullivan:2015:HDI

[1642] Danielle Sullivan and Rebecca Andridge. A hot deck imputation procedure for multiply imputing nonignorable missing data: the proxy pattern-mixture hot deck. Computational Statistics & Data Analysis, 82 (??):173–185, February 2015. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947314002692.

Chen:2015:ELR

[1643] Xiwei Chen, Albert Vexler, and Marianthi Markatou. Empirical likelihood ratio confidence interval estimation of best linear combinations of biomarkers. Computational Statistics & Data Analysis, 82(??):186– 198, February 2015. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947314002710.

Dalmau:2015:KMA

[1644] Oscar Dalmau, Teresa E. Alarcón, and Graciela González. Kernel multilogit algorithm for multiclass classification. Computational Statistics & Data Analysis, 82(??):199– 206, February 2015. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947314002679.

Tutz:2015:VSG

[1645] Gerhard Tutz, Wolfgang Pößnecker, and Lorenz Uhlmann. Variable selection in general multinomial logit models. Computational Statistics & Data Analysis, 82(??):207– 222, February 2015. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947314002709.

Gomes:2015:NPR

[1646] M. Ivette Gomes, M. Fátima Brilhante, Frederico Caeiro, and Dinis Pestana. A new partially reducedbias mean-of-order p class of extreme value index estimators. Computational Statistics & Data Analysis, 82 (??):223–237, February 2015. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947314002680.

Anonymous:2015:EBc

[1647] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 83(??):iii–vi, March 2015. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947314003260.

Anonymous:2015:Cc

[1648] Anonymous. Contents. Computational Statistics & Data Analysis, 83 (??):vii–viii, March 2015. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947314003272.

Chung:2015:AEP

[1649] Dongjun Chung and Hyunjoong Kim. Accurate ensemble pruning with PL-bagging. Computational Statistics & Data Analysis, 83(??): 1–13, March 2015. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947314002631.

Hwang:2015:SBS

[1650] Eunju Hwang and Dong Wan Shin. Stationary bootstrapping for semiparametric panel unit root tests. Computational Statistics & Data Analysis, 83(??):14–25, March 2015. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947314002643.

Wang:2015:IWD

[1651] Bing Xing Wang and Zhi-Sheng Ye. Inference on the Weibull distribution based on record values. Computational Statistics & Data Analysis, 83 (??):26–36, March 2015. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947314002655.

DeOliveira:2015:PII

[1652] Victor De Oliveira and Bazoumana Kone. Prediction intervals for integrals of Gaussian random fields. Computational Statistics & Data Analysis, 83(??):37–51, March 2015. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947314002746.

Zhang:2015:CAD

[1653] Jun Zhang, Gaorong Li, and Zhenghui Feng. Checking the adequacy for a distortion errors-in-variables parametric regression model. Computational Statistics & Data Analysis, 83 (??):52–64, March 2015. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947314002795.

Christiansen:2015:MFD

[1654] Marcus C. Christiansen, Andreas Niemeyer, and Lucia Teigiszerová. Modeling and forecasting durationdependent mortality rates. Computational Statistics & Data Analysis, 83(??):65–81, March 2015. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947314002783.

Diaz:2015:PMB

[1655] Mireya Diaz. Performance measures of the bivariate random effects model for meta-analyses of diagnostic accuracy. Computational Statistics & Data Analysis, 83(??): 82–90, March 2015. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947314002825.

Bhattacharya:2015:COR

[1656] Ritwik Bhattacharya, Biswabrata Pradhan, and Anup Dewanji. Computation of optimum reliability acceptance sampling plans in presence of hybrid censoring. Computational Statistics & Data Analysis, 83(??): 91–100, March 2015. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947314002941.

Mbalawata:2015:AMA

[1657] Isambi S. Mbalawata, Simo Särkkä, Matti Vihola, and Heikki Haario. Adaptive Metropolis algorithm using variational Bayesian adaptive Kalman filter. Computational Statistics & Data Analysis, 83(??):101– 115, March 2015. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947314002989.

delCastillo:2015:LIG

[1658] Joan del Castillo and Isabel Serra. Likelihood inference for generalized Pareto distribution. Computational Statistics & Data Analysis, 83(??): 116–128, March 2015. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947314003065.

Hanafi:2015:MWP

[1659] Mohamed Hanafi, Samia Samar Ouertani, Julien Boccard, Gérard Mazerolles, and Serge Rudaz. Multiway PLS regression: Monotony convergence of tri-linear PLS2 and optimality of parameters. Computational Statistics & Data Analysis, 83 (??):129–139, March 2015. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947314002953.

Wang:2015:RAB

[1660] Naichen Wang, Lianming Wang, and Christopher S. McMahan. Regression analysis of bivariate current status data under the gamma-frailty proportional hazards model using the EM algorithm. Computational Statistics & Data Analysis, 83(??):140– 150, March 2015. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947314003053.

Kang:2015:CTH

[1661] Fangyuan Kang, Liuquan Sun, and Xingqiu Zhao. A class of transformed hazards models for recurrent gap times. Computational Statistics & Data Analysis, 83(??):151– 167, March 2015. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947314002977.

White:2015:MCA

[1662] Staci A. White and Radu Herbei.

A Monte Carlo approach to quantifying model error in Bayesian parameter estimation. Computational Statistics & Data Analysis, 83(??): 168–181, March 2015. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947314003004.

Peavoy:2015:SPC

[1663] Daniel Peavoy, Christian L. E. Franzke, and Gareth O. Roberts. Systematic physics constrained parameter estimation of stochastic differential equations. Computational Statistics & Data Analysis, 83(??):182– 199, March 2015. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794731400303X.

Liu:2015:TMZ

[1664] Yin Liu and Guo-Liang Tian. Type I multivariate zero-inflated Poisson distribution with applications. Computational Statistics & Data Analysis, 83 (??):200–222, March 2015. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947314003028.

Wang:2015:MCF

[1665] Wan-Lun Wang. Mixtures of common t-factor analyzers for modeling high-dimensional data with missing values. Computational Statistics & Data Analysis, 83(??):223– 235, March 2015. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947314002990.

Kong:2015:DSV

[1666] Dehan Kong, Howard D. Bondell, and Yichao Wu. Domain selection for the varying coefficient model via local polynomial regression. Computational Statistics & Data Analysis, 83 (??):236–250, March 2015. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947314002965.

Touloumis:2015:NST

[1667] Anestis Touloumis. Nonparametric Stein-type shrinkage covariance matrix estimators in high-dimensional settings. Computational Statistics & Data Analysis, 83(??):251– 261, March 2015. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947314003107.

Maronna:2015:HFS

[1668] Ricardo A. Maronna and Victor J.

Yohai. High finite-sample efficiency and robustness based on distance-constrained maximum likelihood. Computational Statistics & Data Analysis, 83(??):262–274, March 2015. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0167947314003077.

Ma:2015:TPS

[1669] Yingying Ma, Wei Lan, and Hansheng Wang. Testing predictor significance with ultra high dimensional multivariate responses. Computational Statistics & Data Analysis, 83 (??):275–286, March 2015. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947314002813.

Bartolucci:2015:TSE

[1670] Francesco Bartolucci, Giorgio E. Montanari, and Silvia Pandolfi. Threestep estimation of latent Markov models with covariates. Computational Statistics & Data Analysis, 83(??): 287–301, March 2015. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947314003090.

Anonymous:2015:EBd

[1671] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 84(??):iii–vi, April 2015. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947314003508.

Anonymous:2015:Cd

[1672] Anonymous. Contents. Computational Statistics & Data Analysis, 84 (??):vii, April 2015. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016794731400351X.

Haff:2015:NEP

[1673] Ingrid Hobæk Haff and Johan Segers. Nonparametric estimation of paircopula constructions with the empirical pair-copula. Computational Statistics & Data Analysis, 84(??): 1–13, April 2015. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947314003120.

Bee:2015:AML

[1674] Marco Bee, Giuseppe Espa, and Diego Giuliani. Approximate maximum likelihood estimation of the autologistic model. Computational Statistics & Data Analysis, 84(??): 14–26, April 2015. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947314003119.

Klouda:2015:EPT

[1675] Karel Klouda. An exact polynomial time algorithm for computing the least trimmed squares estimate. Computational Statistics & Data Analysis, 84(??):27–40, April 2015. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947314003144.

Fu:2015:GPA

[1676] Liya Fu, You-Gan Wang, and Min Zhu. A Gaussian pseudolikelihood approach for quantile regression with repeated measurements. Computational Statistics & Data Analysis, 84 (??):41–53, April 2015. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947314003156.

Sun:2015:PSM

[1677] Libo Sun, Chihoon Lee, and Jennifer A. Hoeting. A penalized simulated maximum likelihood approach in parameter estimation for stochastic differential equations. Computational Statistics & Data Analysis, 84(??):54–67, April 2015. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016794731400320X.

Mante:2015:IBO

[1678] Claude Manté. Iterated Bernstein operators for distribution function and density estimation: Balancing between the number of iterations and the polynomial degree. Computational Statistics & Data Analysis, 84 (??):68–84, April 2015. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947314003168.

Liu:2015:PLA

[1679] Wei Liu, Bo Zhang, Zhiwei Zhang, Baojiang Chen, and Xiao-Hua Zhou. A pseudo-likelihood approach for estimating diagnostic accuracy of multiple binary medical tests. Computational Statistics & Data Analysis, 84(??):85–98, April 2015. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947314003193.

Jang:2015:SBA

[1680] Woosung Jang and Huixia Judy Wang. A semiparametric Bayesian approach for joint-quantile regression with clustered data. Computational Statistics & Data Analysis, 84(??): 99–115, April 2015. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947314003211.

Poon:2015:CIC

[1681] Wai-Yin Poon, Shi-Fang Qiu, and Man-Lai Tang. Confidence interval construction for the Youden index based on partially validated series. Computational Statistics & Data Analysis, 84(??):116–134, April 2015. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0167947314003363.

Ding:2015:NMA

[1682] Jieli Ding, Guo-Liang Tian, and Kam Chuen Yuen. A new MM algorithm for constrained estimation in the proportional hazards model. Computational Statistics & Data Analysis, 84(??):135–151, April 2015. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947314003181.

Anonymous:2015:EBe

[1683] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 85(??):iii–vi, May 2015. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947315000067.

Anonymous:2015:Ce

[1684] Anonymous. Contents. Computational Statistics & Data Analysis, 85 (??):vii, May 2015. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947315000079.

Gijbels:2015:RNG

[1685] I. Gijbels and I. Vrinssen. Robust nonnegative garrote variable selection in linear regression. Computational Statistics & Data Analysis, 85 (??):1–22, May 2015. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947314003326.

Tan:2015:CGL

[1686] Kean Ming Tan, Daniela Witten, and Ali Shojaie. The cluster graphical lasso for improved estimation of Gaussian graphical models. Computational Statistics & Data Analysis, 85(??):23–36, May 2015. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947314003387.

Bernhardt:2015:FEA

[1687] Paul W. Bernhardt, Daowen Zhang, and Huixia Judy Wang. A fast EM algorithm for fitting joint models of a binary response and multiple longitudinal covariates subject to detection limits. Computational Statistics & Data Analysis, 85(??): 37–53, May 2015. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794731400334X.

Kong:2015:GTI

[1688] Maiying Kong, Sheng Xu, Steven M.

Levy, and Somnath Datta. GEE type inference for clustered zero-inflated negative binomial regression with application to dental caries. Computational Statistics & Data Analysis, 85(??):54–66, May 2015. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947314003375.

Dobbin:2015:SSM

[1689] Kevin K. Dobbin and Alexei C. Ionan. Sample size methods for constructing confidence intervals for the intra-class correlation coefficient. Computational Statistics & Data Analysis, 85(??):67–83, May 2015. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947314003338.

Lee:2015:BTS

[1690] J. Lee, Y. Fan, and S. A. Sisson. Bayesian threshold selection for extremal models using measures of surprise. Computational Statistics & Data Analysis, 85(??): 84–99, May 2015. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947314003429.

Lee:2015:ETR

[1691] Sangyeol Lee and Haejune Oh. Entropy test and residual empirical process for autoregressive conditional duration models. Computational Statistics & Data Analysis, 86(??): 1–12, June 2015. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947314003545.

Hu:2015:DDF

[1692] Qinqin Hu, Peng Zeng, and Lu Lin. The dual and degrees of freedom of linearly constrained generalized lasso. Computational Statistics & Data Analysis, 86(??):13–26, June 2015. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947314003582.

Moores:2015:EFP

[1693] Matthew T. Moores, Catriona E. Hargrave, Timothy Deegan, Michael Poulsen, Fiona Harden, and Kerrie Mengersen. An external field prior for the hidden Potts model with application to cone-beam computed tomography. Computational Statistics & Data Analysis, 86(??): 27–41, June 2015. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947314003399.

Zhu:2015:MCS

[1694] Hong Zhu and Bo Lu. Multiple comparisons for survival data with propensity score adjustment. Computational Statistics & Data Analysis, 86(??):42–51, June 2015. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947315000110.

Bowman:2015:ACI

[1695] Adrian W. Bowman, Stanislav Katina, Joanna Smith, and Denise Brown. Anatomical curve identification. Computational Statistics & Data Analysis, 86(??):52–64, June 2015. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947314003557.

Burgin:2015:TBV

[1696] Reto Bürgin and Gilbert Ritschard.

Tree-based varying coefficient regression for longitudinal ordinal responses. Computational Statistics & Data Analysis, 86(??):65–80, June 2015. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947315000146.

Zhang:2015:BED

[1697] Rong Zhang, Brett A. Inder, and Xibin Zhang. Bayesian estimation of a discrete response model with double rules of sample selection. Computational Statistics & Data Analysis, 86(??):81–96, June 2015. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947315000134.

Chkrebtii:2015:TAB

[1698] Oksana A. Chkrebtii, Erin K. Cameron, David A. Campbell, and Erin M. Bayne. Transdimensional approximate Bayesian computation for inference on invasive species models with latent variables of unknown dimension. Computational Statistics & Data Analysis, 86(??): 97–110, June 2015. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947315000122.

Anonymous:2015:Cf

[1699] Anonymous. Contents. Computational Statistics & Data Analysis, 86 (??):ibc, June 2015. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947315000389.

Anonymous:2015:EBf

[1700] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 86(??):ibc, June 2015. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947315000377.

Yang:2015:KSC

[1701] Zhao Yang and Ming Zhou. Kappa statistic for clustered physician– patients polytomous data. Computational Statistics & Data Analysis, 87 (??):1–17, July 2015. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947315000183.

Barranco-Chamorro:2015:CDD

[1702] I. Barranco-Chamorro, M. D. Jiménez-Gamero, J. A. Mayor-Gallego, and J. L. Moreno-Rebollo. A case-deletion diagnostic for penalized calibration estimators and BLUP under linear mixed models in survey sampling. Computational Statistics & Data Analysis, 87(??):18–33, July 2015. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947315000158.

Chen:2015:RAM

[1703] Yurong Chen, Yanqin Feng, and Jianguo Sun. Regression analysis of multivariate current status data with auxiliary covariates under the additive hazards model. Computational Statistics & Data Analysis, 87(??): 34–45, July 2015. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794731500016X.

Niu:2015:NEE

[1704] Yi Niu and Yingwei Peng. A new estimating equation approach for marginal hazard ratio estimation. Computational Statistics & Data Analysis, 87 (??):46–56, July 2015. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947315000237.

Klimova:2015:FLH

[1705] Anna Klimova, Caroline Uhler, and Tamás Rudas. Faithfulness and learning hypergraphs from discrete distributions. Computational Statistics & Data Analysis, 87(??):57– 72, July 2015. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947315000286.

Gard:2015:BHM

[1706] Charlotte C. Gard and Elizabeth R. Brown. A Bayesian hierarchical model for estimating and partitioning Bernstein polynomial density functions. Computational Statistics & Data Analysis, 87(??):73–83, July 2015. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947315000274.

Tang:2015:MBC

[1707] Yang Tang, Ryan P. Browne, and Paul D. McNicholas. Model based clustering of high-dimensional binary data. Computational Statistics & Data Analysis, 87(??):84– 101, July 2015. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947314003570.

Epifanio:2015:ANA

[1708] Guillermo Vinuéand Irene Epifanio and Sandra Alemany. Archetypoids: a new approach to define representative archetypal data. Computational Statistics & Data Analysis, 87 (??):102–115, July 2015. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947315000298.

Yu:2015:AQM

[1709] Dalei Yu, Peng Bai, and Chang Ding. Adjusted quasi-maximum likelihood estimator for mixed regressive, spatial autoregressive model and its small sample bias. Computational Statistics & Data Analysis, 87(??): 116–135, July 2015. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947315000328.

Anonymous:2015:Cg

[1710] Anonymous. Contents. Computational Statistics & Data Analysis, 87 (??):ibc, July 2015. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947315000663.

Anonymous:2015:EBg

[1711] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 87(??):ibc, July 2015. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947315000651.

Anonymous:2015:EBh

[1712] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 88(??):iii–vi, August 2015. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947315000894.

Anonymous:2015:Ch

[1713] Anonymous. Contents. Computational Statistics & Data Analysis, 88 (??):vii–viii, August 2015. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947315000900.

Rajesh:2015:LLE

[1714] G. Rajesh, E. I. Abdul-Sathar, and R. Maya. Local linear estimation of residual entropy function of conditional distributions. Computational Statistics & Data Analysis, 88(??): 1–14, August 2015. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947315000316.

Yu:2015:TSA

[1715] Wenbao Yu and Taesung Park. Two simple algorithms on linear combination of multiple biomarkers to maximize partial area under the ROC curve. Computational Statistics & Data Analysis, 88(??):15– 27, August 2015. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947314003405.

Stober:2015:CCD

[1716] Jakob Stöber, Hyokyoung Grace Hong, Claudia Czado, and Pulak Ghosh. Comorbidity of chronic diseases in the elderly: Patterns identified by a copula design for mixed responses. Computational Statistics & Data Analysis, 88(??):28–39, August 2015. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947315000304.

Ankinakatte:2015:MDL

[1717] Smitha Ankinakatte and David Edwards. Modelling discrete longitudinal data using acyclic probabilistic finite automata. Computational Statistics & Data Analysis, 88(??): 40–52, August 2015. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794731500050X.

Fu:2015:URT

[1718] Wei Fu and Jeffrey S. Simonoff. Unbiased regression trees for longitudinal and clustered data. Computational Statistics & Data Analysis, 88 (??):53–74, August 2015. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947315000432.

Mahani:2015:SPM

[1719] Alireza S. Mahani and Mansour T. A.

Sharabiani. SIMD parallel MCMC sampling with applications for bigdata Bayesian analytics. Computational Statistics & Data Analysis, 88 (??):75–99, August 2015. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947315000511.

Zhang:2015:NNC

[1720] Lingsong Zhang, Shu Lu, and J. S.

Marron. Nested nonnegative cone analysis. Computational Statistics & Data Analysis, 88(??):100– 110, August 2015. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947315000195.

Ling:2015:BLN

[1721] Bingo Wing-Kuen Ling, Charlotte Yuk-Fan Ho, Wan-Chi Siu, and Qingyun Dai. Best linear near unbiased estimation for nonlinear signal models via semi-infinite programming approach. Computational Statistics & Data Analysis, 88(??):111– 118, August 2015. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947315000481.

Lin:2015:GOC

[1722] Yanzhu Lin, Min Zhang, and Dabao Zhang. Generalized orthogonal components regression for high dimensional generalized linear models. Computational Statistics & Data Analysis, 88 (??):119–127, August 2015. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947315000456.

Conde-Amboage:2015:LFT

[1723] Mercedes Conde-Amboage, César Sánchez-Sellero, and Wenceslao González-Manteiga. A lack-of-fit test for quantile regression models with highdimensional covariates. Computational Statistics & Data Analysis, 88 (??):128–138, August 2015. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947315000572.

Zhao:2015:MMS

[1724] Jianhua Zhao, Libin Jin, and Lei Shi.

Mixture model selection via hierarchical BIC. Computational Statistics & Data Analysis, 88(??):139– 153, August 2015. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794731500047X.

Xie:2015:QRM

[1725] Shangyu Xie, Alan T. K. Wan, and Yong Zhou. Quantile regression methods with varying-coefficient models for censored data. Computational Statistics & Data Analysis, 88(??): 154–172, August 2015. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947315000523.

Chan:2015:MCL

[1726] Moon tong Chan, Dalei Yu, and Kelvin K. W. Yau. Multilevel cumulative logistic regression model with random effects: Application to British social attitudes panel survey data. Computational Statistics & Data Analysis, 88 (??):173–186, August 2015. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947315000596.

Paul:2015:SOE

[1727] Subhadeep Paul and Ayanendranath Basu. On second order efficient robust inference. Computational Statistics & Data Analysis, 88(??):187– 207, August 2015. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947315000493.

Anonymous:2015:EBi

[1728] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 89 (??):iii–v, September 2015. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947315001127.

Anonymous:2015:Ci

[1729] Anonymous. Contents. Computational Statistics & Data Analysis, 89 (??):vi–vii, September 2015. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947315001139.

Swihart:2015:MSF

[1730] Bruce J. Swihart, Naresh M. Punjabi, and Ciprian M. Crainiceanu. Modeling sleep fragmentation in sleep hypnograms: an instance of fast, scalable discrete-state, discrete-time analyses. Computational Statistics & Data Analysis, 89(??):1–11, September 2015. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947315000699.

Vu:2015:VAB

[1731] Duy Vu and Murray Aitkin. Variational algorithms for biclustering models. Computational Statistics & Data Analysis, 89(??):12– 24, September 2015. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947315000560.

Shen:2015:ATM

[1732] Yanfeng Shen and Zhengyan Lin. An adaptive test for the mean vector in large-p-small-n problems. Computational Statistics & Data Analysis, 89 (??):25–38, September 2015. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947315000729.

Ye:2015:SAR

[1733] Peng Ye, Xingqiu Zhao, Liuquan Sun, and Wei Xu. A semiparametric additive rates model for multivariate recurrent events with missing event categories. Computational Statistics & Data Analysis, 89(??): 39–50, September 2015. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947315000705.

Wang:2015:ICV

[1734] Qing Wang and Bruce G. Lindsay. Improving cross-validated bandwidth selection using subsamplingextrapolation techniques. Computational Statistics & Data Analysis, 89 (??):51–71, September 2015. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947315000730.

Hino:2015:NPE

[1735] Hideitsu Hino, Kensuke Koshijima, and Noboru Murata. Non-parametric entropy estimators based on simple linear regression. Computational Statistics & Data Analysis, 89(??): 72–84, September 2015. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947315000791.

Tian:2015:EME

[1736] Yuzhu Tian, Qianqian Zhu, and Maozai Tian. Estimation for mixed exponential distributions under type-II progressively hybrid censored samples. Computational Statistics & Data Analysis, 89(??):85– 96, September 2015. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947315000717.

Tian:2015:GEI

[1737] Guo-Liang Tian, Huijuan Ma, Yong Zhou, and Dianliang Deng. Generalized endpoint-inflated binomial model. Computational Statistics & Data Analysis, 89(??):97–114, September 2015. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947315000778.

Lazariv:2015:BET

[1738] Taras Lazariv, Yarema Okhrin, and Wolfgang Schmid. Behavior of EWMA type control charts for small smoothing parameters. Computational Statistics & Data Analysis, 89(??):115– 125, September 2015. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794731500078X.

Yue:2015:EGI

[1739] Chen Yue, Shaojie Chen, Haris I.

Sair, Raag Airan, and Brian S. Caffo. Estimating a graphical intraclass correlation coefficient (GICC) using multivariate probit-linear mixed models. Computational Statistics & Data Analysis, 89(??):126–133, September 2015. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947315000535.

Zhou:2015:AMS

[1740] Ruoyu Zhou, Lianjie Shu, and Yan Su. An adaptive minimum spanning tree test for detecting irregularlyshaped spatial clusters. Computational Statistics & Data Analysis, 89(??): 134–146, September 2015. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947315000766.

Lee:2015:REM

[1741] Sangin Lee, Yudi Pawitan, and Youngjo Lee. A random-effect model approach for group variable selection. Computational Statistics & Data Analysis, 89(??):147– 157, September 2015. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794731500081X.

Torabi:2015:LIS

[1742] Mahmoud Torabi, Subhash R. Lele, and Narasimha G. N. Prasad. Likelihood inference for small area estimation using data cloning. Computational Statistics & Data Analysis, 89(??): 158–171, September 2015. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947315000821.

nez-Gamero:2015:FGF

[1743] M. Dolores Jiménez-Gamero and Hyoung-Moon Kim. Fast goodnessof-fit tests based on the characteristic function. Computational Statistics & Data Analysis, 89(??):172– 191, September 2015. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947315000845.

Kawano:2015:SPC

[1744] Shuichi Kawano, Hironori Fujisawa, Toyoyuki Takada, and Toshihiko Shiroishi. Sparse principal component regression with adaptive loading. Computational Statistics & Data Analysis, 89(??):192–203, September 2015. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947315000936.

Zhao:2015:OPB

[1745] Yan-Yong Zhao, Jin-Guan Lin, Pei-Rong Xu, and Xu-Guo Ye. Orthogonalityprojection-based estimation for semivarying coefficient models with heteroscedastic errors. Computational Statistics & Data Analysis, 89(??): 204–221, September 2015. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794731500095X.

Gonzalez:2015:BNE

[1746] Jorge González, Andrés F. Barrientos, and Fernando A. Quintana. Bayesian nonparametric estimation of test equating functions with covariates. Computational Statistics & Data Analysis, 89(??):222– 244, September 2015. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947315000808.

Anonymous:2015:EBj

[1747] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 90(??):iii–v, October 2015. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947315001280.

Anonymous:2015:Cj

[1748] Anonymous. Contents. Computational Statistics & Data Analysis, 90 (??):vi, October 2015. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947315001292.

Schliep:2015:DAP

[1749] Erin M. Schliep and Jennifer A. Hoeting. Data augmentation and parameter expansion for independent or spatially correlated ordinal data. Computational Statistics & Data Analysis, 90(??):1–14, October 2015. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947315000973.

Gregorutti:2015:GVI

[1750] Baptiste Gregorutti, Bertrand Michel, and Philippe Saint-Pierre. Grouped variable importance with random forests and application to multiple functional data analysis. Computational Statistics & Data Analysis, 90 (??):15–35, October 2015. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947315000997.

Prater:2015:FDS

[1751] Ashley Prater, Lixin Shen, and Bruce W. Suter. Finding Dantzig selectors with a proximity operator based fixed-point algorithm. Computational Statistics & Data Analysis, 90 (??):36–46, October 2015. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947315001024.

Karunamuni:2015:REE

[1752] Rohana J. Karunamuni, Qingguo Tang, and Bangxin Zhao. Robust and efficient estimation of effective dose. Computational Statistics & Data Analysis, 90(??):47–60, October 2015. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947315000985.

Wraith:2015:LSM

[1753] Darren Wraith and Florence Forbes.

Location and scale mixtures of Gaussians with flexible tail behaviour: Properties, inference and application to multivariate clustering. Computational Statistics & Data Analysis, 90 (??):61–73, October 2015. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794731500105X.

Ahn:2015:SHD

[1754] Jeongyoun Ahn and Yongho Jeon.

Sparse HDLSS discrimination with constrained data piling. Computational Statistics & Data Analysis, 90 (??):74–83, October 2015. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947315001036.

Tutz:2015:IMI

[1755] Gerhard Tutz and Shahla Ramzan.

Improved methods for the imputation of missing data by nearest neighbor methods. Computational Statistics & Data Analysis, 90(??): 84–99, October 2015. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947315001061.

Jovanovic:2015:TEB

[1756] Milan Jovanović, Bojana Milošević, Ya. Yu. Nikitin, Marko Obradović, and K. Yu. Volkova. Tests of exponentiality based on Arnold–Villasenor characterization and their efficiencies. Computational Statistics & Data Analysis, 90 (??):100–113, October 2015. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947315000961.

Tenenhaus:2015:KGC

[1757] Arthur Tenenhaus, Cathy Philippe, and Vincent Frouin. Kernel generalized canonical correlation analysis. Computational Statistics & Data Analysis, 90 (??):114–131, October 2015. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947315001012.

Anonymous:2015:EBk

[1758] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 91(??):iii–v, November 2015. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947315001541.

Anonymous:2015:Ck

[1759] Anonymous. Contents. Computational Statistics & Data Analysis, 91 (??):vi, November 2015. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947315001553.

Jarocinski:2015:NID

[1760] Marek Jarociński. A note on implementing the Durbin and Koopman simulation smoother. Computational Statistics & Data Analysis, 91 (??):1–3, November 2015. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947315001176.

Allassonniere:2015:CSE

[1761] Stéphanie Allassonnière and Estelle Kuhn. Convergent stochastic Expectation Maximization algorithm with efficient sampling in high dimension. Application to deformable template model estimation. Computational Statistics & Data Analysis, 91(??):4– 19, November 2015. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947315001164.

Charlier:2015:CQE

[1762] Isabelle Charlier, Davy Paindaveine, and Jérôme Saracco. Conditional quantile estimation based on optimal quantization: From theory to practice. Computational Statistics & Data Analysis, 91(??):20– 39, November 2015. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947315001358.

Liu:2015:JLC

[1763] Yue Liu, Lei Liu, and Jianhui Zhou.

Joint latent class model of survival and longitudinal data: an application to CPCRA study. Computational Statistics & Data Analysis, 91 (??):40–50, November 2015. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947315001346.

Barreto-Souza:2015:LTS

[1764] Wagner Barreto-Souza. Long-term survival models with overdispersed number of competing causes. Computational Statistics & Data Analysis, 91 (??):51–63, November 2015. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947315001206.

Wang:2015:HMF

[1765] Guochang Wang, Yan Zhou, Xiang-Nan Feng, and Baoxue Zhang. The hybrid method of FSIR and FSAVE for functional effective dimension reduction. Computational Statistics & Data Analysis, 91(??):64– 77, November 2015. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947315001383.

Xu:2015:EMA

[1766] Ganggang Xu and Marc G. Genton. Efficient maximum approximated likelihood inference for Tukey’s g-and-h distribution. Computational Statistics & Data Analysis, 91(??):78– 91, November 2015. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947315001401.

Zhao:2015:JEL

[1767] Yichuan Zhao, Xueping Meng, and Hanfang Yang. Jackknife empirical likelihood inference for the mean absolute deviation. Computational Statistics & Data Analysis, 91(??):

92–101, November 2015. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947315001395.

Anonymous:2015:EBl

[1768] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 92(??):iii–v, December 2015. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947315001826.

Anonymous:2015:Cl

[1769] Anonymous. Contents. Computational Statistics & Data Analysis, 92 (??):vi–vii, December 2015. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947315001838.

Friedrich:2015:FIV

[1770] Ulf Friedrich, Ralf Münnich, Sven de Vries, and Matthias Wagner. Fast integer-valued algorithms for optimal allocations under constraints in stratified sampling. Computational Statistics & Data Analysis, 92(??):1– 12, December 2015. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947315001413.

Pesonen:2015:CME

[1771] Maiju Pesonen, Henri Pesonen, and Jaakko Nevalainen. Covariance matrix estimation for left-censored data. Computational Statistics & Data Analysis, 92(??):13–25, December 2015. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947315001437.

Ryan:2015:SBF

[1772] Elizabeth G. Ryan, Christopher C. Drovandi, and Anthony N. Pettitt. Simulation-based fully Bayesian experimental design for mixed effects models. Computational Statistics & Data Analysis, 92(??):26– 39, December 2015. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947315001450.

Yang:2015:SAN

[1773] Hu Yang and Danhui Yi. Studies of the adaptive network-constrained linear regression and its application. Computational Statistics & Data Analysis, 92 (??):40–52, December 2015. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947315001462.

Kwon:2015:MCL

[1774] Sunghoon Kwon, Sangin Lee, and Yongdai Kim. Moderately clipped LASSO. Computational Statistics & Data Analysis, 92(??):53– 67, December 2015. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947315001589.

Zhu:2015:EIC

[1775] Liping Zhu and Wei Zhong. Estimation and inference on central mean subspace for multivariate response data. Computational Statistics & Data Analysis, 92 (??):68–83, December 2015. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947315001231.

Rashid:2015:UMD

[1776] S. Rashid, R. Mitra, and R. J. Steele.

Using mixtures of t densities to make inferences in the presence of missing data with a small number of multiply imputed data sets. Computational Statistics & Data Analysis, 92 (??):84–96, December 2015. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794731500136X.

Liu:2015:TET

[1777] Zhi Liu, Xiaochao Xia, and Wang Zhou. A test for equality of two distributions via jackknife empirical likelihood and characteristic functions. Computational Statistics & Data Analysis, 92(??):97– 114, December 2015. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947315001425.

Wu:2015:MCD

[1778] Tong Tong Wu and Kenneth Lange.

Matrix completion discriminant analysis. Computational Statistics & Data Analysis, 92(??):115–125, December 2015. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0167947315001449.

Lin:2015:DMR

[1779] Kuo-Chin Lin and Yi-Ju Chen. Detecting misspecification in the randomeffects structure of cumulative logit models. Computational Statistics & Data Analysis, 92(??):126–133, December 2015. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947315001590.

Salehabadi:2015:RUC

[1780] Sedigheh Mirzaei Salehabadi and Debasis Sengupta. Regression under Cox’s model for recall-based time-to-event data in observational studies. Computational Statistics & Data Analysis, 92 (??):134–147, December 2015. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947315001620.

Funke:2015:NDE

[1781] Benedikt Funke and Rafael Kawka. Nonparametric density estimation for multivariate bounded data using two non-negative multiplicative bias correction methods. Computational Statistics & Data Analysis, 92(??):148– 162, December 2015. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947315001632.

Schafer:2015:BMM

[1782] Martin Schäfer, Yvonne Radon, Thomas Klein, Sabrina Herrmann, Holger Schwender, Peter J. Verveer, and Katja Ickstadt. A Bayesian mixture model to quantify parameters of spatial clustering. Computational Statistics & Data Analysis, 92(??): 163–176, December 2015. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947315001619.

Anonymous:2016:EBa

[1783] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 93(??):iii–v, January 2016. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947315002157.

Anonymous:2016:Ca

[1784] Anonymous. Contents. Computational Statistics & Data Analysis, 93 (??):vi–viii, January 2016. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947315002169.

Kontoghiorghes:2016:CSI

[1785] Erricos J. Kontoghiorghes. CSDA special issues. Computational Statistics & Data Analysis, 93(??):1, January 2016. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0167947315002285.

Hinde:2016:TSI

[1786] John Hinde, Salvatore Ingrassia, Tsung-I Lin, and Paul McNicholas. The third special issue on advances in mixture models. Computational Statistics & Data Analysis, 93(??): 2–4, January 2016. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947315002029.

Scrucca:2016:ICC

[1787] Luca Scrucca. Identifying connected components in Gaussian finite mixture models for clustering. Computational Statistics & Data Analysis, 93(??):5–17, January 2016. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947315000171.

OHagan:2016:CMN

[1788] Adrian O’Hagan, Thomas Brendan Murphy, Isobel Claire Gormley, Paul D. McNicholas, and Dimitris Karlis. Clustering with the multivariate normal inverse Gaussian distribution. Computational Statistics & Data Analysis, 93(??):18– 30, January 2016. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947314002667.

Melnykov:2016:MBB

[1789] Volodymyr Melnykov. Model-based biclustering of clickstream data. Computational Statistics & Data Analysis, 93(??):31–45, January 2016. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947314002771.

Fernandez:2016:MBC

[1790] D. Fernández, R. Arnold, and S. Pledger. Mixture-based clustering for the ordered stereotype model. Computational Statistics & Data Analysis, 93(??):46–75, January 2016. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794731400317X.

Nguyen:2016:MSS

[1791] Hien D. Nguyen, Geoffrey J. McLachlan, and Ian A. Wood. Mixtures of spatial spline regressions for clustering and classification. Computational Statistics & Data Analysis, 93 (??):76–85, January 2016. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794731400022X.

Ciarleglio:2016:WBS

[1792] Adam Ciarleglio and R. Todd Ogden.

Wavelet-based scalar-on-function finite mixture regression models. Computational Statistics & Data Analysis, 93 (??):86–96, January 2016. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947314003454.

Papastamoulis:2016:EMP

[1793] Panagiotis Papastamoulis, Marie-Laure Martin-Magniette, and Cathy Maugis-Rabusseau. On the estimation of mixtures of Poisson regression models with large number of components. Computational Statistics & Data Analysis, 93(??):97– 106, January 2016. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947314002096.

Chee:2016:NES

[1794] Chew-Seng Chee and Yong Wang. Nonparametric estimation of species richness using discrete k-monotone distributions. Computational Statistics & Data Analysis, 93(??):107– 118, January 2016. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947314003132.

Lu:2016:FMN

[1795] Xiaosun Lu, Yangxin Huang, and Yiliang Zhu. Finite mixture of nonlinear mixed-effects joint models in the presence of missing and mismeasured covariate, with application to AIDS studies. Computational Statistics & Data Analysis, 93(??):119– 130, January 2016. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947314000991.

Bolano:2016:GFM

[1796] Danilo Bolano and André Berchtold. General framework and model building in the class of Hidden Mixture Transition Distribution models. Computational Statistics & Data Analysis, 93 (??):131–145, January 2016. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947314002722.

Morgan:2016:LPA

[1797] Grant B. Morgan, Kari J. Hodge, and Aaron R. Baggett. Latent profile analysis with nonnormal mixtures: a Monte Carlo examination of model selection using fit indices. Computational Statistics & Data Analysis, 93 (??):146–161, January 2016. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947315000602.

Wu:2016:MQR

[1798] Qiang Wu and Weixin Yao. Mixtures of quantile regressions. Computational Statistics & Data Analysis, 93 (??):162–176, January 2016. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947314001212.

Nguyen:2016:LML

[1799] Hien D. Nguyen and Geoffrey J.

McLachlan. Laplace mixture of linear experts. Computational Statistics & Data Analysis, 93(??):177– 191, January 2016. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947314003089.

Cheam:2016:MRO

[1800] Amay S. M. Cheam and Paul D. Mc-Nicholas. Modelling receiver operating characteristic curves using Gaussian mixtures. Computational Statistics & Data Analysis, 93(??):192– 208, January 2016. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947315001073.

Gonzalez:2016:MLE

[1801] M. González, C. Minuesa, and I. del Puerto. Maximum likelihood estimation and expectation-maximization algorithm for controlled branching processes. Computational Statistics & Data Analysis, 93(??):209– 227, January 2016. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947315000262.

Dou:2016:EAE

[1802] Xiaoling Dou, Satoshi Kuriki, Gwo Dong Lin, and Donald Richards. EM algorithms for estimating the Bernstein copula. Computational Statistics & Data Analysis, 93(??):228– 245, January 2016. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947314000206.

McGrory:2016:TSM

[1803] C. A. McGrory, A. N. Pettitt, D. M. Titterington, C. L. Alston, and M. Kelly. Transdimensional sequential Monte Carlo using variational Bayes — SMCVB. Computational Statistics & Data Analysis, 93(??): 246–254, January 2016. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947315000742.

Datta:2016:SIA

[1804] Somnath Datta, María del Carmen Pardo, Thomas Scheike, and Kam C. Yuen. Special issue on advances in survival analysis. Computational Statistics & Data Analysis, 93 (??):255–256, January 2016. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947315002030.

Hu:2016:PLT

[1805] Tao Hu and Liming Xiang. Partially linear transformation cure models for interval-censored data. Computational Statistics & Data Analysis, 93 (??):257–269, January 2016. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947314002461.

Bremhorst:2016:FEC

[1806] Vincent Bremhorst and Philippe Lambert. Flexible estimation in cure survival models using Bayesian P -splines. Computational Statistics & Data Analysis, 93(??):270– 284, January 2016. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947314001492.

Shen:2016:ELC

[1807] Junshan Shen, Kam Chuen Yuen, and Chunling Liu. Empirical likelihood confidence regions for one- or two- samples with doubly censored data. Computational Statistics & Data Analysis, 93 (??):285–293, January 2016. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947315000213.

Moreira:2016:NRD

[1808] C. Moreira, J. de Uña-Álvarez, and L. Meira-Machado. Nonparametric regression with doubly truncated data. Computational Statistics & Data Analysis, 93(??):294– 307, January 2016. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947314000942.

Talamakrouni:2016:PGN

[1809] Majda Talamakrouni, Ingrid Van Keilegom, and Anouar El Ghouch. Parametrically guided nonparametric density and hazard estimation with censored data. Computational Statistics & Data Analysis, 93(??):308– 323, January 2016. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947315000201.

Holst:2016:LTM

[1810] Klaus K. Holst, Thomas H. Scheike, and Jacob B. Hjelmborg. The liability threshold model for censored twin data. Computational Statistics & Data Analysis, 93(??):324– 335, January 2016. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947315000250.

Luo:2016:CMA

[1811] Peng Luo and Zhi Geng. Causal mediation analysis for survival outcome with unobserved mediator-outcome confounders. Computational Statistics & Data Analysis, 93(??):336– 347, January 2016. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947314003442.

Sun:2016:GFT

[1812] Yanqing Sun, Mei Li, and Peter B.

Gilbert. Goodness-of-fit test of the stratified mark-specific proportional hazards model with continuous mark. Computational Statistics & Data Analysis, 93(??):348– 358, January 2016. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947314003351.

Arbel:2016:FBI

[1813] Julyan Arbel, Antonio Lijoi, and Bernardo Nipoti. Full Bayesian inference with hazard mixture models. Computational Statistics & Data Analysis, 93(??):359–372, January 2016. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0167947314003417.

Rancoita:2016:BND

[1814] Paola M. V. Rancoita, Marco Zaffalon, Emanuele Zucca, Francesco Bertoni, and Cassio P. de Campos. Bayesian network data imputation with application to survival tree analysis. Computational Statistics & Data Analysis, 93 (??):373–387, January 2016. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947314003569.

Berry:2016:SIA

[1815] Michael W. Berry, Jung Jin Lee, Giovanni Montana, Stefan Van Aelst, and Ruben H. Zamar. Special issue on advances in data mining and robust statistics. Computational Statistics & Data Analysis, 93(??):388– 389, January 2016. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947315002297.

Cui:2016:SEH

[1816] Ying Cui, Chenlei Leng, and Defeng Sun. Sparse estimation of highdimensional correlation matrices. Computational Statistics & Data Analysis, 93(??):390–403, January 2016. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S016794731400293X.

Tarr:2016:REP

[1817] G. Tarr, S. Müller, and N. C. Weber. Robust estimation of precision matrices under cellwise contamination. Computational Statistics & Data Analysis, 93 (??):404–420, January 2016. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947315000444.

Alfons:2016:RGL

[1818] Andreas Alfons, Christophe Croux, and Sarah Gelper. Robust groupwise least angle regression. Computational Statistics & Data Analysis, 93 (??):421–435, January 2016. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947315000468.

Salibian-Barrera:2016:RTL

[1819] Matias Salibian-Barrera, Stefan Van Aelst, and Víctor J. Yohai. Robust tests for linear regression models based on τ-estimates. Computational Statistics & Data Analysis, 93 (??):436–455, January 2016. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947314002734.

Kirschstein:2016:MVP

[1820] T. Kirschstein, S. Liebscher, G. C.

Porzio, and G. Ragozini. Minimum volume peeling: a robust nonparametric estimator of the multivariate mode. Computational Statistics & Data Analysis, 93(??):456– 468, January 2016. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947315001188.

Hamalainen:2016:NUB

[1821] Wilhelmiina Hämäläinen. New upper bounds for tight and fast approximation of Fisher’s exact test in dependency rule mining. Computational Statistics & Data Analysis, 93 (??):469–482, January 2016. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947315001747.

Martinez:2016:NPM

[1822] Waldyn Martinez and J. Brian Gray. Noise peeling methods to improve boosting algorithms. Computational Statistics & Data Analysis, 93(??): 483–497, January 2016. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947315001486.

Anonymous:2016:EBb

[1823] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 94(??):iii–v, February 2016. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947315002534.

Anonymous:2016:Cb

[1824] Anonymous. Contents. Computational Statistics & Data Analysis, 94 (??):vi–vii, February 2016. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947315002546.

Kim:2016:NER

[1825] Minjo Kim and Sangyeol Lee. Nonlinear expectile regression with application to Value-at-Risk and expected shortfall estimation. Computational Statistics & Data Analysis, 94 (??):1–19, February 2016. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947315001681.

Gardner-Lubbe:2016:TMC

[1826] Sugnet Gardner-Lubbe. A triplot for multiclass classification visualisation. Computational Statistics & Data Analysis, 94(??):20– 32, February 2016. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947315001711.

Zhou:2016:PMS

[1827] Jingke Zhou and Lixing Zhu. Principal minimax support vector machine for sufficient dimension reduction with contaminated data. Computational Statistics & Data Analysis, 94 (??):33–48, February 2016. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947315001498.

Tang:2016:FAC

[1828] Qingguo Tang and Rohana J. Karunamuni. Fast and accurate computation for kernel estimators. Computational Statistics & Data Analysis, 94 (??):49–62, February 2016. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947315001723.

Hasegawa:2016:LFF

[1829] Takanori Hasegawa, Atsushi Niida, Tomoya Mori, Teppei Shimamura, Rui Yamaguchi, Satoru Miyano, Tatsuya Akutsu, and Seiya Imoto. A likelihoodfree filtering method via approximate Bayesian computation in evaluating biological simulation models. Computational Statistics & Data Analysis, 94 (??):63–74, February 2016. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947315001760.

Martin:2016:SSM

[1830] Ryan Martin and Zhen Han. A semiparametric scale-mixture regression model and predictive recursion maximum likelihood. Computational Statistics & Data Analysis, 94(??): 75–85, February 2016. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947315001929.

Ng:2016:DCV

[1831] F. C. Ng, W. K. Li, and Philip L. H. Yu. Diagnostic checking of the vector multiplicative error model. Computational Statistics & Data Analysis, 94 (??):86–97, February 2016. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947315001693.

Jiang:2016:PSA

[1832] Depeng Jiang, Puying Zhao, and Niansheng Tang. A propensity score adjustment method for regression models with nonignorable missing covariates. Computational Statistics & Data Analysis, 94(??):98– 119, February 2016. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947315001930.

Wheatley:2016:HPR

[1833] Spencer Wheatley, Vladimir Filimonov, and Didier Sornette. The Hawkes process with renewal immigration & its estimation with an EM algorithm. Computational Statistics & Data Analysis, 94(??):120– 135, February 2016. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947315001954.

Bagheri:2016:GMW

[1834] S. F. Bagheri, E. Bahrami Samani, and M. Ganjali. The generalized modified Weibull power series distribution: Theory and applications. Computational Statistics & Data Analysis, 94 (??):136–160, February 2016. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947315001966.

Chesneau:2016:NEQ

[1835] Christophe Chesneau, Isha Dewan, and Hassan Doosti. Nonparametric estimation of a quantile density function by wavelet methods. Computational Statistics & Data Analysis, 94 (??):161–174, February 2016. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947315001942.

Lavancier:2016:GPC

[1836] F. Lavancier and P. Rochet. A general procedure to combine estimators. Computational Statistics & Data Analysis, 94(??):175– 192, February 2016. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947315001735.

Marino:2016:GQA

[1837] Maria Francesca Marino and Marco Alfó. Gaussian quadrature approximations in mixed hidden Markov models for longitudinal data: a simulation study. Computational Statistics & Data Analysis, 94(??):193– 209, February 2016. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947315001759.

Bebu:2016:PCE

[1838] Ionut Bebu, George Luta, Thomas Mathew, Paul A. Kennedy, and Brian K. Agan. Parametric costeffectiveness inference with skewed data. Computational Statistics & Data Analysis, 94(??):210–220, February 2016. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947315002054.

Qian:2016:SPR

[1839] Junyang Qian and Jinzhu Jia. On stepwise pattern recovery of the fused lasso. Computational Statistics & Data Analysis, 94(??):221– 237, February 2016. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947315002017.

Zhang:2016:SLR

[1840] Xin Zhang, Daniel R. Jeske, Jun Li, and Vance Wong. A sequential logistic regression classifier based on mixed effects with applications to longitudinal data. Computational Statistics & Data Analysis, 94(??): 238–249, February 2016. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947315001978.

Svenson:2016:MOE

[1841] Joshua Svenson and Thomas Santner. Multiobjective optimization of expensive-to-evaluate deterministic computer simulator models. Computational Statistics & Data Analysis, 94 (??):250–264, February 2016. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947315001991.

Uno:2016:FFA

[1842] Kohei Uno, Hironori Satomura, and Kohei Adachi. Fixed factor analysis with clustered factor score constraint. Computational Statistics & Data Analysis, 94(??):265– 274, February 2016. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794731500198X.

Lee:2016:MLQ

[1843] Sangin Lee, Sunghoon Kwon, and Yongdai Kim. A modified local quadratic approximation algorithm for penalized optimization problems. Computational Statistics & Data Analysis, 94(??):275– 286, February 2016. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947315002078.

Lin:2016:ACF

[1844] Lu Lin and Jing Sun. Adaptive conditional feature screening. Computational Statistics & Data Analysis, 94 (??):287–301, February 2016. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947315002108.

Zang:2016:JEL

[1845] Yangguang Zang, Sanguo Zhang, Qizhai Li, and Qingzhao Zhang. Jackknife empirical likelihood test for high-dimensional regression coefficients. Computational Statistics & Data Analysis, 94(??):302– 316, February 2016. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947315002005.

Usset:2016:IMF

[1846] Joseph Usset, Ana-Maria Staicu, and Arnab Maity. Interaction models for functional regression. Computational Statistics & Data Analysis, 94(??): 317–329, February 2016. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794731500208X.

Hron:2016:SPC

[1847] K. Hron, A. Menafoglio, M. Templ, K. Hruzová, and P. Filzmoser. Simplicial principal component analysis for density functions in Bayes spaces. Computational Statistics & Data Analysis, 94(??):330–350, February 2016. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947315001644.

Aykroyd:2016:CMT

[1848] Robert G. Aykroyd, Stuart Barber, and Luke R. Miller. Classification of multiple time signals using localized frequency characteristics applied to industrial process monitoring. Computational Statistics & Data Analysis, 94 (??):351–362, February 2016. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947315001668.

Lee:2016:ALS

[1849] Keunbaik Lee, Insuk Sohn, and Donguk Kim. Analysis of long series of longitudinal ordinal data using marginalized models. Computational Statistics & Data Analysis, 94 (??):363–371, February 2016. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794731500167X.

Benavent:2016:MFH

[1850] Roberto Benavent and Domingo Morales. Multivariate Fay–Herriot models for small area estimation. Computational Statistics & Data Analysis, 94(??):372–390, February 2016. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S016794731500170X.

Anonymous:2016:EBc

[1851] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 95(??):iii–v, March 2016. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947315002777.

Anonymous:2016:Cc

[1852] Anonymous. Contents. Computational Statistics & Data Analysis, 95 (??):vi–vii, March 2016. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947315002789.

Lee:2016:BVS

[1853] Kuo-Jung Lee, Ray-Bing Chen, and Ying Nian Wu. Bayesian variable selection for finite mixture model of linear regressions. Computational Statistics & Data Analysis, 95(??): 1–16, March 2016. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947315002303.

Xu:2016:LCI

[1854] Ping-Feng Xu, Jubo Sun, and Na Shan.

Local computations of the iterative proportional scaling procedure for hierarchical models. Computational Statistics & Data Analysis, 95(??): 17–23, March 2016. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947315002649.

Lee:2016:MIM

[1855] Min Cherng Lee and Robin Mitra.

Multiply imputing missing values in data sets with mixed measurement scales using a sequence of generalised linear models. Computational Statistics & Data Analysis, 95(??): 24–38, March 2016. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947315001772.

Wang:2016:BHM

[1856] Yixin Wang and Mike K. P. So.

A Bayesian hierarchical model for spatial extremes with multiple durations. Computational Statistics & Data Analysis, 95(??):39– 56, March 2016. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947315002091.

Koblents:2016:NPM

[1857] Eugenia Koblents, Joaquín Míguez, Marco A. Rodríguez, and Alexandra M. Schmidt. A nonlinear population Monte Carlo scheme for the Bayesian estimation of parameters of α-stable distributions. Computational Statistics & Data Analysis, 95: 57–74, March 2016. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947315002340.

Qin:2016:TOP

[1858] Yingli Qin and Weiming Li. Testing the order of a population spectral distribution for high-dimensional data. Computational Statistics & Data Analysis, 95(??):75–82, March 2016. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947315002364.

Zhang:2016:MIC

[1859] Tonglin Zhang and Ge Lin. On Moran’s I coefficient under heterogeneity. Computational Statistics & Data Analysis, 95(??):83–94, March 2016. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947315002376.

Ikeda:2016:CLS

[1860] Yuki Ikeda, Tatsuya Kubokawa, and Muni S. Srivastava. Comparison of linear shrinkage estimators of a large covariance matrix in normal and nonnormal distributions. Computational Statistics & Data Analysis, 95(??): 95–108, March 2016. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947315002388.

Stoklosa:2016:QLR

[1861] Jakub Stoklosa, Yih-Huei Huang, Elise Furlan, and Wen-Han Hwang. On quadratic logistic regression models when predictor variables are subject to measurement error. Computational Statistics & Data Analysis, 95 (??):109–121, March 2016. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794731500239X.

Kahle:2016:CIR

[1862] David J. Kahle, Phil D. Young, Brandi A. Greer, and Dean M. Young. Confidence intervals for the ratio of two Poisson rates under one-way differential misclassification using double sampling. Computational Statistics & Data Analysis, 95(??):122– 132, March 2016. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947315002406.

Secchi:2016:HIC

[1863] Piercesare Secchi, Simone Vantini, and Paolo Zanini. Hierarchical independent component analysis: a multi-resolution non-orthogonal datadriven basis. Computational Statistics & Data Analysis, 95(??):133– 149, March 2016. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947315002418.

Bandyopadhyay:2016:CCS

[1864] Dipankar Bandyopadhyay and M. Amalia Jácome. Comparing conditional survival functions with missing population marks in a competing risks model. Computational Statistics & Data Analysis, 95(??):150–160, March 2016. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0167947315002467.

Gutierrez:2016:TDB

[1865] Luis Gutiérrez, Ramsés H. Mena, and Matteo Ruggiero. A time dependent Bayesian nonparametric model for air quality analysis. Computational Statistics & Data Analysis, 95 (??):161–175, March 2016. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947315002479.

Shi:2016:CDD

[1866] Lei Shi, Jun Lu, Jianhua Zhao, and Gemai Chen. Case deletion diagnostics for GMM estimation. Computational Statistics & Data Analysis, 95 (??):176–191, March 2016. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947315002480.

Cheng:2016:EFT

[1867] Cheng Cheng. Exploratory failure time analysis in large scale genomics. Computational Statistics & Data Analysis, 95(??):192–206, March 2016. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947315002595.

Ollier:2016:SAF

[1868] Edouard Ollier, Adeline Samson, Xavier Delavenne, and Vivian Viallon. A SAEM algorithm for fused lasso penalized NonLinear Mixed Effect Models: Application to group comparison in pharmacokinetics. Computational Statistics & Data Analysis, 95 (??):207–221, March 2016. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947315002613.

He:2016:RQR

[1869] Qianchuan He, Linglong Kong, Yanhua Wang, Sijian Wang, Timothy A. Chan, and Eric Holland. Regularized quantile regression under heterogeneous sparsity with application to quantitative genetic traits. Computational Statistics & Data Analysis, 95 (??):222–239, March 2016. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947315002625.

Hirukawa:2016:CSN

[1870] Masayuki Hirukawa. Corrigendum to “Nonparametric multiplicative bias correction for kernel-type density estimation on the unit interval” [Comput. Statist. Data Anal. 54 (2010) 473–495]. Computational Statistics & Data Analysis, 95(??):240– 242, March 2016. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947315002352. See

Raghupathi:2016:FCL

[1871] Laks Raghupathi, David Randell, Kevin Ewans, and Philip Jonathan. Fast computation of large scale marginal extremes with multi-dimensional covariates. Computational Statistics & Data Analysis, 95(??):243– 258, March 2016. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947315002315.

Anonymous:2016:EBd

[1872] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 96(??):iii–v, April 2016. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947315003102.

Anonymous:2016:Cd

[1873] Anonymous. Contents. Computational Statistics & Data Analysis, 96 (??):vi, April 2016. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947315003114.

Zhao:2016:EMA

[1874] Kaifeng Zhao and Heng Lian. The expectation-maximization approach for Bayesian quantile regression. Computational Statistics & Data Analysis, 96(??):1–11, April 2016. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947315002844.

Wickramarachchi:2016:HOD

[1875] D. C. Wickramarachchi, B. L. Robertson, M. Reale, C. J. Price, and J. Brown. HHCART: an oblique decision tree. Computational Statistics & Data Analysis, 96(??):12– 23, April 2016. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947315002856.

Qin:2016:SMC

[1876] Guoyou Qin, Jiajia Zhang, and Zhongyi Zhu. Simultaneous mean and covariance estimation of partially linear models for longitudinal data with missing responses and covariate measurement error. Computational Statistics & Data Analysis, 96(??):24–39, April 2016. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947315002698.

Zhao:2016:EVS

[1877] Weihua Zhao, Heng Lian, Riquan Zhang, and Peng Lai. Estimation and variable selection for proportional response data with partially linear single-index models. Computational Statistics & Data Analysis, 96(??):40–56, April 2016. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947315002728.

Janitza:2016:RFO

[1878] Silke Janitza, Gerhard Tutz, and Anne-Laure Boulesteix. Random forest for ordinal responses: Prediction and variable selection. Computational Statistics & Data Analysis, 96(??):57–73, April 2016. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947315002601.

Stoklosa:2016:ESC

[1879] Jakub Stoklosa, Peter Dann, Richard M. Huggins, and Wen-Han Hwang. Estimation of survival and capture probabilities in open population capturerecapture models when covariates are subject to measurement error. Computational Statistics & Data Analysis, 96(??):74–86, April 2016. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947315002650.

Yen:2016:SVS

[1880] Tso-Jung Yen and Yu-Min Yen. Structured variable selection via prior-induced hierarchical penalty functions. Computational Statistics & Data Analysis, 96(??):87– 103, April 2016. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947315002662.

Xia:2016:REL

[1881] Xiaochao Xia, Zhi Liu, and Hu Yang. Regularized estimation for the least absolute relative error models with a diverging number of covariates. Computational Statistics & Data Analysis, 96(??):104–119, April 2016. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947315002674.

Hart:2016:FNG

[1882] Jeffrey D. Hart, Taeryon Choi, and Seongbaek Yi. Frequentist nonparametric goodness-of-fit tests via marginal likelihood ratios. Computational Statistics & Data Analysis, 96 (??):120–132, April 2016. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947315002686.

Tu:2016:BAT

[1883] Shiyi Tu, Min Wang, and Xiaoqian Sun. Bayesian analysis of two-piece location-scale models under reference priors with partial information. Computational Statistics & Data Analysis, 96(??):133–144, April 2016. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947315002704.

Petrie:2016:GTM

[1884] Adam Petrie. Graph-theoretic multisample tests of equality in distribution for high dimensional data. Computational Statistics & Data Analysis, 96(??):145–158, April 2016. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947315002716.

Anonymous:2016:EBe

[1885] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 97(??):iii–v, ???? 2016. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947316000086.

Anonymous:2016:Ce

[1886] Anonymous. Contents. Computational Statistics & Data Analysis, 97 (??):vi–vii, ???? 2016. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947316000098.

Hyndman:2016:FCR

[1887] Rob J. Hyndman, Alan J. Lee, and Earo Wang. Fast computation of reconciled forecasts for hierarchical and grouped time series. Computational Statistics & Data Analysis, 97(??):16–32, ???? 2016. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016794731500290X.

Michoel:2016:NCD

[1888] Tom Michoel. Natural coordinate descent algorithm for L^{1}-penalised regression in generalised linear models. Computational Statistics & Data Analysis, 97(??):60–70, ???? 2016. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947315002923.

Ruggieri:2016:EAB

[1889] Eric Ruggieri and Marcus Antonellis. An exact approach to Bayesian sequential change point detection. Computational Statistics & Data Analysis, 97(??):71–86, ???? 2016. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947315002935.

Golchi:2016:SCM

[1890] Shirin Golchi and David A. Campbell. Sequentially Constrained Monte Carlo. Computational Statistics & Data Analysis, 97(??):98– 113, ???? 2016. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947315002960.

Wang:2016:MFL

[1891] Shaoli Wang, Mian Huang, Xing Wu, and Weixin Yao. Mixture of functional linear models and its application to CO_{2}-GDP functional data. Computational Statistics & Data Analysis, 97(??):1–15, ???? 2016. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947315002911.

Hirotsu:2016:UAS

[1892] Chihiro Hirotsu, Shoichi Yamamoto, and Harukazu Tsuruta. A unifying approach to the shape and changepoint hypotheses in the discrete univariate exponential family. Computational Statistics & Data Analysis, 97(??):33–46, ???? 2016. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947315002959.

Yao:2016:SRA

[1893] Bin Yao, Lianming Wang, and Xin He. Semiparametric regression analysis of panel count data allowing for within-subject correlation. Computational Statistics & Data Analysis, 97(??):47–59, ???? 2016. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016794731500300X.

Zhang:2016:HDT

[1894] Jie Zhang and Meng Pan. A highdimension two-sample test for the mean using cluster subspaces. Computational Statistics & Data Analysis, 97 (??):87–97, ???? 2016. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947315003047.

Joe:2016:MMD

[1895] Harry Joe and Peijun Sang. Multivariate models for dependent clusters of variables with conditional independence given aggregation variables. Computational Statistics & Data Analysis, 97(??):114–132, ???? 2016. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0167947315003011.

Morris:2016:CCD

[1896] Katherine Morris and Paul D. Mc-Nicholas. Clustering, classification, discriminant analysis, and dimension reduction via generalized hyperbolic mixtures. Computational Statistics & Data Analysis, 97(??):

133–150, ???? 2016. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947315002637.

Boente:2016:RTS

[1897] Graciela Boente and Juan Carlos Pardo-Fernández. Robust testing for superiority between two regression curves. Computational Statistics & Data Analysis, 97(??):151– 168, ???? 2016. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947315003023.

Vilca:2016:BBS

[1898] Filidor Vilca, Renata G. Romeiro, and N. Balakrishnan. A bivariate Birnbaum–Saunders regression model. Computational Statistics & Data Analysis, 97(??):169– 183, ???? 2016. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947315003035.

Anonymous:2016:EBf

[1899] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 98(??):iii–v, ???? 2016. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794731630010X.

Anonymous:2016:Cf

[1900] Anonymous. Contents. Computational Statistics & Data Analysis, 98 (??):vi, ???? 2016. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947316300111.

Dyckerhoff:2016:ECH

[1901] Rainer Dyckerhoff and Pavlo Mozharovskyi. Exact computation of the halfspace depth. Computational Statistics & Data Analysis, 98(??):19– 30, ???? 2016. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947315003199.

Gallardo:2016:DWP

[1902] Diego I. Gallardo, Heleno Bolfarine, and Antonio Carlos Pedroso de Lima. Destructive weighted Poisson cure rate models with bivariate random effects: Classical and Bayesian approaches. Computational Statistics & Data Analysis, 98(??):31–45, ???? 2016. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794731500314X.

Ikemoto:2016:STA

[1903] Hiroki Ikemoto and Kohei Adachi. Sparse Tucker2 analysis of three-way data subject to a constrained number of zero elements in a core array. Computational Statistics & Data Analysis, 98(??):1–18, ???? 2016. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947315003151.

Bak:2016:HDC

[1904] Britta Anker Bak and Jens Ledet Jensen. High dimensional classifiers in the imbalanced case. Computational Statistics & Data Analysis, 98 (??):46–59, ???? 2016. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947315003175.

Zhao:2016:PFC

[1905] Haibing Zhao and Wing Kam Fung.

A powerful FDR control procedure for multiple hypotheses. Computational Statistics & Data Analysis, 98 (??):60–70, ???? 2016. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947315003217.

Takahashi:2016:RCC

[1906] Akihito Takahashi and Takeshi Kurosawa. Regression correlation coefficient for a Poisson regression model. Computational Statistics & Data Analysis, 98(??):71–78, ???? 2016. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947315003205.

Oedekoven:2016:UHC

[1907] C. S. Oedekoven, R. King, S. T.

Buckland, M. L. Mackenzie, K. O. Evans, and L. W. Burger, Jr. Using hierarchical centering to facilitate a reversible jump MCMC algorithm for random effects models. Computational Statistics & Data Analysis, 98(??):79–90, ???? 2016. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947315003187.

Park:2016:EET

[1908] Myung Hyun Park and Joseph H. T. Kim. Estimating extreme tail risk measures with generalized Pareto distribution. Computational Statistics & Data Analysis, 98(??):91– 104, ???? 2016. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947315003163.

Anonymous:2016:EBg

[1909] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 99(??):iii–v, July 2016. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947316300330.

Anonymous:2016:Cg

[1910] Anonymous. Contents. Computational Statistics & Data Analysis, 99 (??):vi–vii, July 2016. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947316300342.

Leung:2016:RRE

[1911] Andy Leung, Hongyang Zhang, and Ruben Zamar. Robust regression estimation and inference in the presence of cellwise and casewise contamination. Computational Statistics & Data Analysis, 99(??):1–11, July 2016. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794731600013X.

Fang:2016:IPS

[1912] Kuangnan Fang, Xiaoyan Wang, Ben-Chang Shia, and Shuangge Ma. Identification of proportionality structure with two-part models using penalization. Computational Statistics & Data Analysis, 99(??):12– 24, July 2016. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947316000037.

Li:2016:THS

[1913] Hui-Qiong Li, Guo-Liang Tian, Xue-Jun Jiang, and Nian-Sheng Tang. Testing hypothesis for a simple ordering in incomplete contingency tables. Computational Statistics & Data Analysis, 99(??):25–37, July 2016. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947316000128.

Kundu:2016:BIW

[1914] Debasis Kundu and Debanjan Mitra. Bayesian inference of Weibull distribution based on left truncated and right censored data. Computational Statistics & Data Analysis, 99(??):38–50, July 2016. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947316000025.

Chen:2016:GPA

[1915] Cathy W. S. Chen and Sangyeol Lee. Generalized Poisson autoregressive models for time series of counts. Computational Statistics & Data Analysis, 99(??):51–67, July 2016. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947316000189.

Sellers:2016:FZI

[1916] Kimberly F. Sellers and Andrew Raim. A flexible zero-inflated model to address data dispersion. Computational Statistics & Data Analysis, 99 (??):68–80, July 2016. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947316000165.

Safo:2016:GSM

[1917] Sandra E. Safo and Jeongyoun Ahn. General sparse multi-class linear discriminant analysis. Computational Statistics & Data Analysis, 99(??): 81–90, July 2016. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947316000207.

Zhao:2016:GLR

[1918] Junguang Zhao and Xingzhong Xu. A generalized likelihood ratio test for normal mean when p is greater than n. Computational Statistics & Data Analysis, 99(??):91– 104, July 2016. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947316000153.

Zhou:2016:MIA

[1919] Jie Zhou, Jiajia Zhang, Alexander C.

McLain, and Bo Cai. A multiple imputation approach for semiparametric cure model with interval censored data. Computational Statistics & Data Analysis, 99(??): 105–114, July 2016. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947316000220.

Elfadaly:2016:PEA

[1920] Fadlalla G. Elfadaly, Paul H. Garthwaite, and John R. Crawford. On point estimation of the abnormality of a Mahalanobis index. Computational Statistics & Data Analysis, 99 (??):115–130, July 2016. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947316000232.

Garcia-Escudero:2016:JRT

[1921] Luis Angel García-Escudero, Alfonso Gordaliza, Francesca Greselin, Salvatore Ingrassia, and Agustín Mayo-Iscar. The joint role of trimming and constraints in robust estimation for mixtures of Gaussian factor analyzers. Computational Statistics & Data Analysis, 99(??):131– 147, July 2016. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947316000141.

Mount:2016:PAA

[1922] David M. Mount, Nathan S. Netanyahu, Christine D. Piatko, Angela Y. Wu, and Ruth Silverman. A practical approximation algorithm for the LTS estimator. Computational Statistics & Data Analysis, 99 (??):148–170, July 2016. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947316300020.

Emily:2016:SND

[1923] Mathieu Emily, Christophe Hitte, and Alain Mom. SMILE: a novel dissimilarity-based procedure for detecting sparse-specific profiles in sparse contingency tables. Computational Statistics & Data Analysis, 99(??): 171–188, July 2016. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947316300032.

Stegeman:2016:NMS

[1924] Alwin Stegeman. A new method for simultaneous estimation of the factor model parameters, factor scores, and unique parts. Computational Statistics & Data Analysis, 99(??): 189–203, July 2016. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947316000219.

Bongiorno:2016:CMH

[1925] Enea G. Bongiorno and Aldo Goia. Classification methods for Hilbert data based on surrogate density. Computational Statistics & Data Analysis, 99(??):204–222, July 2016. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947316300056.

Fabrizi:2016:SAE

[1926] Enrico Fabrizi and Carlo Trivisano.

Small area estimation of the Gini concentration coefficient. Computational Statistics & Data Analysis, 99 (??):223–234, July 2016. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947316000190.

Serfling:2016:LSD

[1927] Robert Serfling and Yunfei Wang.

On Liu’s simplicial depth and randles’ interdirections. Computational Statistics & Data Analysis, 99(??): 235–247, July 2016. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947316300184.

Anonymous:2016:EBh

[1928] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 100(??):iii–v, August 2016. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947316300615.

Anonymous:2016:Ch

[1929] Anonymous. Contents. Computational Statistics & Data Analysis, 100 (??):vi–ix, August 2016. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947316300627.

Kontoghiorghes:2016:CAC

[1930] Erricos J. Kontoghiorghes and Herman K. Van Dijk. CFEnetwork: the Annals of Computational and Financial Econometrics, 3rd issue. Computational Statistics & Data Analysis, 100(??):1–3, August 2016. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947316300214.

Bailey:2016:SAP

[1931] Natalia Bailey and Liudas Giraitis. Spectral approach to parameter-free unit root testing. Computational Statistics & Data Analysis, 100(??): 4–16, August 2016. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794731500119X.

Bauwens:2016:EEP

[1932] Luc Bauwens, Lyudmila Grigoryeva, and Juan-Pablo Ortega. Estimation and empirical performance of nonscalar dynamic conditional correlation models. Computational Statistics & Data Analysis, 100(??):17– 36, August 2016. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947315000547.

Billio:2016:EGS

[1933] Monica Billio, Roberto Casarin, and Anthony Osuntuyi. Efficient Gibbs sampling for Markov switching GARCH models. Computational Statistics & Data Analysis, 100(??): 37–57, August 2016. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947314001182.

Blasques:2016:SSD

[1934] Francisco Blasques, Jiangyu Ji, and André Lucas. Semiparametric score driven volatility models. Computational Statistics & Data Analysis, 100 (??):58–69, August 2016. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947315001000.

Bisaglia:2016:BNF

[1935] Luisa Bisaglia and Antonio Canale.

Bayesian nonparametric forecasting for INAR models. Computational Statistics & Data Analysis, 100(??): 70–78, August 2016. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947315000109.

Caldeira:2016:PYC

[1936] João F. Caldeira, Guilherme V. Moura, and André A. P. Santos. Predicting the yield curve using forecast combinations. Computational Statistics & Data Analysis, 100(??): 79–98, August 2016. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947314001480.

Creel:2016:SSA

[1937] Michael Creel and Dennis Kristensen. On selection of statistics for approximate Bayesian computing (or the method of simulated moments). Computational Statistics & Data Analysis, 100(??):99–114, August 2016. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016794731500122X.

Dissanayake:2016:SSM

[1938] G. S. Dissanayake, M. S. Peiris, and T. Proietti. State space modeling of Gegenbauer processes with long memory. Computational Statistics & Data Analysis, 100(??):115– 130, August 2016. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947314002758.

Fengler:2016:MRR

[1939] Matthias R. Fengler and Ostap Okhrin. Managing risk with a realized copula parameter. Computational Statistics & Data Analysis, 100(??):131– 152, August 2016. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947314002151.

Fiorentini:2016:SKM

[1940] Gabriele Fiorentini, Christophe Planas, and Alessandro Rossi. Skewness and kurtosis of multivariate Markovswitching processes. Computational Statistics & Data Analysis, 100(??): 153–159, August 2016. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947315001474.

Franses:2016:STB

[1941] Philip Hans Franses. A simple test for a bubble based on growth and acceleration. Computational Statistics & Data Analysis, 100(??):160– 169, August 2016. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947314001777.

Fresoli:2016:UCR

[1942] Diego E. Fresoli and Esther Ruiz.

The uncertainty of conditional returns, volatilities and correlations in DCC models. Computational Statistics & Data Analysis, 100(??):170– 185, August 2016. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947315000948.

vanGiersbergen:2016:ACB

[1943] Noud P. A. van Giersbergen. The ability to correct the bias in the stable AD(1,1) model with a feedback effect. Computational Statistics & Data Analysis, 100(??):186– 204, August 2016. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947315001048.

Giorgi:2016:CMS

[1944] Emanuele Giorgi and Alexander J. Mc-Neil. On the computation of multivariate scenario sets for the skew-t and generalized hyperbolic families. Computational Statistics & Data Analysis, 100 (??):205–220, August 2016. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947314002035.

Groen:2016:RUA

[1945] Jan J. J. Groen and George Kapetanios. Revisiting useful approaches to data-rich macroeconomic forecasting. Computational Statistics & Data Analysis, 100(??):221– 239, August 2016. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947315002972.

Hayakawa:2016:IGE

[1946] Kazuhiko Hayakawa. Improved GMM estimation of panel VAR models. Computational Statistics & Data Analysis, 100(??):240– 264, August 2016. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947315001218.

Hayakawa:2016:BGE

[1947] Kazuhiko Hayakawa and Shuichi Nagata. On the behaviour of the GMM estimator in persistent dynamic panel data models with unrestricted initial conditions. Computational Statistics & Data Analysis, 100(??):265– 303, August 2016. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947315000754.

Hendrych:2016:CCM

[1948] R. Hendrych and T. Cipra. On conditional covariance modelling: an approach using state space models. Computational Statistics & Data Analysis, 100(??):304–317, August 2016. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0167947314002801.

Holzmann:2016:TNS

[1949] Hajo Holzmann and Florian Schwaiger.

Testing for the number of states in hidden Markov models. Computational Statistics & Data Analysis, 100 (??):318–330, August 2016. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947314001832.

Ishihara:2016:MES

[1950] Tsunehiro Ishihara, Yasuhiro Omori, and Manabu Asai. Matrix exponential stochastic volatility with cross leverage. Computational Statistics & Data Analysis, 100(??):331– 350, August 2016. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947314003041.

Jondeau:2016:ATD

[1951] Eric Jondeau. Asymmetry in tail dependence in equity portfolios. Computational Statistics & Data Analysis, 100 (??):351–368, August 2016. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947315000559.

Kapetanios:2016:FIG

[1952] George Kapetanios, Massimiliano Marcellino, and Fotis Papailias. Forecasting inflation and GDP growth using heuristic optimisation of information criteria and variable reduction methods. Computational Statistics & Data Analysis, 100(??):369– 382, August 2016. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947315000584.

Laurent:2016:TJC

[1953] Sébastien Laurent, Christelle Lecourt, and Franz C. Palm. Testing for jumps in conditionally Gaussian ARMA–GARCH models, a robust approach. Computational Statistics & Data Analysis, 100(??):383– 400, August 2016. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947314001650.

Li:2016:SGE

[1954] Rui Li, Alan T. K. Wan, and Jinhong You. Semiparametric GMM estimation and variable selection in dynamic panel data models with fixed effects. Computational Statistics & Data Analysis, 100 (??):401–423, August 2016. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947316300147.

Li:2016:GNS

[1955] Degui Li, Léopold Simar, and Valentin Zelenyuk. Generalized nonparametric smoothing with mixed discrete and continuous data. Computational Statistics & Data Analysis, 100(??): 424–444, August 2016. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947314001741.

Okou:2016:HET

[1956] Cédric Okou and Éric Jacquier. Horizon effect in the term structure of longrun risk-return trade-offs. Computational Statistics & Data Analysis, 100 (??):445–466, August 2016. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947314002084.

Pan:2016:BPI

[1957] Li Pan and Dimitris N. Politis. Bootstrap prediction intervals for Markov processes. Computational Statistics & Data Analysis, 100(??):467– 494, August 2016. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947315001371.

Panopoulou:2016:FEP

[1958] Ekaterini Panopoulou and Theologos Pantelidis. The Fisher effect in the presence of time-varying coefficients. Computational Statistics & Data Analysis, 100(??):495– 511, August 2016. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947314002473.

Pohlmeier:2016:SSS

[1959] Winfried Pohlmeier, Ruben Seiberlich, and Selver Derya Uysal. A simple and successful shrinkage method for weighting estimators of treatment effects. Computational Statistics & Data Analysis, 100(??):512– 525, August 2016. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016794731400276X.

Sarafidis:2016:NGE

[1960] Vasilis Sarafidis. Neighbourhood GMM estimation of dynamic panel data models. Computational Statistics & Data Analysis, 100(??):526– 544, August 2016. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947315002984.

Spierdijk:2016:CIA

[1961] Laura Spierdijk. Confidence intervals for ARMA–GARCH Value-at-Risk: the case of heavy tails and skewness. Computational Statistics & Data Analysis, 100(??):545– 559, August 2016. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947314002436.

Stojanovic:2016:SSM

[1962] Vladica S. Stojanović, Biljana C. Popović, and Gradimir V. Milovanović. The Split-SV model. Computational Statistics & Data Analysis, 100(??): 560–581, August 2016. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947314002424.

Sucarrat:2016:EIU

[1963] Genaro Sucarrat, Steffen Grønneberg, and Alvaro Escribano. Estimation and inference in univariate and multivariate log-GARCH-X models when the conditional density is unknown. Computational Statistics & Data Analysis, 100 (??):582–594, August 2016. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947315003059.

Vijverberg:2016:LTL

[1964] Chu-Ping C. Vijverberg, Wim P. M.

Vijverberg, and Süleyman Taspinar. Linking Tukey’s legacy to financial risk measurement. Computational Statistics & Data Analysis, 100(??): 595–615, August 2016. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947315002066.

Vosseler:2016:BMS

[1965] Alexander Vosseler. Bayesian model selection for unit root testing with multiple structural breaks. Computational Statistics & Data Analysis, 100 (??):616–630, August 2016. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947314002485.

Boswijk:2016:SIT

[1966] H. Peter Boswijk, Christian Francq, Marc Hallin, and Robert Taylor.

Special issue on time series econometrics. Computational Statistics & Data Analysis, 100(??):631– 632, August 2016. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947316300226.

Alj:2016:EGL

[1967] Abdelkamel Alj, Kristján Jónasson, and Guy Mélard. The exact Gaussian likelihood estimation of timedependent VARMA models. Computational Statistics & Data Analysis, 100 (??):633–644, August 2016. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947314002102.

Arteche:2016:BAD

[1968] Josu Arteche and Jesus Orbe. A bootstrap approximation for the distribution of the Local Whittle estimator. Computational Statistics & Data Analysis, 100(??):645– 660, August 2016. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947315000833.

Clements:2016:RTF

[1969] Michael P. Clements. Real-time factor model forecasting and the effects of instability. Computational Statistics & Data Analysis, 100(??):661– 675, August 2016. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947315000225.

Horvath:2016:ABS

[1970] Lajos Horváth, Gregory Rice, and Stephen Whipple. Adaptive bandwidth selection in the long run covariance estimator of functional time series. Computational Statistics & Data Analysis, 100(??):676– 693, August 2016. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947314001790.

Lin:2016:IVT

[1971] Wei Lin and Gloria González-Rivera.

Interval-valued time series models: Estimation based on order statistics exploring the Agriculture Marketing Service data. Computational Statistics & Data Analysis, 100(??):694– 711, August 2016. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947315001656.

McCulloch:2016:MRE

[1972] J. Huston McCulloch. Moment ratio estimation of autoregressive/unit root parameters and autocorrelationconsistent standard errors. Computational Statistics & Data Analysis, 100 (??):712–733, August 2016. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947315001607.

Phillips:2016:ARB

[1973] Garry D. A. Phillips and Gareth Liu-Evans. Approximating and reducing bias in 2SLS estimation of dynamic simultaneous equation models. Computational Statistics & Data Analysis, 100 (??):734–762, August 2016. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947315002947.

Shelef:2016:GBU

[1974] Amit Shelef. A Gini-based unit root test. Computational Statistics & Data Analysis, 100(??):763– 772, August 2016. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947314002448.

Ziel:2016:IRA

[1975] Florian Ziel. Iteratively reweighted adaptive lasso for conditional heteroscedastic time series with applications to AR–ARCH type processes. Computational Statistics & Data Analysis, 100(??):773–793, August 2016. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0167947315002996.

Bauwens:2016:SIB

[1976] Luc Bauwens, Gary Koop, John Maheu, and Yasuhiro Omori. Special issue on Bayesian econometrics. Computational Statistics & Data Analysis, 100(??):794, August 2016. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947316300238.

Kurose:2016:DES

[1977] Yuta Kurose and Yasuhiro Omori.

Dynamic equicorrelation stochastic volatility. Computational Statistics & Data Analysis, 100(??):795– 813, August 2016. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947315000249.

Virbickaite:2016:BNP

[1978] Audrone Virbickaite, M. Concepción Ausín, and Pedro Galeano. A Bayesian non-parametric approach to asymmetric dynamic conditional correlation model with application to portfolio selection. Computational Statistics & Data Analysis, 100(??):814– 829, August 2016. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947314003430.

Lubrano:2016:IID

[1979] Michel Lubrano and Abdoul Aziz Junior Ndoye. Income inequality decomposition using a finite mixture of log-normal distributions: a Bayesian approach. Computational Statistics & Data Analysis, 100(??):830– 846, August 2016. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947314003016.

Chan:2016:FCD

[1980] Joshua C. C. Chan and Angelia L.

Grant. Fast computation of the deviance information criterion for latent variable models. Computational Statistics & Data Analysis, 100(??): 847–859, August 2016. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947314002321.

Anonymous:2016:EBi

[1981] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 101 (??):iii–v, September 2016. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947316300998.

Anonymous:2016:Ci

[1982] Anonymous. Contents. Computational Statistics & Data Analysis, 101 (??):vi–vii, September 2016. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947316301001.

Ding:2016:EAG

[1983] Wei Ding and Peter X.-K. Song. EM algorithm in Gaussian copula with missing data. Computational Statistics & Data Analysis, 101(??): 1–11, September 2016. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947316000177.

Davies:2016:SAS

[1984] Tilman M. Davies, Khair Jones, and Martin L. Hazelton. Symmetric adaptive smoothing regimens for estimation of the spatial relative risk function. Computational Statistics & Data Analysis, 101(??):

12–28, September 2016. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794731630024X.

Hosseini:2016:IMM

[1985] Reshad Hosseini, Suvrit Sra, Lucas Theis, and Matthias Bethge. Inference and mixture modeling with the Elliptical Gamma Distribution. Computational Statistics & Data Analysis, 101 (??):29–43, September 2016. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947316300251.

McElroy:2016:CAT

[1986] Tucker S. McElroy and Scott H. Holan.

Computation of the autocovariances for time series with multiple longrange persistencies. Computational Statistics & Data Analysis, 101(??): 44–56, September 2016. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947316300202.

Lee:2016:CAQ

[1987] Minjung Lee and Junhee Han.

Covariate-adjusted quantile inference with competing risks. Computational Statistics & Data Analysis, 101(??): 57–63, September 2016. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947316300287.

Cipolli:2016:BNM

[1988] William Cipolli III, Timothy Hanson, and Alexander C. McLain. Bayesian nonparametric multiple testing. Computational Statistics & Data Analysis, 101(??):64–79, September 2016. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0167947316300421.

Kim:2016:CSS

[1989] Yongku Kim and L. Mark Berliner. Change of spatiotemporal scale in dynamic models. Computational Statistics & Data Analysis, 101(??):80– 92, September 2016. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947316300391.

Zhou:2016:DGG

[1990] Lixing Zhou, Yoshio Takane, and Heungsun Hwang. Dynamic GSCANO (Generalized Structured Canonical Correlation Analysis) with applications to the analysis of effective connectivity in functional neuroimaging data. Computational Statistics & Data Analysis, 101(??):93– 109, September 2016. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947316300433.

Korobilis:2016:PSP

[1991] Dimitris Korobilis. Prior selection for panel vector autoregressions. Computational Statistics & Data Analysis, 101(??):110–120, September 2016. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947316300275.

Wu:2016:BMS

[1992] Zizhen Wu and David B. Hitchcock.

A Bayesian method for simultaneous registration and clustering of functional observations. Computational Statistics & Data Analysis, 101(??): 121–136, September 2016. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947316300263.

Hu:2016:MLE

[1993] Hao Hu, Yichao Wu, and Weixin Yao. Maximum likelihood estimation of the mixture of log-concave densities. Computational Statistics & Data Analysis, 101(??):137– 147, September 2016. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947316300445.

OBrien:2016:FOM

[1994] Travis A. O’Brien, Karthik Kashinath, Nicholas R. Cavanaugh, William D. Collins, and John P. O’Brien. A fast and objective multidimensional kernel density estimation method: fastKDE. Computational Statistics & Data Analysis, 101(??):148– 160, September 2016. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947316300408.

Bedair:2016:MFM

[1995] Khaled Bedair, Yili Hong, Jie Li, and Hussein R. Al-Khalidi. Multivariate frailty models for multi-type recurrent event data and its application to cancer prevention trial. Computational Statistics & Data Analysis, 101(??): 161–173, September 2016. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947316300044.

Chen:2016:ULP

[1996] Qixuan Chen, Myunghee Cho Paik, Minjin Kim, and Cuiling Wang. Using link-preserving imputation for logistic partially linear models with missing covariates. Computational Statistics & Data Analysis, 101(??):174– 185, September 2016. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947316300469.

Haff:2016:SLB

[1997] Ingrid Hobæk Haff, Kjersti Aas, Arnoldo Frigessi, and Virginia Lacal. Structure learning in Bayesian networks using regular vines. Computational Statistics & Data Analysis, 101(??):186–208, September 2016. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0167947316300457.

Li:2016:RCF

[1998] Qi Li, Heng Lian, and Fukang Zhu. Robust closed-form estimators for the integer-valued GARCH(1, 1) model. Computational Statistics & Data Analysis, 101(??):209–225, September 2016. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947316300494.

Gross:2016:DSL

[1999] Samuel M. Gross and Robert Tibshirani. Data shared lasso: a novel tool to discover uplift. Computational Statistics & Data Analysis, 101(??): 226–235, September 2016. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794731630041X.

Hatjispyros:2016:RDF

[2000] Spyridon J. Hatjispyros, Theodoros Nicoleris, and Stephen G. Walker. Random density functions with common atoms and pairwise dependence. Computational Statistics & Data Analysis, 101(??):236–249, September 2016. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947316300512.

Lee:2016:BSE

[2001] Namgil Lee, Hyemi Choi, and Sung-Ho Kim. Bayes shrinkage estimation for high-dimensional VAR models with scale mixture of normal distributions for noise. Computational Statistics & Data Analysis, 101(??): 250–276, September 2016. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947316300500.

Kleiber:2016:HRS

[2002] William Kleiber. High resolution simulation of nonstationary Gaussian random fields. Computational Statistics & Data Analysis, 101(??):277– 288, September 2016. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947316300482.

Mandal:2016:RMI

[2003] B. N. Mandal and Jun Ma. l_{1} regularized multiplicative iterative path algorithm for non-negative generalized linear models. Computational Statistics & Data Analysis, 101(??): 289–299, September 2016. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947316300524.

Wilson:2016:APC

[2004] Huon Wilson and Uri Keich. Accurate pairwise convolutions of non-negative vectors via FFT. Computational Statistics & Data Analysis, 101(??):300– 315, September 2016. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947316300536.

Anonymous:2016:EBj

[2005] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 102(??):iii–v, October 2016. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947316301220.

Anonymous:2016:Cj

[2006] Anonymous. Contents. Computational Statistics & Data Analysis, 102 (??):vi, October 2016. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947316301232.

Bryner:2016:ELT

[2007] Darshan Bryner, Fred Huffer, Michael Rosenthal, J. Derek Tucker, and Anuj Srivastava. Estimation of linear target-layer trajectories using cluttered point cloud data. Computational Statistics & Data Analysis, 102 (??):1–22, October 2016. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794731630055X.

Nguyen:2016:MLE

[2008] Hien D. Nguyen and Geoffrey J.

McLachlan. Maximum likelihood estimation of triangular and polygonal distributions. Computational Statistics & Data Analysis, 102(??):23– 36, October 2016. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947316300561.

Costantini:2016:STP

[2009] Mauro Costantini and Amit Sen. A simple testing procedure for unit root and model specification. Computational Statistics & Data Analysis, 102 (??):37–54, October 2016. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947316300548.

Marques:2016:INE

[2010] Filipe J. Marques and Florence Loingeville. Improved near-exact distributions for the product of independent Generalized Gamma random variables. Computational Statistics & Data Analysis, 102(??):55– 66, October 2016. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947316300809.

Sreevani:2016:BSU

[2011] Sreevani and C. A. Murthy. On bandwidth selection using minimal spanning tree for kernel density estimation. Computational Statistics & Data Analysis, 102(??):67– 84, October 2016. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947316300810.

Xia:2016:FSG

[2012] Xiaochao Xia, Hu Yang, and Jialiang Li. Feature screening for generalized varying coefficient models with application to dichotomous responses. Computational Statistics & Data Analysis, 102(??):85–97, October 2016. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947316300846.

Feng:2016:NNC

[2013] Sanying Feng, Heng Lian, and Liugen Xue. A new nested Cholesky decomposition and estimation for the covariance matrix of bivariate longitudinal data. Computational Statistics & Data Analysis, 102(??):98– 109, October 2016. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947316300822.

Anonymous:2016:EBk

[2014] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 103(??):iii–v, November 2016. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947316301633.

Anonymous:2016:Ck

[2015] Anonymous. Contents. Computational Statistics & Data Analysis, 103 (??):vi–vii, November 2016. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947316301645.

Chauveau:2016:NMM

[2016] Didier Chauveau and Vy Thuy Lynh Hoang. Nonparametric mixture models with conditionally independent multivariate component densities. Computational Statistics & Data Analysis, 103 (??):1–16, November 2016. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947316300925.

Wu:2016:DMP

[2017] K. Y. K. Wu and W. K. Li. On a dispersion model with Pearson residual responses. Computational Statistics & Data Analysis, 103(??):17– 27, November 2016. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947316300949.

Zilko:2016:CMM

[2018] Aurelius A. Zilko and Dorota Kurowicka. Copula in a multivariate mixed discrete-continuous model. Computational Statistics & Data Analysis, 103 (??):28–55, November 2016. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947316300895.

Hazelton:2016:BSK

[2019] Martin L. Hazelton and Murray P. Cox. Bandwidth selection for kernel log-density estimation. Computational Statistics & Data Analysis, 103 (??):56–67, November 2016. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947316301050.

Alhamzawi:2016:BMS

[2020] Rahim Alhamzawi. Bayesian model selection in ordinal quantile regression. Computational Statistics & Data Analysis, 103(??):68– 78, November 2016. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947316300937.

Ma:2016:CRA

[2021] Ling Ma, Tao Hu, and Jianguo Sun. Cox regression analysis of dependent interval-censored failure time data. Computational Statistics & Data Analysis, 103(??):79– 90, November 2016. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947316300901.

Santos:2016:MLI

[2022] Vera Lúcia F. Santos, Fernando A. S.

Moura, Dalton F. Andrade, and Kelly C. M. Gonçalves. Multidimensional and longitudinal item response models for non-ignorable data. Computational Statistics & Data Analysis, 103 (??):91–110, November 2016. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947316301049.

Fang:2016:IIE

[2023] Fang Fang and Jun Shao. Iterated imputation estimation for generalized linear models with missing response and covariate values. Computational Statistics & Data Analysis, 103(??): 111–123, November 2016. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794731630086X.

Groenitz:2016:CNR

[2024] Heiko Groenitz. A covariate nonrandomized response model for multicategorical sensitive variables. Computational Statistics & Data Analysis, 103 (??):124–138, November 2016. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947316300834.

Feng:2016:RRR

[2025] Sanying Feng, Heng Lian, and Fukang Zhu. Reduced rank regression with possibly non-smooth criterion functions: an empirical likelihood approach. Computational Statistics & Data Analysis, 103(??):139– 150, November 2016. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947316300913.

Nguyen:2016:LMM

[2026] Hien D. Nguyen and Geoffrey J. McLachlan. Linear mixed models with marginally symmetric nonparametric random effects. Computational Statistics & Data Analysis, 103(??): 151–169, November 2016. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947316301074.

Qiu:2016:CIO

[2027] Shi-Fang Qiu, Wai-Yin Poon, and Man-Lai Tang. Confidence intervals for an ordinal effect size measure based on partially validated series. Computational Statistics & Data Analysis, 103 (??):170–192, November 2016. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947316301086.

Xue:2016:ESD

[2028] Yuan Xue, Xiangrong Yin, and Xiaolin Jiang. Ensemble sufficient dimension folding methods for analyzing matrix-valued data. Computational Statistics & Data Analysis, 103(??): 193–205, November 2016. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947316301037.

ElAssaad:2016:VEM

[2029] Hani El Assaad, Allou Samé, Gérard Govaert, and Patrice Aknin. A variational expectation-maximization algorithm for temporal data clustering. Computational Statistics & Data Analysis, 103(??):206–228, November 2016. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947316301098.

Rodrigues:2016:FRA

[2030] G. S. Rodrigues, David J. Nott, and S. A. Sisson. Functional regression approximate Bayesian computation for Gaussian process density estimation. Computational Statistics & Data Analysis, 103(??):229– 241, November 2016. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947316301116.

Chen:2016:MIA

[2031] Ling Chen, Jianguo Sun, and Chengjie Xiong. A multiple imputation approach to the analysis of clustered intervalcensored failure time data with the additive hazards model. Computational Statistics & Data Analysis, 103(??): 242–249, November 2016. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794731630113X.

Hao:2016:REB

[2032] Meiling Hao, Yunyuan Lin, and Xingqiu Zhao. A relative errorbased approach for variable selection. Computational Statistics & Data Analysis, 103(??):250–262, November 2016. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947316301153.

Zhu:2016:HTR

[2033] Xuehu Zhu, Fei Chen, Xu Guo, and Lixing Zhu. Heteroscedasticity testing for regression models: a dimension reduction-based model adaptive approach. Computational Statistics & Data Analysis, 103(??):263– 283, November 2016. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947316300858.

vanWieringen:2016:REI

[2034] Wessel N. van Wieringen and Carel F. W. Peeters. Ridge estimation of inverse covariance matrices from high-dimensional data. Computational Statistics & Data Analysis, 103(??): 284–303, November 2016. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947316301141.

Zimmer:2016:TLU

[2035] Zachary Zimmer, DoHwan Park, and Thomas Mathew. Tolerance limits under normal mixtures: Application to the evaluation of nuclear power plant safety and to the assessment of circular error probable. Computational Statistics & Data Analysis, 103(??): 304–315, November 2016. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947316301177.

He:2016:FAI

[2036] Heping He and Thomas A. Severini. A flexible approach to inference in semiparametric regression models with correlated errors using Gaussian processes. Computational Statistics & Data Analysis, 103(??): 316–329, November 2016. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947316301128.

Zhang:2016:ASE

[2037] Shibin Zhang. Adaptive spectral estimation for nonstationary multivariate time series. Computational Statistics & Data Analysis, 103(??):330– 349, November 2016. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947316301372.

Hunt:2016:CCC

[2038] Lynette A. Hunt and Kaye E. Basford. Comparing classical criteria for selecting intra-class correlated features in Multimix. Computational Statistics & Data Analysis, 103(??): 350–366, November 2016. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794731630130X.

Garbuno-Inigo:2016:GPH

[2039] A. Garbuno-Inigo, F. A. DiazDelaO, and K. M. Zuev. Gaussian process hyper-parameter estimation using Parallel Asymptotically Independent Markov Sampling. Computational Statistics & Data Analysis, 103(??): 367–383, November 2016. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947316301311.

Huang:2016:RSE

[2040] Lele Huang, Junlong Zhao, Huiwen Wang, and Siyang Wang. Robust shrinkage estimation and selection for functional multiple linear model through LAD loss. Computational Statistics & Data Analysis, 103(??): 384–400, November 2016. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947316301293.

Kwon:2016:URE

[2041] Sunghoon Kwon, Seungyoung Oh, and Youngjo Lee. The use of randomeffect models for high-dimensional variable selection problems. Computational Statistics & Data Analysis, 103(??): 401–412, November 2016. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947316301281.

Xiang:2016:SMC

[2042] Sijia Xiang, Weixin Yao, and Byungtae Seo. Semiparametric mixture: Continuous scale mixture approach. Computational Statistics & Data Analysis, 103 (??):413–425, November 2016. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947316301384.

Hook:2016:ECQ

[2043] Lars Josef Höök and Erik Lindström.

Efficient computation of the quasi likelihood function for discretely observed diffusion processes. Computational Statistics & Data Analysis, 103(??): 426–437, November 2016. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947316301165.

Anonymous:2016:EBl

[2044] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 104(??):iii–v, December 2016. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947316301979.

Anonymous:2016:Cl

[2045] Anonymous. Contents. Computational Statistics & Data Analysis, 104 (??):vi–vii, December 2016. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947316301980.

Lyubchich:2016:DFB

[2046] Vyacheslav Lyubchich, Xingyu Wang, Andrew Heyes, and Yulia R. Gel. A distribution-free m-out-of-n bootstrap approach to testing symmetry about an unknown median. Computational Statistics & Data Analysis, 104 (??):1–9, December 2016. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947316301062.

Lin:2016:PCH

[2047] Ruitao Lin, Zhongying Liu, Shurong Zheng, and Guosheng Yin. Power computation for hypothesis testing with high-dimensional covariance matrices. Computational Statistics & Data Analysis, 104(??):10– 23, December 2016. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947316301104.

Epifanio:2016:FAA

[2048] Irene Epifanio. Functional archetype and archetypoid analysis. Computational Statistics & Data Analysis, 104 (??):24–34, December 2016. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016794731630144X.

Singh:2016:BCD

[2049] Satya Prakash Singh and Siuli Mukhopadhyay. Bayesian crossover designs for generalized linear models. Computational Statistics & Data Analysis, 104(??):35–50, December 2016. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0167947316301396.

Gruenhage:2016:VEC

[2050] Gina Gruenhage, Manfred Opper, and Simon Barthelme. Visualizing the effects of a changing distance on data using continuous embeddings. Computational Statistics & Data Analysis, 104 (??):51–65, December 2016. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947316301438.

Pan:2016:IWG

[2051] Qing Pan and Yunpeng Zhao. Integrative weighted group lasso and generalized local quadratic approximation. Computational Statistics & Data Analysis, 104(??):66– 78, December 2016. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947316301414.

Ahfock:2016:PIS

[2052] Daniel Ahfock, Saumyadipta Pyne, Sharon X. Lee, and Geoffrey J. McLachlan. Partial identification in the statistical matching problem. Computational Statistics & Data Analysis, 104 (??):79–90, December 2016. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947316301426.

Li:2016:CDE

[2053] Shuo Li and Yundong Tu. \sqrt{n}- consistent density estimation in semiparametric regression models. Computational Statistics & Data Analysis, 104 (??):91–109, December 2016. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947316301505.

Marra:2016:SPC

[2054] Giampiero Marra and Karol Wyszynski. Semi-parametric copula sample selection models for count responses. Computational Statistics & Data Analysis, 104(??):110– 129, December 2016. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947316301402.

Barbeito:2016:SSB

[2055] Inés Barbeito and Ricardo Cao. Smoothed stationary bootstrap bandwidth selection for density estimation with dependent data. Computational Statistics & Data Analysis, 104(??): 130–147, December 2016. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947316301529.

Ameraoui:2016:BET

[2056] Abdelkader Ameraoui, Kamal Boukhetala, and Jean-François Dupuy. Bayesian estimation of the tail index of a heavy tailed distribution under random censoring. Computational Statistics & Data Analysis, 104(??):148– 168, December 2016. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947316301463.

Ippel:2016:ERI

[2057] L. Ippel, M. C. Kaptein, and J. K.

Vermunt. Estimating random-intercept models on data streams. Computational Statistics & Data Analysis, 104 (??):169–182, December 2016. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947316301451.

Le:2016:UBS

[2058] Tri Le and Bertrand Clarke. Using the Bayesian Shtarkov solution for predictions. Computational Statistics & Data Analysis, 104(??):183– 196, December 2016. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947316301554.

Li:2016:CSH

[2059] Chenxi Li. Cause-specific hazard regression for competing risks data under interval censoring and left truncation. Computational Statistics & Data Analysis, 104(??):197– 208, December 2016. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794731630158X.

Atkinson:2016:RMH

[2060] Anthony C. Atkinson, Marco Riani, and Francesca Torti. Robust methods for heteroskedastic regression. Computational Statistics & Data Analysis, 104 (??):209–222, December 2016. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947316301578.

Lu:2016:MCT

[2061] Tong-Yu Lu, Wai-Yin Poon, and Siu Hung Cheung. Multiple comparisons of treatments with skewed ordinal responses. Computational Statistics & Data Analysis, 104(??):223– 232, December 2016. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947316301682.

Li:2016:MNC

[2062] Yang Li and Zhengyuan Zhu. Modeling nonstationary covariance function with convolution on sphere. Computational Statistics & Data Analysis, 104 (??):233–246, December 2016. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947316301566.

Dehghan:2016:EPS

[2063] Mohammad Hossein Dehghan and Thierry Duchesne. Erratum to “On the performance of some non-parametric estimators of the conditional survival function with interval-censored data” [Comput. Statstic Data Anal. 55 (12) (2011) 3355–3364]. Computational Statistics & Data Analysis, 104(??): 247, December 2016. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947316301487. See

[607].

Anonymous:2017:EBa

[2064] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 105(??):iii–v, January 2017. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947316302195.

Anonymous:2017:Ca

[2065] Anonymous. Contents. Computational Statistics & Data Analysis, 105 (??):vi–vii, January 2017. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947316302201.

Hui:2017:MBS

[2066] Francis K. C. Hui. Model-based simultaneous clustering and ordination of multivariate abundance data in ecology. Computational Statistics & Data Analysis, 105(??):1– 10, January 2017. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947316301724.

Wilhelm:2017:QSS

[2067] Matthieu Wilhelm, Yves Tillé, and Lionel Qualité. Quasi-systematic sampling from a continuous population. Computational Statistics & Data Analysis, 105(??):11– 23, January 2017. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794731630175X.

Serfling:2017:DBN

[2068] Robert Serfling and Uditha Wijesuriya. Depth-based nonparametric description of functional data, with emphasis on use of spatial depth. Computational Statistics & Data Analysis, 105 (??):24–45, January 2017. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947316301712.

Gertheiss:2017:NMS

[2069] Jan Gertheiss, Jeff Goldsmith, and Ana-Maria Staicu. A note on modeling sparse exponential-family functional response curves. Computational Statistics & Data Analysis, 105 (??):46–52, January 2017. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947316301748.

Xiao:2017:FAT

[2070] Yuanhui Xiao. A fast algorithm for two-dimensional Kolmogorov–Smirnov two sample tests. Computational Statistics & Data Analysis, 105(??): 53–58, January 2017. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947316301785.

Schmidt:2017:FLS

[2071] Paul Schmidt, Mark Mühlau, and Volker Schmid. Fitting large-scale structured additive regression models using Krylov subspace methods. Computational Statistics & Data Analysis, 105(??):59–75, January 2017. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947316301700.

Nicosia:2017:GHS

[2072] Aurélien Nicosia, Thierry Duchesne, Louis-Paul Rivest, and Daniel Fortin. A general hidden state random walk model for animal movement. Computational Statistics & Data Analysis, 105(??):76–95, January 2017. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947316301736.

Kim:2017:NGB

[2073] Seongho Kim, Hyejeong Jang, Imhoi Koo, Joohyoung Lee, and Xiang Zhang. Normal-gamma–Bernoulli peak detection for analysis of comprehensive two-dimensional gas chromatography mass spectrometry data. Computational Statistics & Data Analysis, 105 (??):96–111, January 2017. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947316301797.

Silva:2017:CIT

[2074] Ivair R. Silva. Confidence intervals through sequential Monte Carlo. Computational Statistics & Data Analysis, 105(??):112–124, January 2017.

CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0167947316301815.

Wang:2017:AOD

[2075] WenWu Wang and Ping Yu. Asymptotically optimal differenced estimators of error variance in nonparametric regression. Computational Statistics & Data Analysis, 105(??):125– 143, January 2017. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947316301761.

Lopez-Cheda:2017:NIE

[2076] Ana López-Cheda, Ricardo Cao, M. Amalia Jácome, and Ingrid Van Keilegom. Nonparametric incidence estimation and bootstrap bandwidth selection in mixture cure models. Computational Statistics & Data Analysis, 105 (??):144–165, January 2017. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947316301839.

Picchini:2017:AML

[2077] Umberto Picchini and Rachele Anderson. Approximate maximum likelihood estimation using datacloning ABC. Computational Statistics & Data Analysis, 105(??):166– 183, January 2017. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947316301876.

Zhang:2017:BLI

[2078] Yan-Qing Zhang and Nian-Sheng Tang.

Bayesian local influence analysis of general estimating equations with nonignorable missing data. Computational Statistics & Data Analysis, 105 (??):184–200, January 2017. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947316301918.

Lai:2017:MFF

[2079] Peng Lai, Yiming Liu, Zhi Liu, and Yi Wan. Model free feature screening for ultrahigh dimensional data with responses missing at random. Computational Statistics & Data Analysis, 105 (??):201–216, January 2017. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794731630189X.

Arias-Castro:2017:SAS

[2080] Ery Arias-Castro and Xiao Pu. A simple approach to sparse clustering. Computational Statistics & Data Analysis, 105(??):217–228, January 2017. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0167947316301840.

Kiatsupaibul:2017:RCD

[2081] Seksan Kiatsupaibul, Anthony J.

Hayter, and Wei Liu. Rank constrained distribution and moment computations. Computational Statistics & Data Analysis, 105(??):229– 242, January 2017. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947316301906.

Song:2017:RES

[2082] Junmo Song, Dong hyun Oh, and Jiwon Kang. Robust estimation in stochastic frontier models. Computational Statistics & Data Analysis, 105 (??):243–267, January 2017. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947316301864.

Lombard:2017:SRC

[2083] F. Lombard, Douglas M. Hawkins, and Cornelis J. Potgieter. Sequential rank CUSUM charts for angular data. Computational Statistics & Data Analysis, 105(??):268– 279, January 2017. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947316301827.

Persson:2017:DDA

[2084] Emma Persson, Jenny Häggström, Ingeborg Waernbaum, and Xavier de Luna. Data-driven algorithms for dimension reduction in causal inference. Computational Statistics & Data Analysis, 105(??):280– 292, January 2017. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947316302018.

Anonymous:2017:EBb

[2085] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 106(??):iii–v, February 2017. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947316302547.

Anonymous:2017:Cb

[2086] Anonymous. Contents. Computational Statistics & Data Analysis, 106 (??):vi, February 2017. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947316302559.

Kwon:2017:GEE

[2087] Yongchan Kwon, Young-Geun Choi, Taesung Park, Andreas Ziegler, and Myunghee Cho Paik. Generalized estimating equations with stabilized working correlation structure. Computational Statistics & Data Analysis, 106 (??):1–11, February 2017. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947316302079.

Wang:2017:PMC

[2088] Dan Wang and Lili Tian. Parametric methods for confidence interval estimation of overlap coefficients. Computational Statistics & Data Analysis, 106 (??):12–26, February 2017. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794731630202X.

Gramacki:2017:FBF

[2089] Artur Gramacki and Jaroslaw Gramacki. FFT-based fast bandwidth selector for multivariate kernel density estimation. Computational Statistics & Data Analysis, 106(??):27– 45, February 2017. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947316302092.

Bianco:2017:REP

[2090] Ana M. Bianco and Paula M. Spano. Robust estimation in partially linear errors-in-variables models. Computational Statistics & Data Analysis, 106 (??):46–64, February 2017. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947316302109.

Schaarschmidt:2017:SCI

[2091] Frank Schaarschmidt, Daniel Gerhard, and Charlotte Vogel. Simultaneous confidence intervals for comparisons of several multinomial samples. Computational Statistics & Data Analysis, 106 (??):65–76, February 2017. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947316302122.

Li:2017:EAB

[2092] J. Li, D. J. Nott, Y. Fan, and S. A. Sisson. Extending approximate Bayesian computation methods to high dimensions via a Gaussian copula model. Computational Statistics & Data Analysis, 106(??): 77–89, February 2017. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947316301694.

Alvarez-Iglesias:2017:APB

[2093] Alberto Alvarez-Iglesias, John Hinde, John Ferguson, and John Newell. An alternative pruning based approach to unbiased recursive partitioning. Computational Statistics & Data Analysis, 106(??):90– 102, February 2017. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794731630192X.

Baek:2017:SSP

[2094] Changryong Baek, Richard A. Davis, and Vladas Pipiras. Sparse seasonal and periodic vector autoregressive modeling. Computational Statistics & Data Analysis, 106(??):103– 126, February 2017. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947316302134.

McGinnity:2017:CVW

[2095] K. McGinnity, R. Varbanov, and E. Chicken. Cross-validated wavelet block thresholding for non-Gaussian errors. Computational Statistics & Data Analysis, 106(??):127–137, February 2017. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947316302262.

Panagiotelis:2017:MSD

[2096] Anastasios Panagiotelis, Claudia Czado, Harry Joe, and Jakob Stöber. Model selection for discrete regular vine copulas. Computational Statistics & Data Analysis, 106(??):138– 152, February 2017. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947316302237.

Liu:2017:EFL

[2097] Baisen Liu, Liangliang Wang, and Jiguo Cao. Estimating functional linear mixed-effects regression models. Computational Statistics & Data Analysis, 106(??):153–164, February 2017. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0167947316302250.

Anonymous:2017:EBc

[2098] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 107(??):iii–v, March 2017. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016794731630278X.

Anonymous:2017:Cc

[2099] Anonymous. Contents. Computational Statistics & Data Analysis, 107 (??):vi, March 2017. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947316302791.

Berry:2017:DEM

[2100] Tyrus Berry and Timothy Sauer. Density estimation on manifolds with boundary. Computational Statistics & Data Analysis, 107(??):1– 17, March 2017. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947316302274.

Giorno:2017:ENH

[2101] Virginia Giorno, Patricia Román-Román, Serena Spina, and Francisco Torres-Ruiz. Estimating a non-homogeneous Gompertz process with jumps as model of tumor dynamics. Computational Statistics & Data Analysis, 107(??):18– 31, March 2017. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947316302328.

Boubeta:2017:PMM

[2102] Miguel Boubeta, María José Lombardía, and Domingo Morales. Poisson mixed models for studying the poverty in small areas. Computational Statistics & Data Analysis, 107 (??):32–47, March 2017. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947316302419.

Rocha:2017:NCD

[2103] Ricardo Rocha, Saralees Nadarajah, Vera Tomazella, and Francisco Louzada. A new class of defective models based on the Marshall– Olkin family of distributions for cure rate modeling. Computational Statistics & Data Analysis, 107(??):48– 63, March 2017. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947316302286.

Cui:2017:IEG

[2104] Xia Cui, Jianhua Guo, and Guangren Yang. On the identifiability and estimation of generalized linear models with parametric nonignorable missing data mechanism. Computational Statistics & Data Analysis, 107(??):64–80, March 2017. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947316302444.

Doove:2017:DOD

[2105] Lisa L. Doove, Tom F. Wilderjans, Antonio Calcagnì, and Iven Van Mechelen. Deriving optimal data-analytic regimes from benchmarking studies. Computational Statistics & Data Analysis, 107(??):81–91, March 2017. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947316302432.

Agostinelli:2017:REA

[2106] Claudio Agostinelli, Isabella Locatelli, Alfio Marazzi, and Víctor J. Yohai. Robust estimators of accelerated failure time regression with generalized log-gamma errors. Computational Statistics & Data Analysis, 107(??): 92–106, March 2017. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947316302390.

Li:2017:VSU

[2107] Hanning Li and Debdeep Pati. Variable selection using shrinkage priors. Computational Statistics & Data Analysis, 107(??):107–119, March 2017. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947316302353.

Deng:2017:SRA

[2108] Shirong Deng, Kin yat Liu, and Xingqiu Zhao. Semiparametric regression analysis of multivariate longitudinal data with informative observation times. Computational Statistics & Data Analysis, 107(??):120– 130, March 2017. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794731630233X.

Tao:2017:CKD

[2109] Chenyang Tao and Jianfeng Feng.

Canonical kernel dimension reduction. Computational Statistics & Data Analysis, 107(??):131–148, March 2017. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0167947316302304.

Krivitsky:2017:UCD

[2110] Pavel N. Krivitsky. Using contrastive divergence to seed Monte Carlo MLE for exponential-family random graph models. Computational Statistics & Data Analysis, 107(??): 149–161, March 2017. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947316302420.

Aeberhard:2017:STA

[2111] William H. Aeberhard, Eva Cantoni, and Stephane Heritier. Saddlepoint tests for accurate and robust inference on overdispersed count data. Computational Statistics & Data Analysis, 107 (??):162–175, March 2017. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947316302365.

Anonymous:2017:EBd

[2112] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 108(??):iii–v, April 2017. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947316303036.

Anonymous:2017:Cd

[2113] Anonymous. Contents. Computational Statistics & Data Analysis, 108 (??):vi, April 2017. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947316303048.

Maurya:2017:NTS

[2114] Vishal Maurya, A. N. Gill, and Aarti Goyal. A new two-stage multiple comparison procedure for comparing several exponential populations with a control under heteroscedasticity. Computational Statistics & Data Analysis, 108(??):1–11, April 2017. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947316302481.

Kundu:2017:ALT

[2115] Debasis Kundu, Debanjan Mitra, and Ayon Ganguly. Analysis of left truncated and right censored competing risks data. Computational Statistics & Data Analysis, 108(??): 12–26, April 2017. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794731630247X.

Seo:2017:DSM

[2116] Byungtae Seo. The doubly smoothed maximum likelihood estimation for location-shifted semiparametric mixtures. Computational Statistics & Data Analysis, 108(??):27–39, April 2017. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0167947316302675.

Monbet:2017:SVM

[2117] Valérie Monbet and Pierre Ailliot. Sparse vector Markov switching autoregressive models. Application to multivariate time series of temperature. Computational Statistics & Data Analysis, 108(??):40– 51, April 2017. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947316302584.

Wang:2017:TCD

[2118] Li-Yu Wang, Cheolwoo Park, Kyupil Yeon, and Hosik Choi. Tracking concept drift using a constrained penalized regression combiner. Computational Statistics & Data Analysis, 108(??):52–69, April 2017. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947316302663.

Yang:2017:APS

[2119] Lianqiang Yang and Yongmiao Hong. Adaptive penalized splines for data smoothing. Computational Statistics & Data Analysis, 108(??):70– 83, April 2017. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947316302493.

Bee:2017:AML

[2120] Marco Bee, Roberto Benedetti, and Giuseppe Espa. Approximate maximum likelihood estimation of the Bingham distribution. Computational Statistics & Data Analysis, 108(??): 84–96, April 2017. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947316302687.

Park:2017:NED

[2121] Byeong U. Park, Léopold Simar, and Valentin Zelenyuk. Nonparametric estimation of dynamic discrete choice models for time series data. Computational Statistics & Data Analysis, 108(??):97–120, April 2017. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947316302596.

Zhang:2017:CRS

[2122] Jing Zhang, Yanyan Liu, and Yuanshan Wu. Correlation rank screening for ultrahigh-dimensional survival data. Computational Statistics & Data Analysis, 108(??):121– 132, April 2017. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947316302699.

Kaffine:2017:MRD

[2123] Daniel T. Kaffine and Graham A.

Davis. A multi-row deletion diagnostic for influential observations in small-sample regressions. Computational Statistics & Data Analysis, 108 (??):133–145, April 2017. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947316302341.

Rekabdarkolaee:2017:REV

[2124] Hossein Moradi Rekabdarkolaee, Edward Boone, and Qin Wang. Robust estimation and variable selection in sufficient dimension reduction. Computational Statistics & Data Analysis, 108(??):146–157, April 2017. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947316302717.

Anonymous:2017:EBe

[2125] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 109(??):iii–v, May 2017. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947317300117.

Anonymous:2017:Ce

[2126] Anonymous. Contents. Computational Statistics & Data Analysis, 109 (??):vi–vii, May 2017. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947317300129.

Hilafu:2017:PPR

[2127] Haileab Hilafu and Wenbo Wu. Partial projective resampling method for dimension reduction: With applications to partially linear models. Computational Statistics & Data Analysis, 109(??):1–14, May 2017. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947316302924.

Zhao:2017:CQR

[2128] Weihua Zhao, Heng Lian, and Xinyuan Song. Composite quantile regression for correlated data. Computational Statistics & Data Analysis, 109(??):15–33, May 2017. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947316302894.

Chee:2017:MMB

[2129] Chew-Seng Chee. A mixture modelbased nonparametric approach to estimating a count distribution. Computational Statistics & Data Analysis, 109(??):34–44, May 2017. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947316302869.

Pan:2017:ESP

[2130] Jia-Chiun Pan, Yufen Huang, and J. T. Gene Hwang. Estimation of selected parameters. Computational Statistics & Data Analysis, 109(??): 45–63, May 2017. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947316302651.

Maronna:2017:REE

[2131] Ricardo A. Maronna and Victor J.

Yohai. Robust and efficient estimation of multivariate scatter and location. Computational Statistics & Data Analysis, 109(??):64– 75, May 2017. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947316302705.

Fujita:2017:CBG

[2132] André Fujita, Daniel Yasumasa Takahashi, Joana Bisol Balardin, Maciel Calebe Vidal, and João Ricardo Sato. Correlation between graphs with an application to brain network analysis. Computational Statistics & Data Analysis, 109(??): 76–92, May 2017. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947316302900.

Han:2017:ABN

[2133] Shengtong Han, Hongmei Zhang, Wilfried Karmaus, Graham Roberts, and Hasan Arshad. Adjusting background noise in cluster analyses of longitudinal data. Computational Statistics & Data Analysis, 109(??): 93–104, May 2017. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947316302730.

Zhu:2017:IBT

[2134] Xuehu Zhu, Tao Wang, Junlong Zhao, and Lixing Zhu. Inference for biased transformation models. Computational Statistics & Data Analysis, 109(??):105–120, May 2017. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947316302729.

Das:2017:BQR

[2135] Priyam Das and Subhashis Ghosal. Bayesian quantile regression using random B-spline series prior. Computational Statistics & Data Analysis, 109(??):121–143, May 2017. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947316302882.

Han:2017:BRP

[2136] Ningning Han, Yumeng Song, and Zhanjie Song. Bayesian robust principal component analysis with structured sparse component. Computational Statistics & Data Analysis, 109 (??):144–158, May 2017. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794731630295X.

Farooq:2017:SLA

[2137] Muhammad Farooq and Ingo Steinwart. An SVM-like approach for expectile regression. Computational Statistics & Data Analysis, 109(??): 159–181, May 2017. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947316302821.

Richardson:2017:FID

[2138] Robert Richardson, Athanasios Kottas, and Bruno Sansó. Flexible integro-difference equation modeling for spatio-temporal data. Computational Statistics & Data Analysis, 109 (??):182–198, May 2017. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947316302833.

Lian:2017:OBA

[2139] Yongqiang Lian, Yincai Tang, and Yijun Wang. Objective Bayesian analysis of JM model in software reliability. Computational Statistics & Data Analysis, 109(??):199– 214, May 2017. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947316302961.

Couellan:2017:USL

[2140] Nicolas Couellan and Wenjuan Wang.

Uncertainty-safe large scale support vector machines. Computational Statistics & Data Analysis, 109(??): 215–230, May 2017. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947316302985.

Bedard:2017:HML

[2141] Mylène Bédard. Hierarchical models: Local proposal variances for RWMwithin-Gibbs and MALA-within-Gibbs. Computational Statistics & Data Analysis, 109(??):231–246, May 2017. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0167947316302973.

Anonymous:2017:EBf

[2142] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 110(??):iii–v, June 2017. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947317300348.

Anonymous:2017:Cf

[2143] Anonymous. Contents. Computational Statistics & Data Analysis, 110 (??):vi, June 2017. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016794731730035X.

Kraus:2017:VCB

[2144] Daniel Kraus and Claudia Czado. Dvine copula based quantile regression. Computational Statistics & Data Analysis, 110(??):1–18, June 2017. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947316303073.

Crane:2017:HMM

[2145] Harry Crane. A hidden Markov model for latent temporal clustering with application to ideological alignment in the U.S. Supreme Court. Computational Statistics & Data Analysis, 110(??):19–36, June 2017. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947316303085.

Huang:2017:PPC

[2146] Zhendong Huang, Davide Ferrari, and Guoqi Qian. Parsimonious and powerful composite likelihood testing for group difference and genotypephenotype association. Computational Statistics & Data Analysis, 110(??): 37–49, June 2017. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947316302948.

Lin:2017:ETV

[2147] Hongmei Lin, Riquan Zhang, Wenchao Xu, and Yuedong Wang. Estimating time-varying treatment switching effects via local linear smoothing and quasi-likelihood. Computational Statistics & Data Analysis, 110(??):50–63, June 2017. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947316303127.

Doksum:2017:NNE

[2148] Kjell A. Doksum, Jiancheng Jiang, Bo Sun, and Shuzhen Wang. Nearest neighbor estimates of regression. Computational Statistics & Data Analysis, 110(??):64–74, June 2017. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947317300014.

Li:2017:RAC

[2149] Shuwei Li, Tao Hu, Peijie Wang, and Jianguo Sun. Regression analysis of current status data in the presence of dependent censoring with applications to tumorigenicity experiments. Computational Statistics & Data Analysis, 110(??):75–86, June 2017. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947316303115.

Ranalli:2017:MMM

[2150] Monia Ranalli and Roberto Rocci. Mixture models for mixed-type data through a composite likelihood approach. Computational Statistics & Data Analysis, 110(??):87–102, June 2017. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0167947317300038.

Geronimi:2017:VSM

[2151] J. Geronimi and G. Saporta. Variable selection for multiply-imputed data with penalized generalized estimating equations. Computational Statistics & Data Analysis, 110(??): 103–114, June 2017. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947317300051.

Mallick:2017:BGB

[2152] Himel Mallick and Nengjun Yi.

Bayesian group bridge for bi-level variable selection. Computational Statistics & Data Analysis, 110(??): 115–133, June 2017. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947317300063.

Liu:2017:PCA

[2153] Zhongkai Liu, Rui Song, Donglin Zeng, and Jiajia Zhang. Principal components adjusted variable screening. Computational Statistics & Data Analysis, 110(??):134– 144, June 2017. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947317300026.

Dlugosz:2017:GPL

[2154] Stephan Dlugosz, Enno Mammen, and Ralf A. Wilke. Generalized partially linear regression with misclassified data and an application to labour market transitions. Computational Statistics & Data Analysis, 110(??): 145–159, June 2017. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947317300166.

Anonymous:2017:EBg

[2155] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 111(??):iii–v, July 2017. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016794731730066X.

Anonymous:2017:Cg

[2156] Anonymous. Contents. Computational Statistics & Data Analysis, 111 (??):vi–vii, July 2017. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947317300671.

Gong:2017:RIC

[2157] Joonho Gong and Hyunjoong Kim. RHSBoost: Improving classification performance in imbalance data. Computational Statistics & Data Analysis, 111(??):1–13, July 2017. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016794731730018X.

Hu:2017:RET

[2158] Hao Hu, Weixin Yao, and Yichao Wu. The robust EM-type algorithms for log-concave mixtures of regression models. Computational Statistics & Data Analysis, 111(??): 14–26, July 2017. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947317300178.

Maire:2017:OEF

[2159] Florian Maire, Eric Moulines, and Sidonie Lefebvre. Online EM for functional data. Computational Statistics & Data Analysis, 111(??):

27–47, July 2017. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947317300191.

Shin:2017:PPL

[2160] Seung Jun Shin and Andreas Artemiou.

Penalized principal logistic regression for sparse sufficient dimension reduction. Computational Statistics & Data Analysis, 111(??):48– 58, July 2017. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947316302936.

Leung:2017:MLS

[2161] Andy Leung, Victor Yohai, and Ruben Zamar. Multivariate location and scatter matrix estimation under cellwise and casewise contamination. Computational Statistics & Data Analysis, 111(??):59–76, July 2017. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947317300270.

Qian:2017:SWE

[2162] Lianfen Qian and Suojin Wang.

Subject-wise empirical likelihood inference in partial linear models for longitudinal data. Computational Statistics & Data Analysis, 111(??): 77–87, July 2017. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794731730021X.

Zhu:2017:MFR

[2163] Hongxiao Zhu, Jeffrey S. Morris, Fengrong Wei, and Dennis D. Cox. Multivariate functional response regression, with application to fluorescence spectroscopy in a cervical pre-cancer study. Computational Statistics & Data Analysis, 111(??): 88–101, July 2017. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947317300245.

Ouyang:2017:BLI

[2164] Ming Ouyang, Xiaodong Yan, Ji Chen, Niansheng Tang, and Xinyuan Song. Bayesian local influence of semiparametric structural equation models. Computational Statistics & Data Analysis, 111(??):102–115, July 2017. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0167947317300208.

Smucler:2017:RSE

[2165] Ezequiel Smucler and Victor J. Yohai. Robust and sparse estimators for linear regression models. Computational Statistics & Data Analysis, 111(??): 116–130, July 2017. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947317300221.

Al-Najjar:2017:SDR

[2166] Elias Al-Najjar and Kofi P. Adragni. Sufficient dimension reduction constrained through sub-populations. Computational Statistics & Data Analysis, 111(??):131–144, July 2017.

CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0167947317300282.

Kouritzin:2017:RSB

[2167] Michael A. Kouritzin. Residual and stratified branching particle filters. Computational Statistics & Data Analysis, 111(??):145– 165, July 2017. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947317300233.

Belalia:2017:SCD

[2168] Mohamed Belalia, Taoufik Bouezmarni, and Alexandre Leblanc. Smooth conditional distribution estimators using Bernstein polynomials. Computational Statistics & Data Analysis, 111(??):166–182, July 2017. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947317300257.

Alam:2017:PGO

[2169] M. Iftakhar Alam, Barbara Bogacka, and D. Stephen Coad. Pharmacokinetically guided optimum adaptive dose selection in early phase clinical trials. Computational Statistics & Data Analysis, 111(??):183– 202, July 2017. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947317300294.

Ueckert:2017:NME

[2170] Sebastian Ueckert and France Mentré. A new method for evaluation of the Fisher information matrix for discrete mixed effect models using Monte Carlo sampling and adaptive Gaussian quadrature. Computational Statistics & Data Analysis, 111(??): 203–219, July 2017. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947316302389.

Anonymous:2017:EBh

[2171] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 112(??):iii–v, August 2017. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947317300853.

Anonymous:2017:Ch

[2172] Anonymous. Contents. Computational Statistics & Data Analysis, 112 (??):vi–vii, August 2017. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947317300865.

Lee:2017:VSH

[2173] Kyu Ha Lee, Sounak Chakraborty, and Jianguo Sun. Variable selection for high-dimensional genomic data with censored outcomes using group lasso prior. Computational Statistics & Data Analysis, 112(??): 1–13, August 2017. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947317300440.

Bermudez:2017:FMM

[2174] Lluís Bermúdez, Dimitris Karlis, and Miguel Santolino. A finite mixture of multiple discrete distributions for modelling heaped count data. Computational Statistics & Data Analysis, 112(??):14–23, August 2017. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947317300439.

Kojadinovic:2017:SCI

[2175] Ivan Kojadinovic. Some copula inference procedures adapted to the presence of ties. Computational Statistics & Data Analysis, 112(??):24– 41, August 2017. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947317300269.

Wang:2017:ODO

[2176] Xue Wang and Stephen G. Walker.

An optimal data ordering scheme for Dirichlet process mixture models. Computational Statistics & Data Analysis, 112(??):42–52, August 2017. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0167947317300385.

Fang:2017:BMR

[2177] Tian Fang, William Mackillop, Wenyu Jiang, Allan Hildesheim, Sholom Wacholder, and Bingshu E. Chen. A Bayesian method for risk window estimation with application to HPV vaccine trial. Computational Statistics & Data Analysis, 112(??):53– 62, August 2017. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947317300397.

Nancy:2017:IMV

[2178] Jane Y. Nancy, Nehemiah H. Khanna, and Kannan Arputharaj. Imputing missing values in unevenly spaced clinical time series data to build an effective temporal classification framework. Computational Statistics & Data Analysis, 112(??):63– 79, August 2017. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947317300403.

Garcia-Escudero:2017:FPN

[2179] Luis A. García-Escudero, Agustín Mayo-Iscar, and Clara I. Sánchez-Gutiérrez. Fitting parabolas in noisy images. Computational Statistics & Data Analysis, 112(??):80– 87, August 2017. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794731730052X.

Goessling:2017:LAN

[2180] Marc Goessling. LogitBoost autoregressive networks. Computational Statistics & Data Analysis, 112(??): 88–98, August 2017. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947317300543.

Marra:2017:BCA

[2181] Giampiero Marra and Rosalba Radice.

Bivariate copula additive models for location, scale and shape. Computational Statistics & Data Analysis, 112 (??):99–113, August 2017. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947317300488.

Zhang:2017:EHT

[2182] Jun Zhang, Yan Zhou, Bingqing Lin, and Yao Yu. Estimation and hypothesis test on partial linear models with additive distortion measurement errors. Computational Statistics & Data Analysis, 112(??):114– 128, August 2017. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947317300531.

Lv:2017:MAC

[2183] Jing Lv, Chaohui Guo, Hu Yang, and Yalian Li. A moving average Cholesky factor model in covariance modeling for composite quantile regression with longitudinal data. Computational Statistics & Data Analysis, 112 (??):129–144, August 2017. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947317300567.

Jeong:2017:ABL

[2184] Seonghyun Jeong, Minjae Park, and Taeyoung Park. Analysis of binary longitudinal data with timevarying effects. Computational Statistics & Data Analysis, 112(??):145– 153, August 2017. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947317300518.

Shang:2017:GME

[2185] Wenpeng Shang and Xiao Wang. The generalized moment estimation of the additive-multiplicative hazard model with auxiliary survival information. Computational Statistics & Data Analysis, 112(??):154– 169, August 2017. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947317300580.

Chapple:2017:BVS

[2186] Andrew G. Chapple, Marina Vannucci, Peter F. Thall, and Steven Lin. Bayesian variable selection for a semi-competing risks model with three hazard functions. Computational Statistics & Data Analysis, 112 (??):170–185, August 2017. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947317300464.

He:2017:AFC

[2187] Haijin He, Deng Pan, Liuquan Sun, Yimei Li, Leslie L. Robison, and Xinyuan Song. Analysis of a fixed center effect additive rates model for recurrent event data. Computational Statistics & Data Analysis, 112(??): 186–197, August 2017. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947317300476.

Carfora:2017:QCS

[2188] M. F. Carfora, L. Cutillo, and A. Orlando. A quantitative comparison of stochastic mortality models on Italian population data. Computational Statistics & Data Analysis, 112(??): 198–214, August 2017. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947317300579.

Lai:2017:CRM

[2189] Xin Lai, Kelvin K. W. Yau, and Liu Liu. Competing risk model with bivariate random effects for clustered survival data. Computational Statistics & Data Analysis, 112(??):215– 223, August 2017. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947317300555.

Chen:2017:SML

[2190] Xuerong Chen, Tao Hu, and Jianguo Sun. Sieve maximum likelihood estimation for the proportional hazards model under informative censoring. Computational Statistics & Data Analysis, 112 (??):224–234, August 2017. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947317300506.

Xia:2017:TCR

[2191] Qiang Xia, Rubing Liang, and Jianhong Wu. Transformed contribution ratio test for the number of factors in static approximate factor models. Computational Statistics & Data Analysis, 112(??):235– 241, August 2017. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794731730049X.

Lansangan:2017:SDR

[2192] Joseph Ryan G. Lansangan and Erniel B. Barrios. Simultaneous dimension reduction and variable selection in modeling high dimensional data. Computational Statistics & Data Analysis, 112(??):242– 256, August 2017. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947317300609.

Zamanzade:2017:EPP

[2193] Ehsan Zamanzade and Xinlei Wang. Estimation of population proportion for judgment post-stratification. Computational Statistics & Data Analysis, 112(??):257–269, August 2017. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0167947317300610.

Anonymous:2017:EBi

[2194] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 113 (??):iii–v, September 2017. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794731730110X.

Anonymous:2017:Ci

[2195] Anonymous. Contents. Computational Statistics & Data Analysis, 113 (??):vi–viii, September 2017. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947317301111.

Lee:2017:AMS

[2196] Jae Won Lee, Nan Lin, and Martina Mittlböck. Advances in medical statistics. Computational Statistics & Data Analysis, 113(??):1– 2, September 2017. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794731730124X.

Bertrand:2017:REM

[2197] A. Bertrand, C. Legrand, D. Léonard, and I. Van Keilegom. Robustness of estimation methods in a survival cure model with mismeasured covariates. Computational Statistics & Data Analysis, 113(??):3– 18, September 2017. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947316302870.

Blagus:2017:GBH

[2198] Rok Blagus and Lara Lusa. Gradient boosting for high-dimensional prediction of rare events. Computational Statistics & Data Analysis, 113 (??):19–37, September 2017. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947316301803.

Friedrich:2017:WBA

[2199] Sarah Friedrich, Frank Konietschke, and Markus Pauly. A wild bootstrap approach for nonparametric repeated measurements. Computational Statistics & Data Analysis, 113(??): 38–52, September 2017. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947316301530.

Frumento:2017:EEC

[2200] Paolo Frumento and Matteo Bottai. An estimating equation for censored and truncated quantile regression. Computational Statistics & Data Analysis, 113 (??):53–63, September 2017. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947316302067.

Ghebremichael-Weldeselassie:2017:SCC

[2201] Yonas Ghebremichael-Weldeselassie, Heather J. Whitaker, Ian J. Douglas, Liam Smeeth, and C. Paddy Farrington. Self-controlled case series with multiple event types. Computational Statistics & Data Analysis, 113 (??):64–72, September 2017. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947316302377.

Ghosh:2017:NIT

[2202] Santu Ghosh, Arpita Chatterjee, and Samiran Ghosh. Non-inferiority test based on transformations for nonnormal distributions. Computational Statistics & Data Analysis, 113(??):

73–87, September 2017. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947316302316.

Kim:2017:CIO

[2203] Gi-Soo Kim, Myunghee Cho Paik, and Hongsoo Kim. Causal inference with observational data under cluster-specific non-ignorable assignment mechanism. Computational Statistics & Data Analysis, 113(??): 88–99, September 2017. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947316302298.

Schmidli:2017:MAP

[2204] Heinz Schmidli, Beat Neuenschwander, and Tim Friede. Meta-analytic-predictive use of historical variance data for the design and analysis of clinical trials. Computational Statistics & Data Analysis, 113(??):100– 110, September 2017. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947316301888.

Slavtchova-Bojkova:2017:BPC

[2205] Maroussia Slavtchova-Bojkova, Plamen Trayanov, and Stoyan Dimitrov. Branching processes in continuous time as models of mutations: Computational approaches and algorithms. Computational Statistics & Data Analysis, 113(??):111– 124, September 2017. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947316303139.

Wang:2017:MBC

[2206] Zheyu Wang, Krisztian Sebestyen, and Sarah E. Monsell. Model-based clustering for assessing the prognostic value of imaging biomarkers and mixed type tests. Computational Statistics & Data Analysis, 113(??): 125–135, September 2017. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947316302614.

Williamson:2017:BAD

[2207] S. Faye Williamson, Peter Jacko, Sofía S. Villar, and Thomas Jaki. A Bayesian adaptive design for clinical trials in rare diseases. Computational Statistics & Data Analysis, 113(??): 136–153, September 2017. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947316302146.

Won:2017:CSD

[2208] Joong-Ho Won, Xiao Wu, Sang Han Lee, and Ying Lu. Cross-sectional design with a short-term follow-up for prognostic imaging biomarkers. Computational Statistics & Data Analysis, 113(??):154–176, September 2017. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S016794731730004X.

Muff:2017:BTC

[2209] Stefanie Muff, Manuela Ott, Julia Braun, and Leonhard Held. Bayesian two-component measurement error modelling for survival analysis using INLA — a case study on cardiovascular disease mortality in Switzerland. Computational Statistics & Data Analysis, 113(??):177– 193, September 2017. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947317300452.

Donev:2017:SID

[2210] Alexander N. Donev, Jesús López-Fidalgo, and Douglas P. Wiens. Special issue on design of experiments. Computational Statistics & Data Analysis, 113(??):194–195, September 2017. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947317301251.

Dette:2017:ODD

[2211] Holger Dette, Viatcheslav B. Melas, and Petr Shpilev. T -optimal discriminating designs for Fourier regression models. Computational Statistics & Data Analysis, 113(??):196– 206, September 2017. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947316301475.

McGree:2017:DTE

[2212] J. M. McGree. Developments of the total entropy utility function for the dual purpose of model discrimination and parameter estimation in Bayesian design. Computational Statistics & Data Analysis, 113(??):207– 225, September 2017. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947316301323.

Woods:2017:MSB

[2213] David C. Woods, James M. McGree, and Susan M. Lewis. Model selection via Bayesian information capacity designs for generalised linear models. Computational Statistics & Data Analysis, 113(??):226– 238, September 2017. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947316302602.

Konstantinou:2017:MRD

[2214] Maria Konstantinou, Stefanie Biedermann, and Alan Kimber. Model robust designs for survival trials. Computational Statistics & Data Analysis, 113(??):239–250, September 2017. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0167947316302407.

Smucker:2017:RCO

[2215] Byran J. Smucker, Willis Jensen, Zichen Wu, and Bo Wang. Robustness of classical and optimal designs to missing observations. Computational Statistics & Data Analysis, 113(??): 251–260, September 2017. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947316302912.

daSilva:2017:FRS

[2216] Marcelo A. da Silva, Steven G.

Gilmour, and Luzia A. Trinca. Factorial and response surface designs robust to missing observations. Computational Statistics & Data Analysis, 113(??): 261–272, September 2017. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947316301359.

Dette:2017:ODC

[2217] Holger Dette, Kirsten Schorning, and Maria Konstantinou. Optimal designs for comparing regression models with correlated observations. Computational Statistics & Data Analysis, 113(??): 273–286, September 2017. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947316301542.

Rodriguez-Diaz:2017:CIO

[2218] Juan M. Rodríguez-Díaz. Computation of c-optimal designs for models with correlated observations. Computational Statistics & Data Analysis, 113(??): 287–296, September 2017. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947316302468.

Atkinson:2017:ORC

[2219] Anthony C. Atkinson and Atanu Biswas. Optimal response and covariate-adaptive biased-coin designs for clinical trials with continuous multivariate or longitudinal responses. Computational Statistics & Data Analysis, 113(??):297– 310, September 2017. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947316301347.

Kobilinsky:2017:AGG

[2220] André Kobilinsky, Hervé Monod, and R. A. Bailey. Automatic generation of generalised regular factorial designs. Computational Statistics & Data Analysis, 113(??):311– 329, September 2017. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947316302110.

Masoudi:2017:AIC

[2221] Ehsan Masoudi, Heinz Holling, and Weng Kee Wong. Application of imperialist competitive algorithm to find minimax and standardized maximin optimal designs. Computational Statistics & Data Analysis, 113(??): 330–345, September 2017. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947316301517.

Leatherman:2017:DCP

[2222] Erin R. Leatherman, Angela M. Dean, and Thomas J. Santner. Designing combined physical and computer experiments to maximize prediction accuracy. Computational Statistics & Data Analysis, 113(??):346– 362, September 2017. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947316301773.

Rivas-Lopez:2017:OED

[2223] M. J. Rivas-López, R. C. Yu, J. López-Fidalgo, and G. Ruiz. Optimal experimental design on the loading frequency for a probabilistic fatigue model for plain and fibrereinforced concrete. Computational Statistics & Data Analysis, 113(??): 363–374, September 2017. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947316302031.

Gauthier:2017:CRI

[2224] B. Gauthier and L. Pronzato. Convex relaxation for IMSE optimal design in random-field models. Computational Statistics & Data Analysis, 113(??): 375–394, September 2017. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947316302456.

Mueller:2017:SIR

[2225] Samuel Mueller, Graciela Boente, Christophe Croux, Juan Romo, and Stefan Van Aelst. 2nd special issue on robust analysis of complex data. Computational Statistics & Data Analysis, 113(??):395– 397, September 2017. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947317301263.

Dhaene:2017:MBE

[2226] Geert Dhaene and Yu Zhu. Medianbased estimation of dynamic panel models with fixed effects. Computational Statistics & Data Analysis, 113(??):398–423, September 2017. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0167947316301335.

Bali:2017:REU

[2227] Juan Lucas Bali and Graciela Boente. Robust estimators under a functional common principal components model. Computational Statistics & Data Analysis, 113(??):424– 440, September 2017. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947316302080.

Chiancone:2017:SSI

[2228] Alessandro Chiancone, Florence Forbes, and Stéphane Girard. Student Sliced Inverse Regression. Computational Statistics & Data Analysis, 113(??): 441–456, September 2017. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947316301852.

He:2017:HDG

[2229] Yong He, Xinsheng Zhang, Pingping Wang, and Liwen Zhang. High dimensional Gaussian copula graphical model with FDR control. Computational Statistics & Data Analysis, 113(??): 457–474, September 2017. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947316301499.

Maruotti:2017:MBT

[2230] Antonello Maruotti and Antonio Punzo. Model-based time-varying clustering of multivariate longitudinal data with covariates and outliers. Computational Statistics & Data Analysis, 113(??):475–496, September 2017. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947316301360.

Anonymous:2017:EBj

[2231] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 114(??):iii–v, October 2017. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947317301329.

Anonymous:2017:Cj

[2232] Anonymous. Contents. Computational Statistics & Data Analysis, 114 (??):vi, October 2017. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947317301330.

Shults:2017:SLV

[2233] Justine Shults. Simulating longer vectors of correlated binary random variables via multinomial sampling. Computational Statistics & Data Analysis, 114(??):1–11, October 2017. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947317300750.

Yu:2017:HDC

[2234] Philip L. H. Yu, Xiaohang Wang, and Yuanyuan Zhu. High dimensional covariance matrix estimation by penalizing the matrix-logarithm transformed likelihood. Computational Statistics & Data Analysis, 114 (??):12–25, October 2017. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947317300774.

Carzolio:2017:WPT

[2235] Marcos Carzolio and Scotland Leman. Weighted particle tempering. Computational Statistics & Data Analysis, 114(??):26–37, October 2017. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947317300786.

Gorynin:2017:FSS

[2236] Ivan Gorynin, Stéphane Derrode, Emmanuel Monfrini, and Wojciech Pieczynski. Fast smoothing in switching approximations of non-linear and non-Gaussian models. Computational Statistics & Data Analysis, 114(??): 38–46, October 2017. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947317300804.

Sugasawa:2017:TRV

[2237] Shonosuke Sugasawa and Tatsuya Kubokawa. Transforming response values in small area prediction. Computational Statistics & Data Analysis, 114(??):47–60, October 2017. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947317300737.

Jeon:2017:HDH

[2238] Jong-June Jeon, Sunghoon Kwon, and Hosik Choi. Homogeneity detection for the high-dimensional generalized linear model. Computational Statistics & Data Analysis, 114(??): 61–74, October 2017. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947317300749.

Baddeley:2017:TSM

[2239] Adrian Baddeley, Andrew Hardegen, Thomas Lawrence, Robin K. Milne, Gopalan Nair, and Suman Rakshit. On two-stage Monte Carlo tests of composite hypotheses. Computational Statistics & Data Analysis, 114(??): 75–87, October 2017. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947317300762.

Li:2017:UDF

[2240] Xingxiang Li, Guosheng Cheng, Liming Wang, Peng Lai, and Fengli Song. Ultrahigh dimensional feature screening via projection. Computational Statistics & Data Analysis, 114(??): 88–104, October 2017. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947317300798.

Yu:2017:PME

[2241] Chang Yu and Daniel Zelterman. A parametric model to estimate the proportion from true null using a distribution for p-values. Computational Statistics & Data Analysis, 114(??): 105–118, October 2017. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947317300919.

Sahoo:2017:THI

[2242] Shyamsundar Sahoo and Debasis Sengupta. Testing the hypothesis of increasing hazard ratio in two samples. Computational Statistics & Data Analysis, 114(??):119– 129, October 2017. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947317300920.

Marbac:2017:FBW

[2243] Matthieu Marbac and Mohammed Sedki. A family of block-wise onefactor distributions for modeling highdimensional binary data. Computational Statistics & Data Analysis, 114 (??):130–145, October 2017. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947317300932.

Wang:2017:THM

[2244] Chunlin Wang, Paul Marriott, and Pengfei Li. Testing homogeneity for multiple nonnegative distributions with excess zero observations. Computational Statistics & Data Analysis, 114 (??):146–157, October 2017. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947317300944.

Anonymous:2017:EBk

[2245] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 115(??):iii–v, November 2017. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947317301603.

Anonymous:2017:EBl

[2246] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 115 (??):vi–vii, November 2017. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947317301615.

Wang:2017:SEL

[2247] Dongliang Wang, Lili Tian, and Yichuan Zhao. Smoothed empirical likelihood for the Youden index. Computational Statistics & Data Analysis, 115(??):1–10, November 2017. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947317300592.

Li:2017:PEP

[2248] Song Li, Geoffrey K. F. Tso, and Lufan Long. Powered embarrassing parallel MCMC sampling in Bayesian inference, a weighted average intuition. Computational Statistics & Data Analysis, 115 (??):11–20, November 2017. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947317301007.

Li:2017:FDC

[2249] Pai-Ling Li, Jeng-Min Chiou, and Yu Shyr. Functional data classification using covariate-adjusted subspace projection. Computational Statistics & Data Analysis, 115(??):21– 34, November 2017. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947317300981.

Hayes:2017:SWI

[2250] Timothy Hayes and John J. McArdle. Should we impute or should we weight? examining the performance of two CART-based techniques for addressing missing data in small sample research with nonnormal variables. Computational Statistics & Data Analysis, 115 (??):35–52, November 2017. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947317301019.

Bilton:2017:CTP

[2251] Penny Bilton, Geoff Jones, Siva Ganesh, and Steve Haslett. Classification trees for poverty mapping. Computational Statistics & Data Analysis, 115 (??):53–66, November 2017. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947317301044.

Xue:2017:SDR

[2252] Yuan Xue, Nan Zhang, Xiangrong Yin, and Haitao Zheng. Sufficient dimension reduction using Hilbert–Schmidt independence criterion. Computational Statistics & Data Analysis, 115 (??):67–78, November 2017. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947317300968.

Leonard:2017:BOS

[2253] Robert D. Leonard and David J.

Edwards. Bayesian D-optimal screening experiments with partial replication. Computational Statistics & Data Analysis, 115(??):79– 90, November 2017. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947317301275.

Frolich:2017:FSP

[2254] Markus Frölich, Martin Huber, and Manuel Wiesenfarth. The finite sample performance of semi- and nonparametric estimators for treatment effects and policy evaluation. Computational Statistics & Data Analysis, 115 (??):91–102, November 2017. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947317301020.

Cui:2017:EPL

[2255] Xia Cui, Ying Lu, and Heng Peng.

Estimation of partially linear regression models under the partial consistency property. Computational Statistics & Data Analysis, 115(??):103– 121, November 2017. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947317300993.

Svendova:2017:NME

[2256] Vendula Svendová and Michael G. Schimek. A novel method for estimating the common signals for consensus across multiple ranked lists. Computational Statistics & Data Analysis, 115 (??):122–135, November 2017. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947317301056.

Wang:2017:TLE

[2257] Shangshan Wang and Liming Xiang. Two-layer EM algorithm for ALD mixture regression models: a new solution to composite quantile regression. Computational Statistics & Data Analysis, 115(??):136– 154, November 2017. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794731730141X.

Willems:2017:OSS

[2258] S. J. W. Willems, M. Fiocco, and J. J. Meulman. Optimal scaling for survival analysis with ordinal data. Computational Statistics & Data Analysis, 115 (??):155–171, November 2017. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947317301032.

Bhuyan:2017:ERS

[2259] Prajamitra Bhuyan and Debasis Sengupta. Estimation of reliability with semi-parametric modeling of degradation. Computational Statistics & Data Analysis, 115(??):172– 185, November 2017. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947317301470.

Hoff:2017:LFN

[2260] Peter D. Hoff. Lasso, fractional norm and structured sparse estimation using a Hadamard product parametrization. Computational Statistics & Data Analysis, 115(??): 186–198, November 2017. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947317301469.

Ledoit:2017:NIQ

[2261] Olivier Ledoit and Michael Wolf. Numerical implementation of the QuEST function. Computational Statistics & Data Analysis, 115(??):199– 223, November 2017. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947317301433.

Cuervo:2017:AFG

[2262] Daniel Palhazi Cuervo, Peter Goos, and Kenneth Sörensen. An algorithmic framework for generating optimal twostratum experimental designs. Computational Statistics & Data Analysis, 115 (??):224–249, November 2017. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947317301457.

Tsai:2017:HTN

[2263] Arthur C. Tsai, Michelle Liou, Maria Simak, and Philip E. Cheng. On hyperbolic transformations to normality. Computational Statistics & Data Analysis, 115(??):250–266, November 2017. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947317301408.

Lee:2017:ACF

[2264] Keunbaik Lee, Changryong Baek, and Michael J. Daniels. ARMA Cholesky factor models for the covariance matrix of linear models. Computational Statistics & Data Analysis, 115(??): 267–280, November 2017. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947317300956.

Kuk:2017:FCA

[2265] Anthony Y. C. Kuk. Function compositional adjustments of conditional quantile curves. Computational Statistics & Data Analysis, 115(??):281– 293, November 2017. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016794731730097X.

Zhao:2017:TRM

[2266] Junlong Zhao, Lu Niu, and Shushi Zhan. Trace regression model with simultaneously low rank and row(column) sparse parameter. Computational Statistics & Data Analysis, 116(??):1–18, December 2017. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947317301482.

Bin:2017:DIP

[2267] Riccardo De Bin, Anne-Laure Boulesteix, and Willi Sauerbrei. Detection of influential points as a byproduct of resampling-based variable selection procedures. Computational Statistics & Data Analysis, 116(??):19– 31, December 2017. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947317301494.

Rolls:2017:MDE

[2268] David A. Rolls and Garry Robins. Minimum distance estimators of population size from snowball samples using conditional estimation and scaling of exponential random graph models. Computational Statistics & Data Analysis, 116 (??):32–48, December 2017. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947317301524.

Zhang:2017:CTE

[2269] Feipeng Zhang and Qunhua Li.

A continuous threshold expectile model. Computational Statistics & Data Analysis, 116(??):49–66, December 2017. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947317301536.

Chu:2017:DEK

[2270] Chi-Yang Chu, Daniel J. Henderson, and Christopher F. Parmeter. On discrete Epanechnikov kernel functions. Computational Statistics & Data Analysis, 116(??):79– 105, December 2017. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947317301512.

Hao:2017:CCR

[2271] Peng Hao and Junpeng Guo. Constrained center and range joint model for interval-valued symbolic data regression. Computational Statistics & Data Analysis, 116(??):106– 138, December 2017. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947317301445.

Gaynor:2017:IRS

[2272] Sheila Gaynor and Eric Bair. Identification of relevant subtypes via preweighted sparse clustering. Computational Statistics & Data Analysis, 116 (??):139–154, December 2017. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947317301421.

Hirose:2017:NAS

[2273] Masayo Yoshimori Hirose. Non-area-specific adjustment factor for secondorder efficient empirical Bayes confidence interval. Computational Statistics & Data Analysis, 116(??):67– 78, December 2017. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947317301500.

Anonymous:2018:EBa

[2274] Anonymous. Editorial board. Computational Statistics & Data Analysis, 117(??):iii–v, January 2018. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947317302141.

Anonymous:2018:Ca

[2275] Anonymous. Contents. Computational Statistics & Data Analysis, 117 (??):vi–vii, January 2018. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947317302153.

Shen:2018:QRP

[2276] Yu Shen and Han-Ying Liang. Quantile regression for partially linear varyingcoefficient model with censoring indicators missing at random. Computational Statistics & Data Analysis, 117(??):1–18, January 2018. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947317301548.

Park:2018:SEB

[2277] Junyong Park. Simultaneous estimation based on empirical likelihood and general maximum likelihood estimation. Computational Statistics & Data Analysis, 117(??):19– 31, January 2018. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016794731730172X.

Cai:2018:NLD

[2278] Wei Cai, Guoyu Guan, Rui Pan, Xuening Zhu, and Hansheng Wang. Network linear discriminant analysis. Computational Statistics & Data Analysis, 117 (??):32–44, January 2018. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016794731730155X.

Roozbeh:2018:OQB

[2279] Mahdi Roozbeh. Optimal QR-based estimation in partially linear regression models with correlated errors using GCV criterion. Computational Statistics & Data Analysis, 117(??): 45–61, January 2018. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947317301718.

Samanta:2018:ORI

[2280] Debashis Samanta and Debasis Kundu. Order restricted inference of a multiple step-stress model. Computational Statistics & Data Analysis, 117 (??):62–75, January 2018. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947317301706.

Ma:2018:IDM

[2281] Hua Ma, Andriy I. Bandos, and David Gur. Informativeness of diagnostic marker values and the impact of data grouping. Computational Statistics & Data Analysis, 117 (??):76–89, January 2018. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947317301664.

Kang:2018:IDP

[2282] Kai Kang, Vasileios Maroulas, Ioannis Schizas, and Feng Bao. Improved distributed particle filters for tracking in a wireless sensor network. Computational Statistics & Data Analysis, 117 (??):90–108, January 2018. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947317301676.

Barthel:2018:VCB

[2283] Nicole Barthel, Candida Geerdens, Matthias Killiches, Paul Janssen, and Claudia Czado. Vine copula based likelihood estimation of dependence patterns in multivariate event time data. Computational Statistics & Data Analysis, 117(??):109– 127, January 2018. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947317301688.

Mao:2018:TIH

[2284] Guangyu Mao. Testing independence in high dimensions using Kendall’s tau. Computational Statistics & Data Analysis, 117(??):128– 137, January 2018. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947317301743.

Marchetti:2018:SDC

[2285] Yuliya Marchetti, Hai Nguyen, Amy Braverman, and Noel Cressie. Spatial data compression via adaptive dispersion clustering. Computational Statistics & Data Analysis, 117(??): 138–153, January 2018. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947317301731.

Craiu:2018:SEC

[2286] Radu V. Craiu and Thierry Duchesne. A scalable and efficient covariate selection criterion for mixed effects regression models with unknown random effects structure. Computational Statistics & Data Analysis, 117 (??):154–161, January 2018. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016794731730169X.

Sun:2018:DSR

[2287] Peng Sun, Inyoung Kim, and Ki-Ahm Lee. Dual-semiparametric regression using weighted Dirichlet process mixture. Computational Statistics & Data Analysis, 117(??):162– 181, January 2018. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947317301755.

Holland-Letz:2018:OED

[2288] T. Holland-Letz and A. Kopp-Schneider. Optimal experimental designs for estimating the drug combination index in toxicology. Computational Statistics & Data Analysis, 117 (??):182–193, January 2018. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947317301767.

Deng:2018:RAP

[2289] Lifeng Deng, Jieli Ding, Yanyan Liu, and Chengdong Wei. Regression analysis for the proportional hazards model with parameter constraints under casecohort design. Computational Statistics & Data Analysis, 117(??):194– 206, January 2018. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947317301937.

Anonymous:2018:EBb

[2290] Anonymous. Editorial board. Computational Statistics & Data Analysis, 118(??):iii–v, February 2018. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947317302268.

Anonymous:2018:Cb

[2291] Anonymous. Contents. Computational Statistics & Data Analysis, 118 (??):vi, February 2018. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794731730227X.

Drovandi:2018:APM

[2292] Christopher C. Drovandi, Matthew T.

Moores, and Richard J. Boys. Accelerating pseudo-marginal MCMC using Gaussian processes. Computational Statistics & Data Analysis, 118 (??):1–17, February 2018. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947317301950.

Huang:2018:NNP

[2293] Kai Huang and Jie Mi. A new non-parametric estimator for instant system availability. Computational Statistics & Data Analysis, 118(??): 18–29, February 2018. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947317301949.

Lombard:2018:SSR

[2294] F. Lombard and C. van Zyl. Signed sequential rank CUSUMs. Computational Statistics & Data Analysis, 118 (??):30–39, February 2018. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947317301871.

Buonocore:2018:GRV

[2295] A. Buonocore, A. G. Nobile, and E. Pirozzi. Generating random variates from PDF of Gauss–Markov processes with a reflecting boundary. Computational Statistics & Data Analysis, 118 (??):40–53, February 2018. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947317301883.

Jiang:2018:SSD

[2296] Binyan Jiang and Jialiang Li. Sample size determination for high dimensional parameter estimation with application to biomarker identification. Computational Statistics & Data Analysis, 118 (??):54–65, February 2018. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947317301901.

Li:2018:ETT

[2297] Jianbo Li, Heng Lian, Xuejun Jiang, and Xinyuan Song. Estimation and testing for time-varying quantile single-index models with longitudinal data. Computational Statistics & Data Analysis, 118(??):66– 83, February 2018. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947317301913.

Chen:2018:OTL

[2298] Ping-Yang Chen, Ray-Bing Chen, and C. Devon Lin. Optimizing two-level orthogonal arrays for simultaneously estimating main effects and pre-specified two-factor interactions. Computational Statistics & Data Analysis, 118 (??):84–97, February 2018. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947317301925.

Lin:2018:RMB

[2299] Chang-Yun Lin and Po Yang. Robust multistratum baseline designs. Computational Statistics & Data Analysis, 118 (??):98–111, February 2018. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947317301895.

Sun:2018:CTP

[2300] Zhihua Sun, Xue Ye, and Liuquan Sun. Consistent test for parametric models with right-censored data using projections. Computational Statistics & Data Analysis, 118(??):112– 125, February 2018. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947317301998.

Li:2018:SVS

[2301] Haocheng Li, Di Shu, Yukun Zhang, and Grace Y. Yi. Simultaneous variable selection and estimation for multivariate multilevel longitudinal data with both continuous and binary responses. Computational Statistics & Data Analysis, 118(??):126– 137, February 2018. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947317301974.

Anonymous:2018:EBc

[2302] Anonymous. Editorial board. Computational Statistics & Data Analysis, 119(??):iii–v, March 2018. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947317302463.

Anonymous:2018:Cc

[2303] Anonymous. Contents. Computational Statistics & Data Analysis, 119 (??):vi, March 2018. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947317302475.

Hatjispyros:2018:DMG

[2304] Spyridon J. Hatjispyros, Christos Merkatas, Theodoros Nicoleris, and Stephen G. Walker. Dependent mixtures of geometric weights priors. Computational Statistics & Data Analysis, 119(??):1–18, March 2018. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947317302001.

Lloyd-Jones:2018:GCA

[2305] Luke R. Lloyd-Jones, Hien D. Nguyen, and Geoffrey J. McLachlan. A globally convergent algorithm for lassopenalized mixture of linear regression models. Computational Statistics & Data Analysis, 119(??):19– 38, March 2018. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947317301962.

Karimi:2018:JMS

[2306] M. Karimi, G. Rey, and A. Latouche. A joint modelling of socioprofessional trajectories and causespecific mortality. Computational Statistics & Data Analysis, 119(??): 39–54, March 2018. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947317302207.

Hartigan:2018:AHC

[2307] Luke Hartigan. Alternative HAC covariance matrix estimators with improved finite sample properties. Computational Statistics & Data Analysis, 119(??):55–73, March 2018. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947317302074.

Liu:2018:NNS

[2308] Yanyan Liu, Jing Zhang, and Xingqiu Zhao. A new nonparametric screening method for ultrahigh-dimensional survival data. Computational Statistics & Data Analysis, 119(??):74– 85, March 2018. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947317302219.

Ghosh:2018:NEL

[2309] Sujit K. Ghosh, Christopher B. Burns, Daniel L. Prager, Li Zhang, and Glenn Hui. On nonparametric estimation of the latent distribution for ordinal data. Computational Statistics & Data Analysis, 119(??): 86–98, March 2018. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947317302098.

Duarte:2018:ABS

[2310] Belmiro P. M. Duarte, Guillaume Sagnol, and Weng Kee Wong. An algorithm based on semidefinite programming for finding minimax optimal designs. Computational Statistics & Data Analysis, 119(??):99– 117, March 2018. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947317302086.

Chen:2018:RFS

[2311] Xiaolin Chen, Xiaojing Chen, and Hong Wang. Robust feature screening for ultra-high dimensional right censored data via distance correlation. Computational Statistics & Data Analysis, 119(??):118– 138, March 2018. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947317302323.

Perrin:2018:DDK

[2312] G. Perrin, C. Soize, and N. Ouhbi.

Data-driven kernel representations for sampling with an unknown block dependence structure under correlation constraints. Computational Statistics & Data Analysis, 119(??):139– 154, March 2018. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947317302335.

Anonymous:2018:EBd

[2313] Anonymous. Editorial board. Computational Statistics & Data Analysis, 120(??):ii–iv, April 2018. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947317302657.

Carissimo:2018:VCR

[2314] Annamaria Carissimo, Luisa Cutillo, and Italia De Feis. Validation of community robustness. Computational Statistics & Data Analysis, 120 (??):1–24, April 2018. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947317302347.

Cederbaum:2018:FSA

[2315] Jona Cederbaum, Fabian Scheipl, and Sonja Greven. Fast symmetric additive covariance smoothing. Computational Statistics & Data Analysis, 120(??):25–41, April 2018. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947317302372.

Loperfido:2018:SBP

[2316] Nicola Loperfido. Skewness-based projection pursuit: A computational approach. Computational Statistics & Data Analysis, 120(??):42– 57, April 2018. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947317302360.

Yang:2018:SJE

[2317] Hanfang Yang and Yichuan Zhao. Smoothed jackknife empirical likelihood for the one-sample difference of quantiles. Computational Statistics & Data Analysis, 120(??):

58–69, April 2018. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947317302359.

Bergmeir:2018:NVC

[2318] Christoph Bergmeir, Rob J. Hyndman, and Bonsoo Koo. A note on the validity of cross-validation for evaluating autoregressive time series prediction. Computational Statistics & Data Analysis, 120(??): 70–83, April 2018. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947317302384.

Fattore:2018:LSA

[2319] Marco Fattore, Matteo Pelagatti, and Giorgio Vittadini. A least squares approach to latent variables extraction in formative–reflective models. Computational Statistics & Data Analysis, 120(??):84–97, April 2018. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947317302414.

Suesse:2018:MML

[2320] Thomas Suesse. Marginal maximum likelihood estimation of SAR models with missing data. Computational Statistics & Data Analysis, 120 (??):98–110, April 2018. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947317302396.

Hartog:2018:NBL

[2321] Jarno Hartog and Harry van Zanten. Nonparametric Bayesian label prediction on a graph. Computational Statistics & Data Analysis, 120 (??):111–131, April 2018. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947317302529.

Anonymous:2018:EBe

[2322] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 121(??):ii–iv, May 2018. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947318300185.

Chao:2018:MFE

[2323] Shih-Kang Chao, Wolfgang K. Härdle, and Chen Huang. Multivariate factorizable expectile regression with application to fMRI data. Computational Statistics & Data Analysis, 121 (??):1–19, May 2018. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947317302542.

Guadarrama:2018:SAE

[2324] María Guadarrama, Isabel Molina, and J. N. K. Rao. Small area estimation of general parameters under complex sampling designs. Computational Statistics & Data Analysis, 121(??):20–40, May 2018. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947317302517.

Song:2018:PCT

[2325] Junmo Song and Jiwon Kang. Parameter change tests for ARMA– GARCH models. Computational Statistics & Data Analysis, 121(??): 41–56, May 2018. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947317302578.

Tsai:2018:CCC

[2326] Miao-Yu Tsai and Chao-Chun Lin.

Concordance correlation coefficients estimated by variance components for longitudinal normal and Poisson data. Computational Statistics & Data Analysis, 121(??):57– 70, May 2018. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794731730258X.

Chatla:2018:EEC

[2327] Suneel Babu Chatla and Galit Shmueli.

Efficient estimation of COM–Poisson regression and a generalized additive model. Computational Statistics & Data Analysis, 121(??):71– 88, May 2018. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947317302608.

Shan:2018:JEM

[2328] Liang Shan and Inyoung Kim.

Joint estimation of multiple Gaussian graphical models across unbalanced classes. Computational Statistics & Data Analysis, 121(??):89– 103, May 2018. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947317302530.

Lin:2018:ALD

[2329] Huiming Lin, Guoyou Qin, Jiajia Zhang, and Zhongyi Zhu. Analysis of longitudinal data with covariate measurement error and missing responses: An improved unbiased estimating equation. Computational Statistics & Data Analysis, 121(??): 104–112, May 2018. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947317302554.

Xu:2018:ABM

[2330] Hang Xu, Mayer Alvo, and Philip L. H. Yu. Angle-based models for ranking data. Computational Statistics & Data Analysis, 121(??): 113–136, May 2018. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947317302591.

Amorim:2018:SSI

[2331] G. Amorim, O. Thas, K. Vermeulen, S. Vansteelandt, and J. De Neve. Small sample inference for probabilistic index models. Computational Statistics & Data Analysis, 121(??): 137–148, May 2018. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947317302402.

Grazian:2018:JPM

[2332] Clara Grazian and Christian P. Robert.

Jeffreys priors for mixture estimation: Properties and alternatives. Computational Statistics & Data Analysis, 121(??):149–163, May 2018. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794731730261X.

Souza:2018:EDG

[2333] M. A. O. Souza, H. S. Migon, and J. B. M. Pereira. Extended dynamic generalized linear models: The twoparameter exponential family. Computational Statistics & Data Analysis, 121(??):164–179, May 2018. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947316302249.

Cai:2018:SCR

[2334] Qingyun Cai. A scoring criterion for rejection of clustered p-values. Computational Statistics & Data Analysis, 121(??):180–189, May 2018. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947316300196.

Zhu:2018:MTF

[2335] Xuwen Zhu and Volodymyr Melnykov. Manly transformation in finite mixture modeling. Computational Statistics & Data Analysis, 121 (??):190–208, May 2018. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947316300019.

Anonymous:2018:EBf

[2336] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 122(??):ii–iv, June 2018. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947318300367.

Ristl:2018:OET

[2337] Robin Ristl, Dong Xi, Ekkehard Glimm, and Martin Posch. Optimal exact tests for multiple binary endpoints. Computational Statistics & Data Analysis, 122(??):1–17, June 2018. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947318300021.

Yan:2018:FMV

[2338] Xiaodong Yan, Niansheng Tang, Jinhan Xie, Xianwen Ding, and Zhiqiang Wang. Fused mean-variance filter for feature screening. Computational Statistics & Data Analysis, 122(??): 18–32, June 2018. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947317302566.

Alzahrani:2018:MST

[2339] Naif Alzahrani, Peter Neal, Simon E. F. Spencer, Trevelyan J. McKinley, and Panayiota Touloupou. Model selection for time series of count data. Computational Statistics & Data Analysis, 122(??):33–44, June 2018. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947318300033.

Sun:2018:FEE

[2340] Ying Sun, Xiaohui Chang, and Yongtao Guan. Flexible and efficient estimating equations for variogram estimation. Computational Statistics & Data Analysis, 122(??): 45–58, June 2018. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794731830001X.

Lee:2018:AGS

[2341] Unkyung Lee, Yanqing Sun, Thomas H.

Scheike, and Peter B. Gilbert. Analysis of generalized semiparametric regression models for cumulative incidence functions with missing covariates. Computational Statistics & Data Analysis, 122(??):59– 79, June 2018. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947318300045.

Feng:2018:ANM

[2342] Long Feng and Lee H. Dicker. Approximate nonparametric maximum likelihood for mixture models: A convex optimization approach to fitting arbitrary multivariate mixing distributions. Computational Statistics & Data Analysis, 122(??):80– 91, June 2018. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947318300070.

Biscarri:2018:SFM

[2343] William Biscarri, Sihai Dave Zhao, and Robert J. Brunner. A simple and fast method for computing the Poisson binomial distribution function. Computational Statistics & Data Analysis, 122(??):92–100, June 2018. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947318300082.

Park:2018:JDF

[2344] So Young Park, Luo Xiao, Jayson D. Willbur, Ana-Maria Staicu, and N. L’ntshotsholé Jumbe. A joint design for functional data with application to scheduling ultrasound scans. Computational Statistics & Data Analysis, 122(??):101–114, June 2018. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947318300100.

Levi:2018:BIC

[2345] Evgeny Levi and Radu V. Craiu. Bayesian inference for conditional copulas using Gaussian Process single index models. Computational Statistics & Data Analysis, 122(??): 115–134, June 2018. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947318300148.

Hidalgo:2018:GFT

[2346] Sebastian J. Teran Hidalgo, Michael C. Wu, Stephanie M. Engel, and Michael R. Kosorok. Goodness-of-fit test for nonparametric regression models: Smoothing spline ANOVA models as example. Computational Statistics & Data Analysis, 122(??): 135–155, June 2018. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947318300057.

Wang:2018:GTC

[2347] Dewei Wang, Christopher S. McMahan, Joshua M. Tebbs, and Christopher R. Bilder. Group testing case identification with biomarker information. Computational Statistics & Data Analysis, 122(??):156– 166, June 2018. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947318300069.

Anonymous:2018:EBg

[2348] Anonymous. Editorial board. Computational Statistics & Data Analysis, 123(??):ii–iv, July 2018. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947318300616.

Vencalek:2018:DWB

[2349] Ondrej Vencalek and Oleksii Pokotylo.

Depth-weighted Bayes classification. Computational Statistics & Data Analysis, 123(??):1–12, July 2018. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947318300124.

Huttunen:2018:AEA

[2350] J. M. J. Huttunen, J. P. Kaipio, and H. Haario. Approximation error approach in spatiotemporally chaotic models with application to Kuramoto– Sivashinsky equation. Computational Statistics & Data Analysis, 123(??): 13–31, July 2018. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947318300240.

Ye:2018:CSB

[2351] Mao Ye, Peng Zhang, and Lizhen Nie. Clustering sparse binary data with hierarchical Bayesian Bernoulli mixture model. Computational Statistics & Data Analysis, 123(??):32– 49, July 2018. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794731830029X.

Lee:2018:TDV

[2352] Kyeongeun Lee, Young K. Lee, Byeong U. Park, and Seong J. Yang. Time-dynamic varying coefficient models for longitudinal data. Computational Statistics & Data Analysis, 123(??):50–65, July 2018. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947318300252.

Talska:2018:CRF

[2353] R. Talská, A. Menafoglio, J. Machalová, K. Hron, and E. Fiserová. Compositional regression with functional response. Computational Statistics & Data Analysis, 123(??):66– 85, July 2018. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947318300276.

Bartolucci:2018:DRD

[2354] Francesco Bartolucci, Maria Francesca Marino, and Silvia Pandolfi. Dealing with reciprocity in dynamic stochastic block models. Computational Statistics & Data Analysis, 123(??): 86–100, July 2018. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947318300112.

Jacques:2018:MBC

[2355] Julien Jacques and Christophe Biernacki. Model-based co-clustering for ordinal data. Computational Statistics & Data Analysis, 123(??): 101–115, July 2018. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947318300239.

Zammit-Mangion:2018:SLA

[2356] Andrew Zammit-Mangion and Jonathan Rougier. A sparse linear algebra algorithm for fast computation of prediction variances with Gaussian Markov random fields. Computational Statistics & Data Analysis, 123 (??):116–130, July 2018. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794731830032X.

Stindl:2018:LBI

[2357] Tom Stindl and Feng Chen. Likelihood based inference for the multivariate renewal Hawkes process. Computational Statistics & Data Analysis, 123(??):131–145, July 2018. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947318300306.

Zhang:2018:SSM

[2358] Yue Zhang and Bin Zhang. Semiparametric spatial model for intervalcensored data with time-varying covariate effects. Computational Statistics & Data Analysis, 123(??):146– 156, July 2018. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947318300264.

Anonymous:2018:EBh

[2359] Anonymous. Editorial board. Computational Statistics & Data Analysis, 124(??):ii–iv, August 2018. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947318300896.

Kheifets:2018:MST

[2360] Igor L. Kheifets. Multivariate specification tests based on a dynamic Rosenblatt transform. Computational Statistics & Data Analysis, 124(??): 1–14, August 2018. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947318300318.

Guo:2018:TES

[2361] Jia Guo, Bu Zhou, and Jin-Ting Zhang. Testing the equality of several covariance functions for functional data: A supremum-norm based test. Computational Statistics & Data Analysis, 124(??):15–26, August 2018. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947318300392.

Yuen:2018:CEG

[2362] T. P. Yuen, H. Wong, and K. F. C.

Yiu. On constrained estimation of graphical time series models. Computational Statistics & Data Analysis, 124(??):27–52, August 2018. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947318300288.

Peng:2018:SRA

[2363] Mengjiao Peng, Liming Xiang, and Shanshan Wang. Semiparametric regression analysis of clustered survival data with semi-competing risks. Computational Statistics & Data Analysis, 124(??):53–70, August 2018. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947318300409.

Warne:2018:MRS

[2364] David J. Warne, Ruth E. Baker, and Matthew J. Simpson. Multilevel rejection sampling for approximate Bayesian computation. Computational Statistics & Data Analysis, 124(??):71–86, August 2018. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947318300483.

Ogasawara:2018:FIC

[2365] Haruhiko Ogasawara. A family of the information criteria using the phidivergence for categorical data. Computational Statistics & Data Analysis, 124(??):87–103, August 2018. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947318300495.

Daniel:2018:PCL

[2366] Jeffrey Daniel, Julie Horrocks, and Gary J. Umphrey. Penalized composite likelihoods for inhomogeneous Gibbs point process models. Computational Statistics & Data Analysis, 124 (??):104–116, August 2018. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947318300446.

OShaughnessy:2018:BLD

[2367] P. Y. O’Shaughnessy and A. H. Welsh. Bootstrapping longitudinal data with multiple levels of variation. Computational Statistics & Data Analysis, 124 (??):117–131, August 2018. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947318300434.

He:2018:VSH

[2368] Yong He, Xinsheng Zhang, and Liwen Zhang. Variable selection for high dimensional Gaussian copula regression model: An adaptive hypothesis testing procedure. Computational Statistics & Data Analysis, 124 (??):132–150, August 2018. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947318300513.

Gressani:2018:FBI

[2369] Oswaldo Gressani and Philippe Lambert. Fast Bayesian inference using Laplace approximations in a flexible promotion time cure model based on P -splines. Computational Statistics & Data Analysis, 124(??):151– 167, August 2018. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794731830046X.

Xu:2018:RAT

[2370] Wenfu Xu, Jingya Gao, Feifang Hu, and Siu Hung Cheung. Responseadaptive treatment allocation for noninferiority trials with heterogeneous variances. Computational Statistics & Data Analysis, 124(??):168– 179, August 2018. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947318300537.

Kawano:2018:SPC

[2371] Shuichi Kawano, Hironori Fujisawa, Toyoyuki Takada, and Toshihiko Shiroishi. Sparse principal component regression for generalized linear models. Computational Statistics & Data Analysis, 124(??):180– 196, August 2018. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947318300562.

Villa:2018:OPN

[2372] Cristiano Villa and Francisco J. Rubio. Objective priors for the number of degrees of freedom of a multivariate t distribution and the t-copula. Computational Statistics & Data Analysis, 124 (??):197–219, August 2018. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947318300707.

Papastamoulis:2018:OBM

[2373] Panagiotis Papastamoulis. Overfitting Bayesian mixtures of factor analyzers with an unknown number of components. Computational Statistics & Data Analysis, 124(??):220– 234, August 2018. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947318300550.

Shi:2018:FBS

[2374] Xingjie Shi, Yuan Huang, Jian Huang, and Shuangge Ma. A Forward and Backward Stagewise algorithm for nonconvex loss functions with adaptive Lasso. Computational Statistics & Data Analysis, 124(??):235– 251, August 2018. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947318300549.

Giraldi:2018:OPO

[2375] Loïc Giraldi, Olivier P. Le Maître, Ibrahim Hoteit, and Omar M. Knio. Optimal projection of observations in a Bayesian setting. Computational Statistics & Data Analysis, 124(??): 252–276, August 2018. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947318300501.

Dehideniya:2018:OBD

[2376] Mahasen B. Dehideniya, Christopher C. Drovandi, and James M. McGree. Optimal Bayesian design for discriminating between models with intractable likelihoods in epidemiology. Computational Statistics & Data Analysis, 124(??):277– 297, August 2018. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947318300525.

Anonymous:2018:EBi

[2377] Anonymous. Editorial board. Computational Statistics & Data Analysis, 125 (??):ii–iv, September 2018. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947318301063.

Wang:2018:SAA

[2378] Shuying Wang, Chunjie Wang, Peijie Wang, and Jianguo Sun. Semiparametric analysis of the additive hazards model with informatively intervalcensored failure time data. Computational Statistics & Data Analysis, 125 (??):1–9, September 2018. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947318300719.

Cleveland:2018:RTE

[2379] Jason Cleveland, Weilong Zhao, and Wei Wu. Robust template estimation for functional data with phase variability using band depth. Computational Statistics & Data Analysis, 125 (??):10–26, September 2018. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947318300574.

Jung:2018:CDS

[2380] Sungkyu Jung. Continuum directions for supervised dimension reduction. Computational Statistics & Data Analysis, 125(??):27– 43, September 2018. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947318300756.

Xu:2018:ULM

[2381] Peirong Xu, Heng Peng, and Tao Huang. Unsupervised learning of mixture regression models for longitudinal data. Computational Statistics & Data Analysis, 125(??):44– 56, September 2018. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947318300720.

Vuollo:2018:SSA

[2382] Ville Vuollo and Lasse Holmström. A scale space approach for exploring structure in spherical data. Computational Statistics & Data Analysis, 125 (??):57–69, September 2018. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947318300744.

Alam:2018:IOU

[2383] Md. Ashad Alam, Vince D. Calhoun, and Yu-Ping Wang. Identifying outliers using multiple kernel canonical correlation analysis with application to imaging genetics. Computational Statistics & Data Analysis, 125(??): 70–85, September 2018. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947318300732.

Luca:2018:PPM

[2384] Stijn E. Luca, Marco A. F. Pimentel, Peter J. Watkinson, and David A. Clifton. Point process models for novelty detection on spatial point patterns and their extremes. Computational Statistics & Data Analysis, 125 (??):86–103, September 2018. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947318300811.

Berckmoes:2018:SMA

[2385] Ben Berckmoes, Anna Ivanova, and Geert Molenberghs. On the sample mean after a group sequential trial. Computational Statistics & Data Analysis, 125(??):104– 118, September 2018. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947318300768.

Qu:2018:ILS

[2386] Lianqiang Qu, Xinyuan Song, and Liuquan Sun. Identification of local sparsity and variable selection for varying coefficient additive hazards models. Computational Statistics & Data Analysis, 125(??):119– 135, September 2018. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947318300847.

Forzani:2018:SDR

[2387] Liliana Forzani, Rodrigo García Arancibia, Pamela Llop, and Diego Tomassi. Supervised dimension reduction for ordinal predictors. Computational Statistics & Data Analysis, 125(??): 136–155, September 2018. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016794731830080X.

Marchese:2018:JRA

[2388] Scott Marchese and Guoqing Diao. Joint regression analysis of mixedtype outcome data via efficient scores. Computational Statistics & Data Analysis, 125(??):156–170, September 2018. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947318300471.

Anonymous:2018:EBj

[2389] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 126(??):ii–iv, October 2018. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947318301300.

Castro:2018:PLB

[2390] Luis E. Castro and Nazrul I. Shaikh.

A particle-learning-based approach to estimate the influence matrix of online social networks. Computational Statistics & Data Analysis, 126(??): 1–18, October 2018. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947318300094.

Chandra:2018:SAE

[2391] Hukum Chandra, Nicola Salvati, and Ray Chambers. Small area estimation under a spatially nonlinear model. Computational Statistics & Data Analysis, 126(??):19– 38, October 2018. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947318300835.

Chen:2018:SFL

[2392] Lu-Hung Chen and Ci-Ren Jiang.

Sensible functional linear discriminant analysis. Computational Statistics & Data Analysis, 126(??):39– 52, October 2018. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794731830094X.

Rodrigues:2018:RPP

[2393] G. S. Rodrigues, D. Prangle, and S. A. Sisson. Recalibration: a postprocessing method for approximate Bayesian computation. Computational Statistics & Data Analysis, 126(??): 53–66, October 2018. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947318300859.

Dong:2018:SDR

[2394] Yuexiao Dong, Qi Xia, Cheng Yong Tang, and Zeda Li. On sufficient dimension reduction with missing responses through estimating equations. Computational Statistics & Data Analysis, 126(??):67– 77, October 2018. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947318300951.

Zheng:2018:BEH

[2395] Zemin Zheng, Yang Li, Chongxiu Yu, and Gaorong Li. Balanced estimation for high-dimensional measurement error models. Computational Statistics & Data Analysis, 126 (??):78–91, October 2018. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947318300987.

Bernardi:2018:BQR

[2396] Mauro Bernardi, Marco Bottone, and Lea Petrella. Bayesian quantile regression using the skew exponential power distribution. Computational Statistics & Data Analysis, 126(??): 92–111, October 2018. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947318300975.

Price:2018:INS

[2397] David J. Price, Nigel G. Bean, Joshua V. Ross, and Jonathan Tuke. An induced natural selection heuristic for finding optimal Bayesian experimental designs. Computational Statistics & Data Analysis, 126(??): 112–124, October 2018. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947318301002.

Lee:2018:SPB

[2398] Sangin Lee, Youngjo Lee, and Yudi Pawitan. Sparse pathway-based prediction models for high-throughput molecular data. Computational Statistics & Data Analysis, 126(??):125– 135, October 2018. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947318301014.

Russo:2018:BIG

[2399] Massimiliano Russo, Daniele Durante, and Bruno Scarpa. Bayesian inference on group differences in multivariate categorical data. Computational Statistics & Data Analysis, 126 (??):136–149, October 2018. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947318300999.

Igari:2018:BDC

[2400] Ryosuke Igari and Takahiro Hoshino.

A Bayesian data combination approach for repeated durations under unobserved missing indicators: Application to interpurchase-timing in marketing. Computational Statistics & Data Analysis, 126(??):150– 166, October 2018. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947318300823.

Anonymous:2018:EBk

[2401] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 127(??):ii–iv, November 2018. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947318301518.

Wojtys:2018:CBG

[2402] Malgorzata Wojty’s, Giampiero Marra, and Rosalba Radice. Copula based generalized additive models for location, scale and shape with nonrandom sample selection. Computational Statistics & Data Analysis, 127 (??):1–14, November 2018. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947318301026.

Dai:2018:TCS

[2403] Xinjie Dai, Cuizhen Niu, and Xu Guo. Testing for central symmetry and inference of the unknown center. Computational Statistics & Data Analysis, 127 (??):15–31, November 2018. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947318301166.

Zhao:2018:TVQ

[2404] Weihua Zhao, Yan Zhou, and Heng Lian. Time-varying quantile singleindex model for multivariate responses. Computational Statistics & Data Analysis, 127(??):32– 49, November 2018. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947318301154.

Chen:2018:WBS

[2405] Nan Chen, Bradley P. Carlin, and Brian P. Hobbs. Web-based statistical tools for the analysis and design of clinical trials that incorporate historical controls. Computational Statistics & Data Analysis, 127 (??):50–68, November 2018. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947318301099.

Li:2018:ANI

[2406] Huiqiong Li, Guoliang Tian, Niansheng Tang, and Hongyuan Cao. Assessing non-inferiority for incomplete paired-data under non-ignorable missing mechanism. Computational Statistics & Data Analysis, 127(??):69– 81, November 2018. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794731830118X.

Chen:2018:ELC

[2407] Shuo Chen, Jian Kang, Yishi Xing, Yunpeng Zhao, and Donald K. Milton. Estimating large covariance matrix with network topology for highdimensional biomedical data. Computational Statistics & Data Analysis, 127 (??):82–95, November 2018. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947318301178.

Giguelay:2018:TMD

[2408] J. Giguelay and S. Huet. Testing k-monotonicity of a discrete distribution. application to the estimation of the number of classes in a population. Computational Statistics & Data Analysis, 127(??):96– 115, November 2018. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947318300458.

Lee:2018:IDE

[2409] Kyoungjae Lee, Jaeyong Lee, and Sarat C. Dass. Inference for differential equation models using relaxation via dynamical systems. Computational Statistics & Data Analysis, 127(??): 116–134, November 2018. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947318301233.

deToledo:2018:MCS

[2410] Pablo Álvarez de Toledo, Fernando Núñez, and Carlos Usabiaga. Matching and clustering in square contingency tables. who matches with whom in the Spanish labour market. Computational Statistics & Data Analysis, 127 (??):135–159, November 2018. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794731830121X.

Andrews:2018:AOU

[2411] Jeffrey L. Andrews. Addressing overfitting and underfitting in Gaussian model-based clustering. Computational Statistics & Data Analysis, 127(??): 160–171, November 2018. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947318301245.

Das:2018:BNP

[2412] Priyam Das and Subhashis Ghosal.

Bayesian non-parametric simultaneous quantile regression for complete and grid data. Computational Statistics & Data Analysis, 127(??):172– 186, November 2018. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947318300963.

Bhattacharya:2018:SBI

[2413] Arnab Bhattacharya and Simon P.

Wilson. Sequential Bayesian inference for static parameters in dynamic state space models. Computational Statistics & Data Analysis, 127(??): 187–203, November 2018. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947318301336.

Zhao:2018:CTM

[2414] Xin Zhao, Stuart Barber, Charles C.

Taylor, and Zoka Milan. Classification tree methods for panel data using wavelet-transformed time series. Computational Statistics & Data Analysis, 127(??):204–216, November 2018. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947318301348.

Yu:2018:ODD

[2415] Jun Yu, Xiangshun Kong, Mingyao Ai, and Kwok Leung Tsui. Optimal designs for dose-response models with linear effects of covariates. Computational Statistics & Data Analysis, 127 (??):217–228, November 2018. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947318301269.

Zhang:2018:RPB

[2416] Michael Minyi Zhang, Henry Lam, and Lizhen Lin. Robust and parallel Bayesian model selection. Computational Statistics & Data Analysis, 127 (??):229–247, November 2018. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947318301257.

Wong:2018:AND

[2417] Tony S. T. Wong, Kwok Fai Lam, and Victoria X. Zhao. Asymptotic null distribution of the modified likelihood ratio test for homogeneity in finite mixture models. Computational Statistics & Data Analysis, 127(??): 248–257, November 2018. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947318301191.

Chen:2018:JEL

[2418] Sixia Chen and David Haziza. Jackknife empirical likelihood method for multiply robust estimation with missing data. Computational Statistics & Data Analysis, 127(??):258– 268, November 2018. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947318301208.

Zhao:2018:PVC

[2419] Weihua Zhao, Xuejun Jiang, and Heng Lian. A principal varyingcoefficient model for quantile regression: Joint variable selection and dimension reduction. Computational Statistics & Data Analysis, 127(??): 269–280, November 2018. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947318301385.

Hubin:2018:MJM

[2420] Aliaksandr Hubin and Geir Storvik.

Mode jumping MCMC for Bayesian variable selection in GLMM. Computational Statistics & Data Analysis, 127 (??):281–297, November 2018. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794731830135X.

Liu:2018:WBS

[2421] Ying Liu, Yawen Liu, and Tonglin Zhang. Wald-based spatial scan statistics for cluster detection. Computational Statistics & Data Analysis, 127 (??):298–310, November 2018. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947318301403.

Anonymous:2018:EBl

[2422] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 128(??):ii–iv, December 2018. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947318302135.

Chauvet:2018:EBR

[2423] Guillaume Chauvet and Wilfried Do Paco. Exact balanced random imputation for sample survey data. Computational Statistics & Data Analysis, 128(??):1–16, December 2018. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947318301440.

Chumnaul:2018:GCI

[2424] Jularat Chumnaul and Mohammad Sepehrifar. Generalized confidence interval for the scale parameter of the power-law process with incomplete failure data. Computational Statistics & Data Analysis, 128(??): 17–33, December 2018. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947318301452.

Du:2018:FAM

[2425] Lilun Du, Wei Lan, Ronghua Luo, and Pingshou Zhong. Factor-adjusted multiple testing of correlations. Computational Statistics & Data Analysis, 128 (??):34–47, December 2018. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947318301397.

Lenis:2018:MMM

[2426] David Lenis, Benjamin Ackerman, and Elizabeth A. Stuart. Measuring model misspecification: Application to propensity score methods with complex survey data. Computational Statistics & Data Analysis, 128 (??):48–57, December 2018. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947318301105.

Zhang:2018:QLE

[2427] Hongfan Zhang. Quasi-likelihood estimation of the single index conditional variance model. Computational Statistics & Data Analysis, 128 (??):58–72, December 2018. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947318301464.

Beh:2018:CAF

[2428] Eric J. Beh, Rosaria Lombardo, and Gianmarco Alberti. Correspondence analysis and the Freeman– Tukey statistic: a study of archaeological data. Computational Statistics & Data Analysis, 128(??):73– 86, December 2018. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947318301580.

Zhang:2018:EHT

[2429] Jun Zhang, Zhenghui Feng, and Heng Peng. Estimation and hypothesis test for partial linear multiplicative models. Computational Statistics & Data Analysis, 128(??):87– 103, December 2018. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947318301634.

Das:2018:IZI

[2430] Ujjwal Das and Kalyan Das. Inference on zero inflated ordinal models with semiparametric link. Computational Statistics & Data Analysis, 128 (??):104–115, December 2018. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947318301622.

Chen:2018:STJ

[2431] Chyong-Mei Chen, Pao sheng Shen, and Yi-Kuan Tseng. Semiparametric transformation joint models for longitudinal covariates and intervalcensored failure time. Computational Statistics & Data Analysis, 128(??): 116–127, December 2018. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947318301671.

Lee:2018:AMS

[2432] Xing Ju Lee, Markus Hainy, James P. McKeone, Christopher C. Drovandi, and Anthony N. Pettitt. ABC model selection for spatial extremes models applied to South Australian maximum temperature data. Computational Statistics & Data Analysis, 128 (??):128–144, December 2018. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947318301658.

Niu:2018:PDB

[2433] Cuizhen Niu, Xu Guo, Yong Li, and Lixing Zhu. Pairwise distancebased tests for conditional symmetry. Computational Statistics & Data Analysis, 128(??):145–162, December 2018. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947318301646.

Santitissadeekorn:2018:SDA

[2434] N. Santitissadeekorn, M. B. Short, and D. J. B. Lloyd. Sequential data assimilation for 1D self-exciting processes with application to urban crime data. Computational Statistics & Data Analysis, 128(??):163– 183, December 2018. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947318301609.

Archimbaud:2018:IMO

[2435] Aurore Archimbaud, Klaus Nordhausen, and Anne Ruiz-Gazen. ICS for multivariate outlier detection with application to quality control. Computational Statistics & Data Analysis, 128 (??):184–199, December 2018. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947318301579.

Celisse:2018:NEA

[2436] A. Celisse, G. Marot, M. Pierre-Jean, and G. J. Rigaill. New efficient algorithms for multiple change-point detection with reproducing kernels. Computational Statistics & Data Analysis, 128 (??):200–220, December 2018. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947318301683.

Bouranis:2018:MCG

[2437] Lampros Bouranis, Nial Friel, and Florian Maire. Model comparison for Gibbs random fields using noisy reversible jump Markov chain Monte Carlo. Computational Statistics & Data Analysis, 128(??):221– 241, December 2018. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947318301713.

Lu:2018:FSM

[2438] Jun Lu and Lu Lin. Feature screening for multi-response varying coefficient models with ultrahigh dimensional predictors. Computational Statistics & Data Analysis, 128(??): 242–254, December 2018. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947318301476.

Sewell:2018:VDT

[2439] Daniel K. Sewell. Visualizing data through curvilinear representations of matrices. Computational Statistics & Data Analysis, 128(??):255– 270, December 2018. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947318301762.

Ong:2018:LFI

[2440] Victor M.-H. Ong, David J. Nott, Minh-Ngoc Tran, Scott A. Sisson, and Christopher C. Drovandi. Likelihoodfree inference in high dimensions with synthetic likelihood. Computational Statistics & Data Analysis, 128(??): 271–291, December 2018. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947318301749.

Li:2018:EAS

[2441] Peili Li and Yunhai Xiao. An efficient algorithm for sparse inverse covariance matrix estimation based on dual formulation. Computational Statistics & Data Analysis, 128(??): 292–307, December 2018. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947318301774.

Chiou:2018:PTG

[2442] Sy Han Chiou, Jing Qian, Elizabeth Mormino, and Rebecca A. Betensky. Permutation tests for general dependent truncation. Computational Statistics & Data Analysis, 128(??):

308–324, December 2018. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947318301786.

Guo:2018:SDR

[2443] Xu Guo, Yun Fang, Xuehu Zhu, Wangli Xu, and Lixing Zhu. Semiparametric double robust and efficient estimation for mean functionals with response missing at random. Computational Statistics & Data Analysis, 128 (??):325–339, December 2018. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794731830183X.

Fry:2018:GBS

[2444] J. T. Fry, Matt Slifko, and Scotland Leman. Generalized biplots for stress-based multidimensionally scaled projections. Computational Statistics & Data Analysis, 128(??):340– 353, December 2018. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947318301865.

Gamage:2018:GFP

[2445] Prabhashi W. Withana Gamage, Christopher S. McMahan, Lianming Wang, and Wanzhu Tu. A Gamma-frailty proportional hazards model for bivariate interval-censored data. Computational Statistics & Data Analysis, 128(??):354–366, December 2018. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947318301828.

Weiss:2018:GCA

[2446] Christian H. Weiß, Detlef Steuer, Carsten Jentsch, and Murat Caner Testik. Guaranteed conditional ARL performance in the presence of autocorrelation. Computational Statistics & Data Analysis, 128(??):367– 379, December 2018. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947318301798.

Sikora:2018:RSA

[2447] Grzegorz Sikora, Agnieszka Wylomańska, and Diego Krapf. Recurrence statistics for anomalous diffusion regime change detection. Computational Statistics & Data Analysis, 128(??): 380–394, December 2018. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947318301804.

Wang:2018:OBC

[2448] Ketong Wang and Michael D. Porter.

Optimal Bayesian clustering using nonnegative matrix factorization. Computational Statistics & Data Analysis, 128 (??):395–411, December 2018. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947318301853.

Anonymous:2019:EBa

[2449] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 129(??):ii–iv, January 2019. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016794731830224X.

Lin:2019:BMS

[2450] L. Lin and D. K. H. Fong. Bayesian multidimensional scaling procedure with variable selection. Computational Statistics & Data Analysis, 129 (??):1–13, January 2019. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947318301737.

Li:2019:BFJ

[2451] Kan Li and Sheng Luo. Bayesian functional joint models for multivariate longitudinal and time-to-event data. Computational Statistics & Data Analysis, 129(??):14–29, January 2019. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947318301816.

Dickhaus:2019:SSI

[2452] Thorsten Dickhaus and Natalia Sirotko-Sibirskaya. Simultaneous statistical inference in dynamic factor models: Chi-square approximation and model-based bootstrap. Computational Statistics & Data Analysis, 129 (??):30–46, January 2019. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947318301956.

Manghi:2019:GAP

[2453] Roberto F. Manghi, Francisco José A. Cysneiros, and Gilberto A. Paula. Generalized additive partial linear models for analyzing correlated data. Computational Statistics & Data Analysis, 129(??):47–60, January 2019. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947318301889.

Hinoveanu:2019:BLB

[2454] Laurentiu C. Hinoveanu, Fabrizio Leisen, and Cristiano Villa. Bayesian loss-based approach to change point analysis. Computational Statistics & Data Analysis, 129(??):61– 78, January 2019. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947318301919.

Zang:2019:RTG

[2455] Yong Zang, Wing Kam Fung, Sha Cao, Hon Keung Tony Ng, and Chi Zhang. Robust tests for gene-environment interaction in case-control and caseonly designs. Computational Statistics & Data Analysis, 129(??):79– 92, January 2019. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794731830197X.

Yu:2019:REC

[2456] Dalei Yu, Chang Ding, Na He, Ruiwu Wang, Xiaohua Zhou, and Lei Shi. Robust estimation and confidence interval in meta-regression models. Computational Statistics & Data Analysis, 129 (??):93–118, January 2019. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947318301932.

Hsu:2019:GAL

[2457] Hsiang-Ling Hsu, Yuan chin Ivan Chang, and Ray-Bing Chen. Greedy active learning algorithm for logistic regression models. Computational Statistics & Data Analysis, 129(??):119– 134, January 2019. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947318301968.

Ma:2019:QRF

[2458] Haiqiang Ma, Ting Li, Hongtu Zhu, and Zhongyi Zhu. Quantile regression for functional partially linear model in ultra-high dimensions. Computational Statistics & Data Analysis, 129 (??):135–147, January 2019. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947318301439.

Anonymous:2019:EBb

[2459] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 130(??):ii–iv, February 2019. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947318302391.

Flores-Agreda:2019:BEU

[2460] Daniel Flores-Agreda and Eva Cantoni. Bootstrap estimation of uncertainty in prediction for generalized linear mixed models. Computational Statistics & Data Analysis, 130 (??):1–17, February 2019. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947318301890.

Wei:2019:MGH

[2461] Yuhong Wei, Yang Tang, and Paul D.

McNicholas. Mixtures of generalized hyperbolic distributions and mixtures of skew-t distributions for model-based clustering with incomplete data. Computational Statistics & Data Analysis, 130(??):18–41, February 2019. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947318301993.

Cui:2019:PDS

[2462] Xia Cui, Weihua Zhao, Heng Lian, and Hua Liang. Pursuit of dynamic structure in quantile additive models with longitudinal data. Computational Statistics & Data Analysis, 130 (??):42–60, February 2019. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947318302007.

Samejima:2019:DWC

[2463] Kim Samejima, Pedro A. Morettin, and João Ricardo Sato. Directed wavelet covariance. Computational Statistics & Data Analysis, 130(??): 61–79, February 2019. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947318302093.

Bolin:2019:LGR

[2464] David Bolin, Jonas Wallin, and Finn Lindgren. Latent Gaussian random field mixture models. Computational Statistics & Data Analysis, 130(??): 80–93, February 2019. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947318301907.

Hu:2019:TSE

[2465] Lixia Hu, Tao Huang, and Jinhong You. Two-step estimation of timevarying additive model for locally stationary time series. Computational Statistics & Data Analysis, 130(??): 94–110, February 2019. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947318302068.

Trevino:2019:RWF

[2466] E. S. García Treviño, V. Alarcón Aquino, and J. A. Barria. The radial wavelet frame density estimator. Computational Statistics & Data Analysis, 130(??):111– 139, February 2019. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947318302044.

Anonymous:2019:EBc

[2467] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 131(??):ii–iv, ???? 2019. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947318302615.

Lee:2019:IPS

[2468] Jae C. Lee, Erricos John Kontoghiorghes, Ana Colubi, and Byeong Park. Interview of Professor Stanley Azen, founder of Computational Statistics and Data Analysis (CSDA). Computational Statistics & Data Analysis, 131(??):1, ???? 2019. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947318302512.

Afifi:2019:SAM

[2469] Abdelmonem A. Afifi. Stan azen: a master of numbers and notes. Computational Statistics & Data Analysis, 131(??):2–9, ???? 2019. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794731830207X.

Ferraty:2019:ESI

[2470] Frederic Ferraty, Piotr Kokoszka, Jane-Ling Wang, and Yichao Wu. Editorial for the special issue on high-dimensional and functional data analysis. Computational Statistics & Data Analysis, 131(??):10– 11, ???? 2019. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947318302548.

Abpeykar:2019:EDF

[2471] Shadi Abpeykar, Mehdi Ghatee, and Hadi Zare. Ensemble decision forest of RBF networks via hybrid feature clustering approach for highdimensional data classification. Computational Statistics & Data Analysis, 131(??):12–36, ???? 2019. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947318301981.

Park:2019:RPC

[2472] Yeonjoo Park and Douglas G. Simpson. Robust probabilistic classification applicable to irregularly sampled functional data. Computational Statistics & Data Analysis, 131(??): 37–49, ???? 2019. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947318301841.

Dai:2019:DOM

[2473] Wenlin Dai and Marc G. Genton. Directional outlyingness for multivariate functional data. Computational Statistics & Data Analysis, 131(??): 50–65, ???? 2019. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016794731830077X.

Martinez-Hernandez:2019:RDB

[2474] Israel Martínez-Hernández, Marc G. Genton, and Graciela González-Farías. Robust depth-based estimation of the functional autoregressive model. Computational Statistics & Data Analysis, 131(??):66–79, ???? 2019. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947318301415.

Wang:2019:GPM

[2475] Bo Wang and Aiping Xu. Gaussian process methods for nonparametric functional regression with mixed predictors. Computational Statistics & Data Analysis, 131(??):80– 90, ???? 2019. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947318301750.

Febrero-Bande:2019:EIP

[2476] Manuel Febrero-Bande, Pedro Galeano, and Wenceslao González-Manteiga. Estimation, imputation and prediction for the functional linear model with scalar response with responses missing at random. Computational Statistics & Data Analysis, 131(??): 91–103, ???? 2019. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947318301725.

Liebl:2019:POF

[2477] Dominik Liebl and Stefan Rameseder.

Partially observed functional data: the case of systematically missing parts. Computational Statistics & Data Analysis, 131(??):104– 115, ???? 2019. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947318301944.

Ahmad:2019:STR

[2478] M. Rauf Ahmad. A significance test of the RV coefficient in high dimensions. Computational Statistics & Data Analysis, 131(??):116– 130, ???? 2019. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947318302536.

Wong:2019:NOR

[2479] Raymond K. W. Wong and Xiaoke Zhang. Nonparametric operatorregularized covariance function estimation for functional data. Computational Statistics & Data Analysis, 131 (??):131–144, ???? 2019. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947318301221.

Zhang:2019:MFD

[2480] Yaohua Zhang, Jian Zou, Nalini Ravishanker, and Aerambamoorthy Thavaneswaran. Modeling financial durations using penalized estimating functions. Computational Statistics & Data Analysis, 131(??):145– 158, ???? 2019. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947318302032.

Fu:2019:MBC

[2481] Eric Fu and Nancy Heckman. Modelbased curve registration via stochastic approximation EM algorithm. Computational Statistics & Data Analysis, 131(??):159–175, ???? 2019. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947318301543.

French:2019:QRH

[2482] Joshua French, Piotr Kokoszka, Stilian Stoev, and Lauren Hall. Quantifying the risk of heat waves using extreme value theory and spatiotemporal functional data. Computational Statistics & Data Analysis, 131 (??):176–193, ???? 2019. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947318301701.

Sang:2019:WEL

[2483] Peijun Sang, Liangliang Wang, and Jiguo Cao. Weighted empirical likelihood inference for dynamical correlations. Computational Statistics & Data Analysis, 131(??):194– 206, ???? 2019. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947318301695.

Shen:2019:SML

[2484] Keren Shen, Jianfeng Yao, and Wai Keung Li. On a spiked model for large volatility matrix estimation from noisy high-frequency data. Computational Statistics & Data Analysis, 131 (??):207–221, ???? 2019. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947318301427.

Yue:2019:TSS

[2485] Mu Yue, Jialiang Li, and Ming-Yen Cheng. Two-step sparse boosting for high-dimensional longitudinal data with varying coefficients. Computational Statistics & Data Analysis, 131 (??):222–234, ???? 2019. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947318302470.

Anonymous:2019:EBd

[2486] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 132(??):ii–iv, ???? 2019. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016794731930012X.

Cheng:2019:SIB

[2487] Ming-Yen Cheng, Michele Guindani, Jae Won Lee, Yi Li, and Catherine Chunling Liu. Special issue on biostatistics. Computational Statistics & Data Analysis, 132(??):1– 2, ???? 2019. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947319300052.

Zhang:2019:NTF

[2488] Jin-Ting Zhang, Ming-Yen Cheng, Hau-Tieng Wu, and Bu Zhou. A new test for functional one-way ANOVA with applications to ischemic heart screening. Computational Statistics & Data Analysis, 132(??):3– 17, ???? 2019. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947318301130.

Sambucini:2019:BPM

[2489] Valeria Sambucini. Bayesian predictive monitoring with bivariate binary outcomes in phase II clinical trials. Computational Statistics & Data Analysis, 132(??):18–30, ???? 2019. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947318301610.

Godolphin:2019:RAT

[2490] P. J. Godolphin and E. J. Godolphin. Robust assessment of twotreatment higher-order cross-over designs against missing values. Computational Statistics & Data Analysis, 132(??):31–45, ???? 2019. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794731830166X.

Banerjee:2019:SCS

[2491] Sayantan Banerjee, Rehan Akbani, and Veerabhadran Baladandayuthapani. Spectral clustering via sparse graph structure learning with application to proteomic signaling networks in cancer. Computational Statistics & Data Analysis, 132(??): 46–69, ???? 2019. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947318301920.

Chowdhury:2019:NIT

[2492] Shrabanti Chowdhury, Ram C. Tiwari, and Samiran Ghosh. Noninferiority testing for risk ratio, odds ratio and number needed to treat in three-arm trial. Computational Statistics & Data Analysis, 132(??): 70–83, ???? 2019. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947318302019.

Heyard:2019:QPM

[2493] Rachel Heyard and Leonhard Held. The quantile probability model. Computational Statistics & Data Analysis, 132(??):84–99, ???? 2019. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947318302056.

He:2019:CIS

[2494] Kevin He, Jian Kang, Hyokyoung G. Hong, Ji Zhu, Yanming Li, Huazhen Lin, Han Xu, and Yi Li. Covarianceinsured screening. Computational Statistics & Data Analysis, 132(??): 100–114, ???? 2019. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947318302202.

Yu:2019:ELI

[2495] Xue Yu and Yichuan Zhao. Empirical likelihood inference for semiparametric transformation models with length-biased sampling. Computational Statistics & Data Analysis, 132 (??):115–125, ???? 2019. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947318302573.

Overstall:2019:BPP

[2496] Antony M. Overstall, David C. Woods, and Kieran J. Martin. Bayesian prediction for physical models with application to the optimization of the synthesis of pharmaceutical products using chemical kinetics. Computational Statistics & Data Analysis, 132 (??):126–142, ???? 2019. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794731830272X.

Einbeck:2019:ESI

[2497] Jochen Einbeck, John Hinde, Salvatore Ingrassia, Tsung-I. Lin, and Paul D. McNicholas. Editorial for the 4th special issue on advances in mixture models. Computational Statistics & Data Analysis, 132(??): 143–144, ???? 2019. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947318302858.

Morris:2019:ACO

[2498] Katherine Morris, Antonio Punzo, Paul D. McNicholas, and Ryan P. Browne. Asymmetric clusters and outliers: Mixtures of multivariate contaminated shifted asymmetric Laplace distributions. Computational Statistics & Data Analysis, 132(??):145– 166, ???? 2019. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947318302809.

Marbac:2019:TMP

[2499] Matthieu Marbac and Vincent Vandewalle. A tractable multi-partitions clustering. Computational Statistics & Data Analysis, 132(??):167– 179, ???? 2019. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947318301592.

Wichitchan:2019:HTF

[2500] Supawadee Wichitchan, Weixin Yao, and Guangren Yang. Hypothesis testing for finite mixture models. Computational Statistics & Data Analysis, 132(??):180–189, ???? 2019. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947318301142.

Xia:2019:BAM

[2501] Ye-Mao Xia and Nian-Sheng Tang. Bayesian analysis for mixture of latent variable hidden Markov models with multivariate longitudinal data. Computational Statistics & Data Analysis, 132(??):190–211, ???? 2019. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947318301877.

Ahonen:2019:PFF

[2502] Ilmari Ahonen, Jaakko Nevalainen, and Denis Larocque. Prediction with a flexible finite mixture-of-regressions. Computational Statistics & Data Analysis, 132(??):212–224, ???? 2019. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947318300136.

Anonymous:2019:EBe

[2503] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 133(??):ii–iv, May 2019. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947319300428.

Zhang:2019:NVB

[2504] Chun-Xia Zhang, Shuang Xu, and Jiang-She Zhang. A novel variational Bayesian method for variable selection in logistic regression models. Computational Statistics & Data Analysis, 133(??):1–19, May 2019. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947318302081.

Bogomolov:2019:ORU

[2505] Marina Bogomolov and Ori Davidov. Order restricted univariate and multivariate inference with adjustment for covariates in partially linear models. Computational Statistics & Data Analysis, 133(??):20– 27, May 2019. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947318302020.

Lee:2019:MSS

[2506] Myeonggyun Lee and Inkyung Jung.

Modified spatial scan statistics using a restricted likelihood ratio for ordinal outcome data. Computational Statistics & Data Analysis, 133 (??):28–39, May 2019. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947318302330.

Zhu:2019:BSR

[2507] Rui Zhu and Subhashis Ghosal. Bayesian semiparametric ROC surface estimation under verification bias. Computational Statistics & Data Analysis, 133(??):40–52, May 2019. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0167947318302317.

Kirschstein:2019:FSR

[2508] Thomas Kirschstein, Steffen Liebscher, Giuseppe Pandolfo, Giovanni C. Porzio, and Giancarlo Ragozini. On finite-sample robustness of directional location estimators. Computational Statistics & Data Analysis, 133(??): 53–75, May 2019. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947318302196.

Chen:2019:SGI

[2509] Ying Chen, Linlin Niu, Ray-Bing Chen, and Qiang He. Sparse-Group Independent Component Analysis with application to yield curves prediction. Computational Statistics & Data Analysis, 133(??):76–89, May 2019. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947318302184.

Wisniowski:2019:HMF

[2510] Arkadiusz Wiśniowski, Jakub Bijak, Jonathan J. Forster, and Peter W. F. Smith. Hierarchical model for forecasting the outcomes of binary referenda. Computational Statistics & Data Analysis, 133(??): 90–103, May 2019. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947318302354.

Lee:2019:CMD

[2511] Hangsuck Lee, Jae Youn Ahn, and Bangwon Ko. Construction of multiple decrement tables under generalized fractional age assumptions. Computational Statistics & Data Analysis, 133(??):104–119, May 2019. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947318302329.

Godolphin:2019:TLF

[2512] J. D. Godolphin. Two-level factorial and fractional factorial replicates in blocks of size two. Computational Statistics & Data Analysis, 133 (??):120–137, May 2019. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947318302342.

Sheng:2019:SVS

[2513] Ying Sheng and Qihua Wang. Simultaneous variable selection and class fusion with penalized distance criterion based classifiers. Computational Statistics & Data Analysis, 133(??): 138–152, May 2019. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947318302305.

Li:2019:CST

[2514] Shuwei Li, Tao Hu, Xingqiu Zhao, and Jianguo Sun. A class of semiparametric transformation cure models for interval-censored failure time data. Computational Statistics & Data Analysis, 133(??):153– 165, May 2019. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947318302445.

Zhang:2019:BCS

[2515] Shibin Zhang. Bayesian copula spectral analysis for stationary time series. Computational Statistics & Data Analysis, 133(??):166– 179, May 2019. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947318302457.

Fang:2019:NAC

[2516] Fang Fang and Yuanyuan Chen. A new approach for credit scoring by directly maximizing the Kolmogorov– Smirnov statistic. Computational Statistics & Data Analysis, 133(??): 180–194, May 2019. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947318302494.

Karavarsamis:2019:TSA

[2517] N. Karavarsamis and R. M. Huggins. Two-stage approaches to the analysis of occupancy data II. The heterogeneous model and conditional likelihood. Computational Statistics & Data Analysis, 133(??):195–207, May 2019. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947318302469.

Tang:2019:FSU

[2518] Niansheng Tang, Linli Xia, and Xiaodong Yan. Feature screening in ultrahigh-dimensional partially linear models with missing responses at random. Computational Statistics & Data Analysis, 133(??):208– 227, May 2019. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947318302482.

Jhong:2019:SEQ

[2519] Jae-Hwan Jhong and Ja-Yong Koo.

Simultaneous estimation of quantile regression functions using B-splines and total variation penalty. Computational Statistics & Data Analysis, 133(??):228–244, May 2019. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947318302500.

Boonstra:2019:DPI

[2520] Philip S. Boonstra, Ryan P. Barbaro, and Ananda Sen. Default priors for the intercept parameter in logistic regressions. Computational Statistics & Data Analysis, 133(??): 245–256, May 2019. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947318302755.

Wang:2019:MRS

[2521] Kangning Wang, Shaomin Li, Xiaofei Sun, and Lu Lin. Modal regression statistical inference for longitudinal data semivarying coefficient models: Generalized estimating equations, empirical likelihood and variable selection. Computational Statistics & Data Analysis, 133(??):257–276, May 2019. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016794731830255X.

Lennon:2019:EDG

[2522] Hannah Lennon and Jingsong Yuan. Estimation of a digitised Gaussian ARMA model by Monte Carlo expectation maximisation. Computational Statistics & Data Analysis, 133 (??):277–284, May 2019. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947318302767.

Shi:2019:BMS

[2523] Guiling Shi, Chae Young Lim, and Tapabrata Maiti. Bayesian model selection for generalized linear models using non-local priors. Computational Statistics & Data Analysis, 133(??):285–296, May 2019. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947318302524.

Anonymous:2019:EBf

[2524] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 134(??):ii–iv, June 2019. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947319300489.

Liu:2019:MCD

[2525] Li Liu and Liming Xiang. Missing covariate data in generalized linear mixed models with distributionfree random effects. Computational Statistics & Data Analysis, 134(??): 1–16, June 2019. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947318302561.

Yue:2019:RAT

[2526] Lili Yue, Gaorong Li, Heng Lian, and Xiang Wan. Regression adjustment for treatment effect with multicollinearity in high dimensions. Computational Statistics & Data Analysis, 134(??):17–35, June 2019. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947318302780.

DeCanditiis:2019:SNR

[2527] Daniela De Canditiis and Italia De Feis. Simultaneous nonparametric regression in RADWT dictionaries. Computational Statistics & Data Analysis, 134(??):36–57, June 2019. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947318302792.

Song:2019:DSB

[2528] Junmo Song and Changryong Baek.

Detecting structural breaks in realized volatility. Computational Statistics & Data Analysis, 134(??): 58–75, June 2019. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794731830286X.

Hwang:2019:BPS

[2529] Youngdeok Hwang, Hang J. Kim, Won Chang, Kyongmin Yeo, and Yongku Kim. Bayesian pollution source identification via an inverse physics model. Computational Statistics & Data Analysis, 134(??):76– 92, June 2019. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947318302822.

Rodrigues:2019:SFB

[2530] T. Rodrigues, J.-L. Dortet-Bernadet, and Y. Fan. Simultaneous fitting of Bayesian penalised quantile splines. Computational Statistics & Data Analysis, 134(??):93– 109, June 2019. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947318302883.

Huang:2019:NPL

[2531] Lei Huang, Hui Jiang, and Huixia Wang. A novel partial-linear singleindex model for time series data. Computational Statistics & Data Analysis, 134(??):110–122, June 2019. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947318302913.

Yoshida:2019:RSM

[2532] Takuma Yoshida and Kanta Naito.

Regression with stagewise minimization on risk function. Computational Statistics & Data Analysis, 134(??): 123–143, June 2019. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947318302901.

Agostinelli:2019:IRE

[2533] Claudio Agostinelli, Marina Valdora, and Victor J. Yohai. Initial robust estimation in generalized linear models. Computational Statistics & Data Analysis, 134(??):144– 156, June 2019. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947318302895.

Singh:2019:EBD

[2534] Rakhi Singh and Siuli Mukhopadhyay. Exact Bayesian designs for count time series. Computational Statistics & Data Analysis, 134(??): 157–170, June 2019. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947318302871.

Cevallos-Valdiviezo:2019:FCR

[2535] Holger Cevallos-Valdiviezo and Stefan Van Aelst. Fast computation of robust subspace estimators. Computational Statistics & Data Analysis, 134 (??):171–185, June 2019. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947319300015.

Qiu:2019:SNA

[2536] Zhiping Qiu, Limin Peng, Amita Manatunga, and Ying Guo. A smooth nonparametric approach to determining cut-points of a continuous scale. Computational Statistics & Data Analysis, 134(??):186–210, June 2019. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947318302779.

Anonymous:2019:EBg

[2537] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 135(??):ii–iv, July 2019. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947319300647.

Chen:2019:AMS

[2538] Yichen Chen and Somnath Datta. Adjustments of multi-sample U-statistics to right censored data and confounding covariates. Computational Statistics & Data Analysis, 135(??): 1–14, July 2019. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016794731930026X.

Chaudhuri:2019:FAC

[2539] Arin Chaudhuri and Wenhao Hu. A fast algorithm for computing distance correlation. Computational Statistics & Data Analysis, 135(??): 15–24, July 2019. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947319300313.

Li:2019:VSC

[2540] Haocheng Li, Di Shu, Wenqing He, and Grace Y. Yi. Variable selection via the composite likelihood method for multilevel longitudinal data with missing responses and covariates. Computational Statistics & Data Analysis, 135(??):25–34, July 2019. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947319300258.

Zamanzade:2019:EMR

[2541] Elham Zamanzade, Afshin Parvardeh, and Majid Asadi. Estimation of mean residual life based on ranked set sampling. Computational Statistics & Data Analysis, 135(??):35– 55, July 2019. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947319300064.

Gregory:2019:SAB

[2542] Alastair Gregory. A streaming algorithm for bivariate empirical copulas. Computational Statistics & Data Analysis, 135(??):56– 69, July 2019. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947319300295.

Derumigny:2019:CPV

[2543] Alexis Derumigny and Jean-David Fermanian. A classification point-of-view about conditional Kendall’s tau. Computational Statistics & Data Analysis, 135(??):70–94, July 2019. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947319300271.

Choi:2019:MBB

[2544] Ji-Eun Choi and Dong Wan Shin. Moving block bootstrapping for a CUSUM test for correlation change. Computational Statistics & Data Analysis, 135(??):95–106, July 2019. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947319300301.

Baey:2019:ADL

[2545] Charlotte Baey, Paul-Henry Cournède, and Estelle Kuhn. Asymptotic distribution of likelihood ratio test statistics for variance components in nonlinear mixed effects models. Computational Statistics & Data Analysis, 135(??):107–122, July 2019. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947319300283.

Anonymous:2019:EBh

[2546] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 136(??):ii–iv, August 2019. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947319300829.

Ditzhaus:2019:WBL

[2547] Marc Ditzhaus and Markus Pauly.

Wild bootstrap logrank tests with broader power functions for testing superiority. Computational Statistics & Data Analysis, 136(??):1– 11, August 2019. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947319300362.

Yu:2019:SWE

[2548] Dengdeng Yu, Li Zhang, Ivan Mizera, Bei Jiang, and Linglong Kong. Sparse wavelet estimation in quantile regression with multiple functional predictors. Computational Statistics & Data Analysis, 136(??): 12–29, August 2019. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947318302810.

Geraci:2019:MEN

[2549] Marco Geraci. Modelling and estimation of nonlinear quantile regression with clustered data. Computational Statistics & Data Analysis, 136 (??):30–46, August 2019. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947318302846.

Lee:2019:MML

[2550] Keunbaik Lee and Yongsung Joo.

Marginalized models for longitudinal count data. Computational Statistics & Data Analysis, 136(??):47– 58, August 2019. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947319300027.

Xiu:2019:ADM

[2551] Xianchao Xiu, Wanquan Liu, Ling Li, and Lingchen Kong. Alternating direction method of multipliers for nonconvex fused regression problems. Computational Statistics & Data Analysis, 136(??):59–71, August 2019. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947319300039.

Chakraborty:2019:GLP

[2552] Sounak Chakraborty and Aurelie C. Lozano. A graph Laplacian prior for Bayesian variable selection and grouping. Computational Statistics & Data Analysis, 136(??):72– 91, August 2019. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947319300040.

Golightly:2019:CPM

[2553] Andrew Golightly, Emma Bradley, Tom Lowe, and Colin S. Gillespie. Correlated pseudo-marginal schemes for time-discretised stochastic kinetic models. Computational Statistics & Data Analysis, 136(??): 92–107, August 2019. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947319300076.

Mahmoud:2019:SSM

[2554] Hamdy F. F. Mahmoud and Inyoung Kim. Semiparametric spatial mixed effects single index models. Computational Statistics & Data Analysis, 136 (??):108–122, August 2019. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947319300222.

Wang:2019:BHM

[2555] Xia Wang, Ali Shojaie, and Jian Zou. Bayesian hidden Markov models for dependent large-scale multiple testing. Computational Statistics & Data Analysis, 136(??):123– 136, August 2019. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947319300234.

Anonymous:2019:EBi

[2556] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 137 (??):ii–iv, September 2019. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947319301021.

Castillo-Paez:2019:NBM

[2557] Sergio Castillo-Páez, Rubén Fernández-Casal, and Pilar García-Soidán. A nonparametric bootstrap method for spatial data. Computational Statistics & Data Analysis, 137(??):1– 15, September 2019. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947319300325.

Ippel:2019:OEI

[2558] L. Ippel, M. C. Kaptein, and J. K. Vermunt. Online estimation of individuallevel effects using streaming shrinkage factors. Computational Statistics & Data Analysis, 137(??):16– 32, September 2019. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947319300246.

Bui:2019:EAA

[2559] Anh Tuan Bui and Daniel W. Apley. An exploratory analysis approach for understanding variation in stochastic textured surfaces. Computational Statistics & Data Analysis, 137(??): 33–50, September 2019. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947319300349.

Thaden:2019:MEP

[2560] Hauke Thaden, Nadja Klein, and Thomas Kneib. Multivariate effect priors in bivariate semiparametric recursive Gaussian models. Computational Statistics & Data Analysis, 137 (??):51–66, September 2019. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947318302834.

Azzimonti:2019:HEP

[2561] Laura Azzimonti, Giorgio Corani, and Marco Zaffalon. Hierarchical estimation of parameters in Bayesian networks. Computational Statistics & Data Analysis, 137(??):67– 91, September 2019. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947319300519.

Fuentes-Garcia:2019:MPC

[2562] Ruth Fuentes-García, Ramsés H.

Mena, and Stephen G. Walker. Modal posterior clustering motivated by Hopfield’s network. Computational Statistics & Data Analysis, 137(??):92– 100, September 2019. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794731930057X.

Gromping:2019:AGG

[2563] Ulrike Grömping and Roberto Fontana.

An algorithm for generating good mixed level factorial designs. Computational Statistics & Data Analysis, 137(??):101–114, September 2019. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0167947319300350.

Litvinenko:2019:LAH

[2564] Alexander Litvinenko, Ying Sun, Marc G. Genton, and David E. Keyes. Likelihood approximation with hierarchical matrices for large spatial datasets. Computational Statistics & Data Analysis, 137(??):115– 132, September 2019. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947319300374.

Bischofberger:2019:CSF

[2565] Stephan M. Bischofberger, Munir Hiabu, Enno Mammen, and Jens Perch Nielsen. A comparison of in-sample forecasting methods. Computational Statistics & Data Analysis, 137(??): 133–154, September 2019. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947319300581.

He:2019:MMV

[2566] Shengmei He, Shuangge Ma, and Wangli Xu. A modified meanvariance feature-screening procedure for ultrahigh-dimensional discriminant analysis. Computational Statistics & Data Analysis, 137(??):155– 169, September 2019. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947319300386.

Li:2019:MPL

[2567] Jinqing Li and Jun Ma. Maximum penalized likelihood estimation of additive hazards models with partly interval censoring. Computational Statistics & Data Analysis, 137(??):170– 180, September 2019. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947319300593.

Griffin:2019:LAD

[2568] Maryclare Griffin and Peter D. Hoff. Lasso ANOVA decompositions for matrix and tensor data. Computational Statistics & Data Analysis, 137(??):

181–194, September 2019. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947319300520.

Matsuda:2019:EBM

[2569] Takeru Matsuda and Fumiyasu Komaki. Empirical Bayes matrix completion. Computational Statistics & Data Analysis, 137(??):195– 210, September 2019. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947319300556.

Muller:2019:DMU

[2570] Dominik Müller and Claudia Czado.

Dependence modelling in ultra high dimensions with vine copulas and the graphical lasso. Computational Statistics & Data Analysis, 137(??):211– 232, September 2019. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947319300568.

Liu:2019:BIM

[2571] Baisen Liu, Liangliang Wang, Yunlong Nie, and Jiguo Cao. Bayesian inference of mixed-effects ordinary differential equations models using heavytailed distributions. Computational Statistics & Data Analysis, 137(??): 233–246, September 2019. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794731930060X.

Berge:2019:LTB

[2572] Laurent R. Bergé, Charles Bouveyron, Marco Corneli, and Pierre Latouche. The latent topic block model for the co-clustering of textual interaction data. Computational Statistics & Data Analysis, 137(??):247– 270, September 2019. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947319300726.

Liu:2019:ESI

[2573] Jicai Liu, Peirong Xu, and Heng Lian. Estimation for single-index models via martingale difference divergence. Computational Statistics & Data Analysis, 137(??):271– 284, September 2019. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947319300751.

Marinas-Collado:2019:PGP

[2574] Irene Mariñas-Collado, Adrian Bowman, and Vincent Macaulay. A phylogenetic Gaussian process model for the evolution of curves embedded in d-dimensions. Computational Statistics & Data Analysis, 137(??):285– 298, September 2019. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947319300696.

Anonymous:2019:EBj

[2575] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 138(??):ii–iv, October 2019. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947319301227.

Guo:2019:MCG

[2576] Xu Guo, Lianlian Song, Yun Fang, and Lixing Zhu. Model checking for general linear regression with nonignorable missing response. Computational Statistics & Data Analysis, 138 (??):1–12, October 2019. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947319300763.

Ng:2019:FSM

[2577] Kenyon Ng, Berwin A. Turlach, and Kevin Murray. A flexible sequential Monte Carlo algorithm for parametric constrained regression. Computational Statistics & Data Analysis, 138 (??):13–26, October 2019. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947319300787.

Xie:2019:GFT

[2578] Chuanlong Xie and Lixing Zhu.

A goodness-of-fit test for variableadjusted models. Computational Statistics & Data Analysis, 138(??): 27–48, October 2019. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947319300337.

Wagner:2019:NRL

[2579] Heiko Wagner and Alois Kneip.

Nonparametric registration to lowdimensional function spaces. Computational Statistics & Data Analysis, 138 (??):49–63, October 2019. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947319300714.

Li:2019:MCR

[2580] Lingzhu Li, Sung Nok Chiu, and Lixing Zhu. Model checking for regressions: an approach bridging between local smoothing and global smoothing methods. Computational Statistics & Data Analysis, 138(??): 64–82, October 2019. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947319300702.

Zhang:2019:NRA

[2581] Yuexia Zhang, Guoyou Qin, Zhongyi Zhu, and Wanghong Xu. A novel robust approach for analysis of longitudinal data. Computational Statistics & Data Analysis, 138(??):83– 95, October 2019. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947319300866.

Vexler:2019:ELP

[2582] Albert Vexler, Li Zou, and Alan D. Hutson. The empirical likelihood prior applied to bias reduction of general estimating equations. Computational Statistics & Data Analysis, 138 (??):96–106, October 2019. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947319300854.

LeSage:2019:MCM

[2583] James P. LeSage, Yao-Yu Chih, and Colin Vance. Markov Chain Monte Carlo estimation of spatial dynamic panel models for large samples. Computational Statistics & Data Analysis, 138(??):107– 125, October 2019. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947319300878.

Caterina:2019:LAW

[2584] Claudia Di Caterina and Ioannis Kosmidis. Location-adjusted Wald statistics for scalar parameters. Computational Statistics & Data Analysis, 138 (??):126–142, October 2019. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794731930088X.

Cheng:2019:EIN

[2585] Jing Cheng and Ngai Hang Chan. Efficient inference for nonlinear state space models: an automatic sample size selection rule. Computational Statistics & Data Analysis, 138 (??):143–154, October 2019. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947319300775.

Liang:2019:MEL

[2586] Wei Liang, Hongsheng Dai, and Shuyuan He. Mean empirical likelihood. Computational Statistics & Data Analysis, 138(??):155– 169, October 2019. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794731930091X.

Mazo:2019:CKE

[2587] Gildas Mazo and Yaroslav Averyanov. Constraining kernel estimators in semiparametric copula mixture models. Computational Statistics & Data Analysis, 138(??):170– 189, October 2019. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947319300945.

Johnson:2019:ERW

[2588] Brad C. Johnson and Steve Kirkland. Estimating random walk centrality in networks. Computational Statistics & Data Analysis, 138(??): 190–200, October 2019. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947319300933.

Sinha:2019:EMV

[2589] Shyamalendu Sinha and Jeffrey D. Hart. Estimating the mean and variance of a high-dimensional normal distribution using a mixture prior. Computational Statistics & Data Analysis, 138(??):201–221, October 2019. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0167947319300908.

Lin:2019:DDM

[2590] Chang-Yun Lin and Po Yang. Datadriven multistratum designs with the generalized Bayesian D–D criterion for highly uncertain models. Computational Statistics & Data Analysis, 138 (??):222–238, October 2019. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794731930074X.

Liu:2019:SAH

[2591] Lili Liu and Lu Lin. Subgroup analysis for heterogeneous additive partially linear models and its application to car sales data. Computational Statistics & Data Analysis, 138 (??):239–259, October 2019. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947319300957.

Anonymous:2019:EBk

[2592] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 139(??):ii–iv, November 2019. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947319301306.

Lu:2019:EBP

[2593] Jun Lu, Xuehu Zhu, Lu Lin, and Lixing Zhu. Estimation for biased partial linear single index models. Computational Statistics & Data Analysis, 139 (??):1–13, November 2019. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947319300738.

Bindele:2019:GSR

[2594] Huybrechts F. Bindele and Brice M. Nguelifack. Generalized signedrank estimation for regression models with non-ignorable missing responses. Computational Statistics & Data Analysis, 139(??):14– 33, November 2019. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947319300982.

Li:2019:EPS

[2595] Haoqi Li, Huazhen Lin, Paul S. F. Yip, and Yuan Li. Estimating population size of heterogeneous populations with large data sets and a large number of parameters. Computational Statistics & Data Analysis, 139 (??):34–44, November 2019. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947319301069.

Chang:2019:PBC

[2596] Bo Chang and Harry Joe. Prediction based on conditional distributions of vine copulas. Computational Statistics & Data Analysis, 139 (??):45–63, November 2019. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947319301057.

Wang:2019:MCM

[2597] Zhonglei Wang. Markov chain Monte Carlo sampling using a reservoir method. Computational Statistics & Data Analysis, 139(??):64– 74, November 2019. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947319301124.

Chen:2019:NEM

[2598] Tong Chen and Thomas Lumley.

Numerical evaluation of methods approximating the distribution of a large quadratic form in normal variables. Computational Statistics & Data Analysis, 139(??):75– 81, November 2019. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947319301136.

Yang:2019:WCM

[2599] Guangren Yang, Yiming Liu, and Guangming Pan. Weighted covariance matrix estimation. Computational Statistics & Data Analysis, 139 (??):82–98, November 2019. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947319301070.

Tian:2019:FDD

[2600] Yahui Tian and Yulia R. Gel. Fusing data depth with complex networks: Community detection with prior information. Computational Statistics & Data Analysis, 139(??): 99–116, November 2019. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947319300088.

Cui:2019:DFT

[2601] Hengjian Cui and Wei Zhong. A distribution-free test of independence based on mean variance index. Computational Statistics & Data Analysis, 139 (??):117–133, November 2019. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016794731930115X.

Choi:2019:IRA

[2602] Byeong Yeob Choi and Jae Won Lee. The isotonic regression approach for an instrumental variable estimation of the potential outcome distributions for compliers. Computational Statistics & Data Analysis, 139(??): 134–144, November 2019. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947319300970.

Zhao:2019:ETJ

[2603] Yan-Yong Zhao and Jin-Guan Lin. Estimation and test of jump discontinuities in varying coefficient models with empirical applications. Computational Statistics & Data Analysis, 139(??): 145–163, November 2019. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947319301148.

Skripnikov:2019:RJE

[2604] A. Skripnikov and G. Michailidis. Regularized joint estimation of related vector autoregressive models. Computational Statistics & Data Analysis, 139 (??):164–177, November 2019. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947319301185.

Mohammadi:2019:EDN

[2605] Hossein Mohammadi, Peter Challenor, and Marc Goodfellow. Emulating dynamic non-linear simulators using Gaussian processes. Computational Statistics & Data Analysis, 139(??): 178–196, November 2019. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947319301173.

Anonymous:2019:EBl

[2606] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 140(??):ii–iv, December 2019. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947319301525.

Vicuna:2019:MDE

[2607] M. Ignacia Vicuña, Wilfredo Palma, and Ricardo Olea. Minimum distance estimation of locally stationary moving average processes. Computational Statistics & Data Analysis, 140 (??):1–20, December 2019. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947319301161.

Quessy:2019:GFT

[2608] Jean-François Quessy, Louis-Paul Rivest, and Marie-Hélène Toupin. Goodness-of-fit tests for the family of multivariate chi-square copulas. Computational Statistics & Data Analysis, 140(??):21–40, December 2019. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0167947319300921.

Rathke:2019:FML

[2609] Fabian Rathke and Christoph Schnörr. Fast multivariate log-concave density estimation. Computational Statistics & Data Analysis, 140(??):41– 58, December 2019. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947319300891.

Groll:2019:LTP

[2610] Andreas Groll, Julien Hambuckers, Thomas Kneib, and Nikolaus Umlauf. LASSO-type penalization in the framework of generalized additive models for location, scale and shape. Computational Statistics & Data Analysis, 140 (??):59–73, December 2019. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947319301392.

Shen:2019:SRA

[2611] Pao sheng Shen, Hsin-Jen Chen, Wen-Harn Pan, and Chyong-Mei Chen. Semiparametric regression analysis for left-truncated and intervalcensored data without or with a cure fraction. Computational Statistics & Data Analysis, 140(??):74– 87, December 2019. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947319301409.

deGunst:2019:PED

[2612] Mathisca de Gunst, Bartek Knapik, Michel Mandjes, and Birgit Sollie. Parameter estimation for a discretely observed population process under Markov-modulation. Computational Statistics & Data Analysis, 140(??): 88–103, December 2019. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947319301422.

Wang:2019:MBC

[2613] Wan-Lun Wang, Luis M. Castro, Victor H. Lachos, and Tsung-I Lin. Modelbased clustering of censored data via mixtures of factor analyzers. Computational Statistics & Data Analysis, 140 (??):104–121, December 2019. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947319301355.

Tian:2019:NMA

[2614] Guo-Liang Tian, Yin Liu, Man-Lai Tang, and Tao Li. A novel MM algorithm and the mode-sharing method in Bayesian computation for the analysis of general incomplete categorical data. Computational Statistics & Data Analysis, 140(??):122– 143, December 2019. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947319300969.

Zhang:2019:RSD

[2615] Jia Zhang and Xin Chen. Robust sufficient dimension reduction via ball covariance. Computational Statistics & Data Analysis, 140(??):144– 154, December 2019. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947319301380.